A Complete Bibliography of Publications in Linear [and] Multilinear Algebra

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

27 September 2023 Version 1.21

Title word cross-reference

(0, 1) [DdF22b, FV97, Kim97, Zel85, BF15, BS18d, Cao19, KS83, LN10, Min74d, Zho16]. (0, 1, −1) [GZ99]. (1, −1) [KS84]. (1, 0) [MP13b]. (1, 1) [Zha20a, Che03b]. (1, 2, 3) [GK18]. (2) [Che95a]. (2, 2, 2) [SF17]. (2, p) [YD14]. (A, m) [GJR22, SM23]. (α, β) [YZ13]. (α, β, γ) [KO10]. (AXB, CXD) = (E, F) [YLW15]. (AXB, CXD) = (E, G) [YTL20]. (b, c) [SĆK^{+}20, Boa19, Dra18b, Dra22b, KCICV18]. (C_{1}^{∨}, C_{1}) [Hua22a]. (e(x_{r}y_{s})) [Pre85]. (k) [Cha77, Cha79a]. (k + 1) [BT06]. (k, l) [Rem90]. (k = 2, 3, 4) [Sta15b]. (m, C) [KLL22, MCL20]. (m, m − 1, 0) [HZ21]. (max, +) [HKT15]. (n + 1) [Sch78, WT23, Sch75]. (n, k) = (4, 2) [JLL88]. (n, m) [CM19]. (p, h) [DV18]. (p, q) [KL95]. (p, σ) [ADRP14]. (qr) [Raw81]. (R, S) [fLLyP15, Pen15]. (s, p) − w [Ras17]. (\sqrt{5} − 1)/2 [Sim95]. (T) [Che95b]. (υ, k, λ) [Der83]. (Z_{m})^{s} [How86]. ∗ [SN23, ZCZT22]. −1 [WWWT22]. −2 [Ano16b, BPS16]. 0 [HNS99, HLZ05, HL20, JS96, Wu09]. 0, 1 [MS79, BX07]. 0 < p ≤ 1 [BS22a]. 0 ∉ F (A^{m}) [JL90]. 1 [CM15, DFK19, Fra17, HNS99, HLZ05, HL20, JS96, KP14a, Min77, Sta10, Wu09]. 1, 2, 1 [KHG13]. {1, 2, 3} [ZX10]. {1, 2, 4} [ZX10]. {1, 3} [CIM19]. 10 [LN10]. 12 [EM95]. 2 [AA20, BL10, BL01, BdPN14, BDvW16, BM89, But10a, COK20, DFK19, DP09, DD10, Din15b, – DL92, – Dok93, Ere15, Far94, FGMSS19, Fur08, GS06, Hai88a, KY17, LGA^{+}13, Liu16a, Liu18a, Liu19, Mah82, MP18, NJP19, SG09a, SF17, WYM11, YF18]. {2, 3} [SKP18]. {2, 4} [SKP18]. 2d − 2 [DM19]. 2n − 2r(≥ n) [LM95]. 2s [Lu22]. 2 × 2 [ADJAH22, AY74, Bas19, BB96, CM92, CTW15, Coh88, CZ22, Dre22, FLT22, GR93c, HW06, JR93a, JR93b, Kan87, LWC22, NP91, RSM22, Tau81, UG21, WLB19, Wan19c, Wei98]. 2 × 2× [Bre12a, Mar11, Sta14]. 2 × 2 × 2 [BS13, KJ13]. 2 × 2 × 2 × 2 [BS13, KJ13, Sta14]. 2 × 2 × 2 × 2 × 2 [SLH16, SLH18]. 2 × 2 × 4 [WCS^{+}21]. 2 × ⋯ × 2 [SL21]. 2 × k × (k + 1) [Ait19]. 3 [BWLL12, BW15, BPW16, CCGO11, DJ89b, DBG18, KP12, LK16, LZ12, MSS18, QH17, SF17, Vo83c, Zha15c, dNV19]. 3k < n ≤ 4k [Cha79a]. 3 ≤ m ≤ n [SMS15]. 3 × 3 [BLSV12, Cai93, Che22a, CCL10, CNP06, Dom95, Dom00, DC20, For79d, SvW14, Tia02]. 3 × 3× [LPZ13]. 4 [CM07, GS21a, LHbL^{+}15, MH16, MS16a, TWZ16]. 4 × 4 [CFSRL22, WCS^{+}21]. 5 × 5 [LS21b, Zha20b]. 6 [MC19]. 6 × 6 [Che22a]. 7 [Rom89]. 8 [MN13]. [−1, 1] [dLMO17]. [A, [A, [A, X]]] = 0 [Gra80]. [A, [A, X]] = 0 [Gra80]. [a, b] [CR22b]. ^{d} [GKO22]. _{1} [CY22a]. _{2} [JN14a]. _{β}^{α}Φ_{s} [Lei18]. _{n} [CdM15, INR18, Slo13c]. _{n,m} [SA14]. _{p} [Boz99]. A [CS23, Fan92, GBM21, JT03, NT22, QHB22, RSM21, RSM22, SAK22, Sta15a, WZ22, WM84]. a ∗ a = P (a) [HE20]. A + BX [JL91]. A + XBX^{−1} [dOMD77]. A − I [Fan92]. A = BXC [LT10]. A^{∗} [HS83a]. A^{+} [HS83a]. A^{n} = A ∗ A [DMT21]. A_{α} [CH22a, GZ22b, JWBB22, LY22c, LS22b, LCG^{+}23, XSW22]. A_{T,S}^{(2)} [Boa19, Sta15a, YW09]. AA [CS23]. aa^{+} = bb^{+} [PA10]. AA ≥ 0 [JT03]. {ab − ba} [BLR12]. AB = BX [LR89]. aba = a [BLR12]. ABC [Che22b]. AH + HA^{∗} [AL78]. α [CR18, GX17a, Heo16, LGZ20, LL22b, SWH22, YS23]. α_{β}Ω [LXQ21]. ∗ [BD12, Cir15, HZXL15, Kre16, MRD15, TRD16, WJ16]. A → A^{+} [Gol83]. Au = b [McD92b]. AX + BX^{∗} = 0 [De 13]. Ax + B|x| = b [Roh04]. AX = B [fLLL^{+}17]. AX = C [YZL22]. AXA^{∗} = B [Bak84, Zha04c]. axa^{∗} = c [CIDK07]. AXB + CY D = E [LQ95]. AXB = C [dJfLlZ^{+}21, Tia03, ZLZ^{+}22]. B [BCT23, GCSC20, SC21, WS18, ZZ16, Liu17, ME18b]. B(H) [BJZ13, Dol04, IK18, ZXL16, YJ21]. B(n, k) [LH21b]. B(X) [QC23]. b_{2,2} [ZZ00]. B_{ran} [JM13]. B_{row} [JM13]. B_{v}^{(m)} [MM20b]. bab = b [BLR12]. β [LLF18]. bv_{0} [ESAJ18]. C [Abo20, And96, BD88, BPD88, CDPS23, CT96, DHKSH08, DvE20a, DvE20b, LXQ21, Li87a, LT89b, Li94, LR04, LZ19b, LP22c, LP22b, Man91, Man92, NNT95, NY21, PMP22, RRN22, RRN23, ShDHG08, SA21, WZZ22, BCNG23, BD94, BD96, Ilk20, Jon92, JR93c, Lei91, Lei92, Li86, LT88a, LT88c, LST89, LMR94b, LDL17]. C(X) [BG16]. C∗ [BBX20, DM20, FMX18, GS20a, GNO20, HJ19b, HJ20b, Lud19, MW21, TH18, HJ19a, Xia20]. C^{∗} [AR15, BCNG23, BBK22, BBBS09, BR10, BM17b, CIDK07, DM17, DKS09, DKM10, EM15, EPR16a, EPR16b, FXY22, FMXZ22, MR10, PRA22, Son14, Tag12, WZ23]. c^{n} [Gol90, Shk12]. C^{⋆} [BB17b]. C_{0} [Ber02, Ilk20]. C_{0}(Z) [AO22]. C_{m} [YWO07]. B_{s}(H) [AKS18]. G [Mit92]. H [HM19]. L(l_{m}^{2}, l_{n}^{p}) [Grz89]. charF ≠ 2 [Uhl79]. cp [BX07]. CXC^{∗} = D [Zha04c]. D [JvdD88, LN10, NW02, Sch22, SA21, vE22, Cai93]. D_{α} [ABP22, LXS21]. δ [MC17]. Δ_{i}^{3} [TRM22]. det X^{t}X [Sha94]. E [Arm17, HQ14, XSH18, Ano22a, MH20]. e_{11} − e_{22} [Jul21]. ℓ_{p}(F) [Ogu20]. ℓp^{2} [MZ23]. EP_{r} [Mee80]. η [HW14, TKT19, YLW15]. F [Bea86, Uhl79, KD16]. F_{1}(A, B, B′, C; z, w) [AEAAES20]. F_{a} [LH22]. f → f^{+} [Gol90]. G [Hof75, MS82a, MW22, Nie07, Nie22b, Nie22c, Pab22, BCNG23, CS21a, CS22, FZ21, GK18, GLL21, HLY22, LLHS18, LL20, LD22, LLL09, LW08, LDL17, MAT22, Xia20, XQZ13, XQ16b]. {g, h} [Bre16]. G_{2} [Sun21]. γ [JKD16]. Γ_{0}(n) [Cha96]. GC_{2} [ZY08]. GF(2) [Bot97, Cla92, Tra89, Tra90]. GF(3) [Ell95]. GF(q) [Ger86]. GL(n) [KW93, Ner22]. GL(n, F) [Sch78, Sch75]. GL(n, R) [Mer82]. GQ(2, 4) [BLSV12]. GSp(V) [Vin04]. H [AM13, CHL18, FH89, HJ20a, Jin17, JKB17, LZ16, LZX19, LW20, Neu84, WLM21, YZH19, ESAJ18, GR91, GR93b, TRM22]. H > 0 [AL78]. I [LRD21]. i(XAX^{−1} + B) [Mac95]. ind(E) ≤ 2 [XSH18]. ∞ [BB19b, CD21, MW15, WT16]. I × I [Esh22b]. J [Heo16, MAH18, MAH21]. K [BCNG23, Poo80b, SLW21, Slo13c, AAK16, Atk84, Bal22b, BLZ10, Bom02, BM18c, BM19b, CLP22, Che14b, CKT20, CL21, FGHL22, FNBR96, FLT20, Gro17, HHSY22, HQ18, HLZ11, HL20, IAM23, JMS19, JLL88, JR94, KW21, KKP13, KLPK17, LT91, Li92, LLS10, LF15, LCC17, LP23, LDL17, LWL21, LW22a, Loe93a, Loe93b, MW15, MPD22, MDP22, Mej19, Mes90, Šem97, SSG12, SHS17, Sta15b, TB01, WT16, WDJ10, ZL19]. K(A) [Mer75]. K(∫↕_{∈}, ∥) [LY22a]. K^{d} ⊗ K^{d} ⊗ K^{d} [DM19]. K_{3,3} [FK91]. K_{s,t} [CZ21]. L [BB19b, CLX03, Geh15, Tau74, BY19]. L(l^{p}, l^{r}) [Grz91]. L(X) [Cos21]. L^{2} [Lla96]. L^{2}(Σ) [JS19]. l^{p} [Rho94]. l_{1} [AG96]. L_{2}(0, ∞) [FD14]. L_{p} [BS22a, BPR12, MM09, Boz98a, Gol87, PS20, RF19b, Zen22]. l_{p}(1 < p < ∞) [MM20b]. l_{p}(Δ^{n}) [RF19b]. l_{p}(w, F) [DTS14]. l_{p}^{n} [CSS21]. L_{r}⊗ [YLX20]. l_{u}(l_{v}^{n}|n ∈ N) [Bad19]. λ [CM19, LGG19, PZGL22]. λ − τ [MS15a]. λ2(G) = d2(G) [PH03]. λI − N [Che94]. LDL^{∗} [Sta15a]. lim_{λ→0} X(λI_{p} + Y AX)^{−1}Y [YZ16]. log [KKY22]. LR [Mar78, TT15]. M [And80b, BCN05, BCEM12, BC94, CS23, CKN95, DLL19, DJL94, Fan92, FS83, GN02, Ha88, HS89, HS91, HXL16, Ibr00, Joh77b, JOvdD84, Lee91, Li95, LLZ13, LCLL22, MSZ04, Mia93a, Nab98, NP80, NS92, RP22, vD96, Ber12, Den14, DK21, DK22, JOvdD85, LOO22, MRT16, NS93, yT86, WD17, Zhu19, ZCP18]. M_{1,1}(E) [NT22]. M_{2,A,B} [BM22a]. M_{n}() [Foš05b]. M_{X} [CIP17]. Mat_{n}(K) [dB16]. F_{2} [BS13, SLH16, SLH18, SL21]. F_{3} [BS13]. GL_{1} [TS21]. GL_{2} [TS21]. GL_{3} [TS21]. Z [Ger04]. L [Tur18]. Q [Tur18]. ⇕ [WLL21]. ∫↕_{\} [SZ19]. ∫↕_{\} [CdSS17]. ∫↕_{\}(C) [MS08]. m × n × (m − 1)n [SMS15]. μ [Dia90a, dF05, dF19]. N [Shr90, AS20a, An18b, Are22, BL10, BGLZ18, Bed11, CKS19, CLX22, CR18, Cha79a, CQ22, Chi78, Elg22, EÖ15, GLOT13, GR10, HMK22, HW75, JO23, KP16, LZ13, LDL17, MCL20, MP19, MS21a, Mos21b, Pop17, Qi15, SV14, Uhl14, Wer87, YLL90, dB16]. n − 1 [Esc93]. n − 2 [HW75]. n/2 [KMS02]. N_{0} [Sin90, Lee91]. n ≤ 3k [Cha77]. n × n [CZY12, Gai77, SvW14]. O [Lop11a]. O(2, 1) [CT22]. O(4) [Lop18]. O(N) [Uhl99]. O(N^{2}) [Uhl99]. O(p, q) [Les93]. Ω [Poz22, EV77, TAK22]. Ω(r, s) [Har87]. Ω_{n} [DH91]. P [AEdF13a, AEdF13b, DdF15, ES98a, EMS02, KS14a, Tau74, TZ21, ACO21, BHK21, BB96, BKMMG22, CFLL23, Han17, HS93b, KKY22, Kra06, Kul22, KV22, LL22b, LHGL22, NMAB22, Roh91b, Sin77, So90, Sun21]. P^{14} [BFM22]. P_{+} [KS14a]. P_{5} [Gho19]. P_{n} [RP90]. ∂W_{c}(A) [Beb86]. per(I − A) [Mal86]. Φ [HSM17]. Π [LH13]. Π_{2}(l_{p}, l_{q}), 1 ≤ p, q < infty [Bad19]. PSL_{n}(F) [Gur97]. Q [GZ22a, LLC21, ND05, SST22, YGW15, YWZ21, ZXL22, CNP05, CN05, CNP06, CN07, Gar79, KHG13, Li98, LN98, PV00, Sal12, Sal14, SAK22, WHG17, YLC19b, Zha97, ZZ23, Zho22]. Q_{#} [SST22]. Q_{+} [SST22]. QN [GHL20a]. QR [LL17]. QX [LZ22]. R [AW13, LZ15a, LD22, LZ19b, BL93a, BT15, CK19, HH14, McC85, Mou22, Wan22b, ZL19]. R^{n} [AGP22, Shk12]. S [AJL22, FJK22, GMZ11, GZH19, XL17]. S(1^{a}, b)[P_{k}] [CRY95]. S(n, p, 2, 2) [Hai92a, Hai92b]. S(n; 1) [CC19]. s(ϕ) [GW98a, GW98b]. S_{m} [SS98]. S_{n} [Di 88, Reg87a]. Σ [BCKP10, Ben16, Ben22, Lee16, Wan18]. SL(2, R) [Sch79]. SL(3) [Raj09]. SL(n, F) [New82b]. SL(n, K) [Wat84b]. sl_{2} [AC13, Aln15, Gau76]. Sp(2n, Z) [New85b]. ⋆ [RD12]. T [NY21, WBZ14, BT08b, CX23, GL16, LM93, SRD19, vBW95]. T^{∗} [LZ12, LZ13]. τ [AR22, Han17, Meh84, Sha20]. θ [ZT12]. U(n) [– DM79]. U_{q}(sl_{2}) [Aln11]. UT_{m} [UF22]. W [Arm17, GCP20, SKM17, WfMR16, ZT12, Ras19, ZWM23]. W (12, 11) [KMS07]. W (2, 2) [ZT12]. W (m, n; t) [CCC21]. W (n; 1) [CC19]. X [Tia17]. X + M^{∗}X^{−1}M − N^{∗}X^{−1}N = 1 [Zha16b]. X − AX̂B = C [ZWLZ22]. X = (x_{ij})_{N×n} [WM84]. X^{∗}Y + Y X^{∗} [LJ11]. X^{p} = A#_{t}(M^{T} (X^{−1} + B)^{−1}M) [LM22]. X^{p} = A + ∑_{i=1}^{m} M_{i}^{∗}(B + X − 1)^{−1}M_{i} [JZ22]. X^{T} AX = B [BCT23]. x_{ij} = −1, 0, 1 [WM84]. XA^{t}Y = A [Wal75a]. XA^{t}Y = A^{t} [Wal75a]. XAX = BX [WHL22]. XB = D [YZL22]. XC = D [fLLL^{+}17]. ξ [LLF14, YZ13]. Xm + A ∗ X − nA = I [DH03]. X → AX + XA [Bou14a, Bou14b]. XY [CD22]. Y X [CD22]. Z [JDD21, MNN^{+}98, Smi94, SH98, Stu89]. Z_{1} [CPZ17]. Z_{2} [DB14b]. Z_{n} [PRAC22]. Z_{p^{r}}Z_{p^{s}} [AS15]. ||A||_{2} [Mer81]. ||A||_{∞}1 [Roh00]. ||f(A)|| = f(||A||) [BR22].

* [FdSV16, FA19, GV95, GdSV16, IT16, LWZ15, Ng14, NY21, QZ16, FMXZ22, HJ20b]. *-algebras [Ng14]. *-crossed [NY21]. *-derivations [FA19, LWZ15, QZ16]. *-identities [GV95]. *-isomorphisms [IT16]. *-modules [FMXZ22, HJ20b]. *-superalgebras [FdSV16, GdSV16]. *congruence [HS09].

-additive [AS15]. -adic [Kul22, KV22]. -algebra [ZT12]. -algebras [BM18b, CR18, GLOT13, Poz22, BBK22, BBBS09, BR10, BB17b, BBX20, BG16, BM17b, CIDK07, DM17, DM20, DKS09, EM15, EPR16a, EPR16b, Lud19, MW21, PRA22, Tag12, TH18]. -Aluthge [CM19]. -analogues [Gar79, Zha97]. -ary [Bed11, GR10, KP16]. -asymptotic [Geh15]. -based [Mit92]. -bases [HJ20a]. -beta [Sal12]. -bi-Hermitian [YLW15]. -bihermitian [HW14]. -bilinear [MS82a]. -bistochastic [ME18b]. -by- [DD10, DJ89b, GS21a]. -Cayley [Are22]. -central [CR18]. -circulant [Bom02, BT15, Mou22]. -circulants [GK18]. -cocharacters [Di 88]. -colorings [Dia90a, SG09a]. -coloured [KP12]. -commutative [FLT20]. -commutativity [HQ18, QH17]. -congruence [Lei91, LM93]. -connected [SHS17]. -connectivity [Vo83c]. -continuous [ADRP14]. -contractions [FH89]. -convertible [Ber12]. -convex [DV18, Li86, LST89, Li87a]. -core [ZWM23, WLL21]. -cospectral [BB19b]. -cospectrality [Tur18]. -CP [Heo16]. -cyclic [BL93a, LDL17, MW15, WT16]. -deficient [LLF18]. -demicompact [FJK22]. -derivations [Ben16, Ben22, Bre16, Lee16, Qi15, Wan18]. -derivative [ZZ23]. -det [BPD88]. -det-convex [BPD88]. -difference [Sal14]. -Differential [LGG19, PZGL22]. -dimensional [COK20, Far94, HMK22, MC19, WT23, YLL90]. -Divergence [KD16]. -Drazin [CS22, Pab22]. -duals [LD22]. -edge-coverings [But10a]. -edge-Hamiltonian [LP23]. -eigenvalue [GZ22a, YWZ21]. -eigenvalues [GZ22b, Jin17, JKB17, LXQ21, LCLL22, LY22c]. -eigenvector [CPZ17]. -eigenvectors [HQ14]. -eliminated [TB01]. -energy [Tur18]. -Entry [Che03b]. -EP [Zhu19, ZCP18, WD17]. -even [AM13]. -extension [LZ12, LZ13]. -Filiform [CCGO11, ACO21]. -frame [MAT22, Xia20]. -frames [FZ21, GLL21, HLY22, LLHS18, LL20, BCNG23, LH22]. -free [Gho19]. -function [HSM17]. -functions [Hof75]. -Gamma [SAK22, Sal12]. -generalized [Che95a, CIM19, GS06, Liu17]. -generator [Mah82]. -Graded [Din15b, CdM15]. -gradings [FGK23]. -graph [LZ15a]. -graphs [BB19b]. -groups [Cha77, Cha79a, Kra06]. -Hamiltonian [LP23]. -hamiltonicity [Lu22]. -Hirano [CS21a]. -homogeneous [FGMSS19]. -hook [Rem90]. -hypertournaments [KKP13]. -hyponormal [Ras17, Ras19]. -hyponormality [KKY22, GBM21]. -ic [Chi78]. -ideals [BY19]. -identities [NT22]. -inclusion [LXQ21, Lei18]. -increasing [Nie22b]. -index [LS22b, YGW15]. -inner [SN23]. -invariants [Lop11a, Lop18]. -inverse [Boa19, XQ16b]. -inverses [Dra18b, GK18, KCICV18, LW08, SKP18, XQZ13, ZX10]. -invertibility [DK22]. -invertible [Dra22b]. -Inverting [KHG13]. -isometric [MP18, MP19, KLL22, GJR22, MCL20]. -isotropic [JMS19]. -Jordan [MC17, An18b, TRD16]. -Jordan*-derivations [LLF14]. -Jordan-triple [CM19]. -Laguerre [Mej19]. -Laplacian [CFLL23]. -lattices [Ger04]. -left [DK21]. -Lie [AS20a, BL10, BWLL12, BD12, BW15, BGLZ18, DBG18, JO23, LZ12, LZ13, SV14, TKT19, WJ16, YZ13, Zha15c]. -linear [CKS19, Pop17].

-linearity [WZ22]. -Local [Liu16a, AA20, Liu18a, Liu19, YF18]. -majorization [Nie22b, Nie07, Nie22c, vE22]. -Martices [MNN^{+}98]. -Matrices [JT95, Min74d, Ano22a, Cao19, EV77, GZ99, HS93b, KS83, KS84, MS79, MH20, Meh84, NS93, Roh91b, Zho16, EMS02, GN02, And80b, BCN05, CKN95, DJL94, ES98a, FS83, GHL20a, Ha88, Ibr00, JDD21, Joh77b, KS14a, Lee91, Li95, LLZ13, LZ16, MSZ04, Mia93a, Nab98, NP80, Neu84, NS92, SC21, Sin90, SST22, Smi94, SH98, Stu89, TZ21, WS18, YZH19]. -matrix [BF15, BS18d, CQ22, JOvdD85, BC94, Fan92, ABP22, CS23, DLL19, HS89, HS91, HXL16, JOvdD84, NP80, SST22]. -measurable [AR22, Han17, Sha20]. -minor [CZ21]. -minors [Sin77]. -module [HJ19a, FXY22]. -modules [AR15, AO22, BCNG23, DKM10, FMX18, GS20a, GNO20, HJ19b, MS08, MR10, Raj09, Son14, WZ22, WZ23, Xia20]. -multigraphs [BM89]. -nilpotent [Fur08]. -norm [BKMMG22, CD21, Han17, JLL88, QHB22]. -normal [GR91, GR93b, LL22b, MRT16, RRN22, RRN23, WZZ22]. -Norms [So90, Gro17, YLC19b]. -Numerical [BD94, LN98, AAK16, BD96, BHK21, CLP22, CNP05, CN05, CNP06, DvE20a, FNBR96, Jon92, JR93c, JR94, Lei92, LT88a, LT88c, Li98, PV00, DvE20b, LR04, Abo20, BD88, CT96, CN07, DHKSH08, LT89b, Li94, LP22c, LP22b, Man91, Man92, NNT95, QHB22, RSM21, RSM22, ShDHG08, vD96]. -odd [AM13]. -operators [LW20]. -optimal [LN10, NW02, WM84]. -or- [DFK19]. -orbihedra [ZZ00]. -orthogonal [MAH18, MAH21]. -orthogonalities [TAK22]. -outer [MW22]. -parallelism [TAK22]. -partial [SM23]. -partite [Shr90]. -path [SRD19]. -pendant [WDJ10]. -permanent [dF05, dF19]. -point [Sch75, Sch78]. -potent [BT06, BLZ10, Den14, HLZ11, McC85, BT08b]. -potents [Zho22]. -prism [CLX22]. -product [CX23, HZXL15]. -property [BCEM12, ND05, Tau74]. -pseudoalgebras [LW20]. -quasi- [MCL20]. -Racah [WHG17]. -radial [Li86]. -radii [Abo20]. -rank [CHL18, LZX19, WLM21]. -relatively [JKD16]. -representations [Heo16]. -representing [RD12]. -restricted [LOO22]. -rings [BY19, Cir15, Kre16, MRD15, ZCZT22]. -rooted [RP22]. -rose [MH16]. -Schmidt [LF15]. -secant [Bal22b]. -selfadjointness [DK22]. -sets [Li92]. -shape [LH13, WBZ14]. -Simon [Raw81]. -singular [GL16]. -smoothness [MPD22, MDP22]. -solvable [BL10]. -spaces [Atk84, CSS21, Mes90, BPR12]. -spectra [LXS21]. -spectral [Li86, LMR94b, LGZ20, LL22b, SWH22, YS23, CH22a, JWBB22, LCG^{+}23, ZXL22]. -spectrum [AJL22]. -splittings [JM13]. -stability [Cai93, GX17a, JvdD88]. -statistical [SAK22]. -strictly [BCKP10]. -strong [Mos21b]. -subgroups [Sun21]. -symmetric [CDPS23, fLLyP15, Pen15]. -tensor [PMP22]. -tensors [MSS18, GZH19, HM19]. -th [HL20, IAM23, LLS10, vBW95]. -to- [LZ19b]. -Toeplitz [BdPN14]. -torsionfree [BDvW16]. -traceability [Lu22]. -trees [FGHL22]. -tuple [MS21a]. -type [MW15, WT16]. -uniform [LCC17, LWL21, LW22a, Wan22b, ZL19]. -unitary [GR91, ZT12]. -uplets [BB96]. -valued [AW13, BG16]. -variable [KY17]. -vectors [LK16]. -vertices [AEdF13a, AEdF13b, DdF15]. -weighing [LN10]. -Weighted [SKM17, GCP20, LLC21, SLW21, WfMR16]. -Weyl [GCSC20, ZZ16]. -zero [CLX03].

03081087.2015.1020746 [Ano16c]. 03081087.2015.1107020 [Ano16b].

1 [LPH17]. 1-inverse [LPH17]. 10.1080 [Ano16b, Ano16c]. 10.1080/03081087.2015.1020746 [Ano16c]. 10.1080/03081087.2015.1107020 [Ano16b]. 1978/79 [TT78].

2- [SN23]. 2-by-2 [WS23]. 2-connected [YSZW22]. 2-equitable [LHF22]. 2-layered [Nie22a]. 2-linear [DIJ23]. 2-Local [AKY21, SYZ22, AAUN22].

3 [ZLS22]. 3-BiHom [BMN22]. 3-coloured [KS22b]. 3-LieRep [ZLS22]. 3-tensors [LPZ13, SMS21].

6 [TT78]. 63 [Ano16c]. 64 [Ano16b, Sło18b]. 65 [Sło18a]. 67 [HJ19a].

7 [Ano79b].

= [Tia21].

a-Weyl [BZ22a]. A. [Mar76, TR21]. abelian [Are23, CKR18, CKR21, CCS98, ES76b, MN80, MP13b, Nir11, Odo83, QX23, RC18, STME10, SZ19, WL16a, WL16b, WW20b, BF96, Tow22]. Absolute [Fie90a, KP02, DH17, Jam95, MQ21, ST73, SA10, Tag12, TH18, ZZL23]. absolutely [MSS18, PSS15]. absorption [BS18b, LZ23]. Abstract [Sin83, Sal79]. Accelerated [BLX18]. Accretive [BJ75, BKS21, JK22, JKL22, KS19b, LS17, ML21b, RQ22]. accretive-dissipative [JK22, KS19b, ML21b]. accurate [DOP22, MS21b, SM21, Uhl14, Uhl20]. achievements [Abd20]. achieving [GWW20]. acting [Fri97, KMEO12]. action [Liu16b, Mer78b, Sta23]. actions [RP04]. acyclic [AEdF13a, AEdF13b, APEPPS23, DdF15, KS09, Moh16b, dF13]. Addendum [DvE20a, JR93a, Ric87a, Bal76]. adding [LY22c]. addition [EL09, LŠ96a, LŠ96b, NC22, SS22, WTD22]. additions [LLS16, NS93]. Additive [BHD10, BLZ11, Che14a, CK19, DHB07, FLS88, Gha12, JQ15, Kuz05, LP02, OK07, Pet97, ST22, Tag12, TZ06, XCC99, XPY16, XL17, Zha04b, ZS05, ZCCIW12, AGV19, AS15, CdS03, Cho10, CW22, CL21, Fin92, Gar23, MO13, Men22, Smi83, mWyD11, YZ13]. Additivity [DW12b, QH11, AA12, Gro99, Har81b, Tia05a]. adic [Kul22, KV22]. Adjacency [Hua10, Lim10, BR14, But10b, DFPS23, GB22, HA22, JS22, KP22, KFR17, KFR20, LY22b, MKS22, Sah21, WMFW21]. Adjacent [DJK08, CW93]. adjoint [AA20, CGY13, CKL16, CK20, DRW93, DK22, FJS14, HH16, HX21, Kyr08, LLMT05, LT13, Moo95, ST19, TZ06, Tan13b, XS18]. adjointable [FMXZ22]. adjointness [FS19, Kit91b, Sin84a, WW20b]. Adjoints [Mar75, Rao18, BT22b, ER81]. adjugate [Sin82a, dCdSA22]. adjusted [MH12]. Adjustment [Har95]. admissibility [BV16b]. admissible [Aze99]. admitting [Gar75b, MID17]. advances [Mer79c]. affects [Har81a]. affine [BBT07, CY22a, FP17, GX17b, Kes83, dSP16b, SX23, Wit11]. affine-Virasoro [CY22a]. afforded [Mer78b]. after [WW20a]. Aitken [Min74c]. alcoved [dlP20]. Alexander [Min74c]. Algebra [DS02a, HJ19a, Mac09, Sło18a, Sło18b, SK03, Aln11, BPS00, Bed11, BLZ11, BV92, Boa13, Bre12b, BS18c, BF96, BSS12, Byt20, Car88, Car90, Cha90a, CKLP13, CY22a, CHS22, CZ23, CWSZ17, CSY18, Chi22, DK11, De 82b, Din15b, FGK23, FW73, Gai77, GPW22, Ghi22, GG22b, Gor83, GR10, GPK94, GV87, Guz88, IKR12, Jac79, KS15, KLJ23, LY22a, LK23, LCL13, LS22a, Lip96, LZ12, LGB17, LGG19, LW20, Mag22, MS85a, Ma’15b, Mer92, OWX09, Pat86, Pat89, Pow88, Prz92, TZ19b, TS19, UG21, Wad90, Wal79a, WOY06, lWpW14, WLZ22, WT23, Yam18, Yao22, YWO07, YH22, ZT12, ZWW10, ZAK22, ZW14, dBS00, Ano79b, Ano16b, Ano16c, TT78, MT84]. Algebraic [ASZ02, BJ96, BP98a, CIRW21, FKP03, Gur97, KN97, KP02, LZT18, Mil22, Pat86, Tho90, WOW09, Wu05, Zel85, AA21b, BLP12, Bar07, Bar12, BBK22, Bén10, BGLSSS20, CK84, Din00, DLL19, DU21, Epk89a, FN14, Fie90a, FA22, GZY18, GL97, Haw05, HSX16, KS19a, KK22b, KF98, Kir00, KMN01, KN04, Koi21, KN81, LRR87, Lim92b, LGB17, MG22, Ma’15b, New85a, NF19, PT10, Pie92c, Put82a, Put82b, TJ16, Woo22, XSL15, YSZW22]. Algebraically [Had83]. Algebras [Boz97, Bre15, Che03a, DŠ04, Fin92, Kay17, KT21, LR17, Luk74, MŠ02, Poz22, ST15, ARH15, AF11a, AF11b, AF12, ACO21, AM15, AN14, AS20a, AAP82, Ake83, AAB15, AELPM18, AGV19, AC19, AG15, AH13, AXH15, AC21, AG18, AAA22, ABM22, ALPM20, AKOZ20, AKY21, AK21, AAUN22, Bah22, BHD10, BL10, BHD11, BWLL12, BD12, BW15, BCZ22, BB21, Bal23, BCLS23, Bed10, BX11, BBK22, BBBS09, BR10, BB17b, BBX20, BBM15, BRL22, Ben15, Ben16, Ben22, BCS14, BCS18, BM14, BS18a, BDV19, BB14, BDH23, BG16, BH13, BH16, BM17b, BM18b, BTV07, Bre14, BE20, Bre16, BS22d, BH15b, BL11, BFJ12, BFK18, Bun94, CMR20, CMR21, Caf13, CR15, aCC17, CGNT15, COK20, CCGO11, CGOT13, COK19, CKKK22a, CKKK22b, CPS22, CSG22]. algebras [CV94, CLOK13, CR18, CKR21, Cen11, CC19, CW11b, CX12, CW12, CS21a, CY22b, CS22, CZZ22, CZ23, CW23, Chi78, Cho10, CJW19, CAP20, CAP22, CIDK07, DM17, DM20, Di 88, DKS09, – Dok73, DR18, Dra22a, Dre93, DW12b, DW16, Egg95, EM15, EKO14, EL21, EPR16a, EPR16b, Far93b, FJK14, FZ17, FTGSNV22, FdM18, FA19, GLOT13, GP14, Gha12, Gha13, GMV18, GS95, Gol07, GT11, GJ22, Gow22, Gra10, Gra11, Gur80, HK16a, HE20, HKP23, Hil76, Hon19, Hon22, HFS15, Hou17, HQ18, Hua96, HLH18, HZXL15, IM18, Iop19, JP15a, JL22, JLZ12, JQS13, JQ15, JX13, Joh22, JN22, JO23, Kan16, KK75, KKL20, KKL21, KP16, KPPV18, KPS19, KK22b, KLO14, Kol23, KZN22, Kov77, KLO19, Laf80, Laf84, Law19, LZ09, Lee15, LS16]. algebras [LB11, LPS13, LF13, LLF14, LWXF15, Lim77, Lim92c, LJ11, LZ13, LCL14, Liu15, Liu16b, Liu16a, LL18, Liu22a, LG23, Lud19, Lud21a, Lud21b, LBV13, LT22b, MJ07, MDW11, ML12, MPGS21, MM92, MM05, Mar12, MHA17, MW21, Mil15, Mit10, MRS10, Mos21a, NTW23, NR23, Nas74, Ng14, Ng22, NZK21, Nir11, NJP19, Oi20, ORT15, Pan22, Pat96, PZGL22, Per17, Per15, Per19, PW78, dSPdS22, Poo17, PRA22, PV20, QM21, QH11, Qi15, RH11a, RH11b, Reg93, RC22, RZ17, RE22, Rom89, SV14, SH21, SDVWT16, STME10, SN22a, SX15, Shi18, Sin21, SL15, Spi04, ST99, Tag12, TRD16, TH18, TKT19, Tan18c, TWK22, TJG22, Tho92c, Tow22, TNJ22, Vol22, WOY06, WHX08, WOW09, WY11, WJ16, Wan16d, Wan18, WLB19, Wan19c, Wan20, XW12, XWF15, YD14, YZ13]. algebras [YF18, YZ20, Yan21, YW10, YH22, Zha97, Zha15c, ZL15, Zha16a, Zha21, ZWZ10, dTS20, vWZ19, LPS21, Ano22b]. algèbres [MM92]. algorithm [AAB17, DCLW19, FTA20, Fog73, GTLN18, Gre99, HKT15, HRT15, HM20, KV22, fLLyP15, LLC21, LLV18, Mum86b, TT15, Uhl13, XJW17, ZZT23, vD96]. Algorithmic [CSG22]. algorithms [Blo95, CW90, Che95a, HM18, Kul22, Mum86c, Stu92, WX19]. all-derivable [LPS13, ZXL16]. allow [FHL^{+}13, GGH22, LHbL^{+}15]. Almost [BP19, Rum97, CS19, Dag23a, GdSV16, ND05, PR12, WHG17, Zel75]. almost-commuting [Zel75]. alone [De 82b]. along [BB17b, Sza09, WM21, ZCP16, ZPCZ16, ZCP17]. alternate [Kuz05]. Alternating [Bri89, KM17, BS07, FHZ17, GQ06, Hor04, KKS22, MN13, Mis17, Rei75, WZY22]. Alternative [ST80, Bar07, Car20, Mel10, Mel18, Vei83a]. alternatives [Roh04]. Aluthge [And04, CM19, DKY22, JLL07, KKY22, RS08b]. AM [Bak16]. AM-GM [Bak16]. Amitsur [Ari16, DP18]. among [DIJ23, KD23, Tan10, TLL11, TLC16, XL11]. Analog [Mia93a, CH87, Foa79]. analogous [CW91]. Analogs [Kyr08, HINS10, Ter22b, WW76]. Analogue [Bed11, Gut23, KA19, NSW21]. analogues [AdFK15, Gar79, HS96, Hon92, Hon93, LN16, Zha97]. analyses [DS23c]. Analysis [CLX03, BX11, BVW84, Chu84b, DLQ18, FS22, KW89, LYS15, LZ19a, LZ22, MMRR09, Mej19, SRD19, SR22, Ste94a, Tia21, jXL15, ZY13, ZBW12, vD96]. Analytic [FS83, Fla74a, LCC17, Tur02, Dra22a, Fan83, FJS14, GH23]. analyticity [JMS19]. Andelić [KAE22]. Ando [FS15, HSW19, KMS21, Lim23, Tom22]. angle [Bal80, Zha13]. angles [Gal08, RBB22]. Anisotropic [Pie74]. Annihilating [MW74a, DS23a]. annihilator [LT22a]. announcement [Tho90]. answer [GR14]. Answers [BHM76]. Anti [BFJ12, AF11a, AF11b, AF12, AHJ20, AHJ21, AAB17, BGM22, BSZW13, Haj18, HA22, HS12, LK21, LK23, LLL09, LL10a, PSB23, XSH18]. anti- [AAB17, LL10a]. anti-automorphic [LK21]. Anti-automorphisms [BFJ12, LK23]. anti-involution [AF11a, AF11b, AF12]. anti-Loewner [HS12]. anti-reciprocal [AHJ20, AHJ21, BGM22, HA22, PSB23]. anti-reflexive [Haj18]. anti-symmetric [LLL09]. anti-triangular [BSZW13, XSH18]. Anticommutative [Vol22]. anticommutator [DWCD19]. anticommuting [Ito16]. Antiderivations [Lim77]. antidiagonal [AFVL00]. antiinvariant [Tam98]. antilinear [HL21]. Antinorms [Pro22]. antiregular [BZ00]. antisymmetry [Pat95]. any [BL21, Hua10, LB19, RC22]. Anzahl [GL16]. AOR [CLW21]. Appell [AEAAES20, Vei82c, VY21]. appended [MN04]. Application [MR12, Pat81, ABJ16, ALP94, AB22, BHY22, BKL97, BC91, CDJ13, CH06, Che91, Esh22a, Fri75, JKD16, Li21b, lLZjZ19, LC18b, Ma22a, MCIM21, Ng14, PMP22, Paz84, Rau18, Sch22, SR17, So99, Wan79, WSZL17, ZLS11]. Applications [Fur08, Gor83, JKZ20, Mer92, Whi80, AKR23, AAPR19, Ati20, BAN12, BLZ11, BBP21, Bru79a, CGMR11, CV94, CKR18, Che92, CBZ22, Cro08, CCLN16, De 82b, DeV22, DK23, DW12a, DHKSH08, Dra17a, FJS14, FLS88, FJK98, GCP21, GZZF23, GKS01, GHK16, GL16, GLS16, Gur98, HW13, HL18, HM19, Iov15, IAM23, JZL20, JKL22, KM19, KMRS89, KJF23, KT06, LS16, LL16b, LPH17, LZB23, ML21a, Mai11, MH22, MM20a, Mer94, MS12, Mou04, Mou13, Pro22, SRW07, SQH15, ySpW17, SL95, Son14, TZ19a, TZ19b, Ter22b, Ter22a, Tia01, Tia11a, Tia11b, Tia11c, Uhl79, Wan80, WH12, WL20a, WLL21, WY21, WS23, Wit11, YGÇ19, YZZ20, YSY17, YW09, ZZY16, Zha15a, ZMG20, ZPCZ16, vD96]. Approach [ASZ02, ACMR17, Bar07, BS17d, BAN12, BZ22a, Bur11, CKLP13, CZ23, Cla92, DP98, DS22b, GHL20b, Hau12, HRMP95, KMH19, MT22a, MVPS14, MZ22, Mou04, RST22, RS23, SK15b, TX13]. approximable [AGV19]. Approximate [KN22, LK16, RS23, DK23, GLL21, How83, MMR07, Pas15, Sol14]. Approximation [SAK22, TD16, TD19, CFSGL18, Che91, CJ12, CDF21, DS23c, Gog21, Hac22, Hav88, HUZ17, KS17b, MW21, TR21, UZ16]. approximations [DTSI15, DH16, Hal86, SF17]. Araki [TWK22]. Arbitrary [Bom02, Mou22, AAB20, BM12, CH22b, CM07, DFPS23, DM19, Ful76, GS01, HWZ18a, HRT15, HZS22, KW11, LLC22, MY12, Mee80, MW21, Moh16b, NJP19, Pab22, Pea84, Uhl79, WM17, Wan17b, ZHLZ13]. arcs [GW08]. Arctic [BS17c]. area [MR12]. area-preserving [MR12]. Arf [Bae84]. arguments [Neu83, Raï14b, Raï14a]. arising [Dag23b, DWL13, Lon76, MO22]. Arithmetic [Sur98, GHT11, JKL22, ST94, SY23, TLC16, XZ22]. arithmetic-geometric [JKL22, SY23]. arithmetic-geometric-harmonic [XZ22]. Arithmetical [Lor00, BL93b, Hau12, Hon99]. arrangements [CMM22]. arrays [Bre12a, BS13, GP74, MB22b, Sim83]. articles [Ano97b]. articulation [BLP12]. Artinian [PRA22, HPV84, HPV88, LQ95]. ary [Bed11, GR10, KP16]. Aspects [Whi74, MM06, WW01]. assignment [BHH86, CFP10, Gib87]. associated [Aln11, AO97, Ano22a, BCI13, BL93b, BT15, CMR20, Caf13, CLX22, Cha92, Cha93a, CKN95, CKXX23, DPR22, DP00, Dia90b, DJL94, Fal97, Gam18, Han17, Heo16, HS88, HT92, HK16b, Hon99, HH14, JLDS12, KH05, LY22a, LD22, LTT11, MH20, MP73, Moo95, ME18b, NSW21, OWTZ21, RC22, TLC16, TJ21]. associative [Bas17, Bas19, Gol07, Hon19, Hua96, KKL20, KKL21, MP21, RE22]. assuming [Wan20]. Asymmetric [KMN22]. Asymptotic [Ab23, Har78a, Hou09, BSA17, CW17, DHB07, Geh15, Min80a, MFPSS12]. Asymptotically [La 19, BG23a, WS23]. Asymptotics [WS23]. Atkinson [BBBS09]. attached [ACMR17]. attaining [GR23]. attainment [Sai21]. Audenaert [AkK17, Ala18, Zou20]. Author [Bal76]. Automorphic [Pea84, LK21]. Automorphism [HNB22, MW16, XWT22, ZWM17a, AF12, RE22, TW17]. Automorphisms [BL11, DM13, Foš05a, GWZ16, HLZ02, Leg06, Lud19, Maj22, OWX09, Wan16a, WWS19, YWO07, ZWM17b, ACI15, AKOZ20, AAUN22, BFJ12, EL21, HS17, LK23, Liu18a, MB22a, Ter06, Ter09]. Autonne [HZ12]. autonomous [MR12]. average [DH16]. Averaged [DS23b]. averages [Szw18]. axis [HHC12, Nas75]. Azumaya [SX12].

B [Ano79b, TT78, Tia21]. B. [TT74]. BA [AP74]. Backward [AM13, BR04, CLW15, SWL13, MHM22, ZZT23]. Backwards [Pan17]. Bad [Hal91]. Baker [BB14, DST91, NST89, BBM20]. balance [WLF22]. balanced [BBG22c, RBG23, WW01]. balancing [Kir14]. Ball [LCH22, Sch92b]. Ballantine [Rad80]. Banach [Bal23, BB17b, BB14, Boa13, Boa19, BS18c, CS21a, CS22, CZZH22, CFZ23, DPR22, Dra22a, FU20, FGMSS19, GT14, Gha12, LYZW12, LCL14, LL18, Mos21a, Sai21, SRBB22, TN22, Tom19, Vel18]. Band [Elo18, ELG88, EL90, HM76, HM77, HM78, PS20]. banded [BdPN14, DdF22b, Woe93]. bandwidth [HRMP95]. Bannai [HZG14]. Bannai/Ito [HZG14]. Barbara [Ano79c]. Bargmann [CDJ13]. Barthe [LCH22]. Bartholdi [MMS21, Sat11, SMM16]. base [LLS10, LLL09, MM11, Mum86b, WYM11, YMYS13]. base-listing [Mum86b]. based [BNKJ22, CPS22, CX23, CW20, CFZ23, Cro08, DK23, Guo20, GHT16, HZV23, KC16, Kir09, LZ15a, LZ16, Mar18, Mit92, Roo20, SKP18, Sun20, WZY22, ZV19]. Bases [Pet09, Wil99, AGS20, Ami11, ALPM20, BH90, BN88, BD14, CGNT15, CW89, GSS09, HJ20a, HS89, HS91, Hol95, HK13, KCBR87, Knu85, LHS^{+}21, LPS21, MC86, MHA17, MG22, Nor08, Pat17, SSG12, Vo83c, WG91b, WCS^{+}21]. basic [AHJ87, Bra86, CW90, DS19b, GV87, Sal14, Vei82a]. Basis [SLH20, AC13, Aln15, AGP22, CDJ13, DP00, FJS14, Laf80, MID17, NSW21, WG91a]. Basis-free [SLH20]. Batcher [CW91]. Bavula [ML12]. Baxter [ZLS22, DBG18, GG22b, LW20, ML16]. Bayesian [Joh82a]. BCR [Haj18]. be [Gam18, GHL16, LPS09, LLF18]. before [Mac09]. Behavior [Har78a, Phu80, RS08b, Sta84]. behaviour [MFPSS12]. Beidleman [DH20]. being [CGH15, LP23]. Belardo [Ano16b]. bell [Kar23]. Berberian [Far94]. Berezin [HLB20, YK22]. Berkani [AZ22]. Bernoulli [CP16, KS14b, PSA14]. Bernstein [MM92, Wör20]. Bessel [BHY22, DKM10, MS88, Ota20, Xia20]. Bessel-type [DKM10]. Best [Pat03a, CLL14, GKS01, NASR19, SSGP23, SF17]. beta [Sal12]. Bethe [ACMR17]. better [UB22]. Between [Cai93, CD03, BK21, BD12, Bén10, BDvW12, CMR21, CHL18, Cho10, DM20, Der18, GZ76, Gut23, HK16a, Haw05, JS20, sL19b, LT88a, LZX19, LY22b, Lim79b, LLD17, Lon73, MDP22, MP98, Mos19, Pat07, PT87, SH21, Sai21, SRT22, TZ06, TOM88, Tho92c, Vel18, Wal79a, WOY06, Wan19a, Wat86, WL20b, WW12, ZH22, ZS05, ZCZ12, Zha13, ZBW16, ZCY22, ZG08]. Beurling [HLP23]. beyond [BS22d, GLS13]. Bezout [SK15a, SK15b, WY18]. Bezoutian [CK84, GB90, LM85, Ota13, Wim88]. Bhatia [KKS21]. BHH [Ikr15a]. bi [AAB17, BB17a, GS21a, HS16, LW23, Wer86, YLW22, YLW15, YH22]. bi-block [HS16]. bi-Cayley [LW23, YLW22]. bi-complementarity [Wer86]. bi-elliptical [GS21a]. bi-Hermitian [AAB17]. bi-homogeneous [BB17a]. Bialgebra [Sin21]. bialgebras [BGLZ18, DBG18, GZW20, ML16]. Bicommutants [Mag13, Mag14]. bicommutative [BCZ22]. Bicommuting [Dra17b]. biconnected [Ste94a]. Bicyclic [Hou01, FTZ08, FDD14, LW22a, Nat13, Tan12, TL13, XYH22]. bidegreed [Bel15]. Biderivations [CC19, CY22b, CWSZ17, CSY18, Tan18c, TZ19b, Wan16b, WT23, CCC21, CW23, FD17, LS22a]. bideterminant [ySpW17]. bidiagonal [AC13]. Bidiagonalization [CP97]. Biggs [Koo95]. bigraphs [GVS96]. bihermitian [HW14, ZLZ^{+}22]. BiHom [BMN22, LCC19]. Bijections [HS17]. Bijective [Blo95, Rem82, Shc21, Slo13a, WZC11, Bay22, CW16, Hua10, RW84, Rem90, RŠ06, RŠ08a]. Bilinear [MR75, BCT23, Bre14, COK22, CL80, GQ06, HMN91, IKR12, JSY14, MS82a, RS85, Ruc89, Sai21, Wag87]. Bilocal [KZN22, NZK21]. bimodules [SDVWT16]. bimultiplication [BM22a]. Binary [JT04, Bed10, BLSV12, CO12, GS20b, KD23, LLY22, Tau81]. Binet [Bay22, Bun91]. binomial [BCDF09, FW73, Gar73, Rei77, Whi80]. binomial-type [Whi80]. Binormal [TMH21, KKL18]. Biography [Mar03]. biorthogonal [CJ12]. bipartite [BP16, BS93, GLS16, KS22a, KS23, LZ11b, Lu22, MR19, Obo16, PGLM00, SSfW15, zSqSZ16, WHG17, YHS13, ZG08, dLMO17]. bipermanents [Cla82]. biquaternion [YTL20]. Birkhoff [AV22, BBK22, CP13]. bisection [UZ16]. Bishop [DKK^{+}20]. bistochastic [ME18b, OP13]. bisymmetric [CC17a, DH11, WLL19, ZHZ11]. bivariate [BP18, SK15a]. bivectors [Wes81b]. BK [EB17]. BK-space [EB17]. Block [DMQV23, ES98a, EMS02, JKZ20, Liv94, Piz87, Tia02, TSYB82, Wei98, YC19, Zim89, Aud15, BH23, BS11, BLP12, BR14, BGdH92, BLZ11, BZZ10, CM92, CdM15, CTW15, Cho18b, Cho10, CICX09, DM22a, DTS14, DD10, DHS95, Doo16, FS22, FH89, FLT22, FG22, GS21b, Ghi22, GR91, GR93b, HS16, Hua01, Ikr15b, JK22, JK10, JSE90, KT21, LWC22, LP22a, LW12, Liu22b, MHH96, MS10b, Ner22, Ric87b, Tam94, TL12, Tis93, Uhl79, Xio23, XSH18, YLX20, YQ90, Zha20a, ZBW12, ZD18, SYZ22]. block-diagonalization [MS10b]. Block-matrix [Liv94]. block-shift [Tam94]. block-Toeplitz [GR91, GR93b]. block-triangular [CdM15]. blocked [Con90]. Blocks [AY74, CP18, CP21, GS21b, KS19a, Sjo93, Yua98, Zha20a]. BLUPs [TJ16]. BMV [Bur11]. Board [Ano11a, Ano12a, Ano13, Ano14, Ano15, Ano16a, Ano79a, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano84a, Ano84b, Ano84c, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e]. board [Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f]. Bogart [Mer79b]. Bohnenblust [JRMFMASS16, MFPSS12]. Bohr [MR10, SN23]. Bollobás [DKK^{+}20]. Book [Ake80, AR80, AK78, Ano83a, Ano89a, BFMM73, CST75, CS79, Con94, DW76, Doz79, EA77, FAR78, Fan81, Fan84, GMS^{+}76, HUSG74, Hum83, JTH77, JM77, Joh83, Kap80, Keg87, LW78, Lei81, Mer98a, MB76, MS80, MBSL74, MGK^{+}74, Pie92a, PA79, RY73, Sch92a, SB75, Sen81, Sha95, SA77, Sko85, Sta94, TB76, Tho81a, Tho81b, Tho82e, Tho83, Wat84a, WTM78, WNM74, Yaq80, ZF78, MT83, Tho82b, Tho82c, Tho82d, Tho82f]. boolean [Bo01, GPK94, HKS94, KP92, Liu92, LB99, BS22c, KS12, LT14]. Bootstrap [BC06]. border [CJ20, DM19, Hac22]. bordered [Tia10]. Borel [Chu84a, Shc21, WOW09, YWO07]. Borg [KK18]. Borg-type [KK18]. bosonic [Ter22b]. both [DC20]. Bott [CLX03, GZZF23]. bottleneck [AD17]. Bound [Cal03, Pat03a, Zhu15, Bal14, BGV11, BD84, CR15, CDK20, CJL22, Gil17, Guo20, Kir00, KMN01, Kir09, MF20a, Mar97, Pat09, QK92, Raj09, SC21, Sol00, zSqSZ16, TD16, TD19, WS18, WWW20, ZBW16, ZWS20]. boundaries [Ven19a]. Boundary [NS85, jAS15, jAS21, BD88, BGLSSS20, Cha18, Che14b, Che16, CN07, CN08, ERS12, GMZ17, GW08, LLS14, LLS16, NNT95, Nak95, San14, Spi00, TG09, Uhl14]. Bounded [NP73, Bou18, CLN13, Dra18a, DK21, GMV18, HJ20b, Hos22, HL20, JKD16, KK22c, LS22c, Loe01, MRS10, MBB22, Pep19a, Pop17, Rho94, SPMR20, Sai21, SRT22, dNV19]. Boundedness [SXR18]. bounding [MH19]. Bounds [Cla92, GSS09, GRSW92, HHY22, Joh22, KS07, LLS10, LCLL22, Lia23, LL15, LL10b, LSZ86, MT22b, Min80b, NS15, Rad12, Rum97, Sch92b, VM07, Wal11, WMFW21, WLM21, YJ21, YHS13, ADJAH22, AS12, Ben90, Bén11, BR19, CQW16, CW10, CW11a, CR77, CGH14, Dat16, DTS14, GLL18, GHL10, GM21, Hav88, Hon99, JKB17, Kar83, Lás91, LZ11b, LGA^{+}13, LL17, LS22b, LS12, LS13, LZ17a, MC89, Mel13b, Mel19, MZ15, Men22, Mer81, MNS22, Neu84, PLH01, Pat87, Rho87, Rho90, Rho94, Rob75, ST94, ST15, Sha91, SWL13, Sou03, TD14, Tia17, TW19, WfMR16, WW20a, Wan22a, Xu22, YZ14, YSY17, ZYL13, ZVL18, Zho16]. Bourin [DHHK23]. Bourque [KMH19]. Box [BMFPBSS21]. Boyle [KNX95, KNX93]. braces [Rod22, Zhu22]. brackets [Kay17, Zus22]. braid [Abd09, Lip87]. braided [Zhu22]. Branch [JS08]. branching [CPT15]. Brauer [HK13, LLL15, LZL16, Mel10]. Brauer-type [LLL15, LZL16]. Brègman [Sol00]. bridge [Ste94b, Wil93]. brief [BL92, Far93a]. Broadcasting [YGZC22]. broken [BS95]. Brouwer [ABGP21]. Browder [AFM13, BHC16, BCM18, ZZZ16]. Brownian [Suc22]. Bruhat [FF21, TT15]. Bruhat-graph [FF21]. Bruno [Wen06]. Building [RRR^{+}22]. bundles [CP10, KL92, Lla96, Per19, Sch22, SL95]. Bunyakovsky [Dra11]. Burnside [Law19]. Buzano [Dra17a].

C [Ano97b, Day97, Don91, GKO22, GPR84, Min74c, New97, GT08]. C* [JKH20]. C*-module [JKH20]. C. [Rub89]. cacti [SYZL17, WDJ10]. cactoid [DM22b]. cactoid-type [DM22b]. calculating [Tia18]. calculation [Bel83, SW81]. calculations [Har96]. calculus [BBT07, Cro08, Net22, Nie14, Par78]. Callebaut [BM15a]. Campbell [BBM20, BB14, DST91, NST89]. Can [CAP22, GHL16]. Cancellation [Wag87]. Canonical [Afr73, BS13, HS09, Sta14, Sta15b, Wil93, CCU21, CFP21, GAGM18, HL09, HWZ18a, Hor04, LK16, NSW21, PS91, Uhl79, You80]. capability [JN22, PN22]. Capable [NJP19, Joh22]. capacitated [Li85]. Capelli [Reg93, Wil81, Yam18]. caps [Kes83]. Carathéodory [CH12, Sch92b]. cardinality [Hil18]. carrier [Ste94a]. Cartan [CC19, GLOT13, SZ19]. Cartan-type [CC19]. Cartesian [QD85, RRN22]. case [Bal18, BFZ07, Cai93, CH12, DLL19, FJ09, Gut01, GLS13, JLL88, Laf78, Lei83, PA10, PM23, Rai87, SvW14, tHK17]. Cases [JN90a, BCdHP10, Hia94, Laf78]. Cassini [Mel12, Mel13a]. Catalan [Ilk20, LM19, RRC23]. Catalan-like [LM19]. categories [BF96, Zhu22]. category [Nas74, Smi83]. caterpillars [APEPPS23]. Cauchy [Bay22, Boz98a, Boz98b, Boz99, Bun91, Dra11, GT11, HKO16, IK22, Oku91, Sol06, zY11, dCdSA22]. Cayley [Are22, Are23, Bed11, CW81, CF21, CFL^{+}22b, CFL22a, FSFSO^{+}14, LW23, Lud21a, LCXC22, LCWW22, NC22, Paz84, QX23, Sch95, TM19, Van95, WC22a, YLW22]. cb [Gog21]. cb-norm [Gog21]. CD [KK22b]. CD-algebras [KK22b]. Cebotarev [New76a]. Cebysev [Dra10, Dra16]. center [For79d, Joh22, Pat07, Zim98]. Central [BD13a, CKKK22a, Hon22, KKL21, UG21, XZ18, CR18, CN14, Kim13, Lip96, PP22, Per15, ZHZ11, ZWW22]. centrality [REG07, YZB22]. centralizer [ASAS79, dSP11]. centralizers [Bre15, DB14b, Ghi22, Liu16b, Liu22a, LG23]. Centralizing [LWXF15, XWF15, Wan16d]. centrally [LT22b]. centre [AFJK17]. centres [HLYZ22]. centroid [AFJK17, Pat07]. Centrosymmetric [CMP22, LZ22]. Certain [Lev74, MR74, RJ04, ARK22, AL21, AFW22, BK83, BWF83, BP98b, BS18a, BD85, BDH23, BFI22, Cha88, CCW20, CFL22a, CN08, Düf97, FV97, FJM00, FX12, Fei75, Fri90a, Fri97, GZ76, Gol79, HSX16, Hem93, HT92, Hul81, JC22a, JN22, JN00, KM19, Kit91b, LT88a, LT90a, LP93, LLWX17, Lim15, Lon83, LSZ86, MS83, MSO83, MS79, MHA17, MO13, MGA17, MM76, MSS18, Mou13, Ota13, Pat77, Pat89, Raj09, RA21, Rho94, RF19b, Tom19, WHX08, Wea83, vR11]. Cesáro [Geh15, RF19a, Roo20]. CFI [DC23]. Chain [AADdF23, AAKS21, AS11b, MS10a]. chained [LMLL17, SC21]. chains [BC94, Gue17, Gue19, Gur98, HWX22, HH14, Kir14, LT12b, ME18b, NR23, ORT15, PT10, Sie99, Tan10, TLL11, TLC16, XL11]. change [BCdHP10, Har81a, JLDS12, TLC22, TJ21]. channels [Ter22b]. chaotic [FJK98]. Chapter [BLP92, LT92b, LT92c, LT92d, Lim92a, Loe92, McD92a, Pie92c, Pie92b]. character [DF87, FJK14, Fon89, HT91a, SS98]. characterisation [BKS12, Laf84]. Characteristic [Ben74, BH74, Fit94, GB77, Ger04, KNS96, KL92, Mir81, Nom73, Ric81, SL95, Tre87, AFJK17, Ami80, BS15a, BP98a, BG79, CNW18, DP09, DM19, Ful76, GMZ11, Hai88b, HZ21, Li84, Lit17, MR18, Nor95, Nor08, Pat07, Prz92, RC22, Wad85, WW12, Yao22, Zim98, dO75]. characteristics [Abd20, AHK92, HS89, HS91, Koh11, MMRR16]. Characterization [AH13, BP12, CMR20, CSS21, CDPS23, FJK98, Geh15, HDJT97, HOT15, Liu16b, MPD22, MDP22, MB22a, Sed11, Šem97, WM17, AAS14, AXH15, Ben78, BS15b, Cha77, – Dok73, Dom14, FA19, HH89, HLZ11, JK10, Kir04, KL21b, KKL18, Koh11, LM95, LH13, MH16, MS83, MW23, MHA17, McD82, NP80, Nik22, Pan15, SPMR20, Sel81, Shr90, Ste94b, TL13, TG09, TWZ16, WSH^{+}11, WBZ14, WW15, Yua98, ZB13]. Characterizations [AH10, An18b, BT08a, De 78, DOP22, FLT20, FdSV16, FXY22, GWW20, HLH18, JQS13, KCSĆ21, LK21, Li95, LZ11a, LC18a, LG23, ML21a, Ma22a, Pan13, Ruc89, SRW07, SZWB16, WL15, WZD16, Wer94, YZ13, ZXL16, ZP19, ZCP18, BJ96, BM18b, CZWF23, CM22, Dic84, ES98b, FZ21, GZZF23, Iop19, LST89, SSfW15, lWW15, XCM20, ZWL15, Zhu18, ZIC22]. Characterized [EMS02]. Characterizing [GKT97, Joh11, LZ14a, Qi15]. Characters [Leh76, Nie22a, BG79, Bri89, Fon92, Hol04, HK16b, Hua90, Kim91, Mer78b, Mer83b, Mit10, Pat93, Reg87a, Sil84, Sta84, Zha21]. Charles [Ano79b]. Chebyshev [Ma’15b]. checking [RR95]. chicken [Pan20]. Choi [KMN22]. Choices [AP74]. Cholesky [DO20, FS22]. chordal [JL93]. chordality [BJ96]. chromatic [SD20]. circle [BL21, Dat80, LB19, Tho92c]. circles [SX23]. circuit [DP98]. circuits [Mer83b]. Circulant [GN02, Sjo93, Wat85a, BSA17, Bom02, BKMMG22, BT15, BH20, CH22b, DPP17, FL18, GKS01, HFSZ22, MSS19, Mou22, Pal21, XJW17, ZQZ21].

Circulant-style [Wat85a]. circulants [GK18, Ikr15b, New80a]. Circular [BJZ13, GWW16, Tur16, CL13]. Circularity [Tam94]. Clarkson [Kan87]. Class [AHJ20, Ber02, BR04, CL22b, MSZ04, jAS21, BK21, BL10, BH23, Bas19, Bea83a, Bea83b, Bea84, BR01, BM89, BZZ11, Cha78, Cha96, CDPS23, Cla92, CFP21, DD10, DRY22, DS09, DWL14, Ell95, ES76a, Fal97, FNBR96, GMW86, Hon22, HM76, HM78, HM22, IOMA21, Joh22, JN22, JP86, Jun88, KM19, KM17, LT93, Luo19, MRT16, MC82, Mer82, Mou13, PSMA16, Pal21, Phu80, PSB23, Pok87, Rei75, Sed11, SGS05, Sol80, WG91a, Wat82, WY18, Xio23, XW22c]. Classes [GK02, MR74, ARH15, Ano22a, BS18d, CT22, CL86a, CL86b, DP00, DF95, FJM00, Fie95, GCSC20, Hil76, HT92, HK13, HK16b, Hon99, KSMB16, KCBR87, KS76, Kur04, Lim76a, Lim76b, Lim82, Lim90b, MS08, MS83, MSO83, MC86, MH20, MM76, Mer78c, Mer79c, Mos20, ND05, Pop14, RH11b, Rob77b, RC18, Sil88c, Sil90, Sil93, SS96, SMM22, yT86, Tau81, Tra89, Tra90, WG91b, XHL23, Zho16]. Classical [CK20, HZH20, CW81, CGY13, CKL16, CJK^{+}13, Dic84, DBG18, Gra10, GL16, GLLM21, Hua16, Mar89, PLLL20, Shc21, Ste82, Tam97]. Classification [CLOK13, – Dok83, GR93b, Hai88a, Hai92a, JST18, Kar90, LK23, MR17, NR23, RH11a, Rom89, Shi18, Tia21, TJ21, COK20, COK22, CF15, GKO22, Hai97, KK22b, Koo95, Lia80, MC19, Sol80, SLH16, SLH18, dlP20]. Classifying [BDV19]. Cleavages [dlP09]. Clifford [AF11a, AF11b, AF12, Ang00, Chi78, GR93a, LS16, Per19, PV20, Shi18, SL15]. Cline [Mos15]. clique [LL15, LZB23]. cliques [AS11b]. close [KK16]. Closed [BN22, BSA17, AL22, HS83c, Li92, LD19, LP22c, LP22b, Muk20, PY82, RAN22, Wat85a, Zha14b]. Closedness [CLP22, LP22b, LD17, SXR18]. Closure [BZ89, Lim00]. closures [Koi21]. clover [WW15]. clusters [LY22c]. CMP [CZWF23, Ma22a, XCM20]. co [Mil15, SMBB^{+}15]. co-flag [Mil15]. Coalescence [GM90]. coalgebra [HT12, Pan17]. coalgebra-graphs [Pan17]. Coalgebraic [Wis12]. coalgebras [ALP22, Iov15, ML16, Pan19, Pan20]. coapproximations [SSGP23]. coaxial [Vei83b]. cocharacter [Car88]. Cocharacters [GMV18, Car06, Di 88, Rem88]. Cochran [TS05]. cocommutative [LW20]. cocycles [Dra21a]. codes [AS15, CHFC19, CW93, GTLN18, Hou09, KLZ10, Koo95, KP14b, La 19, LL22a, LLY22]. codimension [COK22, dB16]. codimensions [Iop19]. coding [Gor83]. coefficient [BV16a, DH86, HT12, LLC22]. coefficients [AEG19, BB19c, CR22b, – DL92, – Dok93, GPW22, MH22, Sen83, Tan12, ZZ15]. cographs [Gho19, LST23, Tur18]. Cohen [Cam14]. Coherence [RST22, CC23a, CP16, DO19, YGZC22]. Cohomologies [ZLS22]. Cohomology [AC19, Zha15c, BCS18, HH22b, Hou17]. cohomology-decreasing [BCS18]. coincidence [BT22b, Rom22b]. Coinvariants [CNW18, MCIM21]. colength [dNV19]. Collatz [Zar01]. collections [LR89]. color [KP16]. colored [SGS05, SG09b]. colorings [Dia90a, SG09a]. colour [Zha15c]. coloured [KP12, KS22b]. colouring [LZB23]. Column [BZ90, GZ99, SK03, BT13, BCdHP10, FV97, HKS94, LSOvdD95, MSS18, Mum86b, Sal79, SS12]. columns [MNS22]. comb [Gol21]. combination [BB06, BZZ10, For87, Pat83, PUÖ15, RRC23, Xu15]. combinations [BB20, BT08b, BLZ10, BR10, BSÜÖ13, Den14, ES98b, FD12, For88, IK18, KKR12, LD17, LWY11, LWB11a, LWB11b, MM90, Tia11a, UÖÖ15, UPÖ16, ZCW21]. Combinatiorial [Tra90]. Combinatorial [AD17, BM18a, BM18c, HS91, Tra89, CW89, CRY95, ER90, RS18, Reg87a, Wan79, WW01, Wil73, BM19b]. combined [BGGS17]. Comment [Gro96]. comments [TT78]. Common [BDK^{+}05, GKLR82, KD23, Son14, Bar80, BKS12, DMdO92, JP15b, JN90b, TT06, Wil87, XW22a, Zha04c]. commutant [BBEL18, JS94]. Commutation [XHL20]. commutative [Bas17, Bas19, CL22a, CZ23, Coa99, CAP20, DM18, DO20, FLT20, Fre76, Gar21, Gus81, KK22a, KT19, LCL13, LL10a, lLpW16, Mar12, Mas10, OWX09, ySpW17, ySpWQ21, SM19, Son22, SZ19, Tan13a, Tan18a, Tan20a, TZS19, Tia11a, Tia11c, WOY06, WZL13, Wat85b, YLW22, YW05, YWO07, ZWZ10, dTS20]. commutatively [AL22]. Commutativity [DŠ04, JR93a, JR93b, Kun99, BZ89, CY22b, DS13, DK16, Foš05b, GR93c, HQ18, JQ15, LL11, Liu15, MS10a, Mol09, Pet97, QH17, Slo16a, WZL13]. Commutativity-preserving [Kun99, CY22b]. commutator [BLR12, Fei79, GPR84, Gur79, GL83, MW74a, Nof15, Obl12]. Commutators [RR02, Smi76, BKL97, BH15a, CLR81, DP18, Est80, GG22a, Gil17, HS17, Hou22, KL12, MS95, OR09, Rob77a, Sha18, Slo13a, WZC11]. Commute [Ric81, Tur02, BT08b, HS83a, KE82, Sin76b]. Commuting [BD22a, CKT20, CW22, Dra13b, Fra13, SB03, AL21, BH90, Bea78, BCM18, CM19, CC19, CW23, CWSZ17, CSY18, CKL16, CK20, CL21, Cos21, DKY22, EGMS22, Fra17, GJR22, Gur92, KT21, KKY22, MS10b, TCI13, UPÖ16, Wad90, Wan79, Wat77, Zel75]. comonads [EGT16]. comorphic [Bre21]. Compact [DC20, DH17, Kan16, Kit91b, MT11, Pep11, RRN23, Rao18, Ruc89, Tam97, XS18]. companion [BT90, CC21, ERS12, LC20, MR12, Mel19, PH97, dPN81]. compared [Neu74]. comparison [BS92, Blo95]. Comparisons [JM13]. compartmental [GOS22]. compatibility [RST22]. compatible [dSP16b]. Competition [JKY23]. Complement [Zha04a, AA21b, BT06, CPLH17, GAMA23, LM98, Neu81, PRS03, SL95, ZJC^{+}22]. complementable [GS20a]. complementarity [CLW21, DN18, GLL18, HZV23, JDD21, LZ16, Li21a, LLV18, MM06, MN96, MB22a, ND13, SC21, WS18, Wan22a, Wer86, WL16c, WZY22, ZVL18, ZV19]. Complementary [BM15a, FS14, MF20b, Sta20]. complemented [ACO21]. Complements [Nie10, Bap88, BS18b, Bru79b, CS95, FJ15, Gib92, LHZ22, PV23, ySpW22, Smi83, WZ97, YLW22]. completability [JJK95]. Completable [YD14]. Complete [GS95, JN14a, TL13, ADHM19, EM95, HZ11, HWZ18a, HJ77, IS16, Obo16, SPMR20, Shr90, Tan13b, WOW09, ZWL15]. Completely [AKS18, DJL94, AG15, BSM98, DJ98, GP22b, HJ19a, HJ19b, HJ20b, HH14, JP15b, Koo95, Lei95b, LP11, MP16b, PV20, SM22a, SBMP16, Stu89, SL17, ZX16, Zha20b]. Completeness [Den05, LBV13, IS15]. Completion [CJ20, DS22a, BC05, BZ90, dAB21, BR01, BS18c, CKP16, Cao19, CKKY23, CP18, CP21, Der16, Der18, DS22b, DJ98, DJK08, DCLW19, Esh22b, JJK95, JN14a, JTU00]. Completions [GMR81, Lud21a, CKGP21, GKV89, GRSW92, JR84, JW91, Woe93, Woe97]. Complex [Ano79b, Bal78, CKGP21, CN04, ETM22, JKS79, KFR17, LPX23, Mac09, MY12, Mir81, Sah21, Wal73, ABC18, BH23, Bal14, BPR20, Cha90a, Che22a, CT09, CH87, CL74, – DL92, – Dok93, Gór22, Gut23, HK21, HHY22, HZ12, HQ18, JVMCPR19, KKL22, KL19, Li84, LT88d, LM93, LY22b, LHS^{+}21, LZ19b, LT07, Mag22, MF92, MKS22, RH11a, RH11b, RŠ06, Rom89, RS23, Sin82a, Sin84a, ST91, ST96, Soa08, Tan13b, Tan18c, TMH21, Tie16, WL18a, Wes81b, WL23, ZH22, tHK17, KSU22, KFR20]. complexified [Shi18]. complexity [BZ89, Con90, Har85]. complic [Bal79c]. component [BCS14, Das17, Fan73]. components [Bén10, KF98, Lla96, Mum86a, NS92, NS93, Vo83b]. Componentwise [Rum97]. Composite [Que08, KW93]. Composition [Tau81, LPX23, TH80]. composition-differentiation [LPX23]. Compositions [AY93, BM07, Wil73]. compound [dCdSA22]. compounds [BC91, TMOvdD93]. Compressed [DJS21]. compressing [EM15]. Compression [AP04, GW04, JM85]. compressions [Far94]. Computation [CN21, KK16, SKP18, BL21, Ma’15b, Miy22, Uhl14]. computational [Zel85]. computations [DOP22, Net22, Uhl20]. Compute [YZ16, FP17, McF79b, SX22, vD96]. Computer [MP87a, KCBR87]. computer-generated [KCBR87]. Computing [Car06, CD15, HJ20a, Ous15a, PS91, Rem88, Roh00, Sta15a, Uhl13, DK23, FHZ17, GA85, GK18, KPL^{+}09, LL22b, MZ22, SN22b]. comtrans [BE20]. concave [AHK19, AA12, JK22, Tam02]. Concavity [BD85, Mat90, Sra20]. concept [SST22]. concepts [AR15, Shi15b]. Concerning [Har96, Sin80, Sin81a, Sin93, ST99, Tau74, BK87, DS23a, GAMA23, GT79, Gyi78, KNX93, KNX95, Mou13, Sin82b, TT78, TH80, Zou12]. Concomitants [BB96]. condensation [Cal91]. Condition [KK22c, LPS09, LZS23, SLW21, AADD18, CC98, DS13, DLQ18, DKK19, HB83, Jam88, Jam95, LLW^{+}11, LL17, Liu17, Pal79, Rou98, Zim98]. Conditional [CMM22, HS12, NS22, BG16, FHZ17, JS19, JE21, Joh82a, PSW18]. Conditionally [Józ15]. Conditioning [WY21, Que08]. Conditions [AKK^{+}23, Gam18, Ibr00, JKH20, JNT96, Lim92b, PH03, AE10, jAS15, jAS21, BMGMC15, CHQ18, CIRW21, DS23a, DJK08, EOvdD09a, EOvdD09b, GMZ17, Gue17, Gue19, HB83, Hua16, Ikr15a, LCLL22, LP23, LLF18, LT22a, LOO22, Lu22, Mar90, Mat11, Mil22, Oku93, Pet09, SQ15, Tia05a, Wan82, WYSL22, XZW22, ZW17b]. Cone [KS02, LMR00, And96, FU20, Hil07a, Hil11, MB22a, Pat03b, TB92, YH05, Zha20b]. Cone-Systems [KS02]. Cone-transitive [LMR00]. Cones [FH73, HS88, SS79, AGP22, Bar74, Bar78, BC80, BKS12, CLL14, Hil07b, JP17, NS85, NST91, Pro22, SH21, Tam77]. conference [Ano79c, ETM22, LRT08]. configuration [NS09, Nat13, SG06]. configurations [CMM22, TL13]. confluent [MZ22]. conformal [aCC17, CSY18, DH18, GZW20, Hon19, Hon22, TZ19b]. Congruence [LT92a, Lei98, MA75, ACI15, Bot97, FJ01, FJ09, GL97, HH89, HS09, JR09, Lei91, Lei95b, LM93, LPWX10, Tam85, Tho80, Wat90, Wil83]. congruences [HH85, Ikr15b]. Coninvolutory [SX23]. Conjecture [Kit81, Wan07, ABGP21, AYT83, BMD87, BD96, Bri89, Bur11, Cha89, Cha83a, Cha84b, DJ89a, EM95, EF98, Esh17, Fie84, For79c, Fri78, Fri82a, Gib87, Gou13, Gum15, GLS13, HS93b, Hu87, Hu94, HL10, Jia92, KM84, Koo95, KMH19, Law19, Lon80, Lon81, ML12, Mal89, MR07, Min74d, Min82b, Moh15, NT94, Obo22, PS22, Pat89, Pat93, Rad80, Ric85, Ric87a, Rub89, SYZL17, Sil84, Sin81b, Tam86, Wan77, Wri93, Yu96, Zou12, dF05]. conjectured [Tho86b]. conjectures [Ash16, Cha90b, GSZ22, Kop94, KAE22, Meh84, Shi16]. conjugacy [Ell95]. conjugate [CD01, JTW17, YCZ21]. conjugate-Hankel [JTW17]. conjugate-Toeplitz [JTW17]. conjugation [Gal01]. conjugations [CDPS23, HL21]. Connected [BB19b, KS17a, RP90, ABP22, ACMR17, BH20, LS22b, LOO22, MT11, Put82a, Put82b, SS23, SSfW15, SHS17, XZW22, YSZW22, ZZZ23, ZW17b]. Connectedness [MT11]. connection [KP14b, Lon73, MZ17, MP98, MA15c, ZG08]. Connections [Wal79a]. Connectivity [FKP03, KP02, ZG11, AA21b, AL21, BP98a, BLP12, DWL20, FHZ17, Fie90a, GZY18, KS19a, KN97, KF98, Kir00, KMN01, KN04, LCC17, LS21a, Vo83c, Wu05, YSZW22]. consequences [GT11]. conservative [Dag23b, Ilk20]. considerations [AdF21]. consistency [FHW16, lWY19]. consistent [Gib89, Sti87]. consisting [TL12, WC22b, ZWW10]. constacyclic [CHFC19]. Constancy [Bal79a]. constant [BF75, CDK22, FV97, For79b, Fre76, GLR82, KS93, Kut19, LSOvdD95, MH19, Wes90]. constants [CLL14, Che95b, Gol87, JRMFMASS16, Ma’15b, MFPSS12, NASR19]. constrained [KD23, LMR94a, LW10, Nak95, Pap13, SW22, WY21]. constraint [HM22, fLLyP15, ZHZ11, Zha22]. constraints [Din98, XY22]. Construct [GLL21, CG17, Slo14a]. constructed [COK19, LF16]. Constructing [BBCD21, GPM14, LV18, LF15, Sou83, WCS^{+}21, Wri22a, MMG15]. Construction [Ber86, Far22, HK16a, JW22, KW92, KW93, Mai11, MC86, SDVWT16, Smi05, Son22, Sun15, Wal75b, WW80, Bha93, CSG22, CC17b, CPKK22, Far93b, Foa79, GZW20, Hun83, JS08, JSW09, KP22, TT74, TZ21, Zhu22]. Constructions [BW15, FZ21, MTR11]. Constructive [Cro18, DP98]. contact [Zus22]. contain [Bas17]. containing [Che22a, HW75, Pat84]. contents [Ano92g]. context [Boa19, BDvW12]. continuation [CNW21]. continued [CR22b]. Continuity [Cam80, CJK^{+}13, GCP21, Mey15, BBX20, CP13, LLS14, LLS16]. Continuous [LP16, Poo17, Slo16a, TS21, ADRP14, BMFPBSS21, CHZ22, FZ21, GT08, HF18, JK10]. contract [BVW84]. contractions [FH89, Grz91, HS83b, MPD22, WZD16]. Contractive [CKKY23, XXZ11]. control [BL80, CF15, DHKSH08, DHKS16, Har94, Igl99, Lia80, ShDHG08]. Controllability [FKS22, JTU00, Sil95]. controlled [SA21]. conventional [HJ87, JN90b]. Convergence [Doo16, Haj18, NST89, Szw18, BB14, BBM20, CLW21, HK21, KV22, LW17, Li21a, Neu84, SAK22, Xu18, ZV19]. Convergent [CP81, Dag23a, SES21, YZH19]. Converse [JN90b, NS18]. Converses [Mar76, SV14]. Conversion [CA03, CD98]. Convertible [Ber12, CD01, HK92, HKS95, Kim97, KS83]. Convex [And04, ES98b, ES00, EMS02, Fri81, GP22a, JT95, AHK19, Anj17, AA12, Bar78, BPD88, Che17, DW92, DV18, For88, Fri84, Gil12, IOMAN21, IK22, JK22, KJF23, KD16, Li86, Li87a, LST89, LT89a, Li98, MM90, Poo94, Prz92, SSZ93, Tho75a, Tho75b, TI77, TT18, Woe08, dlP20]. Convexity [Hwa91, LP22c, ST96, AY84, Dra21b, IK20, KMRS89, Lék19, Li96, Man87, MFS22, NY23, SP87, Sra20, Tam97]. convolutional [GY22, La 19]. Coperturbation [AFM13]. Copies [Bad19]. copositive [BS17e, FJZ21, Hil18, SM20, SQ15]. coprime [LT12b, Tan10, TLL11, TLC16, XL11]. coproducts [Nas74]. Copson [CW10]. coquaternions [FMSS19]. coranks [Ger86]. Core [BT10, FM22a, Mos21a, PM14, Rom20, BY76, BBX20, CC21, DLD21, FM22b, GCP20, GCP21, LC18a, LCZW19, LC20, LZ23, ML21a, Ma22b, MRT14, Mar18, Mos19, MM22b, SX22, WL15, WLL21, ZCW21, ZC22, ZP19, ZW20, ZWM23]. Core-EP [Mos21a, PM14, DLD21, GCP20, GCP21, Mos19]. Core-Nilpotent [Rom20, Mar18]. corona [CL17c, LZ14b, Ng22, WZ13]. coronae [LPH17]. coronas [BS17a, DP17, LLZ22, Sta10, TYC21, WLW23]. Correction [Ano20a, Ano20b, Ano22a, Cal99, – Dok93, Tra90, CKL^{+}21, MS16b]. Corrections [KNX95, LLS16, NS93]. corrector [Chu84b]. correlated [Sma17, TJ16]. Correlation [JT04, Pat03a, PV20, Dat16, Lov21]. correspondence [Foa79, Wat86]. corresponding [Bén10, CL86a, CL86b, FD14, Har81a, Kyr08, Kyr11, LD19, SLUG04]. Corrigendum [Ano17, Ano22b, Bou14b, BM19b, HJ19a, KFR20, Raï14a, Sło18a, Sło18b, SLH18]. cosets [Hol04, Ner22]. cospectral [BB19b, BBCD21, But10b, CDG23, HS95, KP22, LV18, Sta22]. Cospectrality [Obo16, Tur18]. Cosquares [Bal78, Bal79c, Rad80]. costandard [GMZ11]. cotangents [Cal95a]. countable [dSP22]. countable-dimensional [dSP22]. Counterexample [DJ89a, EF98]. counterexamples [BH74]. counterpart [Lon97]. Counting [KK90, Ota13, LZB23, Sur98, Wri22a]. coupled [jAS15, BS17d, Pen15]. covariance [TJ16]. covariances [Kan94]. covariants [Bed10, BB17c]. covectors [MC19]. cover [BS16, Wan21]. Covering [CS95, ST86, Fer09]. coverings [But10a, MS10c]. covers [CG16, Sch95]. Cowen [Bol12]. cp [SMBB^{+}15, SM22b, GLM^{+}22, Heo16, SM22a, Shi16]. cp-rank [SMBB^{+}15, SM22b, GLM^{+}22, Shi16]. Cramer [Che93b, Che93a, Kyr08, Kyr11, Wer84, WY18]. criteria [GB77, HM19, Joh14, WSZL17]. criterion [BGdH92, Dix79, GPM98, JP15b]. Critical [Cal97, Cal99, Gei78, KOvDD09, AVVÁ18, DLL19, DH16, HJP17, MNS22]. cross [Dat16, Joh14, ZHH09]. crossed [NY21]. crown [Mir23]. cryptology [MZ17]. CSR [MNS22]. cubic [BH20, GR10, HKT15, LR17]. cubics [Wri93]. Curvature [Jac79, MZ91]. curves [CN21, Ma’15b]. cut [Fer10a, HZ03, KM98, WH11, Zhu15]. cut-points [KM98]. cutpoints [Kir00]. cuts [Mer94]. Cutsets [BS93]. Cutting [Ost74]. cycle [BP90, HO17, JJK95, Kir09, Vo83c]. cycle-based [Kir09]. cycle-sums [HO17]. cycles [CT16, DHKS16, FdF09, KK19, LN16, SG09b, Tik22, WSH^{+}11, Wil93]. Cyclic [KW89, BL93a, BS17d, BP09, CP15, CP18, CP21, Hua01, Lew77, LEH96, LDL17, LW23, MW15, MS10b, dSPdS22, SMF95, WT16]. cyclicity [AFA22]. cyclotomic [New93, Zha21].

D. [KM84]. DAHA [Hua22a]. Dam [zY11]. d’amorcage [BC06]. damped [YZC21]. damping [LLLY23]. Dashnic [LHZ22]. data [BCR91]. Davis [Alo23, CNP06, DW23, KMN22, ZMCN19]. – Doković [Lon81]. decidability [GB77]. Decomposability [Wes96, LK16]. Decomposable [MC78, Mar78, dP74, BD97, BLD91, BL92, CL95, Che94, CL94, CT09, Har87, LT88b, LZ01, Lim92b, Lim97, Lim01, Lim10, Loe01, Lov21, MF79, MC80, MC83, MC89, Mer77, Pat90, RS17, yT86, Wan82, WG91a]. Decomposably [GP96]. Decomposing [Sza09]. Decomposition [AJL14, CM85, LP02, YQ90, BDH23, CM92, CH87, DO20, FLS88, Gow22, GHT16, Gut22, HMR22, HT10, HZS22, JW22, Koi21, LK16, Li21b, Mar78, Mar18, QD85, RRN22, Roi99, Rom20, SRT22, Smi83, SZ19, Sta15a, TZ19a, Ter22b, Ter22a, Tia11a, WL20a, jXL15, YZ16, Zha15a, ZZC17]. Decompositions [CKS19, DGL22, Fen10, BB17a, DNY22, Fur08, GCSC20, GVS96, Hor04, HLZ18, HLYZ22, Kru06, LMMR01a, MG22, Sta20, YLX20, YC19]. decrease [Sin76a]. decreasing [BCS18, NY23]. Dedekind [FSvW11, GMR81, Tan18a, Wag87]. Dedication [Tho92a]. deduction [Joh82a]. defect [CK17, SU77]. Defekt [SU77]. deficiency [LS22c]. deficient [LLF18]. defined [Dag23a, ES76a, Erd22, GP06, HSM17, MDP22, MC80, TR21]. Definite [DH03, ARK22, AO78, Bak84, Cha92, Cha93a, Che92, CKS20, DHHK23, DPP17, Dru15, DWL14, ESH23, Fan73, Fan92, Fla75, FG22, FS15, FSY16, GMZ17, GJMW86, Har78b, HBP18, HM18, Hua22c, JN80, JP87, Józ15, LLMT05, LP22a, LL14b, Liu22b, LSX19, Mai11, MK16, Mol15, NMAB22, Pat92, ST94, WL22, Zha04c, Zha16b, ZL21]. definiteness [Cha90a, Ibr00, LL16b, Sun20]. definition [CKP09]. deflating [DLL19]. deflation [Ami11]. Deformations [ZL15, AC19, Bre17, Zha15c]. degenerate [Ait19, ABAH23]. Degenerations [AAA22, KPPV18, AH20, Ano22b]. Degree [FdM18, Yiz03, dFSY22, BZ00, BM89, Cal94, Cal97, Cal99, CD22, Cha77, Cha79a, CPLH17, DTSI15, Ere15, FA22, Gro95, HLYZ22, HLB22, HW75, HLP23, LLP18, LLF18, Mar90, MP73, Mer78c, Mes13, O’R15, zSqSZ16, Wan22b, WY20]. Degree-inverting [FdM18, dFSY22]. Degrees [HS87, MZ17]. delay [CW17]. delayed [DHKS16]. delete [BVW84]. delete-contract [BVW84]. deleted [LS09, ZWL15]. deletion [TLC22]. delta [HMK22]. Demicompact [JKZ20, FJK22]. dendriform [Bre17]. Dénes [Aha76]. dense [LCG^{+}23]. denseness [GR23]. density [BLX18, GHL16, MN12]. dependence [CPKK22, DU21]. dependencies [Knu85]. dependent [GMZ17, dCH21]. depth [Vo83c]. depth-first [Vo83c]. Deree [Cal03]. derivability [CW11b, CW12]. derivable [AC21, Gha12, JL12, LPS13, LF13, Wan17a, Zha16a, ZW14, ZXL16]. derivation [Bah22, Bou06, Hou17, MS87, WHX08]. Derivations [AC21, Che03a, GPW22, Lim92c, LT22a, RC22, Vla22, Vla23, WOY06, WT23, YW10, AADD18, AH10, AH13, AXH15, AWM22, AA20, AKOZ20, AKY21, AK21, AAUN22, BD22a, Ben07, Ben15, Ben16, Ben22, Bre16, CMR20, Cal94, CGOT13, CX12, CZZ22, CQ22, Da 94, DD15, DS23a, EPR16a, EPR16b, FZ17, FA19, Gog21, Hon22, JLZ12, JQS13, KP16, Khr16, KZN22, LWZ15, Lee16, LCL13, LZ11a, LW11a, LB11, LLF14, LL10a, Lin21, LL11, LCL14, Liu16a, Liu17, LL18, Liu19, Lon76, Lud19, MJ07, Mag13, Mag14, MW74b, MA75, NZK21, Qi15, QZ16, ST99, TRD16, WY11, Wan18, Wan20, XW12, YZ13, YF18, Yan21, Yao22, YH22, ZWZ10]. Derivative [Tur02, Chu22, Dec74, DM87, MS81, OvdD77, ZZ23]. derivatives [Car20, KE82, Vei82c, WZ23]. derived [CHFC19, EB17, HS91, Ilk20, NJP19, Ogu20, SK15b]. described [Gon17]. Description [AA20, AAUN22, GMZ11, AC23, Bal18, Hil07a, Hil11, Li92]. descriptions [XHL23]. Design [Chu84b, PS22, WM84]. Designs [GK02, NW02, CPKK22, Der83, GGW76, LN10, Sma17]. detection [Kar90]. Determinant [CD03, Nag17, BS15b, BM07, Cal91, CK20, DM22a, FV97, FLR85, Han19, HS96, Hon91, JOTvdD94, Kua18, KAE22, Lau81, LSOvdD95, Lim79b, Lit17, LC83, McD82, Min77, QK92, RRC23, Rou98, Vei82a, Wen06, ZZ23]. Determinantal [BJ94, CNUV16, Kyr11, Kyr21, YW09, AdF21, BP98b, BM18a, BP18, Dru14, DL14, DdF22b, HL84, Hon92, Hon93, Li92, LZ17b, Lon97, New79, New82a, Vei83a, YLC19b, dC13, dC15, dC18]. Determinants [BT15, HKP23, HZ11, KS22c, LT12b, MN80, Mah82, New80a, PA94, Pri78, SvW14, TT12, Tan14, TT78, WC22b, BMD87, Bla77, ES73, Est80, FW86, Gar75b, GSZ22, Hon99, HH14, JNT96, LPS08, LM19, Loe87, MM22a, MS89, Min80b, Odo83, San14, Sim83, Sto79, TX13, TG23, Tho75a, Vei82a, Vei82c, Vei82b, Vei83a, Vei83b, WW20b, dCdSA22]. determination [CLX22]. determine [LL16b]. determined [Bre12b, BH15b, CSG22, CAP22, Gha13, Gra10, Gra11, LW22b, MDW11, MA15c]. Determining [BMGMC15, Kas85, Vo83a, AS12, Mos20]. developable [BB13]. Developments [CD85]. Deza [AGH^{+}22]. Diagonability [Tan20a]. Diagonal [BGGS17, BD22b, CS18, DV19, Wan76, Bal79b, BBP78, BOvdD91, BKS12, BC20, BS08, Bur95, DHS95, DD12, DW23, FS89, GS21b, HH85, Hon91, Ikr15b, JR09, JLDS12, sL19b, Lee18, LM89, Lon90, Mir92, Moo76, Neu83, Pat82, Pok87, Pri78, SSGP23, Sha91, Sto79, TT78, WOY06, WC12, Zab89]. diagonalisability [Kob06]. diagonalizability [FHL^{+}13]. Diagonalizable [Tur02, DGL22, KS16, Nie14, PG22, WW12]. Diagonalization [AY74, KN22, MS10b]. diagonalizations [Uhl79]. diagonally [BCKP10, LMLL17, SC21]. Diagonals [EM73, Pif87, BS08, Lás91, Web83]. dialgebra [ZCY22]. dialgebras [BPSO12, FSFSO^{+}14, KV13, SDVWT16]. Diameter [JS16, Tsi83, CLN21, CDK22, CJL22, GL11, GWW20, HLW^{+}21, LP07, TJ11, VH95, WH11, XSW21]. diamond [Cir17, GLLM21, LPT14, YJ21]. diassociative [Bre17]. Dickson [FSFSO^{+}14, Lud21a]. dicyclic [CFL^{+}22b, DP00]. Dieudonné [BGLS07]. Diffeological [Per19]. Difference [KR03, AHT85, Bal22a, BB19c, CKXX23, DH17, ESAJ18, FN14, HP90a, Igl99, JvdD91, MM20a, MM20b, Mia93a, Rom22a, Sal14, SES21, Sil88c, Sil88b, SS96, TCI13, TRM22]. differences [TX13, WC22b, Wu22]. Different [PP22, ZIC22, Bén10, SS22]. differentiability [BBX20]. differentiable [CFP10, MWLW22, Nom73]. Differential [DU21, LGG19, BHY22, BBT07, BZZ11, Cam83, CW17, DTSI15, DDA20, Fin92, HSX16, Lon76, NS99, NT91, PZGL22, Wit11, WW20b, XSL15]. differentiation [LPX23]. differing [New93]. diffusion [HFSZ22, ZQZ21]. digit [JT23]. digital [Chu84b, Sto79]. digraph [BBZ15, DK17, KKS22, LM10, MMS21, SZWB16]. Digraphs [AVVÁ18, BP90, BP16, BR19, BPR20, CRW21, CR22a, CJY22, CFGM06, DM22b, FKC19, GB22, GM21, KP12, KS22b, KFR17, LM10, PR08, Rad12, RBG23, Sah21, SWH22, SGS05, SG09a, SG09b, SSG12, XSW21, XSW22, KFR20]. Dihedral [Ano16c, BS18d, Ous15b, OS18, CF21, CP15, CNW18, CFL^{+}22b, HK13, LCXC22, LCWW22, WC22a]. Dilation [GW98a, JE21, Liu22b]. dilations [CL00, GC22]. dilute [Din15a]. Dimension [ASZ02, Gra80, How80, Lim00, Ab23, AS20a, BCZ22, BM14, BM16, Cen11, CL86a, CL86b, CR77, Con77, GPK94, Joh22, LHS^{+}21, MM09, MN13, NJP19, OWTZ21, Pam87, Rob75, Rom89, TH14, Tow22, Wri93, ZCY22]. Dimensional [BR04, ySpW14, ARH15, AELPM18, ALPM20, AAUN22, BC83, Bar07, BB14, COK20, Car88, CDJ13, CW11b, CX12, CL17b, Das17, DH20, – Dok83, Edw78, Far94, Gol79, Gol07, HMK22, Hua22a, JO23, KK75, KL15, LF15, Liv94, Lud21a, MDP22, MC19, ME21, Nag17, NR23, Nir11, Pas15, dSP22, Rai87, RH11a, RH11b, RŠ08a, RŠ09, Sel81, SHC22, Tal17, Tan18c, WC12, WT23, YLL90, dTS20]. Dimensionality [Hil76].

Dimensions [DvE20b, MR74, BMFPBSS21, Cas91, DBG18, Jon78, MO91, MM76, WL16a, WL16b, DvE20a]. dimer [Min80a]. d’intégrales [BC06]. direct [BHK10, Fow80, Gol92, Hol84, Hol85, HLYZ22, JC22a, KLO14, KL13, LW10, LLLY23, RC18, Tho92c, XY22]. directed [BBG22c, Bru82, GH23, Guz88, JKKC22, Kir14, Lee98, Wu05, XQ16a]. direction [KM17]. directional [RS23]. directions [Abd20]. disaggregation [HUZ17]. disc [CPLH17]. Discontinuous [GL01, Fri97]. Discrete [DL22, Dic84, Elo18, FN14, GMZ17, Gue19, HMO11, HMK22, KW89, KK18, LRR87, LRD21, lWY19]. discrete-time [Gue19]. Discriminant [Pie79]. Discriminants [Wad85, Edw80]. discs [AK94, GW12]. discussion [Sch22]. Disjoint [Bru79b, CT16, Tia11a, WSH^{+}11]. disparity [CJL22]. Displacement [Ma22b, Boz97]. Disproof [Gib87]. dissipative [JK22, KS19b, ML21b]. Distance [Aue95, BHD11, BCEM12, CFL^{+}22b, DM22a, DM22b, HW08, LT89a, LLD17, LWL21, LW22a, LH21b, Rum97, Sha15, ZW17a, AAKS21, AW07, ABGP21, AAB20, AH16a, AKB21, Aue98, BBG22b, BBG22a, BS11, BK17, BS17a, CC14, CFGM06, Cos14, DFK19, DAH19, DMV18, DeV22, Dod13, DIF13, EGMS22, GB11, HS95, HLZ22, HS16, JJK95, KSU22, KLPK17, LS21a, LS12, LS13, LL15, LZL18, LZ19c, LT13, LXG17, LY22d, LHH18, MS89, MT22b, Mir23, Mou20, NP13, NC22, Pat07, PRAC22, PGAB22, RP90, Sch92b, SP87, TG09, WZ18, XSW21, XZL14, XZ16, XYH22, XLS20, ZG11, ZBW16]. distance- [DFK19]. distance-2 [DFK19]. Distance-regular [BCEM12, HS95, Mou20]. distance-regularity [CFGM06]. distances [PP22, SWZB15]. distinct [AOTR11, CJ97, HL22, KLPK17, MTR19, Ota13, Ram22, SRK22, Vei82b, WW20a]. distributed [HFSZ22]. distributed-order [HFSZ22]. distribution [Ab23, LN14, Sto79, Til98]. Distributions [FKR93, TL12]. distributive [CZY12, MPGS21]. Divergence [KD16, Hac22]. divergences [DDF22a]. Divided [TX13]. Divisibility [DK11, Tan10, TLL11, TLC16, Wit11, XL11, Zha14b, Hau12, LT12b, Tho88]. Division [RR04, BDH23, BS22d, HWZ18a, Hua10, HZ20, RŠ08a, TNJ22]. divisor [DJS21, HH14, Kri87, LT12b, OWTZ21, PRAC22, RM22, Tan10, TLL11, TLC16, XL11, ZWM17b]. divisors [Bar80, BC91, CG19, DS16, Gau76, GKLR82, MW74b, McF79b, Mcf79a, MP73, Min82a, New79, New82a, Nor08, Tho86a]. Djordjević [ZZY16]. DMP [Zhu19, ZIC22]. DMP-inverse [ZIC22]. Do [MS10a, BT08b, Kuz05]. Dodgson [Bay22]. does [BPW16, BR22, Sta10]. doi [Ano16b, Ano16c]. Domain [TRM22, BJS98, Dix79, Ilk20]. domain-range [BJS98]. domains [BE22, BZ22b, HH22a, NST89, Wag87, Wei93, WY18, YW09]. Dominance [Chu90, GMW86, Hey88, JP87, MS12, Pat89]. dominant [BCKP10, CPLH17, Ghi22, Hav88, Lee94, LMLL17, Pri78, SC21, Sin77, TT78]. Dominated [Pop17, BPR14]. domination [AB22, WWS19, YWZ21]. Double [FL18, BF05, DLL19, DdF22b, Gut23, JS94, Ner22, Sch95, SGM22, SH92, Tal17, YW10]. double-complex [Gut23]. double-one [SH92]. double-zero [SH92]. doubling [HM18]. Doubly [AS95, CW93, Joh81a, Mer98b, Sin73, Sin76a, Sin77, Sin81a, SHC02, SJ02, Ach77, Ben78, Bru85, BS08, BD13b, BD22b, Bur95, CKP16, CK17, CCKL20, Cha83b, Cha84a, Cha88, CCW20, CL05, DJ98, Esh17, Esh22a, Esh22b, Fie90b, For79a, FS86, For88, FM78, Gib76, Gol92, HJP17, Hwa85, Hwa91, Joh73, Joh74a, JNS23, KS76, KS82, LM89, Mal91, MKS84a, MR07, Min75, Min82a, Min84, Mou04, MAM15, MA15c, NMAB22, Phu80, Pif87, Sin76b, Sin80, Sin84a, Wal75a, Wan76, Web83, dF13]. doubly-stochastic [Web83]. downdating [FS22]. Drazin [Boa13, Bou13a, Bou18, BZZ11, BSZW13, Cam80, Cam83, CS22, CICX09, CIP18, Cve18, CM22, DCIW10, DD12, DK22, Gar23, HBM15, KW11, LPT13, Mos15, Mos21b, Pab22, Rom19, Rom21, Rom22b, SPM14, SPK15, SKM17, WfMR16, Wei98, mWyD11, Wer87, XSH18, YZZ20, ZZ16, ZMG20, ZWW22, ZCCIW12, ZZC17]. Drazin-inverse [SPM14]. Drazin-Moore [Rom21]. Drew [Hu94, Moh15]. drifts [WS23]. DSS [EBB18, LRD21]. DSS-weak [EBB18, LRD21]. Dual [IKR12, BBX20, CFZ23, CDK20, De 08, DS23b, DLQ18, DF95, Din15a, GLL21, GC22, Gut22, LC18a, LCZW19, Man91, OP13, Oku93, PJ90, SX22]. dual-numbered [Gut22]. Duality [LT88a, LD22, BL83, Bar78, De 75, Joh14, Pro22, SA21]. duals [HLT22, LD22]. due [TLC22]. Duffin [CLX03, GZZF23]. Duhamel [KT06]. duplication [JS08]. Dyads [dP74, De 78]. dynamical [JL22, MW19]. Dynkin [BB14].

E-efficient [Sma17]. E-optimal [Sma17]. EA [HWL21]. each [BLP12]. Eaton [yT00, Tam02]. eccentricity [MGKA23]. Eckart [Nas75]. Edge [BGK00, DWL20, AAGL22, AMK23, But10a, CT16, CL17c, DJK08, FHZ17, Fer10a, GM90, KN04, LS09, LLZ22, LP23, LPH17, LOO22, LCXC22, Mer89, NP15, TLC22, WZ13, YHS13, ZLS11, ZQC18, ZWL15]. edge-connected [LOO22]. edge-deleted [LS09, ZWL15]. edge-disjoint [CT16]. edge-grafting [ZLS11]. edge-most [YHS13]. edge-path [AAGL22]. edge-pockets [NP15]. edge-Zagreb [ZQC18]. edges [JST18, KS17a, LLF08, LY22c, TJ21, XW19]. editor [Bha91, Cha93b, Roh13, Sch15, dS19a, Ano73, Ram02]. Editorial [Ano79a, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f]. Editorial [Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano11a, Ano12a, Ano13, Ano14, Ano15, Ano16a, LT04, Mar82a, Tho90]. effect [GOS22, LLF08, LY22c]. Effective [CMM17]. Effects [GLS16, TJ16]. efficient [AAB17, DK23, Sma17, ZWLZ22]. Egorycev [Min82b]. eigenbases [AC13, Aln15]. eigendata [GPM14, MMG15]. eigendiscriminants [Bus23]. Eigenfunctions [Red02]. eigenmatrices [Kir11]. eigenpairs [SN22b]. eigenparameter [GMZ17]. eigenparameter-dependent [GMZ17]. Eigenproblem [lWpW14, BOvdD95]. eigenspace [KL21b, NSW21]. eigenspaces [DFPS23, GTM14]. Eigenvalue [AAKS21, Har78a, JSE90, MMRR16, MS21a, AHJ20, AHJ21, ABP22, Ano16b, Ano22a, BNP06, BLP08, BGM22, BMP22, Bel15, BPS16, BPW16, Bén11, BCK15, CC17a, CQW16, Cro18, CCLN16, CR90, Das21, ES76a, Esc93, Fan05, FdF09, Gar85, Guo07, GZ22a, HA22, Har95, Joh84, JL85, JLR^{+}92, JD99, JLS09, JLDS12, JST18, JNS23, KMT96, LP00, LLF08, LLZ13, LLL15, LL16b, LZL16, LW12, LZ15b, LPW17, LXG17, Loe73, LHH18, Mag22, MH20, Mel19, Mel22, Mir18, MMG15, MS15a, Obo22, PLH01, PMP22, dSP12, PSB23, Rea96, Sad23, SL19a, SG96, Shi15a, Sim95, Smi05, Sta10, SW22, TXL21, Tho92b, TJ21, Tre85, Tre86, Ves80, Wal11, WWWT22, XZ16, Yin21, YWZ21, YZC21, ZHZ11, ZCLW22, Zhu15]. eigenvalue-eigenvector [Har95]. Eigenvalue-free [AAKS21]. Eigenvalues [AHT85, AM85, And04, jAS21, BWF83, But10a, Cas91, FJM00, FIS06, GKV89, Gro95, HLP03, MS95, RM22, Rod02, Ter09, VH95, Wea83, AM13, ADHM19, AAGL22, AAD22, ABGP21, AS17, AB78, Ano79b, Ash16, AOTR11, BS21, BF05, BOvdD91, Ben90, BT07, Bru82, CTW15, Cra09, Da 05, DFK19, Das04, DAH19, Deu80, DL22, DIJ23, – DM79, DJ89b, DWL20, DZ22, Edw78, ELG88, EJ91, Esc93, FDLB20, FJ15, FWZ11, FS14, Fer10a, FW13, FS89, GAMA23, GRSW92, GJMW86, GZ90, Gue19, GHK16, GZ22b, Ha89, Her83, HS93b, HUZ17, HL22, HW75, Ilm16, Jin17, JKB17, JKS79, JS08, JSW09, Kir91, Kir09, KLPK17, KNX93, KNX95, KK16, LS10, LST23, LXQ21, Lei18, LL14a, LS21a, LCLL22, LY22c, Liu18b, LH21b, Mac09, MVPS14, MVPSTT23, MS89, MH19, MR07]. eigenvalues [Mer94, Mou13, NS99, Nie10, NTU93, Pal79, PGLM00, Ram22, SRK22, SU77, ST15, Sil88a, Sil93, Sil99, SD20, Ter01, Ter06, Tho74, TL12, WW20a, XZS19, XLS20, dLMO17, dOMD77]. Eigenvector [RP90, Che14a, CPZ17, Har81a, Har95, Hav88, Kul22]. Eigenvectors [BD14, AM13, AGM13, BWF83, BDK^{+}05, CT21, ER81, Fan88, FJS14, Fre76, GF10, HQ14, Lit17, Liu18b, MOvdD90, McD92b, Moh16a, Wea83]. Eigenwerte [SU77]. Eight [Tia10, – Dok83]. eight-dimensional [– Dok83]. Einstein [BM17a, BS22c, HM22, lLZjZ19, MWC22, PM22, SZBW16, XJS20]. elasticity [LCLL22]. element [BB17b, DHS95, Fer09, HW75, LLLY23, ZCP16, ZCP17]. Elementary [BC91, CL13, FL16a, Gau76, MW74b, Mcf79a, MP73, MŠ02, Pie85, Sed04, Aue95, Boa13, DS16, FGK23, For87, Gog21, Jam81, Joh81b, KKR12, Kri87, Li87b, LB21, MS82b, MKS84a, McF79b, Min76, Min82a, Nor08, Oge94, Sch22]. Elements [Ger04, BBBS09, Boa13, BM17b, BM18b, Bre15, BS07, Cro18, Deu80, DKS09, FA22, GG22a, Hon91, HK10, KM85, LR13, sL19b, LPT13, Lim10, LL18, Lon90, MV91, Mir92, dSP16a, Sch79, Slo13c, UF22, Vei82c, WM21, Yam18, Zhu19, ZCCIW12, ZCP18]. eliminated [TB01]. elimination [BJ94, LW10, SX22, Stu92]. ellipsoids [Zen22]. Elliptic [CN15, GW08, GW12, MB22b]. Elliptical [BS04, GS21a]. ellipticity [LCLL22]. EM-based [CW20]. Embeddability [ER22]. embeddable [CFSRL22]. embedded [Fon92]. Embedding [Cra91, SR22, Wal86, Gue17, Gue19, TZ17]. embeddings [BB19a, De 08, Hor04, Pan13, SZ85, VDW85]. empty [LPS09]. enclosing [DM22c, vD96]. endomorphism [ZY08]. Endomorphisms [Bre18, PS21, GR14, Mol15, Pab22, dSP22, Rom19, Rom20, Rom21, Rom22a, Rus16]. Energies [YLW22, GH22, MT22b]. Energy [Mer94, PR08, Ste05, AMK23, BdP21, BP16, BR19, BPR20, BSA17, BBZ15, BSG^{+}12, BH20, CR22a, DLSS18, FKC19, FHT14, GB11, Hou01, KFR17, KFR20, MGKA23, MT22b, NC22, Rad12, Tan12, TW19, TLC22, Tur18, WMFW21, WL20b, WLM21, ZQC18, ZZ15, ZWS20]. Engel [AADD18, DS13, Liu17, Mat11, Vol22]. enhanced [MR18]. entangled [DM22d, WCS^{+}21]. Entanglement [BW23, Aud15, Joh14]. entire [SK15a, SK15b]. entried [dPN81]. Entries [Mol18, Sin81a, BGV11, BGGS17, CJ97, CL14, CKS20, FW07, FW86, Her83, HNS99, HL20, INO15, Ito16, JKD16, KT21, LHS^{+}21, Mee80, Min84, Moo76, Pok87, SG96, Sil98, STZ94, SC22, WC22b, Zel75]. Entropy [Zha14a, BM19a, BdP21, Drn13, Fur11, MN12, MF20b, ZBW16]. Entry [Che03b, Cha84a, JLDS12, KS82, SSZ22]. Enumeration [CJY22, Chu90, WW76, Whi79, Wil83]. enumerative [Whi75]. envelope [Chi98]. enveloping [PZGL22]. EP [Ano22c, BT08a, DLD21, DKS09, Dra22b, GCP20, GCP21, HS83b, JKH20, KDHJ22, MRT16, Mos19, Mos21a, PM14, WD17, WL20a, Zhu19, ZCP18]. EP-nilpotent [WL20a]. Epimorphisms [Slo15b]. epr [MR17]. epr-sequences [MR17]. equal [AW83, Bot87, Cat21, FHT14, HW75, Jul21, Lim90a, Lon90]. Equalities [AC78, BCI13, MS74, FLT20, KT19, MC80, Tia01, Tia05b, Tia11c, Tia21]. Equality [CGW14, Hia94, CL85, DSD83, Din98, HS89, HS91, JN90a, LW11b, LW10, MC83, Oku93, WY21]. Equation [AP74, Che93a, HE20, Tia03, AGR91, Bak84, BGdH92, BB19c, BCT23, Che93b, CK84, CIDK07, Dat80, DS23a, De 13, DWL13, DWL14, FOT23, FS19, GPM98, Haj18, Has18, HW13, Igl99, dJfLlZ^{+}21, JZ22, LRR87, LM22, Lei83, LZ19a, LQ95, LT10, LR89, MK16, Ned98, NF19, PLPC19, Roh04, Sal14, SLH20, SWL13, SPK15, Vei83a, WXD21, WHL22, Wil83, YL11, YLW15, YZL17, YTL20, YCZ21, Zha16b, ZL21, ZWLZ22, ZLZ^{+}22, ZZL23]. Equations [DH03, Fla74a, Pas16, BHY22, BBP78, BS17d, BC06, BZZ11, BSS12, Cam83, CL80, CW17, CCL10, CKXX23, CIRW21, DH86, DTSI15, DH11, DM22c, DMT21, Din00, DLL19, DBG18, EÖ15, ESH23, Esh22a, FHW16, FNW18, FN14, Fin92, Gra80, Har94, HKT15, HW14, HW15, HSX16, He21, Hla17, HBP18, HM18, HM20, HFSZ22, HM22, JvdD91, JSY14, KP11, KM17, KL22, KN81, Kyr21, fLLL^{+}17, LD19, lLZjZ19, Lip96, LB21, Lon76, Mag83, MWC22, Mil22, NS99, New80b, Pen15, PM78, Roh91a, Rom22a, SLUG04, Son14, SK15a, SK15b, ST97b, Tia00, Wal75a, WH12, WZH13a, WZH13b, WX19, WLZ22, XJS20, XSL15, XW22c, YqS16, YW16, YZL22, Zha04c, ZQZ21, lWY19, BC06]. Equiangular [CDK20]. equienergetic [LS09, PV23]. equioriented [Sch22]. equiponderant [Pif87]. Equitable [YZB22, Aln11, AC13, Aln15, Ati20, LHF22]. Equivalence [Gol87, HWZ18b, MH20, New79, Zha01b, Ari16, BB20, CH87, DMV18, Der18, Drn13, EOvdD09a, EOvdD09b, Gow80, Grü76, HWZ18a, HW08, Hou09, New78, Oge94, PR20, Pet00, Ano22a]. equivalences [AKR23]. Equivalent [AXH15, FA19, Bru79a, Hou09, LST23, Vei83b]. erasures [Ano16c, DS23b, Ous15b]. Erdo [LN16]. Erdos [GX17b, Woo22]. Ergodic [MSS13, HWX22, Har84b]. ergodicity [MH22, Sen83]. Ernst [Vei83a]. Erratum [Ano79b, Ano16b, Ano16c, Ano19, Ano22c, JZ95]. Error [Hav88, Wan22a, CKL^{+}21, GLL18, Guo20, MS16b, Ned98, SC21, SWL13, UZ16, UB22, WS18]. errors [AM13, CLW15, MHM22]. Essential [BHC16, JKD16, JKZ20, JLL07, XS18, ABJ16, BC20, JR94, LT22b, Pep19a]. Essentially [Joh73, Joh74a, Joh81b, NST91]. establish [Tia18]. estimable [YC19]. Estimate [Gil93, Bou06, Con77]. Estimates [Fri93, GHK16, LLW^{+}11, MFPSS12, Tom22, AP17, CPLH17, Luo19, OM21]. estimating [CPKK22]. estimation [BLX18, LSX19, Mel22, SLW21, UB22]. estimations [FSWX22, WWW20, ZFSX17]. Estrada [Sha15, ZZZ23]. Euclidean [AW07, AGR91, BBG22a, Gil16, GT11, GJ22, Gow22, JJK95, ST15, TWK22, TJG22, TG09, Zen22, dlP20]. Euler [Prz92, TD14, Tal17]. Eulerian [Gar79, TB01]. evaluated [Mes13]. evaluating [Cal95a, TX13]. even [AM13, CQW16, CCLN16, HH14, Jin17, JKB17, SA14]. Eventual [YHCW14]. Every [Slo17, Vas86, Sło18a]. everywhere [CS19]. Evolution [EL21, JL22, Pan19, Pan20, CMR20, CMR21, CGOT13, FTGSNV22, NR23, ORT15, Pan22, QM21, RC22]. Exact [FP17, HH22b, FHW16, HSX16, HLY22, Hua01, MO93, Tia18, XSL15]. exactly [Esc93, HW75]. Examples [Wes90, JO23, KCBR87]. exceed [BPW16, Sta10]. exceeding [LZ15b]. exceptional [RW22]. excess [DFK19]. exchange [Pre22]. Exclusion [SL19a]. Existence [DHS95, Her83, AE10, CC22b, Fow80, LRR87, Mag22, MHA17, Nas74, Roh91a, Tie16]. expanding [Che95b, EM15]. expansion [CRY95, Hon91, Krä87, MZ22, TSYB82]. expansions [CKXX23, MNS22, RW96, Wil87]. expectations [BG16]. Expected [GLLM21, HL18]. experiments [Con90]. Explicit [BP18, CLX22, CICX09, DM19, Elo18, Mar20, Rom22a, CJL22, XHL23]. explicitly [KS16]. Exploiting [HWSH13]. exponent [HJP17, HL10, LM95, Liu92]. Exponential [MB22b, SW81, BV16b, Bel83, Chu84a, Dra21a, How83, Mar84, Tho86b]. exponentials [Cam78, ST91, ST00]. Exponents [JKKC22, SGS05, SG09a, LB99, SG09b]. exposé [Gar73]. exposed [HK16a, Mar13, YH05]. exposition [Mar89]. expressible [Pat83]. Expressing [GG22a]. Expression [DLW18, Tia11b]. Expressions [BL13, Wei98, FP17, Tia10]. Extended [GRMA13, BHY22, De 75, GHL16, HZG14, Mar78, Mas10, SRS21, WMFW21]. Extendible [BS07]. Extending [Cir16, Hon19, Sol00, ySpW22]. Extension [Hua90, LS16, Lei83, LLW07, PM22, Pat81, ySpWQ21, SL17, BG23a, BR01, BD13b, Far94, Fie73, FH89, GJR97, HH22b, HS75, Hou17, dHS23, Kua18, LZ17b, LZ12, LZ13, Mil15, Ng22, Pop20a, ZZY16, Zha15b, dC13]. Extensions [All14, Atk91, Dji16, Gar75a, Hal86, JKL22, LL14b, SPS15, AKK^{+}23, BWLL12, BM19a, BG83, BCS18, CKKK22a, CR18, DPR22, EL90, Epk89b, Epk89a, Far93a, Hon22, Hua09, KKL21, MV75, MCIM21, Mit92, MP13b, SST22, Suc22, TT12, Wer84, Wes96, ZL15, ZN18]. Exterior [Tam98, AGR91, Bah22, Coa99, Gre99, GV87, Jac79, Lim77]. Extermal [SM94, Lei95a]. extrema [For87, KKR12]. Extremal [APEPPS23, BP98b, CJ97, CH12, FF21, Grz89, GLS13, Hil18, KN00, Tia11b, ZL22, AO78, BSG^{+}12, BK87, CSCH01, Fla75, Li85, RW22, SWH22, WYSL22, ZXL22]. Extremals [YH05]. Extreme [Pat03b, Tho75a, CCW20, Kan87, LKLC22, MPD22]. eztensions [ELG88].

F [Ano16b]. Faà [Wen06]. Faber [ZZZ12]. Face [SHC02, Bru85]. Faces [Bar78, Zha20b, DH91, FS86, Hwa85, YH05]. Facial [So90]. factor [Ami11, HZXL15, JQS13, JQ15, Joh22, LT14, LJ11, Liu16a, Mol09, NZK21, WJ16, YZ20, Zha16a]. factorial [Tan20b]. factoriality [MM05]. factorials [TX13]. Factoring [Eba92, JS96]. factorised [ZWZ23]. Factorization [ADRP14, BT22b, Bot94, DKS09, Erd22, GLR82, JT03, KML95, KLZ96, Pop20a, Roo20, Vin04, Adu15, BG86, Bot96, Che07, Gal01, HZ12, JC22b, KT19, LYS15, LL17, LLC21, LZ22, MZ04, MR12, Neu81, Pop22, Rem90, Sou86, Sou92, Ter22a, Tia11c, Wim88]. Factorizations [Gal03, cc85, CZZH22, CZ22, Joh81b, LR96, MR97, RR91, SM22a, Wal79b]. Factors [Ger74, Aze99, Fow80, HS85, HS87, Liu18a, Liu19, MZ04, NT86, Sil88b, Tho87a, Zab89, Zab90, ZYL13]. faithful [CR15, Man17, WL16b]. Faithfully [RD12]. Families [Vei82b, AAS14, BD14, CHFC19, CM15, HH85, KSU22, MR17, Mer97, Mou20, Vei83b]. Family [KMS02, KS16, Kan87, KMEO12, Nom73, TNJ22]. fan [Adk95, HLZ22, LH21a, Smi94, Gro17, JLL88, MP94, MP95, Mos12]. fan-pall [Adk95]. Farka [Bar07, Bar12]. Farkas [XLL15]. Farkas-type [XLL15]. Fast [DM22c, Uhl20, XJW17]. Faster [Uhl14]. Fastest [AS11b]. Feedback [BZ99, BFZ07, BCdHP10, DHKS16]. feedbacks [HRT15].

Fenessey [MM23]. few [DFK19, MTR19, TDL18]. Fibonacci [BB19c, DH17, EB17, KI16, Ogu20, Sla16, TD16, TD19, dF11]. fibre [TZ21]. fibres [MT11]. fidelity [ZBW16]. Fiedler [BD14, CGMR11, Kua18]. Fiedler-like [BD14]. Field [Ric81, AKK^{+}23, Ber05, Bra86, CJK^{+}13, DB14b, Ell93, Fri90a, Ful76, Gib76, dHS23, Joh75b, JST18, KK90, Kyr11, LLC22, Mau78, Mee80, MN13, New93, NJP19, Ota13, dSP16a, PM81, PM85, RC22, SMS15, Uhl79, Uhl14, Uhl20, Wes96]. Fields [XCC99, Adk95, BM22b, Ben74, BB75, BM12, BD13a, BZZ10, CC17b, DP09, DJS21, Epk89b, Epk89a, Far17, Far20, GWZ14, GWZ16, Hai88b, Haw05, HZS22, KC16, Moh16b, MP13b, Nor95, Nor08, SCL^{+}22, Sta14, Sta15b, Wat89]. fifth [Gar21]. figures [Pok87]. Filbert [KA19]. Filiform [CCGO11, ARH15, ACO21, BS18a, CKKK22a, CLOK13, KKL20, KKL21, KLO14, LBV13, RH11a, RH11b]. Fill [Dji16]. fillings [AW13]. Fillmore [Tan20b]. filters [Chu84b, Pat17]. finally [Uhl13]. find [FTA20]. Finding [CDG23, Vo83b, HKT15, Sol14]. Fine [MG22, RE22, PS20]. finest [Uhl79]. Finitary [DS02b, DW16, BFJ12, BFK18, Khr16]. Finite [BR04, Che95a, Dre93, Hua22a, KLS08, LM03, Tau74, Adu15, AL21, AFW22, AMZ13, AAUN22, BM22b, BRL22, BG16, BM12, Bra86, BD13a, CCS16, CC17b, Car88, CDJ13, CP15, CW11b, CZY12, CX12, CG19, CR22b, Das17, DH20, – DL92, – Dok93, DJS21, Ell93, ES76b, Fal97, Far20, FTA20, Ful76, GG22a, GA85, Gol79, Gol07, GWZ14, GWZ16, GX17b, Hol04, IOMAN21, JC22a, JW22, KC16, Kes83, KK90, KL15, Lew77, Liu16a, LW23, LLLY23, Mat90, MGA17, MN13, Mir09, New80b, Nir11, Pab22, Pas15, dSPdS22, PA94, PM81, PM85, RA21, RŠ08a, RŠ09, RC18, Rom19, Rom20, Rom21, Rom22a, Sel81, SM19, SZ19, Tan18c, Tho92c, Wal79a, Wat89, WW20b, YF18, Yan21]. Finite-dimensional [Hua22a, AAUN22, CDJ13, CW11b, CX12, Das17, Gol07, KL15, Nir11, Pas15, Tan18c]. finitely [BS15b, Ber05, HMO11, Hua96, TLC16]. finiteness [FSvW11, PRA22]. Finsler [RQ22]. First [RP90, AEAAES20, BC06, CNW21, HWX22, KN00, Vo83c, Vol22]. first-order [CNW21]. Fischer [BJ93, Dru14, DL14, ES76a, HB83, MHH96, Mar76]. Fischer-type [Dru14]. Fishkind [Dji16]. five [HM76, HM78, MO91, Vol22]. Fixed [ZX16, Ach77, BOvdD91, BDK^{+}05, COK22, Cha84a, CDK22, DM87, GL11, HKv23, Hon92, JP07, KS17a, LP07, LT93, LS22b, TH14, TJ11, Tia17, Wes87, YWZ21]. fixed-points [Hon92]. Fixing [OWTZ21]. Flag [HLZ02, Mil15]. flagged [Chu90]. Flanders [GJR97]. Flat [CN08, ERS12, Pat84]. flows [Lee98, LT13]. Focal [GS19]. Fonseca [KAE22]. force [CN14]. forcing [WWZ20]. Foregger [Cha84b]. forests [Muk20]. Form [Afr73, New94, Arm17, BK17, Bed11, BCT23, BG82, Cha96, CFP21, Der83, Ger82, GRMA13, HS83c, HWZ18a, HH85, HZH20, Ikr15b, IH11, INO15, IK20, IK22, KH05, KPPR21, LV18, LMR94a, LCH22, LLC22, Mau78, Nak95, NSW21, Nor95, PS91, Pro95, Shr90, Uhl79, Vei83a]. formal [BC06, CQH15, Dre22, MPI17, Reg89, Rys75]. formally [Adk95, BB75]. format [Sta15b]. formation [LH22, dC15, dC18]. forme [Rev81]. formelles [BC06]. forms [AY93, Bae84, BT90, Bed10, BX11, BS13, Bus23, CCU21, Chi78, CO12, Dra17a, Epk89b, Epk89a, Fei75, Fow80, Ful76, Gal79, GP14, GQ06, HL09, Hor04, HS09, HLYZ22, IKR12, KN81, Lee18, Lew96, Lew97, Mar97, MN13, MP13b, O’R15, Per17, Pie74, Raï19, Rob77b, RB22, Sin93, Sol80, SPM14, Sta14, Sta15b, Sun20, Tau81, TS05, Ven19a, Wag87, Wil93]. formula [AH16b, Bed11, BG90, Cal91, Cla82, CD15, Dji16, Ebe80, FZ11, GR93a, GK18, HLZ22, IW95, KMMS22, Krä87, Lei83, Mit10, Mos15, NST89, SC20, SvW14, Tho86b, Vel21, Wen06, ZZ23, Zha21]. formulae [BJ94, BT11, HS83c, Rem82, Van95]. Formulas [Bus23, AYC93, Bay22, BSA17, GRMA13, Mar20, Rom82, SQH15, ySpW14, Tia18, YZB22]. Formulation [CN04]. forward [ZZT23]. forward-backward-half [ZZT23]. Four [Kom10, BTT20, Bre18, Edw78, Gra87, HWZ18a, HLW^{+}21, Tal17, ZMG20]. four-dimensional [Edw78, Tal17]. Four-term [Kom10]. Fourier [– Dok93, DL22, – DL92, Elo18, KW89]. Fourth [KMS02, MWC22, jAS15, jAS21, GM87a]. Fourth-order [MWC22, jAS15, jAS21]. Fraction [BL21, LB19, TSYB82]. Fraction-free [BL21, LB19]. Fractional [GW21, LLGX22, XZS19, Bal22a, HFSZ22, MM20a, Rei77, ZQZ21]. fractions [CM85, CR22b]. Frame [AN18a, Far17, GC22, KL13, MAT22, SA17, Xia20]. Frames [AB17, Ano16c, AB19, BCNG23, CC23a, CDK20, DO19, DS23b, FD14, FM21, Far22, FZ21, GNO20, GLL21, HLY22, JW22, KLS08, LLHS18, LL20, LH22, MO22, Ous15b, OS18, SA21]. framework [Mej19]. Frattini [GLOT13]. Fredholm [Amm18, Bal23, BBBS09, GCSC20, MŠ09, MO13, Yu14]. Free [Alp95, MCIM21, AAKS21, BdP21, BL21, Bre18, CZ21, DMV18, Dre93, DZ22, Egg95, GB77, Gho19, LLMN20, LB19, PA94, SLH20, ySpWQ21, Sla16, Sur98, TW19]. freedom [HS93c]. freely [Sch79]. French [BC06, Del94, MM92]. Friedland [KKS21]. friendly [Koh11]. friendship [SS23]. Frobenius [BOvdD95, BGLS07, Bla77, BDD23, BS22d, Fie73, Gil17, GS83, Gut01, HL84, MHH96, Mit10, Nag17, SPS15, Wan79, Zha21, Zou12]. Fuglede [Ras17]. Full [AB19, CJW19, Har91, MO22, Mum86a, AEdF13a, AEdF13b, DMQV23, Laf84, MSS18, Rub20, Wan17a, ZWM17b, dFSY22]. full-rank [Rub20]. Full-spark [MO22]. Fully [Har87, HL10, Liu92]. function [AEAAES20, AFM13, BB22, BH16, Bou14a, Bou14b, CRY95, CMM20, Cha92, Cha93a, Da 88, EDKA07, Erd22, FD14, Gam18, GP14, GP74, HSM17, HFS15, Hos22, How80, Jam81, JS20, Kap77, KN04, KKS22, KK16, Lei95b, Lim00, LCL22, Lon73, MMS21, Miy22, Pam87, Per17, RW84, Sal14, SMM14, SMM16, SK15a, Tia17, Vei82c, Wan79, WZD16]. function-induced [Da 88]. function-valued [FD14]. Functional [AS20b, CD22, Ere15, Lee16, Nie14, REG07, Wan19b, AGR91, Cro08, KD16, Maj22, Pep11, TB01]. Functionally [Fre76]. functionals [Lee15, MR75, Pap13]. Functions [AMP02, Hon99, Pie85, Tam02, Tur02, You80, Abd20, ABAH23, AHK19, Anj17, AA12, BHY22, BT13, Bal79a, BCN05, BC82, Bea83a, Bea83b, Bea84, BC85, BD85, BMFPBSS21, BL93b, Bra86, BHH86, CSG22, CH12, Che91, CP10, DV18, – DLR91, Dra10, Dra11, Dra16, Dra22a, DS18, DS19b, Fan83, FS83, For87, Fri81, Fri90a, FM22c, Gam00, GR23, GP22a, GR81, Gha21, Gib87, Gil93, Gil12, Gol89, Gol90, HHSY22, Hau12, HBI73, HWL92, HOT15, Hof75, HF18, HX21, dHS23, HH14, HLPW17, HLP23, IK22, JK22, JMS19, Joh75a, KJF23, KD16, KH05, KKR12, LR96, Li87b, LT90a, LT92b, Lim10, LM19, Lla96, LC83, MS10a, Mat90, Mer74, Mer79a, Mer82, Mer83a, MWLW22, Min76, MRS10, MB22b, NY23, Nie22b, OU22, PSMA16, Pat83, Pat93, Pat03b, Pat17]. functions [Pie92c, RR91, RW22, Rem90, RW96, Sal12, Sat11, SW81, SK15b, TN22, TLC16, TG23, Tik98, Tre85, Tre86, VY21, Wat77, ZW19, dCdSA22]. functors [JN22]. Fundamental [Mac09, Far90, Mag22, Rod22]. Further [ARK22, ASR22, BHK78, BM17a, Boa19, CZWF23, DW12a, DIF13, GHL10, How83, MRT14, Mis17, WC15, YK22, YWZ21, ZCP16, JLL11, Zha15b]. Furuta [Dra13a, FJK98]. fusion [SA17]. future [Abd20].

G. [BMD87, TT78]. Gabor [GKO22]. Gaddum [Das21]. Gaddum-type [Das21]. gain [HHY22, LY22c, LY22b, MKS22, WL23, ZH22]. Galilean [CSY18, TZ19b]. Galilei [aCC17]. Galois [Ota13]. Gamma [SAK22, ABAH23, Miy22, Sal12, UB22]. gammoids [Mum86a]. Gaps [YMYS13]. gauge [CFZ23, Che91, – DLR91]. Gauss [Sal14, SX22]. Gaussian [BJ94, CP16, DMV18, Ter22b]. Gaussian/Bernoulli [CP16]. gcd [PB17, TM19, Zha14b, AST05, HINS10, Hau12, Li92, Tan10, TLL11, XL11, Zha14b]. gcd-closed [Zha14b, Li92]. gcd-graphs [PB17, TM19]. GDMP [HLT22]. GDMP-inverses [HLT22]. Gelfand [BCZ22, Cen11, HT10, ZCY22]. Gelfand-Kirillov [BCZ22]. General [CHK00, HSX16, NP80, Uhl99, XSL15, XHL23, BO22b, CT21, Che93b, CW12, CPT15, CP21, DK23, HW15, HWL21, HLB22, Hun83, Hun86, KH05, LXQ21, LCL13, LLW07, MT22a, Pan13, Pen15, Pie92b, Pre22, SMM16, SLH20, SMM22, TJ21, WL16a, Wat89, Wri22b, Xu22, Zha04c]. Generalised [Fri75, FNBR96, HK80]. Generalization [– Dok17, Mou13, BM07, CGMR11, CC92, DMV18, Fie84, Hey92, HZ12, Kan87, KS14a, KPS19, KA19, Kov86, KAE22, MP94, MP95, Rub20, SN22b, TR21, dC13]. Generalizations [BC80, Dra13a, Gut22, MR10, Sha20, Sla16, CDK20, Dat16, DN18, FS83, Fri78, GS83, Liv94, Lu22, MP90, Wu22]. Generalized [BBG22a, BGK99, Bea83a, Bea83b, Bea84, BLdPS04, BdPNdP17, BS22c, Bou18, BCCL22, CM84, Cha96, CLX03, CL05, Cve18, DD15, DP18, Esh17, Far20, FMXZ22, JP15b, KP16, KL12, LW11a, LPS12, LCL14, LL18, MJ07, MW23, MP13a, Mer79a, Mer83a, MM23, PSA14, PUÖ15, Pet00, RS18, Rao81, Sin84a, SM19, Siv06a, SPK15, Tam02, Til77, Wer86, YZZ20, Zar01, ZZC17, AAB17, AADD18, AYT83, AYC93, AyT84, AKY21, BOvdD95, BI95, BF05, Bar80, BCDF09, BHK78, Bay22, BD22a, BC82, BC85, BM17a, BT06, BL13, BG21, Bo01, BT22b, BC94, CC17a, Cha92, Cha93a, CC14, CCC21, Che95a, CKN95, CW11a, CD21, CHZ22, CFZ23, Che17, Chi98, Chi78, CIN15, CIM19, DPvW13, DS23a, DeV22, Dec74, DH11, DTS14, DDA20, DCIW10, DLD22, DMS23, Dra13b]. generalized [Dra17b, Dra23, ESAJ18, EAM19, FT87, FS22, FZ17, FJK22, FN14, FLT20, Fie90b, Fri82a, Fur11, GB22, GS06, GZZF23, GPM98, GH23, Gog21, GB90, HNB22, Haj18, HK79, HJP17, HBI73, HKO16, HBM15, HT91a, HLPW17, HM20, Hug90, Jon92, JNBR93, KM19, KS15, KM14, KI16, KS14b, Kay17, KW17, KR14, KPPR21, KD23, Kyr08, Lei92, Lei93a, Lei95b, Lei95a, LCL13, LTT84, LMR94a, Li96, LLS10, LB11, LPT11, LWXF15, LZ19b, LW11b, LWB11a, LL11, Liu17, LPW17, Liu22a, Man87, MKS84b, MR15, MŠ09, MS18, MM20a, MM20b, MHM22, Mer74, Mer82, Mia93b, MS12, MMS21, MMG15, Mol18, MAT22, Mos15, Mos20, Mos21b, Mou04, Mum86b, Nab98, Nas75, ND13, NT86, NSW21, NF19, Pat03b, Pat08, PS20, Pep19a, PT10, PGAB22, Poo80b, Rad16, RŠ09, Rom20, Roo20, RS08b]. generalized [SMM16, SES21, SM20, SC20, Slo17, Sło18a, Smi74, SCL^{+}22, SY23, Tam86, yT86, TX13, TT06, Tia10, Tia18, Tia21, TCI13, Tsi81, TRM22, UPÖ16, WC15, Wan22a, Wil81, Wim88, XCB18, YLS10, YZ13, YBCW96, YW05, YW09, ZZ16, Zha20a, ZZL23, Zhu18]. generalizing [KS16]. generate [BT17]. generated [Ber05, CKKK22b, GPK94, KM19, KCBR87, KLS08, Law19, MP87a, RS07, Sch79, Tre85, Tre86, Tre87, Wad90, Zen22]. Generating [New85a, BdPNdP17, CSG22, Lew77, LM19, Lop11a, RW84, Spi76]. generation [Dre93]. generator [BTV07, CFSRL22, FTA20, GC22, Mah82]. Generators [TW17, Aln11, Chi22, Gai77, GC22, Lop18, NTW23, SL15, Wat89]. Generic [Wil87, BFZ07, FW86, For79d, Lei83, MMRR16, PS91, dTS20]. genericity [BGLSSS20]. genetic [Pan17]. Genus [BS20, Ger82]. Geodesic [CGNT15, RBB22]. geodesically [KJF23]. Geometric [CKS20, CF15, Sei95, ZZ00, AG22, ASR22, Dra18a, EJD88, EB17, FSY16, FS18b, GHT11, JK10, JKL22, JST18, Lim23, LS17, Pep19a, Pep19b, ST94, Sch22, Smi74, SY23, XZ22]. Geometrical [BD94, Pan15, CL94, MM06, SCL^{+}22]. geometries [Kes83, RQ22]. Geometry [CGHK08, HW06, GL16, Jam88, Lei93c, LWC22, NP91, NF19, Wit11]. German [SU77]. Geronimus [GAGM18]. Gerschgorin [AB78]. Gersgorin [EV77, LLL15]. Gersgorin-type [LLL15]. GIBS [DK23]. girth [LDL17, LOO22]. give [Kap77]. given [BS21, BJS98, CH22a, CFLL23, CJL22, DJ89b, FTZ08, Far22, FLZ07, GWW20, GJMW86, GZY18, HZG14, HLZ05, JST18, KS19a, KK90, LT12a, LC14, LWL21, LY22d, LOO22, Moo76, SM94, SYZL17, Slo14a, SKLS20, SKLS21, WH11, WWWT22, Wu09, XSW21, XSW22, XW19, YS23, ZXL22, ZL22]. glimpse [MGH19]. Global [BVW84, CDF21, DK14, Mil15]. GLT [BG23a]. Gluing [CTY20]. GM [Bak16]. GMRES [Doo16]. Goldberg [MS82b]. Golden [Hia94, TWK22]. Goldman [Bla77]. good [BDV19, Hal91, La 19, WC22a]. Gordon [Mer79b].

Gorenstein [GTLN18]. Graded [Din15b, NT22, ACO21, BS18a, CCGO11, Cen11, CdM15, CdSS17, Egg95, GMV18, dSPdS22, RZ17, dNV19]. Gradient [LT13]. gradients [CG17, Nie22b]. gradings [BDV19, FGK23, RE22]. grafting [LLF08, ZLS11]. gram [Pat82]. grand [BS22a]. Graph [Con90, HZ20, Mer94, AN18a, AA21b, AN14, AFTG19, AADdF23, ABGP21, AM85, ACMR17, Ano79b, AH16a, AAA22, Ari16, AT18, Ash16, BC05, BP98a, BKP01, BS11, BR14, BS20, BHH^{+}11, CT16, CHL18, Cla92, CTY20, CP10, CD85, CR90, Cve95, Da 05, DFK19, Das04, Das17, Das18, Das21, DM22a, DeV22, DIJ23, Dua22, Fan05, FF21, Gon17, Guo07, GZ22a, Guz88, HZ11, HLZ22, HHY22, HM23, HJ77, HS16, HUZ17, HZ21, JS22, JWBB22, JKS79, JD99, JS08, JST18, Kir00, Kir14, KMMS22, Koo95, KL92, LZ15a, LV18, LP00, LLF08, LZX19, LY22b, LPWX10, LXG17, LZB23, MW16, MW18, MF20a, MH22, MT22b, MKS22, Mer95, Mer98b, Mir23, MA15c, Mum86b, OWTZ21, PLH01, PLLL20, PRS03, PGAB22, QX23, RA21, Ric78, Sar20, SMM14, SMM16, SS22, So99, SL95, zSqSZ16, Ter01]. graph [Ter06, Ter09, Tho90, TW17, TLC22, TJ21, UZ16, Wan16a, Wan19a, WWZ20, Wan21, Wat90, WL20b, WLM21, WL23, XWT22, XZS19, YWZ21, ZH22, ZWL15, ZWM17a, ZWM17b, ZWS20, Zhu15, LPS21]. graphic [Kas85]. Graphs [AL22, ASZ02, FKP03, HS95, HLP03, KP02, LS09, LZ15b, LW22b, MSZ04, Moh16a, RP90, TB92, Yiz02, Yiz03, tHK17, AAT17, AHJ20, AHJ21, AGM13, ADHM19, AGS20, AAGL22, AGH^{+}22, AMK23, AAKS21, AADdF23, ABP22, AAB20, AL21, AFW22, AAdFS19, ACMR17, Ano16b, APEPPS23, Are22, Are23, AOTR11, BBG22b, BBG22c, BGK99, BLP12, BP19, Bar08, BGM22, BCLS23, BCDF09, BS22b, Bel15, BPS16, BPW16, BB19b, BBCD21, BVW84, BCEM12, BZ00, BSA17, Bru82, BS93, BSG^{+}12, BH20, But10b, CMR20, CMR21, CF21, CP15, CL17c, Che18, CHQ18, CZ21, CH22a, Che22b, CFLL23, CFL^{+}22b, CFL22a, CKM18, CKMY22, CKKY23, CDG23, CG16, CSCH01, DFK19, DFPS23, DP17, DAH19, DGR08, DIF13, DWL20, DZ22, DJS21, EL21, FTZ08, FDD14, FGHL22, FS14, FS18a, FKS22, FYIS13, FdF09, Fer10a, Fie90b]. graphs [FG06, Fri93, FHT14, GN20, GL11, GWW20, Gho19, Gol06, Gol16, Gol21, GH22, GM90, Gro95, GWZ14, GWZ16, GB11, GLS16, GZY18, GLLM21, GZ22b, HRMP95, Hou01, HW08, HLW^{+}21, HHW15, HL18, HLB22, HL22, HZ03, JC22a, JLS09, JKKC22, KS19a, KS22a, KS23, KP22, KT95, KK19, KOSS23, KF98, KMN01, KADM10, KS17a, LV18, Law19, Lee98, LZ11b, LT11, LLP18, LY22c, LS22b, LP23, LS12, LS13, LLD17, LXS21, LL10b, LZ14a, LZ14b, LDL17, LLF18, LLGX22, LOO22, LH23, LW23, LCG^{+}23, Lor00, LHH18, LHF22, Lu22, LCXC22, LCWW22, MW15, MH16, MGKA23, Maj22, MS16a, MW23, MGA17, Mer97, MS10c, MTR11, MN04, Mol18, MR19, Mou20, NS15, NS09, Nat13, NP15, NC22, NP09, NG15, Obo16, Obo22, Ost74, PB17, Pal21, PP16, Pan17, PGLM00, PRAC22, PSB23, PV23, Pow88, QM21]. graphs [RP22, Ram22, RM22, RC22, RAN22, REG07, RY93, Sat11, SS79, Sch95, SdF12, SS23, Sei95, Sha15, SHS17, Shr90, Sim95, SS07, SRD19, SR22, SM19, Sjo93, Sri20, Sta10, Sta20, Sta22, Ste94a, Ste94b, Ste05, SD20, Tan12, TM19, TWZ16, TW19, VH95, Vo83a, Vo83b, Vo83c, WH11, WZ13, WBZ14, WW15, WT16, WM17, WWS19, WMFW21, WC22a, WY20, WYSL22, Wil93, Wu05, WL18b, XZL14, XYH22, XZW22, XLS20, YW19, YLW22, YZB22, YHS13, YGW15, YSZW22, YS23, ZLS11, ZXL22, ZQC18, ZG11, ZZZ12, ZZ15, ZY17, ZZZ23, ZLZ20, ZG08, ZB13, ZW17a, ZW17b, ZD18, ZZ21, dLMO17, dlP09]. Grassmann [Car88, Car90, Cha79b, DK11, De 08, Die79, Din15b, FGK23, GP15, Gor83, Wan82, Wes74, dB10]. Grassmann-embeddings [De 08]. Grassmannian [LCH22, MR75, PT10]. Grassmannians [Pas16]. gray [CW93]. greater [Ano16b, BPS16, PGLM00]. Greatest [Bar80]. Green [CMM20, Put82a]. Gribov [CDJ13]. grids [BG23a]. Griffin [Ric87a, Ric85]. Griggs [Har84a]. Grothendieck [PSS15]. Group [BT98, CKN95, DS02b, Pat96, SQM99, yT00, ZBW12, Abd09, AF12, AP23, Ano16c, BB06, BAN12, BT06, BLZ10, BLZ11, BSÜÖ13, BH15a, Bot79, BZZ10, Cha90a, CCS98, CNW18, CHZ22, CZZH22, CFL22a, Chu84a, CW93, DP00, Ell93, Ell95, ES76b, Fal97, Far90, FJK14, FHZ17, FP17, Fri85, GA85, GMV18, HNB22, HP18, HK13, Hun83, Hun86, JGK79, KW92, Lew77, Lip87, LWY11, LWB11b, MW16, MN80, Mah82, Mal92, Mar84, MT11, MP16a, Mol18, MS15b, MS16b, Ng22, Odo83, Ous15b, OS18, Pat89, PS21, Put82b, PH97, QX23, Rai87, RP04, Sch75, Sch78, Shc21, SX12, TW17, Wan80, WZL13, Wan16a, XWT22, Yam18, YQ90, ZWM17a, ZCZT22]. groupoids [RD12]. Groups [Gur97, KMEO12, LM03, AF11b, Adu15, AL21, AFW22, AKK^{+}23, Are23, Ber05, Bri89, BS07, CF21, Cha77, Cha79a, CL86a, CL86b, CP15, CFL^{+}22b, Den05, – Dok81, Don91, Dre93, Fon92, Fri97, GL16, GT79, GL97, HNB22, HT92, Hol04, HS17, HW75, JC22a, Kim91, Koh11, Kra05, Kra06, LLT93, LW23, LCXC22, LCWW22, Mah82, Mal95, MO22, MC80, Mar00, MZ17, MGA17, Ner22, NP73, Nie22a, Ota13, Pan13, Pat82, dSPdS22, Pre22, RA21, Rei75, RE22, RC18, Sla16, Slo13b, Sta23, Tam97, Til77, Vin06, Wal79a, WC22a, Wat89, WW20b, Woo22, dB10]. Grover [KMMS22]. growth [CdSS17, FdSV16, GdSV16, IM18, Iop19]. Grüss [Ota20, DM17]. guage [HWL92]. Guessab [TR21]. GUS [LLV18]. GUS-property [LLV18]. GW [Hai88a, Hai88b]. GW-rootsystems [Hai88a, Hai88b]. gyroscopic [YZC21].

H [Cha84b, MP13a, Sun20, WSZL17, Wan77]. H-tensors [Sun20, WSZL17]. H. [Cha83a, For79c, Jia92, Krä87, Mal89, New83]. Haagerup [MW19, MW21]. Haar [OS18]. Hadamard [AHJ87, Aue98, Bap91, BG21, Boz99, CS23, Che22a, Cho18b, Cra91, – Dok17, Du10, EM95, EF98, ES76a, FKR93, Fri75, GA20, GP74, GN02, GS20b, Gór22, GJMW84, GMW86, GM87a, HB83, Hey92, HJ87, Hua08, HXL16, HK17, Joh77a, Joh74b, Joh75a, Joh77b, JM85, JN90a, JN90b, JN00, KLZ10, KMS02, KLJ23, LHS^{+}21, MKS84b, MF92, Mat92, MTR19, Ota20, Pep19a, Pep19b, SAV01, SZZY18, Wal73, Wal86, WZ97, Zha00, Zha89]. hafnians [BFI22]. Hahn [DPR22, TRM22]. Hahn-Banach [DPR22]. half [DS16, ZZT23]. halfplane [Rod08]. hall [LB99, HL10]. Halmos [Zha95]. Hamburger [ACJ17]. Hamilton [LN16, Paz84, Van95, XZW22, ZW17b]. Hamilton-connected [XZW22, ZW17b]. Hamiltonian [CQH15, CHQ18, GN20, GZY18, LP23, Mer83b, NG15, Nof15, NF19, RY93, SHC22, WC12, ZW17a]. hamiltonicity [Lu22, LLP18]. Handelman [KNX95, KNX93]. Hankel [BG83, Boz99, CCLL22, Ikr15b, JTW17, KS16, Lew97, LC20, LM19, MHH96, RRC23, Sol06]. Hankel-symmetric [CCLL22]. hard [GLM^{+}22, Roh00]. Hardy [BS22a, CDPS23, YGÇ19]. Harmonic [KK19, ST94, Web83, XZ22]. harmonious [FXY22]. Harnack [CB22, LZ17b]. Hartwig [DD14]. Hasse [Bah22, Gar75a]. Hausdorff [BB14, BBM20, BMFPBSS21, CW11a, DST91, DTS14, NST89, TD14, Tal17]. having [BCEM12, BCS14, DH91, Est80, GH22, HZH20, JGK79, JS84, JOvdD84, LD19, LCXC22, MPGS21, Šem97, Wan22b]. Hecke [LLT93, Mit10, Zha21]. Height [HS89, AHK92, HS91]. Heinz [ASR22, Ala16, BM15b, GHT11, KM11, KK13, LS15, Man17, NS18, SYK15, Sab19, SG18]. Heinz-type [Sab19]. Heisenberg [AC19, CPS22, CHS22, LS22a, Nie22a]. Henryk [Mar03]. heptadiagonal [Lit17]. Hereditary [BY19, EGT16, KM98]. Hermite [BFZ07, CS18, Sin93]. Hermitian [– Dok93, AM13, AAB17, AY93, AB78, AY74, BT08a, BY76, BL83, BM16, Bur11, CM92, CTW15, CK20, De 08, – DL92, Dra17a, DNY22, Elg22, ESH23, FdF09, FH89, Gro97, GJMW86, HR85, HNS99, HW06, Hua22c, HM95, JN80, JR84, JP86, JL91, JS08, Kim91, KN81, LL16a, LP03, LW17, LS22c, Loe87, LSX19, Mas10, Mir92, NTU93, NT94, Oi20, OK07, Pat92, Pat03b, Sar20, ST73, ST91, Sta15a, Tho92b, TL12, Tis93, Tra84, WZH13a, WZH13b, WL20b, WLM21, Woe97, WL22, YLW15, YTL20, Zha04c, ZL21, Zha22, ZCLW22]. Heron [DDF22a]. Herstein [AEV12, IT16]. Hessenberg [KP14a, MR15, Mar20]. Hiai [HSW19, KMS21, Tom22]. Hiai-type [Tom22]. hidden [JDD21]. hierarchical [BCDF09, GLLM21]. high [LZB23, SWL13, WG91a]. Higher [AM21, CKL^{+}21, EGMS22, HWX22, BdP12, Car20, CGW14, CL17b, CLN21, CGHK08, CLMR09, DTSI15, FHL^{+}14, HLYZ22, Jon78, KP14b, Li87c, LPS09, LPS13, MC86, MP73, Mer78c, MS15b, MS16b, O’R15, Woe08, XW12, Xu18, YLX20]. higher-dimensional [CL17b]. Higher-distance [EGMS22]. higher-order [Xu18]. higher-rank [BdP12, CGW14, CGHK08]. Highly [DM22d, Sma17, Zha20a]. Hilbert [HJ19a, Alo21, Alo23, AJL22, ACMS10, AR15, AB17, AO22, AY84, BDK21, BCNG23, BHK21, BR04, CF23, CDF21, DLD21, DKM10, Dra10, Dra11, Dra16, Dra18a, DK22, FMX18, FZ21, FXY22, FMXZ22, GGGA21, GS20a, GNO20, GJR22, GBM21, GLL21, HJ19b, HJ20b, JKH20, JS15, KS16, LLHS18, LL20, MAT22, Mos19, MR10, OM21, Roo20, ST19, SA10, Son14, Tau73a, WTD22, WZ22, WZ23, Xia20, YGÇ19, ZMCN19]. Hilbert-modules [AB17]. Hilbertian [BFM20, BN22, MSS13, QHB22, RSM21, SM23]. Hille [JRMFMASS16, MFPSS12]. Hirano [CS21a]. historical [Pri78, TT78]. history [HPS83]. Hitting [Che18, GLLM21, HL18, ZL22, ZZ21].

Hochschild [Hou17]. Höder [AO97]. Höder-type [AO97]. Hodge [Sam78]. Hoffman [MV91]. hold [BR22, MS10a]. Hölder [Dra18a, LPW17, Mer81, Mer78a]. holds [MO91]. hollow [FJ15, Lim23]. Holomorphic [BGLSSS20, MRS10, TN22]. Hom [AC19, ABM22, CKR18, CR18, CKR21, CZ23, GZW20, SX15]. Hom-left-symmetric [GZW20]. homeomorphisms [Pan15]. homogeneous [BB17a, BMP10, BPR12, FGMSS19, Fri97, FM22c, HSX16, LW23, LT07, MO93, ME18b, PSW18, Rom22a, Soa08, Sun20, Vel21, XSL15]. Homomorphisms [DK16, Tho92c, AAB15, An18b, Ben05, BG23b, BM16, BMGMC15, EPR16a, EPR16b, HZ20, LZ11a, Lud21a, ML12, ST22, TOM88]. homotheties [RRR^{+}22]. Hook [CW81, Pat95, Hey88, JP87, Mer83b, Rem90, TOM88]. Hopf [Adu15, Egg95, FJK14, LW20, Lla96, MZ04, Rod22, Zhu22]. Hori [Vei83a]. horizontal [HZV23]. Horn [Hua23, Mas10]. HSS [Hua22c]. Hua [XXZ11, And80a, LW11b]. Huang [Uhl13]. Hull [And04, Gil12, IOMAN21, KSMB16, LT89a, Poo94, Tho75a, TI77]. hulls [AS11a, SA07, Tho75b, TT18]. Hungary [BL80]. Hurwitz [Ang00, AYC93, CCL10, Kri87, MGAH19]. hybrid [CKL^{+}21, fLLyP15]. hyper [BT17]. hyper-ideals [BT17]. hyperbolic [CN21, HK10, LL17, RBB22]. hyperbolicity [BV16a, Dra21a]. hypercomplex [Lei93c]. hypercubic [GR10]. Hypercyclic [Shk12]. hyperdeterminant [Bre12a]. Hyperdeterminants [Luq08, WHL16, Sun15]. hypergeneralized [TCI13]. hypergeometric [DS18, DS19b, Sal14, VY21]. hypergraph [SQH15, XQ16a]. Hypergraphs [BG79, Rod02, Rod03, BS08, CT21, GMV23, LCC17, LZL18, LGZ20, LW22a, LL22b, SSfW15, WZ18, WLF22, XW19, ZL19]. hypergroups [BS91, Bha93]. hypermatrices [Ait19, Ilm16]. hypermatrix [CD15]. Hypernormal [Mal94]. hypersurface [BFM22]. Hypersurfaces [Pat84]. Hypertournament [Kir91]. hypertournaments [KKP13]. hypertrees [CBZ22, LWL21, SKLS20, SKLS21]. hyponormal [BK21, CH10, KY17, Ras17, Ras19]. Hyponormality [GHL20b, GBM21, KKY22]. Hypotheses [Red02].

ic [Chi78]. ID* [LC22]. ID*-superderivations [LC22]. Ideal [GS20a, Rao81, AVVÁ18, Bou06, Dix79, DK14, MW16, RM22]. ideal-relation [MW16]. ideal-triangularizability [DK14]. idealizations [DW16]. Ideals [Vel18, AN14, Ake83, BY19, BCI13, BV92, BRL22, BJZ13, BT17, DS23a, DDA20, – Dok73, KM14, LL11, LCL14, Pat89, RS85, Tow22, XWT22]. Idempotence [ZCZ12, TXC08]. idempotence-preserving [TXC08]. Idempotency [BT08b, Pet10, Xu15]. Idempotent [Bot96, CZ22, Ma15a, AS95, AJL18, AH10, BB06, BT22a, BT08b, Cat21, Foš05a, Fur08, GLX22, GH98, Laf83, LWY11, Tia11a, Tia11c, Wri22b, You15]. Idempotents [ÁBvW16, Dol04, Dre22, Hou21, KR02, KR03, RR02, AF11b, AJL14, BB20, Ben15, Bun94, CS10, CG19, FL16a, FD12, Gha12, HP90a, HP90b, HK16b, Hua96, Slo14b, TZS19, XHL23, YQ90]. identical [Che22a]. Identifiability [Bal22b, GOS22, ABC18, YLX20]. identifiable [AC23]. identify [CAP22]. identifying [HM19, Sun20, YZH19]. Identities [AAB20, AELL20, For79a, GdSV16, RZ17, BM18a, BP09, BE20, Car06, CdM15, CdSS17, CD22, Chu22, Din15b, Dom95, Düf97, Ere15, GV95, HNB22, Hon92, Hon93, KV13, Lee16, NT22, Reg87b, Reg89, Reg93, Rem88, TC04, UF22, Wan19b, Wen06, ZZ23, dC13, dC15, dC18]. Identity [Zha04a, Bre14, DDA20, – Dok75, Fei79, Gil14, GV87, Gur83, JOTvdD94, MS85a, Mer74, Pat08, Reg87a, Ste82, Whi80, Wil81]. Ihre [SU77]. II [Ano79b, MS21b, AF11b, AS11a, Bea83b, Ben74, BKL97, CL92, Cha93a, DM14, Fan83, Fon89, Gar75a, GN20, GKLR82, GGL03, GTM14, Gur97, Hil11, HW06, HZ20, JKS79, JP86, KK75, KS07, KL92, Lim90b, Loe74, Loe93b, MM76, Mer78c, Mer98b, MP95, MAH21, NP80, Nie07, NS22, PS22, Rho90, RS10, Sch78, SGM22, Sil88a, SHC02, TLL11, Tau74, Wea83, Wes95]. III [AF12, Bea84, BH74, GL97, MM92, Wes74]. image [Lim00, Pam87]. images [UF22, WLB19, Wan19c]. Imaginärteil [SU77]. imaginary [CCL10, EJ91, GHK16, HHC12, SU77]. imbedding [Adk95]. Immanant [CA03, CD98, CD01, Do 96, Gro85, Ito16]. Immanants [Jam92, Pat90, Gam00, GMW86, GM87a, Har85, JP87, Pat95, RS10, Wat86]. immanent [Fer10b]. Immant [Hey88]. Implicit [JSW09, HZV23]. Imprimitive [MM11, FW07]. imprimitivity [BP92]. improper [Tho75b]. Improved [LW17, ZV19, CPLH17, GB77, Li21a, MF20a]. improvement [GLL18, XZW22]. Improvements [HLB20, BP22]. Improving [Dra19a, GJMW84]. Inaccurate [Ned98]. incidence [AGS20, Bap97, BL11, BFJ12, BFK18, CD21, DW16, HZ11, HM23, Khr16, Kol23, KH05, Mum86b, Rys73, Rys75, Spi04, WLF22, Yan21, ZZ15]. including [LL14a]. Inclusion [GZ76, JR93c, JR94, LS04, PT87, Das18, FMX18, LXQ21, Lei18, LST89, MW18, ZL18, Zha19, ZZT23]. incoherent [YGZC22]. incomplete [GKO22, VY21]. Increasable [Esh22a]. increase [Kuz05]. increasing [Cho10, Nie22b]. Indecomposable [LF16, BB15b, HK16a, Har87, HK16b, HL10, Liu92]. Indefinite [BLdPS04, BdP12, BdPN14, LZ19b, LMMR01a, LMMR01b, Rod81]. Independence [MW18, HHY22, LZX19, SD20, Tan18a]. Independences [JX13]. independent [Bra86, FTA20, Mer78b, Pat83]. Indeterminates [Rys73]. index [BB22, BP92, BS95, dAB17, BG16, BF15, CMM17, CM15, EDKA07, GWW20, GM18, GM21, HRT15, Jun88, Kir93, KS17a, LLT93, LS22b, LLL09, LHbL^{+}15, MC82, MM07, MM09, MQ21, Mer89, MZ23, PJ90, Raj09, RRR93, RY93, Sha15, WL18a, Ye11, YGW15, ZZZ23, dB14a]. indices [AEG19, BFZ07, Bo01, DF95, Dua22, LS22c, MZ04, WL23, ZH22, ZCLW22]. Induced [CL80, Sch95, Tik22, ABM22, BMN22, CKS19, CL85, Da 88, Gau76, HWL92, JN00, KL95, Lei93a, Lei93b, Mal95, MG22, MO94, SS23, Ste94b, Tam88, WLF22]. Inequalites [SH98]. Inequalities [Ala16, And80b, Cho18a, Cho18b, ES73, FLT22, Fur11, Haw05, HMN91, HLPW17, JK22, JL88, JSG94, KS19b, Lee91, Lev74, Li91, MS74, Pat82, Pep19a, Pep19b, SAV01, XZ22, AC78, AR22, AAP82, ANK22, ARK22, ASR22, AkK17, AS20b, AW07, Alo21, Alo23, AHK19, AdFK15, And80a, AO97, Anj17, AA12, BM15a, BM19a, BS22a, BDK09, BDK21, BP98b, BO22a, BCK15, BKL97, BBP21, BP22, Bol12, BHK10, Bun91, CX23, CLL14, CL17a, DM17, DHHK23, DV18, DW23, DKM10, Dra10, Dra16, Dra18a, Dra21b, Dra22a, DL14, Du10, Fan92, FH15b, FH15a, FG22, FJK98, FSY16, FMS19, GA18b, Gha21, GHT11, Gur98, Gyi78, HLB20, Han17, Han19, HB83, Hia94, HSW19, Hla17, HOT15, How80, Hua08, HXL16, HZX22, JKL22, JM85, JN14b, Kan87, KM85, Kan94, KMS18]. inequalities [KMN22, KM22, Kim91, KS07, KM11, KK13, KT83, KW89, KLJ23, Lék19, LM96, LPS12, LLZ13, LH21a, LS15, Lim00, LF21, Man17, ML21b, MM22a, MOGN18, MS21a, MS21b, MFS22, Mos12, Nie07, Nie22b, Nie22c, OU22, Ota20, QHB22, Raï19, RK94, RSM21, RSM22, SYK15, Sab19, SM21, SRS21, Sed11, SG18, Sha18, Sha20, SN23, SA10, SG96, ST86, SZZY18, TJG22, Tho78, Tia05b, Tom22, Tra84, WZ97, WZ14, YK22, YLS10, YLC19b, YSK15, ZW19, ZCLW22, Zou20]. Inequality [Pat81, AFTG19, AMK23, Ala18, AB78, AHJ87, Aud15, BM15a, BM15b, Bak16, Bap88, Bap91, BdP21, CF23, Dia79, Dra11, Dra13a, Dra17a, Dra19a, Dra19b, Dru14, ES76a, Fri75, FL16b, FS18b, GAMA23, GJ22, GJMW84, Gro85, GM87a, Gum15, Ha88, Hey92, HKO16, HM22, HM95, JRMFMASS16, Joh77a, JN90b, Joh90, JZ95, KS14b, KMS21, Kua18, Li87b, fLLyP15, LCH22, LP22a, LW12, LZ17b, LPW17, Lon83, Lon97, Mar76, MS88, MV91, Mer94, Mia93a, MP94, MP95, MP96, MGH19, MF20b, MR10, MFPSS12, NS18, NASR19, Oku91, PSA14, Pat96, Pat09, Pep11, Rea96, SY23, TWK22, Tho89, Tik98, VM18, Wan79, Wim85, Wu22, YGÇ19, zY11, YLC19a, YqS16, Zha01b, ZZZ12, Zou20]. Inertia [GS06, GA18a, GA20, JR84, LRT91, WL23, ZCLW22, AL78, BS97, CM92, CFSGL18, Dua22, GWW20, JP86, Loe92, MF92, ZH22]. inertia-preserving [BS97]. inertial [ZZT23]. Inertially [GS01]. Inertias [KMOV07, JL91, KOvDD09, TL12]. inexact [WZY22]. infimum [Dji16]. Infinite [ALPM20, DvE20b, DS02b, GK02, KK75, KMS02, ME21, WC16, AELPM18, All14, Bar07, BB14, BDH23, BH15a, Cas91, CP10, DvE20a, Esh22a, Far22, Gol16, HLZ18, JO23, KPPR21, LF15, Lim23, Liv94, LRD21, Lud21a, MDP22, Nag17, Ng14, Ng22, Pop14, Rai87, RD12, SHC22, Siv06a, Siv06b, Slo13a, Slo14b, Slo15b, Slo15a, Slo16b, Slo17, Sło18a, Sło18b, Sol06, WC12]. Infinite-dimensional [ALPM20, KK75, AELPM18, Bar07, BB14, JO23, LF15, Liv94, MDP22, SHC22, WC12]. infinity [DM87, HM78, MZ04]. inflations [GW98a]. influence [Sch77]. Influences [CP16]. Information [JS91, BL80, DMS23, DHKSH08, FHLS13, Fur11, MF20b, ShDHG08]. Inheritance [BS17e, BS18b, Pan17]. Inner [BLdPS04, Tan16, Dra19a, Dra19b, FD14, HLP23, LMR94b, LMMR01a, LMMR01b, LB21, Ota20, Pat08, SN23, TZ19a, Tan13b, Uhl82]. Input [Har94]. inputs [GOS22]. insertion [Cas23]. instability [KSU22]. integer [BT07, BM18c, BM19b, CZ22, GA20, GRMA13, LSOvdD95]. integers [New85a]. Integral [CFL22a, Lee18, LYZW12, Yiz02, Yiz03, BHK78, BSG^{+}12, CQ18, CL17a, CFL^{+}22b, Dra21b, GHL16, HK79, HK80, Hsi78, HHW15, IAM23, Kir04, KADM10, Kri87, Kut19, Luo19, Mah82, Mir23, MTR11, New74, Tau81, Tho86a, Tho86c, Tho87a, Tho88, Vas86, Wei93, YW09, cc85]. integrality [MM05]. integrals [BC06, Sim83]. interesting [Bru85]. Interlacing [Mar90, CM84, GT11, ISCS11, LM98, LPS12, NU94]. intermediate [BK21, WOY06]. Intermixing [Dra23]. Internal [BC83]. Interpolated [SYK15, YSK15]. Interpolating [AkK17, CJ12, KK16]. Interpolation [JvdD88, LP11, OU22, Ala18, AG15, BCR91, CH12, DMV18, DWL13, ZZ00]. interpretation [ER90, RS18]. Intersection [JNBR93, AN14, BCS14, HS83c, Tho75b, XWT22]. interval [Aze99, CR22b, Har96, Hla17, HWL21, KSMB16, KP11, RR95, RK94, Roh06, Rub20, Ska17, TXL21, XLL15]. Intervals [AG16, AAKS21]. intraclass [Mai11]. intransitive [CL86a]. Introduction [Abo20, AL94, Row95, KP14b, Pie92b, Tra84]. inv [Gar79]. Invariance [MF77, SSZ93, BK83, Bea78, Gro96, Hem93, LT88a, LFY16, Man91, Min76, WW20b, XQ15]. Invariant [BR04, De 90, Dom14, Ger74, MZ04, NT86, Pat77, Sil98, Tho87a, AE10, Aze99, Bae84, BG86, Ben78, CW81, CDJ13, CLW15, CLR81, Fri90a, GP22b, Heo16, HJ20b, HS87, HJ87, HMN91, JS94, JP15b, JL88, JN92, KMEO12, LT87, LT89a, LT89b, LT90b, LT90c, Li91, LT92c, LT93, LLW07, MKS84b, Mar90, McD92a, MMRR09, NS85, Pok87, RRR93, Ras19, Rod81, Sch22, Sel81, SX23, Sil88a, Sil88b, SS96, SAV01, Sin73, Sol80, Ste82, Wan17b, XFG14, Zab89, Zab90]. Invariants [Gur97, Mau78, PW78, ARH15, AMZ13, BZ99, Bed10, Bed11, BX11, BCdHP10, BB17c, CLX22, CNW18, CN15, – Dok09, Dom00, Dre93, Fog73, GA85, GHV99, GL97, KP92, Lop11a, Lop11b, Lop18, MZ17, MM23, RH11b, Sch75, Sch78, Tho88, Wat83, dTS20]. inventory [Mer83a]. Inverse [BS11, CLX03, HA22, HS16, LLS14, MNN^{+}98, Min82a, Mir18, ZHZ11, AHK92, BT10, BT22a, BST23, Bap97, BCN05, BP19, BAN12, BLZ11, BB17b, BBX20, BS17e, Boa19, Bou18, BDD23, BT98, BZZ10, BZZ11, BSZW13, CC17a, Cam80, Cam83, Che95a, CS19, CS21a, CHZ22, CS22, CZWF23, Chu22, Cro18, CICX09, CIN15, DM22a, DMV18, Dec74, DK23, DD10, DCIW10, DLW18, DS16, DK22, ER90, FS18a, FdF09, FOT23, FSWX22, GCP20, GCP21, GZZF23, Gar23, Gar85, Gro99, HLZ22, HP18, HM23, HM78, HXL16, HPV84, HPV88, JL93, KMRS89, KS22b, KCSĆ21, KW11, KR07, Kyr11, LCZW19, LZ23, LP16, LWY11, LYZW12, LPH17, LL78, ML21a, Ma22a, Ma22b, MP16a, MS18, MZ15, Men22, MMG15, Mol18, MS15a, Mos15, Mos19, MM22b, MZ22, MAM15, PM22, PJ90, PM14, PH97, Rea96, Rom19]. inverse [Rom22b, SSZ22, SM20, SN22b, SX22, SC20, Siv06b, Smi05, SPM14, SPK15, SPKC15, SKM17, SW22, SZBW16, Tia05a, WL15, WfMR16, WLL19, WLL21, WM21, Wei98, mWyD11, XQ16b, Xu93, XSH18, XCM20, YZ14, YW05, YW09, YZC21, ZFSX17, ZMG20, ZBW12, ZCP16, ZPCZ16, ZCP17, ZIC22, vR11, LLS16]. Inverses [BP74, GN02, KS22a, MR15, Rao81, Zha04b, AELL20, BI95, BHK78, BM17a, BS22c, BT06, BL13, BGV11, BT15, BGZW16, CC22a, CKN95, CD21, CC21, CZZH22, CIM19, Dra13b, Dra17b, Dra18b, Dra23, FLT20, GK18, HK79, HK80, HLT22, HM76, HM77, JGK79, KS23, KS15, KM14, KCICV18, KR14, KD23, Kyr08, LPT13, Li11, LC18a, LW11b, LW08, LB21, MR13, MP13a, Mia93b, MS12, Mos20, Mos21a, Mos21b, MW22, Nab98, Pab22, PP16, PBM20, Rad16, Rom20, Rom21, Rom22b, Siv06a, Smi74, Sol14, Sta15a, SKP18, SĆK^{+}20, SMM22, SCL^{+}22, SPS15, Tia01, TC04, TS05, TT06, Tia10, Tia18, Tia21, VM07, WC15, XCC99, XQZ13, XCB18, YZZ20, YLS10, Zar01, ZWW22, ZX10, ZCW21, ZC22, ZCZT22, Zhu18, ZP19, Zhu19, ZW20, ZWM23]. Inversion [JTW17, Elo18, FZ11, JS91, Jon78, lLZjZ19, Mar20, Tis93, Wer86]. Invertibility [BLR12, CH22b, Den14, FSvW11, IK18, KR02, KR03, MAT22, SR17, Wat83, Xia20, Zel75, Aue95, Aue98, BLZ10, Bou13a, BMGMC15, CQH15, CIP18, Cve18, CM22, DW12a, DD12, DK22, GNO20, HC14, LD17, LC20, MŠ09, PA10, SA17, TCI13, WHC22, ZZ16, ZCCIW12]. Invertible [BLSV12, Gur97, GL97, BSÜÖ13, BDH23, Che07, DLD22, Dra22b, DK21, Fra13, Gow80, GHL16, HBM15, lLpW16, LWB11b, Siv06a, Siv06b, Tan13a, Tan18b, XPY16, XZ18, ZCW21, ZZC17]. Inverting [KHG13, FdM18, dFSY22]. involution [AF11a, AF11b, AF12, BCI13, BB17b, HE20, HPV88, Iop19, KR07, Lee15, LC18a, NTW23, QZ16, RD12, UG21, Xu15, ZC22, ZP19, ZWM23, dFSY22, dNV19]. Involutions [Slo13b, BL11, BFJ12, Fie95, FdM18, Hou22, KL21b, LY22a, LK21, TZS19]. involutive [LK23, LWB11a]. involutory [BB06, Bal75, Bal76, Bal77, Bal79c, BT98, PA10, Tia11c, ZW14]. Involving [Pat81, ASR22, AHK19, Du10, FG22, Gha21, GLLM21, JK22, JKD16, Joh77b, JSG94, JR94, Kar23, KMS18, LM96, LW11b, LLF18, Mer74, PLLL20, Raï14a, Raï14b, Raï19, TX13, TG23, XJS20, XHL23]. iota [FKC19]. irrational [DH91]. Irreducibility [MM05, Mer75, Mur89]. Irreducible [CY22a, GT14, Lla96, Min74b, Abd09, BCS14, Bru79a, CCU21, Coh88, DF87, DLL19, Els78, Fon89, Fri12, Hol04, Hua90, Hua22a, Hun83, Hun86, JP15b, KSMB16, Kra05, LZ09, Li95, Min74a, Pat93, Pul75, Roh06, SZWB16, TB92, lWW15, WWW20, YSY17]. irregular [LGZ20, SHS17]. irregularity [JWBB22]. Isolated [BS08, SRW07]. Isometric [KLL22, GJR22, MCL20, MP18, MP19]. Isometries [– DLR91, KL95, LŠ02, MN12, – Dok90, GWW16, HS83b, HFS15, HF18, Hos22, IK19, JS20, LT88d, MWLW22, Mol15, Oi20, SM23, ZCZT22]. isometry [CP16]. Isomorph [Whi74]. Isomorphism [ARH15, Oge94, Cve95, Dix79, RH11b]. Isomorphisms [BDvW12, BD12, BDvW16, BFK18, CMR21, DPvW13, IT16, LWXF15, Pan15, Wan16d, XWF15]. isospectral [Mer97]. isotropic [JMS19]. Isotropy [Sta23]. Issue [KL17, KL21a, LT09]. issues [BBM20]. Itô [AM15]. Iteration [Fan83, CH06, CL22b, HZV23, LZ16, TT15, Xu18, ZV19]. iterations [Mis17]. Iterative [CC17a, DH86, DCLW19, GLM^{+}22, LLC21, WX19, AAB17, ALP94, BS17d, BNKJ22, CW20, DK23, FTA20, Gib89, HM20, HM22, Kul22, LC18b, PLPC19, Xio23, YZH19, ZWZ23]. Ito [HZG14]. IV [Chu90]. Iwahori [Mit10].

J [New83]. J. [Mer79b]. Jacobi [All14, BB21, CKN95, GPM14, HK21, Mir18, MB22b]. Jacobian [Wri93, Yu96]. Jacobians [Li21b, Neu83]. Jacobson [YZZ20]. Jain [KKS21]. James [AV22, BBK22, CP13]. James-orthogonality [CP13]. JB* [CAP22]. JB*-algebras [CAP22]. JBW} [SH21]. JBW}-algebras [SH21]. Jensen [Anj17, BHK10]. Johnson [Ano79b, Moh15, Hu94, Mol18, Rub89]. join [MH12]. joins [CC22b]. Joint [BFM20, GR93c, GBM21, MS15b, BM22c, Cha18, DM14, JX13, KV23, Pep19a]. jointly [ST22]. Jones [Byt20]. Jordan [MC17, Sło18a, AAB15, AEV12, AH10, An18b, AGAG96, ABM22, AM21, AAUN22, BB21, Ben05, Ben16, BDvW16, BDD23, Bre16, BG23b, BN88, BC94, BFK18, BM16, BMGMC15, CCU21, CM19, CL17b, CMDE19, DPvW13, DW12b, Elg22, FKKS11, FJ09, GT11, GJ22, Gow22, Gra11, HKP23, HL09, HS91, HLH18, HZXL15, IT16, Kay17, KPS19, Khr16, KPPR21, KLJ23, LWZ15, Lee16, LZ11a, LB11, LPS13, LF13, Lin21, MJ07, Mol15, NSW21, Nor95, Nor08, Pas15, PT10, PS91, QZ16, Roi99, ST15, SX22, Slo17, TRD16, TS21, TWK22, TJG22, WHX08, Wan18, Yua98, Zha16a, ZW14]. Jordan* [LLF14]. Jordan-like [AM21]. Jordan-multiplicative [AGAG96].

K. [Mer79b]. Kac [CW23, CMDE19, Mar00, dFK21]. Kakutani [Esh17]. Kantor [EKO14]. Kantorovich [MP94, MP95, MGH19, Zha01b]. Karcher [FS15, FS19]. Kato [ZZC17]. Kazhdan [RS10]. Kellogg [Ano79b]. Kemeny [CDK22]. Kernel [BR04, GGGA21, GY22, LM85, LCZW19, Pep19a, Pep19b, YGÇ19]. Kernels [LLWX17, CL17a, Hun86, Józ15, Kut19]. Khinchin [zY11]. Khintchine [NASR19]. kinds [LLWX17, MHM22, SLUG04]. King [Don91]. Kippenhahn [Law19]. Kirchhoff [CMM17, Ye11]. Kirillov [BCZ22, Cen11, ZCY22]. Kittaneh [WZ14]. Knopp [Car19]. knots [Bae84]. known [Krä87]. Ko [GX17b, Woo22]. Kobayashi [Hsu85]. Koethe [CKLP13]. Kolmogoroff [FP17]. König [Kov86, Aha76, HL84]. Kostant [HK10, Tam97]. Kostka [ER90]. Koteljanskii [JN14b]. Köthe [DV19]. Krahn [ZZZ12]. Krawtchouk [KHG15]. Krein [LLMT05]. Kronecker [Cal99, Atk91, Cal97, Cal03, CG17, CK20, De 78, Di 88, HZ11, HS81, HPS83, Lew96, Nor08, RW96, SW81, Tia05b, Wan80, vR11]. Krusemeyer [Ric87a, Ric85]. Krylov [EJS22, ZQZ21]. Kuratowski [HW88]. Kwong [NS22, ZW19]. Ky [Gro17, JLL88, MP94, MP95, Mos12].

L [Mag83, YW16]. L-structured [Mag83, YW16]. labelings [KT95]. labelling [LL22b]. Lagrange [Jon78]. Lagrangian [MMRR09, PT10]. Laguerre [ASAS79, Mej19]. Lambert [JS19]. Landau [BdP21]. Laplace [Bay22]. Laplacian [AAS14, AFTG19, AADdF23, ABGP21, AAB20, AM85, AH16a, Are22, Ash16, AEG19, AKB21, AOTR11, BKP01, Bar08, BLP08, BPW16, But10b, CRW21, CP15, CL17c, CHQ18, CZ21, CFLL23, Con90, Das04, DP17, DAH19, Das21, DK17, FTZ08, FLZ07, FYIS13, FHT14, GHL10, GM90, GZ90, Guo07, GLS16, HMO11, HLP03, HUZ17, HHW15, HL22, JS22, KP22, Kir04, KADM10, LP07, LST23, LP00, LLF08, LZ11b, LC14, LLP18, LLZ22, LPWX10, LL15, LZ15b, LL10b, LPH17, LDL17, LW22b, LY22d, LHH18, MH16, MT22b, Mer97, Mol18, NC22, NP09, Obo22, PLH01, PGLM00, PRAC22, RA21, Rod02, Rod03, Sar20, SYZL17, SHS17, So99, SD20, TJ11, Tan12, TYC21, Vrb86, WSH^{+}11, WH11, WZ13, WBZ14, WW15, WLW23, Wat90, XZL14, XYH22, XLS20, ZLS11, ZZ15, ZZZ23, ZG08, ZW17a]. Laplacian-like [Tan12]. Laplacians [CTY20, Fie95, Lee98, LS21a, Lor00, Mer95, PLLL20]. Larcombe [MM23]. Large [GZ90, Loe01, Mer97, NW02, CW20, Din15a, KMEO12, LLP18, LHH18, Pat03b, dSP16b, Sol14, WY20, Zen22]. largest [AAT17, Bel15, BPW16, Bén11, CR90, Das04, Das21, DdF15, Guo07, GZ22b, LP00, LXG17, SSZ22, Sim95, Sta10, SKLS20, SKLS21, Wal11, XZ16]. Last [KMS02]. latin [SMF95]. Lattice [Cir15, BV92, CZY12, HMK22, JS94, KMH19, MPGS21, RW84, YW10]. lattice-theoretic [KMH19]. lattices [GT14, Ger04, Ye11]. Laurent [JK10]. Lavoie [YLS10]. law [CID16, DD14, LW08, Mal92, PBM20, Rad16, XQ16b, ZCP17]. laws [BS18b, CIN15, CIM19, LZ23, LFY16, XQ15, ZX10, dC15, dC18]. layered [Mur89, Nie22a]. LB [HZH20]. LB-TD [HZH20]. LCD [LLY22]. LCM [HINS10, Tan10, TLL11, XL11, Zha14b]. LCP [LL22a]. LDU [Smi83]. leading [Hol84]. leaks [GOS22]. leaping [Kom10]. Least [WXD21, YL11, YLW15, YZL17, ZLZ^{+}22, AAB17, Ano16b, BPS16, DM19, DLQ18, Din98, Fan05, GW17, GZ22a, GZ22b, JS84, Joh84, KS82, LLW07, LW08, LW10, LZS23, Pen15, Wan15, Wan17b, WLZ22, XQZ13, XQ16b, YWZ21, ZWLZ22, Zhu15]. Least-squares [ZLZ^{+}22, AAB17, WLZ22, ZWLZ22]. leave [Ben78, Sin73]. leaving [LT93]. Lebesgue [BS22a, LRD21]. Lefschetz [MCIM21, Sam78]. Left [Bre21, CQH15, DK22, EKO14, HBM15, Kre16, MW22, Tho86a, DS16, – Dok73, DK21, GMZ17, GZW20, LCL14, LGB17, MVPS14, MVPSTT23, MRD15, WHC22]. Left-star [Kre16, MRD15]. left-symmetric [LGB17]. Legacy [Mer17]. Leibniz [Ano22b, ARH15, ACO21, AM15, AKOZ20, BCLS23, BCS14, CCGO11, COK19, CLOK13, CKR18, CR18, GLOT13, KPPV18, KLO14, KLO19, Mil15, RH11a, RH11b]. lemma [CGMR11, YZZ20, Bar07, Bar12]. length [BP90, CCGO11, CW81, Ell93, Wes80]. lengths [Kol23]. Leonard [Aln11, CGH15, Han21, HZG14, HZH20, KHG15, KC16, WHG17]. Leslie [Har96]. less [Bot87, Lim90a, Sim95, dB16]. Letter [Bha91, Cha93b, Ram02, Roh13, Sch15, dS19a]. letters [Min74c]. Level [SRT22, AHK92, HS89, HS91, Lew90]. levels [Vol22]. Levi [KLO19]. Levitzki [Ari16, DP18]. lexicographic [AAdFS19]. Li [FS15, Lim23]. lie [JN22, Tow22, ACO21, AS20a, AC19, AH20, AAA22, AWM22, AH16b, AA20, AK21, BL10, BWLL12, BD12, BW15, BGLZ18, Bas17, Bas19, BMN22, BV92, BBM15, BRL22, Ben07, Ben15, Ben22, BCS14, BCS18, BM14, BS18a, BB14, BTV07, Bre14, CR15, CGNT15, CCGO11, CPS22, CSG22, CV94, CKR21, CC19, CCC21, CW11b, CX12, CW12, CY22b, CQ22, CZ23, CWSZ17, Che03a, Chi22, Chu84a, CPT15, DH18, DS23a, DDA20, Dol07, DR18, DBG18, FD17, Far93b, FZ17, Ghi22, GS95, Gra10, HH22b, Hon22, HH16, Hou17, HX21, JLZ12, JQS13, JL12, Joh22, Jul21, JO23, KK75, KZN22, LK23, Lee94, LCL13, LW11a, LF13, LCC19, LWXF15, LL10a, LL11, LZ12, LZ13, Liu16a, Liu19, LW20, LC22, Liu22a, LG23, LBV13, Luk74, MC17, MPGS21, MT11, MHA17, Mou04, NZK21, Nir11]. Lie [NJP19, OWX09, PN22, PZGL22, QH11, Qi15, Rom89, SV14, STME10, SN22a, SX15, Shi18, Spi04, ST99, SYZ22, Sun21, ST97b, TKT19, Tan18c, UF22, WOY06, WOW09, WY11, WL16a, Wan16c, WL16b, WJ16, Wan16d, Wan20, XW12, XWF15, YD14, YZ13, YZ20, Yan21, YWO07, YH22, Zha15c, ZL15, vWZ19]. Lie* [Liu18a]. Lie*-automorphisms [Liu18a]. Lie-theoretic [Mou04]. Lie-type [MHA17]. Lieb [TWK22, Byt20, NY23, Pat96]. LieRep [ZLS22]. Liftings [Suc22]. lifts [DFPS23]. Ligh [KMH19]. like [AM21, BD14, KM17, LM19, Nie22c, Tan12, Zha97, ZZL23, ZAK22]. Limit [Ger86, Cam78, Har84b, Kat75, LYZW12, Rei83, Tos79]. Limiting [Sie99, LN14]. limits [Geh15]. Lindström [LN80]. Line [GW08, Cas23, CMM22, HJ77, HT91b, Mir23, QX23, WBZ14, ZB13]. Lineability [BMP10, BMFPBSS21]. linéaires [Del94]. Linear [AS17, ASZ02, Bea78, BP89, BP90, BP92, BL93a, BL01, BBBS09, Bot79, Bot87, CL83, CL92, CL95, CLN13, CLMR09, Coe96, DRY22, DGJ^{+}16, FHLS13, FHL^{+}14, FA22, FJZ21, Ghi22, HJ19a, Hof75, Hua96, HW75, HKS94, JS94, JS84, Jam81, JP86, JLL88, JLR^{+}92, KP92, LL90, LL16a, Lau81, Lev74, Li87c, LT88c, LT88b, LT90b, LT90a, LT90c, LT93, LM93, LP93, LMR94b, LŠS01, LZ01, LM03, LS04, Lim76a, Lim79a, Lim82, Lim90a, Lim90b, Lim97, Lim01, LRD21, Loe93a, Loe93b, Lor14, Luk74, MPI17, MF79, MC80, MŠ09, Min77, MSS19, Mol09, Prz92, RK94, Sin82a, Sło18a, Sło18b, SA14, SB03, Tom19, Wat82, Wes74, Whi74, Zha05, Zha22, ZXF15, ZW14, Abd09, AC13, Aln15, AFM13, Ami11, Amm18, Ano92g, AW83, AB22, BB06, BH23, BB20, Bar07, BS16, BS17c]. linear [BS22b, BNKJ22, BGdH92, BB19c, BT08b, BLZ10, BR10, Ben78, BCM18, Bou18, BZZ10, BMGMC16, BF96, BSS12, CKS19, Cam83, CW90, CW20, CL91, CC19, CS10, CKLP13, Che93b, CW12, CZZH22, CW23, CWSZ17, CSY18, CL05, CKL16, CK20, CL21, Chu84b, CPT15, CF15, CFP21, CLW21, DM17, DM20, DPR22, DPP17, DH86, Del94, Den14, DLQ18, Din00, DIJ23, DKY22, Doo16, Dra21a, Dra18a, Dre93, DN18, EL09, EÖ15, EBB18, Esh22a, FDLB20, FK91, FW73, For87, Foš05b, Fow80, FH15a, GLL18, GP06, GL85, GOS22, Gib89, GJ22, Gur98, HK16a, Had83, HHSY22, HR85, Hla17, HH16, Hua16, HWL21, Hua22c, Hun83, Hun86, ISCS11, IK18, INO15, IKR12, JDD21, Joh11, KP11, Kas85, KMS18, Knu85, Koi21, KT83, KKR12, Kov77, KV22, KK22c, Kun99, Leh76]. linear [LCL13, LT88a, LT92b, LT92c, LT92d, LM96, LLF14, LF16, LZ16, LD17, Li21a, Lia80, LLWX17, LZ19b, Lim10, Lim15, Lip96, LWB11a, MT22a, Mag22, Mag83, Maj22, MM06, Mar97, MR12, MHA17, MID17, Ma’15b, McD92a, Mil22, MN96, MŠ98, MB22a, MOGN18, MBB22, ME18a, Ned98, ND13, NT91, New80b, Nie07, Pan13, Pap13, Pat77, Pat83, PUÖ15, PW78, Pie79, Pop17, Pre22, Raï14a, Raï14b, RRC23, RJ04, Roh91a, SPMR20, SRT22, San18, SC21, SLH20, SXR18, ST86, SSZ93, Sin83, Sla16, Slo15a, SPKC15, Sti87, ST97b, TRD16, Tan18a, Tan13b, Tia00, Tia11a, TJ16, UÖÖ15, UPÖ16, Wal79a, WL16a, WS18, lWY19, Wat83, Wat85b, Wat89, Wei93, Wer84, Wri93, WL16c, WZY22, WL22, XLL15, Xio23, Xu15, YT15, YqS16, YW16, ZWZ23, ZVL18, ZV19, ZAK22, ZCW21, ZL17, dCdSA22, Ano79b, Ano16b]. Linear [Ano16c, TT78, MT84]. linear-algebraic [Bar07]. linearisations [Ves80]. linearity [GR14, Rom07, WZ22]. linearizations [Adh13, BD14, DA22, DMQV23]. Linearizing [Chi78]. Linearly [Mer78b, Rus16, FTA20, Koh11, Pat83, SRK22, YZH19, dCH21]. lines [SZ85]. Liouville [GMZ17]. Lipschitz [Oi20, RP04]. List [Ano12b]. listing [Mum86b]. lists [Wil73]. Literature [KS17b]. little [Krä87]. Littlewood [AW13, AM90, BG90, Wat86]. LMI [Hil07a, Hil11]. Load [Kir14]. Local [AKY21, AK21, Ben74, Bun94, CZZ22, CFP10, Gal79, HSM17, Liu16a, LT22b, SYZ22, Yao22, ZL17, AA20, AAUN22, BEH19, BDS13, BJS15, Ben18, BM15c, Cos14, DK14, EAM19, Epk89b, EPR16a, EPR16b, GSS09, GP22b, Jaa22, Liu16b, Liu18a, Liu19, MWLW22, ShDHG08, YF18, YW10]. localization [Ano22a, JL85, LR96, LLL15, LL16b, LZL16, MH20, PMP22, SL19a, Sen83]. localized [BZ22a]. Location [Fer10a, RW21, HHC12, Jin17, Joh75b, JLR^{+}92, LB19]. loci [Ven19a]. Loewner [HS12, NS18]. Loewy [Moh15]. log [GAMA23, KD16, TJG22]. log-convex [KD16]. log-majorization [GAMA23]. logarithm [CFSRL22]. logarithmic [DM13, OvdD77]. lollipop [LZ14a]. London [KM84, MV91]. look [MC19]. looking [DH91]. Loop [DH18]. loopless [SG09a]. loops [KMP^{+}19, MPI17]. Lorentz [Hil07a, Hil11]. Lorentz-positive [Hil07a, Hil11]. Lorentzian [NP91]. Lotka [ALP22]. Low [DO19, ARH15, Bal18, CFSGL18, Der18, DBG18, FP17, GLM^{+}22, KS17b, Laf78, RH11a, RH11b, WW20a, WW12]. low-dimensional [ARH15, RH11a, RH11b]. low-order [FP17]. low-rank [CFSGL18, Der18, WW20a, WW12]. Lower [Bén11, BR19, Cal03, CW10, DTS14, Hon99, JKB17, MC89, Mer81, PLH01, Pat87, Pat03a, Rho87, Rho90, Rob75, AFM13, CR15, CW11a, Kar83, KMN01, Li11, MF20a, Mar97, Mel13b, Neu83, Pat09, PS20, Ric78, ST94, SSG12, Tia17, Xio23, YSY17, ZBW16, ZWS20]. LU [Neu81]. LU-factorization [Neu81]. Lucas [Chi11, lGyW99, KA19]. Lusztig [RS10]. Lyapunov [AL78, BBP78, BV16a, HS85, LZT18, LZ19a, Loe73, Loe74, Orl18].

M [GT08, SA14, GT08, Wan77, XCC99]. M. [Cha83a, For79c, Hsu85, Krä87, Mal89]. M/M/c [GT08]. MacMahon [BX11]. made [KS19a]. magic [BTT20, Lor14]. Magma [LPS21]. Magnitude [Sin81a]. Main [Ter01, Ter06, AFVL00, Bur95, FS14, FS89]. maj [Gar79]. Majorization [Bap87, BJ93, KPP13, BNKJ22, EBB18, FMX18, GAMA23, GJ22, GM90, JLL11, JMP17, LW12, LRD21, LP22c, Mas10, MSS19, Nie07, Nie22b, Nie22c, SA14, yT00, TJG22, ZL17, vE22]. majorizations [Hua22b]. Malcev [BPSO12, KLO19]. Maligranda [Nie22c]. Man [Mer17]. Manasrah [WZ14]. manifold [MR75]. Manifolds [GH98, Muk20]. many [BS15b, Hua96, TLC16]. map [AF11a, AC13, Aln15, Car20, Chu84a, LM96, LP16, San14]. Mapping [ST08, Gol83, Gol90]. Mappings [Pet10, TH18, BSO13, CL01, CG22, Eba92, FU20, Gha12, HLH18, HZXL15, JL12, Kuz05, LF13, Lim82, Lim90a, Loe93a, Loe93b, MW74a, McD92a, PR12, Pet97, Tag12, Vel21, Wan17a, Wit11, ZL17]. Maps [AN19, BDS13, BE22, BM15c, CGY13, CL17b, Che03b, DŠ04, Dol04, Dol07, EAM19, EM15, FKKS11, GM18, HH22a, LJ11, SH21, Slo14b, TH14, WJ16, XXF22, Zha04b, ZHH09, ZWW10, AA21a, AEV12, AS17, AG15, AC21, AKS18, BHD10, BHD11, BEH19, BBBS09, BH13, Bot87, Bre12b, CKS19, COK22, Car19, CL80, CL85, Cat21, CCL21, CM19, CC19, CW11b, CW12, Che14a, CW16, CY22b, CW23, CWSZ17, CSY18, Cho10, CKL16, CK19, CK20, CKT20, CW22, CL21, Cos14, Cos21, DM17, DM20, DS13, DHB07, DW12b, FDLB20, Foš05b, FHL^{+}14, Fra13, Fra17, FH15a, GP06, GP22b, GJ22, HK16a, HR85, Heo16, HJ19a, HJ19b, HJ20b, Hil07a, Hil07b, Hil11, HH16, Hua10, JS94, JQ15, JP86, JLR^{+}92, Jul21, KMS18, LP11, LF16, Liu15, Mar13, MA75, MR12, MŠ09]. maps [MO13, Mol09, MN12, MOGN18, Nag17, Nie07, OP13, Oge94, Oi20, OK07, Ore19, dSP16b, PR20, PW78, Pie79, Poo17, QH11, QH17, QC23, Rai87, RJ04, RŠ06, RŠ08a, Ruc89, ST22, Shc21, Slo13a, Slo16a, SDL10, SL17, TKT19, TS21, TB92, TXC08, WZC11, WZL13, Wat83, XCC99, XPY16, XL17, XZ18, YZ20, YH05, ZCZ12, ZX16, Zha16a, ZXF15, ZW14]. Marcus [And80a, AYT83, BD96, Cha89, Cha83a, For79c, GLS13, Hu87, Hu89, KL17, Krä87, Mal89, Mer17, ST80, Tam86, Tam91, Wan77]. marginal [Joh82a]. Markham [Kua18]. Markov [AS11b, CFSRL22, Gue17, Gue19, Gur98, HWX22, Kir14, ME18a, ME18b, ME21, Pan22, Sie99]. Martices [MNN^{+}98, Wer87]. Martin [Koo95]. Marvin [Mer17]. mass [BS15b, LLLY23, XY22]. matchability [AKK^{+}23]. Matching [LS21a, CHL18, CH22a, CFLL23, FLZ07, KS22a, LC14, LY22b, LLGX22, LY22d, LW23, MF20a, Mum86c, Rau18, SYZL17, SKLS20, SKLS21, TW19, WL20b, WL18b, WL23, XZS19, YLW22, YS23, ZXL22, ZZ15]. Matchings [KOSS23, AFTG19, FGHL22, GZ22a]. Mathias [FS15, Lim23]. Mathieu [dB16]. matric [PM78]. matrical [AK94]. matrice [Sło18b]. Matrices [AY74, BKL02, BL93b, BS04, BT98, Bru79a, Bru82, BF75, CLX03, CS21b, CA03, CD03, Cra09, De 82a, Dom95, EMS02, Gar21, Gib92, GN02, Gra87, GK77, HS83a, JR93a, JKZ20, Joh81a, JL90, JL93, JT95, JT04, JP07, Kap77, KW21, KK22a, KMS02, Lev74, LP02, MS74, Min74d, Nab98, New80b, Peñ02, RR04, Rao81, Ric81, Ric85, Ric87a, Rob77a, ST73, Sin76b, SHC02, SJ02, SK03, SB03, Stu89, TZS19, Tau74, TDL18, Zab89, Zab90, Zha04b, dO75, AG22, AA21a, ACI15, AKR23, Abo20, Ach77, AS95, AG16, AAK16, AR22, AAB17, AAT16, AGS20, ADJAH22, ANK22, ARK22, AAJ12a, AJL14, AJL18, AAJ12b, AS20b, Ala20, AAKS21, AW07, AAB20, Alp95, AST05, Ami80, AMZ13, AEdF13a, AEdF13b, AdF21, AO78, AB78, And80b]. matrices [AD17, Ang00, Ano22a, AB19, APEPPS23, AG96, AM21, Arm17, Ati20, AKB21, Aue95, Aue98, ALP94, AA20, Aze99, BS15a, BHC16, BLX18, BJS98, BB06, BT08a, BT10, BT22a, BST23, BBG22b, BBG22a, BBG22c, BG86, Bal77, Bal79c, BL83, BDK09, BS91, BPS00, BCN05, BK17, BZ90, BZ99, Bar74, Bar80, BG83, BOvdD91, BH90, BT90, Bas19, BHK78, Bea78, Bea85, BL93a, BL97, BL01, BS16, BS17b, BS17c, BPD88, BLD91, BL92, BP98b, BKS21, BDS13, Ben74, BH74, BB75, BT06, BLZ10, BL13, BSÜÖ13, BBM15, Ben05, Ben07, Ben78, BGV11, Ber17, Ber12, BB15b, BP74, BS97, BSM98, BX07, BDK^{+}05, Ber05, BZ89, dAB17, dAB21, BD85, BBP21, BP22, BDH23, BH15a, BS17e, BLSV12, Bo01, BR01, Bot94, Bot96, Bot97, BB96]. matrices [BB17c, Bou13a, BCCL22, BKMMG22, BZ22b, Boz98a, Boz98b, Boz99, BT15, BFI22, Bra86, BC17, BD13a, BCKP10, Bru85, BS08, BD13b, BM18c, BM19b, BD22b, Bru79b, BK87, BZZ10, BSZW13, BGZW16, BM16, Bur95, But10b, CC17a, CM92, Cal97, Cal99, CQ18, CKGP21, CP81, CCS16, CCU21, CGY13, CK17, Cao19, CCKL20, CMP22, CP97, CDM22, CFSGL18, CFSRL22, Cas23, CdM15, CL92, CCS98, Cha83b, Cha84a, Cha88, CGW14, Che92, CKN95, Che07, CW10, CW11a, CZY12, CP16, CW16, CCW20, Che22a, CZWF23, CTW15, CKMY22, CT96, CL14, Che17, CL05, CNP05, CNP06, CN08, CLNT17, CN21, CC23b, CKS20, CH22b, CLN13, CKL16, CK20, CKT20, CW22, CL01, CKP09, CJW19, Coe96, CR77, CZ22, Cos14, Cra91, CL74, CGH14, CPLH17, CIP18, Cve18, DFPS23, DGL22, DHHK23, DA22]. matrices [DPP17, De 90, De 75, Dec95, DK23, DTS14, DOP22, DD10, DCIW10, DRY22, DHS95, Din15a, DIJ23, Do 96, Dod13, – DL92, – Dok93, – Dok17, Dol07, DO20, Dom14, DC20, DC23, DW23, DMQV23, DGJ^{+}16, Dra22b, DJL94, Dru15, DdF15, DdF22b, DHH22, DH22, DN18, Düf97, ERS12, Elg22, ELG88, EL90, Elo18, Els78, ER81, EJD88, ES98a, ES98b, ES00, ES76a, EV77, EM73, ESH23, Esh22a, Esh22b, Est80, ETM22, FL16a, Fal97, FV97, FJM00, FX12, FT21, Fan73, Fan92, FL18, FW07, Far17, FJ15, FWZ11, Fen10, FdF09, Fer10b, FF21, FM22a, FS83, Fie90b, FP17, Fit94, Fla75, For79a, For79b, FS86, For88, For79d, Foš05a, FHL^{+}14, Fra13, Fra17, Fre76, FH89, FM78, Fri84, FLS88, Fri12, FL16b, FLT22, FSWX22, FG22, FS15, FSY16, FIS06, FJZ21]. matrices [FJ22, FS89, Gai77, GLL18, GHL20a, GPM98, GA18a, GG22a, Gar85, GWW15, Geh15, GJR97, GKS01, Ger86, GS21b, GS21a, GPM14, Ghi22, GV95, Gib76, GZ99, Gil14, Gil17, GKLR82, GZ76, GTZ75, Gol87, Gol92, GH98, GS20b, GR10, GR93c, Gro97, Gro99, GLP92, GP96, Gro17, GHT11, GHK16, GHL16, Gur79, Gur92, Gut22, Gut23, Gyi78, Ha88, HC14, Haj18, HK79, Hal91, HA22, HJP17, HZ11, Har87, Har78b, HS83b, HS83c, HMAD13, HINS10, Hau12, HDJT97, HWZ18a, HR85, Her83, HS88, HS93b, HNS99, HS12, Hil17, Hil18, HLS18, HKO16, Hol85, HH85, Hon99, HJ87, HZ12, HM76, HM77, HM78, HLZ18, Hsi78, HLZ05, HH14, Hua01, HW06, Hua08, Hua10, HLZ11, HZ20, HL20, HZX22, HK17, Hwa85, Hwa91, HK92, HS93c, HKS94, HM95, HKS95, Ibr00]. matrices [IS15, IS16, IAM23, Ito16, IOMAN21, JK22, JGK79, JGS90, JS96, JDD21, JR93b, JM13, JTW17, JK10, Joh73, Joh74a, Joh77a, Joh74b, Joh77b, Joh81b, Joh82b, JR84, JP86, JP87, JSE90, JKS94, JWP98, JS99, JR09, JN14b, JTU00, KM19, KN22, KS16, Kan87, KS12, KS15, KS14a, KP22, KS22c, KMS16, KLZ10, KCSĆ21, KP14a, KT21, Kim97, KS09, Kim13, Kim91, Kir91, KP92, KS93, Kir93, Kir04, Kir09, Kir11, KM11, KS19b, KS76, KS82, KKL18, KL19, Kob06, KNX93, KH05, KT19, KE82, Koz17, KS83, KS84, KM98, KS17b, KAE22, Kuz05, LR17, Laf78, Laf83, LL16a, sL19b, Lau81, Law19, Lee91, Lee18, Lei95a, Lei93c, Lew77, Lew97, Li86, Li87a, Li87c, LT88c, LT88d, LST89, LT90b, LT90a, LT93, LM93, Li95, LSOvdD95, LEH96, Li98, LŠS01, LPS08]. matrices [LLZ13, LL14a, LLL15, fLLyP15, LZ16, LLGL16, LMLL17, LH21a, LWC22, LP22a, LHZ22, LM95, LLL09, LHS^{+}21, lLpW16, Lia23, Lim79a, LT14, Lim15, Lim23, LT12b, LW12, LZ17a, LF21, LMMR01a, LMMR01b, Lit17, Liu92, LB99, LWY11, LWB11a, LWB11b, LLGX22, LHGL22, LCL22, LL78, Loe87, Loe01, LS21b, Lon80, Lon81, Lon83, LSZ86, LM89, Lon90, LR89, Lop11b, Lop18, Lov21, LSX19, LLLY23, LZ22, MM11, MVPS14, MVPSTT23, Mag83, MGKA23, MN80, Mah82, MM90, Mal91, Mal94, MRT16, Man92, MHH96, ML21b, MZ04, MS79, MKS84a, MS89, MR13, MR15, Mar20, MH19, MH20, MSZ04, MR07, Mas10, MH12, MP90, McC85, McD92b, McF79b, MS10b, Mee80, MS18, MMRR09, MMRR16, Meh84, MO93, Mel12, MV91, Mer94, Mer98b, Mes90, Mes13, Mia93a, Min74b, Min74a, Min75, Min77, Min82a, Min84, Mir92, Mir18]. matrices [MS12, Mis17, MS10c, MMG15, Moh16b, MTR19, Moh15, MP16b, MSS19, MŠ98, Mol09, Mol15, Moo76, MS21a, MAH18, MAH21, Mou22, MZ22, Mou04, Mou13, MAM15, MA15c, Mur89, NMAB22, ND05, ND13, Ner22, NS99, Neu79, NP80, Neu84, NST91, NS92, NS93, New74, New75, NT86, Nie14, Nie10, Nof15, Nom73, NP91, NS22, NTU93, OP13, Oml14, OK07, Ore19, Pab22, Pat92, PS20, PG22, dSP12, dSP16a, dSP17, PUÖ15, Pet09, Pet00, Phu80, Pif87, Piz87, Pok87, PT87, Pop14, PV20, Pul75, QLK22, QK92, RBG23, RRR93, RW22, Rau18, Rea96, RW96, RR95, Rho87, Rho90, RS74, Ric78, Ric87b, RQ22, RD12, Rob75, Rod97, Roh91b, Roo20, RW21, RSM22, Rub20, Rys73, ST94, SSGP23, SC21, SS79, SU77, Šem97, SX23, Sen83, SM94]. matrices [SMBB^{+}15, Sha91, SG96, Shi15b, Sil88a, Sil88c, Sil88b, Sil90, Sil93, Sil95, Sil98, Sil99, SP87, Sin90, SAV01, Sin73, Sin76a, Sin77, Sin80, Sin82a, Sin82b, Sin84a, Sin93, SST22, Ska17, Slo13a, Slo14b, Slo15b, Slo15a, Slo16a, Slo16b, Smi05, Smi76, Smi94, SH98, ST91, STZ94, Sol06, Sol14, Sou83, Sou03, Sou86, Sou92, Sta15a, Sta23, Sul08, SH92, SWZB15, SK15a, SK15b, SvW14, TD14, Tal17, Tam86, Tam94, TT12, TZ19a, Tan10, TLL11, Tan13a, Tan14, TLC16, Tan18b, Tan20a, Tan20b, TG09, Tau81, Tho74, Tho75a, Tho75b, Tho86a, Tho86c, Tho87a, Tho88, Tia01, TS05, Tia05a, Tia05b, TT06, Tia11a, Tia11c, TL12, TJ16, Tia18, Tia21, Til77, Til98, Tis93, Ton87, Tra89, Tra90, Tre85, Tre86, Tre87, TGW14, TMOvdD93, TZ21, UÖÖ15, UPÖ16, Uhl79, Uhl14, UG21]. matrices [Vas87, VM07, Wad90, Wal73, Wal75b, Wal86, Wan76, Wan79, Wan80, WZ95, WOY06, WZC11, WYM11, WZL13, lWpW14, lWW15, Wan16a, WC16, Wan17a, WS18, WL18a, WLL19, WS23, Wan05, Wat84b, Wat85a, Wat85b, Wat77, Wat82, Web83, Wes87, Wim85, Wim88, Woe93, WM84, Wör20, Wri22a, Wri22b, Wu09, WHC22, XW22a, XW22b, XCC99, XZ22, XQZ13, Xu93, XXZ11, XL11, Xu15, XPY16, XL17, XZ18, XHL20, XCM20, XY22, YLS10, YLC19a, YLC19b, You80, YBCW96, Yu96, YW16, YZH19, ZZ16, Zel85, Zel75, Zen22, Zha00, Zha01a, Zha89, ZS05, Zha05, Zha14a, ZZZ16, ZFSX17, ZMG20, Zha20a, ZWW10, ZHZ11, Zha14b, Zha22, ZXF15, ZCLW22, ZJC^{+}22, Zho16, ZBW12, ZW14, ZCW21, Zho22, Zim89, Zou12, ZIC22, cc85, dF13, dFSY22, dCH21, dPN81, vBW95, vR11, KNX95]. matrices-nonsingularity [ES98b]. Matricial [Far93a, MRS10, CKL^{+}21, CH12]. matrics [– DDJ76, HK80, NU94]. Matrix [AMP02, AP04, Afr73, AP74, And04, jAS15, AW13, AM90, BD94, Bom02, BC94, Che93a, Dat80, Der16, DS02a, DŠ04, EMS02, Fan92, Fla74a, Foa79, FS18b, Gal03, Gil12, GMR81, Har95, JT03, KM22, Leh91, LT92d, LM96, LP03, Mir81, Net22, OW02, Pat03a, PT02, PY82, Rum97, Sin81a, Tho78, Tia02, Tia03, Tur02, Wei98, Zha04b, Zha04a, AEAAES20, Abd20, Ab23, AFVL00, Adh13, AS11a, AM13, AN14, AAD22, ABAH23, Ake83, AAB15, ABGP21, ABP22, AdFK15, AG15, Ano79b, Ano79c, jAS21, AELL20, AKB21, AY95, Aud15, BY19, BS21, BHY22, BK83, Bak84, BB06, BCR91, BJ75, BY76, Bap97, BGK99, BGK00, BKP01, BS11, BR14, BWF83, Bas17, BC82, Bea83a, Bea83b, Bea84, BC85, BD88, BS17d, BNKJ22, BGdH92, Bel83, Ben90, Bén11, Ben18]. matrix [BT08b, BDV19, BT07, Bot94, Bou13b, Bou14a, Bou14b, BHK10, Boz97, Bra86, Bre21, BGGS17, BC91, BF15, BS18d, BZZ10, Bun91, BSS12, Caf13, Cal95a, CRW21, Cam78, CKP16, CL80, CJ97, Cav10, CS23, Che92, CH12, CLW15, CS18, CD21, CC21, CQ22, Che91, CCL10, CT09, CNP05, CN07, CN14, CM85, Cho10, CR22b, CKKY23, Coh88, Con90, CC98, CC18, CICX09, CIP17, CIRW21, Dag23a, Dag23b, DM22a, DM22b, DPvW13, DH86, DW92, Dec74, DTSI15, DH11, DM22c, Der18, Deu80, DHT94, DDF22a, DS09, DS22a, DS22b, – Dok73, – Dok90, – Dok09, DB14b, Dom00, DLL19, Dre22, DJ89b, DWL13, DWL14, DS23c, DJS21, DS18, DS19b, EM95, EF98, ER90, EOvdD09a, EGT16, EOvdD09b, EB17, ESH23, FZ11, FTZ08, FHW16, FNW18, Fei79, Fie81, FLR85, FSvW11]. matrix [FdM18, FKKS11, Fri97, FH15b, FSY16, GAGM18, GA20, GHL10, Gil93, GKLR82, GLR82, GRSW92, Gol89, GG22b, Gon17, GB90, Gow80, Gra80, GHV99, Gre99, GPK94, Gro96, GJMW86, GTM14, GT79, Gus81, GHT16, Ha89, HK16a, Haj18, HKP23, HL09, HZ11, HLZ22, Har81a, Har94, Har96, HP90a, HP90b, Has18, HKT15, Hav88, HBI73, HW13, HW14, HW15, He21, HZV23, HS81, HLT22, HS88, HS89, HS91, HM23, Hia94, HS87, HMN91, HBP18, HS16, Hou21, How83, HT10, HXL16, HLPW17, HM18, HWSH13, Hul81, HPV84, HPV88, HLP23, Ikr15a, Ilk20, IH11, Iov15, INO15, JS22, JP15a, JT23, JKD16, dJfLlZ^{+}21, JZ22, Joh77a, JKS79, JOvdD84, JOvdD85, JD99, JS08, JST18, KMRS89, KSU22, KS16, KM85, Kar83, Kat75, KM17, KFR17, KFR20, KL12, KA19, KOvDD09]. matrix [KS07, KLPK17, Kol23, KMMS22, KPPR21, Kov77, Kra05, KMS07, KML95, KLZ96, KK16, Kur04, Kyr08, Kyr21, Laf80, Laf84, LRR87, Lás91, LZ09, Lee15, LM22, Lei93a, LRT91, LR96, Lew96, Li84, LT88a, LT91, LP93, LP00, LLS10, LZ11b, fLLyP15, fLLL^{+}17, LC20, LLC21, Li21b, LCH22, LLC22, LPWX10, LWXF15, LW11b, LC18b, LQ95, Lip96, Lip98, Lip87, LW08, LT10, LTT11, LL22a, LLY22, Liu22a, Liu22b, Liv94, LMR00, Lop11a, MJ07, MS10a, MDW11, MW16, MT22a, Mag22, Mag83, Mai11, MS83, MSO83, MRT14, MKS84b, Mar89, MR97, MP98, MS95, Mat90, Mat92, MP87b, McD92a, MS81, MKS22, Mel13b, Mel18, Mel19, Mel20, MK16, MNS22, Mer74, Mer79a, Mer82, Mer83a, Mer89, Miy22, Mol18, MP94, MP95, Moo95, MM23, NNT95, NTW23, Ned98, Neu83, NP80, Neu81, NP73]. matrix [New76b, New78, Nof15, NT94, Obl12, Ogu20, Oi20, OvdD77, Pal79, Pat82, Pat03b, Paz84, dSP11, Pea84, Pen15, PLPC19, PM23, PW78, PGAB22, PM81, PM85, Pre85, PV00, PH97, QD85, RS07, RS18, RR91, RS10, RP04, Roh06, RF19b, RF19a, Roo20, Rou98, Rys75, SRK22, Sal79, Sal12, Sal14, SA07, SSZ22, SQM99, Sar20, SMM14, Sch77, SM20, SM22a, ST80, SLUG04, Shi15a, ySpW17, Siv06a, Siv06b, SST22, Slo14a, Slo17, Sło18a, ST96, ST00, SBMP16, SC22, Sta15a, Ste94a, Ste94b, SX12, TD16, TD19, TZ06, TZ17, Tau73a, Tau81, TB01, Tho87b, Tho89, Tho92b, Tia00, TC04, TS05, Tia10, Tia11b, Tia17, Tik98, TT18, Tos79, Tra84, TSYB82, TRM22, Uhl13, Vas86, VY21, Vla22, Vla23, VM18, Vrb86, Wal11, Wal75a, WZ97, WHX08, WH12, WZH13a]. matrix [WZH13b, Wan19b, WLB19, Wan19c, WW20a, WMFW21, WC22b, Wat83, Wat90, Wea83, WW12, XSH18, XWT22, Xu22, zY11, Yin21, YSY17, YqS16, YL11, YLW15, YW16, YZL17, YTL20, YCZ21, YZL22, bPFZH08, Zha04c, ZCZ12, Zha15a, Zha16b, ZL21, ZWLZ22, ZLZ^{+}22, ZW19, ZX10, ZV19, ZAK22, ZWZ10, ZW14, ZWM17a, ZWM17b, dFK21, dB14a, dOMD77, dBS00, vWZ19]. Matrix-analog [Foa79]. matrix-product [LLY22]. matrix-subadditive [Tik98]. matrix-tree [Mer89]. Matrix-valued [Tur02, DS23c, Gil93, Oi20, Tia17]. matrix-weighted [SMM14]. Matrizen [SU77]. matroid [Dia90a, Kas85, Mum86a, Mum86b, Mum86c]. matroids [BVW84, Mum86a, Mum86c]. Maurey [Pop20a, Pop20b, Pop22]. Max [SK03, BPS00, BSS12, ISCS11, KS15, LL14b, Mer94, Moh15, Shi16, Tia18, lWpW14, lWY19, ZBW16]. max-algebra [BSS12]. max-cuts [Mer94]. max-min [Moh15, Shi16]. max-plus [lWpW14]. max-relative [ZBW16]. Maxima [YGW15]. Maximal [Bal79b, INO15, Mos20, Yiz03, ACO21, BM22a, BS21, BBK22, CH22a, GF10, HKS94, HKS95, JLL07, KMN01, KS17a, Oml14, Tow22, UZ16, WL16a, Wes80, YS23]. Maximality [Fri85]. maximally [Ano16c, Ous15b]. maximization [DJ89a]. Maximizing [CL14, FKP03, FTZ08, GZ22a, LP07, Pat07, WH11, GJMW86, Hu94, Sha94]. Maximum [BdP21, FV97, BI95, BP90, BS22b, CCGO11, Cha88, DLSS18, DJ89b, FTA20, GWW20, JD99, JLS09, LH23, MM11, Mac95, Mal86, Mer80, Sin83, zSqSZ16, SL21, Wan79, Wan22b, XW19, Zha05, ZL19, ZZZ23]. maximum-nullity [BS22b]. Mazur [CAP20, JVMCPR19]. Mean [MS81, AG22, ASR22, Ala16, BP12, CW10, CKS20, Dra18a, EJD88, FP17, FS15, FS19, FS18b, Lim23, LS17, Pep19a, Pep19b, SY23, Wim85, vBW95]. mean-type [SY23]. Means [ZJC^{+}22, CPKK22, DDF22a, FH15b, FSY16, FMS19, GHT11, HSW19, JKL22, KMS18, KMS21, KM22, KK16, PSMA16, Raï14a, Raï14b, Raï19, ST94, Tom22, Tra84, Web83, XZ22]. Measurable [Sim83, AR22, BO22a, Han17, Sha20, SY23]. measurement [CP16]. measures [BPR14, JWBB22, Joh11, KM22]. meet [Ilm16, KH05, MH12, WHL16]. meet-semilattices [WHL16]. meets [CC22b, IH11]. Mehta [DJ89a]. memory [Uhl99]. Menus [HW88]. merge [CW91]. Merris [Pat93]. metabelian [AM15, Mil15]. method [BLX18, CW20, CL22b, Chu90, CLW21, DK23, Gib89, Haj18, KN81, LZ16, Li21a, LZ19b, LC18b, LL22b, LLLY23, Neu84, PLPC19, Sad23, SN22b, SAK22, WLF22, WL16c, WZY22, WL22, Xio23, XY22, YZH19, ZL21, ZWLZ22, ZZL23, ZWZ23]. Methods [SX22, ALP94, BNKJ22, Cro18, EJS22, HK21, HZV23, Hua22c, KM17, LW17, LW10, Sun20, ZV19, ZQZ21]. Metric [AEV12, Rod02, BdPN14, CAP22, OWTZ21, SC20, Tho78]. middle [CJY22]. mimetic [CG17]. min [ISCS11, LL14b, Moh15, NSW21, Shi16, Tia18]. min-max [ISCS11, LL14b]. min-plus [NSW21]. Minc [Cha83a, Els78, For79c, Jia92, Krä87, Mal89, Mar03, Pul75, Sol00, Wan77, Wan07]. Minimal [Cal03, Cha79a, CCS98, CD21, DB14b, EV77, Esh22b, Fie81, Hac22, HMAD13, Lop18, MZ17, NS09, Nat13, SG06, WL16b, Woe93, AGP22, BX07, BM14, BZ89, BJZ13, BD14, Cal94, Cal97, Cal99, CKP16, Cao19, Cha77, Dia90b, – Dok73, DMdO92, GV95, GMD85, GW21, HJ20a, Hil18, JS16, Kuz05, LC14, Lip98, Loe73, Lop11a, MR19, OvdD77, dSPdS22, RY93, TL13, Tsi83, Wer94, WW80, ZG11, Zha13]. Minimality [HKv23]. minimally [SSG12]. Minimax [GL85, Koz17, CM84, Fie90b, Zim98]. minimax-condition [Zim98]. Minimization [HS93c, SPM14, FW13, fLLyP15, Pap13]. Minimizing [FLZ07, KS19a, FS86]. Minimum [BS22b, Bur95, Cha84a, Cha88, Hwa85, JW91, KS82, LSOvdD95, Min84, SHC02, SJ02, Tia02, BHH^{+}11, DLSS18, DIJ23, DH91, FTA20, GR23, GH22, GZY18, Hou01, LLZ13, LGA^{+}13, LLP18, LT13, LLF18, MS16a, NS15, Wan22b, WY20, YSZW22, Zha05, Zho16, dF13]. Minimum-rank [BS22b]. Minkowski [Mer78a, Mia93a]. Minor [IW95, CZ21, Wil87]. Minors [Sar20, AAB20, BI95, BCCL22, BF75, For79b, Hol84, HO17, INO15, JOvdD84, Mia93b, RS10, San14, Sin77, TDL18, TZ21, dCdSA22]. Minus [BMGMC16, KMS16, CGM19, NS92, NS93]. Miranda [Li84]. Mirman [GS77]. miscellaneous [BLP92]. missing [CKS20]. Mixed [Kut19, jAS15, BGK99, CHL18, Fan05, GPM14, GV87, JN14b, LZX19, LY22c, LFY16, LZS23, MMG15, Mur89, Sar20, WL20b, WLM21, XQ15, YZ20, ZD18]. mixed-type [LFY16, XQ15]. mixing [AS11b]. Mn [Leg06]. mod [BM18c, BM19b, HH14, Sma17]. model [JL22, MT22a, Sal79, SW22]. models [GOS22, GLM^{+}22, TJ16]. modification [UB22]. Modified [BH23, Xio23, Zha20a, BS17d, CICX09, ER81, Gei78, PZGL22, SAK22, ZZL23]. Modular [GNO20, Ano22c, CLL14, CN15, KDHJ22, Reg93]. Module [BWLL12, FXY22, HJ19a, How86, JKH20, Lee94, Wis12]. Modules [AELPM18, Gar75b, MID17, AC78, AR15, AB17, AO22, BCNG23, Ber86, Bre18, aCC17, CL22a, CY22a, CHS22, Dix79, DKM10, FMX18, FMXZ22, GPW22, GP22a, GS20a, GNO20, GR93a, GMZ11, HJ20a, HJ19b, HJ20b, Hua22a, KW92, KW93, LGG19, MS08, Mal95, MR10, Pat95, Per19, PA94, Raj09, Rod22, Son14, Tan13b, TOM88, WZ22, WZ23, Xia20, ZY08, dB10]. moduli [dTS20]. modulo [GP06, MP13a, Sri20]. Modulus [HZV23, HNS99, WZY22, LZ16, ZV19]. Modulus-based [HZV23, WZY22, LZ16, ZV19]. Moment [KV23, ACJ17, BS15b]. moments [CBZ22]. monic [Coh88]. mono [KY17]. mono-weakly [KY17]. monogamy [Aud15]. monogenic [CF15]. monoid [Put82a, Put82b]. monomial [Fri85, Yuz84]. Monotone [Bar76, BGV11, Gha21, Gue17, HOT15, Mat90, NY23, PSMA16, ZZT23]. monotonic [Gol89, Mar92, SH98]. Monotonicity [AAJ12b, FM78, Gol92, BJ93, BS17e, ES76a, JGS90, LRR87, Mal91, MF20b, Pat96, Sin80, VM07]. monotonicty [Wer87]. Montgomery [DF82]. Moody [CW23, Mar00]. Moon [LN16]. Moore [Rom22b, BT22a, BST23, Bap97, Boa19, Bom02, BGZW16, CC22a, CHZ22, CFGM06, DK23, DD10, DW12a, DLW18, Fie95, FSWX22, Gro99, HPV84, HPV88, JGK79, KMRS89, KR07, Kyr11, Li11, MZ15, Men22, PBM20, PM22, PA10, Rom21, SC20, Siv06b, SZBW16, SPS15, TC04, Tia05a, YZ14, Zha04b, ZFSX17]. Moore-type [CFGM06]. Morita [BDvW12]. morphism [LCZW19]. Morphisms [Har84a]. Morrison [SC20]. Moser [LN16]. most [AVVÁ18, BTT20, BLP12, HLW^{+}21, LZ15b, TWZ16, YHS13, tHK17]. moving [GOS22]. Moyls [ST80]. mpletely [SMBB^{+}15]. MPW [Bea83a, Bea83b, Bea84, Mer82]. MSMAOR [WL16c]. MSMMAOR [Li21a]. much [BCT23]. Mühlbach [dC13]. Mullin [FW73, Gar73]. Multi [LLL09, ZZT23, BNKJ22, DM22a, DS23c, SX23, WZ22]. Multi- [LLL09, WZ22]. multi-affine [SX23]. multi-block [DM22a]. multi-linear [BNKJ22]. multi-resolution [DS23c]. Multi-step [ZZT23]. multiadditive [Eba92, Fra13, XZ18]. multianticommutators [MR08]. Multicommutators [MR08]. multicone [HLB22]. multicone- [HLB22]. multidimensional [CLL14, Kut19, Min80a]. Multifold [Chu22]. multiform [Poo94]. multigraphs [BM89]. multilevel [dBS00]. Multilinear [Ano79b, Ano16b, Ano16c, HFS15, HF18, LLMN20, MT84, Pop22, Sło18a, Sło18b, TT78, Vel21, Whi74, Whi75, Whi79, ADRP14, AADD18, AP17, BD22a, BT17, Cam14, Car20, CG22, De 82b, DV19, Erd22, FU20, GP22b, Guo20, HKv23, Heo16, HJ19a, HJ19b, HJ20b, Hor04, How80, Lim00, LT22a, Oge94, Pam87, Pat83, Pat17, PR12, PSS15, Pie74, RSPT17, SDL10, WLB19, Wan19c, WW01, Wit11]. Multilinearity [Fon92, DF87, Fon89, LC83]. multipartite [JKY23, KS93]. Multiple [HH14, AAPR19, Cam14, HZXL15, JS08, JSW09, LW08, NASR19, Pop20b, Rad16, RSPT17, Yin21, ZX10]. multiples [GKLR82, Tho86a]. Multiplication [Bre12b, De 75, MO94, Wat85a]. Multiplications [FDLB20, Bun94, Tur16]. Multiplicative [BD12, BC82, CQ22, Che03b, Ger74, LR04, QZ16, AGAG96, CZ23, DLW18, Est80, Hil17, MID17, MZ15, Poo17, QH11, YZ14, ZFSX17]. Multiplicatively [AA21a]. Multiplicativity [New94, Pro95]. multiplicities [Da 05, GMV18, Uhl99]. multiplicity [DIJ23, JD99, JLS09, JLDS12, JST18, LHH18, TJ21, WWWT22, XW22a, XLS20, Yu14]. multiplier [AS20a, PN22]. multipliers [GNO20, JK10, Liv94, Mej19, MAT22, SA17, Xia20]. multipolynomial [Vel21]. multisplitting [CLW21, ZV19]. multivalued [AFM13, Amm18, Gar23]. multivariable [MMR07]. Multivariate [Bah22, Par78, Chu22, HHSY22, HSW19, LLC22, SK15b, ZL17]. multiwavelet [CJ12]. multlinear [Mer92]. mutual [CP16]. mutually [LHS^{+}21].

N [BMD87]. Nagata [Koh11]. Nagumo [FP17]. Naimark [HT10]. Natural [GHL16]. naturally [ACO21, CCGO11]. near [BGLSSS20, DH20, HCS19]. near-vector [DH20, HCS19]. Nearest [Rum97, SRK22]. Nearly [BB15b, Kim13, GSS09, Har87, Sha91]. necessarily [LD19]. Necessary [Gue17, HB83, PH03, SQ15, AW07, Wan82]. negative [Ano79b, BGGS17, CCU21, ELG88, GA20, GR14, GP96, JKS79, Józ15, Kan94, KS17a, Koz17, Pul75, Rau18, RS74, Rou98, Shi15a, Smi74]. negativity [Aud15, HS12, NS22]. neighborhood [CSCH01, Sin81b]. neighbourhood [LZ14b]. neighbourhoods [GHV99]. Nekrasov [GLL18, LLGL16]. NEPS [Ste05, ZLZ20]. NEPSes [FKS22]. nest [BHD10, BHD11, JQ15, KZN22, LF13, Liu16b, Liu19]. nested [KRR18, MRT12]. network [CW91, FHZ17, Li85]. networks [CLX22, CMM17, CMM20, GY22, Uhl20]. Neuberger [– Dok75, Pat08]. Neumann [AC21, BD12, BH13, CC23b, CAP20, Drn13, FA19, HZXL15, JQS13, LLF14, LJ11, Liu15, Liu16b, Liu16a, NZK21, Pas15, San18, TRD16, WJ16, YF18, YZ20, Zha16a]. Neural [Uhl20, GY22]. neutrality [JSE90]. Nevanlinna [CH12]. Newcomb [Raw81]. Newman [Lei83]. Newton [KK16, PLPC19, SvW14, ZL21]. Nichols [AG18]. Nijenhuis [ZLS22]. Nikodým [GP22b, HJ20b]. Nill [GV89]. Nill-polynomials [GV89]. nilpotence [LHbL^{+}15, MO91]. nilpotency [Xu15, dB14a]. Nilpotent [GLdS22, OR09, Rom20, Sou92, Ano22b, BL10, BV92, Bot94, BC17, CR15, CKKK22b, CN14, Chi22, Fur08, GW12, Joh22, JN22, KPPV18, KK22b, LCL14, Mar18, MO93, Nir11, Obl12, OWX09, PN22, Ric78, Rom89, SN22a, Smi76, Sul08, ST97b, Tow22, Wan79, WL20a, YD14, YBCW96, Yu96]. nilpotents [CCL21, CG19, Rob77a]. nilradical [ACO21, CLOK13, KLO14]. nilradicals [ACO21]. nilspaces [Oml14]. nivellateur [HP18]. nnd [YZL22]. No [ZCY22, DKK^{+}20, LPS21, TT78]. node [lWpW14]. Noetherian [PRA22]. noise [DMV18]. noise-free [DMV18]. Non [Ait19, DM20, Das17, Foš05b, HH16, KS09, LLF14, Mas10, PSW18, TRD16, ACO21, Ano79b, BL10, Bal18, BBM20, Bre15, CCU21, CKR18, CKR21, Cho10, DH17, DP09, DM18, DZ22, GA20, Gha12, GP96, GLS16, GM21, Hou09, Hua22c, JKS79, KS22a, KS23, Kan94, KK22a, KADM10, Koz17, LPS09, Lim10, MM11, MF77, MR12, Mar12, ME18a, ME18b, NY21, Nir11, Pal79, PG22, Pul75, Rau18, RS74, Rom22a, Rou98, STME10, SR17, Smi74, Spi00, SDL10, Vei82b, WYM11, WZY22, WL22, dLMO17]. non-abelian [CKR18, CKR21, Nir11, STME10]. non-absolute [DH17]. non-alternating [WZY22]. non-autonomous [MR12]. non-bipartite [GLS16, KS22a, KS23, dLMO17]. Non-commutative [Mas10, DM18, KK22a, Mar12]. non-convergence [BBM20]. Non-degenerate [Ait19]. non-diagonalizable [PG22]. non-distinct [Vei82b]. non-empty [LPS09]. non-equivalent [Hou09]. non-Hermitian [Hua22c, WL22]. Non-homogeneous [PSW18, ME18b, Rom22a]. non-increasing [Cho10]. non-integer [GA20]. non-Lie [ACO21]. Non-linear [DM20, Foš05b, HH16, LLF14, TRD16, ME18a]. non-negative [Ano79b, CCU21, GA20, GP96, JKS79, Kan94, Koz17, Pul75, Rau18, RS74, Rou98, Smi74]. non-nilpotent [BL10]. non-perfect [DP09]. non-powerful [MM11, WYM11]. non-primitive [GM21]. non-round [Spi00]. non-scalar [Bre15]. Non-singular [KS09, SDL10]. non-threshold [KADM10]. non-trivial [Gha12]. non-uniform [SR17]. non-uniqueness [Bal18]. non-unital [NY21]. non-vanishing [MF77]. Non-zero [Das17, DZ22, Lim10, Pal79]. nonabelian [CFL22a, Mah82]. nonassociative [Lud19]. nonbinary [KP92]. noncanonical [BG86]. Nonchordal [GLP92]. Noncircular [GW12]. Noncommutative [Yu96, Gut01, KW21, Nas74, Sha18, dPN81]. Noncommutative-nilpotent [Yu96]. noncommutative-spaces [Sha18]. noncommuting [FW86]. Nonconvexity [Sol97]. Noncorona [AHJ21]. nondegenerate [CH12]. nonderogatory [FW13, Laf78]. Nondifferentiabie [Beb86]. nonegative [EM73]. Nong [Uhl13]. nongraded [GPW22]. nonincreasing [CL01, Loe93a, Loe93b, Ore19]. nonintersecting [AST05, Che22a]. Nonlinear [BBEL18, BEH19, BH13, CM19, CW11b, CX12, CW12, CW16, Cos21, FZ17, HZXL15, JLZ12, QC23, Wan17a, XW12, YZ20, Yan21, Zha16a, BH16, DWL14, ESH23, HBP18, HM18, JS20, JZ22, LM22, Liu22a, MK16, ME21, PLPC19, Sad23, SLUG04, ZL21, KA19]. Nonnegative [Bak84, BN88, CBQW19, Che92, DS16, JGK79, JT03, Lew77, Shi15a, SB03, Zha04c, AAJ12a, AJL18, AAJ12b, BS92, BS17c, BP74, BC94, BSS12, CP81, Cro18, DJ98, DJK08, DCLW19, Els78, EJD88, FW07, FJ15, FLS88, Gar85, HJP17, Har94, Hav88, HK92, JS84, JGS90, Joh75a, JP17, JMP17, JMS15, KSMB16, Kir11, KNX93, KNX95, LEH96, LLC21, LZ17a, LLGX22, LL78, LS21b, Mel13b, MV91, MS12, NST91, Rea96, Rod81, Roh06, Smi05, Sou83, Sou03, SZWB16, SPS15, VM07, WWW20, WS23, Wim85, Xu93, YSY17, vBW95]. Nonnegative-definite [Zha04c]. Nonnegatively [LS10]. Nonnegativity [BFI22, HRT15]. nonperiodic [CKN95]. nonpositive [FJ15]. nonpowerful [GSS09, SSG12, YMYS13]. nonprincipal [Loe87]. nonreal [EJ91, Esc93]. nonregular [Gil12]. nonrelation [Tho88]. nonself [FJS14]. nonself-adjoint [FJS14]. Nonsingular [AEdF13a, AEdF13b, BY76, BNP06, CL22b, FJ09, HT10, HXL16, IS15, JT95, Liu18b]. Nonsingularity [BLZ10, BSÜÖ13, ES98b, LWB11b, LCL22]. Nonstandard [Ore19]. nonsymmetric [BJ94, BT07, BSO17, SW22]. Nonuniform [BV16b, BV16a]. Nonzero [CL86a, CL86b, DF90, CM07, Cro18, Fri12, GGH22, GG22b, KMOV07, MY12, TDL18]. Nordhaus [Das21]. Nörlund [CW10, TD14]. Norm [AHK19, BDK21, DHHK23, Du10, Fla74b, FSWX22, FSY16, Gha21, Joh14, Liu22b, Neu74, Sed04, Sha18, SA10, ZMCN19, ANK22, AE10, AA12, BJZ13, Bou06, BKMMG22, CL17a, CD21, DO19, DM87, FM21, FH73, Gar75a, GAMA23, Gil93, Gil17, Gog21, Han17, Hia94, HM78, JKL22, Joh77a, JLL88, JN90a, JN92, KL95, LTT84, Lon83, LSZ86, Luo19, MO94, MR19, Oku91, Oku93, Pat09, Pep19b, Pre85, QHB22, Roh00, RF19a, Sai21, ST73, WLZ22, WZ23, YSK15, YL11, ZFSX17, Zou12]. Norm-parallelism [ZMCN19]. Norm-preserving [Liu22b]. Normal [AP04, BD94, Far94, GW04, JMS15, KL15, LMMR01a, New94, Arm17, AY95, BT08a, BK17, BD88, CDM22, Cha96, CN05, CM22, Der83, Epk89b, Epk89a, FX12, FJ22, GR91, GR93b, GRMA13, Hua90, IH11, KLS08, Kun99, Li98, LMMR01b, LL22b, MRT16, MV75, McC85, MZ91, NTU93, Pro95, QK92, Rad12, RRN22, RRN23, RW21, STZ94, SPK15, WZZ22, Woe97, Xu93, YC19]. normal-preserving [Kun99]. Normality [Ikr15a, WW20b, BFM20, Kit91a, LST89, Man92, WZ95]. Normalized [Are22, DP17, JT04, SD20, But10b, CL17c, GLS16, Hun83, IAM23, JP87, KP22, LLF08, PLLL20]. Normalizers [Luk74]. normaloid [BK21]. normed [KT83, Ng14, Pan15, Pas15, RS23, SRT22]. Norms [BKL02, HWL92, LT89b, OW02, So90, AGAG96, AG96, Bar76, Boz98a, Boz98b, Boz99, CW11a, CFZ23, CM15, DLQ18, Gol87, Gro17, Hem93, HJ87, HMN91, JL88, JN00, LT87, LT89a, LT90b, LT90c, Li91, LT92c, LLW07, LPT11, MS83, MSO83, MKS84b, MC89, Mar92, MM07, Mel22, MQ21, Mer78a, Mer81, OvdD77, Que08, Rho94, RF19b, RSPT17, SAV01, Sol06, TD14, Tal17, Wan17b, YLC19b, ZHH09]. notation [Car06]. Note [Ano73, Ano82d, Ano83k, Ano84d, DSD83, DS02a, HHC12, Lau81, lLpW16, LY22d, LHGL22, Mar82a, Mir81, Sam78, Tam85, TNJ22, AA21b, Ala18, Ala20, AAPR19, ABJ16, Ano82e, Ano83l, ACMS10, AOTR11, AZ22, Bap91, BOvdD91, Bar12, Ben90, BZ00, dAB17, dAB21, BH15a, But10b, Cha18, Cha83a, Cha84b, Che22b, CTW15, CKL16, Da 05, Da 88, Düf97, Els78, FGK23, For79b, Fri12, Gol89, GJMW86, Gum15, Gur79, Gur92, Har78b, Har81b, HWY17, Hug90, Jaa22, JS96, JS20, Jon92, KP11, KT95, KJ13, KKR12, KK18, LHP75, LN80, Li84, LP00, LPWX10, Lim79b, Lim15, LGB17, LL78, Mal89, Mal92, MS89, MR13, MM09, Min82b, Moo76, MOGN18, Mos15, MAM15, New78, OvdD77, dSP17, PR20, PSB23, Pri78, Pul75, RR95, San14, SGM22, SP87, TKT19, TXC08, TT78]. note [Til98, Uhl82, Vin06, WfMR16, Wan18, lWY19, mWyD11, Wei93, Wör20, WL16c, XQ16b, YZ14, Zha16b, Zou20, dCH21, vBW95]. Notes [Ami11, Cha83b, Joh75a, Wan16d, Zha89, LS21b, Mar92, Sim95]. notion [MS12]. notions [Hou09, SS22]. novel [ADJAH22, Dag23b]. Novikov [CKKK22b, DU21]. Nowhere [AGM13, AGS20]. Nowhere-zero [AGM13, AGS20]. NP [Roh00]. NP-hard [Roh00]. NPHSS [CL22b]. NSSD [FS18a]. null [BT13, Bea85, CLOK13, JVMCPR19, KKL20, KLO14]. null-filiform [CLOK13, KKL20, KLO14]. Nullity [Wan21, BS22b, FDD14, MW15, MF20a, zSqSZ16, WLM21, WL18b]. nullspace [AGS20, BC94]. Nullspaces [TMOvdD93]. Number [NW02, AEdF13a, AEdF13b, AYC93, BS21, BC05, BBG22b, BS16, Ben74, BRL22, BT15, Cha96, CHL18, CH22a, CFLL23, Dat80, DIJ23, DH16, DdF15, Epk89a, FMSS19, FS14, FLZ07, FGK23, HLB20, HHY22, Kir00, KK90, KS17a, LLW^{+}11, LT12a, LC14, LZZ17, LZX19, LS21a, LY22b, LL15, LLGX22, LY22d, MW18, MF20a, Mau78, MP13b, Ota13, OWTZ21, Pat83, PGLM00, Rem82, RW84, SM22a, SYZL17, SS96, Son22, SMS15, SD20, TW19, WW20a, WWZ20, Wan21, WWWT22, WL20b, WL18b, WL23, XW19, XZS19, YK22, YWZ21, YS23, ZXL22, ZZ15]. numbered [Gut22]. numbers [Dag23b, EB17, FW07, Fer09, Haw05, Hou09, HLZ05, Ilk20, Kar23, LL17, LM19, LZS23, Ogu20, Ost74, RRC23, SRT22, SLW21, TB01, WWS19, Wu09, Wu22]. Numerical [AMP02, BB22, BDK09, BB15a, BD94, BdP98, BLdPS04, BKS21, Ber02, BBP21, BM17b, BM18b, BS04, CW17, CN04, CN14, CL00, DTSI15, DvE20b, ERS12, Far02, GW98b, GW04, GWW15, HZX22, HL21, Joh75b, KL21a, LN98, LŠ02, LS04, LR04, MM07, MBB22, Sed04, haYsS99, Abo20, AS11a, AAK16, ADJAH22, Alo21, AM21, AYT83, AY84, AY95, AyT84, ALP94, BM22a, Bal80, BDK21, BD88, BLD91, BL92, BD96, BdP12, BdPN14, BBK22, BHK21, BP22, CRW21, CDM22, Cha92, Cha93a, CC14, Cha18, CLP22, CGW14, CL17b, Che94, CL94, CT96, CT09, CL13, Che14b, CL14, Che16, Che17, CM15, Chi98, CNP05, CN05, CNP06, CN07, CN08, Chi11, CN15, CGL^{+}16, CLN21, CH10, CGHK08, CKP09, CLMR09, CF23, Cro08, Da 94, DHKSH08, DvE20a, DM14, EDKA07, EM15]. numerical [FT87, Far93a, FMZ22, FNBR96, FHL^{+}14, FL16b, GW08, GW12, GWW16, GS21b, GS21a, GTZ75, Gol79, HH16, Hu89, HT91a, HT91b, Hug90, IOMA21, IOMAN21, JLL07, JL88, JSG94, Jon92, JNBR93, JR93c, JR94, KV23, KPL^{+}09, KP14b, LLS14, LLS16, Lei91, Lei92, LT92a, Lei93a, Lei93b, Lei95b, Lei95a, Lei98, LTT84, Li87c, LT88a, LT88c, LT88b, LT89b, LT91, LT92c, LMR94a, Li94, Li96, Li98, LŠS01, LZ01, LPS09, LT09, LPT11, LWC22, LP16, LP22c, LP22b, Man87, Man91, Man92, Mar75, MF79, MS79, MW80, MS82b, MS85b, MP87a, MS87, MR97, MP98, MM09, Mel22, MQ21, MZ23, MS21b, MS15b, MS16b, NNT95, Oku93, OM21, Pep19b, Poo80b, Poo94, PV00, QHB22, Rod08, RW21, RS08b, RSM21, RSM22, Sab19, SM21, SRS21, SA07, ShDHG08, ST80, STZ94, ST96, Spi00, Sza09, Tam85, yT86, Tam88]. numerical [Tam94, Tam98, yT00, TS19, Tho80, Tho87b, TGW14, Tsi83, Tsi81, WZD16, Wes75, Woe08, WC12, YLC19a, YLC19b, YC19, Yua98, bPFZH08, ZHH09, vD96]. numerically [DS23b]. Numerous [SG18].

Objects [NW02, LN10]. oblique [BT13, GT98, Gro99]. observations [Egg95, Sma17]. obstructions [Kar90]. occurring [JWP98]. odd [AM13, AP23, BBG22b, SSfW15]. odd-bipartite [SSfW15]. Off [DHS95, DW23, WC12, Zab89]. Off-diagonal [DHS95, DW23, WC12, Zab89]. Olga [Sch77]. Oliveira [BMD87, GLS13]. Omega [Net22]. One [Bre17, CKKK22b, Cha84a, LW23, ZWW22, ZWM23, Ab23, AVVÁ18, BHD10, BLP08, Ben90, Bén11, dAB17, Bot87, BM16, Cat21, CCL21, Cho10, CL01, CAP22, DH16, Ell95, Esc93, FT21, GMV18, Gur79, KS82, LL16a, Lim90a, LZ15b, MO22, MMRR16, Ore19, Shi15a, So99, SKP18, SH92, TNJ22, WL18a, WM21, Wat85b, YLW22, ZS05]. One-generated [CKKK22b]. One-matching [LW23, YLW22]. One-parameter [Bre17, MO22]. One-sided [ZWW22, ZWM23, WM21]. ones [HLZ05, Liu92, Wan05, Wu09]. only [MO91, Ram22]. open [CFSRL22, dF11, dF13]. operads [Bre17, BSO17]. opérateurs [Del94]. operation [Bre14, YW19]. Operational [Rom82]. operations [Cir15, Joh11, LZ15a, MN04, YGZC22, Zha14a]. Operator [CW11a, Dra19b, Dra21b, DH03, Far02, FZ21, Gus94, JKZ20, JC22b, KJF23, LL20, LB21, MŠ02, RS85, Tal17, WZ14, YSK15, Ab23, AKR23, AAP82, ADJAH22, ASR22, All14, AXH15, AO97, Anj17, BHC16, Bak16, BM19a, Bal22a, BDK09, BJS15, BD84, BBP21, BP22, BH16, Bou13a, BC20, CKGP21, CL91, CDJ13, Che14a, CL17b, CNUV16, CC18, CIP17, CIP18, Cve18, DMT21, DD10, DCIW10, DV18, DD12, DC20, DC23, DW23, ESAJ18, FJS14, FTGSNV22, FH15a, FJK98, FMS19, GR23, Gha21, HC14, HH22a, HS81, HSW19, HOT15, HQ18, HZX22, Ikr15a, JE21, JKH20, JX13, KS14b, KMS18, KD16, KMS21, Kit91a, KK13, KL19, KL22, LY22a, LM85, LWC22, LS15, LD19, LFY16, MO94, MM20a, MM20b, MGH19, MF20b, MFS22, Mos19, NY23, Nik22, PSA14]. operator [PSMA16, dSP16b, Pep19b, Rad16, Raï14a, Raï14b, Raï19, RSM22, SRT22, Sed11, SR17, SG18, Son14, SY23, TR21, TD14, Tom22, TRM22, WHL22, WHC22, XQ16b, XSH18, Yu14, ZZ16, ZHH09, ZZZ16, dPN81]. operator-eigenvector [Che14a]. operator-entried [dPN81]. Operator-norm [YSK15]. operator-valued [GR23]. Operators [BLdPS04, Ber02, DS02a, DKK19, GW17, HT91b, JKZ20, MŠ02, Pop20b, Sed04, SB03, AN19, ADRP14, AR22, AGV19, AAPR19, Alo21, Alo23, AHK19, AFM13, Amm18, AP17, AO22, AFA22, BK21, BHD10, BM15a, BFM20, BN22, BDK21, BP89, BP90, BP92, BL93a, BdPN14, BM22c, BO22a, BLZ11, BHK21, BP22, Boa13, BJZ13, BPR14, BT17, BCM18, Bou18, BE22, BM15c, Cam14, CW81, CDPS23, CQH15, CL17a, CZZH22, CN05, CGL^{+}16, CH10, CF23, CGM19, CIN15, CM22, DH17, Del94, Den14, DLD21, DLD22, DV19, DKY22, DM13, Dra10, Dra11, Dra16, Dra18a, DK21, ER22, EAM19, Erd22, Esh17, Fan83, FMX18, FNBR96, FJK22, FMXZ22, Fur08, GT14, GLX22, GGGA21, Gar23, GCSC20, GR81, GW12, GJR22, Gil16, Gog21, GR91, GR93b, GG22b, GHL20b]. operators [GBM21, GC22, Han17, Han19, HMK22, HBM15, HH16, HX21, HL21, HKS94, ISCS11, IOMA21, IOMAN21, JS19, Jac79, JKH20, JKD16, JX13, JKL22, Jon78, JS15, KN22, Kan16, KKL22, KKL20, KDHJ22, KKY22, KP92, Kit91b, KL15, KLL22, KLS08, KK22c, Kut19, LLMT05, LŠ96a, LŠ96b, Li87c, LT88c, LT88b, LT90b, LT90a, LT90c, LT93, LM93, LP93, LMR94b, LŠS01, LZ01, Li11, LPT11, LD19, LS17, LGG19, LW20, LPX23, LB21, Luo19, MCL20, MSS13, MPD22, MDP22, MF79, MŠ09, MZ91, MP18, MP19, MN12, Mos20, MM22b, MBB22, ME18a, ME18b, ME21, Ogu20, Oi20, OM21, Orl18, PSS15, Pep11, Pep19a, Pep19b, Pet10, Pop17, QLK22, QHB22, RT09, RS85, RRN22, RRN23, Rao18, Ras17, Ras19, Rho94, Rod08, RF19b, RF19a, RSM21, RB22, RSPT17, SPMR20, Sai21, Sam78]. operators [SES21, ST19, Sed11, Sha18, Sha20, SA10, SHC22, Shk12, SAK22, Suc22, TH14, TD16, TD19, Tan13b, TMH21, UB22, Wal79b, WD17, WTD22, WZZ22, Wil81, WZ22, WW20b, WC12, YK22, YC19, ZMCN19, ZZ16, ZZC17, ZZ00, dBS00, Ano22c]. operatorwise [DKK^{+}20]. Oppenheim [LP22a]. optical [FP17]. Optimal [AP17, CC23a, dBS00, BGV11, Blo95, CHFC19, DS23b, Igl99, Knu85, LN10, NW02, PSS15, SS79, Sha91, Sma17, Ves80, lWpW14, WM84, XJS20]. optimal-node [lWpW14]. optimality [CPKK22]. optimally [LOO22]. optimization [BLX18, SW22]. orbihedra [ZZ00]. orbit [CHZ22, Far90, Koi21, LT89a, LP22c, LP22b, RS07, Zha01a]. orbit-closed [LP22c, LP22b]. orbits [CQ18, DRW93, GS19, LPS08, LT13]. Order [AAT17, KW17, Peñ02, vBW95, ACMS10, jAS15, jAS21, BS21, BTT20, BS18b, BH13, BH16, BMGMC16, Cam83, CNW21, Car20, CQW16, CJ12, CDK22, Cir15, Cir16, Cir17, CC22b, CGM19, CM07, CCLN16, CIN15, CID16, CIM19, DTSI15, DLD21, DD14, – DL92, – Dok93, – Dok09, EM95, FTZ08, FF21, FM22b, FP17, FJK98, GLX22, GG22a, GG22b, GZY18, GHT16, HWX22, HS83c, HSX16, HSM17, Hil07b, HK10, HFSZ22, Jin17, JKB17, KMS16, KS17a, Kre16, Kri87, sL19b, LPT14, Leg06, LS22b, LW08, LFY16, LZB23, MRT14, Mat90, MWC22, Mit92, MB22a, New80a, PBM20, Rad16, SWL13, Slo14a, Sta15b, SF17, Tos79, TRM22, WL18a, XQ15, XQ16b, XSL15, Xu18, YLX20, YJ21, YCZ21, ZX10, ZCP17, tHK17]. order- [Sta15b, SF17]. ordered [Leg06]. Ordering [GZ22b, Har78b, OST20, WL15, WC15]. Orderings [MP16b, Gro97, Hey88]. Orders [Mar18, And96, JS15, MRT14, Mar15, MRD15]. ordinary [ZG08]. Ore [BZ22b]. Oriented [MR19, CT16, Gon17, Sta22, Tik22, WLF22]. orthant [Mar92]. orthant-monotonic [Mar92]. Orthochronous [Les93]. Orthogonal [BD97, GP14, GK02, GGW76, HMR22, Hol95, HK13, Hor04, HZ21, Koi21, Lop11b, Per17, ySpW22, SZ19, Sri20, BT11, Bas17, BBM15, CR22b, CH87, DP00, – Dok09, DWCD19, Ell95, FM22a, Gal08, GAGM18, GZ99, GHT16, JWP98, KW92, KMP^{+}19, KLS08, LM93, Lim01, Mal95, MRT12, MGAH19, MR08, MAH18, MAH21, ME18a, SM19, Sta23, TL12, Vin06, WG91a, Xu18, Zha01a]. orthogonalities [TAK22, WZ23]. Orthogonality [AO22, BM22b, DS02b, Gro17, Hol04, RŠ06, RŠ08a, RB22, WZ22, AR15, AV22, BBK22, CP13, Fan88, FMZ22, GP14, ME21, Per17, RŠ09, RS23]. Orthogonally [AGV19, RS17]. Orthonormal [KCBR87, LZ01, MC86, WG91b]. orthostochastic [PT87]. orthosymplectic [Lee94]. oscillatory [CP97]. Osima [HK16b]. Ostrowski [Mar76, QLK22, She77, WWW20]. Ostrowski-type [WWW20]. other [BY76, Bar07, BSS12, MRT14, Pie92c]. otherwise [Bal78]. Oussa [Ano16c]. Outer [KM14, SPKC15, SĆK^{+}20, AB17, Boa19, FLT20, KCSĆ21, LMR94b, LYZW12, MW22, SMM22, Tia01, TS05, YHS13, YGW15]. outer-planar [YHS13, YGW15]. outerplanarity [Vo83a]. outindependent [FG06]. outputs [GOS22]. outstanding [Joh82b]. oval [Mel13a]. Ovals [Mel12]. Overgroups [Gur97, GL97]. overlapping [Vei83b].

P [JT95, TT74, XCC99]. P. [Fie84, Zha95]. P.I. [CV94, Di 88, Per15, Reg93]. packet [Far17]. packing [DWL20]. PageRank [Guo20]. pair [BT13, BZ90, DKY22, GC22, Han21, IS16, LD17, TT06, Uhl79, WZH13a, WZH13b, WLW23, You15]. pairings [Yuz84]. Pairs [BG86, Do 96, Rod08, SB03, Tau74, Aln11, AW13, AO22, BJS98, BZ99, Bea78, BDD23, Cal97, Cal99, CLW15, Che17, CDK20, Dod13, Dra23, Fer10b, FKKS11, FXY22, Gal08, GP06, GTM14, Gur79, Gur92, HZG14, HZH20, KHG13, KC16, KKY22, MP21, dSP22, Pet00, RBB22, SH98, Uhl82, WHG17, Wat77, ZLS22]. pairwise [PP22]. palindromic [PM23]. pall [Adk95]. paper [Mar82b, Tam85, lWY19]. parabolic [CW11b, CX12, CW12, CY22b, Den05, LCL13, Sch79, WY11]. paracontracting [WC16]. parafermion [LY22a]. parallel [Ber17, Dji16, FMXZ22, LSX19, WTD22, You15]. parallelism [TAK22, ZMCN19]. parameter [Adh13, Bre17, CGH15, Hua22c, MT22a, MO22]. Parameters [Rod02, GLLM21, Koo95, LWL21, PLLL20, XSW22, ZL22]. parametric [DM22c, HMAD13, Ska17]. parametrization [AG15]. paraphrases [Leh91]. Pareto [AS17, BS21]. Parity [ASZ02, TI77]. Parseval [CC23a]. part [CS21b, Dic84, Dru15, JN80, Joh84, JL91, MAH21, SU77, Sil99, ST96]. Partial [HS83b, Tia02, BMGMC16, Cal94, CKGP21, CMM20, Cho18a, Cho18b, CKS20, CM85, Cir17, Cla92, DD10, FM22b, GWW16, Gro97, GHT16, Har78b, HK10, JR84, JTU00, JS15, KMS16, LPT14, MT22a, MRT14, Mar15, MRD15, PT10, SM23, TSYB82, WL15, WC15, WL18a, WCS^{+}21, Woe93, YJ21, ZCZT22]. Partially [Bal18, CC17a, Leg06, Sil95]. particular [BS15a, BCdHP10]. partite [Shr90]. partition [Ati20, BX11, DF87, FTA20, Fon89, YZB22]. partitioned [FL16b, JMP17, MT22a, YLC19a]. partitioning [AMK23]. Partitions [Kes83, PM81, PM85, ST97a, Chu90, Cla92, Cve95, Fer09, Fon92, Hey88, IK22, RW84, SS12, TOM88]. Partly [MOvdD90, McD92b]. parts [AS12, CB22, GW17, GHK16, Ha89, KL22, Pal79, PP22]. Pascal [AFVL00, CKM18, CKMY22, TB01, ZAK22]. Pascal-like [ZAK22]. passage [HWX22, KN00]. Path [DP98, JS99, AAGL22, Guz88, Hou17, RP90, SRD19, Wan21]. paths [AST05, KMP^{+}19, RW84, Ste94b, WSH^{+}11]. pattern [Ach77, BJS98, CHK00, LLS10, LLL09, LCL22, MM11, Mer83a, SM94, WYM11]. patterned [NU94, vR11]. Patterns [MOTvdD03, BJS98, BM12, CJ97, Cav10, CJ20, CM07, EJ91, Esc93, FHL^{+}13, GGH22, GS01, GS06, GSS09, Har91, Har95, HS93a, HLOvdD20, JL93, JWP98, KMOV07, KOvDD09, KMT96, Kir11, LGA^{+}13, LHbL^{+}15, Ma15a, MY12, YMYS13, YHCW14, ZHLZ13]. Pauli [MS15b, MS16b]. PC [Shc21]. PC-maps [Shc21]. pd [YZL22]. Pellet [Mel18, Mel20]. pencil [DKY22, Gou13, LC18b]. pencils [Atk91, BY76, DS22a, LL14b]. pendant [LT12a, LDL17, WWZ20, WWWT22, WDJ10, XW19]. pendent [ZL19]. Penrose [BT22a, BST23, Bap97, Boa19, Bom02, BGZW16, CC22a, CHZ22, DK23, DD10, DW12a, DLW18, Fie95, FSWX22, Gro99, HPV84, HPV88, JGK79, KMRS89, KR07, Kyr11, Li11, MZ15, Men22, PBM20, PM22, PA10, Rom21, Rom22b, SC20, Siv06b, SZBW16, SPS15, TC04, Tia05a, YZ14, Zha04b, ZFSX17]. pentadiagonal [AdF21, MMG15]. pentagonal [Sim83]. Per-spectral [ZWL15]. Perfect [Are23, CF21, CG16, KMP^{+}19, LCWW22, PB17, TM19, ZLZ20, AFTG19, BJ94, BTT20, DP09, FGHL22, GZ22a, KS22a, LCXC22, Que08]. period [New80b]. periodic [BB19c, CKN95, CLW15, DdF22b, IOMA21, IOMAN21, Sat11, Sie99]. periodicity [MNS22]. periods [JKY23]. Permanent [BC78, CD03, Hon92, Pat03a, Coe96, DJ89b, Fal97, FS86, GJMW86, Hon91, Hon93, Hu94, Hwa85, Hwa91, HS93c, KS84, Lim79b, Lon73, LM89, MS85a, Mer80, Pat96, Phu80, Sin73, Sin76a, Sin80, Sin81b, Sin84b, Sol97, Sol00, Stu92, lWY19, dF05, dF13, dF19]. Permanental [JP87, LZ11b, Pat89, Bap91, BF75, Cha90b, DJ89a, For79b, GHL10, HT91b, Kop94, LT92a, LZ14a, LW22b, Lon97, Mal89, Sou03, WL18b]. Permanents [Ach77, AAT16, CCKL20, FW86, For88, HZ03, Kit81, KS76, Pat81, SHC02, SJ02, Wan05, BE90, Bur95, Cal95a, Cha83b, Cha84a, Cha88, DH91, ES73, For79a, FM78, Fri82a, FLS88, Fri90b, Gol92, Gyi78, KS82, Krä87, LHP75, Min74d, Min80b, Min83, Min84, Min87, Pat92]. permutation [BG79, Bru79a, Bru79b, CFZ23, EOvdD09a, EOvdD09b, HS81, Hon92, KM19, Mer78b, SS98]. Permutations [Cha79b, Kap77, Leh76, LF16, SRW07, Wan80]. permuting [CL14, Hol85]. Perron [AD17, BOvdD95, BAN12, DM87, EJD88, Fie73, GS83, Har78a, JP17, KNS96, KF98, Lia23, Mel13b, Mey15, Roh06, SPS15, vBW95]. personal [Pri78, TT78]. perspective [BT11, CDM22, Koi21, NS18, Nik22, Woo22]. Persymmetric [Vei82a, Vei82c, Vei82b, Vei83a, Vei83b]. Perturbation [BCM18, CQW16, CLX03, Din00, – DM79, JKZ20, LYS15, LSX19, LZ22, MMRR09, ZY13, ZZZ16, Ben90, DLW18, Din98, FSWX22, GLS16, JLDS12, KN97, LL17, MMRR16, MZ15, Men22, ST15, SWL13, So99, TXL21, WfMR16, YZ14, ZYL13, ZVL18]. Perturbations [Gal03, PG22, ST97a, Amm18, BOvdD91, BCdHP10, BZ22a, Bén11, CC18, DC20, Fer09, Fer10b, KSU22, LS22c, MMRR16, SS12, XS18, ZFSX17]. perturbed [BKP01, CTW15, FS89, GTM14]. Peterson [GTLN18]. Petrov [AP23]. Pfaffian [BFM22, CT09]. Pfaffians [IW95]. phase [FM21, HLY22]. phase-retrievable [HLY22]. Phelps [DKK^{+}20]. physical [HWSH13]. physics [BdP98, Da 94]. PI [MP21, dSPdS22]. PI-algebras [dSPdS22]. Picard [CH06]. Pick [CH12]. Pick-type [CH12]. PID [Hua01]. piecewise [Kut19]. Pietsch [BPR14, Pop20b]. piezoelectric [LXQ21]. piezoelectric-type [LXQ21]. pinches [RRR^{+}22]. pinching [PT87]. Pinchings [BKL02]. Pisanski [Ano16b]. Pivot [KMS02, EF98, KMS07, ST08]. pivoting [EM95]. places [Kir04]. planar [CSY18, TZ19b, TWZ16, VDW85, YHS13, YGW15]. Planarity [CW90, Blo95, Dic84, Kar90, Vo83c]. plane [CBQW19, CM15, Chu90, DS16, MM07, MQ21, Mir09, RW84, ST96]. planes [AG18]. Plethysm [CRY95]. Plethysms [CV94]. plücker [CC92, Tam91]. plus [Ikr15b, NSW21, Sol06, lWpW14, lWY19]. pockets [Bar08, NP15]. Poincaré [Bed10, Bed11, BX11, Car90, DF82]. Point [– Dok81, CL22b, CMM22, CIP17, MHM22, Sch75, Sch78, Wan17a, Zha20a]. Point-sequences [– Dok81]. points [BLP12, Beb86, BS15b, BDK^{+}05, CSS21, CCW20, CJK^{+}13, Gei78, Hon92, KM98, LPS13, LKLC22, MS87, NT91, Rao18, SRBB22, Spi00, TH14, ZX16, ZXL16]. pointwise [GJ22]. Poisson [BE23, CD15, Kay17, Mar12, YH22]. Polar [Jam95, De 08, DP09, Fur08, LMMR01a, Pas16, Pat87, Pat17, SU77, TN22, Ter22b, ZYL13]. Polarem [SU77]. polarization [GR93a, Vel21]. Polarizations [BBT07, Far93b, Wil81]. polarizing [Ous15a]. pole [CFP10]. poles [AMZ13]. Pólya [SG18, WW76, Whi79]. polyhedra [dlP20]. polyhedral [AGP22, MM07, Orl18, Tam77]. polyhedron [FS86]. Polynomial [AS11a, BP09, Cal03, CGH14, Mel19, Mel22, Red02, SA07, Wat77, dTS20, Adh13, ASAS79, Ami11, Ami80, Are22, BB17a, Bra86, Bre14, BM07, Cal94, Cal97, Cal99, CdSS17, Dat80, Dia90b, DMdO92, Düf97, FZ11, Fie81, FP17, FdSV16, GA85, GdSV16, GMD85, HHC12, HMAD13, IS15, IM18, Iop19, JRMFMASS16, KSU22, KLPK17, Kul22, KK16, LLC22, LB19, Lip98, Lit17, Mel13a, Mer79b, Reg89, SRK22, SS07, Sun20, SK15a, SK15b, Uhl99, WC22b, Wan22a, Wim88, WW12, dO75]. polynomially [CM22, Kan16]. Polynomials [AP04, Far02, LM85, Mes13, New76b, PT02, Spi76, AS11a, AM13, AS12, AGV19, AADD18, AGZ00, BS15a, BCR91, BD22a, Ben74, BH74, BG21, BO22b, BG79, BMP10, BPR12, BPR14, BD13a, BP18, CC17b, CNP05, CNUV16, CR22b, Coh88, DP98, De 90, FMSS19, FGMSS19, FW73, Fit94, GAGM18, Gar73, Gar79, GV89, GKLR82, GLR82, Gór22, HHSY22, KMP^{+}19, KS07, KK90, KLPK17, KML95, KLZ96, KL92, LRT91, LZ14a, LT22a, LW22b, Lon73, LT07, MW74a, MF77, MR97, MP98, Mar90, MRT12, MGAH19, Mel18, Mel19, Mel20, Orz84, Ota13, Paz84, PR12, Per15, PM23, Pie92c, PV00, Rei77, SA07, SX23, Sil88a, SS96, Sil98, Sla16, Soa08, SL95, Sra20, Tie16, Tre87, UG21, UF22, Vel18, Vel21, WLB19, Wan19c, Wat86]. Polytope [SHC02, Bru85, BD13b, Hwa85, LZZ17, dF11]. polytopes [CCLL22, CCW20, Har91, Sie99]. Pommerenke [Bol12]. Poncelet [CN04, Mir09]. Pontryagin [Sin83]. Popoviciu [MM22a]. populations [Pan20]. porism [Mir09]. portions [CN08, ERS12]. portraits [LH22]. posed [BHM76]. poset [Foš05a]. posets [AST05, BC91, KC16, Ste94a, WW01]. posinormality [BM22c]. Positive [AB22, BBP78, BR01, Cha90a, CIDK07, DH03, DWL14, Grz91, Hil07b, KRR18, ANK22, ARK22, AS20b, AGP22, AG15, AO78, And96, AB19, Bak84, BSM98, Car19, Cas23, Cha92, Cha93a, CKS20, CS21b, DM17, DM20, DHHK23, DPP17, DH86, DM13, DGJ^{+}16, DJ89b, DJL94, DJ98, Dru15, DN18, ER22, ESH23, Fan73, Fan92, FW07, Fla75, FH15a, FLT22, FG22, FS15, FSY16, GT14, GP22b, GMZ11, GJMW86, GLP92, HK16a, Har78b, HJ19a, HJ19b, HJ20b, HS88, HNS99, Hil07a, Hil11, HBP18, HM18, Hua22c, HM95, Ibr00, JP15b, JN80, JOvdD84, JP87, JR09, KMS18, KMT96, KL22, Koh11, LS10, LSOvdD95, LM96, LP11, LT11, LT12a, LL16b, LF16, LP22a, LL14b, Lia23, LW12, Liu22b, LSX19, Mai11, Mar13, Mat92, MK16, MP16b, Mol15, MOGN18, Mou22, NMAB22, ND13, NY23]. positive [Orl18, Pat92, Pep19a, Pep19b, PV20, SH21, ST94, SM94, SMBB^{+}15, SM22a, SBMP16, SL17, Sun20, TZ19a, TZ21, Wal79b, WZD16, WCS^{+}21, WL22, YH05, ZX16, Zha16b, Zha20b, ZL21]. positive-partial-transpose [WCS^{+}21]. positively [NS85]. Positivities [WLX16]. positivity [CC98, FJM00, JNT96, YHCW14, vE22]. possess [PP16]. Possibilities [SS96, JR84]. Possible [FW07, HLZ05, JL91, Wu09, BF96, Uhl13]. potent [BT06, BT08b, BLZ10, Den14, HLZ11, McC85, Pab22, Rom19, Rom20, Rom21, Rom22a]. potential [HMO11]. potentials [HMK22, KMP^{+}19]. potents [Zho22]. POVMs [GKO22]. Power [CL21, DK21, FJ22, KPP13, MF92, MO93, AFW22, BM07, Bun91, CP15, DS23a, Dre22, Gol07, IAM23, JC22a, JM85, KMS21, KM22, Lei93b, MGA17, New80a, RA21, SN22b, SAK22, Sri20, Tam88, Tan10, TLL11, TLC16, Tom22, XL11, Zha14b]. power-associative [Gol07]. power-compositions [BM07]. Power-exact [MO93]. powerful [MM11, WYM11]. powers [ARK22, AGR91, Ash16, Aue98, Bou13b, CFL^{+}22b, Coa99, Dru15, FKR93, Fri82b, GA20, Gar21, Gór22, HL20, JGS90, KW21, KK22a, Mcf79a, MV91, MNS22, Mou22, MM23, Pat83, Pok87, Stu89, Tik22, TT18, Vas87, HOT15]. pp [Ano16c]. pr [MR17]. pr- [MR17]. practical [WSZL17]. Pre [sL19b, DLD21, LW20]. pre-Lie [LW20]. Pre-order [sL19b, DLD21]. Preaffine [DW92]. preceding [Mar82b]. Preconditioned [BNKJ22, LZ16, WZY22, WL22, ZQZ21]. preconditioner [BH23, HFSZ22, XJS20, Zha20a]. preconditioning [CP16]. predictor [Chu84b]. predictor-corrector [Chu84b]. Preduals [BPR12]. Preface [Ano16d, BLZ12, CK12, KL11, KL17, KL21a, LRT08, LT09, LPW14, LSS16, ŠTZ15]. Prehomogeneous [Ven19b]. premières [BC06]. prescribed [AMZ13, Bot94, BM89, CM92, Cra09, De 82a, DHS95, DS09, Est80, GPM14, GMD85, Her83, HLB22, KM14, KLPK17, Min84, SRK22, Sil88a, Sil88c, Sil88b, Sil90, Sil93, Sil95, Sil98, Sil99, Tho75a, Tho75b, TI77, Zab89, Zab90, Zel85, dO75]. presence [FW13]. preservation [PG22]. Preserve [Lim90a, SB03, BCN05, BP89, BP92, CL17b, HKS94, JS94, JLL88, JP07, LT88c, MZ91, Slo16a, Wat77, Wat82]. preserved [Can80, CKR21]. preserver [Ano92g, BLP92, CL01, Hua16, Joh11, MŠ98, ST08]. Preservers [BJS15, Ben18, Gur97, HX21, LM03, LS04, LR04, RŠ09, SM22b, SK03, SBMP16, BPS00, BS92, BL97, BL01, BSL01, BS16, BS17b, BS17c, BS20, BS22b, BBEL18, BH16, BMGMC16, CL95, CC14, CL05, CLN13, CLMR09, DRY22, DGJ^{+}16, EBB18, FHLS13, FJZ21, GP14, GL97, HQ18, Jaa22, KS12, KP14a, LL16a, Lau81, LT88a, LT92b, LT92c, LT92d, Lim97, Lim01, Lim10, LT14, Lim15, LRD21, Loe92, Min77, MSS19, Per17, Sin82a, Slo15a, SA14, TZ06, ZS05, Zha05]. preserves [Coe96]. Preserving [DŠ04, Dol04, Kob06, Zha04b, AN19, AA21a, AS17, AKS18, AO22, BHD10, BHD11, BP90, BL93a, BEH19, BBBS09, BDS13, BS97, BCS18, BH13, Bot79, Bot87, BE22, BM15c, BSO13, BMGMC15, CGY13, Cat21, CCL21, Che14a, CW16, CY22b, CK20, Cos14, DS13, Dol07, DHB07, EAM19, EL90, Foš05b, FKKS11, FHL^{+}14, FA22, GM18, HH22a, HR85, HS17, HH16, Hua10, HM18, HZXL15, JQ15, JLR^{+}92, Joh11, Jul21, KP92, Kov77, Kun99, Li87c, LT88b, LT90b, LT90a, LT90c, LM93, LP93, LMR94b, LŠS01, LZ01, Lim10, LJ11, LL11, Liu15, LL18, Liu22b, MF79, MR12, MŠ09, MO13, McD92a, Mol09, MN12, ME18a, ME21, Nag17, Pet97, Pet10, PR20, Pie79, QH17, QC23, Rai87, RJ04, RŠ06, RŠ08a, SH21, Slo13a, Slo14b, Tag12, TH14, TH18, TKT19, TS21]. preserving [TB92, TXC08, WZC11, WZL13, WJ16, Wat83, Wat85b, WZ22, XCC99, XXF22, YZ20, ZHH09, ZCZ12, ZX16, ZWW10, ZXF15, ZL17]. Pretty [WC22a]. Price [TT78]. prime [BD22a, Cen11, DS13, Far17, GLdS22, JL12, LWZ15, Lin21, New80a, Nor95, Nor08, Per15, QH17, Sta14, Sta15b, Yao22, DDA20]. primeness [Per15]. Primes [RS74]. Primitive [BS17b, dSP16a, WYM11, AF11b, CC17b, GSS09, GM21, JKKC22, LLS10, LM95, LLL09, Liu92, SSG12, YMYS13]. primitivity [BP89]. Principal [CT21, Dru15, LP03, LMMR15, NS92, NS93, Rao81, RBB22, Vrb86, XFG14, Bén10, BCCL22, CFSRL22, Cho82, Dix79, FT21, Fie81, HKT15, HS87, HK16b, Hol84, HT91a, HO17, Joh84, JOvdD84, LH23, MR18, Nas75, NTU93, Pri78, RM22, ST08, TT78, Tho92b, TDL18, TZ21, ZHZ11]. principle [BdP21, dC13]. principles [LL14b, Sin83]. prism [CLX22]. Private [KP14b]. Probabilistic [BAN12, WW76, Sim83]. probabilities [CMM22, Ger86, Joh82a]. probability [CPZ17, DPP17, KM22, LN14, Sti87, XXF22]. Problem [Ano82f, CLLT93, Tho82a, Adu15, AHK92, AS11b, AG15, AO78, Ang00, Ano83m, Ano85g, Ano86e, Atk84, AB22, BC05, Bap92, Bar88, Bru87, BK87, CC17a, Cal95b, CKLP13, CH12, Che91, CP18, CP21, Cro18, CCLN16, CLW21, Cve95, DLQ18, DS16, DF82, DP18, Dod13, DS22b, DJK08, DWL13, DCLW19, DN18, FS22, FdF09, Fie90b, Fla75, FH89, Gar85, GB77, Gra93, JDD21, JC74, JL85, Joh89, JJK95, JN14a, JNS23, LW96, LLW07, fLLyP15, LMMR01b, LW10, LZS23, LL78, Mer79a, Mil15, Min80a, Min82a, Mir18, MMG15, MS15a, MAM15, Mum86a, ND13, New83, Raw81, Rea96, RP04, SR17, Sin84b, Smi05, Sol06, Sol80, SW22, Sto79, TXL21, Tau73b, Tho85, Tho87b, Tho91, Tre85, Tre86, Ves80, Wan17b, WW01, YZC21, ZXL22, Zha92, ZY13, ZWLZ22, Zha22, ZVL18]. Problems [BL80, AAB17, Ano82g, Ano84e, Ano84f, Ano92g, jAS15, jAS21, BCR91, BLP92, BR01, CQW16, Che92, CS19, CL22b, CL01, CKKY23, CDF21, CID16, De 84, Din98, DJ98, Fri88, GMZ17, GLL18, GGGA21, GKT97, GM87b, HZV23, HRMP95, Hua16, Joh82b, Joh11, JTU00, KPL^{+}09, Li85, LT88a, LZ16, Li21a, LLV18, MM06, Man91, Mer91, MŠ98, MB22a, Sad23, SC21, SQM99, Sch91, Sin82b, TT85, WS18, Wan22a, Wil83, WL16c, WZY22, Xu93, YLW15, Zha00, ZHZ11, ZV19, dF13]. Procesi [Reg87a]. process [DMV18, FSFSO^{+}14]. Procrustes [Rau18]. produced [Tia11a]. Product [AK94, BLdPS04, Cal03, CGL^{+}16, GN02, AN19, AF12, AJL18, AHT85, AHJ87, Aze99, BH23, BCDF09, BD22a, BM17a, BS22c, dAB17, BG21, Boa13, BM15c, BM17b, Boz99, BH15b, BZZ10, BM16, Cam78, Can80, CK17, CX23, CKR18, CKR21, CM19, CS23, CW16, CC21, CY22a, CHS22, Cho18b, Con77, CD15, Di 88, DM13, Dra19a, Dra19b, EAM19, FD14, FKKS11, Fow80, Gha13, Gór22, Gra10, Gra11, GT98, HZ11, HPS83, Hol85, HH16, HXL16, HM22, HWSH13, HZXL15, IS15, JC22a, JKH20, dJfLlZ^{+}21, Joh77b, JN90a, JS99, JN00, KT06, KR14, Koz17, KLJ23, LV18, Lee94, fLLL^{+}17, LH21a, lLZjZ19, Lim90a, Lim23, LMMR01b, LJ11, LL22a, LLY22, MW23, Mar97, MP16a, Mat92, McF79b, MWC22, Mir92, Neu74, Nor95, Nor08, Ota20, PM22, Pat08, Pat09, Rad16, RP22]. product [RW96, Rod81, RC18, SH21, STME10, SN22a, SN23, Sil93, SAV01, SZBW16, SZZY18, TZ19a, Tan13b, Tho74, Tho88, WZD16, WJ16, WCS^{+}21, WS23, XJS20, XQ16b, XXF22, YZ16, Zha00, ZPCZ16, Zus22, vR11]. Products [Bal77, Bal79c, EM87, Ell95, Gib76, Hou22, Laf83, MKS84a, O’R15, Pat81, Rod97, ST91, Sul08, Wil99, Alp95, AAdFS19, AO97, BK83, BJ75, Bar76, BJ93, BW23, BJS15, Ben18, BE22, Bre12b, Bre16, BSO13, CW81, CGY13, CL85, CG17, Cat21, CD22, CL17b, Che95b, CKL16, CK20, CW22, CG16, DSD83, De 78, DF90, Dol07, Du10, Erd22, FL16a, FJM00, FHL^{+}14, FIS06, GG22a, GWW15, Gil16, Gro96, HH22a, Hal91, Har84b, HS81, Hol84, HMN91, HX21, Hua08, HLH18, Joh74b, JOvdD84, JOvdD85, JN90b, Jul21, Kit91a, Kur04, LL16a, Lim15, LMMR01a, LW08, LFY16, Liu18b, Liv94, MKS84b, NS99, Ng14, NY21, Pat82, Pat92, Pat08, dSP22, Pet10, PT87, Poz22, RS10, Sil90, Smi94, Spi04, SW81, Sur98, TH14, TKT19, TS21, Tan16, TC04, Tia05b]. products [Tia18, Uhl82, Wan80, WZ97, WC16, Wu22, YZ20, Zha89, ZHH09, ZX10, ZXF15, Zho22, Zim89, dCdSA22]. profile [BL21, LB19]. programming [Bar07, Che92, Gur98, WY21]. programs [Caf13]. projection [Gro99]. Projections [XCB18, AHT85, BR10, BCI13, DPR22, DWCD19, FXY22, Gal08, KW17, KR07, LD17, MR08, Rei83, Szw18]. Projective [Jam88, Rei75, Den05, LPS21, Ng14, Pan13, RŠ06, RŠ08a, RŠ09, SZ85]. projector [Byt20]. projectors [BT11, BT13, DW12a, GT98, KR14, TL12, TCI13]. prolems [EJ88]. prolongation [BBM20]. Proof [GSZ22, Kit81, Rad80, Tho86b, AY95, Aue95, Bar12, Bay22, BD96, Car20, Fri82a, HL09, Koh11, KAE22, Mag22, Mel18, Min82b, Poo80a, Reg87a, RW84, Rem90, Roi99, Tam91]. proofs [BM18a, CL13, DHT94, MS82b, Rem82, ST80]. propagation [JL22]. proper [AMZ13, Ger82, JC22a, Tho75b, lWW15, ZCZT22]. Properly [Fri97, Mar20]. Properties [BD94, EMS02, GZH19, HM77, HK17, JKZ20, KY17, McC85, WD17, AAD22, BCNG23, BC82, BZ19, BS17e, BS18b, BM18c, BM19b, CKR21, CL94, CKM18, CJK^{+}13, Dag23b, DST91, DLD22, Din15a, Dra13b, Dra17b, Dra22b, DS23c, EB17, Far17, FJK22, GCP20, GN20, Gho19, GS20b, Haj18, HLY22, HLS18, JNBR93, KI16, KW17, KS84, KM98, KD23, LZ14a, Lor00, Mai11, MSS13, MRT14, Mar13, MH22, MH12, MCIM21, MAH18, MAH21, NG15, PG22, PRS03, PRA22, RP04, SMBB^{+}15, SSfW15, SHC22, ST08, SKM17, SCL^{+}22, SL17, TG09, Tra89, Tra90, VM07, WC12, Zel85, ZW17a]. Property [Che95b, CN04, Ger74, NY21, AHJ20, AHJ21, AZ22, BNP06, BGM22, BMP22, BCEM12, CNW21, Can80, CP16, CAP20, DKK^{+}20, DdF22b, DH22, GF10, HA22, Har84a, JVMCPR19, LS17, LFY16, LLV18, LC83, MW19, MW21, MB22a, ME18a, ND05, OS18, Per15, PSB23, Pop20b, Rom07, Tau74, Wer87, XQ15, ZCCIW12]. pros [LLMN20]. Proving [Red02]. PSD [DNY22]. Pseudo [AJL22, FG02, AN19, Bal79c, BS18b, – DM79, Mir18, Mos21b, Per19, SHC22, WW20b, ZZ16, ZC22, ZP19, ZW20]. pseudo-bundles [Per19]. pseudo-differential [WW20b]. pseudo-involutory [Bal79c]. pseudo-reflections [– DM79]. pseudo-Schur [BS18b]. pseudo-spectra [SHC22]. Pseudo-Strong [FG02]. pseudo-symmetric [Mir18]. pseudoalgebras [LW20]. pseudocharacteristic [Mer79b]. Pseudoinverse [Bom02, Gut23, SLW21]. pseudoinvolution [IM18]. pseudospectra [ABJ16]. pseudospectral [KK22c]. Psi [Pat93]. Psi-functions [Pat93]. pure [CCL10, EJ91, XXF22]. purely [Bar07, Ng22]. Putnam [Ras17]. Pythagoras [Haw05]. Pythagorean [YLL90].

Q [MB22a]. Q-property [MB22a]. QR [KV22, LYS15]. QR-algorithm [KV22]. Quadratic [BSO17, PM23, Rob77b, Adh13, Bae84, Bre18, Che92, CZ22, DDA20, Dra18a, Edw80, Fei75, Ful76, Ger82, GSZ22, IK20, IK22, KML95, Lee18, Lew97, LPT11, LWC22, ME18b, Pap13, dSP22, PUÖ15, Raï19, Rus16, SH21, Shr90, Sol80, SPM14, Tau81, TS05, Tia11b, UÖÖ15, UPÖ16, Ves80, WY21, YZC21]. quadraticity [PUÖ15, UÖÖ15, UPÖ16]. Quadric [BFM22]. quadrilateral [HL18]. Qualitative [JvdD91]. quantifying [Aud15]. quantitative [Fie73, IT16]. Quantized [JP15a]. Quantum [KLZ10, MS16b, Pal21, Aln11, AG18, BLX18, CKL^{+}21, CGH15, CHFC19, Da 94, DHKSH08, Drn13, FHLS13, JP15a, Joh14, KMP^{+}19, KPL^{+}09, KP14b, LCWW22, ShDHG08, Ter22a, Zha14a]. quartic [CN21]. quartics [AC23]. Quasi [And96, aCC17, dSP16b, dlP20, GCSC20, Gol06, LBV13, MCL20, MP18, Sar20, TLC16, Tom19]. quasi- [MP18, Sar20]. quasi-Banach [Tom19]. quasi-coprime [TLC16]. Quasi-Euclidean [dlP20]. quasi-filiform [LBV13]. quasi-Fredholm [GCSC20]. Quasi-orders [And96]. Quasi-range-compatible [dSP16b]. quasi-strongly [Gol06]. Quasi-Whittaker [aCC17]. Quasidirect [LŠ96a, LŠ96b]. Quasilinear [IK20]. Quasimonotonic [Gol90]. quasinilpotent [LL18]. Quasinormality [Ras17]. quaternary [BX11, BSO17, Bus23]. quaternion [Abo20, AAK16, CKXX23, Din15a, FWZ11, FNW18, HW13, HW14, He21, HZ12, KL21b, Kyr11, Kyr21, STZ94, Tho87a, Tho89, WXD21, WLZ22, YL11, YLW15, YW16, ZWLZ22, ZLZ^{+}22]. Quaternionic [haYsS99, AAB17, AJL22, AY84, AY95, Bal23, CDM22, CCL10, EÖ15, Fen10, MVPS14, MVPSTT23, MBB22, Rod08, SLH20, ST96]. quaternions [CC22a, CT22, KT19, Kri87, LK21, LK23]. question [DHHK23, GR14, KKS21, KNX93, KNX95, Sin80, Zha95]. questions [BHM76, CKLP13, dF11]. queue [GT08]. quintic [BP18]. Quintuple [Sto79]. quiver [Sch22]. quivers [Sch22]. quotient [MVPSTT23, ZY13]. quotients [BdPNdP17].

R [Ano79b, Fie84, Zha95]. R. [Don91]. Racah [WHG17]. radial [HX21, Li86, Tam86]. radical [COK22, dHS23, MV75, Shc21]. radicals [GLOT13, Kol23]. Radii [Gol07, LR04, Abo20, AO97, BKS21, BHK21, CC14, CL14, CLMR09, FNBR96, GWW15, Gol79, GS19, HLB22, KL21a, Lei91, Lei95a, Lei98, LT89b, LT92c, LMR94b, Li94, LZ01, LDL17, LWL21, LW22a, LL22b, MS82b, MS85b, Mel20, Mel22, Sab19, SS23, SKLS21, Tam98, YHS13, ZXL22]. Radius [Fla74b, LŠ02, PT02, ADJAH22, AL21, Alo21, Alo23, ALP94, BDK09, BDK21, BBP21, BP22, BM17b, BM18b, BBZ15, CCS16, CR22a, Cha92, Cha93a, CW17, CHQ18, CZ21, CH22a, Che22b, CFLL23, CF23, Cos14, DLSS18, DW23, DM14, DIF13, EAM19, FTZ08, FGHL22, FMZ22, FLZ07, FYIS13, Fri93, GB22, GN20, GL11, GTZ75, GB11, GLS16, HH16, HLW^{+}21, HZX22, JS22, JWBB22, JL88, JSG94, KOSS23, Koz17, KK22c, LP07, LTT84, Li87c, LT88a, LT88c, LT88b, LŠS01, LT09, LT11, LT12a, LC14, LLP18, LS12, LS13, LL15, LZ17a, LZL18, LZ19c, LGZ20, LL10b, LLGX22, LY22d, LCG^{+}23, MGKA23, Man91, MO94, MP96, MS21b, NP13, NG15, Obo23, Oku93, OST20, OM21, Pep11, Pep19a, Pep19b, QHB22, RSM21, RSM22, SM21, SRS21, SYZL17, SHS17, SKLS20, SZWB16, yT86, TJ11, Ton87, WH11, WZ18, WWW20, WMFW21]. radius [Wan22b, WY20, WDJ10, XSW21, XSW22, XW19, XZL14, XYH22, YSY17, YS23, ZMCN19, ZQC18, ZHH09, ZG11, ZL19, ZW17a, dlP09]. Rado [GX17b, Woo22]. Radon [AYC93, GP74, GP22b, HJ20b, Lee98]. raising [GR81]. Randić [Yin21]. Random [Tan13b, CMR21, CLX22, Che18, Din15a, GLLM21, HL18, LCL22, Rau18, Sha15, TXL21, TJ16, WS23, Zen22, ZZ21]. random-effects [TJ16]. randomised [ZWZ23]. Range [AMP02, BK83, BD94, CN04, DvE20b, Far02, GW04, KR07, LN98, Sed04, haYsS99, AAK16, AYT83, AY84, AY95, BM22a, BJS98, Bal80, BB15a, BD88, BLD91, BL92, BD96, BdPN14, BBK22, CRW21, CJ97, CDM22, Cha18, CLP22, Che94, CL94, CT96, CT09, CL13, Chi98, CNP05, CNP06, CN07, Chi11, CN14, CGL^{+}16, CH10, CKP09, Cro08, DvE20a, DHH22, EM15, FT87, FMX18, Far93a, GW98b, GW08, GS21b, Gro96, HHSY22, HH16, How80, Hu89, HT91a, Hug90, IOMA21, IOMAN21, Jam88, Jam95, JLL07, JC22b, Jon92, JR93c, JR94, KV23, LLS14, Lei92, LT92a, Lei93b, Lei95b, LT88a, LT88c, LT91, LMR94a, LŠS01, LPS09, LT09, LD17, LWC22, LP16, LTT11, LP22c, LP22b, Man87, Man91, Mar75, MF79, MS79, MW80, MS82a, MS87, MBB22, MS15b, MS16b, NNT95, Nak95, dSP16b, Poo80b]. range [Poo94, PV00, Rod08, RW21, RS08b, ShDHG08, ST80, STZ94, ST96, Spi00, Tam85, Tam88, yT00, TS19, Tho80, Tho87b, Tsi83, WZD16, Wes75, Woe08, WC12, YC19, Yua98, bPFZH08, LLS16]. Ranges [BLdPS04, Ber02, BS04, LS04, LR04, Loe74, Abo20, AM21, AyT84, BdP98, BdP12, CKL^{+}21, CGW14, CL17b, Che14b, Che16, Che17, CN05, CN08, CN15, CLN21, CL00, CGHK08, CLMR09, Da 94, DHKSH08, ERS12, FHL^{+}14, FL16b, GW12, GWW16, GS21a, Gol79, HT91b, HL21, IOMAN21, JNBR93, KL21a, KPL^{+}09, KP14b, Lei93a, Lei98, LT92c, Li94, Li96, Li98, LPT11, LD19, Mag13, Mag14, Man92, MP87a, MR97, MP98, Sza09, Tam94, TGW14, Tsi81, YLC19a, YLC19b, vD96]. Rank [BT11, BPS00, BL97, CM15, Cho10, CL01, Dod13, EL90, Fer09, Gal01, GVS96, HZS22, JT03, KP14a, Mar13, MF92, OW02, Pat03a, ST97a, So99, SLH16, Tia01, Tia02, Wal79b, AM21, AW83, BHD10, Bal14, Bal18, BI95, BOvdD91, BL01, BSL01, BS16, BS22b, BdP12, Ben90, Bén11, BHH^{+}11, Bot87, CKL^{+}21, CFSGL18, Cat21, CCL21, CCS98, CGW14, CT16, CHL18, Che14b, CLN21, CGHK08, CLN13, CK19, CKT20, CL21, CLMR09, CP18, CP21, CO12, Dec95, Der16, Der18, DM18, DM19, DS22a, DMQV23, DR18, DH16, FT21, FW13, FOT23, Fie81, Fra17, Ful76, FS89, GRMA13, Gon17, Gro99, GW21, GLM^{+}22, Gur79, GL83, Hac22, Hai88a, HHSY22, Har81b, HHY22, HKv23, HS87, HLZ05, HKS94, JW91, KS12, KP14b, KRR18, KS17b, Kuz05, Laf78, LL16a, Lau81]. rank [LPZ13, LPS09, LGA^{+}13, LZX19, LY22b, Lim90a, Lim92a, LT14, Loe93a, Loe93b, Loe01, Lon80, Lon81, MS16a, Mar11, MR18, MMRR16, Min77, MSS18, MS15b, MS16b, NS15, Ned98, Ore19, Orl18, dSP16a, Pet09, PV20, Rub20, SdF12, Šem97, SMBB^{+}15, SM22b, Shi15a, Shi16, SMF95, Sil88c, Sil88b, SS12, SBMP16, SKP18, SL21, SF17, Tia05a, Tia05b, TWZ16, Tia17, Ven19a, Wan79, Wan15, WM17, Wan19a, WW20a, Wat85b, WLM21, Wes80, Wes81a, Wes87, Wes90, Woe93, Woe08, Wu09, WW12, XL17, YLX20, Yan21, ZS05, Zha05, ZY17, ZXF15, Zho16, SLH18]. Rank- [CM15, KP14a, BL01, Che14b, CL21, Fra17, Loe93a, Loe93b, SF17, XL17]. rank-1 [YLX20]. rank-2 [Hac22]. rank-additivity [Har81b]. Rank-one [Cho10, CL01, BHD10, Cat21, CCL21, DH16, FT21, MMRR16, Ore19, SKP18, ZS05]. Rank-preserving [EL90]. rank-subtractivity [Zha05]. rank-sum-maximum [Zha05]. rank-sum-minimum [Zha05]. rank-two [CLN13]. Ranked [PM78, BC91]. Ranking [VDW85]. Ranks [HS93a, LPS08, MS74, SK03, Tia03, AC78, BB13, Bal14, BB19a, BS92, BS17c, Bot94, FT21, FKR93, HK79, Haw05, KJ13, LGA^{+}13, Lim15, MSS18, MRS10, Pat87, SMS15, SMS21, Tia11b, Tia18]. rate [Doo16]. ratio [LH23]. Rational [AMP02, Afr73, JT23, AMZ13, Bri89, CM85, DA22, DMQV23, Fei75, dHS23, KLL22, LR96, MZ04, Ota13, RR91, Sil84, Tre85, Tre86, TSYB82, Uhl79]. Rationality [Fri90a]. rationally [DH91, Tre87]. rationals [MV75, New93]. ratios [Sra20]. ray [BGZW16, IK20, IK22, LCL22, ZHLZ13]. Rayleigh [AB78, BdPNdP17, MVPSTT23]. rays [Pat03b]. Re [YZL22]. Re-nnd [YZL22]. Re-pd [YZL22]. reachability [NS85, NST91]. Real [ABC18, Joh81b, LP02, LHS^{+}21, Slo13c, SB03, Ven19a, Wes81a, haYsS99, AF11a, AF11b, AF12, Adk95, AS12, AGZ00, Bal14, BS92, BB75, BPR20, CS21b, CO12, CIDK07, DH86, Deu80, DIJ23, – Dok83, Dru15, ER22, EJ91, Esc93, Fan73, GP14, GW17, Gon17, Gue19, GHK16, Ha89, HW13, HW14, HS88, HLZ11, Jin17, JKB17, Joh84, KL22, KML95, sL19b, LS16, LM93, MMRR16, Mes90, Mir09, Moh16a, NR23, NP91, ORT15, Ota13, Pal79, Pas15, Per17, PM23, RŠ06, RŠ09, RBB22, Rom89, SU77, SL15, SMS15, SMS21, SK15b, TT12, Tho75b, Tie16, TNJ22, Wes80, Wes81b, Wu22, Zha01a]. real-positive [CIDK07]. realizability [JMP17, PM23]. realizable [LS10]. realization [APEPPS23, JMS15]. Realizations [NU94, AM90]. realizing [CJK^{+}13]. Realteil [SU77]. rearrangements [TI77]. reas [KLZ96]. Reciprocal [AH16b, AHJ20, AHJ21, BNP06, BGM22, BMP22, HA22, HINS10, PSB23, MT22b]. recognising [Gre99]. recognize [Han21]. reconstruction [FM21, SS07]. rectangles [Lor14]. rectangular [AA21a, BCR91, Car19, CKP09, Cra91, HZ20, LT88d, MHH96, Mis17, MSS19, ZS05, ZL18, Zha19]. recurrence [CR22b, Kom10]. recurrences [PSW18]. recursion [BVW84]. recursive [TZ21]. reduced [Car06, RA21, YTL20]. reduces [For79a]. reducibility [OR09]. reducible [Cav10, CNP06, CN07, GSS09, Ikr15b, Stu89]. Reducing [OW02]. Reduction [BZ22b, BT90, CS18, FN14, Gal01, GRMA13, HH85]. reductive [Koh11]. redundancy [Far22]. Referees [Ano11b, Ano10]. Refined [KK13, WW20a, BM15a, lGyW99, Sha20]. Refinement [Mel20, YLC19a, You15]. Refinements [Alo21, Bak16, Dra11, LS15, TGW14, Dra19b, SG18, YK22]. Refining [Bap88]. Reflecting [AFVL00]. Reflection [LM03, Ell93]. reflections [– DM79, EM87, Ell95]. Reflexive [Mia93b, AH13, CIN15, Had83, Haj18, Qi15]. réflexivité [Del94]. Reflexivity [Hos22, CHK00, Del94]. regarding [DJ89a]. regional [JLR^{+}92]. Regions [LS04, LL14a, MH20, Ano22a]. Regular [CZZH22, KS15, Peñ02, AG16, BCEM12, CP81, CLW15, CC23b, CP18, DFK19, DP17, DGR08, EB17, FG06, Gol06, HS95, KMS16, Koo95, LR13, LM10, MS10c, Mou20, NS99, PH15, PV23, Ram22, Rod03, SR22, Sta10, Sta20, TW17, Wri22a, Zho16]. Regularity [FG02, BV16a, CFGM06, HRT15, RR95, Ska17]. regularization [GY22]. regularized [DPP17]. Rejection [Whi74]. related [AG22, ABAH23, AkK17, Ala16, AGR91, AV22, CT22, Car90, CKLP13, CKMY22, CN14, Cho18b, CL01, CR22b, CKKY23, DHHK23, DP18, For79a, Gow22, GSZ22, GX17a, Hon92, KM22, KLJ23, LPT13, LT87, LFY16, MGAH19, McD92a, Mcf79a, Obo22, SMBB^{+}15, SHC22, Sin84b, Sto79, SK15a, Tia01, Tra84, WZ14, XQ15, YGÇ19, ZW19, ZBW12, Zou20]. relating [JL88, Li91, Pat82]. Relation [CD03, Wan19a, WL20b, ZH22, AV22, AB22, Bén10, Byt20, CHL18, CR22b, Dji16, DK16, DHH22, DH22, Fie90b, LZX19, LY22b, Lim79b, MW16, Wan16a]. Relations [DS02b, BCM18, CC92, Che14a, Chu22, GZ76, JR93c, JR94, Kom10, KD23, LST89, LT92d, LD22, PT87, Put82a, San18, SXR18, Sim83, Tam91, XS18]. relations* [JC22b]. Relationship [WW12, DMV18, Gut23]. relationships [SPS15]. Relative [DHKSH08, Dom00, For87, Fur11, Hol04, MCIM21, MN12, MF20b, ZBW16, ZLS22]. relatively [Dre93, JKD16, LS22c, XS18]. relaxation [CLW21]. relaxed [Szw18]. relevant [Tik98]. Remark [DLD21, KL19, CH10, Fan88, Fri90b, GT79, Hul81, Mar82b]. remarkable [JLL88]. Remarks [BSM98, For79c, KKL22, LW11b, Ng22, WZZ22, dF13, Gro97, Odo83, YT15]. Remarques [Rev81]. removal [SS22]. Rényi [DDF22a]. Replacing [dCdSA22]. report [LRT08]. representable [CG22]. Representation [Adu15, JT04, Tur02, DMV18, Gau76, HCS19, IAM23, RRN23, Sch22, SX12, Wes81b, YW09]. Representations [BZZ11, Che93a, DD10, Don91, Elg22, FOT23, Ful76, GCP20, KO10, Kra06, LCC19, Mer82, SKM17, SCL^{+}22, XSH18, ZFSX17, Abd09, jAS15, BS18a, BP18, Byt20, CR15, CV94, CICX09, – DDJ76, GZZF23, Heo16, Hun83, Hun86, KCSĆ21, Kut19, Kyr11, Kyr21, LLMN20, LL20, Lip87, LYZW12, ML21a, Ma22a, MPI17, MR13, SQM99, Sei95, Til77, WL16b, ZT12]. representative [MN96, WW80]. representing [RD12]. Reproducing [BR04, GGGA21, YGÇ19]. require [BJS98, EJ91, Esc93, KMT96]. requirements [JS91]. Research [Ano82f, Ano82g, Ano83m, Ano84e, Ano84f, Ano85g, Ano86e, BC05, Bap92, Bar88, Cal95b, CLLT93, De 84, EJ88, Fri88, Gra93, GM87b, JC74, Joh89, KPL^{+}09, LMMR01b, Mer91, RP04, Sch91, Sol06, TT85, Tau73b, Tho82a, Tho85, Tho87b, Tho91, WW01, Zha92, Ano97b, LW96, Mer79a, Zha00]. Reserch [Bru87]. residual [CIP17, Guo20]. residual-based [Guo20]. residue [WY18]. residues [GSZ22]. Resistance [BBG22c, AAB20, RBG23, SWZB15]. resistances [CMM17]. resolution [DS23c, PA94]. Resolvents [Gil16, Gil14]. resonances [Ab23]. respect [AC13, Aln15, Che91, FH73, HHC12, HPV88, JST18, KMMS22, LT89a, MO94, Ner22, Oku93, Pop17]. Restricted [MPGS21, Pap13, Bal79a, CRW21, CC19, Che93b, Che93a, CP16, LOO22, Ned98, SG96, SPK15, SPKC15, Wer84]. restricted-rank [Ned98]. restriction [DS22a, Wan15]. restrictions [BZ99, Bou18, Far94, MT22a]. result [AkK17, BB17a, Das21, Har84b, JLL11, New74, New82a, Poo80a, Ric87b, Tam91, ZZY16]. resultant [Bar80, BL21, GKLR82, Orz84, XW22a]. resulting [MP87a]. Results [Gór22, JKZ20, LN98, Tau74, TH80, AG22, AEAAES20, AAK16, AN14, Amm18, Ano22c, AZ22, BL83, BHK78, BM17a, BLZ11, Boa19, BZZ10, BSZW13, CL13, De 82b, DCIW10, DW12a, DHT94, DIF13, FM22b, Gar23, GHL10, GAMA23, GX17a, Guz88, He21, KDHJ22, KCICV18, Li84, LT87, LZT18, MR97, Mou13, NS85, New75, Nie10, PS22, SMBB^{+}15, SWH22, Sin82b, Sra20, SWZB15, Tam77, WC15, mWyD11, Wer87, YGÇ19, YWZ21, ZCP16]. Retarded [DGR08]. retrievable [FM21, HLY22]. reversal [Kir93]. Reverse [CIN15, CIM19, KM11, LW08, Obo23, PBM20, Rad16, ZCP17, BS18b, CID16, DD14, HKO16, LFY16, Ota20, RW84, Sha20, WZ14, XQ15, XQ16b, ZX10]. Reverse-order [PBM20]. Reversed [DL14, Bak16]. Reverses [BM15b]. Review [Fan81, Kap80, Lei81, Tho83, Yaq80, AK78, Ano83a, Ano89a, BFMM73, Con94, Fan84, HS81, Hum83, Joh83, Keg87, MT83, Mer98a, Pie92a, Sch92a, Sha95, Sko85, Sta94, Tho82b, Tho82c, Tho82d, Tho82f, Wat84a]. Reviewers [Ano12b]. Reviews [Ake80, AR80, CST75, DW76, Doz79, EA77, FAR78, HUSG74, JM77, MB76, MS80, MBSL74, MGK^{+}74, RY73, SB75, Sen81, SA77, TB76, Tho81a, Tho81b, Tho82e, WNM74, CS79, GMS^{+}76, JTH77, LW78, PA79, WTM78, ZF78]. revisited [CBQW19, CC23b, DD14, Ere15, FMSS19, FHL^{+}13, Joh90, JZ95, LZS23, Per17, SRT22, Vel18, Wan19b, dF19]. Revisiting [Lip87]. Ribbon [FJK14]. Riccati [CK84, DLL19, FN14, Igl99, LRR87, MWC22, NF19]. Richardson [AW13, AM90, BG90]. Richman [BHM76]. Rickart [Cir15, Cir16, Cir17, CC22b, Kre16, Mar15, MRD15]. Riemann [CN21, Red02]. Riesz [BZ22a, BCM18, CDJ13, CG22, DMV18, FJS14, HSM17, HFSZ22, ZQZ21, ZZC17]. Riesz-representable [CG22]. Riesz-type [BZ22a]. Right [CC21, Ake83, AJL22, Bal80, – Dok73, HBM15, MRD15, MW22, MBB22, Pat89, Tho86a, WHC22]. right-star [MRD15]. rigid [BCS18, COK20, CSG22]. Rigidity [BCS18]. Rigidity-preserving [BCS18]. ring [Bas17, BH15a, – DL92, – Dok93, For79d, HWZ18a, Hua10, HPV84, HPV88, KCSĆ21, LPT13, LCL13, Lew90, LQ95, MW16, Mil22, OWX09, OWTZ21, Piz87, ST22, Slo15b, SZ19, SX12, TZS19, TW17, WOY06, WZL13, Wri22b, XWT22, YWO07, ZWM17a, ZWM17b, ZC22, Zhu18, ZP19, ZCCIW12]. Rings [GMR81, RR04, Rao81, Zho22, AH10, ÁBvW16, AW13, AELL20, AA20, BY19, Bas17, Bas19, BD22a, BLR12, BCI13, BB17b, BDvW12, BDvW16, Bre21, CL22a, CS18, CQ22, CG19, Cir15, Cir16, Cir17, CC22b, Coa99, CZ22, DPvW13, DS13, DD15, Dre22, DJS21, Ere15, FJK14, Fog73, FSvW11, Gar21, GV89, Gus81, Haw05, HLZ18, Hou21, HLH18, HZ20, Jam88, Jam95, JL12, KW21, KK22a, Khr16, KL12, KR07, Kre16, LPT14, LWZ15, LC18a, LZ23, LL10a, Lin21, LCL14, Mar15, MRD15, Mar18, MP13a, MP87b, MCIM21, McD92a, Mos21b, New85a, PA10, Pre22, QZ16, QH17, RM22, RŠ08a, SM19, ST97b, Tan20b, TZ17, Tau73a, Vla23, Wan16b, Wan19b, WM21, Wat85b, WY18, YLW22, YW05, YQ90, ZWZ10, ZCZT22, ZY08, ZCP16, Zhu19, ZW20, ZWM23, ZCP18, DDA20]. Riordan [MB22b, PS21]. Ritz [RW21]. Robert [Ano97b, Day97, GPR84, New97]. Roberts [FMZ22]. Robertson [JZ95, Joh90]. Robinson [Foa79]. robust [Ano16c, Ous15b, Sol14]. robustness [AT18]. Rohn [Rub20]. role [MN96]. Root [Hai92b, BAN12, CGH15, EJD88, FZ11, Hai92a, Lia23, Mel13b, Mey15, vBW95]. rooted [MW23, Muk20, RP22]. roots [AAT16, CL74, HHC12, IK19, JN80, KN81, NMAB22, Nom73, Ota13, Tie16, Uhl99]. rootsystems [Hai88a, Hai88b, Hai97]. rose [MH16]. Rota [FW73, Gar73, GG22b, KKL20, LW20, ML16, ZLS22]. Rota-type [KKL20]. rotation [GTM14]. rotational [Sjo93]. Rotations [Edw80, Edw78]. Rotfel’d [FL16b, HWY17, HWZ18b, YLC19a, Zha15b, ZN18]. Roth [FW13]. round [Spi00]. Row [Joh81a, JP17, SS12, CPKK22, FV97, Har81a, Kap77, LSOvdD95, MH19, MH12]. row-adjusted [MH12]. row-sums [MH19]. rows [MNS22]. RS [GTLN18]. RSVD [LT10]. rule [BG90, CRY95, Che93b, CPT15, Wer84, WY18]. rules [Che93a, Kyr08, Kyr11, MS10a]. Rutishauser [TT15]. Ryser [MS85a, New83, PS22].

S [Ano16b]. S. [DF82]. Sachs [AAT17]. saddle [CL22b, MHM22, Zha20a]. Safety [GHV99]. same [BCCL22, CW90, Do 96, Fer10b, JGK79]. Sampling [CKXX23, BG23a, KKP13, SR17]. Samuel [Yu14]. Sandy [Tam86]. Santa [Ano79c]. satisfied [Wan82]. satisfy [McD92b]. satisfying [CW11b, CW12, Düf97, DKK19, KL22, Pet09]. Scaffoldings [Cas23]. Scalar [Est80, AF12, Bre15, GS21b, PY82, Rod81]. Scaled [BKL02]. Scaling [JR09, FLS88, HS85, Sha91]. Scalings [HS85, SS79]. Scarpis [– Dok17]. scattering [MS10c]. Schatten [FNBR96, So90, YLC19b]. scheme [HZV23]. Schmidt [Bah22, LF15]. Schneider [BHM76]. Schröder [Dag23a, Dag23b, RS18]. Schrödinger [HMK22, Ab23, WT23]. Schultz [DK23]. Schur [AS20a, AO97, Bap88, BT06, BS18b, Bun91, CRY95, CPLH17, Dia79, DS02b, Dra21b, ES98b, ES00, FJ15, Gór22, GMZ11, Gur83, HS75, Hua23, JK10, Kim91, LM98, LHZ22, Liv94, Mas10, MO94, Neu81, NT86, PN22, Rem90, SV14, Smi83, Tam02, WZ97, Wat86, Wu22, Zha04a, ZJC^{+}22]. Schur-concave [Tam02]. Schwarz [Dra11, Dra19a, Dra19b, IK22, Oku91, Ota20]. science [FHLS13]. score [Kir91]. scrambling [GM18, GM21]. SDB} [JLL18]. SDB}-tensors [JLL18]. SDD [LHGL22]. search [CSCH01]. Searching [FK91]. secant [BB19a, Bal22b]. second [Ano79c, BPW16, Cam83, Das21, DTSI15, DdF15, Gro85, HSX16, Hil07b, Jam81, LP00, LLF08, Lim90a, LXG17, MB22a, Sim95, Sta10, SKLS21, VY21, Wal11, XZ16, XSL15, YZ20, YCZ21]. second-order [DTSI15, HSX16, Hil07b, MB22a, XSL15, YCZ21]. sector [FL16b, HWY17, LF21, XZ22, YLC19a, YLC19b, ZJC^{+}22]. sectorial [Ala20]. seeds [JP15a]. Segment [Vo83c, Vo83a]. segments [GW08]. Segre [BB19a, Bal22b, BB13, FU20]. Seidel [AAD22, TLC22]. Selberg [SN23]. selections [LP16]. Self [BL83, OP13, BP19, Din15a, FS14, FS19, Gil17, HH16, HX21, Kit91b, LLMT05, Sin84a, Tan13b, WW20b, XS18]. self-adjoint [HH16, HX21, LLMT05, Tan13b, XS18]. self-adjointness [FS19, Kit91b, WW20b]. self-commutators [Gil17]. self-complementary [FS14]. Self-dual [OP13, Din15a]. Self-duality [BL83]. self-inverse [BP19]. Selfadjoint [JR93a, Dra10, Dra11, ELG88, GLR82, GR93c, JR93b, KML95, KLZ96, MP98, Rod08, TZ19a]. selfadjointness [DK22]. Semi [Knu85, AFM13, Amm18, Are23, BFM20, BN22, BCM18, Cha90a, Cha92, Cha93a, CH10, Chu84a, CF23, DGJ^{+}16, FG22, GCSC20, GH23, Gol89, GBM21, Har78b, HPV84, HPV88, dJfLlZ^{+}21, JP87, LL16b, fLLL^{+}17, LL14b, Liu16a, LSX19, LCWW22, MSS13, MŠ09, MO13, NMAB22, Pat92, QHB22, RC18, RSM21, SM23, SMS21, WC22a, Yu14]. semi- [LL16b]. semi-Browder [AFM13, BCM18]. semi-Cayley [Are23]. semi-definite [Cha92, Cha93a, FG22, Har78b, JP87, LL14b, LSX19, NMAB22, Pat92]. semi-definiteness [Cha90a]. semi-dihedral [LCWW22, WC22a]. semi-direct [RC18]. semi-finite [Liu16a]. semi-Fredholm [Amm18, MŠ09, MO13, Yu14]. semi-Hilbert [CF23, GBM21]. semi-Hilbertian [BFM20, BN22, MSS13, QHB22, RSM21, SM23]. semi-hyponormal [CH10]. semi-monotonic [Gol89]. Semi-optimal [Knu85]. semi-positive [DGJ^{+}16]. semi-simple [Chu84a, HPV84, HPV88]. semi-tall [SMS21]. semi-tensor [dJfLlZ^{+}21, fLLL^{+}17]. semi-trees [GH23]. semibounded [Uhl82]. semicirculant [Mou22]. semidefinite [ANK22, BHH^{+}11, Caf13, CS21b, DM13, DJ89b, FLT22, GLP92, HS88, HNS99, HM95, KRR18, LW12, Liu22b, MM06, Mat92, NS15, PV20, SU77, SM94, TZ19a, Tra84]. Semidefinitem [SU77]. semidirect [Che95b]. semifinite [BZ19]. semigroup [RS07, RS74, cc85]. Semigroups [BC78, AAJ12a, Ber05, CS10, DW92, GT14, LMR00, MP13a, OR09, PY82, RD12, ZCP16]. semilattices [Luq08, WHL16]. semilinear [Pan13, Pan15, ySpW14]. Semimodules [CL22a, LT14, ySpW22, Tan16]. Semipositivity [JKS94, Wer94]. semiprime [DS13, DD15]. Semiregular [Pow88, DGR08]. semiring [KW11]. semirings [AAJ12a, BL93a, BGLS07, BSS12, CL22a, DRY22, DO20, lLpW16, Moh15, MP16b, Shi15b, ySpW14, ySpW17, ySpWQ21, Tan13a, Tan14, Tan18a, Tan18b, Tan20a, Vla22]. Semisimple [AKOZ20, DR18, FW13]. semisimplity [MHA17]. Semitransitive [RT09, BM14]. Sensitivity [FS22, LZ19a, jXL15]. Separability [DNY22, Joh11]. separable [Lud21b, Raj09]. Separating [Kan94, ST22]. separation [KN81]. separator [HRMP95]. sepr [HLOvdD20]. sepr-sets [HLOvdD20]. sequence [BT15, BM89, Car88, CFZ23, Dag23a, DTS14, DV19, ESAJ18, Har84b, HSM17, Hua01, HLB22, KI16, Kar23, MR18, MM20b, Ogu20, PS20, RF19a, TD14, TD16, Tal17, TD19, Tom19, TRM22]. Sequences [Rei77, AADdF23, ACJ17, AM90, BG23a, BB19c, BS15b, – Dok81, Gro95, HH22b, JVMCPR19, KL13, MR17, MGAH19, MS10b, SLUG04, Zar01]. sequential [CW91]. Sergeev [BG90]. series [Bed10, Bed11, BX11, BB14, Car90, Chu22, DST91, DF82, – DL92, – Dok93, Dre22, For79a, SES21, SAK22, Spi76, Yuz84]. Sesquilinear [Lee15, LMR94a, Nak95, RB22]. Set [BS97, FG02, AFJK17, ASAS79, BP98a, CCS16, CFSRL22, DM22c, FGMSS19, Fra17, HM77, LT11, LT12a, LL16b, LZL16, LRD21, Mel10, Pat07, TJ11, Wan79, WL18a, YZB22, YMYS13, Zel85]. Sets [DMdO92, EMS02, HHSY22, AL22, BT90, BC85, Bén10, BDK^{+}05, BMP10, Bra86, BS08, CP13, CAP22, ES00, EV77, FTA20, Fri84, GS06, GZ76, HS91, HLOvdD20, Hos22, JT95, KOvDD09, Law19, LXQ21, Lei18, LT88a, Li92, LP93, LLL15, New85a, Pep19a, Pie92c, Prz92, RW22, SL19a, SRW07, ySpWQ21, SSZ93, Soa08, Sza09, TN22, Wan22a, Zha14b, ZL18, Zha19, vD96]. setting [FW73, GKT97, Suc22]. Seven [NW02, Hai88b, Wes80, Wes81a]. seventh [Gar21]. Several [Orz84, ZHLZ13, Zhu18, CF23, De 75, GJR97, GHL20b, dHS23, Paz84, dO75]. SGB} [JZL20]. SGB}-tensors [JZL20]. shape [Hug90, Jon92, KHG13, LH13, Tsi81, WBZ14]. shaped [CT96]. Sharp [JRMFMASS16, LS12, LS13, Luo19, Rho94, Rum97, YSY17, FMS19, Gil17, LZ17a, Mit92]. sharpness [KMN01]. shell [CNP06]. Shemesh [JP15b]. Sherman [SC20]. Shift [BR04, BG86, CNUV16, Tam94, TGW14, WL22]. shifted [RRC23]. shifts [CB22, GH23, GHL20b, KY17]. Shoda [BS18c]. Shoda-completion [BS18c]. Short [MT83, Tho81a, Tho81b, Tho82e, Tho82b, Tho82c, Tho82d, Tho83, Tho82f, AY95, Bar12, KP11, Roi99]. Shorted [CGM19]. shuffle [Lew96]. sided [EÖ15, Kyr21, Tur16, WM21, ZWW22, ZWM23]. Sigma [LY22a]. Sign [BJS98, EJ91, Esc93, FHL^{+}13, JWP98, KS02, KMT96, Kir11, LGA^{+}13, MOTvdD03, Peñ02, Reg87b, AG16, Cai93, Car06, GS06, GSS09, Hem93, HS93a, HLOvdD20, JvdD88, JP07, Kim13, LLS10, LLL09, LHbL^{+}15, MM11, Ma15a, MMRR16, Rem88, SR22, TI77, WYM11, WW01, YMYS13, YHCW14, ZBW12]. sign-balanced [WW01]. sign-central [Kim13]. Sign-patterns [KMT96]. sign-regular [SR22]. Sign-Solvable [KS02]. sign-stability [Cai93]. Signal [FM21]. Signature [HR85, GLR82, WT16]. signatures [PT10]. Signed [HLP03, ADHM19, BPW16, BB19b, BBCD21, BP16, Dua22, FDD14, FKC19, GVS96, KS17a, MR18, Ram22, SSG12, SRD19, SR22, Sta22, TI77, TWZ16, Wan19a, WLF22, Wan05]. significance [ShDHG08]. Signless [LLP18, MH16, NC22, WBZ14, ZZ15, ABGP21, AH16a, Ash16, AOTR11, CHQ18, CZ21, CFLL23, DAH19, Das21, FYIS13, HL22, JS22, LZ11b, LS21a, LL15, LZ15b, LDL17, LW22b, LY22d, LHH18, Obo22, PRAC22, SYZL17, SHS17, TYC21, WH11, WZ13, XZL14, XLS20, ZW17a]. Simić [Ano16b]. Similar [Bou13b, Joh81a, CLNT17, Liu18b, Lon90, Slo17, Sło18a]. similarities [Bal75, Bal76, CT22, JP17, LT89b, Li91]. Similarity [Bea86, BB75, DS09, GP06, Gur97, New82b, New85b, RJ04, Wat84b, BGdH92, BG82, CR77, DHB07, DKK19, FW13, GB77, GHV99, GL97, Gus81, HX21, KL22, KT19, Kur04, LT89a, LT90c, LT92c, LM93, MW74a, QC23, Rob75, SS22, Sil88c, Sil90, Sil93, SS96, Sta23, Tia11c, Wat82]. Similarity-preserving [RJ04]. similitude [Vin06]. Simon [Raw81]. Simple [Mag22, MP87b, ALPM20, BHH^{+}11, BZ22b, Che07, CW11b, CX12, CY22b, Chu84a, GT11, HPV84, HPV88, LQ95, Lip96, RZ17, Tan18c, WY11]. simplexes [ST86]. simplic [Rad80]. simplices [BS08, Fie95]. simplicial [Tam77]. simplicity [Mar00]. Simultaneous [AY74, BT90, Gow22, Grü76, Laf78, Bed10, HH85, LLLY23, MS10b, Uhl79]. Single [Mer83b, DS22a, – Dok09, JP87, JNS23, Mel13a]. Single-hook [Mer83b, JP87]. Singular [ANK22, BI95, GHT11, Hil17, LT92b, MS85b, Mir81, NS99, RS17, Rum97, ST00, Ter22a, ZL18, AG22, AJL14, AHJ87, Bap87, Bea85, Bot94, BC94, BZZ11, CS23, CQW16, CS19, CL22b, CH87, CZ22, DLL19, DdF15, Düf97, Fra13, GA18b, GAMA23, GRSW92, GL16, Gut22, HRT15, Hua22b, JW22, JN90b, Kar83, KS09, sL19b, Lék19, Li84, LT90a, LT93, LP93, Li95, Li21b, Mat92, MS12, Nie10, NF19, RS07, Ric78, Sad23, SRT22, SDL10, Tho75a, Tho75b, YZ16, Zha00, Zha20a, Zha19, Zou20]. singularity [BL21, LB19, Shi15b]. singularly [CDG23]. Sinkhorn [Car19]. Six [Min74c, DZ22, Hua01, LHS^{+}21]. six-term [Hua01]. sixth [Wan07]. size [CH22b, KS93, KS17a, LS22b, SKLS20, SKLS21, ZXL22]. Skew [CS10, CT16, CW23, AN19, AM13, AADD18, Arm17, AA20, Ben18, BCT23, BE22, BZZ10, CT09, Chu90, DD15, DS23a, DH11, DLSS18, DMS23, ETM22, Fur11, Gog21, GTLN18, Gon17, Kyr11, LW17, LTT11, LCL14, Lop11b, Lop18, MS16a, Mat92, Neu83, Nof15, PM85, Rod97, TKT19, Ter22a, UF22, Zha16a, ZLZ^{+}22, ZQZ21]. skew-adjoint [AA20]. skew-circulant [ZQZ21]. skew-Hamiltonian [Nof15]. skew-Hermitian [AM13, LW17]. Skew-rank [CT16]. Skew-symmetric [CW23, BCT23, DH11, ETM22, Lop11b, Lop18, Mat92, Neu83, Rod97, UF22]. skew-Takagi [Ter22a]. skewsymmetric [LR13]. slice [LKLC22]. slice-stochastic [LKLC22]. Small [Fer10b, ST97a, Bre15, CGH14, – Dok09, Fer09, LLT93, Mes13, Pat83, dSP11, PGLM00, SS12]. smallest [ABP22, CL86a, CL86b, Dec95, LLF08, LHH18, Obo22, PLH01, Sad23]. Smith [Arm17, BK17, Cha96, Der83, GRMA13, HS96, IH11, Koo95, LLC22, LT12b, New94, Pro95, Tho88]. smooth [Rao18]. Smoothness [Sai21, MPD22, MDP22, SPMR20]. Sn* [GGH22]. sole [dSP12]. Solution [JSY14, AHK92, ALP94, BZZ11, Che93b, Che93a, De 13, DM22c, DWL14, FOT23, FS19, Gra80, HKT15, HW14, HBP18, HM18, HM22, JS84, JN14a, MK16, Min80a, SLH20, SPM14, Vei83a, WZH13a, WZH13b, WXD21, Wan22a, WLZ22, YL11, Zha04c, Zha16b, ZL21]. Solutions [DH03, Fla74a, Tia03, YCZ21, Bak84, BBP78, CC17a, CH12, CCL10, CDF21, CIDK07, CID16, DH86, DH11, FHW16, FNW18, HW15, HSX16, HWL21, dJfLlZ^{+}21, KL22, Kyr21, Li85, LD19, Pen15, PM78, Rom22a, Son14, Wan15, WHL22, Wer84, XSL15, XW22c, YZL17, YTL20, YZL22, ZLZ^{+}22]. Solvability [Lev74, fLLL^{+}17, VM18, YqS16, AG15, CIRW21, HWL21, JvdD91, KP11, Lip96, Mil22, Tia00, vWZ19]. Solvable [ACO21, AK21, KS02, BL10, BCS18, CSG22, CLOK13, Far93b, KLO14]. solve [BS17d, HM20, LT10, PLPC19]. Solving [BB19c, ESH23, Kul22, ST97b, CW20, CL22b, KM17, fLLyP15, LZ16, LZ19b, LC18b, LW10, SW22, WX19, XJW17, ZZL23, ZWZ23, ZV19, ZZT23]. Some [AEAAES20, AAK16, AN14, AAD22, AW07, Amm18, AdF21, AyT84, BS15a, BM19a, Bal75, Bal76, BCNG23, BG83, BD94, BZ19, BKL97, BP22, BZZ10, BSZW13, CT22, Cha89, CHFC19, CL94, CF23, DH17, DST91, DCIW10, DV18, DH91, Dra16, Dra18a, Dra22a, EMS02, EB17, Fan92, FM22b, FH15a, GLOT13, Gai77, GA18b, Gro97, GX17a, Han19, HS83c, HS96, He21, Hua08, Hua09, HXL16, HM19, Iop19, JKZ20, Joh82b, KI16, KDHJ22, KMS18, KS84, Lék19, LT88d, Li96, LM98, LN98, LS22b, LS21b, Lu22, MM06, MW80, MR97, Mar92, MH22, MH12, MTR11, MŠ98, MFS22, New75, Odo83, Oku93, PS22, PH15, Pop14, Raï19, Ram22, Ras19, RSM21, SRS21, SWH22, SQH15, SSfW15, SG96, Sim95, Sin82b, SWZB15, SZZY18, Tam77, TJG22, TT78, Tau74, TC04, Tia05b, Wan82, WH12, WWW20]. Some [Wer87, YT15, Zha00, ZN18, ZW19, ZW17b, dF11, AKR23, AE10, Alo21, AAdFS19, AA20, AZ22, BK17, Bel15, BH74, BM18a, BZ89, BSA17, BE22, BD14, Cal94, CGOT13, CX23, CGMR11, CKR21, CB22, CL13, CM15, CL01, CIRW21, DP17, DLD22, Deu80, DHT94, Din15a, – DDJ76, Dra21b, EL09, FD12, Far22, Fen10, FLT20, FH15b, GGGA21, GA18a, GS21b, GS21a, GP15, GT11, GLLM21, GC22, GL97, HE20, HH22a, HW15, Hua16, HLH18, HWZ18b, IAM23, IOMAN21, JKD16, JN22, KMRS89, KN22, KS19a, KSMB16, KS22c, KT06, Kop94, KAE22, KL92, Lei95a, LT87, LT88a, LPS13, LH21a, LD19, LF21, Mah82, MSS13, ML21b, MP87a, Mar89, Mar00, MT11, MM07, Meh84, MOGN18, Mos20, Mou13, Mou20, New80b, Nie10, PN22, PLLL20, PP16, PM78, Reg93, Rho87, Rho90, SRK22]. some [Son14, Sta84, SL17, Sur98, TKT19, TL12, Tra89, Tra90, WX19, Ye11, ZWL15, ZL22, ZLZ20, ZBW12, dC15, dC18, Ano22c]. SOR [ZZL23]. SOR-like [ZZL23]. sorting [CW91]. Soules [Pat89]. Sp [Ber86]. Sp-modules [Ber86]. Space [BLdPS04, ACI15, ACO21, Adh13, Alo21, Alo23, AGR91, AJL22, BDK21, Bea85, BHK21, Byt20, CDJ13, Cla92, CR77, CF23, Dag23a, Das17, Das18, DTS14, DLD21, – Dok83, EB17, Fla75, GGGA21, GJR22, HT12, How80, HFSZ22, IK20, IK22, JVMCPR19, JC22b, JS15, LYZW12, LRD21, MW18, MS88, MM09, MC19, MM20b, MN13, OM21, Pat77, PS20, dSP22, QHB22, Rob75, RBB22, RF19a, RSM21, SPMR20, ST19, SA10, TD14, TRM22, WTD22, WT23, YGÇ19, ZMCN19]. space-fractional [HFSZ22]. space-time [WT23]. Spaces [AW83, BR04, HLZ02, LT92d, NTW23, dSP12, Wes74, Wes87, Zha04b, AF12, ACMS10, Atk84, AY84, BB22, BFM20, BN22, BT13, BS22a, BBT07, Bar07, Bea85, BL97, BdPN14, BZ89, BPR12, Boz97, CMR20, CKS19, Cha79b, CL83, CSS21, CDPS23, CZZH22, CFZ23, Che91, CLN13, Con77, CDF21, CAP22, DH17, De 08, DP09, DV19, DH20, – Dok90, Dra10, Dra11, Dra16, Dra18a, Dra19a, Dra19b, DK22, EDKA07, Edw80, ESAJ18, Erd22, EBB18, FU20, FGMSS19, FLR85, Fri97, FZ21, Gil16, Gra80, GBM21, GX17b, GLL21, HNB22, HSM17, HS85, Hol95, HF18, Hos22, HCS19, Ilk20, IKR12, Jam95, JS20, KC16, KI16, Kar23, KL15, KT83, Kru06, KMEO12, LLMT05, LLHS18, LL20, Lim90a, Loe01, MSS13, Maj22, MDP22, MID17, MZ91, Mer78b, Mes90]. spaces [Mol15, MAT22, Mos19, Ogu20, Ota20, Pan15, Pas15, Pat17, dSP16a, dSP16b, Pet09, RT09, RS85, Rao18, Rod81, RŠ06, RŠ08a, RŠ09, RF19b, RB22, RS23, SM23, Sai21, SRT22, SRBB22, SES21, SZ85, Sel81, Sha18, SN23, TD16, Tal17, TD19, TZ19a, TZ06, Til77, Tom19, Vel18, Ven19b, Wan82, WWS19, Wes90, Wes95, YQ90, ZS05, ZCZ12]. Span [Zha01a]. spanned [Bun94]. spanning [CJY22, DWL20, FGHL22, FTA20]. Spans [How86]. spark [AB19, MO22]. sparse [CW20, DK23, Sol14]. sparsity [SM94, tHK17]. spatial [LMMR15, ZQZ21]. Specht [MMR07]. Special [Abd20, BE20, LPT13, AG16, Bas17, CL91, CL17a, – Dok09, Elg22, GA18a, HS83c, Joh81b, JLL88, KS22c, KL21a, KS76, KV13, LT09, Ma15a, Mal95, New80b, Phu80, Pop14]. specializations [MF77]. specific [Shi18]. Specified [Che03b, Ano79b, JKS79, Kap77, Kir14, PRS03]. specifying [Joh77a]. Spectra [AGH^{+}22, BP16, DKY22, DWCD19, ESAJ18, Gol16, Gol21, JKZ20, LZ15a, LPT11, LM10, LZ14b, MOTvdD03, MGA17, RP22, YW19, AADdF23, APEPPS23, BHC16, Bar08, BS17a, BDS13, BJS15, Ber05, BPR20, CL17c, CTY20, CD15, Cro18, CD85, DFPS23, DK17, Fri82b, Fri12, GR93c, GM90, JKD16, JP17, JMS15, Kra06, LS10, LLD17, LXS21, LS21b, Meh84, Mir18, MA15c, NP15, NP09, Sah21, SES21, SHC22, WZ13, XS18, ZLS11, ZG08]. spectrahedra [KRR18]. Spectral [CBZ22, CKM18, DH22, FJK22, Fla74b, FJ01, GMV23, GN20, Gho19, GS20b, IS15, LN16, LLF18, NG15, OST20, RW22, SS23, SRD19, TZ19a, Tos79, Tur02, WY20, WYSL22, WC12, XQ16a, Yiz02, Yiz03, ZB13, AAS14, AL21, ACMR17, BPD88, BEH19, BD84, BKL97, BE22, BBZ15, CNW21, CCS16, CR22a, CHQ18, CZ21, CH22a, Che22b, CFLL23, Cos14, DFK19, DLSS18, Din15a, DIF13, EAM19, FTZ08, FGHL22, Far94, FLZ07, FYIS13, Fri81, Fri93, GB22, GL11, GTZ75, Gow22, GB11, GLS16, HH22a, HRMP95, HLW^{+}21, HLB22, Jaa22, JS22, JWBB22, Joh77a, JN90a, KOSS23, Kir04, Koz17, LP07, LLW^{+}11, LTT84, Li86, LMR94b, LT11, LT12a, LC14, LLP18, LP23, LS12, LS13, LL15, LZ17a, LZL18, LZ19c, LGZ20, LL10b, LH13, LDL17, LWL21, LW22a, LLGX22, LY22d, LL22b, LCG^{+}23, Lon83, LSZ86, Lu22, MH16]. spectral [MGKA23, MW23, MH22, MP96, MAM15, MA15c, NP13, Obo23, PG22, Pep11, Pep19a, Pep19b, PRS03, PGAB22, SQM99, SWH22, SQH15, SSfW15, SYZL17, SHS17, SR22, SKLS20, SKLS21, SZWB16, Sza09, TJ11, Tan13b, Til98, Ton87, UZ16, WSH^{+}11, WH11, WBZ14, WW15, WZ18, WWW20, WMFW21, Wan22b, WLF22, WW20b, WDJ10, XSW21, XSW22, XW19, XZL14, XYH22, XZW22, Xu22, YSY17, YHS13, YS23, ZXL22, ZQC18, ZG11, ZWL15, ZL19, ZW17a, ZW17b, dPN81, dlP09]. Spectrally [BM12, CM07, GHT16, IS16, MY12, ZHLZ13]. Spectrum [GMZ17, JC22a, KLZ96, LLMT05, PS20, Rod03, Sen83, TN22, Zen22, AN19, AA21a, AJL22, Bal22a, Ben18, BR10, Boa13, BC20, BM15c, CP15, CIP17, DP17, DeV22, Gil12, GT08, HMO11, JX13, KP12, Kra05, KML95, LZB23, MM20b, NC22, PRAC22, QX23, RA21, RAN22, RP90, So99, WW12, Xu93, ZZZ16, dFK21]. spectrums [Sin82b]. Sperner [Can80, Har84a]. spheres [– Dok81]. Spherical [Gam00, KKY22, CLL14, Da 88, DKY22, Gam18]. spin [– Dok81, KLJ23]. spinor [AF12, Ger82]. splines [DTSI15]. Split [KADM10, CC22a, CT22, EÖ15, KL21b, Mar12, Poz22]. split-quaternion [KL21b]. splitting [HZV23, LW17, Pop20b, WZY22, WL22, Xio23, ZV19]. splittings [CP81, JM13, SGM22, SPS15]. spread [AFTG19, FX12, NT94, PGAB22, SdF12]. spreads [Tho92b]. SQB} [JLL18]. SQB}-tensors [JLL18]. Square [Bom02, CL74, JN80, MDW11, ZWZ10, AKS18, BL13, DGL22, Din15b, DS22b, Gre99, HL09, IK19, Kol23, LLW07, Mag22, Nir11, Pab22, dSP17, ST80, Slo16b, Sło18b, Sol80, Sta15a, XQ16b, ZWW10, dFSY22]. Square-zero [MDW11, ZWZ10, AKS18, DGL22, dSP17, Slo16b, Sło18b, ZWW10]. squares [AS95, AAB17, BTT20, Bur11, Caf13, DLQ18, Din98, Gra87, GK77, JS84, LW08, LW10, LZS23, Pen15, Ric85, Ric87a, SMF95, Sin73, Sin76a, TH80, Vas86, Wan15, Wan17b, WXD21, WLZ22, Wu09, XQZ13, YL11, YLW15, YZL17, ZWLZ22, ZLZ^{+}22]. squares-total [LZS23]. SR [jXL15]. SSOR [Neu84]. Stability [DPR22, FD12, JOvdD85, KSU22, Lee94, Lev74, LS22c, PT02, RRR93, TT18, BC83, BV16b, BKS12, BG21, Cai93, CCS16, CW17, Den14, DHKS16, Dra21a, ES98b, GPM98, Gór22, GX17a, HRT15, HS88, JvdD91, JvdD88, LLHS18, LZT18, Neu79]. stabilizable [Raï14a, Raï14b]. stabilizer [AF11b, JP86]. Stable [Fri84, GGL03, Raï14b, ES00, Jam88, Jam95, LC18b, ME18b, Sta84, Raï14a]. staircase [Hwa85]. Stand [De 82b]. Stand-alone [De 82b]. Standard [MŠ02, CW11b, HQ18, Rem82, YWO07]. Stanley [Har84a]. Star [BH16, Cve95, MRD15, Sta20, ACMS10, BS16, BH13, CT96, Cir15, Cir16, CC22b, DSD83, DF90, GS06, GLX22, HS83c, Kre16, Leg06, YZB22, dNV19]. star-graded [dNV19]. star-shaped [CT96]. starlike [OST20, ZB13]. stars [BF05, IK20]. State [CFP21, Are23, BLX18, BFZ07, CF21, CG16, DPP17, Gue17, Gue19, HRT15, KMP^{+}19, LLZ22, LCXC22, LCWW22, PB17, Pal21, TM19, TYC21, WC22a, WLW23, ZLZ20, ZQZ21]. state-feedbacks [HRT15]. state-wise [Gue17]. states [Drn13, HWSH13, Man17, WCS^{+}21, XXF22, ZBW16]. Statistical [DHT94, Mai11, SLW21, TS05, SAK22]. Steady [DPP17, ZQZ21]. steady-state [ZQZ21]. Stein [LZT18]. step [YD14, ZV19, ZZT23]. Stephens [Ano79b]. stiffness [LLLY23, XY22]. Stinespring [HJ19a, HJ19b]. Stochastic [Bar74, Gib92, Joh81a, Sin81a, SHC02, SJ02, Ach77, AS95, BWF83, Ben78, Bru85, BS08, BD13b, BD22b, Bur95, CMP22, Cha83b, Cha84a, Cha88, CBQW19, CL05, DPP17, Esh17, Fie90b, For79a, FS86, For88, FM78, Gib76, Gol92, GLP92, Har81a, Hwa85, Hwa91, Joh73, Joh74a, Joh81b, JNS23, Kir09, KS76, KS82, LLL15, LZZ17, LKLC22, LM89, Mal91, MKS84a, MR07, Mer98b, Min75, Min82a, Min84, Mou04, MAM15, MA15c, ME18b, NMAB22, PG22, Phu80, Pif87, Sen83, Sin73, Sin76a, Sin76b, Sin77, Sin80, Sin82b, Sin84a, SW22, Wal75a, Wan76, Wea83, Web83, Zha14a, dF13]. stochastically [Lon90]. stopping [Zar01]. Størmer [HOT15]. straight [HT91b]. Straightening [Wil99, CW89, Cla82, Mal92]. Straus [MS82b]. strength [BO22b]. Strict [Mum86c, vE22, AMK23, CC98, Mum86a]. strictly [BCKP10, CW16, Józ15, Ner22, Neu83, RQ22, ZCY22]. Striped [MOTvdD03]. Strong [DS13, FG02, HQ18, LL11, Liu15, MO91, QH17, Ska17, AHJ20, AHJ21, AV22, BGM22, BS20, CY22b, Cir17, CC22b, DA22, HA22, Har84a, HM19, LCLL22, Mos21b, Pan13, PSB23, SSG12, TKT19, XW22a, dB14a]. stronger [RST22]. strongly [ADRP14, BP89, BP90, BP92, BL93a, Cam14, DFK19, FM22a, FG06, Gol06, Sta20, Tan18b, YW10, YBCW96]. structurable [EKO14]. structural [AAB15, BDV19, DPvW13, FSvW11, GPM98, HLZ11]. Structure [Gur80, AMZ13, Boz97, CM22, DH20, FJ15, FJ09, GL85, GH23, Grz89, Heo16, HSM17, HM18, Kra05, KMS07, Kre16, LWZ15, LEH96, Ma22b, MHH96, Mar12, Mer78c, Min74a, NP09, RJ04, Tom19, WZZ22, ZZ00]. structure-preserving [HM18]. Structured [DA22, MHM22, BR01, GC22, Mag83, MMRR16, MS15a, SRK22, YW16]. structures [ÁBvW16, DLD22, Hon19, LGB17, Sin21, So90, WZH13a, WZH13b, Wil73, Wis12, YH22, ZLS22, vR11]. study [CW89, Fri78, HRT15, Mis17, SRBB22, TJ16]. Sturm [GMZ17]. style [Wat85a]. Sub [CK17, AA12, BZZ10, WM17]. sub-additivity [AA12]. sub-block [BZZ10]. Sub-defect [CK17]. sub-long [WM17]. subadditive [Tik98]. subalgebra [NJP19, SZ19, YWO07]. Subalgebras [Ake83, DR18, BZ19, BTV07, CW11b, CX12, CW12, CY22b, GLOT13, LZ09, LCL13, Liu16b, Liu19, LT22b, MPGS21, Ous15a, dSP11, Tow22, WOW09, WY11, WL16a]. subclass [RH11a, RH11b]. Subclasses [Dra18b]. Subconstituents [GWZ14, GWZ16, HZ21]. subdefinite [DN18, ND13]. Subdirect [GHL20a, LLGL16, LMLL17, LHGL22]. subdiscriminants [Dom14]. subdivision [CMM17, CMM20, Che18, DP17, LPH17]. subdivision-coronas [DP17]. subdivision-edge [LPH17]. subdivision-vertex [LPH17]. subdominant [Kir09]. subgraph [YZB22]. subgraphs [AAT17, FG06, HW88, LS09, SS23, Tik22, ZWL15]. subgroup [MS16b, Shc21, Sur98, Wan16a]. subgroup-relation [Wan16a]. Subgroups [LLT93, Sch79, Chu84a, Da 88, Den05, GL97, Hua90, Les93, Ner22, Sun21]. subject [Wan15, YqS16]. Submajorization [AR22, Han17, BCK15, LP22b]. submanifolds [RBB22]. Submatrices [LP03, Che22a, Cho82, Cra09, De 82a, FJM00, FT21, Fie81, GKV89, HS87, HK17, Joh84, Loe87, MN96, NTU93, SP87, Tho92b, Vei82b, Zab89, dO75]. submatrix [DS09, XY22, ZHZ11]. submodules [GS20a]. submultiplicative [Kra05, Kra06]. submultiplicativity [HJ87, JN90a, JN92, JN00, Mer78a]. submultiplicity [BB15a]. Subnorms [GGL03, GL01, Gol07].

Subpermanents [Min75, HM95, MM90, Mal91, Vrb86]. subrings [MP87b]. subsequences [ACJ17]. subset [Bal79a]. subsets [MO13, ySpW22, Tom19]. Subspace [Pat17, BEH19, Das18, DH20, EJS22, GP06, Jaa22, KV23, LH22, LC18b, MW18, YW10]. subspaced [RRR93]. Subspaces [BR04, CD01, Gal08, haYsS99, BG86, CDJ13, CLW15, CLR81, CHK00, CS95, CF15, Dat16, Far94, GQ06, JS94, JP15b, KMEO12, LS22c, MMRR09, Pat77, Pat84, Ras19, Rod81, SSGP23, ySpW14, Tam98, WG91b, XFG14, Zha13, dB16]. substochastic [CKP16, CK17, CCKL20, CCW20, Esh22a, Esh22b, Gib92, Har84b]. subtends [Bal80]. subtraction [WTD22]. subtractivity [Zha05]. subunitary [ZYL13]. subvarieties [GP15]. such [Fan92]. Sufficient [CHQ18, CC98, EOvdD09a, EOvdD09b, Gue19, LP23, LOO22, AW07, HB83, Pal79, Rou98, SQ15, Wan82, ZW17b]. Sum [KR02, KL13, Ami80, Ash16, BST23, Ber17, BP09, Caf13, Cal97, Cal99, CS22, Dji16, Fow80, FMXZ22, GR14, Gra87, Gro99, HP90b, HLYZ22, KM85, KLO14, LPS08, LLHS18, LT13, LSX19, MW18, MM90, Mal91, MV91, Nor08, QK92, Sil88a, Sin90, TH18, TCI13, Vas86, Vas87, XQZ13, Zel85, Zha05, ZMG20]. sum-of-squares [Caf13]. summation [IW95]. summing [AAPR19, AP17, BMP10, BPR14, Cam14, PR12, PSS15, RSPT17]. Sums [BC17, BCKP10, Bur11, CQ18, DRW93, FT21, dSP22, Sil90, Slo16b, Tho86c, Bal79b, BHK10, Bre18, BD22b, De 90, De 78, DHS95, FV97, GHL20a, Gar21, Gol92, GK77, HO17, Kar83, KW21, KK22a, Kur04, Lew96, LSOvdD95, LLGL16, LMLL17, LHGL22, MRT12, MH19, PH15, dSP17, Ric85, Ric87a, TT12, TZS19, TH80, Tho92c, Tia05a, Wan76, Wu22, dP74, vR11, Sło18b]. Super [CCC21, CPT15, FD17, KV22, LS22a, AA12, LOO22, MZ17, Mit10, Zha21]. super-additivity [AA12]. Super-biderivations [CCC21, FD17, LS22a]. Super-linear [KV22]. superalgebra [CCC21, CPT15, CMDE19, DH18, GMZ11, HH22b, SYZ22]. Superalgebras [LR13, dNV19, AH20, BMN22, CPS22, CV94, FD17, FdSV16, GLdS22, GdSV16, HLS18, Lee94, LCC19, LC22, PN22, WL16a, Wan16c, WL16b]. Supercharacter [Sun21]. supercommuting [HDJT97]. Superconvex [AFA22]. Superconvex-cyclicity [AFA22]. superderivations [GLdS22, LC22, SYZ22, Wan16c]. supereigenvectors [lWW15]. supereulerian [WYSL22]. superinvolution [LR13]. superoperators [JP15b]. Superoptimal [DS02a]. Superposition [Ogu20]. supersolvable [Tow22]. supertree [Wan22b]. Supertropical [INR18, IKR12, IK20, IK22]. support [Cao19, WZD16]. supported [HMO11]. supports [Hil18]. surfaces [BFM22, FYIS13]. Surjective [Cos14]. surjectivity [BGLSSS20, ME18a]. survey [Ano92g, BL92, CR90, Far93a, KS17b, Mer95]. SVD [Gut23]. SVEP [BZ22a, DC20]. sweep [WL22]. switch [GKT97]. switch-setting [GKT97]. Sylow [Sun21]. Sylvester [Bar80, Bay22, Bed11, BS17d, DH11, DM22c, GPM98, Haj18, HKT15, He21, HM20, HK17, KM17, Lei83, SWL13, WXD21, XW22a, dFK21, dCdSA22].

Sylvester-like [KM17]. Sylvester-transpose [BS17d]. Sylvester-type [He21]. Symbiosis [NT91]. Symbols [GN02, KLL22]. Symmetric [BS17c, BBM15, CCLL22, CKMY22, CA03, CD03, DS02b, GR81, JT04, JMP17, LR96, Pat81, Pie85, Zha04b, All14, BH23, Bal14, Bal18, BL97, BS16, BS17b, Ben90, Bén11, Ben74, BH74, BB75, BB15b, BLSV12, BCT23, CKGP21, CL85, CFSGL18, CL92, CSS21, CDPS23, CFZ23, CW23, Che91, CT09, CLNT17, CLN13, CK19, CKKY23, Coe96, CW93, DH11, Deu80, DIJ23, – Dok75, – DLR91, DCLW19, Elg22, ETM22, Fan73, FJ15, Fit94, For87, FM22c, GS06, Gib76, Gon17, GZW20, HHC12, HWL92, HJ20b, HZS22, IAM23, JKB17, JR09, KKL22, KL19, KKR12, KN81, LR13, Lei95b, Li87b, fLLyP15, LM95, LLL09, LL10a, LZ19b, Lim79a, Lim90a, Lim92b, Lim92c, Lim97, Lim10, LT14, Liu92, LTT11, LGB17, LPX23, Loe01, LS21b, Lop11b, Lop18, MM05, MS85a, Mat92, Mat11, MQ21, Mer94, Min76]. symmetric [Mir18, MMG15, MAM15, MA15c, Neu74, Neu83, New74, Nom73, NTU93, OP13, Ore19, Pat83, Pat86, Pat92, Pat08, Pat09, Pat17, Pen15, PM81, RR91, Rei75, RW96, Rod97, SRBB22, SG09a, SN22b, ST91, Sou83, Sra20, Sta23, Sta14, Sta15b, SvW14, TXL21, TMH21, Til77, Tre87, Uhl79, UF22, Wag87, Wim85, Zha05, dCH21]. symmetric-indecomposable [BB15b]. symmetries [HWSH13, XHL23]. Symmetrization [HO17]. symmetrize [Pat90]. symmetrized [BJ93, BD97, Gam18, Hol95, HT12, MC78, Mar78, MC83, MC89, Mer77, WG91a]. symmetrizers [CdS03, Uhl13]. Symmetry [DF95, HT92, HK16b, MS08, MR74, RC18, Wil81, Cha78, CL86a, CL86b, CL13, DP00, Der83, Fin92, HLS18, Hil76, HK13, Jin17, Jun88, KCBR87, Lim76a, Lim76b, Lim82, Lim90b, MC82, MC86, MM76, Mer78c, Mer79c, Pat95, Rob77b, yT86, WG91a, WG91b]. symmetry-antisymmetry [Pat95]. Symplectic [BB21, BB17c, Ter22b, Chi22, DP09, Den05, Don91, FP17, Hou22, LYS15, Mal92, MC19, MMRR09, New75, Shc21, SM19, Zha22]. synchronous [ZV19]. system [DH11, ESH23, FHW16, HW14, He21, HM18, KO10, Lop18, Mil22, Ned98, NF19, Pop17, SPKC15, Sti87, WLZ22]. Systems [Gib89, JR93a, KS02, KT83, Lev74, BH23, BC83, BNKJ22, BE23, Cam83, CNW21, CW20, CJ12, CF15, CFP21, CIRW21, DPP17, DH86, Din00, Doo16, EKO14, Elg22, FNW18, Gow22, GR93c, GC22, Hai92a, Hai92b, HMAD13, HW15, HRT15, Hla17, HWL21, Hua22c, JS84, JL22, JT23, JR93b, JSY14, KP11, Kul22, Kyr21, LF15, Lia80, LZ19b, Lop11a, MC17, MG22, MW19, MHM22, NT91, Roh91a, ST86, Sin83, Son22, ST97b, Wer84, WW80, WL22, XLL15, Xia20, XJW17, Xio23, YqS16, YZC21, ZWZ23]. syzygy [HJ20a].

T [Ano16b, Slo13c]. T. [Cha84b]. T=AB [AP74]. T=AB-BA [AP74]. table [Sin21]. tableaux [Ber86, Chu90, CPT15, Don91, KW92, KW93, Mal95, Rem82]. tails [Gol16, Gol21]. Takagi [HZ12, Ter22a]. taken [Odo83]. tall [SMS21]. tangent [BB13, TG23, Til77]. Tate [AST07, GR14, Rom07]. Taussky [JZ95, Joh90, Sch77]. Taussky-Todd [Sch77]. Taylor [MZ22]. TD [HZH20]. team [KS93]. technique [DLL19, FN14]. Techniques [BT17, Whi75, Whi79]. Temperley [Byt20]. Tensor [BB13, CHS22, lLZjZ19, LLV18, Pat92, Wil73, Wil99, Bal14, Bal18, Bar76, Boa13, Bre16, Byt20, CX23, CL85, CKR18, CKR21, CL83, CQW16, CY22a, CKL16, CW22, Der16, Der18, DM18, Din15b, DH16, DCLW19, EJS22, FHL^{+}14, Fow80, Fri82b, Gam18, GPW22, GWW15, Gil16, GLM^{+}22, Hac22, HMR22, Hol95, HT12, HM20, HM22, HWSH13, IAM23, dJfLlZ^{+}21, Koi21, Kru06, LL16a, Lee94, LXQ21, Lei18, LN14, LZZ17, fLLL^{+}17, LZ19a, Lim92a, Lim15, Liu18b, MC80, MS88, Mar11, McF79b, Mcf79a, Mer78b, MWC22, Neu74, Ng14, Nir11, Nor95, Nor08, PM22, Pat87, RSPT17, STME10, SN22a, SWL13, Ven19b, WX19, WXD21, WLZ22, Wes95, XJW17, XJS20, XW22c, YLX20, YQ90, ZXF15, Zhu22, Zus22, PMP22]. Tensoring [Abd09]. Tensors [Lim01, MR74, BD97, BM17a, BS22c, Cha78, CL95, CBQW19, CK19, CCLN16, CPZ17, DP00, DM19, DM22d, DF95, DL22, – Dok75, DNY22, FDLB20, GMV23, Gau76, GR10, GZH19, GY22, HKv23, Hil76, HT92, HK13, HK16b, HQ14, HM19, JLL18, JZL20, Jin17, JKB17, Jun88, KSMB16, KJ13, LPZ13, LZ01, LL16b, LZL16, LCLL22, LZT18, LKLC22, Lim76a, Lim82, Lim90b, Lim92b, Lim97, Lim10, Liu18b, MS08, MC78, Mar78, MC82, MC83, MC86, MC89, Mat11, MM76, Mer77, Mer78c, Mer79c, MSS18, PBM20, Pat77, Pat90, Pat08, RS17, Rob77b, RC18, SL19a, SQH15, SN22b, SQ15, SĆK^{+}20, Sta14, Sta15b, SLH16, SLH18, SL21, SF17, SMS15, SMS21, SZWB16, SZBW16, SZZY18, Sun20, yT86, Wan82, WG91a, WG91b, WLX16, WSZL17, WWW20, XQ16a, XHL20, ZL18, Zha19, dP74]. Term [KS12, BSL01, BS16, CR22b, Hua01, Kom10, YLX20]. terms [Bal14, CHQ18, Dua22, HHY22, Kir00, LL15, MF20a, Pan13, zSqSZ16, TW19, WWZ20, Wan21, WL23, ZBW16]. Ternary [ABM22, BX11, BP09, Bus23, KPS19, LLY22]. Test [BC85, LB19]. tetracyclic [GH22]. th [HL20, IAM23, KW21, LT91, LLS10, NMAB22, SSG12, yT86, vBW95]. Thanks [Ano10, Ano11b]. Their [SK03, AdFK15, ACJ17, AKOZ20, BS15a, BS92, BS16, BS17b, BS17c, BS22b, Che92, CBZ22, CFZ23, CL14, CTY20, CAP22, DL22, GKS01, GL16, GLS16, GZ22b, HLT22, Hon93, HL18, Ilm16, Iop19, KJF23, KS15, KT19, KE82, KAE22, Kut19, Lei98, Man91, MP90, MMRR16, Mia93b, Moh16a, MA15c, NS85, Nor08, PP22, PV23, Pow88, QM21, Ric81, SLW21, SU77, SS23, SA21, SLUG04, Sin93, SL95, Sra20, TT12, TZ19b, Ter22b, Tia11c, Tia18, TMOvdD93, Tur02, Uhl99, Zha97, ZBW16, ZY08]. Theorem [BBM20, LMMR15, Mac09, Pop20a, Adk95, AM15, Aha76, Ari16, AY95, BGK99, BGK00, BZ22a, Bla77, BS22d, Bun91, Car19, Chi11, CPLH17, CC18, DFK19, DMV18, De 75, DP18, – Dok17, DC20, Els78, FW13, Fie73, Gar75a, lGyW99, GJR97, GP22b, GS77, GS83, GT11, GMW86, GPR84, GX17b, HL84, HJ19a, HJ19b, HJ20b, HS75, Hsu85, Hua09, Hua23, IT16, JLL11, Kat75, Kov86, Koz17, Krä87, KLO19, Laf80, LRR87, LN80, Lim76b, Mag22, MMR07, Mar78, MS82b, Mel18, Mer89, Nag17, Nas75, New76a, Pas15, Paz84, PSS15, Pop22, Pul75, QLK22, Ras17, Rei83, Rem90, Rod22, Roh91a, Roh91b, Roh04, Roi99, Rub20, SV14, San18, ST80, She77, Sin84b, Sou86, Tam97, Tan18a, Tan20b, TZ17, Tau73a, TS05, Tos79, Wes75, Wes81b, YLL90, ZLS11, Zha15b, ZN18, dPN81]. Theorem [HWY17]. Theorems [Jun88, AEV12, AyT84, Bar07, BGLS07, BT22b, CMR20, CM84, EL09, FD12, GL85, GL16, Gut01, Hon91, Hsi78, Hua16, HWZ18b, ISCS11, KK18, LRT91, Li96, LM98, LN16, LW17, Liu22b, MP87a, Mar89, Mas10, MS81, ST86, Tan13b, XLL15, XFG14, ZV19]. theoretic [Dom14, KMH19, Mou04]. Theoretical [Far17, Ari16, Cla92]. theories [Sun21]. theorists [Sch77]. Theory [Min83, Min87, Adu15, jAS21, Atk91, Bal23, BOvdD95, BG23a, BL80, CdS03, CNW21, CW81, CD85, FGK23, FW73, GMV23, Gar73, Gau76, Gor83, HCS19, Joh82b, JSY14, KN81, Kri87, MH22, MMRR16, Mej19, MN96, MP21, MA15c, PT10, Sch22, Sch77, SQH15, Ste82, Tho90, WY21, WW76, Whi79, Wis12, XQ16a]. There [DKK^{+}20, Stu92]. Thin [Mou20]. third [BB19a, GM87a, Guo07, PLH01]. Thirring [TWK22]. Thompson [Ano97b, Day97, GPR84, Hia94, New97, NT94, Tam85, TWK22]. those [Pat83, WW12, ZH22]. Three [CR22b, GN02, MS82b, Mir81, Slo15a, Wil83, Abd09, AR15, AOTR11, BSÜÖ13, CLX22, Cro18, GG22b, GSZ22, Gue19, Hou09, HL22, Li84, LT12b, LZ15b, Mil22, MMG15, NR23, dSP17, PGLM00, Sch79, TZS19, TH80, TRM22, Vas86, Wri93, XL11]. three-dimensional [NR23]. three-state [Gue19]. Three-term [CR22b]. Threshold [HMK22, LCL22, AAKS21, KADM10]. thresholding [GLM^{+}22]. tight [DO19, JW22, Kim13]. tightness [MNS22]. Time [Uhl99, CW90, FK91, Gue19, HKT15, Kas85, WT23, ZZ21]. times [Che18, GLLM21, HWX22, HL18, KN00, Zar01, ZL22]. TN [CJ20]. Todd [Sch77]. Toeplitz [ACI15, BG83, BdPN14, Boz98a, Boz98b, Boz99, CDPS23, CN08, CLNT17, DOP22, Doo16, Elo18, FH89, GR91, GR93b, Gou13, GHL20b, Hil17, HM76, HM78, Ikr15b, JTW17, JS91, JKKC22, KN22, KKL22, KT21, KL15, KLL22, KAE22, LL14a, Mar20, Mel12, Poo17, Sol06, Til98, Tis93, Tre85, Tre86, Tre87, YZL17, Zim89]. Toeplitz [?]plus-CauchySolak:2006:RPN. Toeplitz-plus-Hankel [Ikr15b]. tomography [Rau18]. Top [Coa99]. topheavy [BC80, FH73]. topics [McD92a]. Topological [MAH18, MAH21, BGLSSS20, EB17, GP22a, Hau12, MVPS14, MRS10]. topologically [GS20a]. toroidal [Ye11]. torsionfree [BDvW16]. Total [Wan17b, XW22a, CC23a, Cao19, CC98, LLW07, LZS23, YC19, Zar01, ZWM17a]. Totally [CJ12, WHG17, AB19, BGGS17, CCU21, Cas23, DJK08, Gar85, RBB22]. Tournament [MP90, BCCL22, Dec95, KS93, Kir93]. tournaments [DLSS18, JKY23]. TP [CJ20, JN14a]. Trace [BF15, Epk89b, Epk89a, FG22, GR14, Gur98, Kov77, Kru06, Lew96, MP13b, AW07, AST07, Car06, Cho18b, Dra16, Ebe80, Gal79, Gei78, IAM23, Kap77, KMMS22, Mau78, MR19, Reg87b, Rem88, Rev81, Rob77a, Rom07, SQH15, SW81, Tik98, YSK15, ZY13, ZX16, ZBW16]. trace-distance [ZBW16]. trace-form [Mau78]. traceability [Lu22, WY20]. traceable [ZW17b].

Traced [RSPT17]. traceless [LT14]. traces [Cho18a, Fra13, LWXF15, Wan16d, XWF15, XZ18]. tracial [BZ19, Man17]. tragedy [Die79]. trains [HMR22, HWSH13]. transfer [Are23, CF21, CG16, KMP^{+}19, LLZ22, LCXC22, LCWW22, PB17, Pal21, TM19, TYC21, WC22a, WLW23, ZLZ20]. transference [FP17]. transform [AL78, CM19, DL22, DKY22, Elo18, JLL07, KKY22, Loe73, RS08b, ST08]. transformation [BG82, CS10, CR77, Dia90b, Ell93, GAGM18, Lim76a, LM10, Rob75, WWS19]. Transformations [Hla17, Loe74, Wes74, Wes95, AW83, Bea78, Ben78, BP12, Bot79, CL83, CL92, CL21, DK17, FA22, Gau76, Had83, Jam81, JLL88, Kov77, Kun99, LL90, Lau81, LLWX17, Lim79a, Lim90b, MHH96, Mcf79a, MP73, Min77, Neu83, PLLL20, Pat77, Pea84, SSZ93, TRM22, Wat85b, Wat82, Wei93]. Transforms [And04, Far20, Lee98]. transients [MNS22]. Transition [RW96, Tau74, CPZ17, LN14, Wal11, XXF22]. transitive [CL86b, LZ09, LMR00]. translation [Sel81]. translations [PY82]. transmission [jAS21]. transportation [BS18d, CCLL22, Har91]. transpose [BS17d, Cho18a, Gow80, WCS^{+}21]. transposed [CR77, Rob75, YH22]. Transposes [Ric81, Sin93]. transposition [AF11a, AF11b, AF12]. transversal [Kas85, Mum86a]. treatment [Smi74]. Tree [AT18, LHbL^{+}15, AFJK17, Bap97, BGK99, BGK00, BCDF09, CD21, DIJ23, GGH22, JD99, JS08, Mer89, NP09, NU94, Pat07, PJ90, WWWT22, Zim98]. tree-patterned [NU94]. Trees [BMP22, MSZ04, Moh16b, AAS14, AD17, AEG19, AKB21, BGK00, BNP06, BLP08, Bru79b, CJY22, CDK22, CJL22, DWL20, FGHL22, FLZ07, Fie90a, For79a, GHL10, GH23, GF10, HLW^{+}21, JSW09, JS16, KNS96, KN97, KN00, LP07, LT12a, LC14, LZ19c, LH13, MN04, Moo95, Mou20, NP13, Obo23, OST20, PP22, SG06, TJ11, WBZ14, Yin21, ZL22, ZB13]. tri [MC19]. tri-covectors [MC19]. Triangle [AAP82, DZ22, Tho89, TW19, YW10]. Triangle-free [DZ22, TW19]. Triangles [FS18a]. Triangular [BKL02, Che03a, Gal03, Iov15, MP19, Wri22b, BHC16, BT90, Ben05, Ben07, Ben16, Ben22, BGV11, BH15a, BDvW16, Bou13a, BSZW13, CdM15, CW16, CQ22, Cho10, CKT20, CL01, CC18, CIP18, DTS14, Dol07, DC20, DC23, DW12b, DHH22, DH22, EGT16, Ere15, FZ17, Foš05a, GG22a, Gha13, Ghi22, HC14, HH85, HLZ18, Hou21, Hua10, JLZ12, JSE90, JTU00, Kob06, KPPR21, LŠS01, LB11, Li11, LG23, LT22b, MJ07, MŠ98, Ner22, Neu83, PS20, QH11, Ric78, Sil99, Slo13a, Slo13b, Slo14a, Slo14b, Slo15b, Slo15a, Slo16a, Slo17, Sło18a, TZ17, UG21, VM07, Vla22, Vla23, WOY06, WZC11, WZL13, Wan16a, Wan16b, Wan16d, Wan18, Wan19b, WLB19, Wan19c, Wan20, Wri22a, XW12, XWF15, Xio23, XSH18, YZ13, ZZZ16, ZWW10, dFSY22]. triangularizability [DK14, YT15]. triangularizable [Rus16]. triangularization [Laf78]. triangularizatlon [Wri93]. Triangularizing [Bot97]. triangulated [Muk20]. triangulation [Che18]. triconnected [Vo83b]. tricyclic [GL11, HHW15, KS23]. Tricylic [RY93]. Tridiagonal [SJ02, AC13, Aln15, BG82, DOP22, Deu80, Doo16, IOMA21, IOMAN21, MR13, MNN^{+}98, SR17, YHCW14]. trigonometry [Gus94]. Trilinear [MN13]. tripartite [HJ77, TLC22]. Triple [LCL13, AN19, AWM22, Ben15, BM15c, BE23, BM16, BMGMC15, CM19, CX12, DD14, EKO14, Elg22, JLZ12, KO10, LW11a, LF13, LWXF15, MC17, Mol15, PS20, SN22a, TS21, WY11, Wan16d, XWF15, YZ20]. Triples [Tam02, CGH15, Dom00, GPM98, HZG14, KHG15, yT00, WHG17]. tripotent [BSÜÖ13, LWB11b, Tia11a]. trivectors [– Dok83, Wes80, Wes81a]. trivial [AVVÁ18, BCS14, Gha12, dSP11]. triviality [Cha78]. Tropical [XW22b, MNS22, Shi16, TS19]. Trotter [AH16b]. true [BR22]. Trumping [KPP13]. Truncated [FM22c, CB22, Har91, KL15]. Tsallis [Fur11, MF20b]. tubal [EJS22]. tubal-Krylov [EJS22]. tuple [MS21a]. tuples [GJR22, Shk12]. Turán [TLC22]. Twisted [Ito16, CHS22]. Two [AY74, AYC93, BTV07, CC14, Egg95, EÖ15, Guz88, Hon91, Hsi78, KM85, MWLW22, PH03, Pat03a, Peñ02, Ric81, SLUG04, Shi15b, Tia11c, Wat89, Zha19, AAS14, Adh13, AS11b, APEPPS23, Ash16, BB06, BLP12, BC83, BH90, Bed10, Bén10, BLZ10, BR10, BCI13, dAB17, CW90, CGY13, Cha90b, CL17c, Che22a, CLN13, Con77, CAP22, Das04, DP17, DWCD19, DR18, Dra17a, DS18, DS19b, FD12, FJ15, FNW18, Ful76, GR14, Gro99, GC22, HS83c, HP90a, Hil18, Hon93, IOMAN21, Joh22, JN22, JLS09, Kir04, Kyr21, LS10, LT14, LWY11, LWB11b, LZ14b, Lon80, Lon81, Lop18, MM09, McF79b, MHM22, Mia93a, Mir18, MRS10, Nof15, Orz84, dSP16a, PUÖ15, PGLM00, Pok87, QK92, Ram22, SGS05, SG09b, SS96, Sin90, SST22, TT12, Tan10, TLL11, Tia00, Tia11a, TJ16, Tia18]. two [Tia21, Tur16, UPÖ16, Wad85, Wad90, WZ13, WM21, Wan22b, WL22, XW22a, Xio23, Xu15, XHL23, YD14, ZBW16, ZV19, ZCW21, dLMO17, dTS20, tHK17]. two-by-two [Xio23]. two-colored [SGS05, SG09b]. two-dimensional [dTS20]. Two-local [MWLW22]. two-parameter [Adh13]. Two-sided [EÖ15, Kyr21, Tur16]. two-step [ZV19]. two-sweep [WL22]. type [ABGP21, AdFK15, AO97, Anj17, BM19a, BS22a, BO22a, BZ22a, CC19, CW10, CH12, CY22a, CFGM06, DM17, DH17, Das21, DM22b, DKM10, Dra10, Dra16, Dra17a, Dra18a, Dru14, FW73, FJK98, Gar73, Gut01, Hai92b, He21, HJ19a, HJ19b, HJ20b, Hia94, HSW19, HOT15, HZG14, HWZ18b, Hua22a, Hua23, HZ21, Hwa85, JS19, KHG15, KKL20, Kim91, KK18, LLMT05, LXQ21, LLL15, LZL16, LP22a, LHZ22, LZ17b, LFY16, LPW17, MW15, MM22a, MHA17, Mer78a, MP13b, ND05, NY23, Oku91, Rei77, Sab19, Sch22, SGM22, Shi18, SAK22, Spi76, SYZ22, Sun21, SY23, Tho78, Tom22, Wal75b, WT16, WHG17, WWW20, Whi80, XLL15, XQ15, XXZ11, YGÇ19, YD14, YLC19a, YWO07, ZZZ12, dFK21]. Types [Mir81, AAdFS19, CL17a, Fen10, Li84, Ma15a, MTR19]. Typical [MSS18, SMS15, SMS21, CO12].

Ulam [CAP20, JVMCPR19]. ‘Ultra [lWY19]. umbral [Jon78, Par78]. unbiased [LHS^{+}21]. unbounded [BHC16]. uncontrollable [CFP21]. underlying [CHL18, Dua22, LZX19, LY22b, Wan19a, WL20b, ZH22]. unextendible [WCS^{+}21]. Unicyclic [FKP03, FHT14, Fan05, GZ22a, HM23, KS22a, KS22b, KS23, LZ11b, LT11, NS09, SS07, ZZZ12, ZZ15, ZZ21]. Unified [ISCS11, Mar89, TX13]. Uniform [BX07, KC16, KKP13, BG23a, LCC17, LZL18, LGZ20, LWL21, LW22a, LL22b, SR17, SSfW15, SKLS20, SKLS21, Wan22b, XW19, ZL19]. uniformly [LCL22, WC16]. unilateral [FMSS19]. Unimodular [Ger04, HW08, LPWX10, Wat90]. union [AS11b, Das17, WSH^{+}11]. uniparametric [CF15]. unipotent [GT79, Shc21]. unique [BGZW16, CNW21, DPR22, FS19, KS22a, SS07]. Uniqueness [BY76, BB17a, Bal18, CPZ17, HHSY22, HS85, Lim76b]. unit [BL21, CGH15, Dat80, DO19, FZ11, HHY22, LY22c, LY22b, LB19, MKS22, RAN22, Sch92b, WZC11, Wan16a, WL23, ZH22]. unit-circle [BL21, LB19]. unital [Ben15, BG16, CAP22, EKO14, NY21]. unitaries [CAP22]. Unitarily [JN92, MKS84b, Pok87, AE10, CLNT17, HJ87, HMN91, JL88, LT87, LT90b, LLW07, SAV01, Wan17b]. Unitary [ACI15, AKR23, AP23, BL83, dAB21, BD84, Bot79, CL22a, Drn13, Ell93, Far90, GR91, Grü76, GWZ14, GWZ16, GC22, GL97, HINS10, HH85, HH89, HS09, HX21, Ikr15b, KC16, Kim91, LT89a, LT89b, LT90c, Li91, LT92c, LM93, LPS08, NC22, RS07, Rai87, Sed11, SC22, Tho74, TM19, Xia20, YLW22, Zha22, dB10, ZT12]. unitriangular [HS17, Slo13a]. Units [New93, CG19, Put82b]. unity [AAT16, Wan20]. Universal [CR18, KT19, PZGL22, DFPS23, Hua22a, Tia11c]. unknowns [EÖ15]. unoriented [FTZ08]. unranking [VDW85]. unresolved [Min74d]. unsolvability [CS19]. update [WW20a, WW12]. updates [SKP18]. updating [LLLY23, XY22]. uplets [BB96]. upon [Deu80]. Upper [CR77, GM21, Lás91, Mel13b, ST94, Sha91, TD14, Tia17, TW19, ADJAH22, BHC16, Bal14, Bou13a, CW16, CJL22, CC18, CIP18, Dol07, DC20, DC23, DHH22, DH22, Foš05a, GG22a, Ghi22, HC14, HLZ18, JKB17, JTU00, Kob06, KPPR21, LGA^{+}13, MŠ98, dSP16a, Raj09, Rho94, SSG12, Sil99, ST96, Sol00, zSqSZ16, Sou03, TD16, TD19, UG21, Vla22, WOY06, WZC11, WZL13, Wan16a, Wan19b, WLB19, Wan19c, Xio23, Xu22, YSY17, ZZZ16, ZWW10, Zho16]. use [LT10]. user [Koh11]. Using [CG17, AFTG19, jAS21, Bap88, DTSI15, DLQ18, EB17, GPM14, KW92, KW93, KMP^{+}19, LLW07, MS16b, Ogu20, SPM14, Vo83a]. Uzawa [CL22b].

V [Ano16c]. valid [Sin81b]. valuation [AW13, Piz87]. valuations [GM90]. value [ANK22, jAS15, jAS21, CQW16, CS19, CH87, GA18b, GHT11, Gut22, JW22, JN90b, Lék19, MS81, Sad23, SRT22, SA10, Ter22a, YZ16, ZZL23, ZL18, Zha19, Zou20]. valued [AW13, BB22, BG16, Bri89, DD10, DS23c, EDKA07, FD14, GR23, Gil93, HSM17, HF18, JS20, MS81, MWLW22, Oi20, Sil84, TN22, Tho89, Tia17, Tur02]. Values [Mir81, AG22, AHJ87, AD17, Bap87, BI95, CJK^{+}13, DS23a, GAMA23, GRSW92, Hil17, Hua22b, JRMFMASS16, Joh75b, Kap77, Kar83, KNS96, sL19b, Li84, LT90a, LT92b, LT93, LCL14, MS85b, Mat92, Nie10, Odo83, RW21, ST00, Tag12, TH18, Tho75a, Tho75b, Uhl14, Uhl20, Zha00]. Vandermonde [Bra86, GKLR82, MZ22, Ric87b, SSZ22, WLX16, Wör20, XW22b]. Vanishing [BS18b, CD22, GQ06, MF77, Mer77, Pok87]. Variable [CSCH01, Hua22c, KY17]. Variable-parameter [Hua22c]. variables [DS18, DS19b, Est80, GHL20b, dHS23, Paz84]. Variants [KMS21, Pop22]. Variation [Fri82b, FJ09, Ben90, BD84, BE90, BKL97, Deu80, Fri90b, FJ01, Hos22, JP07, Kir04, Xu22]. Variations [Yiz02, Yiz03, BM15b, MW80]. Varieties [IM18, BB13, Bal22b, EGMS22, Hul81, Pat86, dSPdS22]. variety [BB19a, Bal22b, BCS14, GP15, Lim92b]. Various [CID16, KD23]. vec [HS81]. vec-permutation [HS81]. Vector [Adh13, BH90, SL15, BB22, Cla92, Con77, Das17, Das18, DPP17, DM87, DH20,

– Dok83, EDKA07, Guo20, HNB22, HS85, HF18, How80, HCS19, JS20, KL95, KMEO12, MW18, Maj22, MWLW22, MN13, Pan15, dSP22, RŠ09, WWS19]. vector-valued [BB22, EDKA07, HF18, JS20, MWLW22]. Vectors [GMD85, BdPNdP17, CJ97, CJK^{+}13, DMdO92, JP07, Kir91, LK16, Nom73, RS17, Roh06]. Vein [Fie84]. verbally [Cen11, Per15]. Vergne [Ous15a]. Verified [Miy22]. Veronese [BB19a, Bal22b]. version [BHY22, BGK00, BS17d, DKK^{+}20, HL84, Hua23, IT16, KS14b, LMMR15, MMR07, Mer89, MP94, MP95, Nag17, PSA14, Pas15, Tho89]. versions [AEV12, EL09, Orz84]. versus [Mel22, Vel21]. Vertex [AAdFS19, KP02, MS87, SS22, LY22a, LPH17, WLW23]. vertex-addition [SS22]. Vertex-removal [SS22]. vertical [MN96]. vertices [AEdF13a, AEdF13b, ACMR17, BBG22b, DdF15, HZ03, JST18, KNS96, LT12a, LZZ17, LDL17, SS22, WH11, WWZ20, WWWT22, Wan22b, WDJ10, ZL19, Zhu15]. via [BM17a, BS22c, BB19c, CMM22, DMV18, FDLB20, FS18b, GMV23, Gau76, HJ20a, HRT15, HM22, Ikr15b, JW22, JMP17, Joh11, KNS96, Kul22, LV18, Lék19, MFS22, MZ22, NS18, PSMA16, PM22, RW84, SAK22, Sta15a, SZWB16, SZBW16, Tau81, Uhl13, XQ16a, YGZC22]. Viennot [Foa79]. viewpoint [Dic84]. views [JNS23]. Virasoro [CY22a, CHS22, DH18, LS22a, LGG19]. Visualization [AP23]. Volterra [ALP22, QM21]. volume [CC23a]. vs [ABC18, ME18a].

w.r.t [Cha96]. Waerden [Fri78, Fri82a, Lon80, Min82b, Sin81b, Sin84b]. Walk [Rod03, ZZ21]. Walk-regular [Rod03]. walks [CMR21, CLX22, Che18, FS14, GLLM21, HL18, WS23]. Wang [Cha89, Hu87, Hu89]. Watkins [Pat93]. wave [Far17]. wave-packet [Far17]. Wavelet [Mej19, Far20]. Wavelet-multipliers [Mej19]. ways [How83]. Weak [EPR16a, EPR16b, Lin21, MW19, Neu79, YF18, ZC22, ZCZT22, CFGM06, EBB18, FSWX22, GJ22, IS15, JGS90, LD22, LRD21, MNS22, MM22b, ND13, SGM22, TJG22, Wer86]. Weak- [YF18]. Weak-local [EPR16a, EPR16b]. weak-majorization [GJ22]. Weakening [Hua16, You15]. Weakly [Fra17, IS16, Lud21b, FJK22, KSMB16, KY17, LMLL17, SC21, SZWB16, WWW20]. Wedderburn [LMMR15].

Weierstrass [Atk91]. weighing [KMS07, LN10]. Weighings [NW02]. weight [GG22b, KN04, LT11, LT12a, TJ11]. Weighted [AG96, BG23b, MS10c, Mos19, MM22b, SKM17, ZW20, AHJ20, ASR22, BR19, BPR20, BSG^{+}12, CW10, CL17a, Che16, CNUV16, DM22b, DTS14, DLQ18, Dra18a, EJD88, Fer10a, FOT23, GMV23, GCP20, GH23, GHL20b, KS16, KY17, KNS96, KN97, KF98, KMP^{+}19, KKS22, LT11, LT12a, LLC21, LZ23, LPX23, Mos20, Pep19b, PSB23, RBG23, RF19a, SLW21, SMM14, SMM22, Szw18, TD14, TD16, TD19, TJ11, TGW14, WfMR16, Wim85, YLS10, ZY17, vBW95]. weighting [DMV18]. weights [AAB20, AKB21, Bén10, Lee94]. Welch [CDK20, Dat16]. Wenzl [Byt20]. Westwick [Atk84, Poo80a]. Weyl [BHC16, BZ22a, CC18, DC20, GCSC20, ML12, TOM88, ZZ16]. WG [FOT23]. wheel [BBG22b]. wheels [BS95]. Where [Bal80, BLR12, CD01]. which [Ano16c, BB75, Ben78, Bre15, BF96, FTA20, FH73, HS83a, JL90, JLS09, Kap77, KMT96, KE82, Loe93a, Loe93b, MZ91, MN04, Ous15b, PP16, Ric81, Sin76b, Sin81b, Zho22]. Whittaker [aCC17]. whose [AS95, ACO21, Ano16b, APEPPS23, AOTR11, BT22a, BP90, BPS16, BPW16, BGZW16, CCGO11, CFSRL22, DFK19, DH86, DM87, DIJ23, FdF09, HNS99, HL20, JD99, JS08, KLO14, LR13, LHH18, MSZ04, Nab98, Sim95, Sin73, Sin76a, Sta10, Stu89]. WHS [EOvdD09a, EOvdD09b]. WHS-matrix [EOvdD09a, EOvdD09b]. width [CLN21, HM76, HM78, Tsi83]. Wielandt [Alo23, CNP06, DW23, FS18b, Gum15, LL14b, ZMCN19, Zha01b]. Wiener [Adu15, MZ04, Mer89, Obo23]. Wigner [DMS23]. Williamson [Wal75b]. wise [Gue17]. without [ALPM20, CNW21, Cao19, FM21, GL85, HLW^{+}21, MS10a, New79, Pok87, Tam91, Wan20]. witnesses [LF15]. Witt [AKY21, COK19, CCC21, CZZ22, GPW22, Hai97, Haw05, Hua09, Lew90, LGB17, Yao22, Zha97]. Witt-like [Zha97]. Woodbury [SC20]. Words [WZ95]. work [New97]. worm [JL22].

x [Bea86].

Yamaguti [BBM15, Bre14, ZL15]. Yamazaki [Vei83a]. Yanase [DMS23]. Yang [DBG18]. Yound [Rem82]. Young [AdFK15, BO22a, CdS03, JKL22, KM11, Man17, Nas75, SYK15, Sha20]. Young-type [AdFK15, BO22a].

Z [BSS12, CdM15, HMO11]. Z-matrix [BSS12]. Zagreb [AEG19, ZQC18]. Zero [BCR91, BSL01, BS16, GGH22, Gha13, Gra10, Gra11, Soa08, WWZ20, Ach77, AGM13, AGS20, AKS18, Bre12b, BH15b, Bur95, BSO13, CD22, CLX03, CW16, CG19, CM07, DGL22, Das17, DZ22, DJS21, EAM19, FGMSS19, FLR85, FKKS11, Gam18, HS93a, Hil18, HLH18, JL93, KMOV07, KS82, Kol23, LB19, Lim10, LM89, MDW11, MOvdD90, McD92b, MY12, Min84, OWTZ21, Pal79, Pat90, dSP17, PRAC22, RM22, Rob77a, Slo16b, Sło18b, SC22, SH92, Wan17a, WS23, ZWW10, ZWZ10, ZWM17b]. zero-divisor [DJS21, OWTZ21, ZWM17b]. zero-divisors [CG19]. Zero-nonzero [GGH22, CM07, KMOV07, MY12]. zero-patterns [JL93]. zero-set [FGMSS19]. Zero-sets [Soa08]. zero-star [BS16]. Zero-term [BSL01, BS16]. zeroes [LR96]. Zeros [AGZ00, LT07, AS12, Dat80, FMSS19, KS07, KK90, Mel13a]. Zeta [CP10, KKS22, MMS21, Sat11, SMM14, SMM16]. Zhan [Zou20]. Zhang [Uhl20]. Zierler [GTLN18]. Zinbiel [Ano22b, CKKK22a, KPPV18]. Zusmanovich [LHZ22]. Zwahlen [TT74].

References

Audenaert:2012:NSA

[AA12] Koenraad M. R. Audenaert and Jaspal Singh Aujla. On norm sub-additivity and super-additivity inequalities for concave and convex functions. Linear Multilinear Algebra, 60(11–12): 1369–1389, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ayupov:2020:DLL

[AA20] Sh. A. Ayupov and F. N. Arzikulov. Description of 2-local and local derivations on some Lie rings of skew-adjoint matrices. Linear Multilinear Algebra, 68(4):764–780, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdelali:2021:MSP

[AA21a] Zine El Abidine Abdelali and Bouchra Aharmim. Multiplicatively spectrum preserving maps on rectangular matrices. Linear Multilinear Algebra, 69(16):3099–3111, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Afshari:2021:NAC

[AA21b] B. Afshari and S. Akbari. A note on the algebraic connectivity of a graph and its complement. Linear Multilinear Algebra, 69(7):1248–1254, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arancibia:2022:DGL

[AAA22] B. Alfaro Arancibia, M. A. Alvarez, and Y. Anza. Degenerations of graph Lie algebras. Linear Multilinear Algebra, 70(1): 91–100, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akkurt:2015:JHS

[AAB15] E. Akkurt, M. Akkurt, and G. P. Barker. Jordan homomorphisms of the structural matrix algebras. Linear Multilinear Algebra, 63(12):2518–2525, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ahmadi-Asl:2017:EIA

[AAB17] Salman Ahmadi-Asl and Fatemeh Panjeh Ali Beik. An efficient iterative algorithm for quaternionic least-squares problems over the generalized -(anti-)bi-Hermitian matrices. Linear Multilinear Algebra, 65(9):1743–1769, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ali:2020:IML

[AAB20] Patrick Ali, Fouzul Atik, and Ravindra B. Bapat. Identities for minors of the Laplacian, resistance and distance matrices of graphs with arbitrary weights. Linear Multilinear Algebra, 68(2):323–336, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2022:SPE

[AAD22] Saieed Akbari, Jalal Askari, and Kinkar Chandra Das. Some properties of eigenvalues of the Seidel matrix. Linear Multilinear Algebra, 70(11):2150–2161, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Albas:2018:ECG

[AADD18] E. Albas, N. Argaç, V. De Filippis, and Ç. Demir. An Engel condition with generalized skew derivations on multilinear polynomials. Linear Multilinear Algebra, 66(10):1925–1938, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alazemi:2023:CGS

[AADdF23] Abdullah Alazemi, Milica Andelić, Kinkar Chandra Das, and Carlos M. da Fonseca. Chain graph sequences and Laplacian spectra of chain graphs. Linear Multilinear Algebra, 71(4): 569–585, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Andelic:2019:VTS

[AAdFS19] Milica Andelić, Firouzeh Ashraf, Carlos M. da Fonseca, and Slobodan K. Simić. Vertex types in some lexicographic products of graphs. Linear Multilinear Algebra, 67(11):2282–2296, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2022:EPE

[AAGL22] Saieed Akbari, Seyran Azizi, Modjtaba Ghorbani, and Xueliang Li. On edge-path eigenvalues of graphs. Linear Multilinear Algebra, 70(15):2998–3008, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alahmadi:2012:SSN

[AAJ12a] Adel N. Alahmadi, Yousef Alkhamees, and S. K. Jain. On semigroups and semirings of nonnegative matrices. Linear Multilinear Algebra, 60(5):595–598, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alahmedi:2012:MNM

[AAJ12b] Adel Alahmedi, Yousef Alkhamees, and S. K. Jain. Monotonicity of nonnegative matrices. Linear Multilinear Algebra, 60(7): 855–864, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aghamollaei:2016:SRN

[AAK16] Gholamreza Aghamollaei, Narges Haj Aboutalebi, and Hossein Momenaee Kermani. Some results on the k-numerical range of quaternion matrices. Linear Multilinear Algebra, 64 (12):2419–2430, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alazemi:2021:EFI

[AAKS21] Abdullah Alazemi, Milica Andelić, Tamara Koledin, and Zoran Stanić. Eigenvalue-free intervals of distance matrices of threshold and chain graphs. Linear Multilinear Algebra, 69 (16):2959–2975, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akemann:1982:TIO

[AAP82] Charles A. Akemann, Joel Anderson, and Gertk Pedersen. Triangle inequalities in operator algebras. Linear and Multilinear

Algebra, 11(2):167–178, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Albuquerque:2019:NMS

[AAPR19] N. Albuquerque, G. Araújo, D. Pellegrino, and P. Rueda. A note on multiple summing operators and applications. Linear Multilinear Algebra, 67(4):660–671, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aalipour:2014:LSC

[AAS14] G. Aalipour, S. Akbari, and N. Shajari. Laplacian spectral characterization of two families of trees. Linear Multilinear Algebra, 62(7):965–977, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2016:PMR

[AAT16] S. Akbari, M. Ariannejad, and Z. Tajfirouz. Permanents of matrices over roots of unity. Linear Multilinear Algebra, 64(9): 1769–1775, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aaghabali:2017:OLS

[AAT17] M. Aaghabali, S. Akbari, and Z. Tajfirouz. Order of the largest Sachs subgraphs in graphs. Linear Multilinear Algebra, 65(1): 204–209, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ayupov:2022:DLD

[AAUN22] Sh. A. Ayupov, F. N. Arzikulov, N. M. Umrzaqov, and O. O. Nuriddinov. Description of 2-local derivations and automorphisms on finite-dimensional Jordan algebras. Linear Multilinear Algebra, 70(18):3525–3542, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anderson:1978:GRI

[AB78] Ned Anderson and Gilbert G. Best. A Gerschgorin–Rayleigh inequality for the eigenvalues of Hermitian matrices. Linear and Multilinear Algebra, 6(3):219–222, 1978. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arambasic:2017:FOF

[AB17] Lj. Arambasić and D. Bakić. Frames and outer frames for Hilbert-modules. Linear Multilinear Algebra, 65(2):381–431,

2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arambasic:2019:FSF

[AB19] Lj. Arambasić and D. Bakić. Full spark frames and totally positive matrices. Linear Multilinear Algebra, 67(8):1685–1700, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ayadi:2022:PLR

[AB22] Mohamed Ayadi and Hamadi Baklouti. Positive linear relation and application to domination problem. Linear Multilinear Algebra, 70(19):3780–3789, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdelmoula:2023:ADR

[Ab23] Salmine Abdelmoula and Hamadi baklouti. Asymptotic distribution of resonances for matrix Schrödinger operator in one dimension. Linear Multilinear Algebra, 71(11):1875–1887, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akel:2023:DGM

[ABAH23] Mohamed Akel, Ahmed Bakhet, Mohamed Abdalla, and Fuli He. On degenerate gamma matrix functions and related functions. Linear Multilinear Algebra, 71(4):673–691, 2023. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Angelini:2018:RIV

[ABC18] Elena Angelini, Cristiano Bocci, and Luca Chiantini. Real identifiability vs. complex identifiability. Linear Multilinear Algebra, 66(6):1257–1267, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdulrahim:2009:TTI

[Abd09] Mohammad N. Abdulrahim. Tensoring three irreducible linear representations of the braid group. Linear Multilinear Algebra, 57(8):749–758, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdalla:2020:SMF

[Abd20] Mohamed Abdalla. Special matrix functions: characteristics, achievements and future directions. Linear Multilinear Algebra, 68(1):1–28, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alhevaz:2021:BTC

[ABGP21] A. Alhevaz, M. Baghipur, Hilal A. Ganie, and S. Pirzada. Brouwer type conjecture for the eigenvalues of distance signless Laplacian matrix of a graph. Linear Multilinear Algebra, 69(13):2423–2440, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ammar:2016:NEP

[ABJ16] Aymen Ammar, Bilel Boukettaya, and Aref Jeribi. A note on the essential pseudospectra and application. Linear Multilinear Algebra, 64(8):1474–1483, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arfa:2022:THJ

[ABM22] Anja Arfa, Abdelkader Ben Hassine, and Sami Mabrouk. Ternary Hom–Jordan algebras induced by Hom–Jordan algebras. Linear Multilinear Algebra, 70(19):3944–3968, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aboutalebi:2020:INR

[Abo20] Narges Haj Aboutalebi. Introduction to C-numerical ranges and C-radii of quaternion matrices. Linear Multilinear Algebra, 68(6):1230–1238, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alhevaz:2022:SEM

[ABP22] Abdollah Alhevaz, Maryam Baghipur, and Shariefuddin Pirzada. On the smallest eigenvalue of D_{α}-matrix of connected graphs. Linear Multilinear Algebra, 70(21):6478–6499, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anh:2016:ISR

[ÁBvW16] P. N. Ánh, G. F. Birkenmeier, and L. van Wyk. Idempotents and structures of rings. Linear Multilinear Algebra, 64(10):

2002–2029, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Agrawala:1978:EIR

[AC78] Vishnu K. Agrawala and Chong-Yun Yun Chao. Equalities and inequalities for ranks of modules. Linear and Multilinear Algebra, 6(4):307–315, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alnajjar:2013:BTL

[AC13] Hasan Alnajjar and Brian Curtin. A bidiagonal and tridiagonal linear map with respect to eigenbases of equitable basis of sl_{2}. Linear Multilinear Algebra, 61(12):1668–1674, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alvarez:2019:CDH

[AC19] María Alejandra Alvarez and Francisco Cartes. Cohomology and deformations for the Heisenberg Hom–Lie algebras. Linear Multilinear Algebra, 67(11):2209–2229, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

An:2021:DDM

[AC21] Runling An and Yaru Cai. Derivations and derivable maps on von Neumann algebras. Linear Multilinear Algebra, 69(15): 2806–2812, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Angelini:2023:DIQ

[AC23] Elena Angelini and Luca Chiantini. On the description of identifiable quartics. Linear Multilinear Algebra, 71(7):1098– 1126, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cai:2017:QWM

[aCC17] Yan an Cai and Qiufan Chen. Quasi-Whittaker modules over the conformal Galilei algebras. Linear Multilinear Algebra, 65(2):313–324, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Achilles:1977:PDS

[Ach77] Eva Achilles. Permanents of doubly stochastic matrices with fixed zero pattern. Linear and Multilinear Algebra, 5(1):63–70,

Abdikalykov:2015:UCA

[ACI15] A. K. Abdikalykov, V. N. Chugunov, and Kh. D. Ikramov. Unitary congruence automorphisms of the space of Toeplitz matrices. Linear Multilinear Algebra, 63(6):1195–1203, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aryal:2017:HMS

[ACJ17] Saroj Aryal, Hayoung Choi, and Farhad Jafari. Hamburger moment sequences and their moment subsequences. Linear Multilinear Algebra, 65(9):1838–1851, 2017. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Andrade:2017:BGA

[ACMR17] Enide Andrade, Domingos M. Cardoso, Luis Medina, and Oscar Rojo. Bethe graphs attached to the vertices of a connected graph — a spectral approach. Linear Multilinear Algebra, 65(4):857–868, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Antezana:2010:NSO

[ACMS10] J. Antezana, C. Cano, I. Mosconi, and D. Stojanoff. A note on the star order in Hilbert spaces. Linear Multilinear Algebra, 58(8):1037–1051, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Adashev:2021:SLA

[ACO21] J. Q. Adashev, L. M. Camacho, and B. A. Omirov. Solvable Leibniz algebras with naturally graded non-Lie p-filiform nilradicals whose maximal complemented space of its nilradical. Linear Multilinear Algebra, 69(8):1500–1520, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Andrade:2017:CPV

[AD17] Enide Andrade and Geir Dahl. Combinatorial Perron values of trees and bottleneck matrices. Linear Multilinear Algebra, 65(12):2387–2405, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Andelic:2021:SDC

[AdF21] Milica Andelić and Carlos M. da Fonseca. Some determinantal considerations for pentadiagonal matrices. Linear Multilinear Algebra, 69(16):3121–3129, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alzer:2015:YTI

[AdFK15] Horst Alzer, Carlos M. da Fonseca, and Alexander Kovacec. Young-type inequalities and their matrix analogues. Linear Multilinear Algebra, 63(3):622–635, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Adhikari:2013:VSL

[Adh13] Bibhas Adhikari. Vector space of linearizations for the quadratic two-parameter matrix polynomial. Linear Multilinear Algebra, 61(5):603–616, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2019:ESC

[ADHM19] S. Akbari, S. Dalvandi, F. Heydari, and M. Maghasedi. On the eigenvalues of signed complete graphs. Linear Multilinear Algebra, 67(3):433–441, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Al-Dolat:2022:NNR

[ADJAH22] Mohammed Al-Dolat, Imad Jaradat, and Baráa Al-Husban. A novel numerical radius upper bounds for 2 × 2 operator matrices. Linear Multilinear Algebra, 70(6):1173–1184, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Adkins:1995:FPI

[Adk95] William A. Adkins. The fan-pall imbedding theorem over formally real fields. Linear and Multilinear Algebra, 39(3):273– 278, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Achour:2014:FSC

[ADRP14] D. Achour, E. Dahia, P. Rueda, and E. A. Sánchez Pérez. Factorization of strongly (p, σ)-continuous multilinear operators. Linear Multilinear Algebra, 62(12):1649–1670, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Adukov:2015:RTF

[Adu15] Victor Adukov. Representation theory of finite groups in Wiener–Hopf factorization problem. Linear Multilinear Algebra, 63(9):1724–1736, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alizadeh:2010:EUI

[AE10] R. Alizadeh and G. H. Esslamzadeh. On the existence of unitarily invariant norm under some conditions. Linear Multilinear Algebra, 58(3):367–375, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abd-Elmageed:2020:SRF

[AEAAES20] H. Abd-Elmageed, M. Abdalla, M. Abul-Ez, and Nasser Saad. Some results on the first Appell matrix function F_{1}(A, B, B′, C; z, w). Linear Multilinear Algebra, 68(2):278– 292, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Andelic:2013:NAMa

[AEdF13a] Milica Andelić, Aleksandra Erić, and C. M. da Fonseca. Nonsingular acyclic matrices with full number of P -vertices. Linear Multilinear Algebra, 61(1):49–57, 2013. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Andelic:2013:NAMb

[AEdF13b] Milica Andelic, Aleksandra Erić, and C. M. da Fonseca. Nonsingular acyclic matrices with full number of P -vertices. Linear Multilinear Algebra, 61(8):1159–1160, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ashrafi:2019:LCZ

[AEG19] A. R. Ashrafi, M. Eliasi, and A. Ghalavand. Laplacian coefficients and Zagreb indices of trees. Linear Multilinear Algebra, 67(9):1736–1749, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Argac:2020:IIM

[AELL20] N. Argaç, M. P. Eroğlu, T.-K. Lee, and J.-H. Lin. Identities with inverses on matrix rings. Linear Multilinear Algebra, 68(3):635–651, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See correction [Ano20a].

Al-Essa:2018:MID

[AELPM18] L. M. Al-Essa, S. R. López-Permouth, and N. M. Muthana. Modules over infinite-dimensional algebras. Linear Multilinear Algebra, 66(3):488–496, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alaminos:2012:MVH

[AEV12] J. Alaminos, J. Extremera, and A. R. Villena. Metric versions of Herstein’s theorems on Jordan maps. Linear Multilinear Algebra, 60(8):941–954, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ablamowicz:2011:TAIa

[AF11a] Rafal Ablamowicz and Bertfried Fauser. On the transposition anti-involution in real Clifford algebras i: the transposition map. Linear Multilinear Algebra, 59(12):1331–1358, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ablamowicz:2011:TAIb

[AF11b] Rafal Ablamowicz and Bertfried Fauser. On the transposition anti-involution in real Clifford algebras II: stabilizer groups of primitive idempotents. Linear Multilinear Algebra, 59(12): 1359–1381, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ablamowicz:2012:TAI

[AF12] Rafal Ablamowicz and Bertfried Fauser. On the transposition anti-involution in real Clifford algebras III: the automorphism group of the transposition scalar product on spinor spaces. Linear Multilinear Algebra, 60(6):621–644, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Azadikhouy:2022:SCO

[AFA22] Masoumeh Azadikhouy and Masoumeh Faghih-Ahmadi. Superconvex-cyclicity of operators. Linear Multilinear Algebra, 70(21):6989–6999, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abreu:2017:CSC

[AFJK17] Nair Abreu, Eliseu Fritscher, Claudia Justel, and Steve Kirkland. On the characteristic set, centroid, and centre for a tree.

Linear Multilinear Algebra, 65(10):2046–2063, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alvarez:2013:CFL

[AFM13] Teresa Alvarez, Fatma Fakhfakh, and Maher Mnif. Coperturbation function and lower semi-Browder multivalued linear operators. Linear Multilinear Algebra, 61(4):494–516, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Afriat:1973:RCF

[Afr73] S. N. Afriat. On the rational canonical form of a matrix. Linear and Multilinear Algebra, 1(3):185–186, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2019:IUP

[AFTG19] S. Akbari, G. H. Fath-Tabar, and E. Ghasemian. An inequality using perfect matchings and Laplacian spread of a graph. Linear Multilinear Algebra, 67(3):442–447, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abrams:2000:RPM

[AFVL00] Lowell Abrams, Donniell E. Fishkind, and Silvia Valdes-Leon. Reflecting the Pascal matrix about its main antidiagonal. Linear and Multilinear Algebra, 47(2):129–136, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ali:2022:PGC

[AFW22] Fawad Ali, Saba Fatima, and Wei Wang. On the power graphs of certain finite groups. Linear Multilinear Algebra, 70(19): 3803–3817, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arens:1996:WNM

[AG96] Richard Arens and Moshe Goldberg. Weighted l_{1} norms for matrices. Linear and Multilinear Algebra, 40(3):229–234, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ambrozie:2015:IPC

[AG15] Calin-Grigore Ambrozie and Aurelian Gheondea. An interpolation problem for completely positive maps on matrix algebras: solvability and parametrization. Linear Multilinear

Adm:2016:ISS

[AG16] Mohammad Adm and Jürgen Garloff. Intervals of special sign regular matrices. Linear Multilinear Algebra, 64(7):1424–1444, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Andruskiewitsch:2018:NAQ

[AG18] Nicolás Andruskiewitsch and João Matheus Jury Giraldi. Nichols algebras that are quantum planes. Linear Multilinear Algebra, 66(5):961–991, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abbas:2022:MRR

[AG22] Hassane Abbas and Mohammad M. Ghabries. More results related to geometric mean and singular values for matrices. Linear Multilinear Algebra, 70(5):878–887, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arens:1996:JMN

[AGAG96] Richard Arens, Moshe Goldberg, R. Arens, and M. Goldberg. Jordan-multiplicative norms. Linear and Multilinear Algebra, 41(1):9–18, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2022:SDG

[AGH^{+}22] S. Akbari, A. H. Ghodrati, M. A. Hosseinzadeh, V. V. Kabanov, E. V. Konstantinova, and L. V. Shalaginov. Spectra of Deza graphs. Linear Multilinear Algebra, 70(2):310–321, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2013:NZE

[AGM13] S. Akbari, E. Ghorbani, and A. Mahmoodi. Nowhere-zero eigenvectors of graphs. Linear Multilinear Algebra, 61(2):273– 279, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alves:2022:MPB

[AGP22] M. B. Alves, J. B. Gomes, and K. M. Pedroso. On minimal positive basis for polyhedral cones in R^{n}. Linear Multilinear

Alsina:1991:FER

[AGR91] C. Alsina and J. L. Garcia-Roig. On a functional equation related to the exterior powers of Euclidean space. Linear and Multilinear Algebra, 30(1–2):77–79, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2020:NZB

[AGS20] Saieed Akbari, Amir Hossein Ghodrati, and Shahriar Shahriari. Nowhere-zero bases for the nullspace of the incidence matrices of graphs. Linear Multilinear Algebra, 68(8):1642–1654, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alaminos:2019:OAP

[AGV19] J. Alaminos, M. L. C. Godoy, and A. R. Villena. Orthogonally additive polynomials on the algebras of approximable operators. Linear Multilinear Algebra, 67(9):1922–1936, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aron:2000:ZRP

[AGZ00] Richard Aron, Raquel Gonzalo, and Andriy Zagorodnyuk. Zeros of real polynomials. Linear and Multilinear Algebra, 48(2): 107–115, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

An:2010:CJD

[AH10] Runling An and Jinchuan Hou. Characterizations of Jordan derivations on rings with idempotent. Linear Multilinear Algebra, 58(6):753–763, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

An:2013:CDR

[AH13] Runling An and Jinchuan Hou. Characterization of derivations on reflexive algebras. Linear Multilinear Algebra, 61(10):1408– 1418, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aouchiche:2016:DSL

[AH16a] Mustapha Aouchiche and Pierre Hansen. On the distance signless Laplacian of a graph. Linear Multilinear Algebra, 64(6): 1113–1123, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Audenaert:2016:RLT

[AH16b] Koenraad M. R. Audenaert and Fumio Hiai. Reciprocal Lie– Trotter formula. Linear Multilinear Algebra, 64(6):1220–1235, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alvarez:2020:DLS

[AH20] María Alejandra Alvarez and Isabel Hernández. On degenerations of Lie superalgebras. Linear Multilinear Algebra, 68 (1):29–44, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aharoni:1976:TDK

[Aha76] Ron Aharoni. On theorem of Dénes König. Linear and Multilinear Algebra, 4(1):31–32, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ando:1987:SVH

[AHJ87] T. Ando, Roger A. Horn, and Charles R. Johnson. The singular values of a Hadamard product: a basic inequality. Linear and Multilinear Algebra, 21(4):345–365, 1987. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ahmad:2020:CWG

[AHJ20] Uzma Ahmad, Saira Hameed, and Shaista Jabeen. Class of weighted graphs with strong anti-reciprocal eigenvalue property. Linear Multilinear Algebra, 68(6):1129–1139, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ahmad:2021:NGS

[AHJ21] Uzma Ahmad, Saira Hameed, and Shaista Jabeen. Noncorona graphs with strong anti-reciprocal eigenvalue property. Linear Multilinear Algebra, 69(10):1878–1888, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Agha:1992:SIH

[AHK92] Nader Agha, Daniel Hershkowitz, and Nataly Kogan. A solution to an inverse height and level characteristics problem. Linear and Multilinear Algebra, 31(1–4):139–146, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alrimawi:2019:NII

[AHK19] Fadi Alrimawi, Omar Hirzallah, and Fuad Kittaneh. Norm inequalities involving convex and concave functions of operators. Linear Multilinear Algebra, 67(9):1757–1772, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anderson:1985:EDP

[AHT85] William N. Anderson, Jr., E. James Harner, and George E. Trapp. Eigenvalues of the difference and product of projections. Linear and Multilinear Algebra, 17(3–4):295–299, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aitken:2019:NDH

[Ait19] Colin Aitken. Non-degenerate 2 × k × (k + 1) hypermatrices. Linear Multilinear Algebra, 67(4):697–704, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alahmadi:2014:DSM

[AJL14] Adel Alahmadi, S. K. Jain, and André Leroy. Decomposition of singular matrices into idempotents. Linear Multilinear Algebra, 62(1):13–27, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alahmadi:2018:WNM

[AJL18] A. Alahmadi, S. K. Jain, and A. Leroy. When are nonnegative matrices product of nonnegative idempotent matrices? Linear Multilinear Algebra, 66(4):748–758, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ammar:2022:PSR

[AJL22] Aymen Ammar, Aref Jeribi, and Nawrez Lazrag. Pseudo Sspectrum in a right quaternionic Hilbert space. Linear Multilinear Algebra, 70(4):581–605, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Amiad:1978:BR

[AK78] Rivca Amiad and Henry F. Kaiser. Book review. Linear and Multilinear Algebra, 6(2):157–159, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ando:1994:PMD

[AK94] T. Ando and F. Kubo. Product of matrical discs. Linear and Multilinear Algebra, 37(1–3):125–129, 1994. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ayupov:2021:LDS

[AK21] Sh. A. Ayupov and A. Kh. Khudoyberdiyev. Local derivations on solvable Lie algebras. Linear Multilinear Algebra, 69(7): 1286–1301, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Atik:2021:DLM

[AKB21] Fouzul Atik, M. Rajesh Kannan, and Ravindra B. Bapat. On distance and Laplacian matrices of trees with matrix weights. Linear Multilinear Algebra, 69(14):2607–2619, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akemann:1980:BRa

[Ake80] Charles A. Akemann. Book reviews. Linear and Multilinear Algebra, 8(4):357–358, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akemann:1983:SRI

[Ake83] Charles A. Akemann. Subalgebras and right ideals in matrix algebras. Linear and Multilinear Algebra, 12(4):311–313, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Al-khlyleh:2017:IIR

[AkK17] Mohammad Al-khlyleh and Fuad Kittaneh. Interpolating inequalities related to a recent result of Audenaert. Linear Multilinear Algebra, 65(5):922–929, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aliabadi:2023:CMG

[AKK^{+}23] Mohsen Aliabadi, Jack Kinseth, Christopher Kunz, Haris Serdarevic, and Cole Willis. Conditions for matchability in groups and field extensions. Linear Multilinear Algebra, 71(7):1182– 1197, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ayupov:2020:SLA

[AKOZ20] Shavkat Ayupov, Karimbergen Kudaybergenov, Bakhrom Omirov, and Kaiming Zhao. Semisimple Leibniz algebras, their derivations and automorphisms. Linear Multilinear Algebra, 68(10):2005–2019, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdollahi:2023:NUE

[AKR23] O. Abdollahi, S. Karami, and J. Rooin. New unitary equivalences for some operator matrices with applications. Linear Multilinear Algebra, 71(12):1957–1978, 2023. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Asadi:2018:CSZ

[AKS18] Mohammad B. Asadi, Hossein Khodaiemehr, and Fereshteh Sady. Completely square-zero preserving maps on B_{s}(H). Linear Multilinear Algebra, 66(5):937–941, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ayupov:2021:LDG

[AKY21] Shavkat Ayupov, Karimbergen Kudaybergenov, and Baxtiyor Yusupov. 2-local derivations on generalized Witt algebras. Linear Multilinear Algebra, 69(16):3130–3140, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Avraham:1978:ILT

[AL78] Sara Avraham and Raphael Loewy. On the inertia of the Lyapunov transform AH+HA^{∗} for H > 0. Linear and Multilinear Algebra, 6(1):1–21, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ando:1994:I

[AL94] T. Ando and C. K. Li. Introduction. Linear and Multilinear Algebra, 37(1–3):1–2, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ali:2021:CSR

[AL21] Fawad Ali and Yongjin Li. The connectivity and the spectral radius of commuting graphs on certain finite groups. Linear

Multilinear Algebra, 69(16):2945–2958, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdi:2022:GCC

[AL22] M. Abdi and A. Leroy. Graphs of commutatively closed sets. Linear Multilinear Algebra, 70(21):6965–6977, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alakhrass:2016:IRH

[Ala16] Mohammad Alakhrass. Inequalities related to Heinz mean. Linear Multilinear Algebra, 64(8):1562–1569, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alakhrass:2018:NAI

[Ala18] Mohammad Alakhrass. A note on Audenaert interpolation inequality. Linear Multilinear Algebra, 66(9):1909–1916, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alakhrass:2020:NSM

[Ala20] Mohammad Alakhrass. A note on sectorial matrices. Linear Multilinear Algebra, 68(11):2228–2238, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Allahverdiev:2014:ESI

[All14] Bilender P. Allahverdiev. Extensions of symmetric infinite Jacobi operator. Linear Multilinear Algebra, 62(9):1146–1152, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alnajjar:2011:LPA

[Aln11] Hasan Alnajjar. Leonard pairs associated with the equitable generators of the quantum algebra U_{q}(sl_{2}). Linear Multilinear Algebra, 59(10):1127–1142, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alnajjar:2015:TLM

[Aln15] Hasan Alnajjar. A tridiagonal linear map with respect to eigenbases of equitable basis of sl_{2}. Linear Multilinear Algebra, 63(7):1458–1467, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alomari:2021:RSN

[Alo21] Mohammad W. Alomari. Refinements of some numerical radius inequalities for Hilbert space operators. Linear Multilinear Algebra, 69(7):1208–1223, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alomari:2023:DWR

[Alo23] M. W. Alomari. On the Davis–Wielandt radius inequalities of Hilbert space operators. Linear Multilinear Algebra, 71(11): 1804–1828, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Axelsson:1994:NRM

[ALP94] O. Axelsson, H. Lu, and B. Polman. On the numerical radius of matrices and its application to iterative solution methods. Linear and Multilinear Algebra, 37(1–3):225–238, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alperin:1995:FPM

[Alp95] Roger C. Alperin. Free products of matrices. Linear and Multilinear Algebra, 38(3):215–223, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arenas:2022:LVC

[ALP22] Manuel Arenas, Alicia Labra, and Irene Paniello. Lotka– Volterra coalgebras. Linear Multilinear Algebra, 70(19):4483– 4497, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aydogdu:2020:IDA

[ALPM20] Pinar Aydogdu, Sergio R. López-Permouth, and Rebin A. Muhammad. Infinite-dimensional algebras without simple bases. Linear Multilinear Algebra, 68(12):2390–2407, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anderson:1985:ELG

[AM85] William N. Anderson, Jr. and Thomas D. Morley. Eigenvalues of the Laplacian of a graph. Linear and Multilinear Algebra, 18(2):141–145, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Azenhas:1990:MRL

[AM90] Olga Azenhas and Eduardo Marques De Sá. Matrix realizations of Littlewood–Richardson sequences. Linear and Multilinear Algebra, 27(4):229–242, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ahmad:2013:BEE

[AM13] Sk. Safique Ahmad and Volker Mehrmann. Backward errors for eigenvalues and eigenvectors of Hermitian, skew-Hermitian, H-even and H-odd matrix polynomials. Linear Multilinear Algebra, 61(9):1244–1266, 2013. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Agore:2015:ITM

[AM15] A. L. Agore and G. Militaru. Itô’s theorem and metabelian Leibniz algebras. Linear Multilinear Algebra, 63(11):2187– 2199, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Argerami:2021:HRN

[AM21] Martín Argerami and Saleh Mustafa. Higher rank numerical ranges of Jordan-like matrices. Linear Multilinear Algebra, 69(5):807–826, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Amitsur:1980:CPS

[Ami80] S. A. Amitsur. On the characteristic polynomial of a sum of matrices. Linear and Multilinear Algebra, 8(3):177–182, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Amiraslani:2011:NLF

[Ami11] A. Amiraslani. Notes on linear factor polynomial deflation in polynomial bases. Linear Multilinear Algebra, 59(12):1419– 1431, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2023:SIE

[AMK23] Saieed Akbari, Kasra Masoudi, and Sina Kalantarzadeh. A strict inequality on the energy of edge partitioning of graphs. Linear Multilinear Algebra, 71(11):1858–1862, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ammar:2018:SRS

[Amm18] Aymen Ammar. Some results on semi-Fredholm perturbations of multivalued linear operators. Linear Multilinear Algebra, 66(7):1311–1332, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Adam:2002:NRR

[AMP02] M. Adam, J. Maroulas, and P. Psarrakos. On the numerical range of rational matrix functions. Linear Multilinear Algebra, 50(1):75–89, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Amparan:2013:SIP

[AMZ13] A. Amparan, S. Marcaida, and I. Zaballa. On the structure invariants of proper rational matrices with prescribed finite poles. Linear Multilinear Algebra, 61(11):1464–1486, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbari:2014:SRI

[AN14] S. Akbari and R. Nikandish. Some results on the intersection graph of ideals of matrix algebras. Linear Multilinear Algebra, 62(2):195–206, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdollahi:2018:FG

[AN18a] Farshid Abdollahi and Hashem Najafi. Frame graph. Linear Multilinear Algebra, 66(6):1229–1243, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

An:2018:CJH

[An18b] Guangyu An. Characterizations of n-Jordan homomorphisms. Linear Multilinear Algebra, 66(4):671–680, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdelali:2019:MPP

[AN19] Zine El Abidine Abdelali and Hamid Nkhaylia. Maps preserving the pseudo spectrum of skew triple product of operators. Linear Multilinear Algebra, 67(11):2297–2306, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ando:1980:HMI

[And80a] T. Ando. Hua–Marcus inequalities. Linear and Multilinear Algebra, 8(4):347–352, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ando:1980:IM

[And80b] T. Ando. Inequalities for M-matrices. Linear and Multilinear Algebra, 8(4):291–316, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ando:1996:QOP

[And96] T. Ando. Quasi-orders on the positive cone of a C. Linear and Multilinear Algebra, 41(1):81–94, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ando:2004:ATC

[And04] T. Ando. Aluthge transforms and the convex hull of the eigenvalues of a matrix. Linear Multilinear Algebra, 52(3–4):281– 292, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anghel:2000:CMP

[Ang00] Nicolae Anghel. Clifford matrices and a problem of Hurwitz. Linear and Multilinear Algebra, 47(2):105–117, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anjidani:2017:OIJ

[Anj17] Ehsan Anjidani. On operator inequalities of Jensen type for convex functions. Linear Multilinear Algebra, 65(7):1493– 1502, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Al-Natoor:2022:SVN

[ANK22] Ahmad Al-Natoor and Fuad Kittaneh. Singular value and norm inequalities for positive semidefinite matrices. Linear Multilinear Algebra, 70(19):4498–4507, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1973:EN

[Ano73] Anonymous. Editor’s note. Linear and Multilinear Algebra, 1(3):284, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1979:EB

[Ano79a] Anonymous. Editorial board. Linear and Multilinear Algebra, 8(1):ebi, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1979:ECR

[Ano79b] Anonymous. Erratum: Charles R. Johnson, R. B. Kellogg, and A. B. Stephens. Complex eigenvalues of a non-negative matrix with a specified graph II, Linear and Multilinear Algebra 7(2) 129–143 (1979). Linear and Multilinear Algebra, 8 (2):171, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). Correction of Figure 6 in [JKS79].

Anonymous:1979:SSB

[Ano79c] Anonymous. A second Santa Barbara matrix conference. Linear and Multilinear Algebra, 7(4):265, 1979. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1980:EBa

[Ano80a] Anonymous. Editorial board. Linear and Multilinear Algebra, 8(3):ebi, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1980:EBb

[Ano80b] Anonymous. Editorial board. Linear and Multilinear Algebra, 8(4):ebi, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1980:EBc

[Ano80c] Anonymous. Editorial board. Linear and Multilinear Algebra, 9(1):ebi, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1980:EBd

[Ano80d] Anonymous. Editorial board. Linear and Multilinear Algebra, 9(2):ebi, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1980:EBe

[Ano80e] Anonymous. Editorial board. Linear and Multilinear Algebra, 9(3):ebi, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1981:EBa

[Ano81a] Anonymous. Editorial board. Linear and Multilinear Algebra, 9(4):ebi, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1981:EBb

[Ano81b] Anonymous. Editorial board. Linear and Multilinear Algebra, 10(1):ebi, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1981:EBc

[Ano81c] Anonymous. Editorial board. Linear and Multilinear Algebra, 10(2):ebi, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1981:EBd

[Ano81d] Anonymous. Editorial board. Linear and Multilinear Algebra, 10(3):ebi, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1981:EBe

[Ano81e] Anonymous. Editorial board. Linear and Multilinear Algebra, 10(4):ebi, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1982:EBa

[Ano82a] Anonymous. Editorial board. Linear and Multilinear Algebra, 12(1):ebi, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1982:EBb

[Ano82b] Anonymous. Editorial board. Linear and Multilinear Algebra, 12(2):ebi, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1982:EBc

[Ano82c] Anonymous. Editorial board. Linear and Multilinear Algebra, 12(3):ebi, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1982:ENa

[Ano82d] Anonymous. Editorial note. Linear and Multilinear Algebra, 12(2):79, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1982:ENb

[Ano82e] Anonymous. Editorial note. Linear and Multilinear Algebra, 12(3):157, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1982:RPa

[Ano82f] Anonymous. Research problem. Linear and Multilinear Algebra, 11(1):101–102, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1982:RPb

[Ano82g] Anonymous. Research problems. Linear and Multilinear Algebra, 11(3):301, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:BR

[Ano83a] Anonymous. Book review. Linear and Multilinear Algebra, 13(3):293–294, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBa

[Ano83b] Anonymous. Editorial board. Linear and Multilinear Algebra, 12(4):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBb

[Ano83c] Anonymous. Editorial board. Linear and Multilinear Algebra, 13(1):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBc

[Ano83d] Anonymous. Editorial board. Linear and Multilinear Algebra, 13(2):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBd

[Ano83e] Anonymous. Editorial board. Linear and Multilinear Algebra, 13(3):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBe

[Ano83f] Anonymous. Editorial board. Linear and Multilinear Algebra, 13(4):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBf

[Ano83g] Anonymous. Editorial board. Linear and Multilinear Algebra, 14(1):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBg

[Ano83h] Anonymous. Editorial board. Linear and Multilinear Algebra, 14(2):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBh

[Ano83i] Anonymous. Editorial board. Linear and Multilinear Algebra, 14(3):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:EBi

[Ano83j] Anonymous. Editorial board. Linear and Multilinear Algebra, 14(4):ebi, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:ENa

[Ano83k] Anonymous. Editorial note. Linear and Multilinear Algebra, 13(1):1, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:ENb

[Ano83l] Anonymous. Editorial note. Linear and Multilinear Algebra, 13(3):183, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1983:RP

[Ano83m] Anonymous. Research problem. Linear and Multilinear Algebra, 13(2):177–178, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1984:EBa

[Ano84a] Anonymous. Editorial board. Linear and Multilinear Algebra, 15(2):ebi, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1984:EBb

[Ano84b] Anonymous. Editorial board. Linear and Multilinear Algebra, 15(3–4):ebi, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1984:EBc

[Ano84c] Anonymous. Editorial board. Linear and Multilinear Algebra, 16(1–4):ebi, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1984:EN

[Ano84d] Anonymous. Editorial note. Linear and Multilinear Algebra, 16(1–4):1, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1984:RPa

[Ano84e] Anonymous. Research problems. Linear and Multilinear Algebra, 15(3–4):279–280, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1984:RPb

[Ano84f] Anonymous. Research problems. Linear and Multilinear Algebra, 15(3–4):345–350, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1985:EBa

[Ano85a] Anonymous. Editorial board. Linear and Multilinear Algebra, 17(2):ebi, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1985:EBb

[Ano85b] Anonymous. Editorial board. Linear and Multilinear Algebra, 17(3–4):ebi, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1985:EBc

[Ano85c] Anonymous. Editorial board. Linear and Multilinear Algebra, 18(1):ebi, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1985:EBd

[Ano85d] Anonymous. Editorial board. Linear and Multilinear Algebra, 18(2):ebi, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1985:EBe

[Ano85e] Anonymous. Editorial board. Linear and Multilinear Algebra, 18(3):ebi, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1985:EBf

[Ano85f] Anonymous. Editorial board. Linear and Multilinear Algebra, 18(4):ebi, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1985:RP

[Ano85g] Anonymous. Research problem. Linear and Multilinear Algebra, 18(4):355–356, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1986:EBa

[Ano86a] Anonymous. Editorial board. Linear and Multilinear Algebra, 19(2):ebi, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1986:EBb

[Ano86b] Anonymous. Editorial board. Linear and Multilinear Algebra, 19(3):ebi, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1986:EBc

[Ano86c] Anonymous. Editorial board. Linear and Multilinear Algebra, 19(4):ebi, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1986:EBd

[Ano86d] Anonymous. Editorial board. Linear and Multilinear Algebra, 20(1):ebi, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1986:RP

[Ano86e] Anonymous. Research problem. Linear and Multilinear Algebra, 19(4):405, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1987:EBa

[Ano87a] Anonymous. Editorial board. Linear and Multilinear Algebra, 20(2):ebi, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1987:EBb

[Ano87b] Anonymous. Editorial board. Linear and Multilinear Algebra, 20(3):ebi, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1987:EBc

[Ano87c] Anonymous. Editorial board. Linear and Multilinear Algebra, 20(4):ebi, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1987:EBd

[Ano87d] Anonymous. Editorial board. Linear and Multilinear Algebra, 21(1):ebi, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1987:EBe

[Ano87e] Anonymous. Editorial board. Linear and Multilinear Algebra, 21(2):ebi, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1987:EBf

[Ano87f] Anonymous. Editorial board. Linear and Multilinear Algebra, 21(3):ebi, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1987:EBg

[Ano87g] Anonymous. Editorial board. Linear and Multilinear Algebra, 21(4):ebi, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1988:EBa

[Ano88a] Anonymous. Editorial board. Linear and Multilinear Algebra, 23(1):ebi, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1988:EBb

[Ano88b] Anonymous. Editorial board. Linear and Multilinear Algebra, 23(2):ebi, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1988:EBc

[Ano88c] Anonymous. Editorial board. Linear and Multilinear Algebra, 23(3):ebi, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1988:EBd

[Ano88d] Anonymous. Editorial board. Linear and Multilinear Algebra, 23(4):ebi, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1988:EBe

[Ano88e] Anonymous. Editorial board. Linear and Multilinear Algebra, 24(1):ebi, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1989:BR

[Ano89a] Anonymous. Book review. Linear and Multilinear Algebra, 24 (2):151, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano89b] Anonymous. Editorial board. Linear and Multilinear Algebra, 24(2):ebi, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1989:EBb

[Ano89c] Anonymous. Editorial board. Linear and Multilinear Algebra, 24(3):ebi, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1989:EBc

[Ano89d] Anonymous. Editorial board. Linear and Multilinear Algebra, 24(4):ebi, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1989:EBd

[Ano89e] Anonymous. Editorial board. Linear and Multilinear Algebra, 25(1):ebi, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1989:EBe

[Ano89f] Anonymous. Editorial board. Linear and Multilinear Algebra, 25(2):ebi, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1989:EBf

[Ano89g] Anonymous. Editorial board. Linear and Multilinear Algebra, 25(3):ebi, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1989:EBg

[Ano89h] Anonymous. Editorial board. Linear and Multilinear Algebra, 25(4):ebi, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1990:EBa

[Ano90a] Anonymous. Editorial board. Linear and Multilinear Algebra, 26(1–2):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano90b] Anonymous. Editorial board. Linear and Multilinear Algebra, 26(3):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1990:EBc

[Ano90c] Anonymous. Editorial board. Linear and Multilinear Algebra, 26(4):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1990:EBd

[Ano90d] Anonymous. Editorial board. Linear and Multilinear Algebra, 27(1):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1990:EBe

[Ano90e] Anonymous. Editorial board. Linear and Multilinear Algebra, 27(2):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1990:EBf

[Ano90f] Anonymous. Editorial board. Linear and Multilinear Algebra, 27(3):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1990:EBg

[Ano90g] Anonymous. Editorial board. Linear and Multilinear Algebra, 27(4):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1990:EBh

[Ano90h] Anonymous. Editorial board. Linear and Multilinear Algebra, 28(1–2):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1990:EBi

[Ano90i] Anonymous. Editorial board. Linear and Multilinear Algebra, 28(3):ebi, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano91a] Anonymous. Editorial board. Linear and Multilinear Algebra, 28(4):ebi, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial board. Linear and Multilinear Algebra, 29(1):ebi, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial board. Linear and Multilinear Algebra, 29(2):ebi, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1991:EBd

[Ano91d] Anonymous. Editorial board. Linear and Multilinear Algebra, 29(3–4):ebi, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1991:EBe

[Ano91e] Anonymous. Editorial board. Linear and Multilinear Algebra, 30(1–2):ebi, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1991:EBf

[Ano91f] Anonymous. Editorial board. Linear and Multilinear Algebra, 30(3):ebi, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1991:EBg

[Ano91g] Anonymous. Editorial board. Linear and Multilinear Algebra, 30(4):ebi, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1992:EBa

[Ano92a] Anonymous. Editorial board. Linear and Multilinear Algebra, 31(1–4):ebi, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano92b] Anonymous. Editorial board. Linear and Multilinear Algebra, 32(1):ebi, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1992:EBc

[Ano92c] Anonymous. Editorial board. Linear and Multilinear Algebra, 32(2):ebi, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1992:EBd

[Ano92d] Anonymous. Editorial board. Linear and Multilinear Algebra, 32(3–4):ebi, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1992:EBe

[Ano92e] Anonymous. Editorial board. Linear and Multilinear Algebra, 33(1–2):ebi, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1992:EBf

[Ano92f] Anonymous. Editorial board. Linear and Multilinear Algebra, 33(3–4):ebi, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1992:SLP

[Ano92g] Anonymous. A survey of linear preserver problems contents. Linear and Multilinear Algebra, 33(1–2):1–2, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial board. Linear and Multilinear Algebra, 34(1):ebi, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial board. Linear and Multilinear Algebra, 34(2):ebi, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano93c] Anonymous. Editorial board. Linear and Multilinear Algebra, 34(3–4):ebi, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1993:EBd

[Ano93d] Anonymous. Editorial board. Linear and Multilinear Algebra, 35(1):ebi, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1993:EBe

[Ano93e] Anonymous. Editorial board. Linear and Multilinear Algebra, 35(2):ebi, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1993:EBf

[Ano93f] Anonymous. Editorial board. Linear and Multilinear Algebra, 35(3–4):ebi, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1993:EBg

[Ano93g] Anonymous. Editorial board. Linear and Multilinear Algebra, 36(1):ebi, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1993:EBh

[Ano93h] Anonymous. Editorial board. Linear and Multilinear Algebra, 36(2):ebi, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1994:EBa

[Ano94a] Anonymous. Editorial board. Linear and Multilinear Algebra, 36(3):ebi, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1994:EBb

[Ano94b] Anonymous. Editorial board. Linear and Multilinear Algebra, 36(4):ebi, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano94c] Anonymous. Editorial board. Linear and Multilinear Algebra, 37(1–3):ebi, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1994:EBd

[Ano94d] Anonymous. Editorial board. Linear and Multilinear Algebra, 37(4):ebi, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1994:EBe

[Ano94e] Anonymous. Editorial board. Linear and Multilinear Algebra, 38(1–2):ebi, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial board. Linear and Multilinear Algebra, 38(3):ebi, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial board. Linear and Multilinear Algebra, 38(4):ebi, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1995:EBc

[Ano95c] Anonymous. Editorial board. Linear and Multilinear Algebra, 39(1–2):ebi, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1995:EBd

[Ano95d] Anonymous. Editorial board. Linear and Multilinear Algebra, 39(3):ebi, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1995:EBe

[Ano95e] Anonymous. Editorial board. Linear and Multilinear Algebra, 39(4):ebi, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano95f] Anonymous. Editorial board. Linear and Multilinear Algebra, 40(1):ebi, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1995:EBg

[Ano95g] Anonymous. Editorial board. Linear and Multilinear Algebra, 40(2):ebi, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1996:EBa

[Ano96a] Anonymous. Editorial board. Linear and Multilinear Algebra, 40(3):ebi, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1996:EBb

[Ano96b] Anonymous. Editorial board. Linear and Multilinear Algebra, 40(4):ebi, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1996:EBc

[Ano96c] Anonymous. Editorial board. Linear and Multilinear Algebra, 41(1):ebi, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1996:EBd

[Ano96d] Anonymous. Editorial board. Linear and Multilinear Algebra, 41(2):ebi, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1996:EBe

[Ano96e] Anonymous. Editorial board. Linear and Multilinear Algebra, 41(3):ebi, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1996:EBf

[Ano96f] Anonymous. Editorial board. Linear and Multilinear Algebra, 41(4):ebi, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano97a] Anonymous. Editorial board. Linear and Multilinear Algebra, 43(1–3):ebi, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1997:RCT

[Ano97b] Anonymous. Robert C. Thompson research articles. Linear and Multilinear Algebra, 43(1–3):7–11, 1997. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1998:EBa

[Ano98a] Anonymous. Editorial board. Linear and Multilinear Algebra, 43(4):ebi, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1998:EBb

[Ano98b] Anonymous. Editorial board. Linear and Multilinear Algebra, 44(1):ebi, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1998:EBc

[Ano98c] Anonymous. Editorial board. Linear and Multilinear Algebra, 44(2):ebi, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1998:EBd

[Ano98d] Anonymous. Editorial board. Linear and Multilinear Algebra, 44(3):ebi, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1998:EBe

[Ano98e] Anonymous. Editorial board. Linear and Multilinear Algebra, 44(4):ebi, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1998:EBf

[Ano98f] Anonymous. Editorial board. Linear and Multilinear Algebra, 45(1):ebi, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1998:EBg

[Ano98g] Anonymous. Editorial board. Linear and Multilinear Algebra, 45(2–3):ebi, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1999:EBa

[Ano99a] Anonymous. Editorial board. Linear and Multilinear Algebra, 45(4):ebi, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1999:EBb

[Ano99b] Anonymous. Editorial board. Linear and Multilinear Algebra, 46(1–2):ebi, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1999:EBc

[Ano99c] Anonymous. Editorial board. Linear and Multilinear Algebra, 46(3):ebi, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:1999:EBd

[Ano99d] Anonymous. Editorial board. Linear and Multilinear Algebra, 46(4):ebi, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2000:EBa

[Ano00a] Anonymous. Editorial board. Linear and Multilinear Algebra, 47(1):ebi, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2000:EBb

[Ano00b] Anonymous. Editorial board. Linear and Multilinear Algebra, 47(2):ebi, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2000:EBc

[Ano00c] Anonymous. Editorial board. Linear and Multilinear Algebra, 47(3):ebi, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano00d] Anonymous. Editorial board. Linear and Multilinear Algebra, 47(4):ebi, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2000:EBe

[Ano00e] Anonymous. Editorial board. Linear and Multilinear Algebra, 48(1):ebi, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2000:EBf

[Ano00f] Anonymous. Editorial board. Linear and Multilinear Algebra, 48(2):ebi, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2001:EBa

[Ano01a] Anonymous. Editorial board. Linear and Multilinear Algebra, 48(3):ebi, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2001:EBb

[Ano01b] Anonymous. Editorial board. Linear and Multilinear Algebra, 48(4):ebi, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2001:EBc

[Ano01c] Anonymous. Editorial board. Linear and Multilinear Algebra, 49(1):ebi, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2001:EBd

[Ano01d] Anonymous. Editorial board. Linear and Multilinear Algebra, 49(2):ebi, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2001:EBe

[Ano01e] Anonymous. Editorial board. Linear and Multilinear Algebra, 49(3):ebi, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano01f] Anonymous. Editorial board. Linear and Multilinear Algebra, 49(4):ebi, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2010:TR

[Ano10] Anonymous. Thanks to referees. Linear Multilinear Algebra, 58(4):535, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2011:EB

[Ano11a] Anonymous. Editorial board. Linear Multilinear Algebra, 59 (12):ebi, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2011:TR

[Ano11b] Anonymous. Thanks to referees. Linear Multilinear Algebra, 59(3):355–356, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2012:EB

[Ano12a] Anonymous. Editorial board. Linear Multilinear Algebra, 60 (11–12):ebi, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2012:LR

[Ano12b] Anonymous. List of reviewers. Linear Multilinear Algebra, 60(6):757–759, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2013:EB

[Ano13] Anonymous. Editorial board. Linear Multilinear Algebra, 61 (12):ebi, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2014:EB

[Ano14] Anonymous. Editorial board. Linear Multilinear Algebra, 62 (12):ebi, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

[Ano15] Anonymous. Editorial board. Linear Multilinear Algebra, 63 (12):ebi, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2016:EB

[Ano16a] Anonymous. Editorial board. Linear Multilinear Algebra, 64 (12):ebi, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2016:EBF

[Ano16b] Anonymous. Erratum: Belardo, F., Pisanski, T. and Simić, S. K. (2015). On graphs whose least eigenvalue is greater than −2, Linear and Multilinear Algebra, 64(8) 1570–1582. doi:10.1080/03081087.2015.1107020. Linear Multilinear Algebra, 64(8):i, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [BPS16].

Anonymous:2016:EOV

[Ano16c] Anonymous. Erratum: Oussa, V. (2015). Dihedral group frames which are maximally robust to erasures, Linear and Multilinear Algebra, 63(12), pp. 2494–2510. doi:10.1080/03081087.2015.1020746. Linear Multilinear Algebra, 64(3):572, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Ous15b].

Anonymous:2016:P

[Ano16d] Anonymous. Preface. Linear Multilinear Algebra, 64(1):1, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2017:C

[Ano17] Anonymous. Corrigendum. Linear Multilinear Algebra, 65 (10):2152, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2019:E

[Ano19] Anonymous. Erratum. Linear Multilinear Algebra, 67(3):646, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anonymous:2020:Ca

[Ano20a] Anonymous. Correction. Linear Multilinear Algebra, 68(3):i, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [?].

Anonymous:2020:Cb

[Ano20b] Anonymous. Correction. Linear Multilinear Algebra, 68(4):ii, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [BFM20].

Anonymous:2022:CEC

[Ano22a] Anonymous. Correction: Equivalence classes of e-matrices and associated eigenvalue localization regions. Linear Multilinear Algebra, 70(6):1191, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [MH20].

Anonymous:2022:CDZ

[Ano22b] Anonymous. Corrigendum to ‘Degenerations of Zinbiel and nilpotent Leibniz algebras’. Linear Multilinear Algebra, 70(5): 993–995, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [KPPV18].

Anonymous:2022:ESR

[Ano22c] Anonymous. Erratum to ‘Some results about EP modular operators’. Linear Multilinear Algebra, 70(22):7384–7386, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [KDHJ22].

Anderson:1978:EPP

[AO78] T. W. Anderson and I. Olkin. An extremal problem for positive definite matrices. Linear and Multilinear Algebra, 6(4): 257–262, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ando:1997:HTI

[AO97] T. Ando and K. Okubo. Höder-type inequalities associated with operator radii and Schur products. Linear and Multilinear Algebra, 43(1–3):53–61, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Asadi:2022:OPP

[AO22] Mohammad B. Asadi and Fatemeh Olyaninezhad. Orthogonality preserving pairs of operators on Hilbert C_{0}(Z)-modules. Linear Multilinear Algebra, 70(16):3151–3158, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ayoobi:2011:NGW

[AOTR11] F. Ayoobi, G. R. Omidi, and B. Tayfeh-Rezaie. A note on graphs whose signless Laplacian has three distinct eigenvalues. Linear Multilinear Algebra, 59(6):701–706, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Anderson:1974:CME

[AP74] Joel Anderson and Joe Parker Jr. Choices for a in the matrix equation T=AB-BA. Linear and Multilinear Algebra, 2(3): 203–209, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Adam:2004:CNM

[AP04] Maria Adam and Panayiotis Psarrakos. On a compression of normal matrix polynomials. Linear Multilinear Algebra, 52(3– 4):251–263, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Araujo:2017:OES

[AP17] Gustavo Araújo and Daniel Pellegrino. Optimal estimates for summing multilinear operators. Linear Multilinear Algebra, 65(5):930–942, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ambily:2023:VPO

[AP23] A. A. Ambily and V. K. Aparna Pradeep. Visualization for Petrov’s odd unitary group. Linear Multilinear Algebra, 71 (10):1707–1719, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arela-Perez:2023:ERS

[APEPPS23] S. Arela-Pérez, J. Egaña, G. Pastén, and H. Pickmann-Soto. Extremal realization spectra by two acyclic matrices whose graphs are caterpillars. Linear Multilinear Algebra, 71(10): 1657–1680, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akemann:1980:BRb

[AR80] Charles Akemann and James B. Robertson. Book reviews. Linear and Multilinear Algebra, 9(3):237–239, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arambasic:2015:TCO

[AR15] Lj. Arambasić and R. Rajić. On three concepts of orthogonality in Hilbert C^{∗}-modules. Linear Multilinear Algebra, 63(7): 1485–1500, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ahat:2022:SIM

[AR22] Raxida Ahat and Madi Raikhan. Submajorization inequalities for matrices of τ-measurable operators. Linear Multilinear Algebra, 70(16):3159–3170, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arezoomand:2022:NLP

[Are22] Majid Arezoomand. Normalized Laplacian polynomial of n-Cayley graphs. Linear Multilinear Algebra, 70(11):2078–2087, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arezoomand:2023:PST

[Are23] Majid Arezoomand. Perfect state transfer on semi-Cayley graphs over abelian groups. Linear Multilinear Algebra, 71 (14):2337–2353, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Abdulkareem:2015:ICI

[ARH15] A. O. Abdulkareem, I. S. Rakhimov, and S. K. Said Husain. Isomorphism classes and invariants of low-dimensional filiform Leibniz algebras. Linear Multilinear Algebra, 63(11): 2254–2274, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ariannejad:2016:GTE

[Ari16] M. Ariannejad. A graph theoretical equivalence for Amitsur– Levitzki theorem. Linear Multilinear Algebra, 64(12):2486– 2490, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Al-Rawabdeh:2022:FIC

[ARK22] Amer Al-Rawabdeh and Fuad Kittaneh. Further inequalities for certain powers of positive definite matrices. Linear Multilinear Algebra, 70(21):7236–7249, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Armario:2017:SNF

[Arm17] José Andrés Armario. On the Smith normal form of skew E– W matrices. Linear Multilinear Algebra, 65(2):375–380, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Achilles:1995:DSM

[AS95] Eva Achilles and Richard Sinkhorn. Doubly stochastic matrices whose squares are idempotent. Linear and Multilinear Algebra, 38(4):343–349, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aghamollaei:2011:PNH

[AS11a] Gholamreza Aghamollaei and Abbas Salemi. Polynomial numerical hulls of matrix polynomials, II. Linear Multilinear Algebra, 59(3):291–302, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Allison:2011:FMM

[AS11b] Mary Allison and Bryan L. Shader. Fastest mixing Markov chain problem for the union of two cliques. Linear Multilinear Algebra, 59(7):801–823, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Al-Saket:2012:DBR

[AS12] Amal Al-Saket. On determining bounds for the real parts of the zeros of polynomials. Linear Multilinear Algebra, 60(3): 377–382, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aydogdu:2015:AC

[AS15] Ismail Aydogdu and Irfan Siap. On Z_{p^{r}}Z_{p^{s}}-additive codes. Linear Multilinear Algebra, 63(10):2089–2102, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alizadeh:2017:LMP

[AS17] R. Alizadeh and F. Shakeri. Linear maps preserving Pareto eigenvalues. Linear Multilinear Algebra, 65(5):1053–1061, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Akbarossadat:2020:DSM

[AS20a] Seyedeh Nafiseh Akbarossadat and Farshid Saeedi. On the dimension of the Schur multiplier of n-Lie algebras. Linear Multilinear Algebra, 68(7):1465–1499, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alakhrass:2020:FIP

[AS20b] M. Alakhrass and M. Sababheh. Functional inequalities for positive matrices. Linear Multilinear Algebra, 68(12):2462– 2472, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Al-Salam:1979:CLP

[ASAS79] Nadhla A. Al-Salam and Waleed A. Al-Salam. The centralizer of the Laguerre polynomial set. Linear and Multilinear Algebra, 7(2):149–154, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ashraf:2016:TCS

[Ash16] Firouzeh Ashraf. On two conjectures on sum of the powers of signless Laplacian eigenvalues of a graph. Linear Multilinear Algebra, 64(7):1314–1320, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Al-Subaihi:2022:FII

[ASR22] Ibrahim Ahmed Al-Subaihi and Mustapha Raïssouli. Further inequalities involving the weighted geometric operator mean and the Heinz operator mean. Linear Multilinear Algebra, 70 (19):4432–4454, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Altinisik:2005:GMP

[AST05] Ercan Altinisik, Bruce E. Sagan, and Naim Tuglu. GCD matrices, posets, and nonintersecting paths. Linear Multilinear Algebra, 53(2):75–84, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Argerami:2007:TT

[AST07] Martin Argerami, Fernando Szechtman, and Ryan Tifenbach. On Tate’s trace. Linear Multilinear Algebra, 55(6):515–520, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Amin:2002:PDG

[ASZ02] Ashok T. Amin, Peter J. Slater, and Guo-Hui Zhang. Parity dimension for graphs — a linear algebraic approach. Linear Multilinear Algebra, 50(4):327–342, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ariannejad:2018:TRG

[AT18] M. Ariannejad and R. Tusserkani. Tree robustness of a graph. Linear Multilinear Algebra, 66(9):1757–1766, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Atik:2020:EPM

[Ati20] Fouzul Atik. On equitable partition of matrices and its applications. Linear Multilinear Algebra, 68(11):2143–2156, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Atkinson:1984:PWS

[Atk84] M. D. Atkinson. A problem of Westwick on k-spaces. Linear and Multilinear Algebra, 16(1–4):263–273, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Atkinson:1991:EKW

[Atk91] M. D. Atkinson. Extensions to the Kronecker–Weierstrass theory of pencils. Linear and Multilinear Algebra, 29(3–4): 235–241, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Audenaert:2015:BMI

[Aud15] Koenraad M. R. Audenaert. On a block matrix inequality quantifying the monogamy of the negativity of entanglement. Linear Multilinear Algebra, 63(12):2526–2536, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Auer:1995:EPI

[Aue95] John W. Auer. An elementary proof of the invertibility of distance matrices. Linear and Multilinear Algebra, 40(2):119– 124, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Auer:1998:IHP

[Aue98] John W. Auer. On the invertibility of Hadamard powers of distance matrices. Linear and Multilinear Algebra, 45(2–3): 235–245, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Arambasic:2022:RRS

[AV22] Ljiljana Arambasić and Anda Valent. On a relation related to strong Birkhoff–James orthogonality. Linear Multilinear Algebra, 70(11):2088–2096, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alfaro:2018:DMO

[AVVÁ18] Carlos A. Alfaro, Carlos E. Valencia, and Adrián Vázquez-Ávila. Digraphs with at most one trivial critical ideal. Linear Multilinear Algebra, 66(10):2036–2048, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Atkinson:1983:SLT

[AW83] M. D. Atkinson and R. Westwick. Spaces of linear transformations of equal rank. Linear and Multilinear Algebra, 13(3): 231–239, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alfakih:2007:SNS

[AW07] A. Y. Alfakih and Henry Wolkowicz. Some necessary and some sufficient trace inequalities for Euclidean distance matrices. Linear Multilinear Algebra, 55(5):499–506, 2007. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03081080701241570.

Appleby:2013:MPV

[AW13] Glenn D. Appleby and Tamsen Whitehead. Matrix pairs over valuation rings and R-valued Littlewood–Richardson fillings.

Asif:2022:LTD

[AWM22] Sania Asif, Zhixiang Wu, and Mobeen Munir. On the Lie triple derivations. Linear Multilinear Algebra, 70(21):6084– 6095, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

An:2015:ECD

[AXH15] Runling An, Chunhui Xue, and Jinchuan Hou. Equivalent characterization of derivations on operator algebras. Linear Multilinear Algebra, 63(1):107–119, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Au-Yeung:1974:SDT

[AY74] Yik-Hoi Au-Yeung. Simultaneous diagonalization of two Hermitian matrices into 2×2 blocks. Linear and Multilinear Algebra, 2(3):249–252, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Au-Yeung:1984:CNR

[AY84] Yik-Hoi Au-Yeung. On the convexity of numerical range in quaternionic Hilbert spaces. Linear and Multilinear Algebra, 16(1–4):93–100, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Alarcon:1993:CHF

[AY93] Ignacio Alarcon and Paul Yiu. Compositions of Hermitian forms. Linear and Multilinear Algebra, 36(2):141–145, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Au-Yeung:1995:SPT

[AY95] Y. H. Au-Yeung. A short proof of a theorem on the numerical range of a normal quaternionic matrix. Linear and Multilinear Algebra, 39(3):279–284, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Au-Yeung:1993:TFG

[AYC93] Yik-Hoi Au-Yeung and Che-Man Cheng. Two formulas for the generalized Radon–Hurwitz number. Linear and Multilinear

Au-Yeung:1983:CMG

[AYT83] Yik-Hoi Au-Yeung and Nam-Kiu Tsing. A conjecture of Marcus on the generalized numerical range. Linear and Multilinear Algebra, 14(3):235–239, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Au-yeung:1984:STG

[AyT84] Yik-Hoi Au-yeung and Nam-Kiu Tsing. Some theorems on the generalized numerical ranges. Linear and Multilinear Algebra, 15(1):3–11, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Aznay:2022:BPN

[AZ22] Zakariae Aznay and Hassan Zariouh. The Berkani’s property and a note on some recent results. Linear Multilinear Algebra, 70(20):5913–5921, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Azenhas:1999:AII

[Aze99] Olga Azenhas. The admissible interval for the invariant factors of a product of matrices. Linear and Multilinear Algebra, 46(1– 2):51–99, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Badea:2019:C

[Bad19] Gabriela Badea. Copies of l_{u}(l_{v}^{n}|n ∈ N) in Π_{2}(l_{p}, l_{q}), 1 ≤ p, q < infty. Linear Multilinear Algebra, 67(3):573–582, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baeza:1984:AIQ

[Bae84] Ricardo Baeza. On the Arf invariant of quadratic forms and of knots. Linear and Multilinear Algebra, 16(1–4):247–252, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bahadorykhalily:2022:MHS

[Bah22] Fereshteh Bahadorykhalily. Multivariate Hasse–Schmidt derivation on exterior algebras. Linear Multilinear Algebra,

70(14):2642–2658, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:1984:NDP

[Bak84] Jerzy K. Baksalary. Nonnegative definite and positive definite solutions to the matrix equation AXA^{∗} = B. Linear and Multilinear Algebra, 16(1–4):133–139, 1984. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bakherad:2016:RRG

[Bak16] Mojtaba Bakherad. Refinements of a reversed AM-GM operator inequality. Linear Multilinear Algebra, 64(9):1687–1695, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballantine:1975:SIS

[Bal75] C. S. Ballantine. Some involutory similarities. Linear and Multilinear Algebra, 3(1–2):19–23, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See addendum [Bal76].

Ballantine:1976:AAS

[Bal76] C. S. Ballantine. Author’s addendum to “Some involutory similarities”. Linear and Multilinear Algebra, 4(1):69, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Bal75].

Ballantine:1977:PIM

[Bal77] C. S. Ballantine. Products of involutory matrices. I. Linear and Multilinear Algebra, 5(1):53–62, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballantine:1978:CCO

[Bal78] C. S. Ballantine. Cosquares: Complex and otherwise. Linear and Multilinear Algebra, 6(3):201–217, 1978. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Balasubramanian:1979:CFR

[Bal79a] K. Balasubramanian. Constancy of functions restricted to a subset. Linear and Multilinear Algebra, 7(3):253–255, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Balasubramanian:1979:MDS

[Bal79b] K. Balasubramanian. Maximal diagonal sums. Linear and Multilinear Algebra, 7(3):249–251, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballantine:1979:PCC

[Bal79c] C. S. Ballantine. Products of complic cosquares and pseudoinvolutory matrices. Linear and Multilinear Algebra, 8(1):73– 78, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballantine:1980:WNR

[Bal80] C. S. Ballantine. Where the numerical range subtends a right angle. Linear and Multilinear Algebra, 9(2):93–97, 1980. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballico:2014:UBR

[Bal14] E. Ballico. An upper bound for the real tensor rank and the real symmetric tensor rank in terms of the complex ranks. Linear Multilinear Algebra, 62(11):1546–1552, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballico:2018:PST

[Bal18] E. Ballico. Partially symmetric tensor rank: the description of the non-uniqueness case for low rank. Linear Multilinear Algebra, 66(3):608–624, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baliarsingh:2022:SFD

[Bal22a] Pinakadhar Baliarsingh. On the spectrum of fractional difference operator. Linear Multilinear Algebra, 70(21):6568–6580, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballico:2022:ISV

[Bal22b] E. Ballico. Identifiability for the k-secant variety of the Segre– Veronese varieties. Linear Multilinear Algebra, 70(21):6441– 6451, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baloudi:2023:FTQ

[Bal23] Hatem Baloudi. Fredholm theory in quaternionic Banach algebras. Linear Multilinear Algebra, 71(6):889–910, 2023. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ben-Ari:2012:PAP

[BAN12] Iddo Ben-Ari and Michael Neumann. Probabilistic approach to Perron root, the group inverse, and applications. Linear Multilinear Algebra, 60(1):39–63, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1987:MSV

[Bap87] R. B. Bapat. Majorization and singular values. Linear and Multilinear Algebra, 21(3):211–214, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1988:RSI

[Bap88] R. B. Bapat. Refining Schur’s inequality using Schur complements. Linear and Multilinear Algebra, 23(1):55–62, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1991:NPH

[Bap91] R. B. Bapat. A note on the permanental Hadamard inequality. Linear and Multilinear Algebra, 30(3):205–208, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1992:RP

[Bap92] R. B. Bapat. Research problem. Linear and Multilinear Algebra, 32(3–4):335–337, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1997:MPI

[Bap97] R. B. Bapat. Moore–Penrose inverse of the incidence matrix of a tree. Linear and Multilinear Algebra, 42(2):159–167, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barker:1974:SMC

[Bar74] George Phillip Barker. Stochastic matrices over cones. Linear and Multilinear Algebra, 1(4):279–287, 1974. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barker:1976:MNT

[Bar76] George Phillip Barker. Monotone norms and tensor products. Linear and Multilinear Algebra, 4(3):191–199, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barker:1978:FDC

[Bar78] George Phillip Barker. Faces and duality in convex cones. Linear and Multilinear Algebra, 6(3):161–169, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barnett:1980:GCD

[Bar80] Stephen Barnett. Greatest common divisors from generalized Sylvester resultant matrices. Linear and Multilinear Algebra, 8(4):271–279, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barker:1988:RP

[Bar88] G. P. Barker. Research problem. Linear and Multilinear Algebra, 23(4):369–371, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bartl:2007:FLO

[Bar07] David Bartl. Farkas’ Lemma, other theorems of the alternative, and linear programming in infinite-dimensional spaces: a purely linear-algebraic approach. Linear Multilinear Algebra, 55(4):327–353, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barik:2008:LSG

[Bar08] Sasmita Barik. On the Laplacian spectra of graphs with pockets. Linear Multilinear Algebra, 56(5):481–490, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bartl:2012:NSA

[Bar12] David Bartl. A note on the short algebraic proof of Farkas’ Lemma. Linear Multilinear Algebra, 60(8):897–901, 2012. CO-

DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bashkirov:2017:MLR

[Bas17] Evgenii L. Bashkirov. On matrix Lie rings over an associative commutative ring with 1 that contain the special orthogonal Lie ring. Linear Multilinear Algebra, 65(9):1878–1893, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bashkirov:2019:CLR

[Bas19] Evgenii L. Bashkirov. On a class of Lie rings of 2 × 2 matrices over associative commutative rings. Linear Multilinear Algebra, 67(3):456–478, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bayat:2022:BPG

[Bay22] M. Bayat. A bijective proof of generalized Cauchy–Binet, Laplace, Sylvester and Dodgson formulas. Linear Multilinear Algebra, 70(16):3220–3230, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bender:1975:SSM

[BB75] Edward A. Bender and J. V. Brawley. Similarity to symmetric matrices over fields which are not formally real. Linear and Multilinear Algebra, 3(1–2):3–8, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boularas:1996:CUM

[BB96] D. Boularas and Z. Bouzar. Concomitants et p-uplets de matrices 2 × 2. Linear and Multilinear Algebra, 41(2):161–173, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:2006:WLC

[BB06] Jerzy K. Baksalary and Oskar Maria Baksalary. When is a linear combination of two idempotent matrices the group involutory matrix? Linear Multilinear Algebra, 54(6):429–435, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballico:2013:TRT

[BB13] E. Ballico and A. Bernardi. Tensor ranks on tangent developable of Segrè varieties. Linear Multilinear Algebra, 61(7): 881–894, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Biagi:2014:CCB

[BB14] S. Biagi and A. Bonfiglioli. On the convergence of the Campbell–Baker–Hausdorff–Dynkin series in infinitedimensional Banach–Lie algebras. Linear Multilinear Algebra, 62(12):1591–1615, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barraa:2015:NRS

[BB15a] Mohamed Barraa and Mohamed Boumazgour. Numerical range submultiplicity. Linear Multilinear Algebra, 63(11): 2311–2317, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berliner:2015:NSI

[BB15b] Adam H. Berliner and Richard A. Brualdi. Nearly symmetricindecomposable matrices. Linear Multilinear Algebra, 63(7): 1468–1484, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballico:2017:URD

[BB17a] Edoardo Ballico and Alessandra Bernardi. A uniqueness result on the decompositions of a bi-homogeneous polynomial. Linear Multilinear Algebra, 65(4):677–698, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benitez:2017:IAE

[BB17b] Julio Benítez and Enrico Boasso. The inverse along an element in rings with an involution, Banach algebras and C^{⋆}-algebras. Linear Multilinear Algebra, 65(2):284–299, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boularas:2017:SIC

[BB17c] D. Boularas and Z. Bouzar. Symplectic invariants and covariants of matrices. Linear Multilinear Algebra, 65(8):1503–1528, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballico:2019:RTS

[BB19a] Edoardo Ballico and Alessandra Bernardi. On the ranks of the third secant variety of Segre–Veronese embeddings. Linear Multilinear Algebra, 67(3):583–597, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Belardo:2019:CSG

[BB19b] Francesco Belardo and Maurizio Brunetti. Connected signed graphs L-cospectral to signed ∞-graphs. Linear Multilinear Algebra, 67(12):2410–2426, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

BenTaher:2019:SLD

[BB19c] R. Ben Taher and H. Benkhaldoun. Solving the linear difference equation with periodic coefficients via Fibonacci sequences. Linear Multilinear Algebra, 67(12):2549–2564, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barraa:2020:ELC

[BB20] Mohamed Barraa and El Hassan Benabdi. On equivalence of linear combinations of idempotents. Linear Multilinear Algebra, 68(5):983–990, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baklouti:2021:SJJ

[BB21] Amir Baklouti and Saïd Benayadi. Symplectic Jacobi–Jordan algebras. Linear Multilinear Algebra, 69(8):1557–1578, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baker:2022:NIV

[BB22] Abdullah Bin Abu Baker and Fernanda Botelho. Numerical index of vector-valued function spaces. Linear Multilinear Algebra, 70(11):2117–2126, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bendaoud:2009:LMP

[BBBS09] M. Bendaoud, A. Bourhim, M. Burgos, and M. Sarih. Linear maps preserving Fredholm and Atkinson elements of C^{∗}- algebras. Linear Multilinear Algebra, 57(8):823–838, 2009.

CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Belardo:2021:CCS

[BBCD21] Francesco Belardo, Maurizio Brunetti, Matteo Cavaleri, and Alfredo Donno. Constructing cospectral signed graphs. Linear Multilinear Algebra, 69(14):2717–2732, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benbouziane:2018:NCP

[BBEL18] H. Benbouziane, Y. Bouramdane, M. Ech-Cherif El Kettani, and A. Lahssaini. Nonlinear commutant preservers. Linear Multilinear Algebra, 66(3):593–601, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Balaji:2022:GED

[BBG22a] R. Balaji, R. B. Bapat, and Shivani Goel. Generalized Euclidean distance matrices. Linear Multilinear Algebra, 70(21): 6908–6929, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Balaji:2022:DMW

[BBG22b] R. Balaji, R. B. Bapat, and Shivani Goel. On distance matrices of wheel graphs with an odd number of vertices. Linear Multilinear Algebra, 70(17):3370–3401, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Balaji:2022:RMB

[BBG22c] R. Balaji, R. B. Bapat, and Shivani Goel. Resistance matrices of balanced directed graphs. Linear Multilinear Algebra, 70(5): 787–808, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benabdi:2022:BJO

[BBK22] El Hassan Benabdi, Abderrahim Baghdad, and Mohamed Chraibi Kaadoud. Birkhoff–James orthogonality and algebraic maximal numerical range in C^{∗}-algebras. Linear Multilinear Algebra, 70(21):6523–6534, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benito:2015:SMO

[BBM15] Pilar Benito, Murray Bremner, and Sara Madariaga. Symmetric matrices, orthogonal Lie algebras and Lie–Yamaguti algebras. Linear Multilinear Algebra, 63(6):1257–1281, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Biagi:2020:BCH

[BBM20] Stefano Biagi, Andrea Bonfiglioli, and Marco Matone. On the Baker–Campbell–Hausdorff theorem: non-convergence and prolongation issues. Linear Multilinear Algebra, 68(7):1310– 1328, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barker:1978:PDS

[BBP78] G. P. Barker, A. Berman, and R. J. Plemmons. Positive diagonal solutions to the Lyapunov equations. Linear and Multilinear Algebra, 5(4):249–256, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhunia:2021:NRI

[BBP21] Pintu Bhunia, Santanu Bag, and Kallol Paul. Numerical radius inequalities of operator matrices with applications. Linear Multilinear Algebra, 69(9):1635–1644, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barile:2007:PDC

[BBT07] Margherita Barile, Fiorella Barone, and Wlodzimierz M. Tulczyjew. Polarizations and differential calculus in affine spaces. Linear Multilinear Algebra, 55(2):121–146, 2007. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benitez:2020:CDD

[BBX20] Julio Benítez, Enrico Boasso, and Sanzhang Xu. On the continuity and differentiability of the (dual) core inverse in C∗-algebras. Linear Multilinear Algebra, 68(4):686–709, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bozkurt:2015:SRE

[BBZ15] S. Burcu Bozkurt, Durmus Bozkurt, and Xiao-Dong Zhang. On the spectral radius and the energy of a digraph. Linear Multilinear Algebra, 63(10):2009–2016, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1978:PS

[BC78] Leroy B. Beasley and Larry J. Cummings. Permanent semigroups. Linear and Multilinear Algebra, 5(4):297–302, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barker:1980:GTC

[BC80] G. P. Barker and Dave Carlson. Generalizations of topheavy cones. Linear and Multilinear Algebra, 8(3):219–230, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1982:MPG

[BC82] Leory B. Beasley and Larry J. Cummings. Multiplicative properties of generalized matrix functions. Linear and Multilinear Algebra, 11(1):23–31, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barker:1983:IST

[BC83] George Phillip Barker and Stephen L. Campbell. Internal stability of two dimensional systems. Linear and Multilinear Algebra, 14(4):365–369, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1985:TSG

[BC85] L. B. Beasley and L. J. Cummings. Test sets for generalized matrix functions. Linear and Multilinear Algebra, 17(3–4): 301–307, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:1991:EDR

[BC91] Ricard A. Brualdi and Keith L. Chavey. Elementary divisors and ranked posets with application to matrix compounds. Linear and Multilinear Algebra, 29(2):81–90, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bru:1994:NJC

[BC94] Rafael Bru and Rafael Cantó. On nonnegative Jordan chains for the generalized nullspace of a singular M-matrix. Linear and Multilinear Algebra, 36(4):267–274, 1994. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bakonyi:2005:RPC

[BC05] M. Bakonyi and T. Constantinescu. Research problem: The completion number of a graph. Linear Multilinear Algebra, 53(3):189–192, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boularas:2006:EAI

[BC06] D. Boularas and A. Chouikrat. Équations d’amorçage d’intégrales premières formelles. (French) [Bootstrap equations of formal first integrals]. Linear Multilinear Algebra, 54(3):219–233, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Breaz:2017:SNM

[BC17] Simion Breaz and Grigore Calugareanu. Sums of nilpotent matrices. Linear Multilinear Algebra, 65(1):67–78, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boumazgour:2020:ESD

[BC20] Mohamed Boumazgour and Ezzeddine Chafai. On the essential spectrum of the diagonal of an operator. Linear Multilinear Algebra, 68(7):1302–1309, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boussairi:2022:GTM

[BCCL22] A. Boussaïri, A. Chaïchaâ, B. Chergui, and S. Lakhlifi. Generalized tournament matrices with the same principal minors. Linear Multilinear Algebra, 70(20):5433–5444, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barriere:2009:HPG

[BCDF09] L. Barrière, F. Comellas, C. Dalfó, and M. A. Fiol. On the hierarchical product of graphs and the generalized binomial tree. Linear Multilinear Algebra, 57(7):695–712, 2009. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beitia:2010:CFI

[BCdHP10] M. Asunción Beitia, Albert Compta, Inmaculada de Hoyos, and Marta Peña. The change of feedback invariants under column perturbations: particular cases. Linear Multilinear Algebra, 58(1):45–59, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bendito:2012:DRG

[BCEM12] E. Bendito, A. Carmona, A. M. Encinas, and M. Mitjana. Distance-regular graphs having the M-property. Linear Multilinear Algebra, 60(2):225–240, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benitez:2013:EIA

[BCI13] Julio Benítez and Dragana Cvetković-Ilić. Equalities of ideals associated with two projections in rings with involution. Linear Multilinear Algebra, 61(10):1419–1435, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhat:2015:SEI

[BCK15] B. V. Rajarama Bhat, Arup Chattopadhyay, and G. Sankara Raju Kosuru. On submajorization and eigenvalue inequalities. Linear Multilinear Algebra, 63(11):2245–2253, 2015. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bru:2010:SSD

[BCKP10] Rafael Bru, Ljiljana Cvetković, Vladimir Kostić, and Francisco Pedroche. Sums of Σ-strictly diagonally dominant matrices. Linear Multilinear Algebra, 58(1):75–78, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barreiro:2023:LAG

[BCLS23] Elisabete Barreiro, Antonio J. Calderón, Samuel A. Lopes, and José M. Sánchez. Leibniz algebras and graphs. Linear Multilinear Algebra, 71(12):1994–2007, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bouaniza:2018:PSB

[BCM18] Hafsa Bouaniza, Yosra Chamkha, and Maher Mnif. Perturbation of semi-Browder linear relations by commuting Riesz operators. Linear Multilinear Algebra, 66(2):285–308, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2005:FPM

[BCN05] R. B. Bapat, M. Catral, and Michael Neumann. On functions that preserve M-matrices and inverse M-matrices. Linear Multilinear Algebra, 53(3):193–201, 2005. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baradaran:2023:SPF

[BCNG23] Javad Baradaran, Jahangir Cheshmavar, Fereshte Nikahd, and Zahra Ghorbani. Some properties of c–K–g-frames in Hilbert C^{∗}-modules. Linear Multilinear Algebra, 71(8):1323–1337, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ball:1991:ZDI

[BCR91] Joseph A. Ball, Nir Cohen, and Leiba Rodman. Zero data and interpolation problems for rectangular matrix polynomials. Linear and Multilinear Algebra, 29(1):53–78, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bermudez:2014:ICV

[BCS14] J. M. Ancochea Bermúdez and R. Campoamor-Stursberg. An irreducible component of the variety of Leibniz algebras having trivial intersection with the variety of Lie algebras. Linear Multilinear Algebra, 62(11):1450–1459, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bermudez:2018:RPC

[BCS18] J. M. Ancochea Bermúdez and R. Campoamor-Stursberg. Rigidity-preserving and cohomology-decreasing extensions of solvable rigid Lie algebras. Linear Multilinear Algebra, 66(3): 525–539, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Borobia:2023:ESS

[BCT23] Alberto Borobia, Roberto Canogar, and Fernando De Terán. The equation X^{T} AX = B with B skew-symmetric: how much of a bilinear form is skew-symmetric? Linear Multilinear Algebra, 71(13):2114–2143, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2022:GKD

[BCZ22] Yuxiu Bai, Yuqun Chen, and Zerui Zhang. Gelfand-Kirillov dimension of bicommutative algebras. Linear Multilinear Algebra, 70(22):7623–7649, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhatia:1984:BSV

[BD84] Rajendra Bhatia and Chandler Davis. A bound for the spectral variation of a unitary operator. Linear and Multilinear Algebra, 15(1):71–76, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhatia:1985:CCF

[BD85] Rajendra Bhatia and Chandler Davis. Concavity of certain functions of matrices. Linear and Multilinear Algebra, 17(2): 155–164, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1988:BNR

[BD88] Natália Bebiano and Joáo Da Providencia. On the boundary of the C-numerical range of a normal matrix. Linear and Multilinear Algebra, 23(2):145–157, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1994:SGP

[BD94] Natália Bebiano and Joao Da Providência. Some geometrical properties of the c-numerical range of a normal matrix. Linear and Multilinear Algebra, 37(1–3):83–92, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1996:APC

[BD96] Natália Bebiano and João Da Providência. Another proof of a conjecture of Marcus on the c-numerical range. Linear and Multilinear Algebra, 41(1):35–40, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1997:ODS

[BD97] Leroy B. Beasley and J. A. Dias Da Silva. Orthogonal decomposable symmetrized tensors. Linear and Multilinear Algebra, 42(1):1–21, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2012:MLI

[BD12] Zhaofang Bai and Shuanping Du. Multiplicative ∗-Lie isomorphisms between von Neumann algebras. Linear Multilinear Algebra, 60(3):311–322, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bresar:2013:CPM

[BD13a] Matej Bresar and Vesselin Drensky. Central polynomials for matrices over finite fields. Linear Multilinear Algebra, 61(7): 939–944, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:2013:EPD

[BD13b] Richard A. Brualdi and Geir Dahl. An extension of the polytope of doubly stochastic matrices. Linear Multilinear Algebra, 61(3):393–408, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bueno:2014:EMB

[BD14] María I. Bueno and Fernando De Terán. Eigenvectors and minimal bases for some families of Fiedler-like linearizations. Linear Multilinear Algebra, 62(1):39–62, 2014. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

BaydarYarbil:2022:CPG

[BD22a] Nihan Baydar Yarbil and Vincenzo De Filippis. Commuting product of generalized derivations on multilinear polynomials in prime rings. Linear Multilinear Algebra, 70(9):1697–1716, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:2022:DSD

[BD22b] Richard A. Brualdi and Geir Dahl. Diagonal sums of doubly stochastic matrices. Linear Multilinear Algebra, 70(20):4946– 4972, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bozzo:2023:JFP

[BDD23] Enrico Bozzo, Piero Deidda, and Carmine Di Fiore. The Jordan and Frobenius pairs of the inverse. Linear Multilinear Algebra, 71(10):1730–1735, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bien:2023:CDI

[BDH23] Mai Hoang Bien, Truong Huu Dung, and Nguyen Thi Thai Ha. A certain decomposition of infinite invertible matrices over division algebras. Linear Multilinear Algebra, 71(12):1948– 1956, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bernik:2005:CFP

[BDK^{+}05] J. Bernik, R. Drnovšek, T. Košir, T. Laffey, G. MacDonald, R. Meshulam, M. Omladič, and H. Radjavi. Common fixed points and common eigenvectors for sets of matrices. Linear Multilinear Algebra, 53(2):137–146, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bani-Domi:2009:NRI

[BDK09] Wathiq Bani-Domi and Fuad Kittaneh. Numerical radius inequalities for operator matrices. Linear Multilinear Algebra, 57(4):421–427, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bani-Domi:2021:NNR

[BDK21] Watheq Bani-Domi and Fuad Kittaneh. Norm and numerical radius inequalities for Hilbert space operators. Linear Multilinear Algebra, 69(5):934–945, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1998:NRP

[BdP98] Natália Bebiano and João da Providência. Numerical ranges in physics. Linear and Multilinear Algebra, 43(4):327–337, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:2012:IHR

[BdP12] N. Bebiano and J. da Providência. Indefinite higher-rank numerical ranges. Linear Multilinear Algebra, 60(9):1009–1026, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:2021:MEP

[BdP21] N. Bebiano and J. da Providência. Maximum entropy principle and Landau free energy inequality. Linear Multilinear Algebra, 69(6):1020–1034, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:2014:NRB

[BdPN14] N. Bebiano, J. da Providência, and A. Nata. The numerical range of banded 2-Toeplitz operators in spaces with an indefinite metric. Linear Multilinear Algebra, 62(5):659–673, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:2017:GRQ

[BdPNdP17] N. Bebiano, J. da Providência, A. Nata, and J. P. da Providência. Generalized Rayleigh quotients and generating vectors. Linear Multilinear Algebra, 65(1):1–23, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bendaoud:2013:MMP

[BDS13] M. Bendaoud, M. Douimi, and M. Sarih. Maps on matrices preserving local spectra. Linear Multilinear Algebra, 61(7): 871–880, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Besleaga:2019:CGG

[BDV19] F. Besleaga, S. Dascalescu, and L. Van Wyk. Classifying good gradings on structural matrix algebras. Linear Multilinear Algebra, 67(10):1948–1957, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boboc:2012:IBM

[BDvW12] C. Boboc, S. Dascalescu, and L. van Wyk. Isomorphisms between Morita context rings. Linear Multilinear Algebra, 60(5): 545–563, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boboc:2016:JIT

[BDvW16] C. Boboc, S. Dascalescu, and L. van Wyk. Jordan isomorphisms of 2-torsionfree triangular rings. Linear Multilinear Algebra, 64(2):290–296, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhatia:1990:VP

[BE90] Rajendra Bhatia and Ludwig Elsner. On the variation of permanents. Linear and Multilinear Algebra, 27(2):105–110, 1990.

Bremner:2020:SIC

[BE20] Murray R. Bremner and Hader A. Elgendy. Special identities for comtrans algebras. Linear Multilinear Algebra, 68(6):1140– 1159, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bouramdane:2022:MPS

[BE22] Y. Bouramdane and M. Ech-Chérif El Kettani. Maps preserving some spectral domains of skew products of operators. Linear Multilinear Algebra, 70(21):6978–6988, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bremner:2023:PTS

[BE23] Murray R. Bremner and Hader A. Elgendy. Poisson triple systems. Linear Multilinear Algebra, 71(7):1145–1157, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1978:LTM

[Bea78] Leroy B. Beasley. Linear transformations on matrices: The invariance of commuting pairs of matrices. Linear and Multilinear Algebra, 6(3):179–183, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1983:GMFa

[Bea83a] Leroy B. Beasley. Generalized matrix functions of class MPW. Linear and Multilinear Algebra, 12(4):273–280, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1983:GMFb

[Bea83b] Leroy B. Beasley. Generalized matrix functions of class MPW, II. Linear and Multilinear Algebra, 13(2):87–95, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1984:GMF

[Bea84] Leroy B. Beasley. Generalized matrix functions of class MPW, III. Linear and Multilinear Algebra, 15(2):175–186, 1984. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1985:NSS

[Bea85] Leroy B. Beasley. The null space of spaces of singular matrices. Linear and Multilinear Algebra, 18(3):249–254, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beard:1986:SX

[Bea86] Jacob T. B. Beard Jr. Similarity over F [x]. Linear and Multilinear Algebra, 19(3):259–265, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1986:NP

[Beb86] Natália Bebiano. Nondifferentiabie points of ∂W_{c}(A). Linear and Multilinear Algebra, 19(3):249–257, 1986. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bedratyuk:2010:PSA

[Bed10] Leonid Bedratyuk. The Poincaré series of the algebras of simultaneous invariants and covariants of two binary forms. Linear Multilinear Algebra, 58(6):789–803, 2010. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bedratyuk:2011:ACS

[Bed11] Leonid Bedratyuk. Analogue of the Cayley–Sylvester formula and the Poincaré series for the algebra of invariants of nary form. Linear Multilinear Algebra, 59(11):1189–1199, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benbouziane:2019:NMP

[BEH19] Hassane Benbouziane, Mustapha Ech-Cherif El Kettani, and Imane Herrou. Nonlinear maps preserving the local spectral subspace. Linear Multilinear Algebra, 67(1):29–38, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bellman:1983:CME

[Bel83] Richard Bellman. On the calculation of matrix exponential. Linear and Multilinear Algebra, 13(1):73–79, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Belardo:2015:LES

[Bel15] Francesco Belardo. On the largest eigenvalue of some bidegreed graphs. Linear Multilinear Algebra, 63(1):166–184, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bender:1974:CPSa

[Ben74] Edward A. Bender. Characteristic polynomials of symmetric matrices II: Local number fields. Linear and Multilinear Algebra, 2(1):55–63, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benson:1978:CLT

[Ben78] Diane Benson. A characterization of linear transformations which leave the doubly stochastic matrices invariant. Linear and Multilinear Algebra, 6(1):65–72, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benasseni:1990:NBV

[Ben90] J. Benasseni. A note on bounds to the variation of eigenvalues in symmetric matrix perturbation of rank one. Linear and Multilinear Algebra, 27(2):111–116, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benkovic:2005:JHT

[Ben05] Dominik Benkovic. Jordan homomorphisms on triangular matrices. Linear Multilinear Algebra, 53(5):345–356, 2005. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benkovic:2007:LDT

[Ben07] Dominik Benkovič. Lie derivations on triangular matrices. Linear Multilinear Algebra, 55(6):619–626, 2007. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benasseni:2010:ARB

[Bén10] Jacques Bénasséni. On the algebraic relation between principal components corresponding to two different sets of weights. Linear Multilinear Algebra, 58(4):425–429, 2010. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benasseni:2011:LBL

[Bén11] Jacques Bénasséni. Lower bounds for the largest eigenvalue of a symmetric matrix under perturbations of rank one. Linear Multilinear Algebra, 59(5):565–569, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benkovic:2015:LTD

[Ben15] Dominik Benkovic. Lie triple derivations of unital algebras with idempotents. Linear Multilinear Algebra, 63(1):141–165, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benkovic:2016:JDT

[Ben16] Dominik Benkovic. Jordan σ-derivations of triangular algebras. Linear Multilinear Algebra, 64(2):143–155, 2016. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bendaoud:2018:PLS

[Ben18] Mohamed Bendaoud. Preservers of local spectrum of matrix skew products. Linear Multilinear Algebra, 66(8):1530–1537, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benkovic:2022:LDT

[Ben22] Dominik Benkovic. Lie σ-derivations of triangular algebras. Linear Multilinear Algebra, 70(15):2966–2983, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berele:1986:CSM

[Ber86] Allan Berele. Construction of Sp-modules by tableaux. Linear and Multilinear Algebra, 19(4):299–307, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bercovici:2002:NRO

[Ber02] Hari Bercovici. Numerical ranges of operators of class C_{0}. Linear Multilinear Algebra, 50(3):219–222, 2002. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bernik:2005:GSM

[Ber05] Janez Bernik. On groups and semigroups of matrices with spectra in a finitely generated field. Linear Multilinear Alge-

bra, 53(4):259–267, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berliner:2012:CCM

[Ber12] Adam H. Berliner. Convertible and m-convertible matrices. Linear Multilinear Algebra, 60(3):267–283, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berkics:2017:PSM

[Ber17] Peter Berkics. On parallel sum of matrices. Linear Multilinear Algebra, 65(10):2114–2123, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdl:1975:MCP

[BF75] Richard A. Brualdl and Thomas H. Foregger. Matrices with constant permanental minors. Linear and Multilinear Algebra, 3(3):227–243, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Busque:1996:ACW

[BF96] Claudi Busqué and Alberto Facchini. Abelian categories in which linear algebra is possible. Linear and Multilinear Algebra, 41(2):137–156, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barioli:2005:EGD

[BF05] Francesco Barioli and Shaun Fallat. On the eigenvalues of generalized and double generalized stars. Linear Multilinear Algebra, 53(4):269–291, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:2015:TIM

[BF15] Richard A. Brualdi and Eliseu Fritscher. Trace index of a (0, 1)-matrix. Linear Multilinear Algebra, 63(6):1073–1085, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bradler:2022:NHC

[BFI22] Kamil Brádler, Shmuel Friedland, and Robert B. Israel. Nonnegativity for hafnians of certain matrices. Linear Multilinear Algebra, 70(19):4615–4619, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brusamarello:2012:AAI

[BFJ12] Rosali Brusamarello, Érica Z. Fornaroli, and Ednei A. Santulo Jr. Anti-automorphisms and involutions on (finitary) incidence algebras. Linear Multilinear Algebra, 60(2):181–188, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brusamarello:2018:JIF

[BFK18] Rosali Brusamarello, Érica Z. Fornaroli, and Mykola Khrypchenko. Jordan isomorphisms of finitary incidence algebras. Linear Multilinear Algebra, 66(3):565–579, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baklouti:2020:JNO

[BFM20] Hamadi Baklouti, Kais Feki, and Sid Ahmed Ould Ahmed Mahmoud. Joint normality of operators in semi-Hilbertian spaces. Linear Multilinear Algebra, 68(4):845–866, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See correction [Ano20b].

Boralevi:2022:QSP

[BFM22] Ada Boralevi, Maria Lucia Fania, and Emilia Mezzetti. Quadric surfaces in the Pfaffian hypersurface in P^{14}. Linear Multilinear Algebra, 70(19):4675–4694, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barrodale:1973:BR

[BFMM73] Ian Barrodale, John Fogarty, Russell Merris, and B. N. Moyls. Book review. Linear and Multilinear Algebra, 1(2):175–181, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baragana:2007:HIS

[BFZ07] I. Baragaña, V. Fernández, and I. Zaballa. Hermite indices and state feedback: generic case. Linear Multilinear Algebra, 55(2): 113–120, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bogart:1979:HCP

[BG79] Kenneth Bogart and Jean Gordon. Hypergraphs, characteristic polynomials and permutation characters. Linear and Mul-

tilinear Algebra, 7(3):213–236, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bunse-Gerstner:1982:STT

[BG82] Angelika Bunse-Gerstner. On the similarity transformation to tridiagonal form. Linear and Multilinear Algebra, 12(1):51–56, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barnett:1983:SEH

[BG83] S. Barnett and M. J. C. Gover. Some extensions of Hankel and Toeplitz matrices. Linear and Multilinear Algebra, 14(1): 45–65, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ball:1986:PSI

[BG86] Joseph A. Ball and Israel Gohberg. Pairs of shift invariant subspaces of matrices and noncanonical factorization. Linear and Multilinear Algebra, 20(1):27–61, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bergeron:1990:SFL

[BG90] N. Bergeron and A. M. Garsia. Sergeev’s formula and the Littlewood–Richardson rule. Linear and Multilinear Algebra, 27(2):79–100, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Blanchard:2016:UBA

[BG16] Etienne Blanchard and Ilja Gogić. On unital C(X)-algebras and C(X)-valued conditional expectations of finite index. Linear Multilinear Algebra, 64(12):2406–2418, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bialas:2021:GHP

[BG21] Stanisław Białas and Michał Góra. The generalized Hadamard product of polynomials and its stability. Linear Multilinear Algebra, 69(7):1338–1351, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barbarino:2023:ETG

[BG23a] Giovanni Barbarino and Carlo Garoni. An extension of the theory of GLT sequences: sampling on asymptotically uniform grids. Linear Multilinear Algebra, 71(12):2008–2025, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bresar:2023:WJH

[BG23b] M. Bresar and M. L. C. Godoy. Weighted Jordan homomorphisms. Linear Multilinear Algebra, 71(8):1265–1279, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beitia:1992:LME

[BGdH92] M. Asunción Beitia, Juan M. Gracia, and Inmaculada de Hoyos. A linear matrix equation: a criterion for block similarity. Linear and Multilinear Algebra, 31(1–4):93–118, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bru:2017:DEC

[BGGS17] Rafael Bru, Maria T. Gassó, Isabel Giménez, and Máximo Santana. Diagonal entries of the combined matrix of a totally negative matrix. Linear Multilinear Algebra, 65(10): 1971–1984, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1999:GMT

[BGK99] Ravindra B. Bapat, Jerrold W. Grossman, and Devadatta M. Kulkarni. Generalized matrix tree theorem for mixed graphs. Linear and Multilinear Algebra, 46(4):299–312, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2000:EVM

[BGK00] Ravindra B. Bapat, Jerrold W. Grossman, and Devadatta M. Kulkarni. Edge version of the matrix tree theorem for trees. Linear and Multilinear Algebra, 47(3):217–229, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:2007:FDT

[BGLS07] Leroy B. Beasley, Alexander E. Guterman, Sang-Gu Lee, and Seok-Zun Song. Frobenius and Dieudonné theorems over semirings. Linear Multilinear Algebra, 55(1):19–34, 2007. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bernal-Gonzalez:2020:HSN

[BGLSSS20] L. Bernal-González, J. López-Salazar, and J. B. Seoane-Sepúlveda. Holomorphic surjectivity near the boundary: topological and algebraic genericity. Linear Multilinear Algebra, 68(7):1434–1456, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2018:LB

[BGLZ18] Ruipu Bai, Weiwei Guo, Lixin Lin, and Yan Zhang. n-Lie bialgebras. Linear Multilinear Algebra, 66(2):382–397, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barik:2022:GSA

[BGM22] Sasmita Barik, Sattick Ghosh, and Debabrota Mondal. On graphs with strong anti-reciprocal eigenvalue property. Linear Multilinear Algebra, 70(21):6698–6711, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berenhaut:2011:OBI

[BGV11] Kenneth S. Berenhaut, Richard T. Guy, and Nathaniel G. Vish. An optimal bound for inverses of triangular matrices with monotone entries. Linear Multilinear Algebra, 59(4):475– 481, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bu:2016:MWM

[BGZW16] Changjiang Bu, Weiqi Gu, Jiang Zhou, and Yimin Wei. On matrices whose Moore–Penrose inverses are ray unique. Linear Multilinear Algebra, 64(6):1236–1243, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bender:1974:CPSb

[BH74] Edward A. Bender and Norman P. Herzberg. Characteristic polynomials of symmetric matrices III: some counterexamples. Linear and Multilinear Algebra, 2(2):173–178, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barria:1990:VBT

[BH90] José Barría and P. R. Halmos. Vector bases for two commuting matrices. Linear and Multilinear Algebra, 27(3):147–157,

Bohata:2013:NMN

[BH13] Martin Bohata and Jan Hamhalter. Nonlinear maps on von Neumann algebras preserving the star order. Linear Multilinear Algebra, 61(7):998–1009, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bier:2015:NCG

[BH15a] Agnieszka Bier and Waldemar Holubowski. A note on commutators in the group of infinite triangular matrices over a ring. Linear Multilinear Algebra, 63(11):2301–2310, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brice:2015:ZPD

[BH15b] Daniel Brice and Huajun Huang. On zero product determined algebras. Linear Multilinear Algebra, 63(2):326–342, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bohata:2016:SOO

[BH16] Martin Bohata and Jan Hamhalter. Star order on operator and function algebras and its nonlinear preservers. Linear Multilinear Algebra, 64(12):2519–2532, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bulut:2020:EAC

[BH20] Alper Bulut and Ilhan Hacioglu. The energy of all connected cubic circulant graphs. Linear Multilinear Algebra, 68(4):679– 685, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Balani:2023:MBP

[BH23] Fariba Bakrani Balani and Masoud Hajarian. Modified block product preconditioner for a class of complex symmetric linear systems. Linear Multilinear Algebra, 71(9):1521–1535, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhatia:1991:LE

[Bha91] Rajendra Bhatia. Letter to the editor. Linear and Multilinear Algebra, 30(1–2):155, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhate:1993:CH

[Bha93] Hemant Bhate. On the construction of hypergroups. Linear and Multilinear Algebra, 34(3–4):395–402, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2016:EWB

[BHC16] Qingmei Bai, Junjie Huang, and Alatancang Chen. Essential, Weyl and Browder spectra of unbounded upper triangular operator matrices. Linear Multilinear Algebra, 64(8):1583–1594, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2010:AMP

[BHD10] Zhaofang Bai, Jinchuan Hou, and Shuanping Du. Additive maps preserving rank-one operators on nest algebras. Linear Multilinear Algebra, 58(3):269–283, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2011:DPM

[BHD11] Zhaofang Bai, Jinchuan Hou, and Shuanping Du. Distance preserving maps of nest algebras. Linear Multilinear Algebra, 59(5):571–594, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:1986:AF

[BHH86] Richard A. Brualdi, D. J. Hartfiel, and Suk-Geun Hwang. On assignment functions. Linear and Multilinear Algebra, 19(3): 203–219, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Booth:2011:MSR

[BHH^{+}11] Matthew Booth, Philip Hackney, Benjamin Harris, Charles R. Johnson, Margaret Lay, Terry D. Lenker, Lon H. Mitchell, Sivaram K. Narayan, Amanda Pascoe, and Brian D. Sutton. On the minimum semidefinite rank of a simple graph. Linear Multilinear Algebra, 59(5):483–506, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Batigne:1978:FRI

[BHK78] Daniel R. Batigne, Frank J. Hall, and Irving J. Katz. Further results on integral generalized inverses of integral matrices. Linear and Multilinear Algebra, 6(3):233–241, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bourin:2010:JMI

[BHK10] Jean-Christophe Bourin, Omar Hirzallah, and Fuad Kittaneh. Jensen matrix inequalities and direct sums. Linear Multilinear Algebra, 58(5):645–652, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benmakhlouf:2021:NRH

[BHK21] Ahlem Benmakhlouf, Omar Hirzallah, and Fuad Kittaneh. On the p-numerical radii of Hilbert space operators. Linear Multilinear Algebra, 69(15):2813–2829, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Borosh:1976:AQP

[BHM76] I. Borosh, D. J. Hartfiel, and C. J. Maxon. Answers to questions posed by Richman and Schneider. Linear and Multilinear Algebra, 3(4):255–258, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bakhet:2022:MVE

[BHY22] Ahmed Bakhet, Fuli He, and Mimi Yu. On the matrix version of extended Bessel functions and its application to matrix differential equations. Linear Multilinear Algebra, 70(20): 5661–5680, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1995:SVM

[BI95] R. B. Bapat and Adi-Ben Israel. Singular values and maximum rank minors of generalized inverses. Linear and Multilinear Algebra, 40(2):153–161, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballantine:1975:AMP

[BJ75] C. S. Ballantine and C. R. Johnson. Accretive matrix products. Linear and Multilinear Algebra, 3(3):169–185, 1975. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barrett:1993:MMS

[BJ93] Wayne W. Barrett and Charles R. Johnson. Majorization monotonicity of symmetrized Fischer products. Linear and Multilinear Algebra, 34(1):67–74, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bakonyi:1994:DFN

[BJ94] Mihály Bakonyi and Charles R. Johnson. Determinantal formulae and nonsymmetric Gaussian perfect elimination. Linear and Multilinear Algebra, 38(1–2):1–11, 1994. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bakonyi:1996:ACC

[BJ96] Mihály Bakonyi and Charles R. Johnson. Algebraic characterizations of chordality. Linear and Multilinear Algebra, 40(3): 187–191, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bakonyi:1998:SPM

[BJS98] Mihaly Bakonyi, Charles R. Johnson, and David P. Stanford. Sign pattern matrices that require domain-range pairs with given sign patterns. Linear and Multilinear Algebra, 44(2): 165–178, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bendaoud:2015:PLS

[BJS15] M. Bendaoud, M. Jabbar, and M. Sarih. Preservers of local spectra of operator products. Linear Multilinear Algebra, 63(4):806–819, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botelho:2013:COM

[BJZ13] Fernanda Botelho, James Jamison, and Bentuo Zheng. Circular operators on minimal norm ideals of B(H). Linear Multilinear Algebra, 61(10):1339–1347, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:1983:RIC

[BK83] Jerzy K. Baksalary and Radoslaw Kala. Range invariance of certain matrix products. Linear and Multilinear Algebra, 14 (1):89–96, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brujald:1987:EPC

[BK87] Richard A. Brujald and Moshe Katz. An extremal problem concerning matrices of 0’s and 1’s. Linear and Multilinear Algebra, 20(4):325–331, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2017:SNF

[BK17] Ravindra B. Bapat and Masoud Karimi. Smith normal form of some distance matrices. Linear Multilinear Algebra, 65(6): 1117–1130, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baghdad:2021:NCI

[BK21] Abderrahim Baghdad and Mohamed Chraibi Kaadoud. A new class intermediate between hyponormal operators and normaloid operators. Linear Multilinear Algebra, 69(15):2926– 2932, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhatia:1997:SIC

[BKL97] Rajendra Bhatia, Fuad Kittaneh, and Ren Cang Li. Some inequalities for commutators and an application to spectral variation. II. Linear and Multilinear Algebra, 43(1–3):207–219, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhatia:2002:PNS

[BKL02] Rajendra Bhatia, William Kahan, and Ren-Cang Li. Pinchings and norms of scaled triangular matrices. Linear Multilinear Algebra, 50(1):15–21, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bouthat:2022:NCM

[BKMMG22] Ludovick Bouthat, Apoorva Khare, Javad Mashreghi, and Frédéric Morneau-Guérin. The p-norm of circulant matrices. Linear Multilinear Algebra, 70(21):7176–7188, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2001:PLM

[BKP01] R. B. Bapat, S. J. Kirkland, and S. Pati. The perturbed Laplacian matrix of a graph. Linear and Multilinear Algebra, 49(3):

Berman:2012:CCD

[BKS12] Abraham Berman, Christopher King, and Robert Shorten. A characterisation of common diagonal stability over cones. Linear Multilinear Algebra, 60(10):1117–1123, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bedrani:2021:NRA

[BKS21] Yassine Bedrani, Fuad Kittaneh, and Mohammed Sababheh. Numerical radii of accretive matrices. Linear Multilinear Algebra, 69(5):957–970, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barnett:1980:PCI

[BL80] S. Barnett and P. Lancaster. Problems of control and information theory (Hungary). Linear and Multilinear Algebra, 9 (2):99–110, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballantine:1983:SDR

[BL83] C. S. Ballantine and R. Loewy. Self-duality results for unitary and Hermitian matrices. Linear and Multilinear Algebra, 13(3):221–230, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1992:BSD

[BL92] Natália Bebiano and Chi-Kwong Li. A brief survey on the decomposable numerical range of matrices. Linear and Multilinear Algebra, 32(3–4):179–190, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1993:LOS

[BL93a] Leroy B. Beasley and Sang-Gu Lee. Linear operators strongly preserving r-cyclic matrices over semirings. Linear and Multilinear Algebra, 35(3–4):325–337, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bourque:1993:MAA

[BL93b] Keith Bourque and Steve Ligh. Matrices associated with arithmetical functions. Linear and Multilinear Algebra, 34(3–4):

Beasley:1997:RPS

[BL97] Leroy B. Beasley and Raphael Loewy. Rank preservers on spaces of symmetric matrices. Linear and Multilinear Algebra, 43(1–3):63–86, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:2001:LPM

[BL01] Leroy B. Beasley and Thomas J. Laffey. Linear preservers of matrices of rank-2. Linear and Multilinear Algebra, 48(4): 319–331, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2010:CNN

[BL10] Ruipu Bai and Xuewen Liu. A class of non-nilpotent 2-solvable n-Lie algebras. Linear Multilinear Algebra, 58(5):653–658, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brusamarello:2011:AII

[BL11] Rosali Brusamarello and David W. Lewis. Automorphisms and involutions on incidence algebras. Linear Multilinear Algebra, 59(11):1247–1267, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benitez:2013:EGI

[BL13] Julio Benítez and Xiaoji Liu. Expressions for generalized inverses of square matrices. Linear Multilinear Algebra, 61(11): 1536–1554, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bistritz:2021:FFC

[BL21] Yuval Bistritz and Alexander Lifshitz. Fraction-free computation of the unit-circle resultant with any singularity profile. Linear Multilinear Algebra, 69(3):377–393, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Blair:1977:GDT

[Bla77] William D. Blair. Goldman determinants and a theorem of Frobenius. Linear and Multilinear Algebra, 5(3):181–182,

Bebiano:1991:NRD

[BLD91] Natália Bebiano, Chi-Kwong Li, and João Da Providência. The numerical range and decomposable numerical range of matrices. Linear and Multilinear Algebra, 29(3–4):195–205, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:2004:GNR

[BLdPS04] N. Bebiano, R. Lemos, J. da Providência, and G. Soares. On generalized numerical ranges of operators on an indefinite inner product space. Linear Multilinear Algebra, 52(3–4):203– 233, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bloom:1995:BCO

[Blo95] Laura A. Bloom. Bijective comparison of optimal planarity algorithms. Linear and Multilinear Algebra, 38(3):249–266, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1992:CMP

[BLP92] Leroy Beasley, Chi-Kwong Li, and Stephen Pierce. Chapter 9: miscellaneous preserver problems. Linear and Multilinear Algebra, 33(1–2):109–119, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barik:2008:TLE

[BLP08] S. Barik, A. K. Lal, and S. Pati. On trees with Laplacian eigenvalue one. Linear Multilinear Algebra, 56(6):597–610, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2012:ACG

[BLP12] R. B. Bapat, A. K. Lal, and S. Pati. On algebraic connectivity of graphs with at most two points of articulation in each block. Linear Multilinear Algebra, 60(4):415–432, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benitez:2012:IRC

[BLR12] Julio Benítez, Xiaoji Liu, and Vladimir Rakocević. Invertibility in rings of the commutator {ab − ba}, where aba = a and bab = b. Linear Multilinear Algebra, 60(4):449–463, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Blunck:2012:ISB

[BLSV12] Andrea Blunck, Péter Lévay, Metod Saniga, and Péter Vrana. Invertible symmetric 3 × 3 binary matrices and GQ(2, 4). Linear Multilinear Algebra, 60(10):1143–1154, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2018:AMO

[BLX18] Jianchao Bai, Jicheng Li, and Fengmin Xu. Accelerated method for optimization over density matrices in quantum state estimation. Linear Multilinear Algebra, 66(5):869–880, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benitez:2010:NGI

[BLZ10] Julio Benítez, Xiaoji Liu, and Tongping Zhu. Nonsingularity and group invertibility of linear combinations of two k-potent matrices. Linear Multilinear Algebra, 58(8):1023–1035, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benitez:2011:ARG

[BLZ11] Julio Benítez, Xiaoji Liu, and Tongping Zhu. Additive results for the group inverse in an algebra with applications to block operators. Linear Multilinear Algebra, 59(3):279–289, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2012:P

[BLZ12] Ravindra Bapat, Raphael Loewy, and Fuzhe Zhang. Preface. Linear Multilinear Algebra, 60(11–12):1223, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:1989:CMP

[BM89] Richard A. Brualdi and T. S. Michael. The class of 2multigraphs with a prescribed degree sequence. Linear and

Multilinear Algebra, 24(2):81–102, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brunat:2007:PGP

[BM07] Josep M. Brunat and Antonio Montes. A polynomial generalization of the power-compositions determinant. Linear Multilinear Algebra, 55(4):303–313, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bodine:2012:SAP

[BM12] E. J. Bodine and J. J. McDonald. Spectrally arbitrary patterns over finite fields. Linear Multilinear Algebra, 60(3):285–299, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bernik:2014:SLA

[BM14] Janez Bernik and Mitja Mastnak. On semitransitive Lie algebras of minimal dimension. Linear Multilinear Algebra, 62(2): 207–217, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bakherad:2015:CRI

[BM15a] Mojtaba Bakherad and Mohammad Sal Moslehian. Complementary and refined inequalities of Callebaut inequality for operators. Linear Multilinear Algebra, 63(8):1678–1692, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bakherad:2015:RVH

[BM15b] Mojtaba Bakherad and Mohammad Sal Moslehian. Reverses and variations of Heinz inequality. Linear Multilinear Algebra, 63(10):1972–1980, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bourhim:2015:MPL

[BM15c] Abdellatif Bourhim and Javad Mashreghi. Maps preserving the local spectrum of triple product of operators. Linear Multilinear Algebra, 63(4):765–773, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bukovsek:2016:JTP

[BM16] Damjana Kokol Bukovsek and Blaz Mojskerc. Jordan triple product homomorphisms on Hermitian matrices to and from dimension one. Linear Multilinear Algebra, 64(8):1669–1678, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Behera:2017:FRG

[BM17a] Ratikanta Behera and Debasisha Mishra. Further results on generalized inverses of tensors via the Einstein product. Linear Multilinear Algebra, 65(8):1662–1682, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bourhim:2017:NRP

[BM17b] Abdellatif Bourhim and Mohamed Mabrouk. Numerical radius and product of elements in C^{∗}-algebras. Linear Multilinear Algebra, 65(6):1108–1116, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bera:2018:CPS

[BM18a] Sudip Bera and Sajal Kumar Mukherjee. Combinatorial proofs of some determinantal identities. Linear Multilinear Algebra, 66(8):1659–1667, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bourhim:2018:NRC

[BM18b] Abdellatif Bourhim and Mohamed Mabrouk. Numerical radius characterizations of elements in -algebras. Linear Multilinear Algebra, 66(10):1939–1951, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:2018:CPI

[BM18c] Richard A. Brualdi and Seth A. Meyer. Combinatorial properties of integer matrices and integer matrices mod k. Linear Multilinear Algebra, 66(7):1380–1402, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [BM19b].

Bakherad:2019:SEO

[BM19a] Mojtaba Bakherad and Ali Morassaei. Some extensions of the operator entropy type inequalities. Linear Multilinear Algebra, 67(5):871–885, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:2019:CCP

[BM19b] Richard A. Brualdi and Seth A. Meyer. Corrigendum: Combinatorial properties of integer matrices and integer matrices mod k. Linear Multilinear Algebra, 67(3):647–648, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [BM18c].

Baghdad:2022:MNR

[BM22a] Abderrahim Baghdad and Chraibi Kaadoud Mohamed. On the maximal numerical range of the bimultiplication M_{2,A,B}. Linear Multilinear Algebra, 70(20):4907–4914, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Basha:2022:OFF

[BM22b] Aishah Ibraheam Basha and Judi J. McDonald. Orthogonality over finite fields. Linear Multilinear Algebra, 70(22):7277– 7289, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beiba:2022:JPO

[BM22c] El Moctar Ould Beiba and Sid Ahmed Ould Ahmed Mahmoud. On joint posinormality of operators. Linear Multilinear Algebra, 70(19):4819–4834, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1987:CGO

[BMD87] Natália Bebiano, Jorma Kaarlo Merikoski, and João Da Providência. On a conjecture of G. N. de Oliveira on determinants. Linear and Multilinear Algebra, 20(2):167–170, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bonilla:2021:HBD

[BMFPBSS21] A. Bonilla, G. A. Muñoz-Fernández, J. A. Prado-Bassas, and J. B. Seoane-Sepúlveda. Hausdorff and Box dimensions of continuous functions and lineability. Linear Multilinear Algebra, 69(4):593–606, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Burgos:2015:DJT

[BMGMC15] M. Burgos, A. C. Márquez-García, and A. Morales-Campoy. Determining Jordan (triple) homomorphisms by invertibility preserving conditions. Linear Multilinear Algebra, 63(7):1294– 1308, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Burgos:2016:MPO

[BMGMC16] M. Burgos, A. C. Márquez-García, and A. Morales-Campoy. Minus partial order and linear preservers. Linear Multilinear Algebra, 64(9):1814–1832, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

BenHassine:2022:BLS

[BMN22] Abdelkader Ben Hassine, Sami Mabrouk, and Othmen Ncib. 3-BiHom–Lie superalgebras induced by BiHom–Lie superalgebras. Linear Multilinear Algebra, 70(1):101–121, 2022. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botelho:2010:LSS

[BMP10] Geraldo Botelho, Mário C. Matos, and Daniel Pellegrino. Lineability of summing sets of homogeneous polynomials. Linear Multilinear Algebra, 58(1):61–74, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barik:2022:TRE

[BMP22] Sasmita Barik, Debabrota Mondal, and Sukanta Pati. Trees with the reciprocal eigenvalue property. Linear Multilinear Algebra, 70(21):6712–6717, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bru:1988:NJB

[BN88] R. Bru and M. Neumann. Nonnegative Jordan bases. Linear and Multilinear Algebra, 23(2):95–109, 1988. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baklouti:2022:COS

[BN22] Hamadi Baklouti and Sirine Namouri. Closed operators in semi-Hilbertian spaces. Linear Multilinear Algebra, 70(20): 5847–5858, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beik:2022:PIM

[BNKJ22] Fatemeh Panjeh Ali Beik, Mehdi Najafi-Kalyani, and Khalide Jbilou. Preconditioned iterative methods for multi-linear systems based on the majorization matrix. Linear Multilinear Algebra, 70(20):5827–5846, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barik:2006:NTR

[BNP06] S. Barik, M. Neumann, and S. Pati. On nonsingular trees and a reciprocal eigenvalue property. Linear Multilinear Algebra, 54(6):453–465, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bo:2001:GIB

[Bo01] Zhou Bo. On the generalized indices of boolean matrices. Linear and Multilinear Algebra, 49(2):131–141, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bekjan:2022:YTI

[BO22a] Turdebek N. Bekjan and Myrzagali N. Ospanov. On Youngtype inequalities of measurable operators. Linear Multilinear Algebra, 70(20):5606–5617, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bik:2022:SGP

[BO22b] Arthur Bik and Alessandro Oneto. On the strength of general polynomials. Linear Multilinear Algebra, 70(21):6114–6140, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boasso:2013:DST

[Boa13] Enrico Boasso. The Drazin spectrum of tensor product of Banach algebra elements and elementary operators. Linear Multilinear Algebra, 61(3):295–307, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boasso:2019:FRI

[Boa19] Enrico Boasso. Further results on the (b, c)-inverse, the outer inverse A_{T,S}^{(2)} and the Moore–Penrose inverse in the Banach context. Linear Multilinear Algebra, 67(5):1006–1030, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bolotnikov:2012:CPI

[Bol12] Dalia Bolotnikov. On Cowen–Pommerenke inequalities. Linear Multilinear Algebra, 60(2):249–254, 2012. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boman:2002:MPP

[Bom02] Eugene Boman. The Moore–Penrose pseudoinverse of an arbitrary, square, k-circulant matrix. Linear Multilinear Algebra, 50(2):175–179, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botta:1979:LTP

[Bot79] Peter Botta. Linear transformations preserving the unitary group. Linear and Multilinear Algebra, 8(2):89–96, 1979. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botta:1987:LMP

[Bot87] Peter Botta. Linear maps preserving rank less than or equal to one. Linear and Multilinear Algebra, 20(3):197–201, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botha:1994:FSM

[Bot94] J. D. Botha. Factorization of a singular matrix into nilpotent matrices with prescribed ranks. Linear and Multilinear Algebra, 38(1–2):145–150, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botha:1996:IFM

[Bot96] J. D. Botha. Idempotent factorization of matrices. Linear and Multilinear Algebra, 40(4):365–371, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botha:1997:TMC

[Bot97] J. D. Botha. Triangularizing matrices over GF(2) by congruence. Linear and Multilinear Algebra, 42(2):109–158, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boumazgour:2006:END

[Bou06] Mohamed Boumazgour. An estimate for the norm of a derivation on a norm ideal. Linear Multilinear Algebra, 54(5):321– 327, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Boumazgour:2013:DIU

[Bou13a] Mohamed Boumazgour. Drazin invertibility of upper triangular operator matrices. Linear Multilinear Algebra, 61(5): 627–634, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bourgeois:2013:SPM

[Bou13b] Gerald Bourgeois. Similar powers of a matrix. Linear Multilinear Algebra, 61(6):699–709, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bourgeois:2014:AMF

[Bou14a] Gérald Bourgeois. About the matrix function X → AX +XA. Linear Multilinear Algebra, 62(7):978–987, 2014. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [Bou14b].

Bourgeois:2014:CAM

[Bou14b] Gérald Bourgeois. Corrigendum: About the matrix function X → AX +XA. Linear Multilinear Algebra, 62(8):1137, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Bou14a].

Boumazgour:2018:GDI

[Bou18] Mohamed Boumazgour. Generalized Drazin inverse of restrictions of bounded linear operators. Linear Multilinear Algebra, 66(5):894–901, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barrett:1991:NEF

[BOvdD91] W. W. Barrett, D. D. Olesky, and P. van den Driessche. A note on eigenvalues of fixed rank perturbations of diagonal matrices. Linear and Multilinear Algebra, 30(1–2):13–16, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1995:PFT

[BOvdD95] R. B. Bapat, D. D. Olesky, and P. van den Driessche. Perron– Frobenius theory for a generalized eigenproblem. Linear and Multilinear Algebra, 40(2):141–152, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bozzo:1997:AMS

[Boz97] Enrico Bozzo. Algebras in matrix spaces with displacement structure. Linear and Multilinear Algebra, 42(1):23–41, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bozkurt:1998:NCTa

[Boz98a] Durmus Bozkurt. On the l_{p} norms of Cauchy–Toeplitz matrices. Linear and Multilinear Algebra, 44(4):341–346, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bozkurt:1998:NCTb

[Boz98b] Durmus Bozkurt. On the norms of Cauchy–Toeplitz matrices. Linear and Multilinear Algebra, 45(2–3):125–130, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bozkurt:1999:NHP

[Boz99] Durmus Bozkurt. On the _{p} norms of Hadamard product of Cauchy–Toeplitz and Cauchy–Hankel matrices. Linear and Multilinear Algebra, 45(4):333–339, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berman:1974:INM

[BP74] Abraham Berman and Robert J. Plemmons. Inverses of nonnegative matrices. Linear and Multilinear Algebra, 2(2):161– 172, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1989:LOS

[BP89] LeRoy B. Beasley and Norman J. Pullman. Linear operators that strongly preserve primitivity. Linear and Multilinear Algebra, 25(3):205–213, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1990:LOS

[BP90] Leroy B. Beasley and Norman J. Pullman. Linear operators strongly preserving digraphs whose maximum cycle length. Linear and Multilinear Algebra, 28(1–2):111–117, 1990. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1992:LOS

[BP92] LeRoy B. Beasley and Norman J. Pullman. Linear operators that strongly preserve the index of imprimitivity. Linear and Multilinear Algebra, 31(1–4):267–283, 1992. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1998:ACC

[BP98a] R. B. Bapat and Sukanta Pati. Algebraic connectivity and the characteristic set of a graph. Linear and Multilinear Algebra, 45(2–3):247–273, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1998:EMC

[BP98b] Natália Bebiano and Cecília Perdigão. Extremal matrices in certain determinantal inequalities. Linear and Multilinear Algebra, 44(3):261–276, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bremner:2009:PIT

[BP09] Murray R. Bremner and Luiz A. Peresi. Polynomial identities for the ternary cyclic sum. Linear Multilinear Algebra, 57(6): 595–608, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Besenyei:2012:CMT

[BP12] Ádám Besenyei and Dénes Petz. Characterization of mean transformations. Linear Multilinear Algebra, 60(3):255–265, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhat:2016:SEB

[BP16] Mushtaq A. Bhat and S. Pirzada. Spectra and energy of bipartite signed digraphs. Linear Multilinear Algebra, 64(9): 1863–1877, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Buckley:2018:EDR

[BP18] Anita Buckley and Bor Plestenjak. Explicit determinantal representations of up to quintic bivariate polynomials. Linear Multilinear Algebra, 66(11):2266–2285, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2019:ASI

[BP19] R. B. Bapat and S. K. Panda. Almost self-inverse graphs. Linear Multilinear Algebra, 67(10):2065–2076, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhunia:2022:SIN

[BP22] Pintu Bhunia and Kallol Paul. Some improvements of numerical radius inequalities of operators and operator matrices. Linear Multilinear Algebra, 70(10):1995–2013, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bebiano:1988:SCM

[BPD88] Natália Bebiano, Yiu Tung Poon, and Joâo Da Providência. On C-det spectral and C-det-convex matrices. Linear and Multilinear Algebra, 23(4):343–351, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zachlin:2008:NRM

[bPFZH08] Translated by Paul F. Zachlin and Michiel E. Hochstenbach. On the numerical range of a matrix. Linear Multilinear Algebra, 56(1–2):185–225, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botelho:2012:PSH

[BPR12] Geraldo Botelho, Daniel Pellegrino, and Pilar Rueda. Preduals of spaces of homogeneous polynomials on L_{p}-spaces. Linear Multilinear Algebra, 60(5):565–571, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botelho:2014:PMS

[BPR14] Geraldo Botelho, Daniel Pellegrino, and Pilar Rueda. On Pietsch measures for summing operators and dominated polynomials. Linear Multilinear Algebra, 62(7):860–874, 2014. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhat:2020:SRE

[BPR20] Mushtaq A. Bhat, S. Pirzada, and J. Rada. On spectra and real energy of complex weighted digraphs. Linear Multilinear Algebra, 68(5):1064–1076, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2000:RPM

[BPS00] R. B. Bapat, Sukanta Pati, and Seok-Zun Song. Rank preservers of matrices over max algebra. Linear and Multilinear Algebra, 48(2):149–164, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Belardo:2016:GWL

[BPS16] Francesco Belardo, Tomaz Pisanski, and Slobodan K. Simić. On graphs whose least eigenvalue is greater than −2. Linear Multilinear Algebra, 64(8):1570–1582, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See correction of author addresses [Ano16b].

Bremner:2012:MD

[BPSO12] Murray R. Bremner, Luiz A. Peresi, and Juana Sánchez-Ortega. Malcev dialgebras. Linear Multilinear Algebra, 60 (10):1125–1141, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Belardo:2016:SGW

[BPW16] Francesco Belardo, Pawel Petecki, and Jianfeng Wang. On signed graphs whose second largest Laplacian eigenvalue does not exceed 3. Linear Multilinear Algebra, 64(9):1785–1799, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bolotnikov:2001:PEC

[BR01] V. Bolotnikov and L. Rodman. Positive extension and completion problems for a class of structured matrices. Linear and Multilinear Algebra, 48(3):189–235, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bolotnikov:2004:FDB

[BR04] V. Bolotnikov and L. Rodman. Finite dimensional backward shift invariant subspaces of a class of reproducing kernel Hilbert spaces. Linear Multilinear Algebra, 52(5):321–334,

Benitez:2010:SLC

[BR10] Julio Benítez and Vladimir Rakocević. On the spectrum of linear combinations of two projections in C^{∗}-algebras. Linear Multilinear Algebra, 58(6):673–679, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2014:AMB

[BR14] R. B. Bapat and Souvik Roy. On the adjacency matrix of a block graph. Linear Multilinear Algebra, 62(3):406–418, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bhat:2019:LBE

[BR19] Mushtaq A. Bhat and Juan Rada. Lower bounds for the energy of weighted digraphs. Linear Multilinear Algebra, 67(4):743– 755, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bunger:2022:WDH

[BR22] Florian Bünger and Siegfried M. Rump. When does ||f(A)|| = f(||A||) hold true? Linear Multilinear Algebra, 70(21):7250– 7263, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brawley:1986:VIM

[Bra86] J. V. Brawley. Vandermonde independent matrices, basic sets, and polynomial matrix functions over a finite field. Linear and Multilinear Algebra, 19(4):373–386, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bremner:2012:HA

[Bre12a] Murray R. Bremner. On the hyperdeterminant for 2 × 2×3 arrays. Linear Multilinear Algebra, 60(8):921–932, 2012. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bresar:2012:MAM

[Bre12b] Matej Bresar. Multiplication algebra and maps determined by zero products. Linear Multilinear Algebra, 60(7):763–768,

Bremner:2014:PIB

[Bre14] Murray R. Bremner. A polynomial identity for the bilinear operation in Lie–Yamaguti algebras. Linear Multilinear Algebra, 62(12):1671–1682, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bresar:2015:AWN

[Bre15] Matej Bresar. Algebras in which non-scalar elements have small centralizers. Linear Multilinear Algebra, 63(9):1864– 1871, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bresar:2016:JDT

[Bre16] Matej Bresar. Jordan {g, h}-derivations on tensor products of algebras. Linear Multilinear Algebra, 64(11):2199–2207, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bremner:2017:OPD

[Bre17] Murray R. Bremner. One-parameter deformations of the diassociative and dendriform operads. Linear Multilinear Algebra, 65(2):341–350, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Breaz:2018:EFM

[Bre18] Simion Breaz. Endomorphisms of free modules as sums of four quadratic endomorphisms. Linear Multilinear Algebra, 66(11): 2215–2217, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Breaz:2021:LCM

[Bre21] Simion Breaz. Left comorphic matrix rings. Linear Multilinear Algebra, 69(12):2187–2191, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brison:1989:AGC

[Bri89] Owen J. Brison. Alternating groups and a conjecture about rational valued characters. Linear and Multilinear Algebra, 24(3):199–207, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benito:2022:LAF

[BRL22] P. Benito and J. Roldán-López. Lie algebras with a finite number of ideals. Linear Multilinear Algebra, 70(19):3702– 3721, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:1979:MPE

[Bru79a] Richard A. Brualdi. Matrices permutation equivalent to irreducible matrices and applications. Linear and Multilinear Algebra, 7(1):1–12, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdit:1979:DPM

[Bru79b] Richard A. Brualdit. Disjoint permutation matrices in complements of trees. Linear and Multilinear Algebra, 8(2):115–119, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:1982:MED

[Bru82] Richard A. Brualdi. Matrices eigenvalues, and directed graphs. Linear and Multilinear Algebra, 11(2):143–165, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:1985:IFP

[Bru85] Richard A. Brualdi. An interesting face of the polytope of doubly stochastic matrices. Linear and Multilinear Algebra, 17(1):5–18, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bruatdi:1987:RP

[Bru87] Richard A. Bruatdi. Reserch problem. Linear and Multilinear Algebra, 20(4):387–388, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:1991:HM

[BS91] R. B. Bapat and V. S. Sunder. On hypergroups of matrices. Linear and Multilinear Algebra, 29(2):125–140, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:1992:CNR

[BS92] Leroy B. Beasley and Seok-Zun Song. A comparison of nonnegative real ranks and their preservers. Linear and Multilinear Algebra, 31(1–4):37–46, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:1993:CBG

[BS93] Richard A. Brualdi and Bryan L. Shader. Cutsets in bipartite graphs. Linear and Multilinear Algebra, 34(1):51–54, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bell:1995:IBW

[BS95] Francis Bell and Slobodan Simić. On the index of broken wheels. Linear and Multilinear Algebra, 39(1–2):137–152, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berman:1997:SIP

[BS97] Abraham Berman and Dafna Shasha. Set inertia-preserving matrices. Linear and Multilinear Algebra, 43(1–3):169–179, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brown:2004:MEN

[BS04] Ethan S. Brown and Ilya M. Spitkovsky. On matrices with elliptical numerical ranges. Linear Multilinear Algebra, 52(3– 4):177–193, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brison:2007:EEA

[BS07] Owen J. Brison and Wasin So. Extendible elements of the alternating groups. Linear Multilinear Algebra, 55(1):65–70, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:2008:ISD

[BS08] Richard A. Brualdi and Jia-Yu Shao. Isolated sets of diagonals, diagonal hypergraphs, and simplices of doubly stochastic matrices. Linear Multilinear Algebra, 56(4):471–480, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bapat:2011:IDM

[BS11] R. B. Bapat and Sivaramakrishnan Sivasubramanian. Inverse of the distance matrix of a block graph. Linear Multilinear

Bremner:2013:CFA

[BS13] Murray R. Bremner and Stavros G. Stavrou. Canonical forms of 2 × 2 × 2 and 2 × 2 × 2 × 2 arrays over F_{2} and F_{3}. Linear Multilinear Algebra, 61(7):986–997, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bahsi:2015:SPM

[BS15a] Mustafa Bahsi and Süleyman Solak. Some particular matrices and their characteristic polynomials. Linear Multilinear Algebra, 63(10):2071–2078, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berg:2015:DCM

[BS15b] Christian Berg and Ryszard Szwarc. A determinant characterization of moment sequences with finitely many mass points. Linear Multilinear Algebra, 63(8):1568–1576, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:2016:ZTR

[BS16] LeRoy B. Beasley and Seok-Zun Song. Zero-term rank and zero-star cover number of symmetric matrices and their linear preservers. Linear Multilinear Algebra, 64(12):2368–2377, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barik:2017:DSC

[BS17a] Sasmita Barik and Gopinath Sahoo. On the distance spectra of coronas. Linear Multilinear Algebra, 65(8):1617–1628, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:2017:PSM

[BS17b] LeRoy B. Beasley and Seok-Zun Song. Primitive symmetric matrices and their preservers. Linear Multilinear Algebra, 65(1):129–139, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:2017:SAR

[BS17c] LeRoy B. Beasley and Seok-Zun Song. Symmetric Arctic ranks of nonnegative matrices and their linear preservers. Linear

Beik:2017:CIA

[BS17d] Fatemeh Panjeh Ali Beik and Davod Khojasteh Salkuyeh. A cyclic iterative approach and its modified version to solve coupled Sylvester-transpose matrix equations. Linear Multilinear Algebra, 65(12):2406–2423, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bisht:2017:IIM

[BS17e] Kavita Bisht and K. C. Sivakumar. Inheritance and inverse monotonicity properties of copositive matrices. Linear Multilinear Algebra, 65(5):897–908, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bernik:2018:CGR

[BS18a] Janez Bernik and Klemen Sivic. On certain graded representations of filiform Lie algebras. Linear Multilinear Algebra, 66 (11):2305–2327, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bisht:2018:VPS

[BS18b] Kavita Bisht and K. C. Sivakumar. Vanishing pseudo-Schur complements, reverse order laws, absorption laws and inheritance properties. Linear Multilinear Algebra, 66(1):167–183, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brits:2018:SCB

[BS18c] R. Brits and F. Schulz. The Shoda-completion of a Banach algebra. Linear Multilinear Algebra, 66(5):1001–1018, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:2018:DTM

[BS18d] Richard A. Brualdi and Bruce Sagan. Dihedral transportation and (0, 1)-matrix classes. Linear Multilinear Algebra, 66(12): 2557–2568, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:2020:GGS

[BS20] LeRoy B. Beasley and Seok-Zun Song. Genus of a graph and its strong preservers. Linear Multilinear Algebra, 68(8):1655– 1662, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baillon:2021:MNP

[BS21] Jean-Bernard Baillon and Alberto Seeger. On the maximal number of Pareto eigenvalues in a matrix of given order. Linear Multilinear Algebra, 69(7):1185–1207, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bandaliyev:2022:HTI

[BS22a] Rovshan A. Bandaliyev and Kamala H. Safarova. On Hardy type inequalities in grand Lebesgue spaces L_{p} for 0 < p ≤ 1. Linear Multilinear Algebra, 70(21):6053–6066, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:2022:MRM

[BS22b] LeRoy B. Beasley and Seok-Zun Song. Minimum-rank and maximum-nullity of graphs and their linear preservers. Linear Multilinear Algebra, 70(9):1732–1743, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Behera:2022:GIB

[BS22c] Ratikanta Behera and Jajati Keshari Sahoo. Generalized inverses of Boolean tensors via the Einstein product. Linear Multilinear Algebra, 70(3):531–556, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bresar:2022:ABF

[BS22d] Matej Bresar and Victor S. Shulman. On, around, and beyond Frobenius’ theorem on division algebras. Linear Multilinear Algebra, 70(7):1369–1381, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Blazquez-Sanz:2017:CAF

[BSA17] David Blázquez-Sanz and Carlos Alberto Marín Arango. Closed and asymptotic formulas for energy of some circulant graphs. Linear Multilinear Algebra, 65(6):1073–1079, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Brualdi:2012:EEI

[BSG^{+}12] Richard A. Brualdi, Jia-Yu Shao, Shi-Cai Gong, Chang-Qing Xu, and Guang-Hui Xu. On the extremal energy of integral weighted graphs. Linear Multilinear Algebra, 60(11–12):1255– 1264, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Beasley:2001:ZTR

[BSL01] Leroy B. Beasley, Seok-Zun Song, and Sang-Gu Lee. Zeroterm rank preservers. Linear and Multilinear Algebra, 48(4): 313–318, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berman:1998:RCP

[BSM98] Abraham Berman and Naomi Shaked-Monderer. Remarks on completely positive matrices. Linear and Multilinear Algebra, 44(2):149–163, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Burgos:2013:MPZ

[BSO13] María Burgos and Juana Sánchez-Ortega. On mappings preserving zero products. Linear Multilinear Algebra, 61(3):323– 335, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bremner:2017:QNQ

[BSO17] Murray R. Bremner and Juana Sánchez-Ortega. Quadratic nonsymmetric quaternary operads. Linear Multilinear Algebra, 65(8):1683–1703, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Butkovic:2012:ZME

[BSS12] Peter Butkovic, Hans Schneider, and Sergei Sergeev. Zmatrix equations in max-algebra, nonnegative linear algebra and other semirings. Linear Multilinear Algebra, 60(10):1191– 1210, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:2023:MPI

[BST23] Oskar Maria Baksalary, K. C. Sivakumar, and Götz Trenkler. On the Moore–Penrose inverse of a sum of matrices. Linear

Benitez:2013:NCT

[BSÜÖ13] Julio Benítez, Murat Sarduvan, Sedat Ülker, and Halim Özdemir. On nonsingularity of combinations of three group invertible matrices and three tripotent matrices. Linear Multilinear Algebra, 61(4):463–481, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bu:2013:SRD

[BSZW13] Changjiang Bu, Lizhu Sun, Jiang Zhou, and Yimin Wei. Some results on the Drazin inverse of anti-triangular matrices. Linear Multilinear Algebra, 61(11):1568–1576, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bart:1990:SRC

[BT90] H. Bart and G. Ph.A. Thijsse. Simultaneous reduction to companion and triangular forms of sets of matrices. Linear and Multilinear Algebra, 26(4):231–241, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bru:1998:GIG

[BT98] Rafael Bru and Néstor Thome. Group inverse and group involutory matrices. Linear and Multilinear Algebra, 45(2–3): 207–218, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benitez:2006:GSC

[BT06] Julio Benítez and Néstor Thome. The generalized Schur complement in group inverses and (k + 1)-potent matrices. Linear Multilinear Algebra, 54(6):405–413, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bondesson:2007:NMI

[BT07] Lennart Bondesson and Imbi Traat. A nonsymmetric matrix with integer eigenvalues. Linear Multilinear Algebra, 55(3): 239–247, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:2008:CEN

[BT08a] Oskar Maria Baksalary and Götz Trenkler. Characterizations of EP, normal, and Hermitian matrices. Linear Multilinear

Benitez:2008:ILC

[BT08b] J. Benítez and N. Thome. Idempotency of linear combinations of an idempotent matrix and a t-potent matrix that do not commute. Linear Multilinear Algebra, 56(6):679–687, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:2010:CIM

[BT10] Oskar Maria Baksalary and Götz Trenkler. Core inverse of matrices. Linear Multilinear Algebra, 58(6):681–697, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:2011:RFP

[BT11] Oskar Maria Baksalary and Götz Trenkler. Rank formulae from the perspective of orthogonal projectors. Linear Multilinear Algebra, 59(6):607–625, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:2013:CNS

[BT13] Oskar Maria Baksalary and Götz Trenkler. On column and null spaces of functions of a pair of oblique projectors. Linear Multilinear Algebra, 61(8):1116–1129, 2013. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bozkurt:2015:DIC

[BT15] Durmus Bozkurt and Tin-Yau Tam. Determinants and inverses of r-circulant matrices associated with a number sequence. Linear Multilinear Algebra, 63(10):2079–2088, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botelho:2017:TGH

[BT17] Geraldo Botelho and Ewerton R. Torres. Techniques to generate hyper-ideals of multilinear operators. Linear Multilinear Algebra, 65(6):1232–1246, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:2022:MWM

[BT22a] Oskar Maria Baksalary and Götz Trenkler. On matrices whose Moore–Penrose inverse is idempotent. Linear Multilinear Algebra, 70(11):2014–2026, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Botelho:2022:FCT

[BT22b] Geraldo Botelho and Leodan A. Torres. Factorization and coincidence theorems for generalized adjoints. Linear Multilinear Algebra, 70(5):899–908, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baksalary:2020:MPM

[BTT20] Oskar Maria Baksalary, Dietrich Trenkler, and Götz Trenkler. On most perfect magic squares of order four. Linear Multilinear Algebra, 68(7):1411–1423, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bowman:2007:TGS

[BTV07] Kevin Bowman, David A. Towers, and Vicente R. Varea. Two generator subalgebras of Lie algebras. Linear Multilinear Algebra, 55(5):429–438, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bunce:1991:BCT

[Bun91] John W. Bunce. The Binet–Cauchy theorem and inequalities for the Schur power matrix. Linear and Multilinear Algebra, 30 (1–2):5–11, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bunch:1994:LMA

[Bun94] The Hadwin Lunch Bunch. Local multiplications on algebras spanned by idempotents. Linear and Multilinear Algebra, 37(4):259–263, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Burghduff:1995:MPD

[Bur95] John Brian Burghduff. Minimum permanents of doubly stochastic matrices with zero main diagonal. Linear and Multilinear Algebra, 40(2):125–140, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Burgdorf:2011:SHS

[Bur11] Sabine Burgdorf. Sums of Hermitian squares as an approach to the BMV conjecture. Linear Multilinear Algebra, 59(1):1–9, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Buse:2023:FET

[Bus23] Laurent Busé. Formulas for the eigendiscriminants of ternary and quaternary forms. Linear Multilinear Algebra, 71(11): 1755–1774, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Butler:2010:EEC

[But10a] Steve Butler. Eigenvalues of 2-edge-coverings. Linear Multilinear Algebra, 58(4):413–423, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Butler:2010:NAC

[But10b] Steve Butler. A note about cospectral graphs for the adjacency and normalized Laplacian matrices. Linear Multilinear Algebra, 58(3):387–390, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Benito:1992:LIN

[BV92] M. P. Benito and V. R. Varea. The lattice of ideals of a nilpotent Lie algebra. Linear and Multilinear Algebra, 33(3–4):203– 215, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barreira:2016:LCR

[BV16a] Luis Barreira and Claudia Valls. Lyapunov’s coefficient of regularity and nonuniform hyperbolicity. Linear Multilinear Algebra, 64(6):1124–1144, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barreira:2016:NES

[BV16b] Luis Barreira and Claudia Valls. Nonuniform exponential stability and admissibility. Linear Multilinear Algebra, 64(3): 440–465, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bender:1984:GAD

[BVW84] E. Bender, G. Viennot, and S. G. Williamson. Global analysis of the delete-contract recursion for graphs and matroids. Linear and Multilinear Algebra, 15(2):133–160, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2015:CLA

[BW15] Rui-Pu Bai and Yong Wu. Constructions of 3-Lie algebras. Linear Multilinear Algebra, 63(11):2171–2186, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barron:2023:EP

[BW23] Tatyana Barron and Noah Wheatley. Entanglement and products. Linear Multilinear Algebra, 71(5):756–767, 2023. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barnett:1983:EEC

[BWF83] Fred C. Barnett, James R. Weaver, and J. S. Frame. Eigenvalues and eigenvectors of a certain stochastic matrix. Linear and Multilinear Algebra, 13(4):345–350, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2012:MEL

[BWLL12] Ruipu Bai, Wanqing Wu, Ying Li, and Zhenheng Li. Module extensions of 3-Lie algebras. Linear Multilinear Algebra, 60(4): 433–447, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berman:2007:UMM

[BX07] Abraham Berman and Changqing Xu. Uniform and minimal {0, 1}–cp matrices. Linear Multilinear Algebra, 55(5):439–456, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bedratyuk:2011:MPA

[BX11] Leonid Bedratyuk and Guoce Xin. MacMahon partition analysis and the Poincaré series of the algebras of invariants of ternary and quaternary forms. Linear Multilinear Algebra, 59(7):789–799, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ballantine:1976:UNC

[BY76] C. S. Ballantine and E. L. Yip. Uniqueness of the “nonsingular core” for Hermitian and other matrix pencils. Linear and Multilinear Algebra, 4(1):61–67, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bai:2019:HIM

[BY19] Ruiqi Bai and Yichuan Yang. Hereditary l-ideals of matrix rings over l-rings. Linear Multilinear Algebra, 67(12):2540– 2548, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bytsko:2020:RJW

[Byt20] Andrei Bytsko. A relation for the Jones–Wenzl projector and tensor space representations of the Temperley–Lieb algebra. Linear Multilinear Algebra, 68(11):2239–2253, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bevilacqua:1989:CCM

[BZ89] R. Bevilacqua and P. Zellini. Closure, commutativity and minimal complexity of some spaces of matrices. Linear and Multilinear Algebra, 25(1):1–25, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baragana:1990:CCP

[BZ90] Itziar Baragaña and Ion Zaballa. Column completion of a pair of matrices. Linear and Multilinear Algebra, 27(4):243–273, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baragana:1999:FIP

[BZ99] I. Baragaña and I. Zaballa. Feedback invariants of pairs of matrices and restrictions. Linear and Multilinear Algebra, 46(1– 2):101–116, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Berman:2000:NDA

[BZ00] Abraham Berman and Xiao-Dong Zhang. A note on degree antiregular graphs. Linear and Multilinear Algebra, 47(4):307– 311, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bekjan:2019:SPS

[BZ19] Turdebek N. Bekjan and Makpal Zhaxylykova. Some properties of semifinite tracial subalgebras. Linear Multilinear Algebra, 67(6):1190–1197, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

BenOuidren:2022:NAW

[BZ22a] Kaoutar Ben Ouidren and Hassan Zariouh. New approach to a-Weyl’s theorem through localized SVEP and Riesz-type perturbations. Linear Multilinear Algebra, 70(17):3231–3247, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bovdi:2022:RMS

[BZ22b] Victor Bovdi and Bohdan Zabavsky. Reduction of matrices over simple Ore domains. Linear Multilinear Algebra, 70(4): 642–649, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bu:2010:SRG

[BZZ10] Changjiang Bu, Kuize Zhang, and Jiemei Zhao. Some results on the group inverse of the block matrix with a sub-block of linear combination or product combination of matrices over skew fields. Linear Multilinear Algebra, 58(8):957–966, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bu:2011:RDI

[BZZ11] Changjiang Bu, Kuize Zhang, and Jiemei Zhao. Representations of the Drazin inverse on solution of a class singular differential equations. Linear Multilinear Algebra, 59(8):863–877, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Coelho:2003:CIA

[CA03] M. PurificaÇão Coelho and M. Antónia. On the conversion of an immanant into another on symmetric matrices. Linear Multilinear Algebra, 51(2):137–145, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cafuta:2013:MAA

[Caf13] Kristijan Cafuta. On matrix algebras associated to sum-of-squares semidefinite programs. Linear Multilinear Algebra, 61 (11):1496–1509, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cain:1993:BSS

[Cai93] Bryan E. Cain. Between d-stability and sign-stability: the 3 × 3 case. Linear and Multilinear Algebra, 36(1):41–45, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Callan:1991:ADC

[Cal91] David Callan. Another determinant condensation formula. Linear and Multilinear Algebra, 29(2):93–98, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Caldeira:1994:DMP

[Cal94] Cristina Caldeira. On the degree of the minimal polynomial of some partial derivations. Linear and Multilinear Algebra, 38(1–2):83–95, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Callan:1995:EPM

[Cal95a] D. Callan. On evaluating permanents and a matrix of cotangents. Linear and Multilinear Algebra, 38(3):193–205, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Callan:1995:RP

[Cal95b] David Callan. Research problem. Linear and Multilinear Algebra, 39(4):401–403, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Caldeira:1997:CPM

[Cal97] Cristina Caldeira. Critical pairs of matrices for the degree of the minimal polynomial of the Kronecker sum. Linear and Multilinear Algebra, 42(1):73–88, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See correction [Cal99].

Caldeira:1999:CCP

[Cal99] Cristina Caldeira. Correction to “Critical pairs of matrices for the degree of the minimal polynomial of the Kronecker sum”. Linear and Multilinear Algebra, 46(4):277–280, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Cal97].

Caldeira:2003:LBD

[Cal03] Cristina Caldeira. A lower bound for the deree of the minimal polynomial of the Kronecker product. Linear Multilinear Algebra, 51(3):243–262, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Campbell:1978:LPM

[Cam78] Stephen L. Campbell. On the limit of a product of matrix exponentials. Linear and Multilinear Algebra, 6(1):55–59, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Campbell:1980:CDI

[Cam80] Stephen L. Campbell. Continuity of the Drazin inverse. Linear and Multilinear Algebra, 8(3):265–268, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Campbell:1983:DIS

[Cam83] Stephen L. Campbell. The Drazin inverse and systems of second order linear differential equations. Linear and Multilinear Algebra, 14(2):195–198, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Campos:2014:CMC

[Cam14] Jamilson R. Campos. Cohen and multiple Cohen strongly summing multilinear operators. Linear Multilinear Algebra, 62(3):322–346, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Canfield:1980:SPP

[Can80] E. Rodney Canfield. A Sperner property preserved by product. Linear and Multilinear Algebra, 9(2):151–157, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2019:MCM

[Cao19] Lei Cao. A minimal completion of (0, 1)-matrices without total support. Linear Multilinear Algebra, 67(8):1522–1530, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cueto-Avellaneda:2020:MUP

[CAP20] María Cueto-Avellaneda and Antonio M. Peralta. The Mazur– Ulam property for commutative von Neumann algebras. Linear Multilinear Algebra, 68(2):337–362, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cueto-Avellaneda:2022:COI

[CAP22] María Cueto-Avellaneda and Antonio M. Peralta. Can one identify two unital JB*-algebras by the metric spaces determined by their sets of unitaries? Linear Multilinear Algebra, 70(22):7702–7727, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carini:1988:CSF

[Car88] Luisa Carini. The cocharacter sequence of the finite dimensional Grassmann algebra. Linear and Multilinear Algebra, 23(2):131–137, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carini:1990:PSR

[Car90] Luisa Carini. The Poincaré series related to the Grassmann algebra. Linear and Multilinear Algebra, 27(3):199–205, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carini:2006:CCS

[Car06] Luisa Carini. Computing cocharacters of sign trace identities in reduced notation. Linear Multilinear Algebra, 54(2):147– 156, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cariello:2019:SKT

[Car19] D. Cariello. Sinkhorn–Knopp theorem for rectangular positive maps. Linear Multilinear Algebra, 67(11):2345–2365, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carvalho:2020:APH

[Car20] Sónia Carvalho. An alternative proof on higher order derivatives of a multilinear map. Linear Multilinear Algebra, 68(7): 1457–1464, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Caserta:1991:EID

[Cas91] Francesco Caserta. Eigenvalues in infinite dimensions. Linear and Multilinear Algebra, 28(4):195–201, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Casteels:2023:STP

[Cas23] Karel Casteels. Scaffoldings of totally positive matrices and line insertion. Linear Multilinear Algebra, 71(11):1888–1904, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Catalano:2021:MPP

[Cat21] Louisa Catalano. On maps preserving products equal to a rank-one idempotent. Linear Multilinear Algebra, 69(4):673– 680, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cavers:2010:RMP

[Cav10] M. S. Cavers. On reducible matrix patterns. Linear Multilinear Algebra, 58(2):257–267, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cassier:2022:HPS

[CB22] Gilles Cassier and Mohammed Benharrat. Harnack parts for some truncated shifts. Linear Multilinear Algebra, 70(5):974– 992, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Che:2019:NTR

[CBQW19] Maolin Che, Changjiang Bu, Liqun Qi, and Yimin Wei. Nonnegative tensors revisited: plane stochastic tensors. Linear Multilinear Algebra, 67(7):1364–1391, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:SMH

[CBZ22] Lixiang Chen, Changjiang Bu, and Jiang Zhou. Spectral moments of hypertrees and their applications. Linear Multilinear Algebra, 70(21):6297–6311, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

chuan:1985:FSI

[cc85] Jen-Chung chuan and Wai-Fong chuan. Factorizations in a semigroup of integral matrices. Linear and Multilinear Algebra, 18(3):213–223, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cassam-Chenai:1992:GPR

[CC92] P. Cassam-Chenaï. A generalization of the plücker relations. Linear and Multilinear Algebra, 31(1–4):77–79, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Craven:1998:SCS

[CC98] Thomas Craven and George Csordas. A sufficient condition for strict total positivity of a matrix. Linear and Multilinear Algebra, 45(1):19–34, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:2014:TDP

[CC14] Jor-Ting Chan and Kong Chan. Two distance preservers of generalized numerical radii. Linear Multilinear Algebra, 62(5): 674–682, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cai:2017:ISG

[CC17a] Jing Cai and Jianlong Chen. Iterative solutions of generalized inverse eigenvalue problem for partially bisymmetric matrices. Linear Multilinear Algebra, 65(8):1643–1654, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cardell:2017:CPP

[CC17b] Sara D. Cardell and Joan-Josep Climent. A construction of primitive polynomials over finite fields. Linear Multilinear Algebra, 65(12):2424–2431, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cui:2018:WTU

[CC18] Miaomiao Cui and Xiaohong Cao. Weyl’s theorem for upper triangular operator matrix and perturbations. Linear Multilinear Algebra, 66(7):1299–1310, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:2019:BLC

[CC19] Yuan Chang and Liangyun Chen. Biderivations and linear commuting maps on the restricted Cartan-type Lie algebras W (n; 1) and S(n; 1). Linear Multilinear Algebra, 67(8):1625– 1636, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2021:RCI

[CC21] Xiaofeng Chen and Jianlong Chen. Right core inverses of a product and a companion matrix. Linear Multilinear Algebra, 69(12):2245–2263, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2022:MPI

[CC22a] Wensheng Cao and Zhenhu Chang. The Moore–Penrose inverses of split quaternions. Linear Multilinear Algebra, 70(9): 1631–1647, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cirulis:2022:EJM

[CC22b] Janis Cirulis and Insa Cremer. On existence of joins and meets under the star order in strong Rickart rings. Linear Multilinear Algebra, 70(22):7370–7383, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cahill:2023:OPF

[CC23a] Jameson Cahill and Peter G. Casazza. Optimal Parseval frames: total coherence and total volume. Linear Multilinear Algebra, 71(13):2067–2092, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chiru:2023:NRM

[CC23b] Iulia-Elena Chiru and Septimiu Crivei. Von Neumann regular matrices revisited. Linear Multilinear Algebra, 71(8):1352– 1363, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:2021:SBG

[CCC21] Yuan Chang, Liangyun Chen, and Yan Cao. Superbiderivations of the generalized Witt Lie superalgebra W (m, n; t). Linear Multilinear Algebra, 69(2):233–244, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Camacho:2011:FLA

[CCGO11] L. M. Camacho, E. M. Cañete, J. R. Gómez, and B. A. Omirov. 3-filiform Leibniz algebras of maximum length, whose naturally graded algebras are Lie algebras. Linear Multilinear Algebra, 59(9):1039–1058, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2020:PDS

[CCKL20] Lei Cao, Zhi Chen, Selcuk Koyuncu, and Huilan Li. Permanents of doubly substochastic matrices. Linear Multilinear Algebra, 68(3):594–605, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:2010:PIQ

[CCL10] Che-Man Cheng, Kam-Loi Cheok, and Ieng-Tak Leong. 3 × 3 pure imaginary quaternionic solutions of the Hurwitz matrix equations. Linear Multilinear Algebra, 58(7):863–874, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Catalano:2021:MPR

[CCL21] Louisa Catalano and Megan Chang-Lee. On maps preserving rank-one nilpotents. Linear Multilinear Algebra, 69(16):3092– 3098, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2022:SHS

[CCLL22] Lei Cao, Zhi Chen, Qiang Li, and Huilan Li. Symmetric and Hankel-symmetric transportation polytopes. Linear Multilinear Algebra, 70(5):955–973, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cui:2016:EPE

[CCLN16] Lu-Bin Cui, Chuan Chen, Wen Li, and Michael K. Ng. An eigenvalue problem for even order tensors with its applications.

Linear Multilinear Algebra, 64(4):602–621, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1998:MRA

[CCS98] Wai-Kiu Chan, Yip-Cheung Chan, and Man-Keung Siu. Minimal rank of abelian group matrices. Linear and Multilinear Algebra, 44(3):277–285, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Canto:2016:SSR

[CCS16] B. Cantó, C. Coll, and E. Sánchez. On the stability and spectral radius of a finite set of matrices. Linear Multilinear Algebra, 64(3):353–361, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Canto:2021:AJC

[CCU21] Begoña Cantó, Rafael Cantó, and Ana M. Urbano. All Jordan canonical forms of irreducible totally non-negative matrices. Linear Multilinear Algebra, 69(13):2389–2409, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2020:EPC

[CCW20] Zhi Chen, Lei Cao, and Qing-Wen Wang. The extreme points of certain polytopes of doubly substochastic matrices. Linear Multilinear Algebra, 68(10):1956–1971, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic:1985:DTG

[CD85] Dragos Cvetković and Michael Doob. Developments in the theory of graph spectra. Linear and Multilinear Algebra, 18(2): 153–181, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Coelho:1998:CIA

[CD98] M. Purificaçãao Coelho and M. Antónia Duffner. On the conversion of an immanant into another. Linear and Multilinear Algebra, 44(2):111–130, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Coelho:2001:SWI

[CD01] M. Purificação Coelho and M. Antónia Duffner. Subspaces where an immanant is convertible into its conjugate. Linear and Multilinear Algebra, 48(4):383–408, 2001. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Coelho:2003:RBD

[CD03] M. Purificação Coelho and M. Antónia Duffner. On the relation between the determinant and the permanent on symmetric matrices. Linear Multilinear Algebra, 51(2):127–136, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cooper:2015:CHS

[CD15] Joshua Cooper and Aaron Dutle. Computing hypermatrix spectra with the Poisson product formula. Linear Multilinear Algebra, 63(5):956–970, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2021:MNG

[CD21] Sheng Chen and Yi Dai. Minimal ∞-norm of generalized inverses of the incidence matrix of a tree. Linear Multilinear Algebra, 69(7):1224–1235, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cezayirlioglu:2022:FID

[CD22] Nefise Cezayirlioğlu and Çağri Demir. Functional identities of degree 2 vanishing on zero products of XY and Y X. Linear Multilinear Algebra, 70(20):4876–4896, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Contino:2021:GSA

[CDF21] M. Contino, M. E. Di Iorio y Lucero, and G. Fongi. Global solutions of approximation problems in Hilbert spaces. Linear Multilinear Algebra, 69(13):2510–2526, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Conde:2023:FSC

[CDG23] Cristian Conde, Ezequiel Dratman, and Luciano N. Grippo. Finding singularly cospectral graphs. Linear Multilinear Algebra, 71(3):496–512, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Charfi:2013:RBF

[CDJ13] Salma Charfi, Alaeddine Damergi, and Aref Jeribi. On a Riesz basis of finite-dimensional invariant subspaces and application to Gribov operator in Bargmann space. Linear Multilinear Algebra, 61(11):1577–1591, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Christensen:2020:EFG

[CDK20] Ole Christensen, Somantika Datta, and Rae Young Kim. Equiangular frames and generalizations of the Welch bound to dual pairs of frames. Linear Multilinear Algebra, 68(12): 2495–2505, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ciardo:2022:KCT

[CDK22] Lorenzo Ciardo, Geir Dahl, and Steve Kirkland. On Kemeny’s constant for trees with fixed order and diameter. Linear Multilinear Algebra, 70(12):2331–2353, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Centrone:2015:ZGI

[CdM15] Lucio Centrone and Thiago Castilho de Mello. On z_{n}-graded identities of block-triangular matrices. Linear Multilinear Algebra, 63(2):302–313, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carvalho:2022:NPQ

[CDM22] Luís Carvalho, Cristina Diogo, and Sérgio Mendes. A new perspective on the quaternionic numerical range of normal matrices. Linear Multilinear Algebra, 70(20):5068–5074, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chattopadhyay:2023:CST

[CDPS23] Arup Chattopadhyay, Soma Das, Chandan Pradhan, and Srijan Sarkar. Characterization of C-symmetric Toeplitz operators for a class of conjugations in Hardy spaces. Linear Multilinear Algebra, 71(12):2026–2048, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Caldeira:2003:YSA

[CdS03] Cristina Caldeira and J. A. Dias da Silva. Young symmetrizers and additive theory. Linear Multilinear Algebra, 51(4):339– 359, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Centrone:2017:GGP

[CdSS17] Lucio Centrone and Manuela da Silva Souza. On the growth of graded polynomial identities of ∫↕_{\}. Linear Multilinear Algebra, 65(4):752–767, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Centrone:2011:GGK

[Cen11] Lucio Centrone. The graded Gelfand–Kirillov dimension of verbally prime algebras. Linear Multilinear Algebra, 59(12): 1433–1450, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Compta:2015:GCM

[CF15] A. Compta and J. Ferrer. Geometric classification of monogenic subspaces and uniparametric linear control systems. Linear Multilinear Algebra, 63(9):1768–1785, 2015. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2021:PST

[CF21] Xiwang Cao and Keqin Feng. Perfect state transfer on Cayley graphs over dihedral groups. Linear Multilinear Algebra, 69(2): 343–360, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Conde:2023:SNR

[CF23] Cristian Conde and Kais Feki. Some numerical radius inequality for several semi-Hilbert space operators. Linear Multilinear Algebra, 71(6):1054–1071, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Comellas:2006:WDR

[CFGM06] F. Comellas, M. A. Fiol, J. Gimbert, and M. Mitjana. On the weak distance-regularity of Moore-type digraphs. Linear Multilinear Algebra, 54(4):265–284, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:2022:ICG

[CFL22a] Tao Cheng, Lihua Feng, and Weijun Liu. Integral Cayley graphs over a certain nonabelian group. Linear Multilinear Algebra, 70(21):7224–7235, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:2022:DPI

[CFL^{+}22b] Tao Cheng, Lihua Feng, Weijun Liu, Lu Lu, and Dragan Stevanović. Distance powers of integral Cayley graphs over dihedral groups and dicyclic groups. Linear Multilinear Algebra, 70(7):1281–1290, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2023:SLS

[CFLL23] Zhouyang Chen, Lihua Feng, Weijun Liu, and Lu Lu. The signless p-Laplacian spectral radius of graphs with given matching number. Linear Multilinear Algebra, 71(14):2366– 2374, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Compta:2010:LDP

[CFP10] Albert Compta, Josep Ferrer, and Marta Peña. Local differentiable pole assignment. Linear Multilinear Algebra, 58(5): 563–569, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Compta:2021:SCF

[CFP21] Albert Compta, Josep Ferrer, and Marta Peña. State canonical form for a class of uncontrollable linear systems. Linear Multilinear Algebra, 69(13):2441–2468, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Casanellas:2018:ISA

[CFSGL18] Marta Casanellas, Jesús Fernández-Sánchez, and Marina Garrote-López. The inertia of the symmetric approximation for low-rank matrices. Linear Multilinear Algebra, 66(11): 2349–2353, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Casanellas:2022:OSE

[CFSRL22] Marta Casanellas, Jesús Fernández-Sánchez, and Jordi Roca-Lacostena. An open set of 4 × 4 embeddable matrices whose principal logarithm is not a Markov generator. Linear Multilinear Algebra, 70(19):3768–3779, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2023:GBS

[CFZ23] Yanni Chen, Tiantian Fu, and Ye Zhang. The generalized Banach sequence spaces and their dual spaces based on permutation symmetric gauge norms. Linear Multilinear Algebra, 71(13):2249–2261, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Coutinho:2016:PST

[CG16] G. Coutinho and C. Godsil. Perfect state transfer in products and covers of graphs. Linear Multilinear Algebra, 64(2):235– 246, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Castillo:2017:UKP

[CG17] J. E. Castillo and R. D. Grone. Using Kronecker products to construct mimetic gradients. Linear Multilinear Algebra, 65 (10):2031–2045, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheraghpour:2019:INU

[CG19] Hassan Cheraghpour and Nader M. Ghosseiri. On the idempotents, nilpotents, units and zero-divisors of finite rings. Linear Multilinear Algebra, 67(2):327–336, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cilia:2022:RRM

[CG22] Raffaella Cilia and Joaquín M. Gutiérrez. Riesz-representable multilinear mappings. Linear Multilinear Algebra, 70(11): 2027–2046, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Crouzeix:2014:PBS

[CGH14] Michel Crouzeix, Frank Gilfeather, and John Holbrook. Polynomial bounds for small matrices. Linear Multilinear Algebra, 62(5):614–625, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:2015:LTQ

[CGH15] Sha Chang, Suogang Gao, and Bo Hou. The Leonard triples with quantum parameter being not a unit root. Linear Multilinear Algebra, 63(12):2546–2560, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choi:2008:GHR

[CGHK08] Man-Duen Choi, Michael Giesinger, John A. Holbrook, and David W. Kribs. Geometry of higher-rank numerical ranges. Linear Multilinear Algebra, 56(1–2):53–64, 2008. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2016:PON

[CGL^{+}16] Mao-Ting Chien, Hwa-Long Gau, Chi-Kwong Li, Ming-Cheng Tsai, and Kuo-Zhong Wang. Product of operators and numerical range. Linear Multilinear Algebra, 64(1):58–67, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Contino:2019:SOM

[CGM19] M. Contino, J. I. Giribet, and A. Maestripieri. Shorted operators and minus order. Linear Multilinear Algebra, 67(11): 2173–2190, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cardoso:2011:GFL

[CGMR11] Domingos M. Cardoso, Ivan Gutman, Enide Andrade Martins, and María Robbiano. A generalization of Fiedler’s lemma and some applications. Linear Multilinear Algebra, 59(8):929–942, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cairns:2015:GBL

[CGNT15] Grant Cairns, Ana Hinić Galić, Yuri Nikolayevsky, and Ioannis Tsartsaflis. Geodesic bases for Lie algebras. Linear Multilinear Algebra, 63(6):1176–1194, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Camacho:2013:DSE

[CGOT13] L. M. Camacho, J. R. Gómez, B. A. Omirov, and R. M. Turdibaev. The derivations of some evolution algebras. Linear Multilinear Algebra, 61(3):309–322, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:2014:EHR

[CGW14] Chi-Tung Chang, Hwa-Long Gau, and Kuo-Zhong Wang. Equality of higher-rank numerical ranges of matrices. Linear

Cao:2013:MPC

[CGY13] C. G. Cao, Y. L. Ge, and H. M. Yao. Maps preserving classical adjoint of products of two matrices. Linear Multilinear Algebra, 61(12):1593–1604, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choudhury:1987:ASV

[CH87] Dipa Choudhury and Roger A. Horn. An analog of the singular value decomposition for complex orthogonal equivalence. Linear and Multilinear Algebra, 21(2):149–162, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2006:PIA

[CH06] Xuzhou Chen and Robert E. Hartwig. The Picard iteration and its application. Linear Multilinear Algebra, 54(5):329–341, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cho:2010:RNR

[CH10] Muneo Cho and Tadasi Huruya. A remark on numerical range of semi-hyponormal operators. Linear Multilinear Algebra, 58(6):711–714, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2012:ESM

[CH12] Gong-Ning Chen and Yong-Jian Hu. Extremal solutions of matricial Carathéodory problem and Nevanlinna–Pick-type interpolation for Carathéodory matrix functions in the nondegenerate case. Linear Multilinear Algebra, 60(11–12):1271–1294, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:MSR

[CH22a] Qianqian Chen and Qiongxiang Huang. The maximal A_{α}- spectral radius of graphs with given matching number. Linear Multilinear Algebra, 70(20):5193–5206, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choi:2022:ICM

[CH22b] Jeong-Ok Choi and Youngmi Hur. Invertibility of circulant matrices of arbitrary size. Linear Multilinear Algebra, 70(21): 7057–7074, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1977:CMG

[Cha77] G. H. Chan. A characterization of minimal (k)-groups of degree n ≤ 3k. Linear and Multilinear Algebra, 4(4):285–305, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1978:TSC

[Cha78] G. H. Chan. On the triviality of a symmetry class of tensors. Linear and Multilinear Algebra, 6(1):73–82, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1979:MGD

[Cha79a] G. H. Chan. Minimal (k)-groups of degree n, 3k < n ≤ 4k. Linear and Multilinear Algebra, 7(2):155–166, 1979. CODEN LNMLA2. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1979:PGS

[Cha79b] G. H. Chan. Permutations on Grassmann spaces. Linear and Multilinear Algebra, 7(3):241–247, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:1983:NCM

[Cha83a] Derek K. Chang. A note on a conjecture of M. Marcus and H. Minc. Linear and Multilinear Algebra, 13(2):115–117, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:1983:NPD

[Cha83b] Derek K. Chang. Notes on permanents of doubly stochastic matrices. Linear and Multilinear Algebra, 14(4):349–356, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:1984:MPD

[Cha84a] Derek. K. Chang. Minimum permanents of doubly stochastic matrices with one fixed entry. Linear and Multilinear Algebra, 15(3–4):313–317, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:1984:NCH

[Cha84b] Derek K. Chang. A note on a conjecture of T. H. Foregger. Linear and Multilinear Algebra, 15(3–4):341–344, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:1988:MMP

[Cha88] Derek K. Chang. Minimum and maximum permanents of certain doubly stochastic matrices. Linear and Multilinear Algebra, 24(1):39–44, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1989:SMC

[Cha89] Chi Fai Chan. Some more on a conjecture of Marcus and Wang. Linear and Multilinear Algebra, 25(3):231–235, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1990:PSD

[Cha90a] Chi Fai Chan. Positive semi-definiteness in the complex group algebra. Linear and Multilinear Algebra, 27(3):189–198, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:1990:TPC

[Cha90b] Derek K. Chang. On two permanental conjectures. Linear and Multilinear Algebra, 26(3):207–213, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1992:GNR

[Cha92] Chi Fai Chan. The generalized numerical radius associated with a positive semi-definite function. Linear and Multilinear Algebra, 32(3–4):265–281, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1993:GNR

[Cha93a] Chi Fai Chan. The generalized numerical radius associated with a positive semi-definite function II. Linear and Multilinear Algebra, 34(3–4):247–259, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1993:LE

[Cha93b] Chi Fai Chan. Letter to the editor. Linear and Multilinear Algebra, 35(3–4):355–356, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:1996:GSN

[Cha96] Shu-Chu Chang. Generalized Smith normal form class number w.r.t Γ_{0}(n). Linear and Multilinear Algebra, 41(3):211–219, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:2018:NBJ

[Cha18] Jor-Ting Chan. A note on the boundary of the joint numerical range. Linear Multilinear Algebra, 66(4):821–826, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:1991:APR

[Che91] Che-Man Cheng. An approximation problem with respect to symmetric gauge functions with application to matrix spaces. Linear and Multilinear Algebra, 29(3–4):169–180, 1991. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:1992:NDM

[Che92] Yonglin Chen. Nonnegative definite matrices and their applications to matrix quadratic programming problems. Linear and Multilinear Algebra, 33(3–4):189–201, 1992. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:1993:RCR

[Che93a] Yong-Lin Chen. Representations and Cramer rules for the solution of a restricted matrix equation. Linear and Multilinear Algebra, 35(3–4):339–354, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:1993:CRS

[Che93b] Yonglin Chen. A Cramer rule for solution of the general restricted linear equation. Linear and Multilinear Algebra, 34(2): 177–186, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:1994:DNR

[Che94] Che-Man Cheng. On the decomposable numerical range of λI − N. Linear and Multilinear Algebra, 37(1–3):197–205, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:1995:FAG

[Che95a] Yong-Lin Chen. Finite algorithms for the (2)-generalized inverse. Linear and Multilinear Algebra, 40(1):61–68, 1995. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cherix:1995:PEC

[Che95b] Pierre-Alain Cherix. Property (T) and expanding constants for semidirect products. Linear and Multilinear Algebra, 39(1– 2):153–160, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheung:2003:LDT

[Che03a] Wai-Shun Cheung. Lie derivations of triangular algebras. Linear Multilinear Algebra, 51(3):299–310, 2003. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheung:2003:MMS

[Che03b] Wai-Shun Cheung. Multiplicative maps with specified (1, 1)- entry. Linear Multilinear Algebra, 51(2):155–162, 2003. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2007:SFI

[Che07] Huanyin Chen. On simple factorization of invertible matrices. Linear Multilinear Algebra, 55(1):81–92, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2014:AMP

[Che14a] Hung-Yuan Chen. Additive maps preserving operatoreigenvector relations. Linear Multilinear Algebra, 62(6):803– 810, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheung:2014:BRN

[Che14b] Wai-Shun Cheung. On the boundary of rank-k numerical ranges. Linear Multilinear Algebra, 62(9):1266–1271, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheung:2016:BWN

[Che16] Wai-Shun Cheung. On the boundary of weighted numerical ranges. Linear Multilinear Algebra, 64(1):78–86, 2016. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheung:2017:NPM

[Che17] Wai-Shun Cheung. New pairs of matrices with convex generalized numerical ranges. Linear Multilinear Algebra, 65(10): 1962–1970, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2018:HTR

[Che18] Haiyan Chen. Hitting times for random walks on subdivision and triangulation graphs. Linear Multilinear Algebra, 66(1): 117–130, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:CHM

[Che22a] Lin Chen. On 6 × 6 complex Hadamard matrices containing two nonintersecting identical 3 × 3 submatrices. Linear Multilinear Algebra, 70(22):7856–7875, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:NSR

[Che22b] Xiaodan Chen. A note on the ABC spectral radius of graphs. Linear Multilinear Algebra, 70(5):775–786, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2019:SFO

[CHFC19] Jianzhang Chen, Yuanyuan Huang, Chunhui Feng, and Riqing Chen. Some families of optimal quantum codes derived from constacyclic codes. Linear Multilinear Algebra, 67(4):725–742, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Childs:1978:LIF

[Chi78] Lindsay N. Childs. Linearizing of n-ic forms and generalized Clifford algebras. Linear and Multilinear Algebra, 5(4):267– 278, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:1998:EGN

[Chi98] Mao-Ting Chien. The envelope of the generalized numerical range. Linear and Multilinear Algebra, 43(4):363–376, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2011:LTN

[Chi11] Mao-Ting Chien. Lucas’ theorem and numerical range. Linear Multilinear Algebra, 59(6):687–691, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chistopolskaya:2022:NGS

[Chi22] Alisa Chistopolskaya. On nilpotent generators of the symplectic Lie algebra. Linear Multilinear Algebra, 70(3):490–496, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choi:2000:GRP

[CHK00] Kunwook Choi, Don Hadwin, and Bok Ja Kim. General reflexivity and pattern subspaces. Linear and Multilinear Algebra, 47(4):259–280, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2018:RBR

[CHL18] Chen Chen, Jing Huang, and Shuchao Li. On the relation between the H-rank of a mixed graph and the matching number of its underlying graph. Linear Multilinear Algebra, 66(9): 1853–1869, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chollet:1982:PS

[Cho82] John Chollet. On principal submatrices. Linear and Multilinear Algebra, 11(3):283–285, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See remark [Mar82b].

Chooi:2010:RON

[Cho10] Wai Leong Chooi. Rank-one non-increasing additive maps between block triangular matrix algebras. Linear Multilinear Algebra, 58(6):715–740, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choi:2018:IAP

[Cho18a] Daeshik Choi. Inequalities about partial transpose and partial traces. Linear Multilinear Algebra, 66(8):1619–1625, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choi:2018:IRP

[Cho18b] Daeshik Choi. Inequalities related to partial trace and block Hadamard product. Linear Multilinear Algebra, 66(2):280– 284, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2018:SCH

[CHQ18] Xiaodan Chen, Yaoping Hou, and Jianguo Qian. Sufficient conditions for Hamiltonian graphs in terms of (signless Laplacian) spectral radius. Linear Multilinear Algebra, 66(5):919– 936, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:TPM

[CHS22] Haibo Chen, Yanyong Hong, and Yucai Su. Tensor product modules over the twisted Heisenberg–Virasoro algebra. Linear Multilinear Algebra, 70(19):4778–4791, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chu:1984:EMB

[Chu84a] Hsin Chu. On the exponential map of Borel subgroups in a semi-simple Lie group. Linear and Multilinear Algebra, 15(3– 4):285–312, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chui:1984:DAL

[Chu84b] Charles K. Chui. Design and analysis of linear predictorcorrector digital filters. Linear and Multilinear Algebra, 15 (1):47–69, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chu:1990:DMP

[Chu90] W. C. Chu. Dominance method for plane partitions. IV: Enumeration of flagged skew tableaux. Linear and Multilinear Algebra, 28(3):185–191, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chu:2022:MIS

[Chu22] Wenchang Chu. Multifold inverse series relations and multivariate derivative identities. Linear Multilinear Algebra, 70(4): 689–704, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:COG

[CHZ22] Saijie Chen, Qianglian Huang, and Lanping Zhu. A continuous orbit for the generalized inverse, Moore–Penrose inverse and group inverse. Linear Multilinear Algebra, 70(19):4049–4055, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic-Ilic:2009:ERD

[CICX09] Dragana S. Cvetković-Ilić, Jianbiao Chen, and Zhaoliang Xu. Explicit representations of the Drazin inverse of block matrix and modified matrix. Linear Multilinear Algebra, 57(4):355– 364, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic-Ilic:2016:VSR

[CID16] D. S. Cvetković-Ilić and M. Djikić. Various solutions to reverse order law problems. Linear Multilinear Algebra, 64(6):1207– 1219, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic-Ilic:2007:PRP

[CIDK07] D. Cvetković-Ilić, Alegra Dajić, and J. J. Koliha. Positive and real-positive solutions to the equation axa^{∗} = c in C^{∗}- algebras. Linear Multilinear Algebra, 55(6):535–543, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic-Ilic:2019:ROL

[CIM19] Dragana Cvetković-Ilić and Jovana Milosević. Reverse order laws for {1, 3}-generalized inverses. Linear Multilinear Alge-

Cvetkovic-Ilic:2015:ROL

[CIN15] Dragana S. Cvetković-Ilić and Jovana Nikolov. Reverse order laws for reflexive generalized inverse of operators. Linear Multilinear Algebra, 63(6):1167–1175, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic-Ilic:2017:PRS

[CIP17] Dragana S. Cvetković-Ilić and Vladimir Pavlović. On the point and residual spectrum of the operator matrix M_{X}. Linear Multilinear Algebra, 65(1):35–43, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic-Ilic:2018:DIU

[CIP18] D. S. Cvetković-Ilić and V. Pavlović. Drazin invertibility of upper triangular operator matrices. Linear Multilinear Algebra, 66(2):260–267, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cirulis:2015:LOR

[Cir15] J. Cirulis. Lattice operations on Rickart ∗-rings under the star order. Linear Multilinear Algebra, 63(3):497–508, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cirulis:2016:ESO

[Cir16] Janis Cirulis. Extending the star order to Rickart rings. Linear Multilinear Algebra, 64(8):1498–1508, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cirulis:2017:DPO

[Cir17] Janis Cirulis. The diamond partial order for strong Rickart rings. Linear Multilinear Algebra, 65(1):192–203, 2017. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic-Ilic:2021:ACS

[CIRW21] D. S. Cvetković-Ilić, J. Nikolov Radenković, and Qing-Wen Wang. Algebraic conditions for the solvability to some systems of matrix equations. Linear Multilinear Algebra, 69(9):1579– 1609, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carlson:1997:EPD

[CJ97] David Carlson and Charles R. Johnson. Extremal patterns of distinct entries in vectors in the range of a matrix. Linear and Multilinear Algebra, 43(1–3):283–297, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choi:2012:TIB

[CJ12] Youngwoo Choi and Jaewon Jung. Totally interpolating biorthogonal multiwavelet systems with approximation order. Linear Multilinear Algebra, 60(6):697–724, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:2020:CTT

[CJ20] Haoge Chang and C. R. Johnson. Completion of TP and TN border patterns. Linear Multilinear Algebra, 68(9):1753–1766, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Corey:2013:CPV

[CJK^{+}13] Dan Corey, Charles R. Johnson, Ryan Kirk, Brian Lins, and Ilya Spitkovsky. Continuity properties of vectors realizing points in the classical field of values. Linear Multilinear Algebra, 61(10):1329–1338, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cinzori:2022:EUB

[CJL22] Isaac Cinzori, Charles R. Johnson, and Hannah Lang. An explicit upper bound on disparity for trees of a given diameter. Linear Multilinear Algebra, 70(19):4584–4596, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cigler:2019:FAM

[CJW19] Grega Cigler, Marjan Jerman, and Piotr J. Wojciechowski. Full algebras of matrices. Linear Multilinear Algebra, 67(8): 1511–1521, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:EST

[CJY22] Xuemei Chen, Xian’an Jin, and Weigen Yan. Enumeration of spanning trees of middle digraphs. Linear Multilinear Algebra, 70(22):7484–7492, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Clancey:1984:BAR

[CK84] K. F. Clancey and B. A. Kon. The Bezoutian and the algebraic Riccati equation. Linear and Multilinear Algebra, 15(3–4): 265–278, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chebotar:2012:P

[CK12] Mikhail Chebotar and Wen-Fong Ke. Preface. Linear Multilinear Algebra, 60(7):761, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2017:SDP

[CK17] Lei Cao and Selcuk Koyuncu. Sub-defect of product of doubly substochastic matrices. Linear Multilinear Algebra, 65(4):653– 657, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chooi:2019:AMR

[CK19] Wai Leong Chooi and Kiam Heong Kwa. Additive maps of rank r tensors and symmetric tensors. Linear Multilinear Algebra, 67(6):1269–1293, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chooi:2020:CAC

[CK20] Wai Leong Chooi and KiamHeong Kwa. Classical adjoint commuting and determinant preserving linear maps on Kronecker products of Hermitian matrices. Linear Multilinear Algebra, 68(5):869–885, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Camara:2021:CSC

[CKGP21] M. Cristina Câmara, Kamila Kli’s-Garlicka, and Marek Ptak. Complex symmetric completions of partial operator matrices. Linear Multilinear Algebra, 69(8):1446–1467, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Camacho:2022:CEF

[CKKK22a] Luisa M. Camacho, Iqboljon Karimjanov, Ivan Kaygorodov, and Abror Khudoyberdiyev. Central extensions of filiform Zinbiel algebras. Linear Multilinear Algebra, 70(8):1479–1495, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Camacho:2022:OGN

[CKKK22b] Luisa M. Camacho, Iqboljon Karimjanov, Ivan Kaygorodov, and Abror Khudoyberdiyev. One-generated nilpotent Novikov algebras. Linear Multilinear Algebra, 70(2):331–365, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chun:2023:CSM

[CKKY23] Sangmin Chun, In Hyoun Kim, Jaewoong Kim, and Jasang Yoon. Contractive symmetric matrix completion problems related to graphs. Linear Multilinear Algebra, 71(10):1597–1605, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chooi:2016:NCA

[CKL16] Wai Leong Chooi, Kiam Heong Kwa, and Jinting Lau. A note on classical adjoint commuting linear maps of tensor products of matrices. Linear Multilinear Algebra, 64(4):745–766, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2021:HRM

[CKL^{+}21] Ningping Cao, David W. Kribs, Chi-Kwong Li, Mike I. Nelson, Yiu-Tung Poon, and Bei Zeng. Higher rank matricial ranges and hybrid quantum error correction. Linear Multilinear Algebra, 69(5):827–839, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chebotar:2013:LAA

[CKLP13] M. A. Chebotar, W.-F. Ke, P.-H. Lee, and E. R. Puczylowski. A linear algebra approach to Koethe’s problem and related questions. Linear Multilinear Algebra, 61(5):635–645, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheon:2018:SPP

[CKM18] Gi-Sang Cheon, Jang Soo Kim, and Seyed Ahmad Mojallal. Spectral properties of Pascal graphs. Linear Multilinear Algebra, 66(7):1403–1417, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheon:2022:SPM

[CKMY22] Gi-Sang Cheon, Jang Soo Kim, Seyed Ahmad Mojallal, and Meesue Yoo. Symmetric Pascal matrices and related graphs. Linear Multilinear Algebra, 70(21):6345–6357, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:1995:GGI

[CKN95] Younghong Chen, Stephen J. Kirkland, and Michael Neumann. Group generalized inverses of M-matrices associated with periodic and nonperiodic Jacobi matrices. Linear and Multilinear Algebra, 39(4):325–340, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chorianopoulos:2009:DNR

[CKP09] Ch. Chorianopoulos, S. Karanasios, and P. Psarrakos. A definition of numerical range of rectangular matrices. Linear Multilinear Algebra, 57(5):459–475, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2016:MCD

[CKP16] Lei Cao, Selcuk Koyuncu, and Timmothy Parmer. A minimal completion of doubly substochastic matrix. Linear Multilinear Algebra, 64(11):2313–2334, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Casas:2018:NAH

[CKR18] J. M. Casas, E. Khmaladze, and N. Pacheco Rego. A nonabelian Hom–Leibniz tensor product and applications. Linear Multilinear Algebra, 66(6):1133–1152, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Casas:2021:SPP

[CKR21] J. M. Casas, E. Khmaladze, and N. Pacheco Rego. On some properties preserved by the non-abelian tensor product of Hom–Lie algebras. Linear Multilinear Algebra, 69(4):607–626,

Calderon:2019:DLS

[CKS19] Antonio Jesús Calderón, Ivan Kaygorodov, and Paulo Saraiva. Decompositions of linear spaces induced by n-linear maps. Linear Multilinear Algebra, 67(6):1250–1268, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choi:2020:GMP

[CKS20] Hayoung Choi, Sejong Kim, and Yuanming Shi. Geometric mean of partial positive definite matrices with missing entries. Linear Multilinear Algebra, 68(12):2408–2433, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chooi:2020:CMR

[CKT20] Wai Leong Chooi, Kiam Heong Kwa, and Li Yin Tan. Commuting maps on rank k triangular matrices. Linear Multilinear Algebra, 68(5):1021–1030, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:2023:SEA

[CKXX23] Dong Cheng, Kit Ian Kou, Yonghui Xia, and Junfeng Xu. Sampling expansions associated with quaternion difference equations. Linear Multilinear Algebra, 71(13):2180–2203, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cross:1974:SRC

[CL74] G. W. Cross and P. Lancaster. Square roots of complex matrices. Linear and Multilinear Algebra, 1(4):289–293, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carlson:1980:IBM

[CL80] David Carlson and Raphael Loewy. Induced bilinear maps and matrix equations. Linear and Multilinear Algebra, 9(1):17–33, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1983:LTT

[CL83] G. H. Chan and M. M. Lim. Linear transformations on tensor spaces. Linear and Multilinear Algebra, 14(1):3–9, 1983.

Carlson:1985:EMI

[CL85] David Carlson and Raphael Loewy. On equality of maps induced by tensor and symmetric products. Linear and Multilinear Algebra, 17(2):87–100, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1986:NSCa

[CL86a] G. H. Chan and M. H. Lim. Nonzero symmetry classes of smallest dimension corresponding to intransitive groups. Linear and Multilinear Algebra, 19(1):1–10, 1986. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1986:NSCb

[CL86b] G. H. Chan and M. H. Lim. Nonzero symmetry classes of smallest dimension corresponding to transitive groups. Linear and Multilinear Algebra, 19(2):149–165, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chang:1991:SLO

[CL91] Stanley S. Chang and Chi-Kwong Li. A special linear operator on. Linear and Multilinear Algebra, 30(1–2):65–75, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1992:LTS

[CL92] G. H. Chan and M. H. Lim. Linear transformations on symmetric matrices II. Linear and Multilinear Algebra, 32(3–4): 319–325, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:1994:SGP

[CL94] Che-Man Cheng and Chi-Kwong Li. Some geometrical properties of the decomposable numerical range. Linear and Multilinear Algebra, 37(1–3):207–212, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1995:LPD

[CL95] G. H. Chan and M. H. Lim. Linear preservers of decomposable tensors. Linear and Multilinear Algebra, 40(2):171–181,

Choi:2000:NRD

[CL00] Man-Duen Choi and Chi-Kwong Li. Numerical ranges and dilations. Linear and Multilinear Algebra, 47(1):35–48, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chool:2001:RON

[CL01] W. L. Chool and M. H. Lim. Rank-one nonincreasing mappings on triangular matrices and some related preserver problems. Linear and Multilinear Algebra, 49(4):305–336, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chiang:2005:GDS

[CL05] Hanley Chiang and Chi-Kwong Li. Generalized doubly stochastic matrices and linear preservers. Linear Multilinear Algebra, 53(1):1–11, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheung:2013:EPS

[CL13] Wai-Shun Cheung and Chi-Kwong Li. Elementary proofs for some results on the circular symmetry of the numerical range. Linear Multilinear Algebra, 61(5):596–602, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheung:2014:MNR

[CL14] Wai-Shun Cheung and Chi-Kwong Li. Maximizing the numerical radii of matrices by permuting their entries. Linear Multilinear Algebra, 62(5):579–594, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2017:WNI

[CL17a] Chang-Pao Chen and Dah-Chin Luor. The weighted norm inequalities of integral operators with special types of kernels. Linear Multilinear Algebra, 65(9):1734–1742, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2017:MPH

[CL17b] Chaoqun Chen and Fangyan Lu. Maps that preserve higherdimensional numerical ranges with operator Jordan products.

Chen:2017:NLS

[CL17c] Haiyan Chen and Liwen Liao. The normalized Laplacian spectra of the corona and edge corona of two graphs. Linear Multilinear Algebra, 65(3):582–592, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chou:2021:PCA

[CL21] Ping-Han Chou and Cheng-Kai Liu. Power commuting additive maps on rank-k linear transformations. Linear Multilinear Algebra, 69(3):403–427, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chajda:2022:SCS

[CL22a] Ivan Chajda and Helmut Länger. Semimodules over commutative semirings and modules over unitary commutative rings. Linear Multilinear Algebra, 70(7):1329–1344, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:2022:CUN

[CL22b] Guo Cheng and Ji-Cheng Li. Class of Uzawa–NPHSS iteration method for solving nonsingular and singular saddle point problems. Linear Multilinear Algebra, 70(14):2706–2738, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Clausen:1982:SFB

[Cla82] Michael Clausen. A straightening formula for bipermanents. Linear and Multilinear Algebra, 11(1):33–38, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Clark:1992:BCP

[Cla92] W. Edwin Clark. Bounds on a class of partial partitions of a vector space over GF(2): a graph theoretical approach. Linear and Multilinear Algebra, 32(3–4):225–235, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2014:BCM

[CLL14] Chang-Pao Chen, Jin-Wen Lan, and Dah-Chin Luor. The best constants for multidimensional modular inequalities over spherical cones. Linear Multilinear Algebra, 62(5):683–713, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1993:RP

[CLLT93] Shih-Ping Chan, Mong-Lung Lang, Chong-Hai Lim, and Ser-Peow Tan. Research problem. Linear and Multilinear Algebra, 35(1):79–81, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Clark:2009:LPH

[CLMR09] Sean Clark, Chi-Kwong Li, Jennifer Mahle, and Leiba Rodman. Linear preservers of higher rank numerical ranges and radii. Linear Multilinear Algebra, 57(5):503–521, 2009. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chooi:2013:LSP

[CLN13] Wai Leong Chooi, Ming Huat Lim, and Zhen Chuan Ng. Linear spaces and preservers of symmetric matrices of bounded rank-two. Linear Multilinear Algebra, 61(8):1051–1062, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2021:DWH

[CLN21] Mao-Ting Chien, Chi-Kwong Li, and Hiroshi Nakazato. The diameter and width of higher rank numerical ranges. Linear Multilinear Algebra, 69(5):871–887, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2017:TMU

[CLNT17] Mao-Ting Chien, Jianzhen Liu, Hiroshi Nakazato, and Tin-Yau Tam. Toeplitz matrices are unitarily similar to symmetric matrices. Linear Multilinear Algebra, 65(10):2131–2144, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Casas:2013:CSL

[CLOK13] J. M. Casas, M. Ladra, B. A. Omirov, and I. A. Karimjanov. Classification of solvable Leibniz algebras with null-filiform nilradical. Linear Multilinear Algebra, 61(6):758–774, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:2022:CNR

[CLP22] J. T. Chan, C. K. Li, and Y. T. Poon. Closedness of the k-numerical range. Linear Multilinear Algebra, 70(11):2178– 2186, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choi:1981:CIS

[CLR81] M.-D. Choi, C. Laurie, and H. Radjavi. On commutators and invariant subspaces. Linear and Multilinear Algebra, 9(4):329– 340, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2015:BEP

[CLW15] Xiao-Shan Chen, Wen Li, and Xianping Wu. Backward errors of periodic invariant subspaces for regular periodic matrix pairs. Linear Multilinear Algebra, 63(4):852–861, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cui:2021:RCM

[CLW21] Lu-Bin Cui, Cui-Xia Li, and Shi-Liang Wu. The relaxation convergence of multisplitting AOR method for linear complementarity problem. Linear Multilinear Algebra, 69(1):40–47, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2003:PAG

[CLX03] Guoliang Chen, Guoming Liu, and Yifeng Xue. Perturbation analysis of the generalized Bott–Duffin inverse of L-zero matrices. Linear Multilinear Algebra, 51(1):11–20, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2022:EDT

[CLX22] Yanhua Cao, Shuchao Li, and Baogen Xu. Explicit determination of three invariants associated with random walks on nprism networks. Linear Multilinear Algebra, 70(10):1854–1870, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carlson:1984:GMI

[CM84] David Carlson and Eduardo Marques De Sá. Generalized minimax and interlacing theorems. Linear and Multilinear Alge-

Chondros:1985:DRM

[CM85] N. Chondros and J. Maroulas. Decomposition of a rational matrix into partial fractions. Linear and Multilinear Algebra, 18(3):225–248, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cain:1992:IHM

[CM92] Bryan E. Cain and Eduardo Marques De Sá. The inertia of Hermitian matrices with a prescribed 2 × 2 block decomposition. Linear and Multilinear Algebra, 31(1–4):119–130, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Corpuz:2007:SAZ

[CM07] L. Corpuz and J. J. McDonald. Spectrally arbitrary zerononzero patterns of order 4. Linear Multilinear Algebra, 55(3): 249–273, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chica:2015:RNI

[CM15] Mario Chica and Javier Merí. Rank-1 numerical index of some families of norms on the plane. Linear Multilinear Algebra, 63(9):1817–1828, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chabbabi:2019:NMC

[CM19] Fadil Chabbabi and Mostafa Mbekhta. Nonlinear maps commuting with the λ-Aluthge transform under (n, m)-Jordan-triple product. Linear Multilinear Algebra, 67(12):2382–2398, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic:2022:DIC

[CM22] Milos D. Cvetković and Dijana Mosić. Drazin invertibility, characterizations and structure of polynomially normal operators. Linear Multilinear Algebra, 70(20):4932–4945, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cordova-Martinez:2019:KJS

[CMDE19] Alejandra S. Córdova-Martínez, Abbas Darehgazani, and Alberto Elduque. On Kac’s Jordan superalgebra. Linear Multilinear Algebra, 67(3):490–498, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carmona:2017:ERK

[CMM17] Ángeles Carmona, Margarida Mitjana, and Enric Monsó. Effective resistances and Kirchhoff index in subdivision networks. Linear Multilinear Algebra, 65(9):1823–1837, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carmona:2020:GFP

[CMM20] Ángeles Carmona, Margarida Mitjana, and Enric Monsó. Green’s function in partial subdivision networks. Linear Multilinear Algebra, 68(1):94–112, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Clarke:2022:CPL

[CMM22] Oliver Clarke, Fatemeh Mohammadi, and Harshit J. Motwani. Conditional probabilities via line arrangements and point configurations. Linear Multilinear Algebra, 70(20):5268– 5300, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2022:CSM

[CMP22] Lei Cao, Darian McLaren, and Sarah Plosker. Centrosymmetric stochastic matrices. Linear Multilinear Algebra, 70(3): 449–464, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cadavid:2020:CTS

[CMR20] Paula Cadavid, Mary Luz Rodiño Montoya, and Pablo M. Rodriguez. Characterization theorems for the spaces of derivations of evolution algebras associated to graphs. Linear Multilinear Algebra, 68(7):1340–1354, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cadavid:2021:IBE

[CMR21] Paula Cadavid, Mary Luz Rodiño Montoya, and Pablo M. Rodriguez. On the isomorphisms between evolution algebras of graphs and random walks. Linear Multilinear Algebra, 69 (10):1858–1877, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2004:CFP

[CN04] Mao-Ting Chien and Hiroshi Nakazato. Complex formulation of Poncelet property in numerical range. Linear Multilinear Algebra, 52(3–4):159–175, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2005:NRN

[CN05] Mao-Ting Chien and Hiroshi Nakazato. The q-numerical ranges of normal operators. Linear Multilinear Algebra, 53(6): 393–416, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2007:BNR

[CN07] Mao-Ting Chien and Hiroshi Nakazato. The boundary of the q-numerical range of a reducible matrix. Linear Multilinear Algebra, 55(3):275–292, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2008:FPB

[CN08] Mao-Ting Chien and Hiroshi Nakazato. Flat portions on the boundary of the numerical ranges of certain Toeplitz matrices. Linear Multilinear Algebra, 56(1–2):143–162, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2014:NRN

[CN14] Mao-Ting Chien and Hiroshi Nakazato. Numerical range of a nilpotent matrix and related central force. Linear Multilinear Algebra, 62(5):595–613, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2015:EMI

[CN15] Mao-Ting Chien and Hiroshi Nakazato. Elliptic modular invariants and numerical ranges. Linear Multilinear Algebra, 63(8):1501–1519, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2021:CRM

[CN21] M. T. Chien and H. Nakazato. Computation of Riemann matrices of hyperbolic quartic curves. Linear Multilinear Alge-

Chien:2005:NRM

[CNP05] Mao-Ting Chien, Hiroshi Nakazato, and Panayiotis Psarrakos. On the q-numerical range of matrices and matrix polynomials. Linear Multilinear Algebra, 53(5):357–374, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2006:NRD

[CNP06] Mao-Ting Chien, Hiroshi Nakazato, and Panayiotis Psarrakos. The q-numerical range and the Davis–Wielandt shell of reducible 3 × 3 matrices. Linear Multilinear Algebra, 54(2): 79–112, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chien:2016:DPW

[CNUV16] Mao-Ting Chien, Hiroshi Nakazato, Batzorig Undrakh, and Adiyasuren Vandanjav. Determinantal polynomials of a weighted shift operator. Linear Multilinear Algebra, 64(1): 2–13, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2018:CID

[CNW18] Haixian Chen, Jizhu Nan, and Ying Wang. Coinvariants and invariants of the dihedral group in characteristic. Linear Multilinear Algebra, 66(2):224–242, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Campbell:2021:STF

[CNW21] Kevin Campbell, Minh Nguyen, and Rudi Weikard. On the spectral theory for first-order systems without the unique continuation property. Linear Multilinear Algebra, 69(12):2315– 2323, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Comon:2012:TRR

[CO12] Pierre Comon and Giorgio Ottaviani. On the typical rank of real binary forms. Linear Multilinear Algebra, 60(6):657–667, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Coan:1999:TEP

[Coa99] Boyd Coan. Top exterior powers over commutative rings. Linear and Multilinear Algebra, 46(3):249–258, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Coelho:1996:LPP

[Coe96] M. Purificação Coelho. Linear preserves of the permanent on symmetric matrices. Linear and Multilinear Algebra, 41(1): 1–8, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cohen:1988:MIM

[Coh88] Nir Cohen. 2 × 2 monic irreducible matrix polynomials. Linear and Multilinear Algebra, 23(4):325–331, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Camacho:2019:LAC

[COK19] L. M. Camacho, B. A. Omirov, and T. K. Kurbanbaev. Leibniz algebras constructed by Witt algebras. Linear Multilinear Algebra, 67(10):2048–2064, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Calderon:2020:CDR

[COK20] Antonio Jesús Calderón, Amir Fernández Ouaridi, and Ivan Kaygorodov. The classification of 2-dimensional rigid algebras. Linear Multilinear Algebra, 68(4):828–844, 2020. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Calderon:2022:CBM

[COK22] Antonio Jesús Calderón, Amir Fernández Ouaridi, and Ivan Kaygorodov. On the classification of bilinear maps with radical of a fixed codimension. Linear Multilinear Algebra, 70(18): 3553–3576, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Connell:1977:EDP

[Con77] I. Connell. An estimate for the dimension of the product of two vector spaces. Linear and Multilinear Algebra, 4(4):273–275, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Constantine:1990:GCL

[Con90] Gregory M. Constantine. Graph complexity and the Laplacian matrix in blocked experiments. Linear and Multilinear Algebra, 28(1–2):49–56, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Constantine:1994:BR

[Con94] Gregory M. Constantine. Book review. Linear and Multilinear Algebra, 36(3):231–232, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Costara:2014:SMM

[Cos14] Constantin Costara. Surjective maps on matrices preserving the local spectral radius distance. Linear Multilinear Algebra, 62(7):988–994, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Costara:2021:NCM

[Cos21] Constantin Costara. Nonlinear commuting maps on L(X). Linear Multilinear Algebra, 69(3):551–556, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Campbell:1981:CRS

[CP81] Stephen L. Campbell and George D. Poole. Convergent regular splittings for nonnegative matrices. Linear and Multilinear Algebra, 10(1):63–73, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carnicer:1997:BOM

[CP97] J. M. Carnicer and J. M. Pena. Bidiagonalization of oscillatory matrices. Linear and Multilinear Algebra, 42(4):365–376, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cooper:2010:ZFI

[CP10] Samuel Cooper and Stratos Prassidis. Zeta functions of infinite graph bundles. Linear Multilinear Algebra, 58(2):185–201, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chorianopoulos:2013:CBJ

[CP13] Christos Chorianopoulos and Panayiotis J. Psarrakos. On the continuity of Birkhoff–James-orthogonality sets. Linear Multilinear Algebra, 61(11):1447–1454, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chattopadhyay:2015:LSP

[CP15] Sriparna Chattopadhyay and Pratima Panigrahi. On Laplacian spectrum of power graphs of finite cyclic and dihedral groups. Linear Multilinear Algebra, 63(7):1345–1355, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2016:IPM

[CP16] Yingtong Chen and Jigen Peng. Influences of preconditioning on the mutual coherence and the restricted isometry property of Gaussian/Bernoulli measurement matrices. Linear Multilinear Algebra, 64(9):1750–1759, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cohen:2018:CRC

[CP18] Nir Cohen and Edgar Pereira. The cyclic rank completion problem with regular blocks. Linear Multilinear Algebra, 66(5):861–868, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cohen:2021:CRC

[CP21] Nir Cohen and Edgar Pereira. The cyclic rank completion problem with general blocks. Linear Multilinear Algebra, 69 (1):48–73, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chasiotis:2022:OCR

[CPKK22] Vasilis Chasiotis, Katerina Pericleous, Christos Koukouvinos, and Stratis Kounias. On optimality and construction of row designs under dependence for estimating means. Linear Multilinear Algebra, 70(4):714–729, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cui:2017:NIE

[CPLH17] Jingjing Cui, Guohua Peng, Quan Lu, and Zhengge Huang. The new improved estimates of the dominant degree and disc theorem for the Schur complement of matrices. Linear Multilinear Algebra, 65(7):1329–1348, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Campa:2022:LSB

[CPS22] L. Campa, R. Peniche, and G. Salgado. Lie superalgebras based on Heisenberg Lie algebras. Linear Multilinear Algebra, 70(8):1553–1562, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Clark:2015:STB

[CPT15] Sean Clark, Yung-Ning Peng, and S. Kuang Thamrongpairoj. Super tableaux and a branching rule for the general linear Lie superalgebra. Linear Multilinear Algebra, 63(2):274–282, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Culp:2017:UET

[CPZ17] J. Culp, K. Pearson, and T. Zhang. On the uniqueness of the Z_{1}-eigenvector of transition probability tensors. Linear Multilinear Algebra, 65(5):891–896, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Caldeira:2018:SOI

[CQ18] Cristina Caldeira and João Filipe Queiró. Sums of orbits of integral matrices. Linear Multilinear Algebra, 66(12):2421– 2429, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:MLD

[CQ22] Huimin Chen and Xiaofei Qi. Multiplicative Lie derivations on triangular n-matrix rings. Linear Multilinear Algebra, 70(7): 1230–1251, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2015:LIF

[CQH15] Alatancang Chen, Yaru Qi, and Junjie Huang. Left invertibility of formal Hamiltonian operators. Linear Multilinear Algebra, 63(2):235–243, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Che:2016:PBT

[CQW16] Maolin Che, Liqun Qi, and Yimin Wei. Perturbation bounds of tensor eigenvalue and singular value problems with even order. Linear Multilinear Algebra, 64(4):622–652, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Collings:1977:UBD

[CR77] Bruce J. Collings and Donald W. Robinson. Upper bounds on the dimension of the space of matrices transposed by a similarity transformation. Linear and Multilinear Algebra, 5(3): 159–173, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic:1990:LEG

[CR90] D. Cvetković and P. Rowlinson. The largest eigenvalue of a graph: A survey. Linear and Multilinear Algebra, 28(1–2): 3–33, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cagliero:2015:LBF

[CR15] Leandro Cagliero and Nadina Rojas. A lower bound for faithful representations of nilpotent Lie algebras. Linear Multilinear Algebra, 63(11):2135–2150, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Casas:2018:UCE

[CR18] J. M. Casas and N. Pacheco Rego. Universal α-central extensions of Hom–Leibniz n-algebras. Linear Multilinear Algebra, 66(12):2468–2486, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carmona:2022:SRE

[CR22a] Juan R. Carmona and Jonnathan Rodríguez. On the spectral radius and energy of digraphs. Linear Multilinear Algebra, 70 (19):4792–4803, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choque-Rivero:2022:TTR

[CR22b] A. E. Choque-Rivero. Three-term recurrence relation coefficients and continued fractions related to orthogonal matrix polynomials on the finite interval [a, b]. Linear Multilinear

Craigen:1991:ERM

[Cra91] R. Craigen. Embedding rectangular matrices in Hadamard matrices. Linear and Multilinear Algebra, 29(2):91–92, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cravo:2009:MPE

[Cra09] Glória Cravo. Matrices with prescribed eigenvalues and prescribed submatrices. Linear Multilinear Algebra, 57(7):737– 748, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Crouzeix:2008:FCB

[Cro08] Michel Crouzeix. A functional calculus based on the numerical range: applications. Linear Multilinear Algebra, 56(1–2):81– 103, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cronin:2018:CMS

[Cro18] Anthony G. Cronin. Constructive methods for spectra with three nonzero elements in the nonnegative inverse eigenvalue problem. Linear Multilinear Algebra, 66(3):435–446, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cameron:2021:RNR

[CRW21] T. R. Cameron, M. D. Robertson, and A. Wiedemann. On the restricted numerical range of the Laplacian matrix for digraphs. Linear Multilinear Algebra, 69(5):840–854, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Carbonara:1995:CRS

[CRY95] J. O. Carbonara, J. B. Remmel, and M. Yang. A combinatorial rule for the Schur function expansion of the Plethysm S(1^{a}, b)[P_{k}]. Linear and Multilinear Algebra, 39(4):341–373, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chollet:1979:BR

[CS79] John Chollet and J. K. Sengupta. Book reviews. Linear and Multilinear Algebra, 7(1):91–92, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Clark:1995:CCS

[CS95] W. Edwin Clark and Boris Shekhtman. Covering by complements of subspaces. Linear and Multilinear Algebra, 40 (1):1–13, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Changphas:2010:SIL

[CS10] Tawhat Changphas and R. P. Sullivan. Skew idempotents in linear transformation semigroups. Linear Multilinear Algebra, 58(4):399–411, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2018:DMR

[CS18] Huanyin Chen and Marjan Sheibani. Diagonal matrix reduction over Hermite rings. Linear Multilinear Algebra, 66(4): 681–691, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2019:UIS

[CS19] Xiao Shan Chen and Hai-Wei Sun. On the unsolvability of inverse singular value problems almost everywhere. Linear Multilinear Algebra, 67(5):987–994, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2021:HIB

[CS21a] Huanyin Chen and Marjan Sheibani. The g-Hirano inverse in Banach algebras. Linear Multilinear Algebra, 69(7):1352–1362, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Choudhury:2021:MPS

[CS21b] Projesh Nath Choudhury and K. C. Sivakumar. Matrices with positive semidefinite real part. Linear Multilinear Algebra, 69(3):448–470, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:DIS

[CS22] Huanyin Chen and Marjan Sheibani. The g-Drazin inverse of the sum in Banach algebras. Linear Multilinear Algebra, 70 (1):53–65, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chatterjee:2023:HPS

[CS23] Manami Chatterjee and K. C. Sivakumar. On the Hadamard product AA, for a singular M-matrix A. Linear Multilinear Algebra, 71(12):1922–1932, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic:2001:VNS

[CSCH01] D. Cvetković, S. Simić, G. Caporossi, and P. Hansen. Variable neighborhood search for extremal graphs 3. Linear and Multilinear Algebra, 49(2):143–160, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Campoamor-Stursberg:2022:ACS

[CSG22] R. Campoamor-Stursberg and F. Oviaño García. Algorithmic construction of solvable rigid Lie algebras determined by generating functions. Linear Multilinear Algebra, 70(2):280–296, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chattopadhyay:2021:CSP

[CSS21] Arup Chattopadhyay, Debmalya Sain, and Tanusri Senapati. Characterization of symmetric points in l_{p}^{n}-spaces. Linear Multilinear Algebra, 69(16):2998–3009, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cain:1975:BR

[CST75] B. E. Cain, Hans Schneider, and R. C. Thompson. Book reviews. Linear and Multilinear Algebra, 2(4):365–366, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chi:2018:BLC

[CSY18] Lili Chi, Jiancai Sun, and Hengyun Yang. Biderivations and linear commuting maps on the planar Galilean conformal algebra. Linear Multilinear Algebra, 66(8):1606–1618, 2018. CO-

Cheung:1996:NRM

[CT96] Wai-Shun Cheung and Nam-Kiu Tsing. The C-numerical range of matrices is star-shaped. Linear and Multilinear Algebra, 41(3):245–250, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheung:2009:PDN

[CT09] Wai-Shun Cheung and Tin-Yau Tam. Pfaffian and decomposable numerical range of a complex skew symmetric matrix. Linear Multilinear Algebra, 57(5):477–489, 2009. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2016:SRO

[CT16] Li Chen and Fenglei Tian. Skew-rank of an oriented graph with edge-disjoint cycles. Linear Multilinear Algebra, 64(6): 1197–1206, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cardoso:2021:PEG

[CT21] Kauê Cardoso and Vilmar Trevisan. Principal eigenvectors of general hypergraphs. Linear Multilinear Algebra, 69(14): 2641–2656, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2022:SSC

[CT22] Wensheng Cao and Zhe Tang. Some similarities classes related to O(2, 1) in split quaternions. Linear Multilinear Algebra, 70 (19):4244–4265, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:2015:NEP

[CTW15] Guang-Hui Cheng, Qin Tan, and Zhuan-De Wang. A note on eigenvalues of perturbed 2 × 2 block Hermitian matrices. Linear Multilinear Algebra, 63(4):820–825, 2015. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Contreras:2020:GGL

[CTY20] Ivan Contreras, Michael Toriyama, and Chengzheng Yu. Gluing of graph Laplacians and their spectra. Linear Multilinear

Carini:1994:PRL

[CV94] Luisa Carini and Drensky Vesselin. Plethysms for representations of Lie superalgebras with applications to P.I. algebras. Linear and Multilinear Algebra, 37(4):239–258, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic:1995:SPG

[Cve95] Dragos Cvetković. Star partitions and the graph isomorphism problem. Linear and Multilinear Algebra, 39(1–2):109–132, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cvetkovic:2018:GDI

[Cve18] Milos D. Cvetković. Generalized Drazin invertibility of operator matrices. Linear Multilinear Algebra, 66(4):692–703, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Canfield:1981:HLP

[CW81] E. R. Canfield and S. G. Williamson. Hook length products and Cayley operators of classical invariant theory. Linear and Multilinear Algebra, 9(4):289–297, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Canfield:1989:CSS

[CW89] E. R. Canfield and S. G. Williamson. A combinatorial study of straightening bases. Linear and Multilinear Algebra, 24(4): 241–264, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Canfield:1990:TBL

[CW90] E. R. Canfield and S. G. Williamson. The two basic linear time planarity algorithms: Are they the same? Linear and Multilinear Algebra, 26(4):243–265, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Canfield:1991:SSN

[CW91] E. Rodney Canfield and S. Gill Williamson. A sequential sorting network analogous to the Batcher merge. Linear and Mul-

Compton:1993:DAG

[CW93] Robert C. Compton and S. Gill Williamson. Doubly adjacent gray codes for the symmetric group. Linear and Multilinear Algebra, 35(3–4):237–293, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2010:LBC

[CW10] Chang-Pao Chen and Kuo-Zhong Wang. Lower bounds of Copson type for weighted mean matrices and nörlund matrices. Linear Multilinear Algebra, 58(3):343–353, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2011:ONL

[CW11a] Chang-Pao Chen and Kuo-Zhong Wang. Operator norms and lower bounds of generalized Hausdorff matrices. Linear Multilinear Algebra, 59(3):321–337, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2011:NMS

[CW11b] Zhengxin Chen and Dengyin Wang. Nonlinear maps satisfying derivability on standard parabolic subalgebras of finitedimensional simple Lie algebras. Linear Multilinear Algebra, 59(3):261–270, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2012:NMS

[CW12] Zhengxin Chen and Dengyin Wang. Nonlinear maps satisfying derivability on the parabolic subalgebras of the general linear Lie algebras. Linear Multilinear Algebra, 60(2):149–157, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2016:NBM

[CW16] Li Chen and Dengyin Wang. Nonlinear bijective maps on strictly upper triangular matrices preserving zero product. Linear Multilinear Algebra, 64(3):412–425, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2017:NRA

[CW17] Yunkun Chen and Yimin Wei. Numerical radius for the asymptotic stability of delay differential equations. Linear Multilinear Algebra, 65(11):2306–2315, 2017. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chae:2020:EBI

[CW20] Minwoo Chae and Stephen G. Walker. An EM-based iterative method for solving large sparse linear systems. Linear Multilinear Algebra, 68(1):45–62, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chooi:2022:CAM

[CW22] Wai Leong Chooi and Jian Yong Wong. Commuting additive maps on tensor products of matrices. Linear Multilinear Algebra, 70(20):5580–5605, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2023:SSB

[CW23] Zhengxin Chen and Yu Wang. Skew-symmetric biderivations and linear commuting maps of Kac–Moody algebras. Linear Multilinear Algebra, 71(5):867–874, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cheng:2017:BLC

[CWSZ17] Xiao Cheng, Minjing Wang, Jiancai Sun, and Honglian Zhang. Biderivations and linear commuting maps on the Lie algebra. Linear Multilinear Algebra, 65(12):2483–2493, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2012:NLT

[CX12] Zhengxin Chen and Zhankui Xiao. Nonlinear Lie triple derivations on parabolic subalgebras of finite-dimensional simple Lie algebras. Linear Multilinear Algebra, 60(6):645–656, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cao:2023:STI

[CX23] Zhengbang Cao and Pengpeng Xie. On some tensor inequalities based on the t-product. Linear Multilinear Algebra, 71(3): 377–390, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:ITP

[CY22a] Qiu-Fan Chen and Yu-Feng Yao. Irreducible tensor product modules over the affine-Virasoro algebra of type a_{1}. Linear Multilinear Algebra, 70(22):7558–7574, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:BSC

[CY22b] Zhengxin Chen and Yalong Yu. Biderivations and strong commutativity-preserving maps on parabolic subalgebras of simple Lie algebras. Linear Multilinear Algebra, 70(14):2659– 2671, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2021:SLS

[CZ21] Ming-Zhu Chen and Xiao-Dong Zhang. On the signless Laplacian spectral radius of K_{s,t}-minor free graphs. Linear Multilinear Algebra, 69(10):1922–1934, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Cossu:2022:IFS

[CZ22] Laura Cossu and Paolo Zanardo. Idempotent factorizations of singular 2 × 2 matrices over quadratic integer rings. Linear Multilinear Algebra, 70(2):297–309, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2023:CAA

[CZ23] Yin Chen and Runxuan Zhang. A commutative algebra approach to multiplicative Hom–Lie algebras. Linear Multilinear Algebra, 71(7):1127–1144, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2023:FCC

[CZWF23] Yang Chen, Kezheng Zuo, Qingwen Wang, and Zhimei Fu. Further characterizations of the CMP inverse of matrices. Linear Multilinear Algebra, 71(1):100–119, 2023. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2012:MFD

[CZY12] Yizhi Chen, Xianzhong Zhao, and Lin Yang. On n × n matrices over a finite distributive lattice. Linear Multilinear Algebra, 60(2):131–147, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:LDW

[CZZ22] Yang Chen, Kaiming Zhao, and Yueqiang Zhao. Local derivations on Witt algebras. Linear Multilinear Algebra, 70(6): 1159–1172, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chen:2022:RFG

[CZZH22] Saijie Chen, Yayuan Zhao, Lanping Zhu, and Qianglian Huang. Regular factorizations and group inverses of linear operators in Banach spaces. Linear Multilinear Algebra, 70(7): 1252–1270, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DaPurificacaoCoelho:1988:NSF

[Da 88] M. A. Da Purificação Coelho. A note on spherical functioninduced subgroups. Linear and Multilinear Algebra, 23(3): 227–231, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DaProvidencia:1994:NRD

[Da 94] João Da Providência. The numerical ranges of derivations and quantum physics. Linear and Multilinear Algebra, 37(1– 3):213–220, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DaFonseca:2005:NME

[Da 05] C. M. Da Fonseca. A note on the multiplicities of the eigenvalues of a graph. Linear Multilinear Algebra, 53(4):303–307, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2022:SSL

[DA22] Ranjan Kumar Das and Rafikul Alam. Structured strong linearizations of structured rational matrices. Linear Multilinear Algebra, 70(20):6018–6051, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bezerra:2017:NPT

[dAB17] Johanns de Andrade Bezerra. A note on the product of two matrices of index one. Linear Multilinear Algebra, 65(7):1479– 1492, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Bezerra:2021:NCU

[dAB21] Johanns de Andrade Bezerra. A note on completion to the unitary matrices. Linear Multilinear Algebra, 69(10):1825– 1840, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dagli:2023:NAC

[Dag23a] Muhammet Cihat Dagli. A new almost convergent sequence space defined by Schröder matrix. Linear Multilinear Algebra, 71(11):1863–1874, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dagli:2023:NCM

[Dag23b] Muhammet Cihat Dagli. A novel conservative matrix arising from Schröder numbers and its properties. Linear Multilinear Algebra, 71(8):1338–1351, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2019:DLD

[DAH19] Kinkar Ch. Das, Mustapha Aouchiche, and Pierre Hansen. On distance Laplacian and distance signless Laplacian eigenvalues of graphs. Linear Multilinear Algebra, 67(11):2307–2324, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2004:LTL

[Das04] Kinkar Ch. Das. The largest two Laplacian eigenvalues of a graph. Linear Multilinear Algebra, 52(6):441–460, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2017:NZC

[Das17] Angsuman Das. Non-zero component union graph of a finitedimensional vector space. Linear Multilinear Algebra, 65(6): 1276–1287, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2018:SIG

[Das18] Angsuman Das. On subspace inclusion graph of a vector space. Linear Multilinear Algebra, 66(3):554–564, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2021:NGT

[Das21] Kinkar Chandra Das. Nordhaus–Gaddum-type result on the second largest signless Laplacian eigenvalue of a graph. Linear Multilinear Algebra, 69(6):1035–1044, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Datta:1980:MEM

[Dat80] B. N. Datta. Matrix equation, matrix polynomial and the number of zeros of a polynomial inside the unit circle. Linear and Multilinear Algebra, 9(1):63–68, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Datta:2016:WBC

[Dat16] Somantika Datta. Welch bounds for cross correlation of subspaces and generalizations. Linear Multilinear Algebra, 64(8): 1484–1497, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Day:1997:RCT

[Day97] Jane M. Day. Robert C. Thompson, 1931–1995. Linear and Multilinear Algebra, 43(1–3):13–18, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deBruyn:2010:GMU

[dB10] Bart de Bruyn. On the Grassmann modules for the unitary groups. Linear Multilinear Algebra, 58(7):887–902, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deBondt:2014:SNI

[dB14a] Michiel de Bondt. The strong nilpotency index of a matrix. Linear Multilinear Algebra, 62(4):486–497, 2014. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dolzan:2014:MMC

[DB14b] David Dolzan and Damjana Kokol Bukovsek. Minimal matrix centralizers over the field Z_{2}. Linear Multilinear Algebra, 62(8):1114–1126, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deBondt:2016:MSC

[dB16] Michiel de Bondt. Mathieu subspaces of codimension less than n of Mat_{n}(K). Linear Multilinear Algebra, 64(10):2049–2067, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2018:LBL

[DBG18] Chengyu Du, Chengming Bai, and Li Guo. 3-Lie bialgebras and 3-Lie classical Yang–Baxter equations in low dimensions. Linear Multilinear Algebra, 66(8):1633–1658, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

diBenedetto:2000:OMM

[dBS00] Fabio di Benedetto and Stefano Serra Capizzano. Optimal multilevel matrix algebra operators. Linear and Multilinear Algebra, 48(1):35–66, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deCamargo:2013:GME

[dC13] André Pierro de Camargo. A generalization of Mühlbach’s extension principle for determinantal identities. Linear Multilinear Algebra, 61(10):1363–1376, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deCamargo:2015:SDI

[dC15] André Pierro de Camargo. On some determinantal identities formation laws. Linear Multilinear Algebra, 63(9):1760–1767, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deCamargo:2018:SDI

[dC18] André Pierro de Camargo. On some determinantal identities formation laws. Linear Multilinear Algebra, 66(5):1067, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dong:2020:CPB

[DC20] Jiong Dong and Xiaohong Cao. Compact perturbations of both SVEP and Weyl’s theorem for 3 × 3 upper triangular operator matrices. Linear Multilinear Algebra, 68(10):2020– 2033, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dong:2023:CUT

[DC23] Jiong Dong and Xiaohong Cao. CFI upper triangular operator matrices. Linear Multilinear Algebra, 71(7):1217–1227, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deCamargo:2022:RML

[dCdSA22] André Pierro de Camargo and Felipe da Silva Alves. Replacing minors by linear functions and products in the Cauchy– Sylvester compound and adjugate determinants. Linear Multilinear Algebra, 70(5):909–918, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deCamargo:2021:NLD

[dCH21] André Pierro de Camargo and Gabriel Haeser. A note on linearly dependent symmetric matrices. Linear Multilinear Algebra, 69(13):2539–2545, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2010:SRG

[DCIW10] Chunyuan Deng, Dragana S. Cvetković-Ilić, and Yimin Wei. Some results on the generalized Drazin inverse of operator matrices. Linear Multilinear Algebra, 58(4):503–521, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duan:2019:IAS

[DCLW19] Xuefeng Duan, Jianheng Chen, Chunmei Li, and Qingwen Wang. Iterative algorithm for the symmetric and nonnegative tensor completion problem. Linear Multilinear Algebra, 67(8):1579–1595, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2010:RMP

[DD10] Chun Yuan Deng and Hong Ke Du. Representations of the Moore–Penrose inverse for a class of 2-by-2 block operator valued partial matrices. Linear Multilinear Algebra, 58(1): 15–26, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Djordjevic:2012:DID

[DD12] S. V. Djordjević and B. P. Duggal. Drazin invertibility of the diagonal of an operator. Linear Multilinear Algebra, 60 (1):65–71, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dincic:2014:HTR

[DD14] Nebojsa C. Dincić and Dragan S. Djordjević. Hartwig’s triple reverse order law revisited. Linear Multilinear Algebra, 62(7): 918–924, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeFilippis:2015:GSD

[DD15] Vincenzo De Filippis and Onofrio Mario Di Vincenzo. Generalized skew derivations on semiprime rings. Linear Multilinear Algebra, 63(5):927–939, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Demir:2020:QGD

[DDA20]  Çagri Demir, Vincenzo De Filippis, and Nurcan Argaç. A

quadratic generalized differential identity on Lie ideals of Prime Rings. Linear Multilinear Algebra, 68(9):1835–1847, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2015:SAM

[DdF15] Zhibin Du and Carlos M. da Fonseca. The singular acyclic matrices with the second largest number of P -vertices. Linear Multilinear Algebra, 63(10):2103–2120, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dinh:2022:MHM

[DDF22a] Trung Hoa Dinh, Raluca Dumitru, and Jose A. Franco. On the matrix Heron means and Rényi divergences. Linear Multilinear Algebra, 70(8):1442–1450, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2022:PDP

[DdF22b] Zhibin Du and Carlos M. da Fonseca. A periodic determinantal property for (0, 1) double banded matrices. Linear Multilinear Algebra, 70(20):5316–5328, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DePillis:1975:EDT

[De 75] John De Pillis. An extended duality theorem and multiplication of several matrices. Linear and Multilinear Algebra, 3(1–2):161–168, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DePillis:1978:CSD

[De 78] John De Pillis. Characterizations of sums of dyads and of Kronecker products. Linear and Multilinear Algebra, 6(1):23– 31, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeOliveira:1982:MPS

[De 82a] G. N. De Oliveira. Matrices with prescribed submatrices. Linear and Multilinear Algebra, 12(2):125–138, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DePillis:1982:SAR

[De 82b] John De Pillis. Stand-alone results and applications in multilinear algebra. Linear and Multilinear Algebra, 12(2):139–151, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeOliveira:1984:RP

[De 84] Graciano De Oliveira. Research problems. Linear and Multilinear Algebra, 16(1–4):261–262, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeOliveira:1990:IPS

[De 90] G. N. De Oliveira. Invariant polynomials of sums of matrices. Linear and Multilinear Algebra, 27(4):285–288, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeBruyn:2008:GEH

[De 08] Bart De Bruyn. On the Grassmann-embeddings of the Hermitian dual polar spaces. Linear Multilinear Algebra, 56(6): 665–677, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeTeran:2013:SE

[De 13] Fernando De Terán. The solution of the equation AX+BX^{∗} = 0. Linear Multilinear Algebra, 61(12):1605–1628, 2013. CO-

Decell:1974:DGI

[Dec74] Henry P. Decell Jr. On the derivative of the generalized inverse of a matrix. Linear and Multilinear Algebra, 1(4):357–359, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Decaen:1995:TMS

[Dec95] D. Decaen. On the tournament matrices with smallest rank. Linear and Multilinear Algebra, 40(1):45, 1995. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Delai:1994:ROL

[Del94] M. Benlarbi Delaï. Sur la réflexivité des opérateurs linéaires. (French) [On the reflexivity of linear operators]. Linear and Multilinear Algebra, 38(1–2):39–43, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dengyin:2005:CPS

[Den05] Wang Dengyin. Completeness of the parabolic subgroups of projective symplectic groups. Linear Multilinear Algebra, 53(5):375–386, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2014:ISL

[Den14] Chunyuan Deng. Invertibility and stability of linear combinations of m-potent operators. Linear Multilinear Algebra, 62 (10):1365–1376, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deretsky:1983:SSN

[Der83] Z. Deretsky. On the symmetry of the Smith normal form for (υ, k, λ) designs. Linear and Multilinear Algebra, 14(2): 187–193, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Derksen:2016:MCT

[Der16] Harm Derksen. Matrix completion and tensor rank. Linear Multilinear Algebra, 64(4):680–685, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Derksen:2018:EBL

[Der18] Harm Derksen. On the equivalence between low-rank matrix completion and tensor rank. Linear Multilinear Algebra, 66(4): 645–667, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deutsch:1980:VER

[Deu80] Emeric Deutsch. The variation of the eigenvalues of a real symmetric tridiagonal matrix upon the variation of some of its elements. Linear and Multilinear Algebra, 9(1):51–62, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeVille:2022:GDS

[DeV22] Lee DeVille. The generalized distance spectrum of a graph and applications. Linear Multilinear Algebra, 70(13):2425–2458, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dicks:1982:PSP

[DF82] Warren Dicks and Edward Formanek. Poincaré series and a problem of S. Montgomery. Linear and Multilinear Algebra, 12 (1):21–30, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiasDaSilva:1987:MPI

[DF87] J. A. Dias Da Silva and Amélia Fonseca. On the multilinearity partition of an irreducible character. Linear and Multilinear Algebra, 20(3):203–218, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiasDaSilva:1990:NSP

[DF90] J. A. Dias Da Silva and Amélia Fonseca. Nonzero star products. Linear and Multilinear Algebra, 27(1):49–55, 1990. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiasDaSilva:1995:SCT

[DF95] J. A. Dias Da Silva and Amélia Fonseca. Symmetry classes of tensors:dual indices. Linear and Multilinear Algebra, 39(3): 245–253, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

daFonseca:2005:CAP

[dF05] C. M. da Fonseca. On a conjecture about the μ-permanent. Linear Multilinear Algebra, 53(3):225–230, 2005. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

daFonseca:2011:SOQ

[dF11] C. M. da Fonseca. Some open questions on the Fibonacci polytope. Linear Multilinear Algebra, 59(3):317–320, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

daFonseca:2013:ROP

[dF13] C. M. da Fonseca. Remarks and open problems on the minimum permanent of acyclic doubly stochastic matrices. Linear Multilinear Algebra, 61(3):281–293, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

daFonseca:2019:PR

[dF19] Carlos M. da Fonseca. The μ-permanent revisited. Linear Multilinear Algebra, 67(8):1713–1714, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dalfo:2019:SET

[DFK19] C. Dalfó, M. A. Fiol, and J. Koolen. The spectral excess theorem for graphs with few eigenvalues whose distance-2 or distance-1-or-2 graph is strongly regular. Linear Multilinear Algebra, 67(12):2373–2381, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

daFonseca:2021:NTS

[dFK21] Carlos M. da Fonseca and Emrah Kiliç. A new type of Sylvester–Kac matrix and its spectrum. Linear Multilinear Algebra, 69(6):1072–1082, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dalfo:2023:SEU

[DFPS23] C. Dalfó, M. A. Fiol, S. Pavlíková, and J. Sirán. On the spectra and eigenspaces of the universal adjacency matrices of arbitrary lifts of graphs. Linear Multilinear Algebra, 71(5): 693–710, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

daFonseca:2022:DII

[dFSY22] Lais S. da Fonseca, Ednei A. Santulo, Jr., and Felipe Y. Yasumura. Degree-inverting involution on full square and triangular matrices. Linear Multilinear Algebra, 70(10):1980–1994, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dorsey:2016:LPS

[DGJ^{+}16] Jonathan Dorsey, Tom Gannon, Nathaniel Jacobson, Charles R. Johnson, and Morrison Turnansky. Linear preservers of semipositive matrices. Linear Multilinear Algebra, 64(9):1853– 1862, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Danchev:2022:DMD

[DGL22] Peter Danchev, Esther García, and Miguel Gómez Lozano. Decompositions of matrices into diagonalizable and squarezero matrices. Linear Multilinear Algebra, 70(19):4056–4070, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeLaPena:2008:RRG

[DGR08] José A. De La Peña, Ivan Gutman, and Juan Rada. Retarded regular graphs are regular or semiregular. Linear Multilinear Algebra, 56(5):507–509, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Day:1986:ISS

[DH86] Jane M. Day and Melvin Henriksen. Iterative solutions of systems of linear equations whose coefficient matrix is positive real. Linear and Multilinear Algebra, 19(3):267–285, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Do:1991:SRL

[DH91] Seok-Su Do and Suk-Geun Hwang. Some rationally looking faces of Ω_{n} having irrational minimum permanents. Linear and Multilinear Algebra, 30(1–2):145–154, 1991. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2003:PDS

[DH03] Shuanping Du and Jinchuan Hou. Positive definite solutions of operator equations Xm + A ∗ X − nA = I. Linear Multilinear Algebra, 51(2):163–173, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dehghan:2011:GBS

[DH11] Mehdi Dehghan and Masoud Hajarian. On the generalized bisymmetric and skew-symmetric solutions of the system of generalized Sylvester matrix equations. Linear Multilinear Algebra, 59(11):1281–1309, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Draisma:2016:ANC

[DH16] Jan Draisma and Emil Horobet. The average number of critical rank-one approximations to a tensor. Linear Multilinear Algebra, 64(12):2498–2518, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2017:SNF

[DH17] Anupam Das and Bipan Hazarika. Some new Fibonacci difference spaces of non-absolute type and compact operators. Linear Multilinear Algebra, 65(12):2551–2573, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dai:2018:LVL

[DH18] Xiansheng Dai and Jianzhi Han. Loop Virasoro Lie conformal superalgebra. Linear Multilinear Algebra, 66(1):131–146, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Djagba:2020:SSF

[DH20] P. Djagba and K.-T. Howell. The subspace structure of finite dimensional Beidleman near-vector spaces. Linear Multilinear Algebra, 68(11):2316–2336, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2022:SPU

[DH22] Yanyan Du and Junjie Huang. Spectral property of upper triangular relation matrices. Linear Multilinear Algebra, 70(8): 1526–1542, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2007:AMP

[DHB07] Shuanping Du, Jinchuan Hou, and Zhaofang Bai. Additive maps preserving similarity or asymptotic similarity on. Linear Multilinear Algebra, 55(3):209–218, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2022:RUT

[DHH22] Yanyan Du, Junjie Huang, and Ran Huo. On the range of upper triangular relation matrices. Linear Multilinear Algebra, 70(20):5750–5769, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Darweesh:2023:NIP

[DHHK23] Amer Darweesh, Mostafa Hayajneh, Saja Hayajneh, and Fuad Kittaneh. Norm inequalities for positive definite matrices related to a question of Bourin. Linear Multilinear Algebra, 71 (12):1933–1947, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dmitrishin:2016:SCD

[DHKS16] D. Dmitrishin, P. Hagelstein, A. Khamitova, and A. Stokolos. On the stability of cycles by delayed feedback control. Linear Multilinear Algebra, 64(8):1538–1549, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dirr:2008:RBN

[DHKSH08] G. Dirr, U. Helmke, M. Kleinsteuber, and Th. Schulte-Herbrüggen. Relative C-numerical ranges for applications in quantum control and quantum information. Linear Multilinear Algebra, 56(1–2):27–51, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiasDaSilva:1995:EMP

[DHS95] J. A. Dias Da Silva, Daniel Hershkowitz, and Hans Schneider. Existence of matrices with prescribed off-diagonal block element sums. Linear and Multilinear Algebra, 40(1):15–28, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2023:REF

[dHS23] Xiang dong Hou and Christopher Sze. On a radical extension of the field of rational functions in several variables. Linear

Dey:1994:SPS

[DHT94] Aloke Dey, Sayaji Hande, and M. L. Tiku. Statistical proofs of some matrix results. Linear and Multilinear Algebra, 38(1–2): 109–116, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiVincenzo:1988:KPC

[Di 88] O. M. Di Vincenzo. On the Kronecker product of S_{n}- cocharacters for P.I. algebras. Linear and Multilinear Algebra, 23(2):139–143, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiasDaSilva:1979:SI

[Dia79] J. A. Dias Da Silva. On the Schur inequality. Linear and Multilinear Algebra, 7(4):343–357, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiasDaSilva:1990:CM

[Dia90a] J. A. Dias Da Silva. On the μ-colorings of a matroid. Linear and Multilinear Algebra, 27(1):25–32, 1990. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiasDaSilva:1990:MPA

[Dia90b] J. A. Dias Da Silva. On the minimal polynomial of the associated transformation. Linear and Multilinear Algebra, 27(4): 275–283, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dick:1984:DCP

[Dic84] Wayne E. Dick. Discrete characterizations of planarity part i: the classical viewpoint. Linear and Multilinear Algebra, 16(1– 4):39–65, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dieudonne:1979:TG

[Die79] J. Dieudonné. The tragedy of Grassmann. Linear and Multilinear Algebra, 8(1):1–14, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2013:FRD

[DIF13] Zhibin Du, Aleksandar Ilić, and Lihua Feng. Further results on the distance spectral radius of graphs. Linear Multilinear Algebra, 61(9):1287–1301, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ding:2023:MNM

[DIJ23] Wenxuan Ding, Matthew Ingwersen, and Charles R. Johnson. The minimum number of multiplicity 1 eigenvalues among real symmetric matrices whose graph is a 2-linear tree. Linear Multilinear Algebra, 71(6):945–960, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ding:1998:PLS

[Din98] Jiu Ding. On the perturbation of least squares problems with equality constraints. Linear and Multilinear Algebra, 45(1): 35–47, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ding:2000:PSL

[Din00] Jiu Ding. Perturbation of systems of linear algebraic equations. Linear and Multilinear Algebra, 47(2):119–127, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ding:2015:SSP

[Din15a] Xue Ding. On some spectral properties of large self-dual dilute quaternion random matrices. Linear Multilinear Algebra, 63(9):1737–1749, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Diniz:2015:GIT

[Din15b] Diogo Diniz. 2-graded identities for the tensor square of the Grassmann algebra. Linear Multilinear Algebra, 63(4):702– 712, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dixon:1979:ICM

[Dix79] John D. Dixon. An isomorphism criterion for modules over a principal ideal domain. Linear and Multilinear Algebra, 8 (1):69–72, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1976:SRM

[– DDJ76] D. Ž. – Doković, C. R. De Prima, and C. R. Johnson. On some representations of matrics. Linear and Multilinear Algebra, 4 (1):33–40, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drew:1989:CCM

[DJ89a] John H. Drew and Charles R. Johnson. Counterexample to a conjecture of Mehta regarding permanental maximization. Linear and Multilinear Algebra, 25(3):253–254, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drew:1989:MPP

[DJ89b] John H. Drew and Charles R. Johnson. The maximum permanent of a 3-by-3 positive semidefinite matrix, given the eigenvalues. Linear and Multilinear Algebra, 25(3):243–251, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See comments [Hu94].

Drew:1998:CPD

[DJ98] John H. Drew and Charles R. Johnson. The completely positive and doubly nonnegative completion problems. Linear and Multilinear Algebra, 44(1):85–92, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ji:2021:SME

[dJfLlZ^{+}21] Zhen dong Ji, Jiao fen Li, Xue lin Zhou, Fu jian Duan, and Tao Li. On solutions of matrix equation AXB = C under semitensor product. Linear Multilinear Algebra, 69(10):1935–1963, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Djikic:2016:EFF

[Dji16] Marko S. Djikić. Extensions of the Fill–Fishkind formula and the infimum–parallel sum relation. Linear Multilinear Algebra, 64(11):2335–2349, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dryden:2008:AEC

[DJK08] Emily B. Dryden, Charles R. Johnson, and Brenda K. Kroschel. Adjacent edge conditions for the totally nonnegative completion problem. Linear Multilinear Algebra, 56(3):

Drew:1994:CPM

[DJL94] John H. Drew, Charles R. Johnson, and Raphael Loewy. Completely positive matrices associated with M-matrices. Linear and Multilinear Algebra, 37(4):303–310, 1994. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duric:2021:CZD

[DJS21] Alen Durić, Sara Jevdenić, and Nik Stopar. Compressed zerodivisor graphs of matrix rings over finite fields. Linear Multilinear Algebra, 69(11):2012–2039, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dacorogna:2011:DGA

[DK11] B. Dacorogna and O. Kneuss. Divisibility in Grassmann algebra. Linear Multilinear Algebra, 59(11):1201–1220, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drnovsek:2014:LGI

[DK14] Roman Drnovsek and Marko Kandić. From local to global ideal-triangularizability. Linear Multilinear Algebra, 62(12): 1616–1628, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dolinar:2016:HCR

[DK16] Gregor Dolinar and Bojan Kuzma. Homomorphisms of commutativity relation. Linear Multilinear Algebra, 64(5):897– 922, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2017:LSD

[DK17] Aiping Deng and Alexander Kelmans. Laplacian spectra of digraph transformations. Linear Multilinear Algebra, 65(4): 699–730, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duggal:2021:PBL

[DK21] B. P. Duggal and C. S. Kubrusly. Power bounded m-left invertible operators. Linear Multilinear Algebra, 69(3):515–525,

Duggal:2022:LIA

[DK22] B. P. Duggal and I. H. Kim. Left m-invertibility by the adjoint of Drazin inverse and m-selfadjointness of Hilbert spaces. Linear Multilinear Algebra, 70(12):2264–2277, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dehdezi:2023:GGE

[DK23] Eisa Khosravi Dehdezi and Saeed Karimi. GIBS: a general and efficient iterative method for computing the approximate inverse and Moore–Penrose inverse of sparse matrices based on the Schultz iterative method with applications. Linear Multilinear Algebra, 71(12):1905–1921, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duggal:2019:OSS

[DKK19] B. P. Duggal, I. H. Kim, and C. S. Kubrusly. Operators satisfying a similarity condition. Linear Multilinear Algebra, 67(5): 896–910, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dantas:2020:TNO

[DKK^{+}20] Sheldon Dantas, Vladimir Kadets, Sun Kwang Kim, Han Ju Lee, and Miguel Martín. There is no operatorwise version of the Bishop–Phelps–Bollobás property. Linear Multilinear Algebra, 68(9):1767–1778, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2010:BTI

[DKM10] S. S. Dragomir, M. Khosravi, and M. S. Moslehian. Besseltype inequalities in Hilbert C^{∗}-modules. Linear Multilinear Algebra, 58(8):967–975, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Djordjevic:2009:FEE

[DKS09] D. S. Djordjević, J. J. Koliha, and I. Straskraba. Factorization of EP elements in C^{∗}-algebras. Linear Multilinear Algebra, 57(6):587–594, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Djordjevic:2022:SSA

[DKY22] Slavisa V. Djordjević, Jaewoong Kim, and Jasang Yoon. Spectra of the spherical Aluthge transform, the linear pencil, and a commuting pair of operators. Linear Multilinear Algebra, 70(13):2533–2550, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drury:2014:RFD

[DL14] Stephen Drury and Minghua Lin. Reversed Fischer determinantal inequalities. Linear Multilinear Algebra, 62(8):1069– 1075, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Diaz:2022:DFT

[DL22] Steven P. Diaz and Adam Lutoborski. Discrete Fourier transform tensors and their eigenvalues. Linear Multilinear Algebra, 70(15):2934–2947, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2021:RCE

[DLD21] Xiaoli Deng, Chujian Lin, and Chunyuan Deng. Remark on the core-EP pre-order of Hilbert space operators. Linear Multilinear Algebra, 69(11):2006–2011, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2022:SPS

[DLD22] Xiaoli Deng, Chujian Lin, and Chunyuan Deng. The structures and properties of some generalized invertible operators. Linear Multilinear Algebra, 70(19):3818–3839, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dong:2019:DDT

[DLL19] Liqiang Dong, Jicheng Li, and Guo Li. The double deflating technique for irreducible singular M-matrix algebraic Riccati equations in the critical case. Linear Multilinear Algebra, 67(8):1653–1684, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deLima:2017:NBG

[dLMO17] L. S. de Lima, A. Mohammadian, and C. S. Oliveira. The nonbipartite graphs with all but two eigenvalues in [−1, 1]. Linear

delaPena:2009:CGS

[dlP09] José A. de la Peña. Cleavages of graphs: the spectral radius. Linear Multilinear Algebra, 57(7):641–649, 2009. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

delaPuente:2020:QEC

[dlP20] M. J. de la Puente. Quasi-Euclidean classification of alcoved convex polyhedra. Linear Multilinear Algebra, 68(10):2110– 2142, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Diao:2018:CAW

[DLQ18] Huai-An Diao, Liming Liang, and Sanzheng Qiao. A condition analysis of the weighted linear least squares problem using dual norms. Linear Multilinear Algebra, 66(6):1085–1103, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2018:MSS

[DLSS18] Bo Deng, Xueliang Li, Bryan Shader, and Wasin So. On the maximum skew spectral radius and minimum skew energy of tournaments. Linear Multilinear Algebra, 66(7):1434–1441, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2018:EMP

[DLW18] Chunyuan Deng, Rufang Liu, and Xiunan Wang. Expression for the multiplicative perturbation of the Moore–Penrose inverse. Linear Multilinear Algebra, 66(6):1171–1185, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deutsch:1987:DPV

[DM87] Emeric Deutsch and Neumann Michael. On the derivative of the Perron vector whose infinity norm is fixed. Linear and Multilinear Algebra, 21(1):75–85, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dolinar:2013:ALP

[DM13] Gregor Dolinar and Lajos Molnár. Automorphisms for the logarithmic product of positive semidefinite operators. Linear Multilinear Algebra, 61(2):161–169, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drnovsek:2014:JNR

[DM14] Roman Drnovsek and Vladimir Müller. On joint numerical radius II. Linear Multilinear Algebra, 62(9):1197–1204, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dadkhah:2017:GTI

[DM17] Ali Dadkhah and Mohammad Sal Moslehian. Grüss type inequalities for positive linear maps on C^{∗}-algebras. Linear Multilinear Algebra, 65(7):1386–1401, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Derksen:2018:NCR

[DM18] Harm Derksen and Visu Makam. On non-commutative rank and tensor rank. Linear Multilinear Algebra, 66(6):1069–1084, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Derksen:2019:ETB

[DM19] Harm Derksen and Visu Makam. Explicit tensors of border rank at least 2d−2 in K^{d}⊗K^{d}⊗K^{d} in arbitrary characteristic. Linear Multilinear Algebra, 67(10):2104–2116, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dadkhah:2020:NLP

[DM20] Ali Dadkhah and Mohammad Sal Moslehian. Non-linear positive maps between C∗-algebras. Linear Multilinear Algebra, 68(8):1501–1517, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2022:DMM

[DM22a] Joyentanuj Das and Sumit Mohanty. Distance matrix of a multi-block graph: determinant and inverse. Linear Multilinear Algebra, 70(19):3994–4022, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2022:DMW

[DM22b] Joyentanuj Das and Sumit Mohanty. Distance matrix of weighted cactoid-type digraphs. Linear Multilinear Algebra, 70(20):5392–5422, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dehghani-Madiseh:2022:FES

[DM22c] Marzieh Dehghani-Madiseh. Fast enclosing the solution set of the parametric Sylvester matrix equations. Linear Multilinear Algebra, 70(10):1814–1830, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Derksen:2022:HET

[DM22d] Harm Derksen and Visu Makam. Highly entangled tensors. Linear Multilinear Algebra, 70(2):380–393, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duffner:1992:SVC

[DMdO92] M. Antónia Duffner, J. F. Machado, and G. N. de Oliveira. Sets of vectors with a common minimal polynomial. Linear and Multilinear Algebra, 33(3–4):285–294, 1992. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dopico:2023:BFR

[DMQV23] Froilán M. Dopico, Silvia Marcaida, María C. Quintana, and Paul Van Dooren. Block full rank linearizations of rational matrices. Linear Multilinear Algebra, 71(3):391–421, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dhanoa:2023:GWY

[DMS23] Rupinderjit Kaur Dhanoa, Mohammad Sal Moslehian, and Mandeep Singh. A generalized Wigner–Yanase skew information. Linear Multilinear Algebra, 71(10):1720–1729, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dehimi:2021:OE

[DMT21] Souheyb Dehimi, Mohammed Hichem Mortad, and Zsigmond Tarcsay. On the operator equations A^{n} = A ∗ A. Linear Multilinear Algebra, 69(9):1771–1778, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeMulder:2018:GID

[DMV18] Wim De Mulder, Geert Molenberghs, and Geert Verbeke. A generalization of inverse distance weighting and an equivalence relationship to noise-free Gaussian process interpolation via Riesz representation theorem. Linear Multilinear Algebra, 66(5):1054–1066, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dubey:2018:GPS

[DN18] Dipti Dubey and S. K. Neogy. On generalizations of positive subdefinite matrices and the linear complementarity problem. Linear Multilinear Algebra, 66(10):2024–2035, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

doNascimento:2019:SGI

[dNV19] T. S. do Nascimento and A. C. Vieira. Superalgebras with graded involution and star-graded colength bounded by 3. Linear Multilinear Algebra, 67(10):1999–2020, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dressler:2022:SHT

[DNY22] Mareike Dressler, Jiawang Nie, and Zi Yang. Separability of Hermitian tensors and PSD decompositions. Linear Multilinear Algebra, 70(21):6581–6608, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DoRosarioFernandes:1996:PMS

[Do 96] Maria Do Rosário Fernandes. Pairs of matrices that have the same immanant. Linear and Multilinear Algebra, 40(3):193– 201, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deOliveira:1975:MPC

[dO75] G. N. de Oliveira. Matrices with prescribed characteristic polynomial and several prescribed submatrices. Linear and Multilinear Algebra, 2(4):357–364, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Datta:2019:LCU

[DO19] Somantika Datta and Jesse Oldroyd. Low coherence unit norm tight frames. Linear Multilinear Algebra, 67(6):1174–1189,

Dolzan:2020:CDM

[DO20] David Dolzan and Polona Oblak. Cholesky decomposition of matrices over commutative semirings. Linear Multilinear Algebra, 68(5):1057–1063, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dodig:2013:RDP

[Dod13] Marija Dodig. Rank distance problem for pairs of matrices. Linear Multilinear Algebra, 61(2):205–215, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dolinar:2004:MPI

[Dol04] Gregor Dolinar. Maps on B(H) preserving idempotents. Linear Multilinear Algebra, 52(5):335–347, 2004. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dolinar:2007:MUT

[Dol07] Gregor Dolinar. Maps on upper triangular matrices preserving Lie products. Linear Multilinear Algebra, 55(2):191–198, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Domokos:1995:NIM

[Dom95] M. Domokos. New identities for 3 × 3 matrices. Linear and Multilinear Algebra, 38(3):207–213, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Domokos:2000:RIM

[Dom00] Mátyá Domokos. Relative invariants of 3 × 3 matrix triples. Linear and Multilinear Algebra, 47(2):175–190, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Domokos:2014:ITC

[Dom14] M. Domokos. Invariant theoretic characterization of subdiscriminants of matrices. Linear Multilinear Algebra, 62(1):63– 73, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deOliveira:1977:EM

[dOMD77] G. N. de Oliveira, Eduardo Marques De Sá, and J. A. Dias Da Silva. On the eigenvalues of the matrix A + XBX^{−1}. Linear and Multilinear Algebra, 5(2):119–128, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Donkin:1991:RSG

[Don91] Stephen Donkin. Representations of symplectic groups and the symplectic tableaux of R. C. King. Linear and Multilinear Algebra, 29(2):113–124, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Doostaki:2016:CRG

[Doo16] R. Doostaki. Convergence rate of GMRES on tridiagonal block Toeplitz linear systems. Linear Multilinear Algebra, 64(12): 2533–2544, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Delgado:2022:CAC

[DOP22] Jorge Delgado, Héctor Orera, and J. M. Peña. Characterizations and accurate computations for tridiagonal Toeplitz matrices. Linear Multilinear Algebra, 70(19):4508–4527, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dozier:1979:BR

[Doz79] Jeff Dozier. Book reviews. Linear and Multilinear Algebra, 7(3):263–264, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

dePillis:1974:DTS

[dP74] John de Pillis. Decomposable tensors as sums of dyads. Linear and Multilinear Algebra, 1(4):327–335, 1974. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DaFonseca:1998:PPC

[DP98] C. M. Da Fonseca and J. Petronilho. Path polynomials of a circuit: a constructive approach. Linear and Multilinear Algebra, 44(4):313–325, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Darafsheh:2000:OBS

[DP00] M. R. Darafsheh and M. R. Pournaki. On the orthogonal basis of the symmetry classes of tensors associated with the dicyclic group. Linear and Multilinear Algebra, 47(2):137–149, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeBruyn:2009:SPS

[DP09] B. De Bruyn and A. Pasini. On symplectic polar spaces over non-perfect fields of characteristic 2. Linear Multilinear Algebra, 57(6):567–575, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Das:2017:NLS

[DP17] Arpita Das and Pratima Panigrahi. Normalized Laplacian spectrum of some subdivision-coronas of two regular graphs. Linear Multilinear Algebra, 65(5):962–972, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dixon:2018:GCP

[DP18] John D. Dixon and Irwin S. Pressman. Generalized commutators and a problem related to the Amitsur–Levitzki theorem. Linear Multilinear Algebra, 66(11):2199–2207, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

dePillis:1981:NST

[dPN81] John de Pillis and Michael Neumann. A noncommutative spectral theorem for operator-entried companion matrices. Linear and Multilinear Algebra, 10(1):45–51, 1981. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dasari:2017:SSP

[DPP17] Rajaiah Dasari, Rajkumar L. P., and Malla Reddy Perati. Steady state probability vector of positive definite regularized linear systems of circulant stochastic matrices. Linear Multilinear Algebra, 65(1):140–155, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Daptari:2022:SUH

[DPR22] Soumitra Daptari, Tanmoy Paul, and T. S. S. R. K. Rao. Stability of unique Hahn-Banach extensions and associated linear projections. Linear Multilinear Algebra, 70(21):6067– 6083, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dascalescu:2013:JIG

[DPvW13] S. Dascalescu, S. Predut, and L. van Wyk. Jordan isomorphisms of generalized structural matrix rings. Linear Multilinear Algebra, 61(3):369–376, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Douglas:2018:SRT

[DR18] Andrew Douglas and Joe Repka. Subalgebras of the rank two semisimple Lie algebras. Linear Multilinear Algebra, 66(10): 2049–2075, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2010:CTI

[Dra10] S. S. Dragomir. Čebyšev’s type inequalities for functions of selfadjoint operators in Hilbert spaces. Linear Multilinear Algebra, 58(7):805–814, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2011:RCB

[Dra11] S. S. Dragomir. Refinements of the Cauchy–Bunyakovsky– Schwarz inequality for functions of selfadjoint operators in Hilbert spaces. Linear Multilinear Algebra, 59(7):711–717, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2013:GFI

[Dra13a] S. S. Dragomir. Generalizations of Furuta’s inequality. Linear Multilinear Algebra, 61(5):617–626, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drazin:2013:CPG

[Dra13b] Michael P. Drazin. Commuting properties of generalized inverses. Linear Multilinear Algebra, 61(12):1675–1681, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2016:STI

[Dra16] S. S. Dragomir. Some trace inequalities of Čebyšev type for functions of operators in Hilbert spaces. Linear Multilinear

Dragomir:2017:BTI

[Dra17a] S. S. Dragomir. A buzano type inequality for two Hermitian forms and applications. Linear Multilinear Algebra, 65(3):514– 525, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drazin:2017:BPG

[Dra17b] Michael P. Drazin. Bicommuting properties of generalized inverses. Linear Multilinear Algebra, 65(9):1770–1780, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2018:SIH

[Dra18a] S. S. Dragomir. Some inequalities of Hölder type for quadratic weighted geometric mean of bounded linear operators in Hilbert spaces. Linear Multilinear Algebra, 66(2):268–279, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drazin:2018:SI

[Dra18b] Michael P. Drazin. Subclasses of (b, c)-inverses. Linear Multilinear Algebra, 66(1):184–192, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2019:ISI

[Dra19a] Silvestru Sever Dragomir. Improving Schwarz inequality in inner product spaces. Linear Multilinear Algebra, 67(2):337– 347, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2019:ORS

[Dra19b] Silvestru Sever Dragomir. Operator refinements of Schwarz inequality in inner product spaces. Linear Multilinear Algebra, 67(9):1839–1855, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragicevic:2021:ESH

[Dra21a] Davor Dragicević. On the exponential stability and hyperbolicity of linear cocycles. Linear Multilinear Algebra, 69(2):

Dragomir:2021:OSC

[Dra21b] Silvestru Sever Dragomir. Operator Schur convexity and some integral inequalities. Linear Multilinear Algebra, 69(14):2733– 2748, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dragomir:2022:SIA

[Dra22a] Silvestru Sever Dragomir. Some inequalities for analytic functions in Banach algebras. Linear Multilinear Algebra, 70(12): 2211–2226, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drazin:2022:EPI

[Dra22b] Michael P. Drazin. EP properties of (b, c)-invertible matrices. Linear Multilinear Algebra, 70(3):431–437, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drazin:2023:IPG

[Dra23] Michael P. Drazin. Intermixing pairs of generalized inverses. Linear Multilinear Algebra, 71(8):1397–1406, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drensky:1993:FGI

[Dre93] Vesselin Drensky. Finite generation of invariants of finite linear groups on relatively free algebras. Linear and Multilinear Algebra, 35(1):1–10, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drensky:2022:IMR

[Dre22] Vesselin Drensky. Idempotents of 2×2 matrix rings over rings of formal power series. Linear Multilinear Algebra, 70(20): 5207–5213, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drnovsek:2013:NEU

[Drn13] Roman Drnovsek. The von Neumann entropy and unitary equivalence of quantum states. Linear Multilinear Algebra, 61 (10):1391–1393, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drury:2014:FTD

[Dru14] Stephen Drury. A Fischer-type determinantal inequality. Linear Multilinear Algebra, 62(11):1421–1427, 2014. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Drury:2015:PPM

[Dru15] Stephen Drury. Principal powers of matrices with positive definite real part. Linear Multilinear Algebra, 63(2):296–301, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dooley:1993:SAO

[DRW93] A. H. Dooley, J. Repka, and N. J. Wildberger. Sums of adjoint orbits. Linear and Multilinear Algebra, 36(2):79–101, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2022:LPM

[DRY22] Weina Deng, Miaomiao Ren, and Baomin Yu. Linear preservers for matrices over a class of semirings. Linear Multilinear Algebra, 70(21):6830–6845, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiBenedetto:2002:NSM

[DS02a] Fabio Di Benedetto and Stefano Serra Capizzano. A note on the superoptimal matrix algebra operators. Linear Multilinear Algebra, 50(4):343–372, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Donley:2002:SOR

[DS02b] Robert W. Donley and Mark Sepanski. Schur orthogonality relations for the finitary infinite symmetric group. Linear Multilinear Algebra, 50(4):373–377, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dolinar:2004:MMA

[DŠ04] Gregor Dolinar and Peter Šemrl. Maps on matrix algebras preserving commutativity. Linear Multilinear Algebra, 52(1): 69–78, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dodig:2009:SCM

[DS09] M. Dodig and M. Stosić. Similarity class of a matrix with a prescribed submatrix. Linear Multilinear Algebra, 57(3): 217–245, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeFilippis:2013:SCE

[DS13] Vincenzo De Filippis and Giovanni Scudo. Strong commutativity and Engel condition preserving maps in prime and semiprime rings. Linear Multilinear Algebra, 61(7):917–938, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Diaz:2016:NIE

[DS16] Roberto C. Díaz and Ricardo L. Soto. Nonnegative inverse elementary divisors problem in the left half plane. Linear Multilinear Algebra, 64(2):258–268, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dwivedi:2018:HMF

[DS18] Ravi Dwivedi and Vivek Sahai. On the hypergeometric matrix functions of two variables. Linear Multilinear Algebra, 66(9): 1819–1837, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deSa:2019:LE

[dS19a] Eduardo Marques de Sá. Letter to the editor. Linear Multilinear Algebra, 67(8):1711–1712, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dwivedi:2019:BHM

[DS19b] Ravi Dwivedi and Vivek Sahai. On the basic hypergeometric matrix functions of two variables. Linear Multilinear Algebra, 67(1):1–19, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dodig:2022:CMP

[DS22a] Marija Dodig and Marko Stosić. Completion of matrix pencils with a single rank restriction. Linear Multilinear Algebra, 70 (21):7027–7048, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dodig:2022:NAS

[DS22b] Marija Dodig and Marko Stosić. A new approach to square matrix completion problem. Linear Multilinear Algebra, 70 (13):2602–2610, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeFilippis:2023:ECP

[DS23a] Vincenzo De Filippis and Giovanni Scudo. An equation concerning power values of generalized skew derivations and annihilating conditions on Lie ideals. Linear Multilinear Algebra, 71(9):1431–1449, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deepshikha:2023:ANO

[DS23b] Deepshikha and Aniruddha Samanta. Averaged numerically optimal dual frames for erasures. Linear Multilinear Algebra, 71(2):301–316, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dubon:2023:APM

[DS23c] E. Dubon and A. San Antolín. On approximation properties of matrix-valued multi-resolution analyses. Linear Multilinear Algebra, 71(14):2263–2281, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DeOliveira:1983:NES

[DSD83] G. N. De Oliveira, Ana P. Santana, and J. A. Dias Da Silva. Note on the equality of star products. Linear and Multilinear Algebra, 14(2):157–163, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pazzis:2011:SMS

[dSP11] Clément de Seguins Pazzis. On small matrix subalgebras with a trivial centralizer. Linear Multilinear Algebra, 59(7):761– 771, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pazzis:2012:SMS

[dSP12] Clément de Seguins Pazzis. Spaces of matrices with a sole eigenvalue. Linear Multilinear Algebra, 60(10):1165–1190, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pazzis:2016:PSM

[dSP16a] Clément de Seguins Pazzis. Primitive spaces of matrices with upper rank two over the field with two elements. Linear Multilinear Algebra, 64(7):1321–1353, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pazzis:2016:QRC

[dSP16b] Clément de Seguins Pazzis. Quasi-range-compatible affine maps on large operator spaces. Linear Multilinear Algebra, 64(6):1056–1085, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pazzis:2017:NST

[dSP17] Clément de Seguins Pazzis. A note on sums of three squarezero matrices. Linear Multilinear Algebra, 65(4):787–805, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pazzis:2022:SPP

[dSP22] Clément de Seguins Pazzis. Sums and products of pairs of quadratic endomorphisms of a countable-dimensional vector space. Linear Multilinear Algebra, 70(21):6718–6752, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pinto:2022:MVP

[dSPdS22] Marcos Antônio da Silva Pinto and Viviane Ribeiro Tomaz da Silva. On the minimal varieties of PI-algebras graded by finite cyclic groups. Linear Multilinear Algebra, 70(20):5790– 5814, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Day:1991:SPC

[DST91] J. Day, W. So, and Robert C. Thompson. Some properties of the Campbell–Baker–Hausdorff series. Linear and Multilinear Algebra, 29(3–4):207–224, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dehghan:2014:LBG

[DTS14] Mohammad Ali Dehghan, Gholamreza Talebi, and Alireza Shojaeifard. Lower bounds for generalized Hausdorff matrices and lower triangular matrices on the block weighted sequence space l_{p}(w, F). Linear Multilinear Algebra, 62(1):126– 138, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

deTraubenberg:2020:PIM

[dTS20] M. Rausch de Traubenberg and M. J. Slupinski. Polynomial invariants and moduli of generic two-dimensional commutative algebras. Linear Multilinear Algebra, 68(3):472–527, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Defez:2015:NAS

[DTSI15] Emilio Defez, Michael M. Tung, Francisco J. Solis, and Javier Ibáñez. Numerical approximations of second-order matrix differential equations using higher degree splines. Linear Multilinear Algebra, 63(3):472–489, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2010:NII

[Du10] Kun Du. Norm inequalities involving Hadamard products. Linear Multilinear Algebra, 58(2):235–238, 2010. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duisengaliyeva:2021:DAD

[DU21] Bibinur Duisengaliyeva and Ualbai Umirbaev. Differential algebraic dependence and Novikov dependence. Linear Multilinear Algebra, 69(6):1061–1071, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duan:2022:IIS

[Dua22] Fang Duan. The inertia indices of a signed graph in terms of the inertia indices of its underlying graph. Linear Multilinear Algebra, 70(3):497–510, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duffner:1997:NSM

[Düf97] M. Antónia Düffner. A note on singular matrices satisfying certain polynomial identities. Linear and Multilinear Algebra, 42(3):213–219, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dinh:2018:SIO

[DV18] Trung Hoa Dinh and Khue Thi Bich Vo. Some inequalities for operator (p, h)-convex functions. Linear Multilinear Algebra, 66(3):580–592, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dimant:2019:DMO

[DV19] Verónica Dimant and Román Villafañe. Diagonal multilinear operators on Köthe sequence spaces. Linear Multilinear Algebra, 67(2):248–266, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dirr:2020:ANR

[DvE20a] Gunther Dirr and Frederik vom Ende. Addendum: ‘The Cnumerical range in infinite dimensions’. Linear Multilinear Algebra, 68(4):867–868, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [DvE20b].

Dirr:2020:NRI

[DvE20b] Gunther Dirr and Frederik vom Ende. The C-numerical range in infinite dimensions. Linear Multilinear Algebra, 68(4):652– 678, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See addendum [DvE20a].

Davis:1976:BR

[DW76] Chandler Davis and Jennifer Seberry Wallis. Book reviews. Linear and Multilinear Algebra, 4(2):149–154, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Day:1992:PSC

[DW92] Jane M. Day and Donald D. Weddington. Preaffine semigroups and convex matrix semigroups. Linear and Multilinear Algebra, 31(1–4):193–216, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Deng:2012:FRM

[DW12a] Chunyuan Deng and Yimin Wei. Further results on the Moore–Penrose invertibility of projectors and its applications. Linear Multilinear Algebra, 60(1):109–129, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Du:2012:AJM

[DW12b] Yiqiu Du and Yu Wang. Additivity of Jordan maps on triangular algebras. Linear Multilinear Algebra, 60(8):933–940, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dugas:2016:FIA

[DW16] Manfred Dugas and Bradley Wagner. Finitary incidence algebras and idealizations. Linear Multilinear Algebra, 64(10): 1936–1951, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dong:2023:DWR

[DW23] Xiaomei Dong and Deyu Wu. Davis–Wielandt radius inequalities for off-diagonal operator matrices. Linear Multilinear Algebra, 71(7):1158–1181, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dou:2019:SAT

[DWCD19] Yan-Ni Dou, Yue-Qing Wang, Miao-Miao Cui, and Hong-Ke Du. Spectra of anticommutator for two orthogonal projections. Linear Multilinear Algebra, 67(10):2077–2081, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duan:2013:MEA

[DWL13] Xuefeng Duan, Qingwen Wang, and Anping Liao. On the matrix equation arising in an interpolation problem. Linear Multilinear Algebra, 61(9):1192–1205, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duan:2014:PDS

[DWL14] Xue-Feng Duan, Qing-Wen Wang, and Chun-Mei Li. Positive definite solution of a class of nonlinear matrix equation. Linear Multilinear Algebra, 62(6):839–852, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duan:2020:ECP

[DWL20] Cunxiang Duan, Ligong Wang, and Xiangxiang Liu. Edge connectivity, packing spanning trees, and eigenvalues of graphs. Linear Multilinear Algebra, 68(6):1077–1095, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Duan:2022:TFG

[DZ22] Fang Duan and Weijuan Zhang. Triangle-free graphs with six non-zero eigenvalues. Linear Multilinear Algebra, 70(19): 4214–4227, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ernest:1977:BR

[EA77] John Ernest and Charles A. Akemann. Book reviews. Linear and Multilinear Algebra, 5(3):243–247, 1977. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

ElAbidineAbdelali:2019:MPL

[EAM19] Zine El Abidine Abdelali, Abdelali Achchi, and Rabi Marzouki. Maps preserving the local spectral radius zero of generalized product of operators. Linear Multilinear Algebra, 67 (10):2021–2029, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ercan:2017:STG

[EB17] Sinan Ercan and Çigdem A. Bektas. Some topological and geometric properties of a new BK-space derived by using regular matrix of Fibonacci numbers. Linear Multilinear Algebra, 65(5):909–921, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ebanks:1992:FMM

[Eba92] Bruce R. Ebanks. Factoring multiadditive mappings. Linear and Multilinear Algebra, 33(3–4):175–184, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eshkaftaki:2018:DWM

[EBB18] A. Bayati Eshkaftaki, M. Heydari Berenjegani, and F. Bahrami. DSS-weak majorization and its linear preservers on spaces. Linear Multilinear Algebra, 66(10):2076–2088, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eberlein:1980:TF

[Ebe80] Patrick Eberlein. A trace formula. Linear and Multilinear Algebra, 9(3):231–236, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ed-Dari:2007:NIV

[EDKA07] Elmouloudi Ed-Dari, Mohamed Amine Khamsi, and Asuman Güven Aksoy. On the numerical index of vector-valued function spaces. Linear Multilinear Algebra, 55(6):507–513, 2007. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Edwards:1978:EFD

[Edw78] Bruce H. Edwards. The eigenvalues of four-dimensional rotations. Linear and Multilinear Algebra, 5(4):283–287, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Edwards:1980:RDQ

[Edw80] Bruce H. Edwards. Rotations and discriminants of quadratic spaces. Linear and Multilinear Algebra, 8(3):241–246, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Edelman:1998:CHM

[EF98] Alan Edelman and David Friendman. A counterexample to a Hadamard matrix pivot conjecture. Linear and Multilinear Algebra, 44(1):53–56, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eggar:1995:TOA

[Egg95] M. H. Eggar. Two observations about free graded Hopf algebras. Linear and Multilinear Algebra, 38(4):333–341, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elyze:2022:HDC

[EGMS22] Madeleine Elyze, Alexander Guterman, Ralph Morrison, and Klemen Sivic. Higher-distance commuting varieties. Linear Multilinear Algebra, 70(17):3248–3270, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

ElKaoutit:2016:HTM

[EGT16] Laiachi El Kaoutit and José Gómez-Torrecillas. Hereditary triangular matrix comonads. Linear Multilinear Algebra, 64(6): 1032–1055, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eschenbach:1988:RP

[EJ88] C. Eschenbach and C. R. Johnson. Research prolems. Linear and Multilinear Algebra, 24(1):79–80, 1988. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eschenbach:1991:SPR

[EJ91] Carolyn A. Eschenbach and Charles R. Johnson. Sign patterns that require real, nonreal or pure imaginary eigenvalues. Linear and Multilinear Algebra, 29(3–4):299–311, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elsner:1988:PRW

[EJD88] L. Elsner, Charles R. Johnson, and J. A. Dias Da Silva. The Perron root of a weighted geometric mean of nonnegative matrices. Linear and Multilinear Algebra, 24(1):1–13, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

ElIchi:2022:TTK

[EJS22] Alaa El Ichi, Khalide Jbilou, and Rachid Sadaka. On tensor tubal-Krylov subspace methods. Linear Multilinear Algebra, 70(22):7575–7598, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elduque:2014:LUK

[EKO14] Alberto Elduque, Noriaki Kamiya, and Susumu Okubo. Left unital Kantor triple systems and structurable algebras. Linear Multilinear Algebra, 62(10):1293–1313, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ellis:1990:RPE

[EL90] Robert L. Ellis and David C. Lay. Rank-preserving extensions of band matrices. Linear and Multilinear Algebra, 26(3):147– 179, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eliahou:2009:LVS

[EL09] Shalom Eliahou and Cédric Lecouvey. On linear versions of some addition theorems. Linear Multilinear Algebra, 57(8): 759–775, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elduque:2021:EAA

[EL21] Alberto Elduque and Alicia Labra. Evolution algebras, automorphisms, and graphs. Linear Multilinear Algebra, 69(2): 331–342, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ellis:1988:NES

[ELG88] Robert L. Ellis, David C. Lay, and Israel Gohberg. On negative eigenvalues of selfadjoint eztensions of band matrices. Linear and Multilinear Algebra, 24(1):15–25, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elgendy:2022:RSJ

[Elg22] Hader A. Elgendy. Representations of special Jordan triple systems of all symmetric and Hermitian n by n matrices. Linear Multilinear Algebra, 70(21):6863–6875, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ellers:1993:RLT

[Ell93] Erich W. Ellers. The reflection length of a transformation in the unitary group over a finite field. Linear and Multilinear Algebra, 35(1):11–35, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ellers:1995:PRO

[Ell95] Erich W. Ellers. Products of reflections in one conjugacy class in the orthogonal group over GF(3). Linear and Multilinear Algebra, 39(3):209–229, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elouafi:2018:EIB

[Elo18] Mohamed Elouafi. Explicit inversion of band Toeplitz matrices by discrete Fourier transform. Linear Multilinear Algebra, 66(9):1767–1782, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elsner:1978:ANT

[Els78] Ludwig Elsner. Another note on a theorem of Minc on irreducible nonnegative matrices. Linear and Multilinear Algebra, 6(1):61–62, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Erdos:1973:DNM

[EM73] Paul Erdős and Henryk Minc. Diagonals of nonegative matrices. Linear and Multilinear Algebra, 1(2):89–95, 1973. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ellers:1987:PR

[EM87] Erich W. Ellers and Jerry Malzen. Products of reflections in. Linear and Multilinear Algebra, 20(4):281–324, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Edelman:1995:CPC

[EM95] Alan Edelman and Walters Mascarenhas. On the complete pivoting conjecture for a Hadamard matrix of order 12. Linear and Multilinear Algebra, 38(3):181–187, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

ElHarti:2015:MCE

[EM15] Rachid El Harti and Mohamed Mabrouk. Maps compressing and expanding the numerical range on C^{∗}-algebras. Linear Multilinear Algebra, 63(11):2332–2339, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elsner:2002:SPC

[EMS02] L. Elsner, V. Monov, and T. Szulc. On some properties of convex matrix sets characterized by P -matrices and block P - matrices. Linear Multilinear Algebra, 50(3):199–218, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Erdogdu:2015:TSL

[EÖ15] Melek Erdogdu and Mustafa Özdemir. Two-sided linear split quaternionic equations with n unknowns. Linear Multilinear Algebra, 63(1):97–106, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eisner:2009:SCP

[EOvdD09a] Ludwig Eisner, D. D. Olesky, and P. van den Driessche. Sufficient conditions for permutation equivalence to a WHSmatrix. Linear Multilinear Algebra, 57(1):103–110, 2009. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elsner:2009:SCP

[EOvdD09b] Ludwig Elsner, D. D. Olesky, and P. van den Driessche. Sufficient conditions for permutation equivalence to a WHSmatrix. Linear Multilinear Algebra, 57(3):319, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Epkenhans:1989:TFNb

[Epk89a] Martin Epkenhans. Trace forms of normal extensions of algebraic number fields. Linear and Multilinear Algebra, 25(4): 309–320, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Epkenhans:1989:TFNa

[Epk89b] Martin Epkenhans. Trace forms of normal extensions over local fields. Linear and Multilinear Algebra, 24(2):103–116, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Essaleh:2016:WLDa

[EPR16a] Ahlem Ben Ali Essaleh, Antonio M. Peralta, and María Isabel Ramírez. Weak-local derivations and homomorphisms on C^{∗}-algebras. Linear Multilinear Algebra, 64(2):169–186, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Essaleh:2016:WLDb

[EPR16b] Ahlem Ben Ali Essaleh, Antonio M. Peralta, and María Isabel Ramírez. Weak-local derivations and homomorphisms on C^{∗}- algebras. Linear Multilinear Algebra, 64(5):1009–1010, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elsner:1981:EAM

[ER81] L. Elsner and P. Rózsa. On eigenvectors and adjoints of modified matrices. Linear and Multilinear Algebra, 10(3):235–247, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Egecioglu:1990:CII

[ER90] Ömer Eğecioğlu and Jeffrey B. Remmel. A combinatorial interpretation of the inverse Kostka matrix. Linear and Multilinear Algebra, 26(1–2):59–84, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eisner:2022:ERP

[ER22] Tanja Eisner and Agnes Radl. Embeddability of real and positive operators. Linear Multilinear Algebra, 70(19):3747–3767, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Erdogan:2022:FMO

[Erd22] E. Erdogan. Factorization of multilinear operators defined on products of function spaces. Linear Multilinear Algebra, 70(2): 177–202, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eremita:2015:FID

[Ere15] Daniel Eremita. Functional identities of degree 2 in triangular rings revisited. Linear Multilinear Algebra, 63(3):534–553, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eldred:2012:NRC

[ERS12] Jeffrey Eldred, Leiba Rodman, and Ilya Spitkovsky. Numerical ranges of companion matrices: flat portions on the boundary. Linear Multilinear Algebra, 60(11–12):1295–1311, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Engel:1973:IDP

[ES73] Gernot M. Engel and Hans Schneider. Inequalities for determinants and permanents. Linear and Multilinear Algebra, 1(3): 187–201, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Engel:1976:HFI

[ES76a] Gernot M. Engel and Hans Schneider. The Hadamard–Fischer inequality for a class of matrices defined by eigenvalue monotonicity. Linear and Multilinear Algebra, 4(3):155–176, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Erdos:1976:HAF

[ES76b] P. Erdős and E. G. Straus. How abelian is a finite group? Linear and Multilinear Algebra, 3(4):307–312, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elsner:1998:BM

[ES98a] Ludwig Elsner and Tomasz Szulc. Block P -matrices. Linear and Multilinear Algebra, 44(1):1–12, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elsner:1998:CCM

[ES98b] Ludwig Elsner and Tomasz Szulc. Convex combinations of matrices-nonsingularity and Schur stability characterizations. Linear and Multilinear Algebra, 44(4):301–312, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Elsner:2000:CSS

[ES00] Ludwig Elsner and Tomasz Szulc. Convex sets of Schur stable and stable matrices. Linear and Multilinear Algebra, 48(1):1– 19, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

El-Shabrawy:2018:SGD

[ESAJ18] Saad R. El-Shabrawy and Suad H. Abu-Janah. Spectra of the generalized difference operator on the sequence spaces bv_{0} and h. Linear Multilinear Algebra, 66(8):1691–1708, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eschenbach:1993:SPR

[Esc93] Carolyn A. Eschenbach. Sign patterns that require exactly one real eigenvalue and patterns that require n − 1 nonreal eigenvalues. Linear and Multilinear Algebra, 35(3–4):213–223, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eshkaftaki:2017:GKC

[Esh17] Ali Bayati Eshkaftaki. Generalized Kakutani’s conjecture for doubly stochastic operators. Linear Multilinear Algebra, 65(7): 1311–1315, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eshkaftaki:2022:IDS

[Esh22a] Ali Bayati Eshkaftaki. Increasable doubly substochastic matrices with application to infinite linear equations. Linear Multilinear Algebra, 70(20):5902–5912, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Eshkaftaki:2022:MCD

[Esh22b] Ali Bayati Eshkaftaki. Minimal completion of I × I doubly substochastic matrices. Linear Multilinear Algebra, 70(11): 2065–2077, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Erfanifar:2023:SSN

[ESH23] Raziyeh Erfanifar, Khosro Sayevand, and Masoud Hajarian. Solving system of nonlinear matrix equations over Hermitian positive definite matrices. Linear Multilinear Algebra, 71(4): 597–630, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Estes:1980:SMM

[Est80] Dennis R. Estes. Scalar matrices as multiplicative commutators having prescribed determinants for the variables. Linear and Multilinear Algebra, 8(3):213–217, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Et-Taoui:2022:CSS

[ETM22] Boumediene Et-Taoui and Abdenacer Makhlouf. Complex skew-symmetric conference matrices. Linear Multilinear Algebra, 70(21):6648–6663, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Engel:1977:MGS

[EV77] G. M. Engel and R. S. Varga. Minimal Geršgorin sets and ω-matrices. Linear and Multilinear Algebra, 5(1):1–10, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2019:ECD

[FA19] Feiyan Fu and Runling An. Equivalent characterization of *-derivations on von Neumann algebras. Linear Multilinear Algebra, 67(3):527–541, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Franca:2022:LTP

[FA22] Willian Franca and Magno Alves. Linear transformations preserving algebraic elements of degree 2. Linear Multilinear Algebra, 70(19):4191–4213, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Falikman:1997:PCM

[Fal97] Dmitry Falikman. The permanent of a class of matrices associated with a finite group. Linear and Multilinear Algebra, 42(4): 323–347, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:1973:RMP

[Fan73] Ky Fan. On real matrices with positive definite symmetric component. Linear and Multilinear Algebra, 1(1):1–4, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:1981:BR

[Fan81] K. Y. Fan. Book review. Linear and Multilinear Algebra, 10(3): 261–262, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:1983:IAF

[Fan83] Ky Fan. Iteration of analytic functions of operators. II. Linear and Multilinear Algebra, 12(4):295–304, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:1984:BR

[Fan84] K. Fan. Book review. Linear and Multilinear Algebra, 15(2): 189–190, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:1988:ROE

[Fan88] Ky Fan. A remark on orthogonality of eigenvectors. Linear and Multilinear Algebra, 23(4):283–284, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:1992:SIM

[Fan92] Ky Fan. Some inequalities for matrices A such that A − I is positive definite or an M-matrix. Linear and Multilinear

Fan:2005:LEU

[Fan05] Yi-Zheng Fan. On the least eigenvalue of a unicyclic mixed graph. Linear Multilinear Algebra, 53(2):97–113, 2005. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:1978:BR

[FAR78] K. Y. Fan, Edward Azoff, and James B. Robertson. Book reviews. Linear and Multilinear Algebra, 5(4):324–328, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farenick:1990:FGU

[Far90] D. R. Farenick. The fundamental group of a unitary orbit. Linear and Multilinear Algebra, 26(1–2):1–4, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farenick:1993:MEN

[Far93a] D. R. Farenick. Matricial extensions of the numerical range: a brief survey. Linear and Multilinear Algebra, 34(3–4):197–211, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farmer:1993:NCP

[Far93b] Thomas A. Farmer. A new construction of polarizations of solvable Lie algebras. Linear and Multilinear Algebra, 35(3– 4):185–190, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farenick:1994:NCR

[Far94] D. R. Farenick. Normal compressions and restrictions to 2dimensional subspaces, and the Berberian spectral extension. Linear and Multilinear Algebra, 37(1–3):131–138, 1994. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farid:2002:NRO

[Far02] F. Farid. On the numerical range of operator polynomials. Linear Multilinear Algebra, 50(3):223–239, 2002. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farashahi:2017:TFP

[Far17] Arash Ghaani Farashahi. Theoretical frame properties of wave-packet matrices over prime fields. Linear Multilinear Algebra, 65(12):2508–2529, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farashahi:2020:GWT

[Far20] Arash Ghaani Farashahi. Generalized wavelet transforms over finite fields. Linear Multilinear Algebra, 68(8):1585–1604, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fard:2022:CIF

[Far22] Mohammad Ali Hasankhani Fard. Construction of infinite frames with some given redundancy. Linear Multilinear Algebra, 70(21):6852–6862, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fang:2012:STS

[FD12] Li Fang and Hongke Du. Stability theorems for some combinations of two idempotents. Linear Multilinear Algebra, 60(2): 159–165, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fard:2014:NFV

[FD14] M. A. Hasankhani Fard and M. A. Dehghan. A new function-valued inner product and corresponding functionvalued frames in L_{2}(0, ∞). Linear Multilinear Algebra, 62(8): 995–1009, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:2017:SBL

[FD17] Guangzhe Fan and Xiansheng Dai. Super-biderivations of Lie superalgebras. Linear Multilinear Algebra, 65(1):58–66, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:2014:NBS

[FDD14] Yi-Zheng Fan, Wen-Xue Du, and Chun-Long Dong. The nullity of bicyclic signed graphs. Linear Multilinear Algebra, 62(2):242–251, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fernandes:2009:IEP

[FdF09] Rosário Fernandes and C. M. da Fonseca. The inverse eigenvalue problem for Hermitian matrices whose graphs are cycles. Linear Multilinear Algebra, 57(7):673–682, 2009. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:2020:MET

[FDLB20] Zhaobing Fan, Chunli Deng, Haifeng Li, and Changjiang Bu. Multiplications and eigenvalues of tensors via linear maps. Linear Multilinear Algebra, 68(3):606–621, 2020. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fonseca:2018:DII

[FdM18] Luís Felipe Gonçalves Fonseca and Thiago Castilho de Mello. Degree-inverting involutions on matrix algebras. Linear Multilinear Algebra, 66(6):1104–1120, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fonseca:2016:CSP

[FdSV16] Luís Felipe Gonçalves Fonseca, Rafael Bezerra dos Santos, and Ana Cristina Vieira. Characterizations of *-superalgebras of polynomial growth. Linear Multilinear Algebra, 64(7):1379– 1389, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Feit:1975:CRQ

[Fei75] Walter Feit. On certain rational quadratic forms. Linear and Multilinear Algebra, 3(1–2):25–32, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Feinberg:1979:IMC

[Fei79] Robert B. Feinberg. The identity as a matrix commutator. Linear and Multilinear Algebra, 7(2):121–122, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Feng:2010:DST

[Fen10] Lianggui Feng. Decompositions of some types of quaternionic matrices. Linear Multilinear Algebra, 58(4):431–444, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fernandes:2009:RPC

[Fer09] Rosário Fernandes. Rank partitions and covering numbers under small perturbations of an element. Linear Multilinear Algebra, 57(4):387–398, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fernandes:2010:LEW

[Fer10a] Rosário Fernandes. Location of the eigenvalues of weighted graphs with a cut edge. Linear Multilinear Algebra, 58(3): 305–322, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fernandes:2010:SPP

[Fer10b] Rosário Fernandes. Small perturbations and pairs of matrices that have the same immanent. Linear Multilinear Algebra, 58(8):977–991, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fernandes:2021:EMB

[FF21] Rosário Fernandes and Susana Furtado. Extremal matrices for the Bruhat-graph order. Linear Multilinear Algebra, 69(7): 1255–1274, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiol:2002:PSR

[FG02] M. A. Fiol and E. Garriga. Pseudo-strong regularity around a set. Linear Multilinear Algebra, 50(1):33–47, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiol:2006:OSS

[FG06] M. A. Fiol and E. Garriga. On outindependent subgraphs of strongly regular graphs. Linear Multilinear Algebra, 54(2): 123–140, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2022:TII

[FG22] Xiaohui Fu and Mehmet Gumus. Trace inequalities involving positive semi-definite block matrices. Linear Multilinear Algebra, 70(20):5987–5994, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:2022:STP

[FGHL22] Dandan Fan, Sergey Goryainov, Xueyi Huang, and Huiqiu Lin. The spanning k-trees, perfect matchings and spectral radius of graphs. Linear Multilinear Algebra, 70(21):7264– 7275, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fidelis:2023:NGG

[FGK23] Claudemir Fidelis, Alan Guimarães, and Plamen Koshlukov. A note on -gradings on the Grassmann algebra and elementary number theory. Linear Multilinear Algebra, 71(7):1244–1264, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ferrer:2019:ZSH

[FGMSS19] Jesús Ferrer, Domingo García, Manuel Maestre, and Juan B. Seoane-Sepúlveda. On the zero-set of 2-homogeneous polynomials in Banach spaces. Linear Multilinear Algebra, 67(10): 1958–1970, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiedler:1973:CWT

[FH73] Miroslav Fiedler and Emilie Haynsworth. Cones which are topheavy with respect to a norm. Linear and Multilinear Algebra, 1(3):203–211, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Freund:1989:CEP

[FH89] Roland Freund and Thomas Huckle. On H-contractions and the extension problem for Hermitian block Toeplitz matrices. Linear and Multilinear Algebra, 25(1):27–37, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2015:SOI

[FH15a] Xiaohui Fu and Chuanjiang He. Some operator inequalities for positive linear maps. Linear Multilinear Algebra, 63(3): 571–577, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2015:SIM

[FH15b] Xiaohui Fu and Dinh Trung Hoa. On some inequalities with matrix means. Linear Multilinear Algebra, 63(12):2373–2378,

Feng:2013:SPA

[FHL^{+}13] Xin-Lei Feng, Ting-Zhu Huang, Zhongshan Li, Jun Luo, and Yubin Gao. Sign patterns that allow diagonalizability revisited. Linear Multilinear Algebra, 61(9):1223–1233, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fosner:2014:LMP

[FHL^{+}14] Ajda Fošner, Zejun Huang, Chi-Kwong Li, Yiu-Tung Poon, and Nung-Sing Sze. Linear maps preserving the higher numerical ranges of tensor products of matrices. Linear Multilinear Algebra, 62(6):776–791, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fosner:2013:LPQ

[FHLS13] Ajda Fošner, Zejun Huang, Chi-Kwong Li, and Nung-Sing Sze. Linear preservers and quantum information science. Linear Multilinear Algebra, 61(10):1377–1390, 2013. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fritscher:2014:UGE

[FHT14] Eliseu Fritscher, Carlos Hoppen, and Vilmar Trevisan. Unicyclic graphs with equal Laplacian energy. Linear Multilinear Algebra, 62(2):180–194, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farid:2016:CES

[FHW16] F. O. Farid, Zhuo-Heng He, and Qing-Wen Wang. The consistency and the exact solutions to a system of matrix equations. Linear Multilinear Algebra, 64(11):2133–2158, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Feng:2017:CCE

[FHZ17] Yan-Quan Feng, Rong-Xia Hao, and Jin-Xin Zhou. On computing of a conditional edge connectivity of alternating group network. Linear Multilinear Algebra, 65(12):2494–2507, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiedler:1973:QEP

[Fie73] Mirsolav Fiedler. A quantitative extension of the Perron– Frobenius theorem. Linear and Multilinear Algebra, 1(1):81– 88, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiedler:1981:MPR

[Fie81] Miroslav Fiedler. Minimal polynomial and the rank of principal submatrices of a matrix. Linear and Multilinear Algebra, 10(2):85–88, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiedler:1984:CPR

[Fie84] Miroslav Fiedler. On a conjecture of P. R. Vein and its generalization. Linear and Multilinear Algebra, 16(1–4):147–154, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiedler:1990:AAC

[Fie90a] Miroslav Fiedler. Absolute algebraic connectivity of trees. Linear and Multilinear Algebra, 26(1–2):85–106, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiedler:1990:MPG

[Fie90b] Miroslav Fiedler. A minimax problem for graphs and its relation to generalized doubly stochastic matrices. Linear and Multilinear Algebra, 27(1):1–23, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiedler:1995:MPI

[Fie95] Miroslav Fiedler. Moore–Penrose involutions in the classes of Laplacians and simplices. Linear and Multilinear Algebra, 39(1–2):171–178, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Finston:1992:ADE

[Fin92] David R. Finston. Algebras and differential equations with additive symmetry. Linear and Multilinear Algebra, 31(1–4): 235–243, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furtado:2006:EPM

[FIS06] Susana Furtado, Laura Iglésias, and Fernando C. Silva. Eigenvalues of products of matrices. Linear Multilinear Algebra, 54(5):343–353, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fitzgerald:1994:CPS

[Fit94] Robert W. Fitzgerald. Characteristic polynomials of symmetric matrices. Linear and Multilinear Algebra, 36(4):233–237, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furtado:2001:SVU

[FJ01] Susana Furtado and Charles R. Johnson. Spectral variation under congruence. Linear and Multilinear Algebra, 49(3):243– 259, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furtado:2009:VJS

[FJ09] Susana Furtado and Charles R. Johnson. Variation in Jordan structure under congruence: the nonsingular case. Linear Multilinear Algebra, 57(1):29–54, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farber:2015:SSC

[FJ15] Miriam Farber and Charles R. Johnson. The structure of Schur complements in hollow, symmetric nonnegative matrices with two nonpositive eigenvalues. Linear Multilinear Algebra, 63(2): 423–438, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furtado:2022:PNM

[FJ22] Susana Furtado and Charles R. Johnson. Power normal matrices. Linear Multilinear Algebra, 70(20):5423–5432, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fujii:1998:CCO

[FJK98] Masatoshi Fujii, Jian Fei Jiang, and Eizaburo Kamei. Characterization of chaotic order and its applications to Furuta’s type operator inequalities. Linear and Multilinear Algebra, 43(4):

Fauser:2014:RHA

[FJK14] Bertfried Fauser, Peter D. Jarvis, and Ronald C. King. Ribbon Hopf algebras from group character rings. Linear Multilinear Algebra, 62(6):749–775, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ferjani:2022:SPG

[FJK22] Imen Ferjani, Aref Jeribi, and Bilel Krichen. Spectral properties for generalized weakly S-demicompact operators. Linear Multilinear Algebra, 70(4):606–624, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fallat:2000:EPM

[FJM00] Shaun M. Fallat, Charles R. Johnson, and Thomas L. Markham. Eigenvalues of products of matrices and submatrices in certain positivity classes. Linear and Multilinear Algebra, 47(3):235–248, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Feki:2014:RBE

[FJS14] Ines Feki, Aref Jeribi, and Ridha Sfaxi. On a Riesz basis of eigenvectors of a nonself-adjoint analytic operator and applications. Linear Multilinear Algebra, 62(8):1049–1068, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furtado:2021:LPC

[FJZ21] Susana Furtado, C. R. Johnson, and Yulin Zhang. Linear preservers of copositive matrices. Linear Multilinear Algebra, 69(10):1779–1788, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fellows:1991:SLT

[FK91] Michael R. Fellows and Paul A. Kaschube. Searching for K_{3,3} in linear time. Linear and Multilinear Algebra, 29(3–4):279– 290, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farooq:2019:IES

[FKC19] Rashid Farooq, Mehtab Khan, and Sarah Chand. On iota energy of signed digraphs. Linear Multilinear Algebra, 67(4): 705–724, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fosner:2011:MPM

[FKKS11] A. Fošner, B. Kuzma, T. Kuzma, and N.-S. Sze. Maps preserving matrix pairs with zero Jordan product. Linear Multilinear Algebra, 59(5):507–529, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fallat:2003:MAC

[FKP03] Shaun M. Fallat, Steve Kirkland, and Sukanta Pati. Maximizing algebraic connectivity over unicyclic graphs. Linear Multilinear Algebra, 51(3):221–241, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Foster:1993:DRH

[FKR93] J. Foster, K. O. Kidman, and F. B. Richards. Distributions of the ranks of Hadamard powers. Linear and Multilinear Algebra, 35(3–4):299–312, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farrugia:2022:CNG

[FKS22] Alexander Farrugia, Tamara Koledin, and Zoran Stanić. Controllability of NEPSes of graphs. Linear Multilinear Algebra, 70(10):1928–1941, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Facchini:2016:EMP

[FL16a] A. Facchini and A. Leroy. Elementary matrices and products of idempotents. Linear Multilinear Algebra, 64(10):1916–1935, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2016:RIP

[FL16b] Xiaohui Fu and Yang Liu. Rotfel’d inequality for partitioned matrices with numerical ranges in a sector. Linear Multilinear Algebra, 64(1):105–109, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:2018:DCM

[FL18] Yun Fan and Hualu Liu. Double circulant matrices. Linear Multilinear Algebra, 66(10):2119–2137, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Flanders:1974:ASM

[Fla74a] Harley Flanders. Analytic solutions of matrix equations. Linear and Multilinear Algebra, 2(3):241–243, 1974. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Flanders:1974:NSR

[Fla74b] Harley Flanders. On the norm and spectral radius. Linear and Multilinear Algebra, 2(3):239–240, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Flanders:1975:EPS

[Fla75] Harley Flanders. An extremal problem on the space of positive definite matrices. Linear and Multilinear Algebra, 3(1–2):33– 39, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2017:SME

[fLLL^{+}17] Jiao fen Li, Tao Li, Wen Li, Yan mei Chen, and Ru Huang. Solvability of matrix equations AX = B, XC = D, under semi-tensor product. Linear Multilinear Algebra, 65(9):1705– 1733, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2015:HAS

[fLLyP15] Jiao fen Li, Wen Li, and Zhen yun Peng. A hybrid algorithm for solving minimization problem over (R, S)-symmetric matrices with the matrix inequality constraint. Linear Multilinear Algebra, 63(5):1049–1072, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fillmore:1985:MSZ

[FLR85] P. Fillmore, C. Laurie, and H. Radjavi. On matrix spaces with zero determinant. Linear and Multilinear Algebra, 18(3): 255–266, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1988:ADN

[FLS88] Shmuel Friedland, Chi-Kwong Li, and Hans Schneider. Additive decomposition of nonnegative matrices with applications to permanents and scaling. Linear and Multilinear Algebra, 23 (1):63–78, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ferreyra:2020:CCE

[FLT20] D. E. Ferreyra, F. E. Levis, and N. Thome. Characterizations of k-commutative equalities for some outer generalized inverses. Linear Multilinear Algebra, 68(1):177–192, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2022:IBP

[FLT22] Xiaohui Fu, Pan-Shun Lau, and Tin-Yau Tam. Inequalities on 2 × 2 block positive semidefinite matrices. Linear Multilinear Algebra, 70(21):6820–6829, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Feng:2007:MLS

[FLZ07] Lihua Feng, Qiao Li, and Xiao-Dong Zhang. Minimizing the Laplacian spectral radius of trees with given matching number. Linear Multilinear Algebra, 55(2):199–207, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1978:MPD

[FM78] Shmuel Friedland and Henryk Minc. Monotonicity of permanents of doubly stochastic matrices. Linear and Multilinear Algebra, 6(3):227–231, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fard:2021:SRP

[FM21] Mohammad Ali Hasankhani Fard and Saeedeh Moazeni. Signal reconstruction without phase by norm retrievable frames. Linear Multilinear Algebra, 69(8):1484–1499, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ferreyra:2022:CSC

[FM22a] D. E. Ferreyra and Saroj B. Malik. Core and strongly core orthogonal matrices. Linear Multilinear Algebra, 70(20):5052–

5067, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ferreyra:2022:SNR

[FM22b] D. E. Ferreyra and Saroj B. Malik. Some new results on the core partial order. Linear Multilinear Algebra, 70(18):3449– 3465, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2022:THS

[FM22c] Houshan Fu and Zhousheng Mei. Truncated homogeneous symmetric functions. Linear Multilinear Algebra, 70(3):438– 448, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furuichi:2019:NSI

[FMS19] Shigeru Furuichi, Hamid Reza Moradi, and Mohammad Sababheh. New sharp inequalities for operator means. Linear Multilinear Algebra, 67(8):1567–1578, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Falcao:2019:NZU

[FMSS19] M. Irene Falcão, Fernando Miranda, Ricardo Severino, and M. Joana Soares. The number of zeros of unilateral polynomials over coquaternions revisited. Linear Multilinear Algebra, 67(6):1231–1249, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fang:2018:MRI

[FMX18] Xiaochun Fang, Mohammad Sal Moslehian, and Qingxiang Xu. On majorization and range inclusion of operators on Hilbert C∗-modules. Linear Multilinear Algebra, 66(12):2493– 2500, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2022:GPS

[FMXZ22] Chunhong Fu, Mohammad Sal Moslehian, Qingxiang Xu, and Ali Zamani. Generalized parallel sum of adjointable operators on Hilbert C^{∗}*-modules. Linear Multilinear Algebra, 70(12): 2278–2296, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Faryad:2022:RNR

[FMZ22] Elias Faryad, Mohammad Sal Moslehian, and Ali Zamani. Roberts numerical radius orthogonality. Linear Multilinear Algebra, 70(19):4282–4296, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ferrante:2014:RTD

[FN14] Augusto Ferrante and Lorenzo Ntogramatzidis. A reduction technique for discrete generalized algebraic and difference Riccati equations. Linear Multilinear Algebra, 62(11):1460–1474, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Felton:1996:GNR

[FNBR96] Anthony J. Felton, B. Nashvadia-Bakhsh, and H. P. Rogosinski. On generalised k-numerical radii of Schatten class operators. Linear and Multilinear Algebra, 40(3):249–258, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farid:2018:STS

[FNW18] Farid O. Farid, Xiang-Rong Nie, and Qing-Wen Wang. On the solutions of two systems of quaternion matrix equations. Linear Multilinear Algebra, 66(12):2355–2388, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Foata:1979:MAV

[Foa79] Dominique Foata. A matrix-analog for Viennot’s construction of the Robinson correspondence. Linear and Multilinear Algebra, 7(4):281–298, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fogarty:1973:ARI

[Fog73] John Fogarty. An algorithm for rings of invariants. Linear and Multilinear Algebra, 1(2):97–101, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fonseca:1989:MPI

[Fon89] Amélia Fonseca. On the multilinearity partition of an irreducible character II. Linear and Multilinear Algebra, 24(3): 191–198, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fonseca:1992:MPC

[Fon92] Amélia Fonseca. Multilinearity partitions of characters of embedded groups. Linear and Multilinear Algebra, 32(3–4):191– 196, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Foregger:1979:IRP

[For79a] Thomas H. Foregger. Identities related to permanents of doubly stochastic matrices and series reduces trees. Linear and Multilinear Algebra, 7(1):37–41, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Foregger:1979:NMC

[For79b] Thomas H. Foregger. A note on matrices with constant permanental minors. Linear and Multilinear Algebra, 7(2):127–128, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Foregger:1979:RCM

[For79c] Thomas H. Foregger. Remarks on a conjecture of M. Marcus and H. Minc. Linear and Multilinear Algebra, 7(2):123–126, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). The conjecture appears in Proc. Cambridge Philos. Soc. 63, 305–309 (1967).

Formanek:1979:CRG

[For79d] Edward Formanek. The center of the ring of 3 × 3 generic matrices. Linear and Multilinear Algebra, 7(3):203–212, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Foregger:1987:REL

[For87] Thomas H. Foregger. On the relative extrema of a linear combination of elementary symmetric functions. Linear and Multilinear Algebra, 20(4):377–385, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Foregger:1988:PCC

[For88] Thomas H. Foregger. Permanents of convex combinations of doubly stochastic matrices. Linear and Multilinear Algebra, 23 (1):79–90, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fosner:2005:APU

[Foš05a] Ajda Fošner. Automorphisms of the poset of upper triangular idempotent matrices. Linear Multilinear Algebra, 53(1):27–44, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fosner:2005:NLC

[Foš05b] Ajda Fošner. Non-linear commutativity preserving maps on M_{n}(). Linear Multilinear Algebra, 53(5):323–344, 2005. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ferreyra:2023:RWW

[FOT23] D. E. Ferreyra, V. Orquera, and N. Thome. Representations of the weighted WG inverse and a rank equation’s solution. Linear Multilinear Algebra, 71(2):226–241, 2023. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fowler:1980:ELF

[Fow80] G. O. Fowler. The existence of linear factors in the direct sum of forms and in the tensor product of forms. Linear and Multilinear Algebra, 8(4):317–322, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiori:2017:ELO

[FP17] Simone Fiori and Stilian Prifti. Exact low-order polynomial expressions to compute the Kolmogoroff–Nagumo mean in the affine symplectic group of optical transference matrices. Linear Multilinear Algebra, 65(4):840–856, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Franca:2013:CTM

[Fra13] Willian Franca. Commuting traces of multiadditive maps on invertible and singular matrices. Linear Multilinear Algebra, 61(11):1528–1535, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Franca:2017:WCM

[Fra17] Willian Franca. Weakly commuting maps on the set of rank-1 matrices. Linear Multilinear Algebra, 65(3):475–495, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Freedman:1976:FCM

[Fre76] H. I. Freedman. Functionally commutative matrices and matrices with constant eigenvectors. Linear and Multilinear Algebra, 4(2):107–113, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1975:GHI

[Fri75] Samuel Friedland. Generalised Hadamard inequality and its application. Linear and Multilinear Algebra, 2(4):327–333, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1978:SVW

[Fri78] Shmuel Friedland. A study of the van der Waerden conjecture and its generalizations. Linear and Multilinear Algebra, 6(2): 123–143, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1981:CSF

[Fri81] Shmuel Friedland. Convex spectral functions. Linear and Multilinear Algebra, 9(4):299–316, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1982:PGV

[Fri82a] Shmuel Friedland. A proof of a generalized van der Waerden conjecture on permanents. Linear and Multilinear Algebra, 11(2):107–120, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1982:VTP

[Fri82b] Shmuel Friedland. Variation of tensor powers and spectra. Linear and Multilinear Algebra, 12(2):81–98, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1984:SCS

[Fri84] Shmuel Friedland. Stable convex sets of matrices. Linear and Multilinear Algebra, 16(1–4):285–294, 1984. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1985:MMG

[Fri85] Shmuel Friedland. Maximality of the monomial group. Linear and Multilinear Algebra, 18(1):1–7, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1988:RP

[Fri88] Shmuel Friedland. Research problems. Linear and Multilinear Algebra, 23(1):91–93, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1990:RCF

[Fri90a] Shmuel Friedland. Rationality of certain field of invariant functions. Linear and Multilinear Algebra, 27(4):299–301, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1990:RVP

[Fri90b] Shmuel Friedland. A remark on the variation of permanents. Linear and Multilinear Algebra, 27(2):101–103, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1993:ESR

[Fri93] Shmuel Friedland. Estimates of the spectral radius of graphs. Linear and Multilinear Algebra, 34(1):33–40, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:1997:PDG

[Fri97] Shmuel Friedland. Properly discontinuous groups acting on certain matrix homogeneous spaces. Linear and Multilinear Algebra, 43(1–3):151–167, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Friedland:2012:NNS

[Fri12] Shmuel Friedland. A note on the nonzero spectra of irreducible matrices. Linear Multilinear Algebra, 60(11–12):1235–1238, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fiedler:1983:AFM

[FS83] Miroslav Fiedler and Hans Schneider. Analytic functions of M-matrices and generalizations. Linear and Multilinear Alge-

Foregger:1986:MMP

[FS86] Thomas H. Foregger and Richard Sinkhorn. On matrices minimizing the permanent on faces of the polyhedron of the doubly stochastic matrices. Linear and Multilinear Algebra, 19(4): 395–397, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furth:1989:REM

[FS89] Dave Furth and Gerard Sierksma. The rank and eigenvalues of main diagonal perturbed matrices. Linear and Multilinear Algebra, 25(3):191–204, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farrugia:2014:MEN

[FS14] Alexander Farrugia and Irene Sciriha. The main eigenvalues and number of walks in self-complementary graphs. Linear Multilinear Algebra, 62(10):1346–1360, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fujii:2015:ALM

[FS15] Jun Ichi Fujii and Yuki Seo. On the Ando–Li–Mathias mean and the Karcher mean of positive definite matrices. Linear Multilinear Algebra, 63(3):636–649, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farrugia:2018:TIN

[FS18a] Alexander Farrugia and Irene Sciriha. Triangles in inverse NSSD graphs. Linear Multilinear Algebra, 66(3):540–546, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fujimoto:2018:MWI

[FS18b] Masayuki Fujimoto and Yuki Seo. Matrix Wielandt inequality via the matrix geometric mean. Linear Multilinear Algebra, 66(8):1564–1577, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fujii:2019:USK

[FS19] Jun Ichi Fujii and Yuki Seo. The unique solution of the Karcher equation and the self-adjointness of the Karcher mean. Linear Multilinear Algebra, 67(5):976–986, 2019. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farooq:2022:SAG

[FS22] Aamir Farooq and Mahvish Samar. Sensitivity analysis for the generalized Cholesky block downdating problem. Linear Multilinear Algebra, 70(6):997–1022, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Felipe-Sosa:2014:CDP

[FSFSO^{+}14] Raúl Felipe-Sosa, Raúl Felipe, Juana Sánchez-Ortega, Murray R. Bremner, and Michael K. Kinyon. The Cayley–Dickson process for dialgebras. Linear Multilinear Algebra, 62(6):811– 830, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Foldes:2011:IDF

[FSvW11] Stephan Foldes, Jenő Szigeti, and Leon van Wyk. Invertibility and Dedekind finiteness in structural matrix rings. Linear Multilinear Algebra, 59(2):221–227, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2022:NEM

[FSWX22] Chunhong Fu, Chuanning Song, Guorong Wang, and Qingxiang Xu. Norm estimations for the Moore–Penrose inverse of the weak perturbation of matrices. Linear Multilinear Algebra, 70(2):215–237, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fujii:2016:NIM

[FSY16] Jun Ichi Fujii, Yuki Seo, and Takeaki Yamazaki. Norm inequalities for matrix geometric means of positive definite matrices. Linear Multilinear Algebra, 64(3):512–526, 2016. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:1987:GNR

[FT87] M. K. H. Fan and A. L. Tits. On the generalized numerical range. Linear and Multilinear Algebra, 21(3):313–320, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fallat:2021:SRO

[FT21] Shaun M. Fallat and Ryan M. Tifenbach. Sums of rank-one matrices and ranks of principal submatrices. Linear Multilinear Algebra, 69(1):9–18, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farazmanesh:2020:IAF

[FTA20] Davod Farazmanesh, Ali Tavakoli, and Abbas Askari. An iterative algorithm to find maximum spanning sets and minimum linearly independent sets which partition a finite generator. Linear Multilinear Algebra, 68(2):363–376, 2020. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fernandez-Ternero:2022:EOE

[FTGSNV22] Desamparados Fernández-Ternero, Víctor M. Gómez-Sousa, and Juan Núñez-Valdés. The evolution operator of evolution algebras. Linear Multilinear Algebra, 70(19):4377–4396, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fan:2008:MSR

[FTZ08] Yi-Zheng Fan, Bit-Shun Tam, and Jun Zhou. Maximizing spectral radius of unoriented Laplacian matrix over bicyclic graphs of a given order. Linear Multilinear Algebra, 56(4): 381–397, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fernandez-Unzueta:2020:SCB

[FU20] Maite Fernández-Unzueta. The Segrè cone of Banach spaces and multilinear mappings. Linear Multilinear Algebra, 68(3): 575–593, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fulton:1976:RQF

[Ful76] John D. Fulton. Representations by quadratic forms of arbitrary rank in a finite field of characteristic two. Linear and Multilinear Algebra, 4(2):89–101, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furuta:2008:APD

[Fur08] Takayuki Furuta. Applications of polar decompositions of idempotent and 2-nilpotent operators. Linear Multilinear Algebra, 56(1–2):69–79, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Furuichi:2011:ITR

[Fur11] Shigeru Furuichi. Inequalities for Tsallis relative entropy and generalized skew information. Linear Multilinear Algebra, 59 (10):1143–1158, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fallat:1997:MDM

[FV97] Shaun Fallat and P. Van Den Driessche. Maximum determinant of (0, 1) matrices with certain constant row and column sums. Linear and Multilinear Algebra, 42(4):303–318, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fillmore:1973:LAS

[FW73] Jay P. Fillmore and S. G. Williamson. A linear algebra setting for the Rota–Mullin theory of polynomials of binomial type. Linear and Multilinear Algebra, 1(1):67–80, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fillmore:1986:PDG

[FW86] Joy P. Fillmore and S. Gill Williamson. Permanents and determinants with generic noncommuting entries. Linear and Multilinear Algebra, 19(4):321–334, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fang:2007:PNP

[FW07] Maozhong Fang and Anqiang Wang. Possible numbers of positive entries of imprimitive nonnegative matrices. Linear Multilinear Algebra, 55(4):399–404, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ferrante:2013:RST

[FW13] Augusto Ferrante and Harald K. Wimmer. Roth’s similarity theorem and rank minimization in the presence of nonderogatory or semisimple eigenvalues. Linear Multilinear Algebra, 61(2):217–231, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Farid:2011:EQM

[FWZ11] F. O. Farid, Qing-Wen Wang, and Fuzhen Zhang. On the eigenvalues of quaternion matrices. Linear Multilinear Algebra, 59(4):451–473, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fallat:2012:SCN

[FX12] Shaun M. Fallat and YongJun Xing. On the spread of certain normal matrices. Linear Multilinear Algebra, 60(11–12):1391– 1407, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2022:CHP

[FXY22] Chunhong Fu, Qingxiang Xu, and Guanjie Yan. Characterizations of the harmonious pairs of projections on a Hilbert C^{∗}- module. Linear Multilinear Algebra, 70(21):6312–6320, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Feng:2013:SLS

[FYIS13] Lihua Feng, Guihai Yu, Aleksandar Ilić, and Dragan Stevanović. The signless Laplacian spectral radius of graphs on surfaces. Linear Multilinear Algebra, 61(5):573–581, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Faliva:2011:IFM

[FZ11] Mario Faliva and Maria Grazia Zoia. An inversion formula for a matrix polynomial about a (unit) root. Linear Multilinear Algebra, 59(5):541–556, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fei:2017:NGL

[FZ17] Xiuhai Fei and Jianhua Zhang. Nonlinear generalized Lie derivations on triangular algebras. Linear Multilinear Algebra, 65(6):1158–1170, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Fu:2021:OCC

[FZ21] Yanling Fu and Wei Zhang. Operator characterizations and constructions of continuous g-frames in Hilbert spaces. Linear

Gay:1985:CPI

[GA85] David Gay and Edgar Ascher. On computing the polynomial invariants of a finite group. Linear and Multilinear Algebra, 18 (2):91–116, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garg:2018:ISS

[GA18a] Isha Garg and Jaspal Singh Aujla. Inertia of some special matrices. Linear Multilinear Algebra, 66(3):602–607, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garg:2018:SSV

[GA18b] Isha Garg and Jaspal Singh Aujla. Some singular value inequalities. Linear Multilinear Algebra, 66(4):776–784, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garg:2020:INI

[GA20] Isha Garg and Jaspal Singh Aujla. Inertia of non-integer Hadamard powers of a non-negative matrix. Linear Multilinear Algebra, 68(2):410–416, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garcia-Ardila:2018:CGT

[GAGM18] Juan Carlos García-Ardila, Luis E. Garza, and Francisco Marcellán. A canonical Geronimus transformation for matrix orthogonal polynomials. Linear Multilinear Algebra, 66(2):357– 381, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gaines:1977:SGA

[Gai77] Fergus Gaines. Some generators for the algebra of n × n matrices. Linear and Multilinear Algebra, 5(2):95–98, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gallagher:1979:LTF

[Gal79] Vincent P. Gallagher. Local trace forms. Linear and Multilinear Algebra, 7(2):167–174, 1979. CODEN LNMLA2. ISSN 0308-1087 (print), 1563-5139 (electronic).

Galantai:2001:RRF

[Gal01] A. Galántai. Rank reduction, factorization and conjugation. Linear and Multilinear Algebra, 49(3):195–207, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Galantai:2003:PTM

[Gal03] A. Galántai. Perturbations of triangular matrix factorizations. Linear Multilinear Algebra, 51(2):175–198, 2003. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Galantai:2008:SAP

[Gal08] A. Galántai. Subspaces, angles and pairs of orthogonal projections. Linear Multilinear Algebra, 56(3):227–260, 2008. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gamas:2000:SFI

[Gam00] Carlos Gamas. Spherical functions and immanants. Linear and Multilinear Algebra, 47(2):151–173, 2000. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gamas:2018:CST

[Gam18] C. Gamas. Conditions for a symmetrized tensor associated with a spherical function to be zero. Linear Multilinear Algebra, 66(12):2534–2556, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghabries:2023:NLM

[GAMA23] Mohammad M. Ghabries, Hassane Abbas, Bassam Mourad, and Abdallah Assi. New log-majorization results concerning eigenvalues and singular values and a complement of a norm inequality. Linear Multilinear Algebra, 71(7):1228–1243, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garsia:1973:EMR

[Gar73] Adriano M. Garsia. An exposé of the Mullin–Rota theory of polynomials of binomial type. Linear and Multilinear Algebra, 1(1):47–65, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garbanati:1975:EHN

[Gar75a] Dennis A. Garbanati. Extensions of the Hasse norm theorem, II. Linear and Multilinear Algebra, 3(1–2):143–145, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gardner:1975:MAD

[Gar75b] Robert B. Gardner. Modules admitting determinants. Linear and Multilinear Algebra, 3(3):209–214, 1975. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garsia:1979:MIA

[Gar79] Adriano M. Garsia. On the “maj” and “inv” q-analogues of Eulerian polynomials. Linear and Multilinear Algebra, 8(1): 21–34, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garloff:1985:IEP

[Gar85] Jürgen Garloff. An inverse eigenvalue problem for totally nonnegative matrices. Linear and Multilinear Algebra, 17(1):19– 23, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garge:2021:MCR

[Gar21] Anuradha S. Garge. Matrices over commutative rings as sums of fifth and seventh powers of matrices. Linear Multilinear Algebra, 69(12):2220–2227, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garbouj:2023:NAR

[Gar23] Zied Garbouj. New additive results for the Drazin inverse of multivalued operators. Linear Multilinear Algebra, 71(5): 774–796, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gauger:1976:EDI

[Gau76] Michael A. Gauger. Elementary divisors of induced transformations on tensors via the representation theory of sl_{2}. Linear and Multilinear Algebra, 3(4):249–253, 1976. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gauger:1977:CFI

[GB77] Michael A. Gauger and Christopher I. Byrnes. Characteristic free, improved decidability criteria for the similarity problem. Linear and Multilinear Algebra, 5(3):153–158, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gover:1990:GBM

[GB90] M. J. C. Gover and S. Barnett. A generalized Bezoutian matrix. Linear and Multilinear Algebra, 27(1):33–48, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gungor:2011:DSR

[GB11] A. Dilek Güngör and S. Burcu Bozkurt. On the distance spectral radius and the distance energy of graphs. Linear Multilinear Algebra, 59(4):365–370, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ganie:2022:GAS

[GB22] Hilal A. Ganie and Maryam Baghipur. On the generalized adjacency spectral radius of digraphs. Linear Multilinear Algebra, 70(18):3497–3510, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guesba:2021:JHO

[GBM21] Messaoud Guesba, El Moctar Ould Beiba, and Sid Ahmed Ould Ahmed Mahmoud. Joint A-hyponormality of operators in semi-Hilbert spaces. Linear Multilinear Algebra, 69(15): 2888–2907, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2022:SOD

[GC22] Xunxiang Guo and Yonghong Chen. On some operators and dilations of frame generator and dual pair of frame generators of two structured unitary systems. Linear Multilinear Algebra, 70(5):830–842, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2020:RPW

[GCP20] Yuefeng Gao, Jianlong Chen, and Pedro Patrício. Representations and properties of the W -weighted core-EP inverse. Linear Multilinear Algebra, 68(6):1160–1174, 2020. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2021:CCE

[GCP21] Yuefeng Gao, Jianlong Chen, and Pedro Patrício. Continuity of the core-EP inverse and its applications. Linear Multilinear Algebra, 69(3):557–571, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garcia:2020:NDC

[GCSC20] Orlando García, Dairo Causil, José Sanabria, and Carlos Carpintero. New decompositions for the classes of quasi-Fredholm and semi B-Weyl operators. Linear Multilinear Algebra, 68(4):750–763, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Giambruno:2016:ISA

[GdSV16] Antonio Giambruno, Rafael Bezerra dos Santos, and Ana Cristina Vieira. Identities of *-superalgebras and almost polynomial growth. Linear Multilinear Algebra, 64(3):484–501, 2016. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geher:2015:CCA

[Geh15] György Pál Gehér. Characterization of cesáro and Lasymptotic limits of matrices. Linear Multilinear Algebra, 63(4):788–805, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geissinger:1978:CPM

[Gei78] Ladnor Geissinger. Critical points of a modified trace. Linear and Multilinear Algebra, 5(4):321–323, 1978. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gerstein:1974:MPI

[Ger74] Larry J. Gerstein. A multiplicative property of invariant factors. Linear and Multilinear Algebra, 2(2):141–142, 1974. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gerstein:1982:PSG

[Ger82] Larry J. Gerstein. On the proper spinor genus of a quadratic form. Linear and Multilinear Algebra, 11(2):203–208, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gerth:1986:LPC

[Ger86] Frank Gerth III. Limit probabilities for coranks of matrices over GF(q). Linear and Multilinear Algebra, 19(1):79–93, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gerstein:2004:CEU

[Ger04] Larry J. Gerstein. Characteristic elements of unimodular Zlattices. Linear Multilinear Algebra, 52(5):381–383, 2004. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gong:2010:PME

[GF10] Shi-Cai Gong and Yi-Zheng Fan. The property of maximal eigenvectors of trees. Linear Multilinear Algebra, 58(1):105– 111, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gargate:2022:EUT

[GG22a] Ivan Gargate and Michael Gargate. Expressing upper triangular matrices as products of commutators of finite order elements. Linear Multilinear Algebra, 70(20):5571–5579, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goncharov:2022:RBO

[GG22b] Maxim Goncharov and Vsevolod Gubarev. Rota–Baxter operators of nonzero weight on the matrix algebra of order three. Linear Multilinear Algebra, 70(6):1055–1080, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garayev:2021:SPO

[GGGA21] M. T. Garayev, H. Guediri, M. Gürdal, and G. M. Alsahli. On some problems for operators on the reproducing kernel Hilbert space. Linear Multilinear Algebra, 69(11):2059–2077, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gan:2022:ZNT

[GGH22] Luyining Gan, Wei Gao, and Jie Han. Zero-nonzero tree patterns that allow sn*. Linear Multilinear Algebra, 70(22):7347– 7369, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:2003:SSI

[GGL03] Moshe Goldberg, Robert Guralnick, and W. A. J. Luxemburg. Stable subnorms II. Linear Multilinear Algebra, 51(2):209– 219, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geramita:1976:OD

[GGW76] Anthony V. Geramita, Joan Murphy Geramita, and Jennifer Seberry Wallis. Orthogonal designs. Linear and Multilinear Algebra, 3(4):281–306, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gonzalez:1998:MIM

[GH98] Raul Ernesto Gonzalez and D. J. Hartfiel. Manifolds of idempotent matrices. Linear and Multilinear Algebra, 44(3):187– 193, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gong:2022:TGH

[GH22] Shi-Cai Gong and Yao-Ping Hou. On tetracyclic graphs having minimum energies. Linear Multilinear Algebra, 70(21):6265– 6278, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghosh:2023:ASW

[GH23] Gargi Ghosh and Somnath Hazra. On analytic structure of weighted shifts on generalized directed semi-trees. Linear Multilinear Algebra, 71(9):1545–1564, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghahramani:2012:AMD

[Gha12] Hoger Ghahramani. Additive mappings derivable at nontrivial idempotents on Banach algebras. Linear Multilinear Algebra, 60(6):725–742, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghahramani:2013:ZPD

[Gha13] Hoger Ghahramani. Zero product determined triangular algebras. Linear Multilinear Algebra, 61(6):741–757, 2013. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gharakhanlu:2021:NII

[Gha21] Nahid Gharakhanlu. Norm inequalities involving operator monotone functions. Linear Multilinear Algebra, 69(15):2882– 2887, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghimire:2022:LLC

[Ghi22] Prakash Ghimire. Linear Lie centralizers of the algebra of dominant block upper triangular matrices. Linear Multilinear Algebra, 70(20):5040–5051, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gumus:2016:ERI

[GHK16] Ibrahim Halil Gumus, Omar Hirzallah, and Fuad Kittaneh. Estimates for the real and imaginary parts of the eigenvalues of matrices and applications. Linear Multilinear Algebra, 64 (12):2431–2445, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geng:2010:FRP

[GHL10] Xianya Geng, Xia Hu, and Shuchao Li. Further results on permanental bounds for the Laplacian matrix of trees. Linear Multilinear Algebra, 58(5):571–587, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2016:NDI

[GHL16] Xiangqian Guo, Fengdan Hou, and Xuewen Liu. Natural density of integral matrices that can be extended to invertible integral matrices. Linear Multilinear Algebra, 64(9):1878–1886,

Gao:2020:SSM

[GHL20a] Lei Gao, Hui Huang, and Chaoqian Li. Subdirect sums of QN-matrices. Linear Multilinear Algebra, 68(8):1605–1623, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gu:2020:HTO

[GHL20b] Caixing Gu, In Sung Hwang, and Woo Young Lee. Hyponormality of Toeplitz operators in several variables by the weighted shifts approach. Linear Multilinear Algebra, 68(8): 1695–1720, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghorbani:2019:SPC

[Gho19] Ebrahim Ghorbani. Spectral properties of cographs and P_{5}- free graphs. Linear Multilinear Algebra, 67(8):1701–1710, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gumus:2011:SVI

[GHT11] Ibrahim Halil Gumus, Omar Hirzallah, and Necati Taskara. Singular value inequalities for the arithmetic, geometric and Heinz means of matrices. Linear Multilinear Algebra, 59(12): 1383–1392, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guterman:2016:NMP

[GHT16] Alexander Guterman, Alicia Herrero, and Néstor Thome. New matrix partial order based on spectrally orthogonal matrix decomposition. Linear Multilinear Algebra, 64(3):362–374, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gracia:1999:SNI

[GHV99] Juan M. Gracia, Inmaculadade Hoyos, and Franscisco E. Velasco. Safety neighbourhoods for the invariants of the matrix similarity. Linear and Multilinear Algebra, 46(1–2):25–49, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gibson:1976:PSD

[Gib76] P. M. Gibson. Products of symmetric doubly stochastic matrices over a field. Linear and Multilinear Algebra, 4(1):45–54, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gibson:1987:DCA

[Gib87] Peter M. Gibson. Disproof of a conjecture on assignment functions. Linear and Multilinear Algebra, 21(1):87–89, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gibson:1989:IMC

[Gib89] Peter M. Gibson. On an iterative method for consistent linear systems. Linear and Multilinear Algebra, 25(3):215–218, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gibson:1992:SCS

[Gib92] Peter M. Gibson. Stochastic complements in substochastic matrices. Linear and Multilinear Algebra, 32(2):115–130, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gil:1993:ENM

[Gil93] Michael Gil. Estimate for the norm of matrix-valued functions. Linear and Multilinear Algebra, 35(1):65–73, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gil:2012:MFN

[Gil12] Michael Gil’. Matrix functions nonregular on the convex hull of the spectrum. Linear Multilinear Algebra, 60(4):465–473, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gil:2014:NIR

[Gil14] Michael Gil’. A new identity for resolvents of matrices. Linear Multilinear Algebra, 62(6):715–720, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gil:2016:ROT

[Gil16] Michael Gil’. Resolvents of operators on tensor products of Euclidean spaces. Linear Multilinear Algebra, 64(4):699–716, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gil:2017:SBF

[Gil17] Michael Gil’. A sharp bound for the Frobenius norm of selfcommutators of matrices. Linear Multilinear Algebra, 65(11): 2333–2339, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gowda:2022:PWM

[GJ22] M. Seetharama Gowda and J. Jeong. A pointwise weakmajorization inequality for linear maps over Euclidean Jordan algebras. Linear Multilinear Algebra, 70(19):4157–4176, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1984:IHI

[GJMW84] Bob Grone, Charles R. Johnson, Eduardo Marques De Sá, and Henry Wolkowicz. Improving Hadamard’s inequality. Linear and Multilinear Algebra, 16(1–4):305–322, 1984. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03081088408817634.

Grone:1986:NMP

[GJMW86] Robert Grone, Charles R. Johnson, Eduardo Marques de Sá, and Henry Wolkowicz. A note on maximizing the permanent of a positive definite Hermitian matrix, given the eigenvalues. Linear and Multilinear Algebra, 19(4):389–393, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gelonch:1997:EFT

[GJR97] José Gelonch, Charles R. Johnson, and Pere Rubić. An extension of Flanders theorem to several matrices. Linear and Multilinear Algebra, 43(1–3):181–200, 1997. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghribi:2022:ICT

[GJR22] Salima Ghribi, Nader Jeridi, and Rchid Rabaoui. On (A, m)- isometric commuting tuples of operators on a Hilbert space.

Griffin:1977:MSS

[GK77] Malcolm Griffin and Mark Krusemeyer. Matrices as sums of squares. Linear and Multilinear Algebra, 5(1):33–44, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Georgiou:2002:NIC

[GK02] S. Georgiou and C. Koukouvinos. New infinite classes of orthogonal designs. Linear Multilinear Algebra, 50(4):293–300, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gouveia:2018:FCI

[GK18] Maria Celeste Gouveia and Alexander Kovacec. A formula for computing (1, 2, 3)-inverses of g-circulants. Linear Multilinear Algebra, 66(12):2487–2492, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gohberg:1982:CMC

[GKLR82] I. Gohberg, M. A. Kaashoek, L. Lerer, and L. Rodman. Common multiples and common divisors of matrix polynomials, II. Vandermonde and resultant matrices. Linear and Multilinear Algebra, 12(3):159–203, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberger:2022:TCI

[GKO22] Assaf Goldberger, Shujie Kang, and Kasso A. Okoudjou. Towards a classification of incomplete Gabor POVMs in c^{d}. Linear Multilinear Algebra, 70(22):7536–7557, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Georgiou:2001:CBM

[GKS01] S. Georgiou, C. Koukouvinos, and Jennifer Seberry. On circulant best matrices and their applications. Linear and Multilinear Algebra, 48(3):263–274, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldwasser:1997:CSS

[GKT97] John Goldwasser, William Klostermeyer, and George Trapp. Characterizing switch-setting problems. Linear and Multilinear Algebra, 43(1–3):121–135, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gohberg:1989:ECS

[GKV89] I. Gohberg, M. A. Kaashoek, and F. Van Schagen. Eigenvalues of completions of submatrices. Linear and Multilinear Algebra, 25(1):55–70, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guralnick:1983:RC

[GL83] Robert Guralnick and Charles Lanski. The rank of a commutator. Linear and Multilinear Algebra, 13(2):167–175, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geraghty:1985:MTL

[GL85] M. A. Geraghty and Bor-Luh Lin. Minimax theorems without linear structure. Linear and Multilinear Algebra, 17(3–4):171– 180, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guralnick:1997:IPAb

[GL97] Robert M. Guralnick and Chi-Kwong Li. Invertible preservers and algebraic groups III: preservers of unitary similarity (congruence) invariants and overgroups of some unitary subgroups. Linear and Multilinear Algebra, 43(1–3):257–282, 1997. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:2001:DS

[GL01] Moshe Goldberg and W. A. J. Luxemburg. Discontinuous subnorms. Linear and Multilinear Algebra, 49(1):1–24, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geng:2011:SRT

[GL11] Xianya Geng and Shuchao Li. On the spectral radius of tricyclic graphs with a fixed diameter. Linear Multilinear Algebra, 59(1):41–56, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2016:ATG

[GL16] Jun Guo and Fenggao Li. Anzahl theorems in geometry of t-singular classical groups and their applications. Linear Multilinear Algebra, 64(8):1617–1636, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garcia:2022:NSP

[GLdS22] Esther García, Miguel Gómez Lozano, and Guillermo Vera de Salas. Nilpotent superderivations in prime superalgebras. Linear Multilinear Algebra, 70(20):5547–5570, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2018:IEB

[GLL18] Lei Gao, Chaoqian Li, and Yaotang Li. An improvement of the error bounds for linear complementarity problems of Nekrasov matrices. Linear Multilinear Algebra, 66(8):1505–1519, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2021:CAD

[GLL21] Qianping Guo, Jingsong Leng, and Houbiao Li. Construct approximate dual g-frames in Hilbert spaces. Linear Multilinear Algebra, 69(2):245–258, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2021:EHT

[GLLM21] Ziliang Guo, Shuchao Li, Xin Liu, and Xiaoling Mei. Expected hitting times for random walks on the diamond hierarchical graphs involving some classical parameters. Linear Multilinear Algebra, 69(10):1841–1857, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grotheer:2022:IHT

[GLM^{+}22] R. Grotheer, S. Li, A. Ma, D. Needell, and J. Qin. Iterative hard thresholding for low CP-rank tensor models. Linear Multilinear Algebra, 70(22):7452–7468, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gago:2013:SRF

[GLOT13] F. Gago, M. Ladra, B. A. Omirov, and R. M. Turdibaev. Some radicals, Frattini and Cartan subalgebras of Leibniz nalgebras. Linear Multilinear Algebra, 61(11):1510–1527, 2013.

Grone:1992:NPS

[GLP92] Robert Grone, Raphael Loewy, and Stephen Pierce. Nonchordal positive semidefinite stochastic matrices. Linear and Multilinear Algebra, 32(2):107–113, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gohberg:1982:FSM

[GLR82] I. Gohberg, P. Lancaster, and L. Rodman. Factorization of selfadjoint matrix polynomials with constant signature. Linear and Multilinear Algebra, 11(3):209–224, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guterman:2013:ECM

[GLS13] Alexander Guterman, Rute Lemos, and Graça Soares. Extremal case in Marcus–Oliveira conjecture and beyond. Linear Multilinear Algebra, 61(9):1206–1222, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2016:ENL

[GLS16] Ji-Ming Guo, Jianxi Li, and Wai Chee Shiu. Effects on the normalized Laplacian spectral radius of non-bipartite graphs under perturbation and their applications. Linear Multilinear Algebra, 64(11):2177–2187, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2022:SOI

[GLX22] Shuhui Gao, Yuan Li, and Xiaoming Xu. The star order for idempotent operators. Linear Multilinear Algebra, 70(19): 4568–4583, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1987:HIT

[GM87a] Robert Grone and Russell Merris. A Hadamard inequality for the third and fourth immanants. Linear and Multilinear Algebra, 21(3):201–209, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1987:RP

[GM87b] Robert Grone and Russell Merris. Research problems. Linear and Multilinear Algebra, 21(4):419–427, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1990:CME

[GM90] Robert Grone and Russell Merris. Coalescence, majorization, edge valuations and the Laplacian spectra of graphs. Linear and Multilinear Algebra, 27(2):139–146, 1990. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guterman:2018:MPS

[GM18] A. E. Guterman and A. M. Maksaev. Maps preserving scrambling index. Linear Multilinear Algebra, 66(4):840–851, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guterman:2021:UBS

[GM21] A. E. Guterman and A. M. Maksaev. Upper bounds on scrambling index for non-primitive digraphs. Linear Multilinear Algebra, 69(11):2143–2168, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goncalves:1985:VPM

[GMD85] A. D. S. Goncalves, J. F. C. Machado, and G. N. De Oliveira. Vectors with a prescribed minimal polynomial. Linear and Multilinear Algebra, 17(3–4):227–234, 1985. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gustafson:1981:MCD

[GMR81] William H. Gustafson, Marion E. Moore, and Irving Reiner. Matrix completions over Dedekind rings. Linear and Multilinear Algebra, 10(2):141–144, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gerstein:1976:BR

[GMS^{+}76] Larry J. Gerstein, H. Mochizuki, George P. H. Styan, D. Ž. – Doković, Alston Householder, Alan J. Hoffman, and C. J. Maxson. Book reviews. Linear and Multilinear Algebra, 3(4): 313–325, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Giambruno:2018:CGG

[GMV18] A. Giambruno, C. Polcino Milies, and A. Valenti. Cocharacters of group graded algebras and multiplicities bounded by one. Linear Multilinear Algebra, 66(8):1709–1715, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Galuppi:2023:STW

[GMV23] Francesco Galuppi, Raffaella Mulas, and Lorenzo Venturello. Spectral theory of weighted hypergraphs via tensors. Linear Multilinear Algebra, 71(3):317–347, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1986:HDT

[GMW86] Robert Grone, Russell Merris, and William Watkins. A Hadamard dominance theorem for a class of immanants. Linear and Multilinear Algebra, 19(2):167–171, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grishkov:2011:DCM

[GMZ11] A. N. Grishkov, F. Marko, and A. N. Zubkov. Description of costandard modules for Schur superalgebra S (2—1) in positive characteristic. Linear Multilinear Algebra, 59(1):57–64, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2017:SDL

[GMZ17] Chenghua Gao, Ruyun Ma, and Fei Zhang. Spectrum of discrete left definite Sturm–Liouville problems with eigenparameter-dependent boundary conditions. Linear Multilinear Algebra, 65(9):1905–1923, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberger:2002:HPC

[GN02] Assaf Goldberger and Michael Neumann. The Hadamard product of circulant matrices on three symbols that are inverses of M-matrices. Linear Multilinear Algebra, 50(2):143– 149, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ge:2020:SRH

[GN20] Jun Ge and Bo Ning. Spectral radius and Hamiltonian properties of graphs, II. Linear Multilinear Algebra, 68(11):2298–

Ghobadzadeh:2020:MFI

[GNO20] Fatemeh Ghobadzadeh, A. Najati, and Elnaz Osgooei. Modular frames and invertibility of multipliers in Hilbert C∗- modules. Linear Multilinear Algebra, 68(8):1568–1584, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gogic:2021:CNA

[Gog21] Ilja Gogić. The cb-norm approximation of generalized skew derivations by elementary operators. Linear Multilinear Algebra, 69(8):1543–1556, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:1979:CFD

[Gol79] Moshe Goldberg. On certain finite dimensional numerical ranges and numerical radii. Linear and Multilinear Algebra, 7(4):329–342, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:1983:M

[Gol83] Moshe Goldberg. On the mapping A → A^{+}. Linear and Multilinear Algebra, 12(4):285–289, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:1987:ECN

[Gol87] Moshe Goldberg. Equivalence constants for l_{p} norms of matrices. Linear and Multilinear Algebra, 21(2):173–179, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:1989:NMS

[Gol89] Moshe Goldberg. A note on monotonic and semi-monotonic matrix functions. Linear and Multilinear Algebra, 24(3):223– 226, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:1990:QFM

[Gol90] Moshe Goldberg. Quasimonotonic functions on c^{n} and the mapping f → f^{+}. Linear and Multilinear Algebra, 27(1):63–

Goldwasser:1992:MPD

[Gol92] John Goldwasser. Monotonicity of permanents of direct sums of doubly stochastic matrices. Linear and Multilinear Algebra, 33(3–4):185–188, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:2006:QSR

[Gol06] Felix Goldberg. On quasi-strongly regular graphs. Linear Multilinear Algebra, 54(6):437–451, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:2007:RSF

[Gol07] Moshe Goldberg. Radii and subnorms on finite-dimensional power-associative algebras. Linear Multilinear Algebra, 55(5): 405–415, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Golinskii:2016:SIG

[Gol16] L. Golinskii. Spectra of infinite graphs with tails. Linear Multilinear Algebra, 64(11):2270–2296, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Golinskii:2021:SCG

[Gol21] L. Golinskii. Spectra of comb graphs with tails. Linear Multilinear Algebra, 69(15):2850–2864, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gong:2017:RRS

[Gon17] Shi-Cai Gong. On the rank of a real skew symmetric matrix described by an oriented graph. Linear Multilinear Algebra, 65(10):1934–1946, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gordon:1983:AGA

[Gor83] Jean Gordon. Applications of the Grassmann algebra to coding theory. Linear and Multilinear Algebra, 14(2):165–177, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gora:2022:RSS

[Gór22] Michał Góra. Results on the Schur stability of the Hadamard powers and the Hadamard product of complex polynomials. Linear Multilinear Algebra, 70(6):1047–1054, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gerberding:2022:ILC

[GOS22] Seth Gerberding, Nida Obatake, and Anne Shiu. Identifiability of linear compartmental models: the effect of moving inputs, outputs, and leaks. Linear Multilinear Algebra, 70(14):2782– 2803, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gouveia:2013:TPC

[Gou13] M. C. Gouveia. On the Toeplitz pencil conjecture. Linear Multilinear Algebra, 61(6):784–790, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gow:1980:EIM

[Gow80] Roderick Gow. The equivalence of an invertible matrix to its transpose. Linear and Multilinear Algebra, 8(4):329–336, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gowda:2022:SSD

[Gow22] M. Seetharama Gowda. Simultaneous spectral decomposition in Euclidean Jordan algebras and related systems. Linear Multilinear Algebra, 70(21):6535–6547, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geramita:1974:RFH

[GP74] A. V. Geramita and N. J. Pullman. Radon’s function and Hadamard arrays. Linear and Multilinear Algebra, 2(2):147– 150, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1996:DNN

[GP96] Robert Grone and Stephen Pierce. Decomposably nonnegative matrices. Linear and Multilinear Algebra, 41(1):63– 79, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garcia-Planas:2006:SPL

[GP06] María Isabel García-Planas. Similarity of pairs of linear maps defined modulo a subspace. Linear Multilinear Algebra, 54(5): 379–389, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garces:2014:OFO

[GP14] Jorge J. Garcés and Antonio M. Peralta. Orthogonal forms and orthogonality preservers on real function algebras. Linear Multilinear Algebra, 62(3):275–296, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Giuzzi:2015:SSG

[GP15] L. Giuzzi and V. Pepe. On some subvarieties of the Grassmann variety. Linear Multilinear Algebra, 63(11):2121–2134, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garcia-Pacheco:2022:CFT

[GP22a] F. J. García-Pacheco. Convex functions on topological modules. Linear Multilinear Algebra, 70(18):3511–3524, 2022. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghatak:2022:RNT

[GP22b] Anindya Ghatak and Santhosh Kumar Pamula. A Radon– Nikodým theorem for local completely positive invariant multilinear maps. Linear Multilinear Algebra, 70(21):7115–7141, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gregory:1994:DAG

[GPK94] David A. Gregory, Norman J. Pullman, and Stephen J. Kirkland. On the dimension of the algebra generated by a boolean matrix. Linear and Multilinear Algebra, 38(1–2):131–144, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garcia-Planas:1998:GSE

[GPM98] Ma. Isabel Garciá-Planas and Ma. Dolors Magret. A generalized Sylvester equation: a criterion for structural stability of triples of matrices. Linear and Multilinear Algebra, 44(2):

Ghanbari:2014:CJM

[GPM14] K. Ghanbari, F. Parvizpour, and H. Mirzaei. Constructing Jacobi matrices using prescribed mixed eigendata. Linear Multilinear Algebra, 62(6):721–734, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grunenfelder:1984:CTR

[GPR84] L. Grunenfelder, R. Paré, and H. Radjavi. On a commutator theorem of Robert C. Thompson. Linear and Multilinear Algebra, 16(1–4):129–131, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2022:DNW

[GPW22] Shoulan Gao, Yufeng Pei, and Haiying Wang. Derivations of the nongraded Witt algebra with coefficients in tensor modules. Linear Multilinear Algebra, 70(3):465–489, 2022. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gow:2006:VSA

[GQ06] Rod Gow and Rachel Quinlan. On the vanishing of subspaces of alternating bilinear forms. Linear Multilinear Algebra, 54(6):415–428, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Garsia:1981:SFR

[GR81] A. M. Garsia and J. Remmel. Symmetric functions and raising operators. Linear and Multilinear Algebra, 10(1):15–43, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gohberg:1991:UBT

[GR91] Israel Gohberg and Boris Reichstein. On h-unitary and block-Toeplitz h-normal operators. Linear and Multilinear Algebra, 30(1–2):17–48, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Giarrusso:1993:PFC

[GR93a] Dante M. Giarrusso and Joseph D. Romano. A polarization formula for Clifford modules. Linear and Multilinear Algebra, 35(3–4):191–194, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gohberg:1993:CBT

[GR93b] Israel Gohberg and Boris Reichstein. Classification of block-Toeplitz h-normal operators. Linear and Multilinear Algebra, 34(3–4):213–245, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Greiner:1993:JSC

[GR93c] G. Greiner and W. J. Ricker. Joint spectra and commutativity of systems of (2 × 2) selfadjoint matrices. Linear and Multilinear Algebra, 36(1):47–58, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goze:2010:AAT

[GR10] Nicolas Goze and Elisabeth Remm. The n-ary algebra of tensors and of cubic and hypercubic matrices. Linear Multilinear Algebra, 58(8):993–1005, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gonzalez:2014:NAQ

[GR14] Julia Ramos González and Fernando Pablos Romo. A negative answer to the question of the linearity of Tate’s trace for the sum of two endomorphisms. Linear Multilinear Algebra, 62(4): 548–552, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ganesh:2023:DMA

[GR23] Jadav Ganesh and B. Madhav Reddy. On the denseness of minimum attaining operator-valued functions. Linear Multilinear Algebra, 71(2):190–205, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gracia:1980:DSS

[Gra80] Juan M. Gracia. Dimension of the solution spaces of the matrix equations [A, [A, X]] = 0 and [A, [A, [A, X]]] = 0. Linear and Multilinear Algebra, 9(3):195–200, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Granville:1987:MSF

[Gra87] Andrew J. Granville. Matrices as the sum of four squares. Linear and Multilinear Algebra, 20(3):247–251, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grady:1993:RP

[Gra93] Michael Grady. Research problem. Linear and Multilinear Algebra, 35(2):179–183, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grasic:2010:ZPD

[Gra10] Mateja Grasic. Zero product determined classical Lie algebras. Linear Multilinear Algebra, 58(8):1007–1022, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grasic:2011:ZPD

[Gra11] Mateja Grasic. Zero product determined Jordan algebras, I. Linear Multilinear Algebra, 59(6):671–685, 2011. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Greenhill:1999:ARE

[Gre99] Catherine Greenhill. An algorithm for recognising the exterior square of a matrix. Linear and Multilinear Algebra, 46(3): 213–244, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Golpar-Raboky:2013:EIR

[GRMA13] Effat Golpar-Raboky and Nezam Mahdavi-Amiri. Extended integer rank reduction formulas and Smith normal form. Linear Multilinear Algebra, 61(12):1641–1659, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1985:ISI

[Gro85] Bob Grone. An inequality for the second immanant. Linear and Multilinear Algebra, 18(2):147–152, 1985. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1995:EDS

[Gro95] Robert D. Grone. Eigenvalues and the degree sequences of graphs. Linear and Multilinear Algebra, 39(1–2):133–136, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grob:1996:CRI

[Gro96] Jürgen Grob. Comment on range invariance of matrix products. Linear and Multilinear Algebra, 41(2):157–160, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grob:1997:SRP

[Gro97] Jürgen Grob. Some remarks on partial orderings of Hermitian matrices. Linear and Multilinear Algebra, 42(1):53–60, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grob:1999:OPR

[Gro99] Jürgen Grob. On oblique projection, rank additivity and the Moore–Penrose inverse of the sum of two matrices. Linear and Multilinear Algebra, 46(4):265–275, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grover:2017:OMK

[Gro17] Priyanka Grover. Orthogonality of matrices in the Ky Fan k-norms. Linear Multilinear Algebra, 65(3):496–509, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gohberg:1992:BES

[GRSW92] Israel Gohberg, Leiba Rodman, Tamir Shalom, and Hugo J. Woerdeman. Bounds for eigenvalues and singular values of matrix completions. Linear and Multilinear Algebra, 33(3–4): 233–249, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grubaum:1976:SUE

[Grü76] F. Alberto Grübaum. Simultaneous unitary equivalence. Linear and Multilinear Algebra, 4(2):79–84, 1976. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grzaslewicz:1989:ES

[Grz89] Ryszard Grzaślewicz. Extremal structure of L(l_{m}^{2}, l_{n}^{p}). Linear and Multilinear Algebra, 24(2):117–125, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grzaslewicz:1991:PC

[Grz91] Ryszard Grzaślewicz. Positive contractions in L(l^{p}, l^{r}). Linear and Multilinear Algebra, 30(1–2):135–144, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:1977:TM

[GS77] Moshe Goldberg and E. G. Straus. On a theorem by Mirman. Linear and Multilinear Algebra, 5(2):77–78, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:1983:GPF

[GS83] Moshe Goldberg and E. G. Staraus. On generalizations of the Perron–Frobenius theorem. Linear and Multilinear Algebra, 14(2):143–156, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gibson:1995:CLA

[GS95] Gregory Gibson and Ernest Stitzinger. Complete Lie algebras. Linear and Multilinear Algebra, 40(1):29–32, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2001:IAP

[GS01] Yubin Gao and Yanling Shao. Inertially arbitrary patterns. Linear and Multilinear Algebra, 49(2):161–168, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2006:ISS

[GS06] Yubin Gao and Yanling Shao. Inertia sets of symmetric 2generalized star sign patterns. Linear Multilinear Algebra, 54 (1):27–35, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gorodski:2019:FRO

[GS19] Claudio Gorodski and Artur B. Saturnino. Focal radii of orbits. Linear Multilinear Algebra, 67(10):2082–2103, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ghahramani:2020:IST

[GS20a] H. Ghahramani and S. Sattari. Ideal submodules and topologically complementable submodules of Hilbert C∗-modules.

Gonzalez:2020:SPB

[GS20b] Luis González and Antonio Suárez. Spectral properties of the binary Hadamard matrices. Linear Multilinear Algebra, 68(1): 113–132, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geryba:2021:SMB

[GS21a] Titas Geryba and Ilya M. Spitkovsky. On some 4-by-4 matrices with bi-elliptical numerical ranges. Linear Multilinear Algebra, 69(5):855–870, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geryba:2021:NRS

[GS21b] Titas Geryba and Ilya M. Spitkovsky. On the numerical range of some block matrices with scalar diagonal blocks. Linear Multilinear Algebra, 69(5):772–785, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2009:BLB

[GSS09] Yubin Gao, Yanling Shao, and Jian Shen. Bounds on the local bases of primitive nonpowerful nearly reducible sign patterns. Linear Multilinear Algebra, 57(2):205–215, 2009. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grinberg:2022:PTC

[GSZ22] Darij Grinberg, Zhi-Wei Sun, and Lilu Zhao. Proof of three conjectures on determinants related to quadratic residues. Linear Multilinear Algebra, 70(19):3734–3746, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guralnick:1979:RCU

[GT79] R. Guralnick and Olga Taussky. A remark concerning unipotent matrix groups. Linear and Multilinear Algebra, 7(1):87– 89, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grob:1998:POP

[GT98] Jürgen Grob and Götz Trenkler. On the product of oblique projectors. Linear and Multilinear Algebra, 44(3):247–259,

Grassmann:2008:CSM

[GT08] Winfried K. Grassmann and Javad Tavakoli. The continuous spectrum for the M/M/c queue. Linear Multilinear Algebra, 56(3):319–331, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gowda:2011:CIT

[GT11] M. Seetharama Gowda and J. Tao. The Cauchy interlacing theorem in simple Euclidean Jordan algebras and some consequences. Linear Multilinear Algebra, 59(1):65–86, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2014:ISP

[GT14] Niushan Gao and Vladimir G. Troitsky. Irreducible semigroups of positive operators on Banach lattices. Linear Multilinear Algebra, 62(1):74–95, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gomez-Torrecillas:2018:PGZ

[GTLN18] José Gómez-Torrecillas, F. J. Lobillo, and Gabriel Navarro. Peterson–Gorenstein–Zierler algorithm for skew RS codes. Linear Multilinear Algebra, 66(3):469–487, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grubisic:2014:REP

[GTM14] Luka Grubisić, Ninoslav Truhar, and Suzana Miodragović. The rotation of eigenspaces of perturbed matrix pairs II. Linear Multilinear Algebra, 62(8):1010–1031, 2014. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Goldberg:1975:NRS

[GTZ75] Moshe Goldberg, Eitan Tadmor, and Gideon Zwas. The numerical radius and spectral matrices. Linear and Multilinear Algebra, 2(4):317–326, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guerry:2017:NEC

[Gue17] Marie-Anne Guerry. Necessary embedding conditions for state-wise monotone Markov chains. Linear Multilinear Algebra, 65(8):1529–1539, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guerry:2019:SEC

[Gue19] Marie-Anne Guerry. Sufficient embedding conditions for threestate discrete-time Markov chains with real eigenvalues. Linear Multilinear Algebra, 67(1):106–120, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gumus:2015:NCA

[Gum15] Ibrahim Halil Gumus. A note on a conjecture about Wielandt’s inequality. Linear Multilinear Algebra, 63(9):1909– 1913, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2007:TLL

[Guo07] Ji-Ming Guo. On the third largest Laplacian eigenvalue of a graph. Linear Multilinear Algebra, 55(1):93–102, 2007. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2020:RBE

[Guo20] Pei-Chang Guo. A residual-based error bound for the multilinear PageRank vector. Linear Multilinear Algebra, 68(3): 568–574, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guralnick:1979:NPM

[Gur79] Robert M. Guralnick. A note on pairs of matrices with rank one commutator. Linear and Multilinear Algebra, 8(2):97–99, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guralnick:1980:SA

[Gur80] Robert M. Guralnick. Structure of algebras. Linear and Multilinear Algebra, 9(2):133–140, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guralnick:1983:IS

[Gur83] R. M. Guralnick. On an identity of Schur. Linear and Multilinear Algebra, 12(4):281–284, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guralnick:1992:NCP

[Gur92] Robert M. Guralnick. A note on commuting pairs of matrices. Linear and Multilinear Algebra, 31(1–4):71–75, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guralnick:1997:IPAa

[Gur97] Robert M. Guralnick. Invertible preservers and algebraic groups II: Preservers of similarity invariants and overgroups of PSL_{n}(F). Linear and Multilinear Algebra, 43(1–3):221–255, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gurvits:1998:TIL

[Gur98] Lemod Gurvits. Trace inequalities and linear programming (with applications to Markov chains). Linear and Multilinear Algebra, 44(1):13–24, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gustafson:1981:MSC

[Gus81] William H. Gustafson. On matrix similarity over commutative rings. Linear and Multilinear Algebra, 10(3):249–252, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gustafson:1994:OT

[Gus94] Karl Gustafson. Operator trigonometry. Linear and Multilinear Algebra, 37(1–3):139–159, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guterman:2001:FTT

[Gut01] Alexander Guterman. Frobenius type theorems in the noncommutative case. Linear and Multilinear Algebra, 48(4):293– 311, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gutin:2022:GSV

[Gut22] Ran Gutin. Generalizations of singular value decomposition to dual-numbered matrices. Linear Multilinear Algebra, 70(20): 5107–5114, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gutin:2023:ARB

[Gut23] Ran Gutin. An analogue of the relationship between SVD and pseudoinverse over double-complex matrices. Linear Multilinear Algebra, 71(11):1775–1793, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guzman:1988:TRP

[Guz88] Fernando Guzmán. Two results on the path algebra of a directed graph. Linear and Multilinear Algebra, 23(3):251–253, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Greub:1987:BIM

[GV87] W. H. Greub and J. R. Vanstone. A basic identity in mixed exterior algebra. Linear and Multilinear Algebra, 21(1):41–61, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Giambruno:1989:NPR

[GV89] A. Giambruno and A. Valenti. Nill-polynomials in rings. Linear and Multilinear Algebra, 24(2):127–134, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Giambruno:1995:MIM

[GV95] Antonio Giambruno and Angela Valenti. On minimal *- identities of matrices. Linear and Multilinear Algebra, 39(4): 309–323, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gregory:1996:RDS

[GVS96] David A. Gregory, Kevin N. Vander Meulen, and Bryan L. Shader. Rank decompositions and signed bigraphs. Linear and Multilinear Algebra, 40(4):283–301, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:1998:DI

[GW98a] Hwa-Long Gau and Pei Yuan Wu. Dilation to inflations of s(ϕ). Linear and Multilinear Algebra, 45(2–3):109–123, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:1998:NR

[GW98b] Hwa-Long Gau and Pei Yuan Wu. Numerical range of s(ϕ). Linear and Multilinear Algebra, 45(1):49–73, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:2004:NRN

[GW04] Hwa-Long Gau and Pei Yuan Wu. Numerical range of a normal compression. Linear Multilinear Algebra, 52(3–4):195– 201, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:2008:LSE

[GW08] Hwa-Long Gau and Pei Yuan Wu. Line segments and elliptic arcs on the boundary of a numerical range. Linear Multilinear Algebra, 56(1–2):131–142, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:2012:NED

[GW12] Hwa-Long Gau and Pei Yuan Wu. Noncircular elliptic discs as numerical ranges of nilpotent operators. Linear Multilinear Algebra, 60(11–12):1225–1233, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:2017:ORP

[GW17] Hwa-Long Gau and Pei Yuan Wu. Operators with real parts at least. Linear Multilinear Algebra, 65(10):1988–1999, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grossmann:2021:FMR

[GW21] Ben Grossmann and Hugo J. Woerdeman. Fractional minimal rank. Linear Multilinear Algebra, 69(1):19–39, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:2015:NRT

[GWW15] Hwa-Long Gau, Kuo-Zhong Wang, and Pei Yuan Wu. Numerical radii for tensor products of matrices. Linear Multilinear Algebra, 63(10):1916–1936, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:2016:CNR

[GWW16] Hwa-Long Gau, Kuo-Zhong Wang, and Pei Yuan Wu. Circular numerical ranges of partial isometries. Linear Multilinear Algebra, 64(1):14–35, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Geng:2020:CGG

[GWW20] Xianya Geng, Yan Wu, and Long Wang. Characterizations of graphs with given inertia index achieving the maximum diameter. Linear Multilinear Algebra, 68(8):1633–1641, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gu:2014:SUG

[GWZ14] Zhenhua Gu, Zhe-Xian Wan, and Kai Zhou. Subconstituents of unitary graphs over finite fields. Linear Multilinear Algebra, 62(7):925–937, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gu:2016:ASU

[GWZ16] Zhenhua Gu, Zhe-Xian Wan, and Kai Zhou. Automorphisms of subconstituents of unitary graphs over finite fields. Linear Multilinear Algebra, 64(9):1833–1852, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gumus:2017:SNR

[GX17a] Mehmet Gumus and Jianhong Xu. Some new results related to α-stability. Linear Multilinear Algebra, 65(2):325–340, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2017:EKR

[GX17b] Jun Guo and Qiuli Xu. The Erdős–Ko–Rado theorem for finite affine spaces. Linear Multilinear Algebra, 65(3):593–599, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2022:RCK

[GY22] Pei-Chang Guo and Qiang Ye. On the regularization of convolutional kernel tensors in neural networks. Linear Multilinear Algebra, 70(12):2318–2330, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gyires:1978:ICM

[Gyi78] Béla Gyires. On inequalities concerning matrices of permanents. Linear and Multilinear Algebra, 5(4):279–282, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Golberg:1976:IRB

[GZ76] Moshe Golberg and Gideon Zwas. Inclusion relations between certain sets of matrices. Linear and Multilinear Algebra, 4 (1):55–60, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Grone:1990:LEL

[GZ90] Robert Grone and Georg Zimmermann. Large eigenvalues of the Laplacian. Linear and Multilinear Algebra, 28(1–2):45–47, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gibson:1999:COM

[GZ99] Peter M. Gibson and Guo-Hui Zhang. Column orthogonal (0, 1, −1)-matrices. Linear and Multilinear Algebra, 45(4):291– 316, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2022:MLE

[GZ22a] Shu-Guang Guo and Rong Zhang. Maximizing the least Qeigenvalue of a unicyclic graph with perfect matchings. Linear Multilinear Algebra, 70(6):1128–1145, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2022:OGT

[GZ22b] Shu-Guang Guo and Rong Zhang. Ordering graphs by their largest (least) A_{α}-eigenvalues. Linear Multilinear Algebra, 70 (21):7049–7056, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2019:PT

[GZH19] Qi Guo, Meng-Meng Zheng, and Zheng-Hai Huang. Properties of S-tensors. Linear Multilinear Algebra, 67(4):685–696, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2020:CHL

[GZW20] Shuangjian Guo, Xiaohui Zhang, and Shengxiang Wang. The construction of Hom-left-symmetric conformal bialgebras. Linear Multilinear Algebra, 68(6):1257–1276, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Guo:2018:HGG

[GZY18] Shu-Guang Guo, Rong Zhang, and Guanglong Yu. Hamiltonian graphs of given order and minimum algebraic connectivity. Linear Multilinear Algebra, 66(3):459–468, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gao:2023:NCR

[GZZF23] Jiale Gao, Kezheng Zuo, Honglin Zou, and Zhimei Fu. New characterizations and representations of the generalized Bott– Duffin inverse and its applications. Linear Multilinear Algebra, 71(6):1026–1043, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ha:1988:IM

[Ha88] Chung-Wei Ha. An inequality for M-matrices. Linear and Multilinear Algebra, 23(3):263–267, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ha:1989:RPE

[Ha89] Chung-Wei Ha. On the real parts of eigenvalues of a matrix. Linear and Multilinear Algebra, 25(4):255–256, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hameed:2022:IAM

[HA22] Saira Hameed and Uzma Ahmad. Inverse of the adjacency matrices and strong anti-reciprocal eigenvalue property. Linear Multilinear Algebra, 70(14):2739–2764, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hackbusch:2022:MDB

[Hac22] Wolfgang Hackbusch. Minimal divergence for border rank-2 tensor approximation. Linear Multilinear Algebra, 70(20): 4915–4931, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hadwin:1983:ARL

[Had83] Don Hadwin. Algebraically reflexive linear transformations. Linear and Multilinear Algebra, 14(3):225–233, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hailat:1988:CRG

[Hai88a] Mohammad Q. Hailat. Classification of rank 2 GWrootsystems. Linear and Multilinear Algebra, 23(1):15–26, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hailat:1988:GRF

[Hai88b] Mohammad Q. Hailat. GW-rootsystems over fields of characteristic seven. Linear and Multilinear Algebra, 23(3):193–205, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hailat:1992:CRS

[Hai92a] Mohammad Q. Hailat. Classification of S(n, p, 2, 2) root systems. Linear and Multilinear Algebra, 33(3–4):217–231, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hailat:1992:RST

[Hai92b] Mohammad Q. Hailat. Root systems of type S(n, p, 2, 2). Linear and Multilinear Algebra, 31(1–4):1–17, 1992. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hailat:1997:CWR

[Hai97] Mohammad Q. Hailat. On the classification of Witt rootsystems. Linear and Multilinear Algebra, 42(1):61–71, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hajarian:2018:CPB

[Haj18] Masoud Hajarian. Convergence properties of BCR method for generalized Sylvester matrix equation over generalized reflexive and anti-reflexive matrices. Linear Multilinear Algebra, 66 (10):1975–1990, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Halmos:1986:EA

[Hal86] P. R. Halmos. Extensions and approximations. Linear and Multilinear Algebra, 19(3):297–298, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Halmos:1991:BPG

[Hal91] Paul R. Halmos. Bad products of good matrices. Linear and Multilinear Algebra, 29(1):1–20, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Han:2017:SNI

[Han17] Yazhou Han. Submajorization and p-norm inequalities associated with τ-measurable operators. Linear Multilinear Algebra, 65(11):2199–2211, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Han:2019:SDI

[Han19] Yazhou Han. Some determinant inequalities for operators. Linear Multilinear Algebra, 67(1):94–105, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hanson:2021:HRL

[Han21] Edward Hanson. How to recognize a Leonard pair. Linear Multilinear Algebra, 69(1):177–192, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1978:ABP

[Har78a] D. J. Hartfiel. On the asymptotic behavior of the Perron eigenvalue. Linear and Multilinear Algebra, 5(4):289–295, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartwig:1978:NPO

[Har78b] Robert E. Hartwig. A note on the partial ordering of positive semi-definite matrices. Linear and Multilinear Algebra, 6(3): 223–226, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1981:HRC

[Har81a] D. J. Hartfiel. How a row change in a stochastic matrix affects the corresponding stochastic eigenvector. Linear and Multilin-

ear Algebra, 10(3):205–216, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartwig:1981:NRA

[Har81b] Robert E. Hartwig. A note on rank-additivity. Linear and Multilinear Algebra, 10(1):59–61, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Harper:1984:MSS

[Har84a] L. H. Harper. Morphisms for the strong Sperner property of Stanley and Griggs. Linear and Multilinear Algebra, 16(1–4): 323–337, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1984:ERL

[Har84b] D. J. Hartfiel. An ergodic result for the limit of a sequence of substochastic products xA_{1} … A_{k}. Linear and Multilinear Algebra, 15(3–4):193–205, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartmann:1985:CI

[Har85] W. Hartmann. On the complexity of immanants. Linear and Multilinear Algebra, 18(2):127–140, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1987:FIN

[Har87] D. J. Hartfiel. Fully indecomposable and nearly decomposable matrices in Ω(r, s). Linear and Multilinear Algebra, 21(3): 231–237, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1991:FPT

[Har91] D. J. Hartfiel. Full patterns in truncated transportation polytopes. Linear and Multilinear Algebra, 30(3):213–221, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1994:ICN

[Har94] D. J. Hartfiel. Input control in nonnegative matrix equations. Linear and Multilinear Algebra, 36(4):293–304, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1995:MPE

[Har95] D. J. Hartfiel. Matrix patterns for eigenvalue-eigenvector adjustment. Linear and Multilinear Algebra, 40(1):69–80, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1996:CIL

[Har96] D. J. Hartfiel. Concerning interval Leslie matrix calculations. Linear and Multilinear Algebra, 41(2):95–106, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hasanov:2018:ME

[Has18] Vejdi Ismailov Hasanov. On the matrix equation. Linear Multilinear Algebra, 66(9):1783–1798, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Haukkanen:2012:TAD

[Hau12] Pentti Haukkanen. A topological approach to divisibility of arithmetical functions and GCD matrices. Linear Multilinear Algebra, 60(3):301–309, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Haviv:1988:EBA

[Hav88] Moshe Haviv. Error bounds on an approximation to the dominant eigenvector of a nonnegative matrix. Linear and Multilinear Algebra, 23(2):159–163, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hawkins:2005:IBP

[Haw05] Sidney T. Hawkins. Inequalities between Pythagoras numbers and algebraic ranks in Witt rings of fields. Linear Multilinear Algebra, 53(6):455–461, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

auYeung:1999:QNR

[haYsS99] Yik hoi au Yeung and Lok shun Siu. Quaternionic numerical range and real subspaces. Linear and Multilinear Algebra, 45(4):317–327, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hershkowitz:1983:NCS

[HB83] Daniel Hershkowitz and Abraham Berman. Necessary conditions and a sufficient condition for the Fischer–Hadamard inequalities. Linear and Multilinear Algebra, 13(1):67–72, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hawkins:1973:GMF

[HBI73] J. B. Hawkins and A. Ben-Israel. On generalized matrix functions. Linear and Multilinear Algebra, 1(2):163–171, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hocine:2015:LRG

[HBM15] Kouider Miloud Hocine, Mohammed Benharrat, and Bekkai Messirdi. Left and right generalized Drazin invertible operators. Linear Multilinear Algebra, 63(8):1635–1648, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hossein:2018:PDS

[HBP18] Sk. Monowar Hossein, Snehasish Bose, and Kallol Paul. On positive definite solution of nonlinear matrix equations. Linear Multilinear Algebra, 66(5):881–893, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hai:2014:IUT

[HC14] Guojun Hai and Alatancang Chen. On the invertibility of upper triangular operator matrices. Linear Multilinear Algebra, 62(4):538–547, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Howell:2019:RTN

[HCS19] K.-T. Howell, D. S. Chistyakov, and S. P. Sanon. On representation theory and near-vector spaces. Linear Multilinear Algebra, 67(7):1495–1510, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Haulk:1997:CSM

[HDJT97] Claire Cates Haulk, John Drew, Charles R. Johnson, and J. Hunter Tart. Characterization of supercommuting matrices.

Linear and Multilinear Algebra, 43(1–3):35–51, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Haily:2020:ESA

[HE20] Abdelfattah Haily and Mohamed ElBadry. Equation a ∗ a = P (a) in some algebras with involution. Linear Multilinear Algebra, 68(11):2221–2227, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

He:2021:SNR

[He21] Zhuo-Heng He. Some new results on a system of Sylvester-type quaternion matrix equations. Linear Multilinear Algebra, 69 (16):3069–3091, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hemasinha:1993:SIC

[Hem93] Rohan Hemasinha. The sign invariance of certain norms on. Linear and Multilinear Algebra, 35(2):135–151, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Heo:2016:SIC

[Heo16] Jaeseong Heo. The structure of invariant α-CP multilinear maps and associated J-representations. Linear Multilinear Algebra, 64(7):1295–1313, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hershkowitz:1983:EMP

[Her83] Daniel Hershkowitz. Existence of matrices with prescribed eigenvalues and entries. Linear and Multilinear Algebra, 14(4): 315–342, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Heyfron:1988:IDO

[Hey88] Peter Heyfron. Immant dominance orderings for hook partitions. Linear and Multilinear Algebra, 24(1):65–78, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Heyfron:1992:GHI

[Hey92] Peter Heyfron. A generalization of Hadamard’s inequality. Linear and Multilinear Algebra, 32(1):75–84, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hosseini:2018:MIS

[HF18] Maliheh Hosseini and Juan J. Font. Multilinear isometries on spaces of vector-valued continuous functions. Linear Multilinear Algebra, 66(8):1681–1690, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hosseini:2015:MIF

[HFS15] Maliheh Hosseini, Juan J. Font, and Manuel Sanchis. Multilinear isometries on function algebras. Linear Multilinear Algebra, 63(7):1448–1457, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:CPR

[HFSZ22] Xin Huang, Zhi-Wei Fang, Hai-Wei Sun, and Chun-Hua Zhang. A circulant preconditioner for the Riesz distributedorder space-fractional diffusion equations. Linear Multilinear Algebra, 70(16):3081–3096, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hong:1985:SRF

[HH85] Yoo Pyo Hong and Roger A. Horn. On simultaneous reduction of families of matrices to triangular or diagonal form by unitary congruences. Linear and Multilinear Algebra, 17(3–4): 271–288, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hong:1989:CUC

[HH89] Yoopyo Hong and Roger A. Horn. A characterization of unitary congruence. Linear and Multilinear Algebra, 25(2):105– 119, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:2014:MDC

[HH14] Shuangnian Hu and Shaofang Hong. Multiple divisor chains and determinants of matrices associated with completely even functions (mod r ). Linear Multilinear Algebra, 62(9):1240– 1257, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2016:NLM

[HH16] Jinchuan Hou and Kan He. Non-linear maps on self-adjoint operators preserving numerical radius and numerical range of Lie product. Linear Multilinear Algebra, 64(1):36–57, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hajighasemi:2022:MPS

[HH22a] Sepide Hajighasemi and Shirin Hejazian. Maps preserving some spectral domains of operator products. Linear Multilinear Algebra, 70(12):2367–2381, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hazra:2022:ESC

[HH22b] Samir Kumar Hazra and Amber Habib. Exact sequences in the cohomology of a Lie superalgebra extension. Linear Multilinear Algebra, 70(20):5014–5031, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hao:2012:NRL

[HHC12] Hui-Feng Hao, Yong-Jian Hu, and Gong-Ning Chen. Note on roots location of a symmetric polynomial with respect to the imaginary axis. Linear Multilinear Algebra, 60(4):499–510, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Halbeisen:2022:SRU

[HHSY22] Lorenz Halbeisen, Norbert Hungerbühler, Salome Schumacher, and Guo Xian Yau. Sets of range uniqueness for multivariate polynomials and linear functions with rank k. Linear Multilinear Algebra, 70(20):5642–5660, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2015:LIT

[HHW15] Xueyi Huang, Qiongxiang Huang, and Fei Wen. On the Laplacian integral tricyclic graphs. Linear Multilinear Algebra, 63 (7):1356–1371, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

He:2022:BRC

[HHY22] Shengjie He, Rong-Xia Hao, and Aimei Yu. Bounds for the rank of a complex unit gain graph in terms of the independence number. Linear Multilinear Algebra, 70(7):1382–1402, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hiai:1994:ECM

[Hia94] Fumio Hiai. Equality cases in matrix norm inequalities of Golden–Thompson type. Linear and Multilinear Algebra, 36(4):239–249, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hillel:1976:DAS

[Hil76] J. Hillel. Dimensionality of algebras of symmetry classes of tensors. Linear and Multilinear Algebra, 4(1):41–44, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hildebrand:2007:LDC

[Hil07a] Roland Hildebrand. An LMI description for the cone of Lorentz-positive maps. Linear Multilinear Algebra, 55(6):551– 573, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hildebrand:2007:PMS

[Hil07b] Roland Hildebrand. Positive maps of second-order cones. Linear Multilinear Algebra, 55(6):575–597, 2007. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hildebrand:2011:LDC

[Hil11] Roland Hildebrand. An LMI description for the cone of Lorentz-positive maps II. Linear Multilinear Algebra, 59(7): 719–731, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hilberdink:2017:SVM

[Hil17] Titus Hilberdink. Singular values of multiplicative Toeplitz matrices. Linear Multilinear Algebra, 65(4):813–829, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hildebrand:2018:ECM

[Hil18] R. Hildebrand. Extremal copositive matrices with minimal zero supports of cardinality two. Linear Multilinear Algebra, 66(11):2151–2155, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Haukkanen:2010:UAG

[HINS10] Pentti Haukkanen, Pauliina Ilmonen, Ayse Nalli, and Juha Sillanpää. On unitary analogs of GCD reciprocal LCM matrices. Linear Multilinear Algebra, 58(5):599–616, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hoffman:1977:LGC

[HJ77] A. J. Hoffman and Basharat A. Jamil. On the line graph of the complete tripartite graph. Linear and Multilinear Algebra, 5(1):19–25, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Horn:1987:HCS

[HJ87] Roger A. Horn and Charles R. Johnson. Hadamard and conventional submultiplicativity for unitarily invariant norms on matrices. Linear and Multilinear Algebra, 20(2):91–106, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Heo:2019:CST

[HJ19a] Jaeseong Heo and Maria Joita. Corrigendum to ‘A Stinespring type theorem for completely positive multilinear maps on Hilbert C∗-modules’ [Linear Multilinear Algebra 67 (2019), 121–140]. Linear Multilinear Algebra, 67(8):1715–1716, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [HJ19b].

Heo:2019:STT

[HJ19b] Jaeseong Heo and Maria Joita. A Stinespring type theorem for completely positive multilinear maps on Hilbert C∗-modules. Linear Multilinear Algebra, 67(1):121–140, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [HJ19a].

Hashemi:2020:CBM

[HJ20a] Amir Hashemi and Masoumeh Javanbakht. Computing Hbases via minimal bases for syzygy modules. Linear Multilinear Algebra, 68(7):1277–1301, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Heo:2020:RNT

[HJ20b] Jaeseong Heo and Maria Joia. A Radon–Nikodým type theorem for invariant symmetric completely positive and completely bounded multilinear maps on Hilbert C∗*-modules. Linear Multilinear Algebra, 68(9):1878–1893, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Han:2017:CEG

[HJP17] Xuchen Han, Charles R. Johnson, and Pietro Paparella. The critical exponent for generalized doubly nonnegative matrices. Linear Multilinear Algebra, 65(5):1035–1044, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hall:1979:RIG

[HK79] Frak J. Hall and Irving J. Katz. On ranks of integral generalized inverses of integral matrices. Linear and Multilinear Algebra, 7(1):73–85, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hall:1980:MIG

[HK80] Frank J. Hall and Irving J. Katz. More on integral generalised inverses of integral matrics. Linear and Multilinear Algebra, 9(3):201–209, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:1992:CNM

[HK92] Suk-Geun Hwang and Si-Ju Kim. On convertible nonnegative matrices. Linear and Multilinear Algebra, 32(3–4):311–318, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2010:KPO

[HK10] Huajun Huang and Sangjib Kim. On Kostant’s partial order on hyperbolic elements. Linear Multilinear Algebra, 58(6): 783–788, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holmes:2013:OBB

[HK13] Randall R. Holmes and Avantha Kodithuwakku. Orthogonal bases of Brauer symmetry classes of tensors for the dihedral group. Linear Multilinear Algebra, 61(8):1136–1147,

Ha:2016:CEI

[HK16a] Kil-Chan Ha and Seung-Hyeok Kye. Construction of exposed indecomposable positive linear maps between matrix algebras. Linear Multilinear Algebra, 64(11):2188–2198, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holmes:2016:SCT

[HK16b] Randall R. Holmes and Avantha Kodithuwakku. Symmetry classes of tensors associated with principal indecomposable characters and Osima idempotents. Linear Multilinear Algebra, 64(4):574–586, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huckle:2017:PSS

[HK17] Thomas K. Huckle and Christos D. Kravvaritis. Properties of submatrices of Sylvester Hadamard matrices. Linear Multilinear Algebra, 65(8):1629–1642, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hari:2021:CCJ

[HK21] Vjeran Hari and Erna Begović Kovac. On the convergence of complex Jacobi methods. Linear Multilinear Algebra, 69(3): 489–514, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hoa:2016:GRC

[HKO16] Dinh Trung Hoa, Vo Thi Bich Khue, and Hiroyuki Osaka. A generalized reverse Cauchy inequality for matrices. Linear Multilinear Algebra, 64(7):1415–1423, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hamhalter:2023:DJM

[HKP23] Jan Hamhalter, Ondrej F. K. Kalenda, and Antonio M. Peralta. Determinants in Jordan matrix algebras. Linear Multilinear Algebra, 71(6):961–1002, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:1994:LOP

[HKS94] Suk-Geun Hwang, Si-Ju Kim, and Seok-Zun Song. Linear operators that preserve maximal column rank of boolean matrices. Linear and Multilinear Algebra, 36(4):305–313, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:1995:MCM

[HKS95] Suk-Geun Hwang, Si-Ju Kim, and Seok-Zun Song. On maximal convertible matrices. Linear and Multilinear Algebra, 38(3):171–176, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hashemi:2015:CTA

[HKT15] Behnam Hashemi, Mahtab Mirzaei Khalilabadi, and Hanieh Tavakolipour. A cubic time algorithm for finding the principal solution to Sylvester matrix equations over (max, +). Linear Multilinear Algebra, 63(2):283–295, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Heaton:2023:MTF

[HKv23] Alexander Heaton, Khazhgali Kozhasov, and Lorenzo venturello. Minimality of tensors of fixed multilinear rank. Linear Multilinear Algebra, 71(8):1364–1377, 2023. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartfiel:1984:DVF

[HL84] D. J. Hartfiel and Raphael Loewy. A determinantal version of the Frobenius–König theorem. Linear and Multilinear Algebra, 16(1–4):155–165, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Han:2009:NPJ

[HL09] Jen-Yin Han and Chin-Yuan Lin. A new proof of Jordan canonical forms of a square matrix. Linear Multilinear Algebra, 57(4):369–386, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2010:CFI

[HL10] Yufei Huang and Bolian Liu. On a conjecture for fully indecomposable exponent and Hall exponent. Linear Multilinear Algebra, 58(6):699–710, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2018:EHT

[HL18] Jing Huang and Shuchao Li. Expected hitting times for random walks on quadrilateral graphs and their applications. Linear Multilinear Algebra, 66(12):2389–2408, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2020:MWP

[HL20] Zejun Huang and Zhenhua Lyu. 0–1 matrices whose k-th powers have bounded entries. Linear Multilinear Algebra, 68(10): 1972–1982, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hur:2021:NRC

[HL21] Injo Hur and Ji Eun Lee. Numerical ranges of conjugations and antilinear operators. Linear Multilinear Algebra, 69(16): 2990–2997, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:GTD

[HL22] Xueyi Huang and Huiqiu Lin. On graphs with three distinct signless Laplacian eigenvalues. Linear Multilinear Algebra, 70 (8):1451–1466, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hladik:2017:TIL

[Hla17] M. Hladík. Transformations of interval linear systems of equations and inequalities. Linear Multilinear Algebra, 65(2):211– 223, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hajmohamadi:2020:IBN

[HLB20] Monire Hajmohamadi, Rahmatollah Lashkaripour, and Mojtaba Bakherad. Improvements of Berezin number inequalities. Linear Multilinear Algebra, 68(6):1218–1229, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:GSR

[HLB22] Yufei Huang, Muhuo Liu, and Francesco Belardo. The general spectral radii of (multicone-)graphs with prescribed degree sequence. Linear Multilinear Algebra, 70(19):4751–4777, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2018:CJM

[HLH18] Wenbo Huang, Jiankui Li, and Jun He. Characterizations of Jordan mappings on some rings and algebras through zero products. Linear Multilinear Algebra, 66(2):334–346, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hogben:2020:SSS

[HLOvdD20] Leslie Hogben, Jephian C.-H. Lin, D. D. Olesky, and P. van den Driessche. The sepr-sets of sign patterns. Linear Multilinear Algebra, 68(10):2044–2068, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2003:LES

[HLP03] Yaoping Hou, Jiongsheng Li, and Yongliang Pan. On the Laplacian eigenvalues of signed graphs. Linear Multilinear Algebra, 51(1):21–30, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:2023:BDI

[HLP23] In Sung Hwang, Woo Young Lee, and Jaehui Park. The Beurling degree of inner matrix functions. Linear Multilinear Algebra, 71(3):472–479, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2017:IGM

[HLPW17] Shaowu Huang, Chi-Kwong Li, Yiu-Tung Poon, and Qing-Wen Wang. Inequalities on generalized matrix functions. Linear Multilinear Algebra, 65(10):1947–1961, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hill:2018:SMS

[HLS18] S. L. Hill, M. C. Lettington, and K. M. Schmidt. On superalgebras of matrices with symmetry properties. Linear Multilinear Algebra, 66(8):1538–1563, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hernandez:2022:GIM

[HLT22] M. V. Hernández, M. B. Lattanzi, and N. Thome. GDMPinverses of a matrix and their duals. Linear Multilinear Algebra, 70(19):3923–3935, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2021:SRG

[HLW^{+}21] Xinmin Hou, Boyuan Liu, Shicheng Wang, Jun Gao, and Chenhui Lv. The spectral radius of graphs without trees of diameter at most four. Linear Multilinear Algebra, 69(8):1407– 1414, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

He:2022:NPP

[HLY22] Miao He, Jinsong Leng, and Jiali Yu. New properties of phaseretrievable g-frames and exact phase-retrievable g-frames. Linear Multilinear Algebra, 70(19):4117–4132, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:CDS

[HLYZ22] Hua-Lin Huang, Huajun Lu, Yu Ye, and Chi Zhang. On centres and direct sum decompositions of higher degree forms. Linear Multilinear Algebra, 70(22):7290–7306, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Havlicek:2002:AFS

[HLZ02] Hans Havlicek, Klaus List, and Corrado Zanella. On automorphisms of flag spaces. Linear Multilinear Algebra, 50(3): 241–251, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:2005:PNO

[HLZ05] Qi Hu, Yaqin Li, and Xingzhi Zhan. Possible numbers of ones in 0–1 matrices with a given rank. Linear Multilinear Algebra, 53(6):435–443, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2011:SCR

[HLZ11] Rong Huang, Jianzhou Liu, and Li Zhu. A structural characterization of real k-potent matrices. Linear Multilinear Algebra, 59(4):433–439, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2018:DIU

[HLZ18] Xin Hou, Shangzhi Li, and Qingqing Zheng. On the decompositions of infinite upper triangular matrices over rings. Linear Multilinear Algebra, 66(9):1917–1924, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hao:2022:IFD

[HLZ22] Chan Hao, Shuchao Li, and Licheng Zhang. An inverse formula for the distance matrix of a fan graph. Linear Multilinear Algebra, 70(22):7807–7824, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hoskins:1976:ICT

[HM76] W. D. Hoskins and G. E. McMaster. On the inverses of a class of Toeplitz matrices of band width five. Linear and Multilinear Algebra, 4(2):103–106, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hoskins:1977:PIS

[HM77] W. D. Hoskins and G. E. McMaster. Properties of the inverses of a set of band matrices. Linear and Multilinear Algebra, 5(3): 183–196, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hoskins:1978:INI

[HM78] W. D. Hoskins and G. E. McMaster. On the infinity norm of the inverse of a class of Toeplitz matrices of band width five. Linear and Multilinear Algebra, 6(2):153–156, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:1995:ISP

[HM95] Suk-Geun Hwang and Roy Meshulam. An inequality for subpermanents of positive semidefinite Hermitian matrices. Linear and Multilinear Algebra, 38(3):177–180, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2018:SPD

[HM18] Na Huang and Chang-Feng Ma. The structure-preserving doubling algorithms for positive definite solution to a system of nonlinear matrix equations. Linear Multilinear Algebra, 66(4): 827–839, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2019:SCI

[HM19] Baohua Huang and Changfeng Ma. Some criteria for identifying strong H-tensors and its applications. Linear Multilinear Algebra, 67(6):1146–1173, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2020:IAS

[HM20] Baohua Huang and Changfeng Ma. An iterative algorithm to solve the generalized Sylvester tensor equations. Linear Multilinear Algebra, 68(6):1175–1200, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:ISC

[HM22] Baohua Huang and Changfeng Ma. The iterative solution of a class of tensor equations via Einstein product with a tensor inequality constraint. Linear Multilinear Algebra, 70(21): 6321–6344, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hessert:2023:IIM

[HM23] Ryan Hessert and Sudipta Mallik. The inverse of the incidence matrix of a unicyclic graph. Linear Multilinear Algebra, 71(4): 513–527, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hashemi:2013:MPS

[HMAD13] Amir Hashemi, Benyamin M.-Alizadeh, and Mahdi Dehghani Darmian. Minimal polynomial systems for parametric matrices. Linear Multilinear Algebra, 61(2):265–272, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hiroshima:2022:TDS

[HMK22] Fumio Hiroshima, Zahriddin Muminov, and Utkir Kuljanov. Threshold of discrete schrödinger operators with delta potentials on n-dimensional lattice. Linear Multilinear Algebra, 70(5):919–954, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Horn:1991:IUI

[HMN91] Roger A. Horn, Roy Mathias, and Yoshihiro Nakamura. Inequalities for unitarily invariant norms and bilinear matrix products. Linear and Multilinear Algebra, 30(4):303–314, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Higuchi:2011:SDL

[HMO11] Yusuke Higuchi, Tomonori Matsumoto, and Osamu Ogurisu. On the spectrum of a discrete Laplacian on z with finitely supported potential. Linear Multilinear Algebra, 59(8):917– 927, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Halaseh:2022:ODT

[HMR22] Karim Halaseh, Tommi Muller, and Elina Robeva. Orthogonal decomposition of tensor trains. Linear Multilinear Algebra, 70 (21):6609–6639, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hai:2022:AGV

[HNB22] Bui Xuan Hai, Cao Minh Nam, and Mai Hoang Bien. Automorphism groups of vector spaces with generalized group identities. Linear Multilinear Algebra, 70(20):5948–5956, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hershkowitz:1999:HPS

[HNS99] Daniel Hershkowitz, Michael Neumann, and Hans Schneider. Hermitian positive semidefinite matrices whose entries are 0 or 1 in modulus. Linear and Multilinear Algebra, 46(4):259–264, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2017:SPM

[HO17] Huajun Huang and Luke Oeding. Symmetrization of principal minors and cycle-sums. Linear Multilinear Algebra, 65(6): 1194–1219, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hoffman:1975:LF

[Hof75] Alan J. Hoffman. Linear G-functions. Linear and Multilinear Algebra, 3(1–2):45–52, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holmquist:1984:LPM

[Hol84] Björn Holmquist. On the leading principal minors of direct products. Linear and Multilinear Algebra, 16(1–4):125–127, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holmquist:1985:DPP

[Hol85] Björn Holmquist. The direct product permuting matrices. Linear and Multilinear Algebra, 17(2):117–141, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holmes:1995:OBS

[Hol95] Randall R. Holmes. Orthogonal bases of symmetrized tensor spaces. Linear and Multilinear Algebra, 39(3):241–243, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holmes:2004:OCR

[Hol04] Randall R. Holmes. Orthogonality of cosets relative to irreducible characters of finite groups. Linear Multilinear Algebra, 52(2):133–143, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hong:1991:TET

[Hon91] Ho S. Hong. Two expansion theorems by diagonal elements for the permanent and the determinant. Linear and Multilinear Algebra, 29(3–4):181–194, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hong:1992:PAD

[Hon92] Ho S. Hong. Permanent analogues of determinantal identities related to permutation fixed-points. Linear and Multilinear Algebra, 32(1):1–20, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hong:1993:TDI

[Hon93] Ho S. Hong. On two determinantal identities and their permanent analogues. Linear and Multilinear Algebra, 35(2):115– 129, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hong:1999:LBD

[Hon99] Shaofang Hong. Lower bounds for determinants of matrices associated with classes of arithmetical functions. Linear and Multilinear Algebra, 45(4):349–358, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hong:2019:ESA

[Hon19] Yanyong Hong. Extending structures for associative conformal algebras. Linear Multilinear Algebra, 67(1):196–212, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hong:2022:CEC

[Hon22] Yanyong Hong. Central extensions and conformal derivations of a class of Lie conformal algebras. Linear Multilinear Algebra, 70(18):3577–3598, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hora:2004:ODC

[Hor04] Jan Hora. Orthogonal decompositions and canonical embeddings of multilinear alternating forms. Linear Multilinear Algebra, 52(2):121–132, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hosseini:2022:RSI

[Hos22] Maliheh Hosseini. Reflexivity of sets of isometries on bounded variation function spaces. Linear Multilinear Algebra, 70(19): 4405–4415, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hoa:2015:COM

[HOT15] Dinh Trung Hoa, Hiroyuki Osaka, and Jun Tomiyama. Characterization of operator monotone functions by Powers– Størmer type inequalities. Linear Multilinear Algebra, 63(8): 1577–1589, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2001:BGM

[Hou01] Yaoping Hou. Bicyclic graphs with minimum energy. Linear and Multilinear Algebra, 49(4):347–354, 2001. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2009:ANN

[Hou09] Xiang-Dong Hou. Asymptotic numbers of non-equivalent codes in three notions of equivalence. Linear Multilinear Algebra, 57(2):111–122, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2017:LDH

[Hou17] Bo Hou. Lie derivation and Hochschild cohomology of an extension of path algebras. Linear Multilinear Algebra, 65(5): 1022–1034, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2021:ITM

[Hou21] Xin Hou. Idempotents in triangular matrix rings. Linear Multilinear Algebra, 69(2):296–304, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2022:PCS

[Hou22] Xin Hou. Products of commutators of symplectic involutions. Linear Multilinear Algebra, 70(15):2984–2997, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Howard:1980:DIR

[How80] Ralph Howard. Dimension inequalities on the range of a multilinear function with vector space range. Linear and Multilinear Algebra, 8(3):247–252, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Howland:1983:FWA

[How83] James Lucien Howland. Further ways to approximate the exponential of a matrix. Linear and Multilinear Algebra, 14(2): 121–129, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Howell:1986:SM

[How86] John A. Howell. Spans in the module (Z_{m})^{s}. Linear and Multilinear Algebra, 19(1):67–77, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartwig:1990:WMD

[HP90a] Robert E. Hartwig and Mohan S. Putcha. When is a matrix a difference of two idempotents. Linear and Multilinear Algebra, 26(4):267–277, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartwig:1990:WMS

[HP90b] Robert E. Hartwig and Mohan S. Putcha. When is a matrix a sum of idempotents? Linear and Multilinear Algebra, 26(4):

Hartwig:2018:GIN

[HP18] R. E. Hartwig and P. Patrício. The group inverse of the nivellateur. Linear Multilinear Algebra, 66(12):2409–2420, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Henderson:1983:HKP

[HPS83] Harold V. Henderson, Friedrich Pukelsheim, and Shayle R. Searle. On the history of the Kronecker product. Linear and Multilinear Algebra, 14(2):113–120, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huylebrouck:1984:MPI

[HPV84] D. Huylebrouck, R. Puystjens, and J. Van Geel. The Moore– Penrose inverse of a matrix over a semi-simple artinian ring. Linear and Multilinear Algebra, 16(1–4):239–246, 1984. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huylebrouck:1988:MPI

[HPV88] D. Huylebrouck, R. Puystjens, and J. Van Geel. The Moore– Penrose inverse of a matrix over a semi-simple artinian ring with respect to an involution. Linear and Multilinear Algebra, 23(3):269–276, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:2014:ET

[HQ14] Shenglong Hu and Liqun Qi. The E-eigenvectors of tensors. Linear Multilinear Algebra, 62(10):1388–1402, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2018:SCP

[HQ18] Jinchuan Hou and Xiaofei Qi. Strong k-commutativity preservers on complex standard operator algebras. Linear Multilinear Algebra, 66(5):902–918, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Helton:1985:SPL

[HR85] J. W. Helton and L. Rodman. Signature preserving linear maps of Hermitian matrices. Linear and Multilinear Alge-

Helmberg:1995:SAB

[HRMP95] Christoph Helmberg, Franz Rendl, Bojan Mohar, and Svatopluk Poljak. A spectral approach to bandwidth and separator problems in graphs. Linear and Multilinear Algebra, 39(1– 2):73–90, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Herrero:2015:ASN

[HRT15] A. Herrero, F. J. Ramírez, and N. Thome. An algorithm to study the nonnegativity, regularity and stability via statefeedbacks of singular systems of arbitrary index. Linear Multilinear Algebra, 63(5):882–892, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Herstein:1975:ETS

[HS75] I. N. Herstein and Lance W. Small. An extension of a theorem of Schur. Linear and Multilinear Algebra, 3(1–2):41–43, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Henderson:1981:VPM

[HS81] Harold V. Henderson and S. R. Searle. The vec-permutation matrix, the vec operator and Kronecker products: a review. Linear and Multilinear Algebra, 9(4):271–288, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartwig:1983:MWC

[HS83a] Robert E. Hartwig and Klaus Spindelböck. Matrices for which A^{∗} and A^{+} commute. Linear and Multilinear Algebra, 14(3): 241–256, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartwig:1983:PIC

[HS83b] Robert E. Hartwig and Klaus Spindelböck. Partial isometries, contractions and EP matrices. Linear and Multilinear Algebra, 13(4):295–310, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hartwig:1983:SCF

[HS83c] Robert E. Hartwig and Klaus Spindelböck. Some closed form formulae for the intersection of two special matrices under the star order. Linear and Multilinear Algebra, 13(4):323–331, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hershkowitz:1985:SVS

[HS85] Daniel Hershkowitz and Hans Schneider. Scalings of vector spaces and the uniqueness of Lyapunov scaling factors. Linear and Multilinear Algebra, 17(3–4):203–226, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holenda:1987:DIF

[HS87] Jirí Holenda and Milos Schlegel. Degrees of the invariant factors and the rank of principal submatrices of a matrix. Linear and Multilinear Algebra, 21(3):297–301, 1987. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hershkowitz:1988:CRP

[HS88] Daniel Hershkowitz and Dafna Shasha. Cones of real positive semidefinite matrices associated with matrix stability. Linear and Multilinear Algebra, 23(2):165–181, 1988. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hershkowitz:1989:HBL

[HS89] Daniel Hershkowitz and Hans Schneider. Height bases, level bases, and the equality of the height and the level characteristics of an M-matrix. Linear and Multilinear Algebra, 25(2): 149–171, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hershkowitz:1991:CBD

[HS91] Daniel Hershkowitz and Hans Schneider. Combinatorial bases, derived Jordan sets and the equality of the height and level characteristics of an M-matrix. Linear and Multilinear Algebra, 29(1):21–42, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hershkowitz:1993:RZP

[HS93a] Daniel Hershkowitz and Hans Schneider. Ranks of zero patterns and sign patterns. Linear and Multilinear Algebra, 34

Hershkowitz:1993:CEM

[HS93b] Daniel Hershkowitz and Felix Shmidel. On a conjecture on the eigenvalues of p-matrices. Linear and Multilinear Algebra, 36(2):103–110, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:1993:MPM

[HS93c] Suk-Geun Hwang and Jia-Yu Shao. Minimization of the permanent and matrices of freedom. Linear and Multilinear Algebra, 35(3–4):313–323, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Haemers:1995:GCD

[HS95] W. H. Haemers and E. Spence. Graphs cospectral with distance-regular graphs. Linear and Multilinear Algebra, 39(1– 2):91–107, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Haukkanen:1996:SAS

[HS96] Pentti Haukkanen and Juha Sillanpää. Some analogues of Smith’s determinant. Linear and Multilinear Algebra, 41(3): 233–244, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Horn:2009:CFU

[HS09] Roger A. Horn and Vladimir V. Sergeichuk. Canonical forms for unitary congruence and *congruence. Linear Multilinear Algebra, 57(8):777–815, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hidaka:2012:CNA

[HS12] Chikara Hidaka and Takashi Sano. Conditional negativity of anti-Loewner matrices. Linear Multilinear Algebra, 60(11–12): 1265–1270, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2016:IDM

[HS16] Yaoping Hou and Yajing Sun. Inverse of the distance matrix of a bi-block graph. Linear Multilinear Algebra, 64(8):1509–1517,

Holubowski:2017:BPC

[HS17] W. Holubowski and A. V. Stepanov. Bijections preserving commutators and automorphisms of unitriangular groups. Linear Multilinear Algebra, 65(1):24–34, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hsia:1978:TTI

[Hsi78] J. S. Hsia. Two theorems on integral matrices. Linear and Multilinear Algebra, 5(4):257–264, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Herawati:2017:LSR

[HSM17] E. Herawati, Supama, and M. Mursaleen. Local structure of Riesz valued sequence spaces defined by an order Φ-function. Linear Multilinear Algebra, 65(3):545–554, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hsu:1985:TMK

[Hsu85] Chen-Jung Hsu. On a theorem of M. Kobayashi. Linear and Multilinear Algebra, 17(3–4):265–269, 1985. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hiai:2019:AHT

[HSW19] Fumio Hiai, Yuki Seo, and Shuhei Wada. Ando–Hiai type inequalities for multivariate operator means. Linear Multilinear Algebra, 67(11):2253–2281, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

He:2016:GES

[HSX16] Lili He, Chuanning Song, and Qingxiang Xu. General exact solutions of certain second-order homogeneous algebraic differential equations. Linear Multilinear Algebra, 64(6):1011–1031, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:1991:GNR

[HT91a] Shu-An Hu and Tin-Yau Tam. On the generalized numerical range with principal character. Linear and Multilinear Algebra, 30(1–2):93–107, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:1991:OPN

[HT91b] Shu-An Hu and Tin-Yau Tam. Operators with permanental numerical ranges on a straight line. Linear and Multilinear Algebra, 29(3–4):263–277, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holmes:1992:SCT

[HT92] Randall R. Holmes and Tin-Yau Tam. Symmetry classes of tensors associated with certain groups. Linear and Multilinear Algebra, 32(1):21–31, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2010:GND

[HT10] Huajun Huang and Tin-Yau Tam. On the Gelfand–Naimark decomposition of a nonsingular matrix. Linear Multilinear Algebra, 58(1):27–43, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Holmes:2012:CCS

[HT12] Randall R. Holmes and David P. Turner. The coefficient coalgebra of a symmetrized tensor space. Linear Multilinear Algebra, 60(4):385–396, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:1987:MWC

[Hu87] Shu-An Hu. On the Marcus–Wang conjecture. Linear and Multilinear Algebra, 20(3):191–196, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:1989:MWN

[Hu89] Shu-An Hu. On Marcus–Wang’s numerical range. Linear and Multilinear Algebra, 24(4):235–240, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:1994:CDJ

[Hu94] Shu-An Hu. On a conjecture of Drew and Johnson: maximizing the permanent. Linear and Multilinear Algebra, 36(4): 279–285, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [DJ89b].

Huang:1990:EIC

[Hua90] Qiang Huang. Extension of irreducible characters from normal subgroups. Linear and Multilinear Algebra, 27(2):117– 119, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:1996:LAA

[Hua96] Wenxue Huang. Linear associative algebras with finitely many idempotents. Linear and Multilinear Algebra, 40(4):303–309, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2001:CST

[Hua01] Danrun Huang. A cyclic six-term exact sequence for block matrices over a PID. Linear and Multilinear Algebra, 49(2): 91–114, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2008:SIH

[Hua08] Rong Huang. Some inequalities for Hadamard products of matrices. Linear Multilinear Algebra, 56(5):543–554, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2009:SEW

[Hua09] Huajun Huang. Some extensions of Witt’s theorem. Linear Multilinear Algebra, 57(4):321–344, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2010:APB

[Hua10] Li-Ping Huang. Adjacency preserving bijective maps on triangular matrices over any division ring. Linear Multilinear Algebra, 58(7):815–846, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2016:WCS

[Hua16] Zejun Huang. Weakening the conditions in some classical theorems on linear preserver problems. Linear Multilinear Algebra, 64(5):923–928, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:FDI

[Hua22a] Hau-Wen Huang. Finite-dimensional irreducible modules of the universal DAHA of type (C_{1}^{∨}, C_{1}). Linear Multilinear Algebra, 70(15):2856–2883, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:MSV

[Hua22b] Hong Huang. On majorizations and singular values. Linear Multilinear Algebra, 70(17):3304–3309, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:VPH

[Hua22c] Na Huang. Variable-parameter HSS methods for non-Hermitian positive definite linear systems. Linear Multilinear Algebra, 70(21):6664–6681, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2023:NVS

[Hua23] Shaowu Huang. A new version of Schur–Horn type theorem. Linear Multilinear Algebra, 71(1):41–46, 2023. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hughes:1990:NSG

[Hug90] Geraint Hughes. A note on the shape of the generalized numerical range. Linear and Multilinear Algebra, 26(1–2):43–47, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hulek:1981:RCM

[Hul81] Klaus Hulek. A remark on certain matrix varieties. Linear and Multilinear Algebra, 10(3):169–172, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Humphreys:1983:BR

[Hum83] J. H. Humphreys. Book review. Linear and Multilinear Algebra, 14(1):111–112, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hunter:1983:CIN

[Hun83] D. B. Hunter. The construction of irreducible normalized representations of the general linear group. Linear and Multilin-

Hunter:1986:KIR

[Hun86] D. B. Hunter. The kernels of irreducible representations of the general linear group. Linear and Multilinear Algebra, 20 (1):17–26, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Haynsworth:1974:BR

[HUSG74] Emile Haynsworth, E. E. Underwood, Ralph L. Shively, and Robert Todd Gregory. Book reviews. Linear and Multilinear Algebra, 2(2):187–193, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hu:2017:ALE

[HUZ17] Xiaozhe Hu, John C. Urschel, and Ludmil T. Zikatanov. On the approximation of Laplacian eigenvalues in graph disaggregation. Linear Multilinear Algebra, 65(9):1805–1822, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huffman:1975:LGD

[HW75] W. C. Huffman and D. B. Wales. Linear groups of degree n containing an element with exactly n − 2 equal eigenvalues. Linear and Multilinear Algebra, 3(1–2):53–59, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hart:1988:MKS

[HW88] Sue O. Hart and S. Gill Williamson. Menus of Kuratowski subgraphs. Linear and Multilinear Algebra, 23(2):111–129, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2006:GHM

[HW06] Li-Ping Huang and Zhe-Xian Wan. Geometry of 2 × 2 Hermitian matrices II. Linear Multilinear Algebra, 54(1):37–54, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2008:DUE

[HW08] Yaoping Hou and Chingwah Woo. Distance unimodular equivalence of graphs. Linear Multilinear Algebra, 56(6):611–626,

He:2013:RQM

[HW13] Zhuo-Heng He and Qing-Wen Wang. A real quaternion matrix equation with applications. Linear Multilinear Algebra, 61(6): 725–740, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

He:2014:BSS

[HW14] Zhuo-Heng He and Qing-Wen Wang. The η-bihermitian solution to a system of real quaternion matrix equations. Linear Multilinear Algebra, 62(11):1509–1528, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

He:2015:GSS

[HW15] Zhuo-Heng He and Qing-Wen Wang. The general solutions to some systems of matrix equations. Linear Multilinear Algebra, 63(10):2017–2032, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:1985:MPF

[Hwa85] Suk Geun Hwang. Minimum permanent on faces of staircase type of the polytope of doubly stochastic matrices. Linear and Multilinear Algebra, 18(4):271–306, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:1991:CPD

[Hwa91] Suk-Geun Hwang. Convexity of the permanent for doubly stochastic matrices. Linear and Multilinear Algebra, 30(1–2): 129–134, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hemasinha:1992:NIS

[HWL92] Rohan Hemasinha, James R. Weaver, and Chi-Kwong Li. Norms induced by symmetric guage functions. Linear and Multilinear Algebra, 31(1–4):217–224, 1992. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2021:ESE

[HWL21] Jinhua Huang, Cong Wang, and Haohao Li. EA solutions and EA solvability to general interval linear systems. Linear Mul-

Huckle:2013:EMS

[HWSH13] Thomas K. Huckle, Konrad Waldherr, and Thomas Schulte-Herbrüggen. Exploiting matrix symmetries and physical symmetries in matrix product states and tensor trains. Linear Multilinear Algebra, 61(1):91–122, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Han:2022:HOE

[HWX22] Lixing Han, Kelun Wang, and Jianhong Xu. Higher order ergodic Markov chains and first passage times. Linear Multilinear Algebra, 70(21):6772–6779, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2017:NRT

[HWY17] Shaowu Huang, Qing-Wen Wang, and Yaoming Yu. A note on Rotfel’d Theorem in a sector. Linear Multilinear Algebra, 65(8):1655–1661, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

He:2018:CEC

[HWZ18a] Zhuo-Heng He, Qing-Wen Wang, and Yang Zhang. The complete equivalence canonical form of four matrices over an arbitrary division ring. Linear Multilinear Algebra, 66(1):74–95, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2018:ESR

[HWZ18b] Shaowu Huang, Qing-Wen Wang, and Yang Zhang. Equivalence on some Rotfel’d type theorems. Linear Multilinear Algebra, 66(8):1626–1632, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2021:PRU

[HX21] Jinchuan Hou and Qingsen Xu. Preservers of radial unitary similarity functions on Lie products of self-adjoint operators. Linear Multilinear Algebra, 69(15):2779–2805, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2016:SNI

[HXL16] Zhengge Huang, Zhong Xu, and Quan Lu. Some new inequalities for the Hadamard product of a nonsingular M-matrix and its inverse. Linear Multilinear Algebra, 64(7):1362–1378, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hwang:2003:PGC

[HZ03] Suk-Geun Hwang and Xiao-Dong Zhang. Permanents of graphs with cut vertices. Linear Multilinear Algebra, 51(4): 393–404, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hanusa:2011:DKP

[HZ11] Christopher R. H. Hanusa and Thomas Zaslavsky. Determinants in the Kronecker product of matrices: the incidence matrix of a complete graph. Linear Multilinear Algebra, 59(4): 399–411, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Horn:2012:GCA

[HZ12] Roger A. Horn and Fuzhen Zhang. A generalization of the complex Autonne–Takagi factorization to quaternion matrices. Linear Multilinear Algebra, 60(11–12):1239–1244, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2020:GHR

[HZ20] Li-Ping Huang and Kang Zhao. Graph homomorphisms on rectangular matrices over division rings II. Linear Multilinear Algebra, 68(2):391–409, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huo:2021:OGT

[HZ21] Lijun Huo and Gengsheng Zhang. Orthogonal graph of type (m, m−1, 0) of characteristic 2 and its subconstituents. Linear Multilinear Algebra, 69(4):681–701, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2014:LTE

[HZG14] Bo Hou, Liwei Zhang, and Suogang Gao. The Leonard triples extended from given Leonard pairs of Bannai/Ito type. Linear Multilinear Algebra, 62(8):1091–1104, 2014. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hou:2020:CLP

[HZH20] Bo Hou, Juan Zhao, and Lihang Hou. Classical Leonard pairs having LB-TD form. Linear Multilinear Algebra, 68(3):622– 634, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:RDS

[HZS22] Riguang Huang, Baodong Zheng, and Xiaoyu Song. Rank decomposition and symmetric rank decomposition over arbitrary fields. Linear Multilinear Algebra, 70(20):5888–5901, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

He:2023:MBM

[HZV23] Jiewen He, Hua Zheng, and Seakweng Vong. Modulus-based matrix splitting iteration methods with new splitting scheme for horizontal implicit complementarity problems. Linear Multilinear Algebra, 71(14):2392–2408, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huang:2022:NRI

[HZX22] Hong Huang, Zhi-Feng Zhu, and Guo-Jin Xu. Numerical radius inequalities for operator matrices. Linear Multilinear Algebra, 70(20):5362–5372, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Huo:2015:NMP

[HZXL15] Donghua Huo, Baodong Zheng, Jinli Xu, and Hongyu Liu. Nonlinear mappings preserving Jordan multiple ∗-product on factor von Neumann algebras. Linear Multilinear Algebra, 63 (5):1026–1036, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Issa:2023:IRN

[IAM23] Hassan Issa, Hassane Abbas, and Bassam Mourad. On an integral representation of the normalized trace of the k-th symmetric tensor power of matrices and some applications. Linear Multilinear Algebra, 71(7):1198–1216, 2023. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ibragimov:2000:CPD

[Ibr00] Marat Ibragimov. Conditions for positive definiteness of Mmatrices. Linear and Multilinear Algebra, 48(2):93–106, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Iglesias:1999:RDE

[Igl99] Pablo A. Iglesias. On the Riccati difference equation of optimal control. Linear and Multilinear Algebra, 46(4):313–326, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ilmonen:2011:SMS

[IH11] Pauliina Ilmonen and Pentti Haukkanen. Smith meets Smith: Smith normal form of Smith matrix. Linear Multilinear Algebra, 59(5):557–564, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ilic:2018:ILC

[IK18] D. S. Cvetković Ilić and M. Kostadinov. Invertibility of linear combinations in B(H). Linear Multilinear Algebra, 66(11): 2139–2150, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ilisevic:2019:SRI

[IK19] Dijana Ilisević and Bojan Kuzma. On square roots of isometries. Linear Multilinear Algebra, 67(9):1898–1921, 2019. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Izhakian:2020:QCQ

[IK20] Zur Izhakian and Manfred Knebusch. Quasilinear convexity and quasilinear stars in the ray space of a supertropical quadratic form. Linear Multilinear Algebra, 68(12):2347–2389, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Izhakian:2022:CSF

[IK22] Zur Izhakian and Manfred Knebusch. Cauchy–Schwarz functions and convex partitions in the ray space of a supertropical quadratic form. Linear Multilinear Algebra, 70(20):5502–5546,

Izhakian:2012:DSB

[IKR12] Zur Izhakian, Manfred Knebusch, and Louis Rowen. Dual spaces and bilinear forms in supertropical linear algebra. Linear Multilinear Algebra, 60(7):865–883, 2012. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ikramov:2015:NCB

[Ikr15a] Khakim D. Ikramov. Normality conditions for the BHH matrix operator. Linear Multilinear Algebra, 63(9):1901–1908, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ikramov:2015:TPH

[Ikr15b] Khakim D. Ikramov. Toeplitz-plus-Hankel circulants are reducible to block diagonal form via unitary congruences. Linear Multilinear Algebra, 63(4):862–867, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ilkhan:2020:NCM

[Ilk20] Merve Ilkhan. A new conservative matrix derived by Catalan numbers and its matrix domain in the spaces c and c_{0}. Linear Multilinear Algebra, 68(2):417–434, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ilmonen:2016:MHT

[Ilm16] Pauliina Ilmonen. On meet hypermatrices and their eigenvalues. Linear Multilinear Algebra, 64(5):842–855, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ioppolo:2018:VAP

[IM18] Antonio Ioppolo and Fabrizio Martino. Varieties of algebras with pseudoinvolution and polynomial growth. Linear Multilinear Algebra, 66(11):2286–2304, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ito:2015:MMM

[INO15] Hiroya Ito, Atsushi Noma, and Masahiro Ohno. Maximal minors of a matrix with linear form entries. Linear Multilinear Algebra, 63(8):1599–1606, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Izhakian:2018:SS

[INR18] Zur Izhakian, Adi Niv, and Louis Rowen. Supertropical SL_{n}. Linear Multilinear Algebra, 66(7):1461–1483, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Itza-Ortiz:2021:NRC

[IOMA21] Benjamín A. Itzá-Ortiz and Rubén A. Martínez-Avendaño. The numerical range of a class of periodic tridiagonal operators. Linear Multilinear Algebra, 69(5):786–806, 2021. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Itza-Ortiz:2021:NRS

[IOMAN21] Benjamín A. Itzá-Ortiz, Rubén A. Martínez-Avendaño, and Hiroshi Nakazato. The numerical range of some periodic tridiagonal operators is the convex hull of the numerical ranges of two finite matrices. Linear Multilinear Algebra, 69(15):2830– 2849, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ioppolo:2019:SCA

[Iop19] Antonio Ioppolo. Some characterizations of algebras with involution with polynomial growth of their codimensions. Linear Multilinear Algebra, 67(6):1217–1230, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Iovanov:2015:TMC

[Iov15] Miodrag Cristian Iovanov. Triangular matrix coalgebras and applications. Linear Multilinear Algebra, 63(1):46–67, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Iglesias:2015:SWP

[IS15] Laura Iglésias and Fernando C. Silva. Spectral and weak polynomial completeness for the product of nonsingular matrices. Linear Multilinear Algebra, 63(10):1937–1946, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Iglesias:2016:WSC

[IS16] Laura Iglésias and Fernando C. Silva. Weakly spectrally complete pair of matrices. Linear Multilinear Algebra, 64(5):942–

Iglesias:2011:UMM

[ISCS11] Laura Iglésias, Catarina Santa-Clara, and Fernando C. Silva. Unified min-max and interlacing theorems for linear operators. Linear Multilinear Algebra, 59(6):651–669, 2011. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ilisevic:2016:QVH

[IT16] Dijana Ilisević and Aleksej Turnsek. A quantitative version of Herstein’s theorem for Jordan *-isomorphisms. Linear Multilinear Algebra, 64(2):156–168, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Itoh:2016:TIM

[Ito16] Minoru Itoh. Twisted immanant and matrices with anticommuting entries. Linear Multilinear Algebra, 64(8):1637–1653, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ishikawa:1995:MSF

[IW95] Masao Ishikawa and Masato Wakayama. Minor summation formula of Pfaffians. Linear and Multilinear Algebra, 39(3): 285–305, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jaatit:2022:NLS

[Jaa22] Ali Jaatit. A note on local spectral subspace preservers. Linear Multilinear Algebra, 70(19):4146–4156, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jacobowitz:1979:COE

[Jac79] Howard Jacobowitz. Curvature operators on the exterior algebra. Linear and Multilinear Algebra, 7(2):93–105, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

James:1981:LTS

[Jam81] D. G. James. Linear transformations of the second elementary function. Linear and Multilinear Algebra, 10(4):347–349, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

James:1988:PGR

[Jam88] Donald G. James. Projective geometry over rings with stable range condition. Linear and Multilinear Algebra, 23(4):299– 304, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

James:1992:I

[Jam92] Gordon James. Immanants. Linear and Multilinear Algebra, 32(3–4):197–210, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

James:1995:PSR

[Jam95] Donald G. James. Polar spaces over rings with absolute stable range condition. Linear and Multilinear Algebra, 38(4):373– 378, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ao:2015:MRF

[jAS15] Ji jun Ao and Jiong Sun. Matrix representations of fourthorder boundary value problems with coupled or mixed boundary conditions. Linear Multilinear Algebra, 63(8):1590–1598, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ao:2021:ECF

[jAS21] Ji jun Ao and Jiong Sun. Eigenvalues of a class of fourth-order boundary value problems with transmission conditions using matrix theory. Linear Multilinear Algebra, 69(9):1610–1624, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1974:RP

[JC74] Charles R. Johnson and D. Carlson. Research problem. Linear and Multilinear Algebra, 2(2):185, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jafari:2022:SPP

[JC22a] Sayyed Heidar Jafari and Sriparna Chattopadhyay. Spectrum of proper power graphs of the direct product of certain finite groups. Linear Multilinear Algebra, 70(20):5460–5481, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Joia:2022:OFR

[JC22b] Maria Joia and Tania Luminia Costache. Operator factorization of range space relations*. Linear Multilinear Algebra, 70 (19):4397–4404, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1999:MME

[JD99] C. R. Johnson and António Leal Duarte. The maximum multiplicity of an eigenvalue in a matrix whose graph is a tree. Linear and Multilinear Algebra, 46(1–2):139–144, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jana:2021:MHM

[JDD21] R. Jana, A. Dutta, and A. K. Das. More on hidden Z-matrices and linear complementarity problem. Linear Multilinear Algebra, 69(6):1151–1160, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jabbarzadeh:2021:DCO

[JE21] M. R. Jabbarzadeh and H. Emamalipour. On the dilation of a conditional operator. Linear Multilinear Algebra, 69(12): 2204–2219, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jain:1979:NMH

[JGK79] S. K. Jain, V. K. Goel, and Edward K. Kwak. Nonnegative matrices having same nonnegative Moore–Penrose and group inverses. Linear and Multilinear Algebra, 7(1):59–72, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jain:1990:WMP

[JGS90] S. K. Jain, A. D. Gunawardena, and L. Snyder. On weak monotonicity of the powers of nonnegative matrices. Linear and Multilinear Algebra, 26(3):223–230, 1990. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jiao:1992:CHM

[Jia92] Zhengming Jiao. On a conjecture of H. Minc. Linear and Multilinear Algebra, 32(2):103–105, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jin:2017:LER

[Jin17] Hongwei Jin. The location of H-eigenvalues of real even order symmetry tensors. Linear Multilinear Algebra, 65(3):623–634, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1995:EDC

[JJK95] Charles R. Johnson, Catherine A. Jones, and Brenda K. Kroschel. The Euclidean distance completion problem: cycle completability. Linear and Multilinear Algebra, 39(1–2): 195–207, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jocic:2010:SLM

[JK10] Danko R. Jocić and Dorde Krtinić. Schur–Laurent multipliers for block matrices and geometric characterization of continuous matrices. Linear Multilinear Algebra, 58(4):523–534, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jabbarzadeh:2022:IAD

[JK22] Mohammad Reza Jabbarzadeh and Venus Kaleibary. Inequalities for accretive-dissipative block matrices involving convex and concave functions. Linear Multilinear Algebra, 70(3):395– 410, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jin:2017:LUB

[JKB17] Hongwei Jin, M. Rajesh Kannan, and Minru Bai. Lower and upper bounds for H-eigenvalues of even order real symmetric tensors. Linear Multilinear Algebra, 65(7):1402–1416, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jeribi:2016:ESS

[JKD16] Aref Jeribi, Bilel Krichen, and Mohammed Zarai Dhahri. Essential spectra of some matrix operators involving γ-relatively bounded entries and an application. Linear Multilinear Algebra, 64(8):1654–1668, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jalaeian:2020:CPO

[JKH20] Maryam Jalaeian, Mehdi Mohammadzadeh Karizaki, and Mahmoud Hassani. Conditions that the product of operators is an EP operator in Hilbert c*-module. Linear Multilinear Algebra, 68(10):1990–2004, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jung:2022:EPD

[JKKC22] Ji-Hwan Jung, Suh-Ryung Kim, Bumtle Kang, and Gi-Sang Cheon. Exponents of primitive directed Toeplitz graphs. Linear Multilinear Algebra, 70(8):1571–1589, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jocic:2022:EAG

[JKL22] Danko R. Jocić, Dorde Krtinić, and Milan Lazarević. Extensions of the arithmetic-geometric means and Young’s norm inequalities to accretive operators, with applications. Linear Multilinear Algebra, 70(20):4835–4875, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1979:CEN

[JKS79] Charles R. Johnson, R. B. Kellogg, and A. B. Stephens. Complex eigenvalues of a non-negative matrix with a specified graph II. Linear and Multilinear Algebra, 7(2):129–143, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See correction of Figure 6 [Ano79b].

Johnson:1994:SM

[JKS94] Charels R. Johnson, M. Kathleen Kerr, and David P. Stanford. Semipositivity of matrices. Linear and Multilinear Algebra, 37(4):265–271, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jung:2023:CPM

[JKY23] Ji-Hwan Jung, Suh-Ryung Kim, and Hyesun Yoon. Competition periods of multipartite tournaments. Linear Multilinear Algebra, 71(4):586–596, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jeribi:2020:PDO

[JKZ20] Aref Jeribi, Bilel Krichen, and Ali Zitouni. Properties of demicompact operators, essential spectra and some perturbation results for block operator matrices with applications. Linear Multilinear Algebra, 68(12):2506–2522, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1985:ELP

[JL85] Charles R. Johnson and David London. On an eigenvalue localization problem. Linear and Multilinear Algebra, 17(2): 165–168, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1988:IRU

[JL88] Charles R. Johnson and Chi-Kwong Li. Inequalities relating unitarily invariant norms and the numerical radius. Linear and Multilinear Algebra, 23(2):183–191, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1990:MW

[JL90] Charles R. Johnson and Chi-Kwong Li. Matrices for which 0 ∉ F (A^{m}). Linear and Multilinear Algebra, 27(1):57–62, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1991:PIH

[JL91] Charles R. Johnson and Raphael Loewy. Possible inertias for the Hermitian part of A + BX. Linear and Multilinear Algebra, 28(4):225–227, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1993:MCI

[JL93] Charles R. Johnson and Michael Lundquist. Matrices with chordal inverse zero-patterns. Linear and Multilinear Algebra, 36(1):1–17, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jing:2012:LDM

[JL12] Wu Jing and Fangyan Lu. Lie derivable mappings on prime rings. Linear Multilinear Algebra, 60(2):167–180, 2012. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jamilov:2022:EAD

[JL22] Uygun Jamilov and Manuel Ladra. Evolution algebras and dynamical systems of a worm propagation model. Linear Mul-

Johnson:2012:CEM

[JLDS12] Charles R. Johnson, António Leal-Duarte, and Carlos M. Saiago. The change in eigenvalue multiplicity associated with perturbation of a diagonal entry. Linear Multilinear Algebra, 60(5):525–532, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1988:LTP

[JLL88] Charles R. Johnson, Thomas Laffey, and Chi-Kwong Li. Linear transformations on that preserve the Ky Fan k-norm and a remarkable special case when (n, k) = (4, 2). Linear and Multilinear Algebra, 23(4):285–298, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ji:2007:ENR

[JLL07] Guoxing Ji, Ni Liu, and Ze Li. Essential numerical range and maximal numerical range of the Aluthge transform. Linear Multilinear Algebra, 55(4):315–322, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jiang:2011:FRM

[JLL11] Xiaoyan Jiang, Yingluan Liu, and Bolian Liu. A further result on majorization theorem. Linear Multilinear Algebra, 59(9): 957–967, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jiao:2018:STS

[JLL18] Aiquan Jiao, Chaoqian Li, and Yaotang Li. SDB-tensors and SQB-tensors. Linear Multilinear Algebra, 66(10):2107–2118, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1992:LMP

[JLR^{+}92] Charles R. Johnson, Chi-Kwong Li, Leiba Rodman, Ilya Spitkovsky, and Stephen Pierce. Linear maps preserving regional eigenvalue location. Linear and Multilinear Algebra, 32(3–4):253–264, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2009:GWM

[JLS09] Charles R. Johnson, Raphael Loewy, and Paul Anthony Smith. The graphs for which the maximum multiplicity of an eigenvalue is two. Linear Multilinear Algebra, 57(7):713–736, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ji:2012:NLT

[JLZ12] Peisheng Ji, Rongrong Liu, and Yingzi Zhao. Nonlinear Lie triple derivations of triangular algebras. Linear Multilinear Algebra, 60(10):1155–1164, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1977:BRb

[JM77] Charles R. Johnson and Marvin Marcus. Book reviews. Linear and Multilinear Algebra, 5(2):149–152, 1977. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1985:CHP

[JM85] Charles R. Johnson and Thomas L. Markham. Compression and Hadamard power inequalities. Linear and Multilinear Algebra, 18(1):23–34, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jena:2013:CSS

[JM13] Litismita Jena and Debasisha Mishra. Comparisons of B_{row}- splittings and B_{ran}-splittings of matrices. Linear Multilinear Algebra, 61(1):35–48, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2017:SNR

[JMP17] C. R. Johnson, C. Marijuán, and M. Pisonero. Symmetric nonnegative realizability via partitioned majorization. Linear Multilinear Algebra, 65(7):1417–1426, 2017. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Julio:2015:NNR

[JMS15] Ana I. Julio, Cristina B. Manzaneda, and Ricardo L. Soto. Normal nonnegative realization of spectra. Linear Multilinear Algebra, 63(6):1204–1215, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jiang:2019:AIF

[JMS19] Tianpei Jiang, Mehdi S. Mousavi, and Hristo S. Sendov. On the analyticity of k-isotropic functions. Linear Multilinear Algebra, 67(2):228–247, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1980:SRP

[JN80] Charles R. Johnson and Michael Neumann. Square roots with positive definite Hermitian part. Linear and Multilinear Algebra, 8(4):353–355, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1990:CES

[JN90a] Charles R. Johnson and Peter Nylen. Cases of equality for the spectral norm submultiplicativity of the Hadamard product. Linear and Multilinear Algebra, 27(2):121–127, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1990:CSV

[JN90b] Charles R. Johnson and Peter M. Nylen. Converse to a singular value inequality common to the Hadamard and conventional products. Linear and Multilinear Algebra, 27(3):167– 187, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1992:UIN

[JN92] Charles R. Johnson and Peter Nylen. Unitarily invariant norm submultiplicativity. Linear and Multilinear Algebra, 31(1–4): 285–301, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2000:HPS

[JN00] Charles R. Johnson and Peter M. Nylen. Hadamard product submultiplicativity of certain induced norms. Linear and Multilinear Algebra, 48(2):165–178, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2014:CST

[JN14a] C. R. Johnson and S. Nasserasr. Complete solution to the TP_{2} completion problem. Linear Multilinear Algebra, 62(10): 1403–1420, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2014:KIM

[JN14b] Charles R. Johnson and Sivaram K. Narayan. The Koteljanskii inequalities for mixed matrices. Linear Multilinear Algebra, 62(12):1583–1590, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johari:2022:CCF

[JN22] Farangis Johari and Peyman Niroomand. The capability and certain functors of some nilpotent lie algebras of class two. Linear Multilinear Algebra, 70(21):7189–7203, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jones:1993:IPG

[JNBR93] M. S. Jones, B. Nashvadian-Bakhsh, and H. P. Rogosinski. Intersection properties for generalized numerical ranges. Linear and Multilinear Algebra, 34(3–4):377–383, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2023:NVD

[JNS23] Charles Johnson, Stephen Newman, and Ilya Spitkovsky. New views of the doubly stochastic single eigenvalue problem. Linear Multilinear Algebra, 71(9):1462–1469, 2023. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1996:CPD

[JNT96] Charles R. Johnson, Michael Neumann, and Michael J. Tsatsomeros. Conditions for the positivity of determinants. Linear and Multilinear Algebra, 40(3):241–248, 1996. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jumaniyozov:2023:NEI

[JO23] Doston Jumaniyozov and Bakhrom Omirov. New examples of infinite-dimensional n-Lie algebras. Linear Multilinear Algebra, 71(2):206–225, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnsen:1973:EDS

[Joh73] E. C. Johnsen. Essentially doubly stochastic matrices. Linear and Multilinear Algebra, 1(1):33–45, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnsen:1974:EDS

[Joh74a] E. C. Johnsen. Essentially doubly stochastic matrices. Linear and Multilinear Algebra, 1(4):343–356, 1974. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1974:HPM

[Joh74b] Charles R. Johnson. Hadamard products of matrices. Linear and Multilinear Algebra, 1(4):295–307, 1974. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1975:NNH

[Joh75a] Charles R. Johnson. Notes on nonnegative Hadamard functions. Linear and Multilinear Algebra, 3(3):187–192, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1975:NLF

[Joh75b] Charles R. Johnson. Numerical location of the field of values. Linear and Multilinear Algebra, 3(1–2):9–14, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnsen:1977:IMS

[Joh77a] E. C. Johnsen. An inequality on the matrix spectral norm specifying Hadamard matrices. Linear and Multilinear Algebra, 4(4):277–279, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1977:HPI

[Joh77b] Charles R. Johnson. A Hadamard product involving Mmatrices. Linear and Multilinear Algebra, 4(4):261–264, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1981:RSM

[Joh81a] Charles R. Johnson. Row stochastic matrices similar to doubly stochastic matrices. Linear and Multilinear Algebra, 10(2): 113–130, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1981:RES

[Joh81b] E. C. Johnson. Real essentially stochastic matrices: factorizations into special elementary matrices. Linear and Multilinear

Johnson:1982:BDM

[Joh82a] Charles R. Johnson. Bayesian deduction of marginal from conditional probabilities. Linear and Multilinear Algebra, 11(4): 333–350, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1982:SOP

[Joh82b] Charles R. Johnson. Some outstanding problems in the theory of matrices. Linear and Multilinear Algebra, 12(2):99–108, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1983:BR

[Joh83] Charles R. Johnson. Book review. Linear and Multilinear Algebra, 13(2):179–181, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1984:LRP

[Joh84] Charles R. Johnson. The least real part eigenvalue and principal submatrices. Linear and Multilinear Algebra, 16(1–4): 175–177, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1989:RP

[Joh89] Charles R. Johnson. Research problem. Linear and Multilinear Algebra, 25(4):289–290, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1990:RTI

[Joh90] Charles R. Johnson. The Robertson–Taussky inequality revisited. Linear and Multilinear Algebra, 26(1–2):5–8, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See erratum [JZ95].

Johnston:2011:COP

[Joh11] Nathaniel Johnston. Characterizing operations preserving separability measures via linear preserver problems. Linear Multilinear Algebra, 59(10):1171–1187, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnston:2014:NDC

[Joh14] Nathaniel Johnston. Norm duality and the cross norm criteria for quantum entanglement. Linear Multilinear Algebra, 62(5): 648–658, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johari:2022:BDC

[Joh22] Farangis Johari. Bounds for the dimension of the center factor in capable nilpotent Lie algebras of class two. Linear Multilinear Algebra, 70(21):6412–6424, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Joni:1978:LIH

[Jon78] S. A. Joni. Lagrange inversion in higher dimensions and umbral operators. Linear and Multilinear Algebra, 6(2):111–122, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jones:1992:NSG

[Jon92] Mark S. Jones. A note on the shape of the generalized cnumerical range. Linear and Multilinear Algebra, 31(1–4):81– 84, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1994:ID

[JOTvdD94] Charles R. Johnson, D. D. Olesky, Michael Tsatsomeros, and P. van den Driessche. An identity for the determinant. Linear and Multilinear Algebra, 36(3):189–194, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1984:MPH

[JOvdD84] Charles R. Johnson, D. D. Olesky, and P. van den Driessche. M-matrix products having positive principal minors. Linear and Multilinear Algebra, 16(1–4):29–38, 1984. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1985:SMP

[JOvdD85] Charles R. Johnson, D. D. Olesky, and P. van den Driessche. Stability of m-matrix products. Linear and Multilinear Algebra, 18(1):67–76, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Joziak:2015:CSN

[Józ15] Pawel Józiak. Conditionally strictly negative definite kernels. Linear Multilinear Algebra, 63(12):2406–2418, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1986:LMH

[JP86] Charles R. Johnson and Stephen Pierce. Linear maps on Hermitian matrices: the stabilizer of an inertia class, II. Linear and Multilinear Algebra, 19(1):21–31, 1986. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1987:PDN

[JP87] Charles R. Johnson and Stephen Pierce. Permanental dominance of the normalized single-hook immanants on the positive semi-definite matrices. Linear and Multilinear Algebra, 21(3): 215–229, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2007:MPV

[JP07] C. R. Johnson and J. M. Peña. Matrices that preserve vectors of fixed sign variation. Linear Multilinear Algebra, 55(6):521– 533, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jakobsen:2015:QMA

[JP15a] Hans Plesner Jakobsen and Chiara Pagani. Quantized matrix algebras and quantum seeds. Linear Multilinear Algebra, 63(4):713–753, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jamiolkowski:2015:GSC

[JP15b] Andrzej Jamiolkowski and Grzegorz Pastuszak. Generalized Shemesh criterion, common invariant subspaces and irreducible completely positive superoperators. Linear Multilinear Algebra, 63(2):314–325, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2017:RCP

[JP17] Charles R. Johnson and Pietro Paparella. Row cones, Perron similarities, and nonnegative spectra. Linear Multilinear Algebra, 65(10):2124–2130, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jiao:2015:AMP

[JQ15] Meiyan Jiao and Xiaofei Qi. Additive maps preserving commutativity up to a factor on nest algebras. Linear Multilinear Algebra, 63(6):1242–1256, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ji:2013:CLD

[JQS13] Peisheng Ji, Weiqing Qi, and Xiaolu Sun. Characterizations of Lie derivations of factor von Neumann algebras. Linear Multilinear Algebra, 61(3):417–428, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1984:IPC

[JR84] Charles R. Johnson and Leiba Rodman. Inertia possibilities for completions of partial Hermitian matrices. Linear and Multilinear Algebra, 16(1–4):179–195, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jefferies:1993:ACS

[JR93a] B. R. F. Jefferies and W. J. Ricker. Addendum: Commutativity for systems of (2 × 2) selfadjoint matrices. Linear and Multilinear Algebra, 35(3–4):357, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [JR93b].

Jefferies:1993:CSS

[JR93b] B. R. F. Jefferies and W. J. Ricker. Commutativity for systems of (2 × 2) selfadjoint matrices. Linear and Multilinear Algebra, 35(2):107–113, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See addendum [JR93a].

Jones:1993:IRN

[JR93c] M. S. Jones and H. P. Rogosinski. Inclusion relations for the c-numerical range. Linear and Multilinear Algebra, 34(3–4): 279–290, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jones:1994:IRI

[JR94] M. S. Jones and H. P. Rogosinski. Inclusion relations involving the k-numerical range and the essential numerical range. Linear and Multilinear Algebra, 37(1–3):161–173, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2009:SSM

[JR09] Charles R. Johnson and Robert Reams. Scaling of symmetric matrices by positive diagonal congruence. Linear Multilinear Algebra, 57(2):123–140, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jimenez-Rodriguez:2016:SVC

[JRMFMASS16] P. Jiménez-Rodríguez, G. A. Muñoz-Fernández, M. Murillo-Arcila, and J. B. Seoane-Sepúlveda. Sharp values for the constants in the polynomial Bohnenblust–Hille inequality. Linear Multilinear Algebra, 64(9):1731–1749, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jain:1984:LSH

[JS84] S. K. Jain and L. E. Snyder. Linear systems having nonnegative least squares solution. Linear and Multilinear Algebra, 15(2):121–132, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1991:IRT

[JS91] Charles R. Johnson and Tamir Shalom. Information requirements for Toeplitz inversion. Linear and Multilinear Algebra, 29(2):99–111, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jafarian:1994:LMP

[JS94] A. A. Jafarian and A. R. Sourour. Linear maps that preserve the commutant, double commutant or the lattice of invariant subspaces. Linear and Multilinear Algebra, 38(1–2):117–129, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jain:1996:NFM

[JS96] S. K. Jain and L. E. Snyder. A note on factoring 0–1 matrices. Linear and Multilinear Algebra, 41(3):193–198, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1999:PPM

[JS99] Charles R. Johnson and Ronald L. Smith. Path product matrices. Linear and Multilinear Algebra, 46(3):177–191, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2008:BDC

[JS08] Charles R. Johnson and Carlos M. Saiago. Branch duplication for the construction of multiple eigenvalues in an Hermitian matrix whose graph is a tree. Linear Multilinear Algebra, 56(4):357–380, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jose:2015:POH

[JS15] Shani Jose and K. C. Sivakumar. On partial orders of Hilbert space operators. Linear Multilinear Algebra, 63(7):1423–1441, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2016:DMT

[JS16] Charles R. Johnson and Carlos M. Saiago. Diameter minimal trees. Linear Multilinear Algebra, 64(3):557–571, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jabbarzadeh:2019:LCT

[JS19] M. R. Jabbarzadeh and M. H. Sharifi. Lambert conditional type operators on L^{2}(Σ). Linear Multilinear Algebra, 67(10): 2030–2047, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jamshidi:2020:NNI

[JS20] Arya Jamshidi and Fereshteh Sady. A note on nonlinear isometries between vector-valued function spaces. Linear Multilinear Algebra, 68(2):293–303, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jahanbani:2022:SRA

[JS22] A. Jahanbani and S. M. Sheikholeslami. On the spectral radius of the adjacency matrix and signless Laplacian matrix of a graph. Linear Multilinear Algebra, 70(21):6846–6851, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1990:ENB

[JSE90] Charles R. Johnson, Erik A. Schreiner, and Ludwig Elsner. Eigenvalue neutrality in block triangular matrices. Linear and

Johnson:1994:IIN

[JSG94] Charles R. Johnson, I. M. Spitkovsky, and S. Gottlieb. Inequalities involving the numerical radius. Linear and Multilinear Algebra, 37(1–3):13–24, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2018:CVE

[JST18] Charles R. Johnson, Carlos M. Saiago, and Kenji Toyonaga. Classification of vertices and edges with respect to the geometric multiplicity of an eigenvalue in a matrix, with a given graph, over a field. Linear Multilinear Algebra, 66(11):2168– 2182, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2009:ICM

[JSW09] Charles R. Johnson, Brian D. Sutton, and Andrew J. Witt. Implicit construction of multiple eigenvalues for trees. Linear Multilinear Algebra, 57(4):409–420, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2014:STS

[JSY14] Charles R. Johnson, Helena Smigoc, and Dian Yang. Solution theory for systems of bilinear equations. Linear Multilinear Algebra, 62(12):1553–1566, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1995:CSN

[JT95] Charles R. Johnson and Michael J. Tsatsomeros. Convex sets of nonsingular and p:-matrices. Linear and Multilinear Algebra, 38(3):233–239, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jain:2003:NRF

[JT03] S. K. Jain and John Tynan. Nonnegative rank factorization of a nonnegative matrix A with AA ≥ 0. Linear Multilinear Algebra, 51(1):83–95, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:2004:BRN

[JT04] Charles R. Johnson and Pablo Tarazaga. Binary representation of normalized symmetric and correlation matrices. Linear Multilinear Algebra, 52(5):359–366, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jankauskas:2023:RMD

[JT23] Jonas Jankauskas and Jörg M. Thuswaldner. Rational matrix digit systems. Linear Multilinear Algebra, 71(10):1606–1639, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1977:BRa

[JTH77] Charles R. Johnson, R. P. Twearson, and Marshall Hall Jr. Book reviews. Linear and Multilinear Algebra, 5(1):71–75, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jordan:2000:CCP

[JTU00] C. Jordán, J. R. Torregrosa, and A. Urbano. Controllability completion problems of partial upper triangular matrices. Linear and Multilinear Algebra, 47(1):57–75, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jiang:2017:ICT

[JTW17] Zhaolin Jiang, Tin-Yau Tam, and Yanfeng Wang. Inversion of conjugate-Toeplitz matrices and conjugate-Hankel matrices. Linear Multilinear Algebra, 65(2):256–268, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Julius:2021:MPL

[Jul21] Hayden Julius. On maps preserving Lie products equal to e_{11} − e_{22}. Linear Multilinear Algebra, 69(14):2620–2628, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jun:1988:TIS

[Jun88] Wu Jun. Theorems on the index of a symmetry class of tensors. Linear and Multilinear Algebra, 23(3):213–225, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1988:ISS

[JvdD88] Charles R. Johnson and P. van den Driessche. Interpolation of D-stability and sign stability. Linear and Multilinear Algebra, 23(4):363–368, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jeffries:1991:QSS

[JvdD91] Clark Jeffries and P. van den Driessche. Qualitative stability and solvability of difference equations. Linear and Multilinear Algebra, 30(4):275–282, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jimenez-Vargas:2019:MUP

[JVMCPR19] Antonio Jiménez-Vargas, Antonio Morales-Campoy, Antonio M. Peralta, and María Isabel Ramírez. The Mazur–Ulam property for the space of complex null sequences. Linear Multilinear Algebra, 67(4):799–816, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1991:MRC

[JW91] Charles R. Johnson and Glen T. Whitney. Minimum rank completions. Linear and Multilinear Algebra, 28(4):271–273, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ji:2022:CFT

[JW22] Jun Ji and Long Wang. Construction of finite tight frames via singular value decomposition. Linear Multilinear Algebra, 70(19):4558–4567, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ji:2022:SRM

[JWBB22] Yuqing Ji, Jianfeng Wang, Maurizio Brunetti, and Ning Bian. A_{α}-spectral radius and measures of graph irregularity. Linear Multilinear Algebra, 70(20):5247–5267, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1998:SPO

[JWP98] Charles R. Johnson, Charles Waters, and S. Pierce. Sign patterns occurring in orthogonal matrices. Linear and Multilinear Algebra, 44(4):287–299, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jin:2013:IOA

[JX13] Shuilin Jin and Li Xu. Independences of operator algebras and joint spectrum of operators. Linear Multilinear Algebra, 61(5): 565–572, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xie:2015:SAS

[jXL15] Ze jia Xie and Wen Li. Sensitivity analysis for the SR decomposition. Linear Multilinear Algebra, 63(2):222–234, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Johnson:1995:ERT

[JZ95] Charles R. Johnson and Fuzhen Zhang. Erratum: “The Robertson–Taussky inequality revisited”. Linear and Multilinear Algebra, 38(3):281–282, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Joh90].

Jin:2022:NME

[JZ22] Zhixiang Jin and Chengbo Zhai. On the nonlinear matrix equation X^{p} = A + ∑_{i=1}^{m} M_{i}^{∗}(B + X − 1)^{−1}M_{i}. Linear Multilinear Algebra, 70(19):4467–4482, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jiao:2020:STA

[JZL20] Aiquan Jiao, Jianxing Zhao, and Yaotang Li. SGB-tensors and its applications. Linear Multilinear Algebra, 68(6):1113–1128, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kilic:2019:NGF

[KA19] Emrah Kiliç and Talha Arikan. A nonlinear generalization of the Filbert matrix and its Lucas analogue. Linear Multilinear Algebra, 67(1):141–157, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2010:SNT

[KADM10] Stephen Kirkland, Maria Aguieiras Alvarez de Freitas, Renata Raposo Del Vecchio, and Nair Maria Maia de Abreu. Split non-threshold Laplacian integral graphs. Linear Multilinear Algebra, 58(2):221–233, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kurmanbek:2022:PAF

[KAE22] Bakytzhan Kurmanbek, Yerlan Amanbek, and Yogi Erlangga. A proof of Andelić–Fonseca conjectures on the determinant of some Toeplitz matrices and their generalization. Linear Multilinear Algebra, 70(8):1563–1570, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kan:1987:FEM

[Kan87] Charn-Huen Kan. A family of 2 × 2 extreme matrices and a generalization of Clarkson’s inequalities. Linear and Multilinear Algebra, 21(2):191–199, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kanter:1994:SIN

[Kan94] Marek Kanter. Separating inequalities for non-negative covariances. Linear and Multilinear Algebra, 38(1–2):151–162, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kandic:2016:APC

[Kan16] Marko Kandić. On algebras of polynomially compact operators. Linear Multilinear Algebra, 64(6):1185–1196, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kapoor:1977:MWU

[Kap77] J. Kapoor. Matrices which, under row permutations, give specified values of trace function. Linear and Multilinear Algebra, 5(2):129–137, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaplansky:1980:BR

[Kap80] I. Kaplansky. Book review. Linear and Multilinear Algebra, 9(2):165–166, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karanam:1983:LBS

[Kar83] V. R. Karanam. On lower bounds for singular values of matrix sums. Linear and Multilinear Algebra, 13(2):101–103, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karabeg:1990:CDO

[Kar90] Almira Karabeg. Classification and detection of obstructions to planarity. Linear and Multilinear Algebra, 26(1–2):15–38, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karakas:2023:SSI

[Kar23] Murat Karakas. On the sequence spaces involving bell numbers. Linear Multilinear Algebra, 71(14):2298–2309, 2023. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaschube:1985:DWG

[Kas85] Paul A. Kaschube. Determining when a graphic matroid is transversal in linear time. Linear and Multilinear Algebra, 17 (1):41–83, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kato:1975:MLT

[Kat75] Tosio Kato. On a matrix limit theorem. Linear and Multilinear Algebra, 3(1–2):67–71, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaygorodov:2017:AJB

[Kay17] Ivan Kaygorodov. Algebras of Jordan brackets and generalized Poisson algebras. Linear Multilinear Algebra, 65(6):1142–1157, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kang:2016:UPL

[KC16] Na Kang and Liangyun Chen. Uniform posets and Leonard pairs based on unitary spaces over finite fields. Linear Multilinear Algebra, 64(6):1163–1184, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:1987:OBS

[KCBR87] Ohoe Kim, John Chollet, Ralf Brown, and David Rauschenberg. Orthonormal bases of symmetry classes with computergenerated examples. Linear and Multilinear Algebra, 21(1): 91–106, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ke:2018:NRI

[KCICV18] Yuanyuan Ke, Dragana S. Cvetković-Ilić, Jianlong Chen, and Jelena Visnjić. New results on (b, c)-inverses. Linear Multilinear Algebra, 66(3):447–458, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ke:2021:CRO

[KCSĆ21] Yuanyuan Ke, Jianlong Chen, Predrag Stanimirović, and Miroslav Ćirić. Characterizations and representations of outer inverse for matrices over a ring. Linear Multilinear Algebra, 69(1):155–176, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kian:2016:DFO

[KD16] Mohsen Kian and S. S. Dragomir. f-divergence functional of operator log-convex functions. Linear Multilinear Algebra, 64(2):123–135, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kuang:2023:CPA

[KD23] Ruifei Kuang and Chunyuan Deng. Common properties among various generalized inverses and constrained binary relations. Linear Multilinear Algebra, 71(8):1295–1322, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karizaki:2022:SRA

[KDHJ22] M. Mohammadzadeh Karizaki, D. S. Djordjević, A. Hosseini, and M. Jalaeian. Some results about EP modular operators. Linear Multilinear Algebra, 70(18):3490–3496, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See erratum [Ano22c].

Kotin:1982:MWC

[KE82] Leon Kotin and Irving J. Epstein. On matrices which commute with their derivatives. Linear and Multilinear Algebra, 12(1): 57–72, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kegel:1987:BR

[Keg87] O. H. Kegel. Book review. Linear and Multilinear Algebra, 21(1):107–108, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kestenband:1983:PFA

[Kes83] Barbu C. Kestenband. Partitions of finite affine geometries into caps. Linear and Multilinear Algebra, 14(3):257–270, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:1998:PCA

[KF98] Steve Kirkland and Shaun Fallat. Perron components and algebraic connectivity for weighted graphs. Linear and Multilinear Algebra, 44(2):131–148, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Khan:2017:CAM

[KFR17] Mehtab Khan, Rashid Farooq, and Juan Rada. Complex adjacency matrix and energy of digraphs. Linear Multilinear Algebra, 65(11):2170–2186, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [KFR20].

Khan:2020:CCA

[KFR20] Mehtab Khan, Rashid Farooq, and Juan Rada. Corrigendum to ‘Complex adjacency matrix and energy of digraphs’. Linear Multilinear Algebra, 68(1):220–222, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [KFR17].

Korkee:2005:GFM

[KH05] Ismo Korkee and Pentti Haukkanen. On a general form of meet matrices associated with incidence functions. Linear Multilinear Algebra, 53(5):309–321, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kang:2013:IPS

[KHG13] Na Kang, Bo Hou, and Suogang Gao. q-inverting pairs of shape 1, 2, 1. Linear Multilinear Algebra, 61(10):1307–1328, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kang:2015:LTK

[KHG15] Na Kang, Bo Hou, and Suogang Gao. Leonard triples of Krawtchouk type. Linear Multilinear Algebra, 63(12):2341– 2358, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Khrypchenko:2016:JDF

[Khr16] Mykola Khrypchenko. Jordan derivations of finitary incidence rings. Linear Multilinear Algebra, 64(10):2104–2118, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kara:2016:SPG

[KI16] Emrah Evren Kara and Merve Ilkhan. Some properties of generalized Fibonacci sequence spaces. Linear Multilinear Algebra, 64(11):2208–2223, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kimelfeld:1991:IST

[Kim91] Boris Kimelfeld. On inequalities of Schur type for Hermitian matrices and characters of unitary groups. Linear and Multilinear Algebra, 28(4):265–270, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:1997:CM

[Kim97] Si-Ju Kim. On convertible (0, 1) matrices. Linear and Multilinear Algebra, 42(4):319–322, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2013:NTS

[Kim13] Ik-Pyo Kim. Nearly tight sign-central matrices. Linear Multilinear Algebra, 61(1):1–14, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:1991:HMS

[Kir91] Steve Kirkland. Hypertournament matrices, score vectors and eigenvalues. Linear and Multilinear Algebra, 30(4):261–274, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:1993:RIT

[Kir93] Steve Kirkland. A reversal index for tournament matrices. Linear and Multilinear Algebra, 34(3–4):343–351, 1993. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2000:BAC

[Kir00] Steve Kirkland. A bound on the algebraic connectivity of a graph in terms of the number of cutpoints. Linear and Multilinear Algebra, 47(1):93–103, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2004:CSI

[Kir04] Steve Kirkland. A characterization of spectral integral variation in two places for Laplacian matrices. Linear Multilinear Algebra, 52(2):79–98, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2009:CBB

[Kir09] Steve Kirkland. A cycle-based bound for subdominant eigenvalues of stochastic matrices. Linear Multilinear Algebra, 57(3):247–266, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2011:SPE

[Kir11] Steve Kirkland. Sign patterns for eigenmatrices of nonnegative matrices. Linear Multilinear Algebra, 59(9):999–1018, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2014:LBM

[Kir14] Steve Kirkland. Load balancing for Markov chains with a specified directed graph. Linear Multilinear Algebra, 62(11): 1491–1508, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kittappa:1981:PCP

[Kit81] R. Kittappa. Proof of a conjecture of 1881 on permanents. Linear and Multilinear Algebra, 10(1):75–82, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kittaneh:1991:NOP

[Kit91a] Fuad Kittaneh. On the normality of operator products. Linear and Multilinear Algebra, 30(1–2):1–4, 1991. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kittaneh:1991:SAC

[Kit91b] Fuad Kittaneh. On the self-adjointness of certain compact operators. Linear and Multilinear Algebra, 28(4):203–206, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kong:2013:NRT

[KJ13] Xu Kong and Yao-Lin Jiang. A note on the ranks of 2 × 2 × 2 and 2 × 2 × 2 × 2 tensors. Linear Multilinear Algebra, 61(10): 1348–1362, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaleibary:2023:OGC

[KJF23] Venus Kaleibary, Mohammad Reza Jabbarzadeh, and Shigeru Furuichi. Operator geodesically convex functions and their applications. Linear Multilinear Algebra, 71(8):1280–1294, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaplansky:1975:IDL

[KK75] Irving Kaplansky and Robert E. Kibler. Infinite-dimensional Lie algebras. II. Linear and Multilinear Algebra, 3(1–2):61–65, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Knopfmacher:1990:CPG

[KK90] Arnold Knopfmacher and John Knopfmacher. Counting polynomials with a given number of zeros in a finite field. Linear and Multilinear Algebra, 26(4):287–292, 1990. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kittaneh:2013:RHO

[KK13] Fuad Kittaneh and Mario Krnić. Refined Heinz operator inequalities. Linear Multilinear Algebra, 61(8):1148–1157, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kurbatov:2016:CFM

[KK16] V. G. Kurbatov and I. V. Kurbatova. Computation of a function of a matrix with close eigenvalues by means of the Newton interpolating polynomial. Linear Multilinear Algebra, 64(2): 111–122, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kumar:2018:NDB

[KK18] V. B. Kiran Kumar and G. Krishna Kumar. A note on discrete Borg-type theorems. Linear Multilinear Algebra, 66(7):1418– 1433, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2019:HCG

[KK19] Younng-Jin Kim and Woong Kook. Harmonic cycles for graphs. Linear Multilinear Algebra, 67(5):965–975, 2019. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Katre:2022:MNC

[KK22a] S. A. Katre and Deepa Krishnamurthi. Matrices over noncommutative rings as sums of powers. Linear Multilinear Algebra, 70(5):824–829, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaygorodov:2022:ACN

[KK22b] Ivan Kaygorodov and Mykola Khrypchenko. The algebraic classification of nilpotent CD-algebras. Linear Multilinear Algebra, 70(19):3840–3886, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kumar:2022:CPR

[KK22c] G. Krishna Kumar and S. H. Kulkarni. Condition pseudospectral radius of bounded linear operators. Linear Multilinear Algebra, 70(1):27–41, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ko:2018:CBM

[KKL18] Eungil Ko, Hyun-Kyoung Kwon, and Ji Eun Lee. A characterization of binormal matrices. Linear Multilinear Algebra, 66(6):1215–1228, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karimjanov:2020:RTO

[KKL20] Iqboljon Karimjanov, Ivan Kaygorodov, and Manuel Ladra. Rota-type operators on null-filiform associative algebras. Linear Multilinear Algebra, 68(1):205–219, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karimjanov:2021:CEF

[KKL21] Iqboljon Karimjanov, Ivan Kaygorodov, and Manuel Ladra. Central extensions of filiform associative algebras. Linear Multilinear Algebra, 69(6):1083–1101, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kang:2022:RCS

[KKL22] Dong-O Kang, Eungil Ko, and Ji Eun Lee. Remarks on complex symmetric Toeplitz operators. Linear Multilinear Algebra, 70(18):3466–3476, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kayibi:2013:USH

[KKP13] K. K. Kayibi, Muhammad Ali Khan, and S. Pirzada. Uniform sampling of k-hypertournaments. Linear Multilinear Algebra, 61(1):123–138, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kovacec:2012:NEL

[KKR12] Alexander Kovacec, Salma Kuhlmann, and Cordian Riener. A note on extrema of linear combinations of elementary symmetric functions. Linear Multilinear Algebra, 60(2):219–224, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kapil:2021:QBF

[KKS21] Yogesh Kapil, Rupinderjit Kaur, and Mandeep Singh. On a question of Bhatia, Friedland and Jain. Linear Multilinear Algebra, 69(16):3031–3042, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Komatsu:2022:WAZ

[KKS22] Takashi Komatsu, Norio Konno, and Iwao Sato. A weighted alternating zeta function of a digraph. Linear Multilinear Algebra, 70(20):5214–5236, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2022:SAT

[KKY22] Hyung Won Kim, Jaewoong Kim, and Jasang Yoon. Spherical Aluthge transform, spherical p and log-hyponormality of commuting pairs of operators. Linear Multilinear Algebra, 70(11): 2047–2064, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kwak:1992:CPS

[KL92] Jin Ho Kwak and Jaeun Lee. Characteristic polynomials of some graph bundles II. Linear and Multilinear Algebra, 32 (1):61–73, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Klaus:1995:IVN

[KL95] Anne-Louise Klaus and Chi-Kwong Li. Isometries for the vector (p, q) norm and the induced (p, q) norm. Linear and Multilinear Algebra, 38(4):315–332, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2011:P

[KL11] Steve Kirkland and Chi-Kwong Li. Preface. Linear Multilinear Algebra, 59(10):1075, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Khurana:2012:GCM

[KL12] Dinesh Khurana and T. Y. Lam. Generalized commutators in matrix rings. Linear Multilinear Algebra, 60(7):797–827, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koo:2013:SDS

[KL13] Yoo Young Koo and Jae Kun Lim. Sum and direct sum of frame sequences. Linear Multilinear Algebra, 61(7):856–870, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ko:2015:NTT

[KL15] Eungil Ko and Ji Eun Lee. Normal truncated Toeplitz operators on finite-dimensional spaces. Linear Multilinear Algebra, 63(10):1947–1971, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2017:PMI

[KL17] Steve Kirkland and Chi-Kwong Li. Preface to the Marcus issue. Linear Multilinear Algebra, 65(10):1925–1926, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ko:2019:RCS

[KL19] Eungil Ko and Ji Eun Lee. Remark on complex symmetric operator matrices. Linear Multilinear Algebra, 67(6):1198– 1216, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2021:PSI

[KL21a] Steve Kirkland and Chi-Kwong Li. Preface of the special issue on numerical ranges and numerical radii. Linear Multilinear Algebra, 69(5):771, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kizil:2021:NEC

[KL21b] Eyüp Kizil and Jimmie Lawson. A new eigenspace characterization of split-quaternion involutions. Linear Multilinear Algebra, 69(11):1981–1988, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ko:2022:SOE

[KL22] Eungil Ko and YoonKyeong Lee. On solutions of operator equations satisfying the similarity of real and positive parts. Linear Multilinear Algebra, 70(17):3271–3286, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kum:2023:HPR

[KLJ23] Sangho Kum, Yongdo Lim, and Juyoung Jeong. Hadamard product and related inequalities in the Jordan spin algebra. Linear Multilinear Algebra, 71(12):2049–2065, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ko:2022:ITO

[KLL22] Eungil Ko, Ji Eun Lee, and Jongrak Lee. (m, C)-isometric Toeplitz operators with rational symbols. Linear Multilinear Algebra, 70(19):4360–4376, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Khudoyberdiyev:2014:SLA

[KLO14] A. Kh. Khudoyberdiyev, M. Ladra, and B. A. Omirov. On solvable Leibniz algebras whose nilradical is a direct sum of null-filiform algebras. Linear Multilinear Algebra, 62(9):1220– 1239, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kudaybergenov:2019:LMT

[KLO19] Karimbergen Kudaybergenov, Manuel Ladra, and Bakhrom Omirov. On Levi–Malcev theorem for Leibniz algebras. Linear Multilinear Algebra, 67(7):1471–1482, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kokabifar:2017:DMP

[KLPK17] E. Kokabifar, G. B. Loghmani, P. J. Psarrakos, and S. M. Karbassi. On the distance from a matrix polynomial to matrix polynomials with k prescribed distinct eigenvalues. Linear Multilinear Algebra, 65(4):658–676, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koo:2008:FOF

[KLS08] Yoo Young Koo, Jae Kun Lim, and In-Sook Shin. Finite orthogonal frames generated by normal operators. Linear Multilinear Algebra, 56(3):345–356, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krupnik:1996:FSM

[KLZ96] Ilya Krupnik, Peter Lancaster, and Peter Zizler. Factorization of selfadjoint matrix polynomials with reas spectrum. Linear and Multilinear Algebra, 40(4):327–336, 1996. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ke:2010:QCH

[KLZ10] W. F. Ke, K. F. Lai, and R. B. Zhang. Quantum codes from Hadamard matrices. Linear Multilinear Algebra, 58(7):847– 854, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kankaanpaa:1984:CDL

[KM84] Hilkka Kankaanpää and Jorma Kaarlo Merikoski. On a conjecture of D. London. Linear and Multilinear Algebra, 15(2):

Kankaanpaa:1985:TIS

[KM85] Hilkka Kankaanpää and Jorma Kaarlo Merikoski. Two inequalities for the sum of elements of a matrix. Linear and Multilinear Algebra, 18(1):9–22, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krupnik:1998:HPM

[KM98] Mark Krupnik and J. J. McDonald. Hereditary properties of matrices with cut-points. Linear and Multilinear Algebra, 45 (1):1–17, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kittaneh:2011:RYH

[KM11] Fuad Kittaneh and Yousef Manasrah. Reverse Young and Heinz inequalities for matrices. Linear Multilinear Algebra, 59(9):1031–1037, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kantun-Montiel:2014:OGI

[KM14] Gabriel Kantún-Montiel. Outer generalized inverses with prescribed ideals. Linear Multilinear Algebra, 62(9):1187–1196, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ke:2017:ADM

[KM17] Yi-Fen Ke and Chang-Feng Ma. Alternating direction methods for solving a class of Sylvester-like matrix equations. Linear Multilinear Algebra, 65(11):2268–2292, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaddoura:2019:CMG

[KM19] Issam Kaddoura and Bassam Mourad. On a class of matrices generated by certain generalized permutation matrices and applications. Linear Multilinear Algebra, 67(10):2117–2134, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kian:2022:MIR

[KM22] Mohsen Kian and Mohammad Sal Moslehian. Matrix inequalities related to power means of probability measures. Linear

Kurdachenko:2012:GAV

[KMEO12] L. A. Kurdachenko, J. M. Muñoz-Escolano, and J. Otal. Groups acting on vector spaces with a large family of invariant subspaces. Linear Multilinear Algebra, 60(4):487–498, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Korkee:2019:LTA

[KMH19] Ismo Korkee, Mika Mattila, and Pentti Haukkanen. A latticetheoretic approach to the Bourque–Ligh conjecture. Linear Multilinear Algebra, 67(12):2471–2487, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krupnik:1995:FSQ

[KML95] Ilya Krupnik, Alexander Markus, and Peter Lancaster. Factorization of selfadjoint quadratic matrix polynomials with real spectrum. Linear and Multilinear Algebra, 39(3):263–272, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Konno:2022:TFR

[KMMS22] Norio Konno, Hideo Mitsuhashi, Hideaki Morita, and Iwao Sato. The trace formula with respect to the Grover matrix of a graph. Linear Multilinear Algebra, 70(10):1963–1979, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2001:SLB

[KMN01] Stephen J. Kirkland, Jason J. Molitierno, and Michael Neumann. The sharpness of a lower bound on the algebraic connectivity for maximal graphs. Linear and Multilinear Algebra, 48(3):237–246, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kian:2022:ACD

[KMN22] M. Kian, M. S. Moslehian, and R. Nakamoto. Asymmetric Choi–Davis inequalities. Linear Multilinear Algebra, 70(17): 3287–3300, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2007:IZN

[KMOV07] In-Jae Kim, J. J. McDonald, D. D. Olesky, and P. Van Den Driessche. Inertias of zero-nonzero patterns. Linear Multilinear Algebra, 55(3):229–238, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2019:PQS

[KMP^{+}19] Steve Kirkland, Darian McLaren, Rajesh Pereira, Sarah Plosker, and Xiaohong Zhang. Perfect quantum state transfer in weighted paths with potentials (loops) using orthogonal polynomials. Linear Multilinear Algebra, 67(5):1043–1061, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaffes:1989:MCM

[KMRS89] Demetrios G. Kaffes, Thomas Mathew, M. Bhaskara Rao, and K. Subramanyam. On the matrix convexity of the Moore– Penrose inverse and some applications. Linear and Multilinear Algebra, 24(4):265–271, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koukouvinos:2002:IFH

[KMS02] C. Koukouvinos, M. Mitrouli, and Jennifer Seberry. An infinite family of Hadamard matrices with fourth last pivot n/2. Linear Multilinear Algebra, 50(2):167–173, 2002. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kravvaritis:2007:PSW

[KMS07] Christos Kravvaritis, Marilena Mitrouli, and Jennifer Seberry. On the pivot structure for the weighing matrix W (12, 11). Linear Multilinear Algebra, 55(5):471–490, 2007. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karantha:2016:MPO

[KMS16] Manjunatha Prasad Karantha, Mohana, and P. Divya Shenoy. Minus partial order on regular matrices. Linear Multilinear Algebra, 64(5):929–941, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Khosravi:2018:SOI

[KMS18] Maryam Khosravi, Mohammad Sal Moslehian, and Alemeh Sheikhhosseini. Some operator inequalities involving operator means and positive linear maps. Linear Multilinear Algebra, 66(6):1186–1198, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kian:2021:VAH

[KMS21] Mohsen Kian, M. S. Moslehian, and Yuki Seo. Variants of Ando–Hiai inequality for operator power means. Linear Multilinear Algebra, 69(9):1694–1704, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:1996:SPW

[KMT96] S. J. Kirkland, J. J. McDonald, and M. J. Tsatsomeros. Signpatterns which require a positive eigenvalue. Linear and Multilinear Algebra, 41(3):199–210, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krein:1981:MSH

[KN81] M. G. Krein and M. A. Naimark. The method of symmetric and Hermitian forms in the theory of the separation of the roots of algebraic equations. Linear and Multilinear Algebra, 10(4):265–308, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:1997:ACW

[KN97] Steve Kirkland and Michael Neumann. Algebraic connectivity of weighted trees under perturbation. Linear and Multilinear Algebra, 42(3):187–203, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2000:EFP

[KN00] Stephen J. Kirkland and Michael Neumann. Extremal first passage times for trees. Linear and Multilinear Algebra, 48 (1):21–33, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2004:ACF

[KN04] S. J. Kirkland and M. Neumann. On algebraic connectivity as a function of an edge weight. Linear Multilinear Algebra, 52 (1):17–33, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kakavandi:2022:ADS

[KN22] Bijan Ahmadi Kakavandi and Elham Nobari. Approximate diagonalization of some Toeplitz operators and matrices. Linear Multilinear Algebra, 70(21):7093–7114, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:1996:CVW

[KNS96] Steve Kirkland, Michael Neumann, and Bryan L. Shader. Characteristic vertices of weighted trees via Perron values. Linear and Multilinear Algebra, 40(4):311–325, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Knuth:1985:SOB

[Knu85] Donald E. Knuth. Semi-optimal bases for linear dependencies. Linear and Multilinear Algebra, 17(1):1–4, 1985. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koltracht:1993:QBH

[KNX93] I. Koltracht, M. Neumann, and D. Xiao. On a question of Boyle and Handelman concerning eigenvalues of nonnegative matrices. Linear and Multilinear Algebra, 36(2):125–140, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrections [KNX95].

Koltracht:1995:CQB

[KNX95] I. Koltracht, M. Neumann, and D. Xiao. Corrections: On a question of Boyle and Handelman concerning eigenvalues of nonnegative matrices. Linear and Multilinear Algebra, 38(4): 379–380, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [KNX93].

Kamiya:2010:RTS

[KO10] Noriaki Kamiya and Susumu Okubo. Representations of (α, β, γ) triple system. Linear Multilinear Algebra, 58(5):617– 643, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kobal:2006:PDU

[Kob06] Damjan Kobal. Preserving diagonalisability on upper triangular matrices. Linear Multilinear Algebra, 54(3):189–194, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kohls:2011:UFP

[Koh11] Martin Kohls. A user friendly proof of Nagata’s characterization of linearly reductive groups in positive characteristics. Linear Multilinear Algebra, 59(3):271–278, 2011. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koiran:2021:OTD

[Koi21] Pascal Koiran. Orthogonal tensor decomposition and orbit closures from a linear algebraic perspective. Linear Multilinear Algebra, 69(13):2353–2388, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kolegov:2023:LMI

[Kol23] N. A. Kolegov. On the lengths of matrix incidence algebras with radicals of square zero. Linear Multilinear Algebra, 71 (10):1736–1753, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Komatsu:2010:FTL

[Kom10] Takao Komatsu. Four-term leaping recurrence relations. Linear Multilinear Algebra, 58(7):875–886, 2010. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koolen:1995:CMP

[Koo95] J. H. Koolen. On a conjecture of Martin on the parameters of completely regular codes and the classification of the completely regular codes in the Biggs–Smith graph. Linear and Multilinear Algebra, 39(1–2):3–17, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kopotun:1994:SPC

[Kop94] Kirill A. Kopotun. On some permanental conjectures. Linear and Multilinear Algebra, 36(3):205–216, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2023:MGS

[KOSS23] Minjae Kim, Suil O., Wooyong Sim, and Dongwoo Shin. Matchings in graphs from the spectral radius. Linear Multilinear Algebra, 71(11):1794–1803, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kovacs:1977:TPL

[Kov77] Amos Kovacs. Trace preserving linear transformations on matrix algebras. Linear and Multilinear Algebra, 4(4):243–250, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kovacec:1986:GTK

[Kov86] Alexander Kovacec. A generalization of a theorem of könig. Linear and Multilinear Algebra, 19(1):51–56, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2009:CSI

[KOvDD09] In-Jae Kim, D. D. Olesky, and P. van Den Driessche. Critical sets of inertias for matrix patterns. Linear Multilinear Algebra, 57(3):293–306, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kozyakin:2017:MTS

[Koz17] V. Kozyakin. Minimax theorem for the spectral radius of the product of non-negative matrices. Linear Multilinear Algebra, 65(11):2356–2365, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:1992:LOP

[KP92] Stephen Kirkland and Norman J. Pullman. Linear operators preserving invariants of nonbinary boolean matrices. Linear and Multilinear Algebra, 33(3–4):295–300, 1992. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:2002:VCA

[KP02] Steve Kirkland and Sukanta Pati. On vertex connectivity and absolute algebraic connectivity for graphs. Linear Multilinear Algebra, 50(3):253–284, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karademir:2011:SNS

[KP11] Serdar Karademir and Oleg A. Prokopyev. A short note on solvability of systems of interval linear equations. Linear Multilinear Algebra, 59(6):707–710, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kalita:2012:SCD

[KP12] D. Kalita and S. Pati. On the spectrum of 3-coloured digraphs. Linear Multilinear Algebra, 60(6):743–756, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Khachorncharoenkul:2014:RPH

[KP14a] P. Khachorncharoenkul and S. Pianskool. Rank-1 preservers on Hessenberg matrices. Linear Multilinear Algebra, 62(1): 96–113, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kribs:2014:PQC

[KP14b] David W. Kribs and Sarah Plosker. Private quantum codes: introduction and connection with higher rank numerical ranges. Linear Multilinear Algebra, 62(5):639–647, 2014. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaygorodov:2016:GDC

[KP16] Ivan Kaygorodov and Yury Popov. Generalized derivations of (color) n-ary algebras. Linear Multilinear Algebra, 64(6): 1086–1106, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kannan:2022:CCG

[KP22] M. Rajesh Kannan and Shivaramakrishna Pragada. On the construction of cospectral graphs for the adjacency and the normalized Laplacian matrices. Linear Multilinear Algebra, 70(15):3009–3030, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kribs:2009:RPN

[KPL^{+}09] David W. Kribs, Aron Pasieka, Martin Laforest, Colm Ryan, and Marcus P. da Silva. Research problems on numerical ranges in quantum computing. Linear Multilinear Algebra, 57(5):491–502, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kribs:2013:TPM

[KPP13] David W. Kribs, Rajesh Pereira, and Sarah Plosker. Trumping and power majorization. Linear Multilinear Algebra, 61(11):

Kostic:2021:GJF

[KPPR21] Aleksandra Kostić, Zoran Z. Petrović, Zoran S. Pucanović, and Maja Roslavcev. On a generalized Jordan form of an infinite upper triangular matrix. Linear Multilinear Algebra, 69(8):1534–1542, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaygorodov:2018:DZN

[KPPV18] Ivan Kaygorodov, Yury Popov, Alexandre Pozhidaev, and Yury Volkov. Degenerations of Zinbiel and nilpotent Leibniz algebras. Linear Multilinear Algebra, 66(4):704–716, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [Ano22b].

Kaygorodov:2019:TGJ

[KPS19] Ivan Kaygorodov, Alexander Pozhidaev, and Paulo Saraiva. On a ternary generalization of Jordan algebras. Linear Multilinear Algebra, 67(6):1074–1102, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koliha:2002:ISI

[KR02] J. J. Koliha and V. RakoCević. Invertibility of the sum of idempotents. Linear Multilinear Algebra, 50(4):285–292, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koliha:2003:IDI

[KR03] J. J. Koliha and V. Rakocević. Invertibility of the difference of idempotents. Linear Multilinear Algebra, 51(1):97–110, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koliha:2007:RPM

[KR07] J. J. Koliha and V. RakoCević. Range projections and the Moore–Penrose inverse in rings with involution. Linear Multilinear Algebra, 55(2):103–112, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Korporal:2014:PPG

[KR14] Anja Korporal and Georg Regensburger. On the product of projectors and generalized inverses. Linear Multilinear Algebra, 62(12):1567–1582, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krauter:1987:TMM

[Krä87] Arnold R. Kräuter. On a theorem of M. Marcus and H. Minc and a little known expansion formula for permanents. Linear and Multilinear Algebra, 20(4):367–371, 1987. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kramar:2005:SIM

[Kra05] Marjeta Kramar. The structure of irreducible matrix groups with submultiplicative spectrum. Linear Multilinear Algebra, 53(1):13–25, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kramar:2006:RGS

[Kra06] Marjeta Kramar. Representations of p-groups with submultiplicative spectra. Linear Multilinear Algebra, 54(5):313–320, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kremere:2016:LSO

[Kre16] Insa Kremere. Left-star order structure of Rickart ∗-rings. Linear Multilinear Algebra, 64(3):341–352, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krieg:1987:EDT

[Kri87] Aloys Krieg. The elementary divisor theory over the Hurwitz order of integral quaternions. Linear and Multilinear Algebra, 21(4):325–344, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kubjas:2018:PSR

[KRR18] Kaie Kubjas, Elina Robeva, and Richard Z. Robinson. Positive semidefinite rank and nested spectrahedra. Linear Multilinear Algebra, 66(10):1952–1974, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krupka:2006:TDT

[Kru06] Demeter Krupka. Trace decompositions of tensor spaces. Linear Multilinear Algebra, 54(4):235–263, 2006. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Knopp:1976:PSC

[KS76] Paul J. Knopp and Richard Sinkhorn. Permanents of special classes of doubly stochastic matrices. Linear and Multilinear Algebra, 4(2):129–136, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Knopp:1982:MPD

[KS82] Paul Knopp and Richard Sinkhorn. Minimum permanents of doubly stochastic matrices with at least one zero entry. Linear and Multilinear Algebra, 11(4):351–355, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krauter:1983:CM

[KS83] Arnold R. Kräuter and Norbert Seifter. On convertible (0, 1)- matrices. Linear and Multilinear Algebra, 13(4):311–322, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Krauter:1984:SPP

[KS84] Arnold R. Kräuter and Norbert Seifter. Some properties of the permanent of (1, −1)-matrices. Linear and Multilinear Algebra, 15(3–4):207–223, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kirkland:1993:MTM

[KS93] Steve Kirkland and Bryan L. Shader. On multipartite tournament matrices with constant team size. Linear and Multilinear Algebra, 35(1):49–63, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2002:SSC

[KS02] Si-Ju Kim and Bryan Shader. Sign-solvable cone-systems. Linear Multilinear Algebra, 50(1):23–32, 2002. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kittaneh:2007:BZP

[KS07] Fuad Kittaneh and Khalid Shebrawi. Bounds for the zeros of polynomials from matrix inequalities — II. Linear Multilinear Algebra, 55(2):147–158, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2009:NSA

[KS09] In-Jae Kim and Bryan L. Shader. Non-singular acyclic matrices. Linear Multilinear Algebra, 57(4):399–407, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kang:2012:TRP

[KS12] Kyung-Tae Kang and Seok-Zun Song. Term rank preservers of Boolean matrices. Linear Multilinear Algebra, 60(2):241–247, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kannan:2014:MGM

[KS14a] M. Rajesh Kannan and K. C. Sivakumar. P_{+}-matrices: a generalization of P -matrices. Linear Multilinear Algebra, 62 (1):1–12, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kaur:2014:AOV

[KS14b] Rupinderjit Kaur and Mandeep Singh. Another operator version of generalized Bernoulli’s inequality. Linear Multilinear Algebra, 62(8):1127–1136, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kang:2015:RMT

[KS15] Kyung-Tae Kang and Seok-Zun Song. Regular matrices and their generalized inverses over the max algebra. Linear Multilinear Algebra, 63(8):1649–1663, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kalvoda:2016:FED

[KS16] Tomás Kalvoda and Pavel Stovícek. A family of explicitly diagonalizable weighted Hankel matrices generalizing the Hilbert matrix. Linear Multilinear Algebra, 64(5):870–884, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Koledin:2017:CSG

[KS17a] Tamara Koledin and Zoran Stanić. Connected signed graphs of fixed order, size, and number of negative edges with maximal index. Linear Multilinear Algebra, 65(11):2187–2198, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kumar:2017:LSL

[KS17b] N. Kishore Kumar and J. Schneider. Literature survey on low rank approximation of matrices. Linear Multilinear Algebra, 65(11):2212–2244, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kalita:2019:MAC

[KS19a] Debajit Kalita and Kuldeep Sarma. Minimizing algebraic connectivity over graphs made with some given blocks. Linear Multilinear Algebra, 67(5):911–925, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kittaneh:2019:IAD

[KS19b] Fuad Kittaneh and Mona Sakkijha. Inequalities for accretivedissipative matrices. Linear Multilinear Algebra, 67(5):1037– 1042, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kalita:2022:INB

[KS22a] Debajit Kalita and Kuldeep Sarma. Inverses of non-bipartite unicyclic graphs with a unique perfect matching. Linear Multilinear Algebra, 70(14):2765–2781, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kalita:2022:IUC

[KS22b] Debajit Kalita and Kuldeep Sarma. On the inverse of unicyclic 3-coloured digraphs. Linear Multilinear Algebra, 70(21):6223– 6239, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kapil:2022:DSS

[KS22c] Yogesh Kapil and Mandeep Singh. Determinants of some special matrices. Linear Multilinear Algebra, 70(16):3119–3141, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kalita:2023:UNB

[KS23] Debajit Kalita and Kuldeep Sarma. On unicyclic non-bipartite graphs with tricyclic inverses. Linear Multilinear Algebra, 71 (8):1378–1396, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kannan:2016:WIN

[KSMB16] M. Rajesh Kannan, Naomi Shaked-Monderer, and Abraham Berman. On weakly irreducible nonnegative tensors and interval hull of some classes of tensors. Linear Multilinear Algebra, 64(4):667–679, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kalinina:2022:SDI

[KSU22] Elizaveta A. Kalinina, Yurii A. Smol’kin, and Alexei Yu. Uteshev. Stability and distance to instability for polynomial matrix families. Complex perturbations. Linear Multilinear Algebra, 70(7):1291–1314, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Komiya:1983:SLI

[KT83] Hidetoshi Komiya and Wataru Takahashi. Systems of linear inequalities on normed linear spaces. Linear and Multilinear Algebra, 13(3):267–279, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Khosrovshahi:1995:NLG

[KT95] G. B. Khosrovshahi and R. Torabi. A note on labelings of graphs. Linear and Multilinear Algebra, 39(1–2):165–169, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Karaev:2006:SAD

[KT06] M. T. Karaev and H. Tuna. On some applications of Duhamel product. Linear Multilinear Algebra, 54(4):301–311, 2006. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kosal:2019:USF

[KT19] Hidayet Hüda Kösal and Murat Tosun. Universal similarity factorization equalities for commutative quaternions and their matrices. Linear Multilinear Algebra, 67(5):926–938, 2019.

Khan:2021:ABT

[KT21] Muhammad Ahsan Khan and Dan Timotin. Algebras of block Toeplitz matrices with commuting entries. Linear Multilinear Algebra, 69(14):2702–2716, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kuai:2018:EFM

[Kua18] Ligang Kuai. An extension of the Fiedler–Markham determinant inequality. Linear Multilinear Algebra, 66(3):547–553, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kulkarni:2022:SAP

[Kul22] Avinash Kulkarni. Solving p-adic polynomial systems via iterative eigenvector algorithms. Linear Multilinear Algebra, 70(4):650–671, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kunicki:1999:CPN

[Kun99] Catherine Kunicki. Commutativity-preserving and normalpreserving linear transformations on. Linear and Multilinear Algebra, 45(4):341–347, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kurtz:2004:SPM

[Kur04] Cornelius Kurtz. On sums and products of matrix similarity classes I. Linear Multilinear Algebra, 52(1):35–47, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kutsenko:2019:MMI

[Kut19] Anton A. Kutsenko. Mixed multidimensional integral operators with piecewise constant kernels and their representations. Linear Multilinear Algebra, 67(1):186–195, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kuzma:2005:AMD

[Kuz05] B. Kuzma. Additive mappings that do not increase minimal rank of alternate matrices. Linear Multilinear Algebra, 53(4):

Kolesnikov:2013:SID

[KV13] P. S. Kolesnikov and V. Yu. Voronin. On special identities for dialgebras. Linear Multilinear Algebra, 61(3):377–391, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kulkarni:2022:SLC

[KV22] Avinash Kulkarni and Tristan Vaccon. Super-linear convergence in the p -adic QR-algorithm. Linear Multilinear Algebra, 70(22):7778–7806, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Klobouk:2023:MSJ

[KV23] Abel H. Klobouk and Alejandro Varela. Moment of a subspace and joint numerical range. Linear Multilinear Algebra, 71(9): 1470–1503, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kubo:1989:CID

[KW89] Fumio Kubo and Shuhei Wada. Cyclic inequalities and discrete Fourier analysis. Linear and Multilinear Algebra, 25(2):121– 136, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

King:1992:COG

[KW92] R. C. King and T. A. Welsh. Construction of orthogonal group modules using tableaux. Linear and Multilinear Algebra, 33(3– 4):251–283, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

King:1993:CMU

[KW93] R. C. King and T. A. Welsh. Construction of GL(n) modules using composite tableaux. Linear and Multilinear Algebra, 34 (2):99–122, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Khan:2011:DIA

[KW11] Israr Ali Khan and Qing-Wen Wang. The Drazin inverse in an arbitrary semiring. Linear Multilinear Algebra, 59(9):1019–

Khairnar:2017:OPG

[KW17] Anil Khairnar and B. N. Waphare. Order properties of generalized projections. Linear Multilinear Algebra, 65(7):1446–1461, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Katre:2021:MNR

[KW21] S. A. Katre and Kshipra Wadikar. Matrices over noncommutative rings as sums of k th powers. Linear Multilinear Algebra, 69(11):2050–2058, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kim:2017:PMW

[KY17] Jaewoong Kim and Jasang Yoon. Properties of mono-weakly hyponormal 2-variable weighted shifts. Linear Multilinear Algebra, 65(6):1260–1275, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kyrchei:2008:AAM

[Kyr08] I. I. Kyrchei. Analogs of the adjoint matrix for generalized inverses and corresponding Cramer rules. Linear Multilinear Algebra, 56(4):453–469, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kyrchei:2011:DRM

[Kyr11] I. I. Kyrchei. Determinantal representations of the Moore– Penrose inverse over the quaternion skew field and corresponding Cramer’s rules. Linear Multilinear Algebra, 59(4):413–431, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kyrchei:2021:DRS

[Kyr21] Ivan Kyrchei. Determinantal representations of solutions to systems of two-sided quaternion matrix equations. Linear Multilinear Algebra, 69(4):648–672, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kong:2022:BLD

[KZN22] Liang Kong, Jianhua Zhang, and Tong Ning. Bilocal Lie derivations on nest algebras. Linear Multilinear Algebra, 70(9):

Dokovic:1992:HMO

[– DL92] D. Ž. – Doković and L. Legorreta. Hermitian matrices of order 2 over the ring of finite Fourier series with complex coefficients. Linear and Multilinear Algebra, 32(3–4):283–309, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See correction [– Dok93].

LaGuardia:2019:AGC

[La 19] Giuliano G. La Guardia. Asymptotically good convolutional codes. Linear Multilinear Algebra, 67(7):1483–1494, 2019. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laffey:1978:STM

[Laf78] Thomas J. Laffey. Simultaneous triangularization of matrices— low rank cases and the nonderogatory case. Linear and Multilinear Algebra, 6(4):269–305, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laffey:1980:BTM

[Laf80] Thomas J. Laffey. A basis theorem for matrix algebras. Linear and Multilinear Algebra, 8(3):183–187, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laffey:1983:PIM

[Laf83] Thomas J. Laffey. Products of idempotent matrices. Linear and Multilinear Algebra, 14(4):309–314, 1983. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laffey:1984:CFM

[Laf84] Thomas J. Laffey. A characterisation of full matrix algebras. Linear and Multilinear Algebra, 15(2):105–108, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laszlo:1991:UBM

[Lás91] Lajos László. Upper bounds for matrix diagonals. Linear and Multilinear Algebra, 30(4):283–301, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lautemann:1981:LTM

[Lau81] Clemens Lautemann. Linear transformations on matrices: rank preservers and determinant preservers (note). Linear and Multilinear Algebra, 10(4):343–345, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lawrence:2019:BGA

[Law19] Ben Lawrence. Burnside graphs, algebras generated by sets of matrices, and the Kippenhahn conjecture. Linear Multilinear Algebra, 67(1):51–69, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:1999:HEB

[LB99] Bolian Liu and Zhou Bo. On the hall exponents of boolean matrices. Linear and Multilinear Algebra, 46(3):165–175, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2011:JGD

[LB11] Yanbo Li and Dominik Benkovic. Jordan generalized derivations on triangular algebras. Linear Multilinear Algebra, 59(8): 841–849, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lifshitz:2019:FFU

[LB19] Alexander Lifshitz and Yuval Bistritz. A fraction-free unitcircle zero location test for a polynomial with any singularity profile. Linear Multilinear Algebra, 67(7):1420–1459, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lombarkia:2021:OEI

[LB21] F. Lombarkia and A. Boussaid. Operator equations and inner inverses of elementary operators. Linear Multilinear Algebra, 69(11):1989–1996, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Luengo:2013:CQF

[LBV13] E. Almaraz Luengo, J. M. Ancochea Bermúdez, and L. García Vergnolle. Completeness of quasi-filiform Lie algebras. Linear Multilinear Algebra, 61(5):582–595, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lomont:1983:MPD

[LC83] J. S. Lomont and M. S. Cheena. A multilinearity property of determinant functions. Linear and Multilinear Algebra, 14(3): 199–223, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2014:MLS

[LC14] Wei Li and An Chang. The minimal Laplacian spectral radius of trees with given matching number. Linear Multilinear Algebra, 62(2):218–228, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2018:CCD

[LC18a] Tingting Li and Jianlong Chen. Characterizations of core and dual core inverses in rings with involution. Linear Multilinear Algebra, 66(4):717–730, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2018:IMS

[LC18b] Matthew M. Lin and Chun-Yueh Chiang. An iterative method for solving the stable subspace of a matrix pencil and its application. Linear Multilinear Algebra, 66(7):1279–1298, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2020:CIC

[LC20] Tingting Li and Jianlong Chen. The core invertibility of a companion matrix and a Hankel matrix. Linear Multilinear Algebra, 68(3):550–562, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:ISL

[LC22] Wende Liu and Mengmeng Cai. On ID*-superderivations of Lie superalgebras. Linear Multilinear Algebra, 70(8):1513– 1525, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2017:ACU

[LCC17] Wei Li, Joshua Cooper, and An Chang. Analytic connectivity of k-uniform hypergraphs. Linear Multilinear Algebra, 65(6): 1247–1259, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2019:RBL

[LCC19] Juan Li, Liangyun Chen, and Yongsheng Cheng. Representations of Bihom–Lie superalgebras. Linear Multilinear Algebra, 67(2):299–326, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2023:SRD

[LCG^{+}23] Muhuo Liu, Chaohui Chen, Shu-Guang Guo, Jiarong Peng, and Tianyuan Chen. The A_{α}-spectral radius of dense graphs. Linear Multilinear Algebra, 71(6):1044–1053, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:MFG

[LCH22] Ai-Jun Li, Zi-Xin Cao, and Qingzhong Huang. A matrix form of Grassmannian Ball–Barthe inequality. Linear Multilinear Algebra, 70(21):6511–6522, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:2013:TGT

[LCL13] Jie Lei, You’an Cao, and Zhiyong Li. Triple and generalized triple derivations of the parabolic subalgebras of the general linear Lie algebra over a commutative ring. Linear Multilinear Algebra, 61(3):337–353, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2014:GSD

[LCL14] Cheng-Kai Liu, Hung-Yuan Chen, and Pao-Kuei Liau. Generalized skew derivations with nilpotent values on left ideals of rings and Banach algebras. Linear Multilinear Algebra, 62(4): 453–465, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:TFR

[LCL22] Yue Liu, Ailian Chen, and Fenggen Lin. Threshold function of ray nonsingularity for uniformly random ray pattern matrices. Linear Multilinear Algebra, 70(20):5708–5715, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:BES

[LCLL22] Shigui Li, Zhen Chen, Qilong Liu, and Linzhang Lu. Bounds of M-eigenvalues and strong ellipticity conditions for elasticity tensors. Linear Multilinear Algebra, 70(19):4544–4557, 2022.

Luo:2022:PQS

[LCWW22] Gaojun Luo, Xiwang Cao, Dandan Wang, and Xia Wu. Perfect quantum state transfer on Cayley graphs over semi-dihedral groups. Linear Multilinear Algebra, 70(21):6358–6374, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Luo:2022:CGD

[LCXC22] Gaojun Luo, Xiwang Cao, Guangkui Xu, and Yingjie Cheng. Cayley graphs of dihedral groups having perfect edge state transfer. Linear Multilinear Algebra, 70(20):5957–5972, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2019:CDC

[LCZW19] Tingting Li, Jianlong Chen, Mengmeng Zhou, and Dingguo Wang. The core and dual core inverse of a morphism with kernel. Linear Multilinear Algebra, 67(10):1937–1947, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2017:IRC

[LD17] Tengfei Li and Chunyuan Deng. On the invertibility and range closedness of the linear combinations of a pair of projections. Linear Multilinear Algebra, 65(3):613–622, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2019:SSO

[LD19] Wenting Liang and Chunyuan Deng. The solutions to some operator equations with corresponding operators not necessarily having closed ranges. Linear Multilinear Algebra, 67(8): 1606–1624, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:DRA

[LD22] Yun-Zhang Li and Jian Dong. Duality relations associated with weak g-R-duals. Linear Multilinear Algebra, 70(20):5482– 5501, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2017:SLS

[LDL17] Muhuo Liu, Kinkar Ch. Das, and Hong-Jian Lai. The (signless) Laplacian spectral radii of c-cyclic graphs with n vertices, girth g and k pendant vertices. Linear Multilinear Algebra, 65(5):869–881, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:1991:IM

[Lee91] Yew Sing Lee. Inequalities for M and N_{0}-matrices. Linear and Multilinear Algebra, 29(2):149–154, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:1994:SDW

[Lee94] Chanyoung Lee. Stability of dominant weights of the tensor product module for orthosymplectic Lie superalgebras. Linear and Multilinear Algebra, 37(4):283–295, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:1998:RTL

[Lee98] Min Ho Lee. Radon transforms, Laplacians, and flows for directed graphs. Linear and Multilinear Algebra, 45(2–3):219– 233, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:2015:SFM

[Lee15] Tsiu-Kwen Lee. Sesquilinear functionals on matrix algebras with involution. Linear Multilinear Algebra, 63(7):1283–1293, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:2016:FIJ

[Lee16] Tsiu-Kwen Lee. Functional identities and Jordan σderivations. Linear Multilinear Algebra, 64(2):221–234, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:2018:IMD

[Lee18] Jungin Lee. Integral matrices as diagonal quadratic forms. Linear Multilinear Algebra, 66(4):742–747, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Legisa:2006:AMP

[Leg06] Peter Legisa. Automorphisms of Mn, partially ordered by the star order. Linear Multilinear Algebra, 54(3):157–188, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lehmer:1976:CLP

[Leh76] D. H. Lehmer. The characters of linear permutations. Linear and Multilinear Algebra, 4(1):1–16, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lehmer:1991:MP

[Leh91] D. H. Lehmer. Matrix paraphrases. Linear and Multilinear Algebra, 28(4):251–264, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1996:SNC

[LEH96] Zhongshan Li, Carolyn A. Eschenbach, and Frank J. Hall. The structure of nonnegative cyclic matrices. Linear and Multilinear Algebra, 41(1):23–33, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leipnik:1981:BR

[Lei81] Roy B. Leipnik. Book review. Linear and Multilinear Algebra, 10(1):83, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leipnik:1983:ENF

[Lei83] Roy B. Leipnik. Extension of Newman’s formula for the Sylvester equation to the generic case. Linear and Multilinear Algebra, 12(4):305–309, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:1991:CNR

[Lei91] Tian-Gang Lei. On the c-congruence numerical radii. Linear and Multilinear Algebra, 30(3):223–227, 1991. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:1992:GNR

[Lei92] Tian-Gang Lei. On the generalized c-numerical range. Linear and Multilinear Algebra, 31(1–4):251–265, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:1993:GNR

[Lei93a] Tian-Gang Lei. On generalized numerical ranges with the induced matrix. Linear and Multilinear Algebra, 34(3–4):385– 394, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:1993:NRI

[Lei93b] Tian-Gang Lei. On the numerical range of an induced power. Linear and Multilinear Algebra, 35(2):173–178, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leite:1993:GHM

[Lei93c] F. Silva Leite. The geometry of hypercomplex matrices. Linear and Multilinear Algebra, 34(2):123–132, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:1995:MSE

[Lei95a] Tian-Gang Lei. On matrices with some extermal generalized numerical radii. Linear and Multilinear Algebra, 38(3):267– 279, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:1995:GCN

[Lei95b] Tian-Gang Lei. On the generalized congruence numerical range with the completely symmetric function. Linear and Multilinear Algebra, 40(1):47–59, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:1998:CNR

[Lei98] Tian-Gang Lei. Congruence numerical ranges and their radii. Linear and Multilinear Algebra, 43(4):411–427, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:2018:ISE

[Lei18] Xiaoqiang Lei. _{β}^{α}Φ_{s}-inclusion sets for eigenvalues of a tensor. Linear Multilinear Algebra, 66(5):942–960, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leka:2019:SSV

[Lék19] Zoltán Léka. Some singular value inequalities via convexity. Linear Multilinear Algebra, 67(2):360–369, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lester:1993:OS

[Les93] J. A. Lester. Orthochronous subgroups of O(p, q). Linear and Multilinear Algebra, 36(2):111–113, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Levin:1974:SMS

[Lev74] Simon A. Levin. Stability matrices and the solvability of certain systems of linear inequalities. Linear and Multilinear Algebra, 2(3):253–255, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lewin:1977:NMG

[Lew77] Mordechai Lewin. Nonnegative matrices generating a finite cyclic group. Linear and Multilinear Algebra, 5(2):91–94, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lewis:1990:LWR

[Lew90] D. W. Lewis. The level of a Witt ring. Linear and Multilinear Algebra, 27(3):163–165, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lewis:1996:TFK

[Lew96] D. W. Lewis. Trace forms, Kronecker sums, and the shuffle matrix. Linear and Multilinear Algebra, 40(3):221–227, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lewis:1997:HMQ

[Lew97] D. W. Lewis. Hankel matrices and quadratic forms. Linear and Multilinear Algebra, 42(4):349–363, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2013:LTJ

[LF13] Changjing Li and Xiaochun Fang. Lie triple and Jordan derivable mappings on nest algebras. Linear Multilinear Algebra, 61(5):653–666, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2015:CSW

[LF15] Xin Li and Xiaochun Fang. Constructing k-Schmidt witnesses for infinite-dimensional systems. Linear Multilinear Algebra,

Li:2016:IPL

[LF16] Xin Li and Xiaochun Fang. Indecomposable positive linear maps constructed from permutations. Linear Multilinear Algebra, 64(2):297–308, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2021:SIS

[LF21] Song Lin and Xiaohui Fu. On some inequalities for sector matrices. Linear Multilinear Algebra, 69(3):428–437, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2016:IPO

[LFY16] Xiaoji Liu, Shiqin Fu, and Yaoming Yu. An invariance property of operator products related to the mixed-type reverse order laws. Linear Multilinear Algebra, 64(5):885–896, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2023:CLC

[LG23] Lei Liu and Kaitian Gao. Characterizations of Lie centralizers of triangular algebras. Linear Multilinear Algebra, 71(14): 2375–2391, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2013:SPM

[LGA^{+}13] Zhongshan Li, Yubin Gao, Marina Arav, Fei Gong, Wei Gao, Frank J. Hall, and Hein van der Holst. Sign patterns with minimum rank 2 and upper bounds on minimum ranks. Linear Multilinear Algebra, 61(7):895–908, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2017:NLS

[LGB17] Xuewen Liu, Xiangqian Guo, and Dongping Bian. A note on the left-symmetric algebraic structures of the Witt algebra. Linear Multilinear Algebra, 65(9):1793–1804, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2019:DOD

[LGG19] Xuewen Liu, Li Guo, and Xiangqian Guo. λ-differential operators and λ-differential modules for the Virasoro algebra. Linear Multilinear Algebra, 67(7):1308–1324, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Gau:1999:LTR

[lGyW99] Hwa long Gau and Pei yuan Wu. Lucas’ theorem refined. Linear and Multilinear Algebra, 45(4):359–373, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2020:SRI

[LGZ20] Hongying Lin, Haiyan Guo, and Bo Zhou. On the α-spectral radius of irregular uniform hypergraphs. Linear Multilinear Algebra, 68(2):265–277, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2013:SCS

[LH13] Fenjin Liu and Qiongxiang Huang. On the spectral characterization of Π-shape trees. Linear Multilinear Algebra, 61(3): 355–367, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2021:SIF

[LH21a] Jing Li and Han Hai. On some inequalities for the fan product of matrices. Linear Multilinear Algebra, 69(12):2264–2273, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lu:2021:DE

[LH21b] Lu Lu and Qiongxiang Huang. Distance eigenvalues of B(n, k). Linear Multilinear Algebra, 69(11):2078–2092, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:FPS

[LH22] Yun-Zhang Li and Tufail Hussain. The formation and portraits of subspace F_{a}-frames. Linear Multilinear Algebra, 70 (9):1610–1630, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2023:MPR

[LH23] Lele Liu and Changxiang He. The maximum principal ratio of graphs. Linear Multilinear Algebra, 71(10):1691–1706, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Luo:2015:TSP

[LHbL^{+}15] Jun Luo, Ting-Zhu Huang, Hou biao Li, Zhongshan Li, and Lihua Zhang. Tree sign patterns that allow nilpotence of index 4. Linear Multilinear Algebra, 63(5):1009–1025, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lu:2022:EG

[LHF22] Lu Lu, Qiongxiang Huang, and Lihua Feng. On 2-equitable graphs. Linear Multilinear Algebra, 70(11):2187–2197, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:NSS

[LHGL22] Qilong Liu, Jianfeng He, Lei Gao, and Chaoqian Li. Note on subdirect sums of SDD(p) matrices. Linear Multilinear Algebra, 70(13):2582–2601, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lu:2018:GWS

[LHH18] Lu Lu, Qiongxiang Huang, and Xueyi Huang. On graphs whose smallest distance (signless Laplacian) eigenvalue has large multiplicity. Linear Multilinear Algebra, 66(11):2218– 2231, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leahey:1975:NP

[LHP75] W. J. Leahey, W. C. Herndon, and V. T. Phan. A note on permanents. Linear and Multilinear Algebra, 3(3):193–196, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2021:REC

[LHS^{+}21] Mengfan Liang, Mengyao Hu, Yize Sun, Lin Chen, and Xiaoyu Chen. Real entries of complex Hadamard matrices and mutually unbiased bases in dimension six. Linear Multilinear

Li:2022:SCD

[LHZ22] Chaoqian Li, Zhengyu Huang, and Jianxing Zhao. On Schur complements of Dashnic–Zusmanovich type matrices. Linear Multilinear Algebra, 70(19):4071–4096, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1984:NMR

[Li84] Chi-Kwong Li. A note on Miranda’s results about the characteristic values and the three types of singular values of a complex matrix. Linear and Multilinear Algebra, 16(1–4):297–303, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Mir81].

Li:1985:ESC

[Li85] Chi-Kwong Li. On the extremal solutions for capacitated network problems. Linear and Multilinear Algebra, 18(2):117– 125, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1986:SRC

[Li86] Chi-Kwong Li. The c-spectral c-radial and c-convex matrices. Linear and Multilinear Algebra, 20(1):5–15, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1987:CM

[Li87a] Chi-Kwong Li. The C-convex matrices. Linear and Multilinear Algebra, 21(3):303–312, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1987:IES

[Li87b] Chi-Kwong Li. An inequality on elementary symmetric functions. Linear and Multilinear Algebra, 20(4):373–375, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1987:LOP

[Li87c] Chi-Kwong Li. Linear operators preserving the higher numerical radius of matrices. Linear and Multilinear Algebra, 21 (1):63–73, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1991:IRN

[Li91] Chi-Kwong Li. Inequalities relating norms invariant under unitary similarities. Linear and Multilinear Algebra, 29(3–4): 155–167, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1992:DDG

[Li92] Zhongshan Li. A determinantal description of GCD-closed sets and k-sets. Linear and Multilinear Algebra, 31(1–4):245–250, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1994:NRN

[Li94] Chi-Kwong Li. C-numerical ranges and C-numerical radii. Linear and Multilinear Algebra, 37(1–3):51–82, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1995:CSI

[Li95] Wen Li. Characterizations of singular irreducible M-matrices. Linear and Multilinear Algebra, 38(3):241–247, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1996:SCT

[Li96] Chi Kwong Li. Some convexity theorems for the generalized numerical ranges. Linear and Multilinear Algebra, 40(3):235– 240, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1998:NRN

[Li98] Chi-Kwong Li. q-numerical ranges of normal and convex matrices. Linear and Multilinear Algebra, 43(4):377–384, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2011:MPI

[Li11] Yuan Li. The Moore–Penrose inverses of lower triangular operators. Linear Multilinear Algebra, 59(3):229–235, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2021:ICM

[Li21a] Cui-Xia Li. The improved convergence of MSMMAOR method for linear complementarity problems. Linear Multilinear Algebra, 69(1):1–8, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2021:JSM

[Li21b] Fei Li. On the Jacobians of singular matrix decomposition and its application. Linear Multilinear Algebra, 69(8):1521–1533, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:1980:CLC

[Lia80] Bob C. Liang. A classification of linear control systems. Linear and Multilinear Algebra, 8(3):261–264, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liao:2023:BPR

[Lia23] Ping Liao. Bounds for the Perron root of positive matrices. Linear Multilinear Algebra, 71(11):1849–1857, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1976:LTS

[Lim76a] M. H. Lim. Linear transformation on symmetry classes of tensors. Linear and Multilinear Algebra, 3(4):267–280, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1976:UTS

[Lim76b] M. H. Lim. A uniqueness theorem in symmetry classes. Linear and Multilinear Algebra, 4(2):85–88, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1977:AEA

[Lim77] M. H. Lim. Antiderivations on the exterior algebras. Linear and Multilinear Algebra, 4(4):251–259, 1977. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1979:LTS

[Lim79a] M. H. Lim. Linear transformations on symmetric matrices. Linear and Multilinear Algebra, 7(1):47–57, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1979:NRB

[Lim79b] M. H. Lim. A note on the relation between the determinant and the permanent. Linear and Multilinear Algebra, 7(2):145– 147, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1982:LMS

[Lim82] M. H. Lim. Linear mappings on symmetry classes of tensors. Linear and Multilinear Algebra, 12(2):109–123, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1990:LMS

[Lim90a] M. H. Lim. Linear mappings on second symmetric product spaces that preserve rank less than or equal to one. Linear and Multilinear Algebra, 26(3):187–193, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1990:LTS

[Lim90b] M. H. Lim. Linear transformations on symmetry classes of tensors II. Linear and Multilinear Algebra, 26(3):195–205, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1992:CRT

[Lim92a] M. H. Lim. Chapter 2: rank and tensor rank. Linear and Multilinear Algebra, 33(1–2):7–21, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1992:CDS

[Lim92b] M. H. Lim. Conditions on decomposable symmetric tensors as an algebraic variety. Linear and Multilinear Algebra, 32(3–4): 249–252, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1992:DSA

[Lim92c] M. H. Lim. Derivations on the symmetric algebras. Linear and Multilinear Algebra, 33(3–4):137–143, 1992. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:1997:LPD

[Lim97] M. H. Lim. Linear preservers of decomposable symmetric tensors. Linear and Multilinear Algebra, 42(4):291–302, 1997.

Lim:2000:DIC

[Lim00] M. H. Lim. Dimension inequalities for the closure of the image of a multilinear function. Linear and Multilinear Algebra, 47(3):231–234, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:2001:LPO

[Lim01] M. H. Lim. Linear preservers of orthogonal decomposable tensors. Linear and Multilinear Algebra, 48(4):333–354, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:2010:APF

[Lim10] Ming-Huat Lim. Adjacency preserving functions on symmetric elements and linear preservers of non-zero decomposable symmetric tensors. Linear Multilinear Algebra, 58(4):481–496, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:2015:NLP

[Lim15] Ming-Huat Lim. A note on linear preservers of certain ranks of tensor products of matrices. Linear Multilinear Algebra, 63(7):1442–1447, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:2023:ALM

[Lim23] Yongdo Lim. The Ando–Li–Mathias geometric mean and infinite product of hollow matrices. Linear Multilinear Algebra, 71(4):552–568, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2021:WJD

[Lin21] Jheng-Huei Lin. Weak Jordan derivations of prime rings. Linear Multilinear Algebra, 69(8):1422–1445, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lipschutz:1987:RMR

[Lip87] Seymour Lipschutz. Revisiting matrix representations of the braid group. Linear and Multilinear Algebra, 21(2):163–172,

Liping:1996:SLM

[Lip96] Huang Liping. The solvability of linear matrix equations over a central simple algebra. Linear and Multilinear Algebra, 40(4): 353–363, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liping:1998:MPM

[Lip98] Huang Liping. The minimal polynomial of a matrix. Linear and Multilinear Algebra, 45(2–3):99–107, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

LitadaSilva:2017:CPE

[Lit17] João Lita da Silva. On the characteristic polynomial, eigenvectors and determinant of heptadiagonal matrices. Linear Multilinear Algebra, 65(9):1852–1866, 2017. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:1992:FIE

[Liu92] Bolian Liu. On fully indecomposable exponent for primitive boolean matrices with symmetric ones. Linear and Multilinear Algebra, 31(1–4):131–138, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2015:SCP

[Liu15] Lei Liu. Strong commutativity preserving maps on von Neumann algebras. Linear Multilinear Algebra, 63(3):490–496, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2016:LLD

[Liu16a] Lei Liu. 2-local Lie derivations on semi-finite factor von Neumann algebras. Linear Multilinear Algebra, 64(9):1679–1686, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2016:CCN

[Liu16b] Lei Liu. Characterization of centralizers on nest subalgebras of von Neumann algebras by local action. Linear Multilinear Algebra, 64(3):383–392, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2017:ECG

[Liu17] Cheng-Kai Liu. An Engel condition with b-generalized derivations. Linear Multilinear Algebra, 65(2):300–312, 2017. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2018:LLA

[Liu18a] Lei Liu. 2-local Lie*-automorphisms on factors. Linear Multilinear Algebra, 66(11):2208–2214, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2018:EEN

[Liu18b] Xifu Liu. The eigenvalues and eigenvectors of nonsingular tensors, similar tensors and tensor products. Linear Multilinear Algebra, 66(9):1717–1725, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2019:LLD

[Liu19] Lei Liu. 2-local Lie derivations of nest subalgebras of factors. Linear Multilinear Algebra, 67(3):448–455, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:NLC

[Liu22a] Lei Liu. On nonlinear Lie centralizers of generalized matrix algebras. Linear Multilinear Algebra, 70(14):2693–2705, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:NPD

[Liu22b] Xifu Liu. Norm-preserving dilation theorems for a block positive semidefinite (definite) matrix. Linear Multilinear Algebra, 70(22):7770–7777, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Livshits:1994:BMG

[Liv94] Leo Livshits. Block-matrix generalizations of infinitedimensional Schur products and Schur multipliers. Linear and Multilinear Algebra, 38(1–2):59–78, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2011:MPP

[LJ11] Lei Liu and Guo-Xing Ji. Maps preserving product X^{∗}Y + Y X^{∗} on factor von Neumann algebras. Linear Multilinear Algebra, 59(9):951–955, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leventides:2016:ADC

[LK16] John Leventides and Nicos Karcanias. Approximate decomposability in and the canonical decomposition of 3-vectors. Linear Multilinear Algebra, 64(12):2378–2405, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lawson:2021:CAA

[LK21] Jimmie Lawson and Eyüp Kizil. Characterizations of automorphic and anti-automorphic involutions of the quaternions. Linear Multilinear Algebra, 69(11):1975–1980, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lawson:2023:CIA

[LK23] Jimmie Lawson and Eyüp Kizil. Classification of involutive automorphisms and anti-automorphisms of the Lie algebra of quaternions. Linear Multilinear Algebra, 71(9):1536–1544, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2022:EPS

[LKLC22] Yan Liang, Rihuan Ke, Wen Li, and Lu-Bin Cui. On the extreme points of slice-stochastic tensors. Linear Multilinear Algebra, 70(19):3969–3981, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:1978:NIP

[LL78] Raphael Loewy and David London. A note on an inverse problem for nonnegative matrices. Linear and Multilinear Algebra, 6(1):83–90, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laffey:1990:LT

[LL90] Thomas J. Laffey and Raphael Loewy. Linear transformations. Linear and Multilinear Algebra, 26(3):181–186, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liao:2010:LRA

[LL10a] Ping-Bao Liao and Cheng-Kai Liu. Lie rings of (anti- )symmetric derivations of commutative rings. Linear Multilinear Algebra, 58(5):589–598, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2010:BLS

[LL10b] Huiqing Liu and Mei Lu. Bounds for the Laplacian spectral radius of graphs. Linear Multilinear Algebra, 58(1):113–119, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2011:SCP

[LL11] Cheng-Kai Liu and Pao-Kuei Liau. Strong commutativity preserving generalized derivations on Lie ideals. Linear Multilinear Algebra, 59(8):905–915, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2014:NRI

[LL14a] Chaoqian Li and Yaotang Li. New regions including eigenvalues of Toeplitz matrices. Linear Multilinear Algebra, 62(2): 229–241, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2014:EWM

[LL14b] Xin Liang and Ren-Cang Li. Extensions of Wielandt’s minmax principles for positive semi-definite pencils. Linear Multilinear Algebra, 62(8):1032–1048, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2015:BDS

[LL15] Huiqiu Lin and Xiwen Lu. Bounds on the distance signless Laplacian spectral radius in terms of clique number. Linear Multilinear Algebra, 63(9):1750–1759, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lau:2016:LRO

[LL16a] Jinting Lau and Ming Huat Lim. Linear rank one preservers on tensor products of Hermitian matrices. Linear Multilinear Algebra, 64(4):717–726, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2016:ELS

[LL16b] Chaoqian Li and Yaotang Li. An eigenvalue localization set for tensors with applications to determine the positive (semi- )definiteness of tensors. Linear Multilinear Algebra, 64(4):587– 601, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2017:NPB

[LL17] Hanyu Li and Peng Lv. New perturbation bounds and condition numbers for the hyperbolic QR factorization. Linear Multilinear Algebra, 65(8):1540–1553, 2017. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2018:GDP

[LL18] Cheng-Kai Liu and Pao-Kuei Liau. Generalized derivations preserving quasinilpotent elements in Banach algebras. Linear Multilinear Algebra, 66(9):1888–1908, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2020:ORF

[LL20] Dongwei Li and Jinsong Leng. Operator representations of g-frames in Hilbert spaces. Linear Multilinear Algebra, 68(9): 1861–1877, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:LMP

[LL22a] Hualu Liu and Xiusheng Liu. LCP of matrix product codes. Linear Multilinear Algebra, 70(22):7611–7622, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:NLM

[LL22b] Lele Liu and Linyuan Lu. The α-normal labelling method for computing the p-spectral radii of uniform hypergraphs. Linear Multilinear Algebra, 70(9):1648–1672, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Llarull:1996:ICF

[Lla96] Marcelo Llarull. Irreducible components of L^{2} functions on Hopf bundles. Linear and Multilinear Algebra, 41(2):113–135, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2021:IAW

[LLC21] Chun-Mei Li, Lin-Zhang Lu, and Zhen Chen. Iterative algorithm for the Q-weighted nonnegative matrix factorization. Linear Multilinear Algebra, 69(11):2130–2142, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:SFM

[LLC22] Dongmei Li, Jinwang Liu, and Delin Chu. The Smith form of a multivariate polynomial matrix over an arbitrary coefficient field. Linear Multilinear Algebra, 70(2):366–379, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2017:DBD

[LLD17] Huiqiu Lin, Dan Li, and Kinkar Ch. Das. Distance between distance spectra of graphs. Linear Multilinear Algebra, 65(12): 2538–2550, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2008:ESS

[LLF08] Hong-Hai Li, Jiong-Sheng Li, and Yi-Zheng Fan. The effect on the second smallest eigenvalue of the normalized Laplacian of a graph by grafting edges. Linear Multilinear Algebra, 56(6): 627–638, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2014:NLJ

[LLF14] Changjing Li, Fangyan Lu, and Xiaochun Fang. Non-linear ξ-Jordan*-derivations on von Neumann algebras. Linear Multilinear Algebra, 62(4):466–473, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2018:SCG

[LLF18] Weijun Liu, Minmin Liu, and Lihua Feng. Spectral conditions for graphs to be β-deficient involving minimum degree. Linear Multilinear Algebra, 66(4):792–802, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2016:SSN

[LLGL16] Chaoqian Li, Qilong Liu, Lei Gao, and Yaotang Li. Subdirect sums of Nekrasov matrices. Linear Multilinear Algebra, 64(2):

Liu:2022:FMN

[LLGX22] Ruifang Liu, Hong-Jian Lai, Litao Guo, and Jie Xue. Fractional matching number and spectral radius of nonnegative matrices of graphs. Linear Multilinear Algebra, 70(19):4133– 4145, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2018:SSF

[LLHS18] Dongwei Li, Jinsong Leng, Tingzhu Huang, and Guomin Sun. On sum and stability of g-frames in Hilbert spaces. Linear Multilinear Algebra, 66(8):1578–1592, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2009:MBI

[LLL09] Yanting Liang, Bolian Liu, and Hong-Jian Lai. Multi-g base index of primitive anti-symmetric sign pattern matrices. Linear Multilinear Algebra, 57(6):535–546, 2009. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2015:GTB

[LLL15] Chaoqian Li, Qingbing Liu, and Yaotang Li. Geršgorin-type and Brauer-type eigenvalue localization sets of stochastic matrices. Linear Multilinear Algebra, 63(11):2159–2170, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Luo:2023:DMS

[LLLY23] Jiajie Luo, Lina Liu, Sisi Li, and Yongxin Yuan. A direct method for the simultaneous updating of finite element mass, damping and stiffness matrices. Linear Multilinear Algebra, 71(12):1979–1993, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laugerotte:2020:MRF

[LLMN20] É. Laugerotte, J.-G. Luque, L. Mignot, and F. Nicart. Multilinear representations of free pros. Linear Multilinear Algebra, 68(10):1911–1955, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Langer:2005:SDT

[LLMT05] Heinz Langer, Matthias Langer, Alexander Markus, and Christiane Tretter. Spectrum of definite type of self-adjoint operators in Krein spaces. Linear Multilinear Algebra, 53(2): 115–136, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2018:SLS

[LLP18] Yawen Li, Yao Liu, and Xing Peng. Signless Laplacian spectral radius and Hamiltonicity of graphs with large minimum degree. Linear Multilinear Algebra, 66(10):2011–2023, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liao:2016:NIM

[lLpW16] Ya lin Liao and Xue ping Wang. Note on invertible matrices over commutative semirings. Linear Multilinear Algebra, 64(3):477–483, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2010:BGB

[LLS10] Qian Li, Bolian Liu, and Jeffrey Stuart. Bounds on the k-th generalized base of a primitive sign pattern matrix. Linear Multilinear Algebra, 58(3):355–366, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leake:2014:ICB

[LLS14] Timothy Leake, Brian Lins, and Ilya M. Spitkovsky. Inverse continuity on the boundary of the numerical range. Linear Multilinear Algebra, 62(10):1335–1345, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrections and additions [LLS16].

Leake:2016:CAI

[LLS16] Timothy Leake, Brian Lins, and Ilya M. Spitkovsky. Corrections and additions to ‘Inverse continuity on the boundary of the numerical range’. Linear Multilinear Algebra, 64(1): 100–104, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [LLS14].

Lang:1993:SHG

[LLT93] Mong-Lung Lang, Chong-Hai Lim, and Ser-Peow Tan. Subgroups of the Hecke groups with small index. Linear and Multilinear Algebra, 35(1):75–77, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2018:TCP

[LLV18] Dongdong Liu, Wen Li, and Seak-Weng Vong. Tensor complementarity problems: the GUS-property and an algorithm. Linear Multilinear Algebra, 66(9):1726–1749, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2007:ETL

[LLW07] Chi-Kwong Li, Xin-Guo Liu, and Xue-Feng Wang. Extension of the total least square problem using general unitarily invariant norms. Linear Multilinear Algebra, 55(1):71–79, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lam:2011:ESC

[LLW^{+}11] James Lam, Zhao Li, Yimin Wei, Jun e Feng, and Kwok Wai Chung. Estimates of the spectral condition number. Linear Multilinear Algebra, 59(3):249–260, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liao:2017:KCK

[LLWX17] Jun Liao, Heguo Liu, Zuohui Wu, and Huan Xiong. Kernels for certain kinds of linear transformations. Linear Multilinear Algebra, 65(6):1131–1141, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:NBT

[LLY22] Xiusheng Liu, Hualu Liu, and Long Yu. New binary and ternary LCD codes from matrix-product codes. Linear Multilinear Algebra, 70(5):809–823, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2013:NIM

[LLZ13] Chaoqian Li, Yaotang Li, and Ruijuan Zhao. New inequalities for the minimum eigenvalue of M-matrices. Linear Multilinear Algebra, 61(9):1267–1279, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:LST

[LLZ22] Yipeng Li, Xiaogang Liu, and Shenggui Zhang. Laplacian state transfer in edge coronas. Linear Multilinear Algebra, 70(6):1023–1046, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2019:TIA

[lLZjZ19] Mao lin Liang, Bing Zheng, and Rui juan Zhao. Tensor inversion and its application to the tensor equations with Einstein product. Linear Multilinear Algebra, 67(4):843–870, 2019. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lancaster:1985:KBO

[LM85] P. Lancaster and J. Maroulas. The kernel of the Bezoutian for operator polynomials. Linear and Multilinear Algebra, 17(3– 4):181–201, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

London:1989:PDS

[LM89] David London and Henryk Minc. On the permanent of doubly stochastic matrices with zero diagonal. Linear and Multilinear Algebra, 24(4):289–300, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1993:LOPa

[LM93] Chi-Kwong Li and Dennis I. Merino. Linear operators preserving unitary t-congruence (orthogonal similarity) on complex (real) matrices. Linear and Multilinear Algebra, 35(2): 83–105, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:1995:CSP

[LM95] Cai Jun Liang and Zhang Ke Min. The characterization of symmetric primitive matrices with exponent 2n − 2r(≥ n). Linear and Multilinear Algebra, 39(4):391–396, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1996:MII

[LM96] Chi-Kwong Li and Roy Mathias. Matrix inequalities involving a positive linear map. Linear and Multilinear Algebra, 41(3):

Li:1998:SIT

[LM98] Chi-Kwong Li and Roy Mathias. Some interlacing theorems on the Schur complement. Linear and Multilinear Algebra, 44(4): 373–382, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2003:LPF

[LM03] Chi-Kwong Li and Thomas Milligan. Linear preservers of finite reflection groups. Linear Multilinear Algebra, 51(1):49–81, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2010:STD

[LM10] Juan Liu and Jixiang Meng. Spectra of transformation digraphs of a regular digraph. Linear Multilinear Algebra, 58(5): 555–561, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2019:HDG

[LM19] Lily Li Liu and Lili Mu. Hankel determinants for generating functions of Catalan-like numbers. Linear Multilinear Algebra, 67(3):499–509, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:2022:NME

[LM22] Hosoo Lee and Jie Meng. On the nonlinear matrix equation X^{p} = A#_{t}(M^{T} (X^{−1} + B)^{−1}M). Linear Multilinear Algebra, 70(2):250–263, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2017:SSW

[LMLL17] Chaoqian Li, Ruidan Ma, Qilong Liu, and Yaotang Li. Subdirect sums of weakly chained diagonally dominant matrices. Linear Multilinear Algebra, 65(6):1220–1231, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lins:2001:NMP

[LMMR01a] Brian Lins, Patrick Meade, Christian Mehl, and Leiba Rodman. Normal matrices and polar decompositions indefinite inner products. Linear and Multilinear Algebra, 49(1):45–89, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lins:2001:RPI

[LMMR01b] Brian Lins, Patrick Meade, Christian Mehl, and Leiba Rodman. Research problem: indefinite inner product normal matrices. Linear and Multilinear Algebra, 49(3):261–268, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Livshits:2015:SVW

[LMMR15] L. Livshits, G. W. MacDonald, L. W. Marcoux, and H. Radjavi. A spatial version of Wedderburn’s principal theorem. Linear Multilinear Algebra, 63(6):1216–1241, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1994:GNR

[LMR94a] Chi-Kwong Li, Paras P. Mehta, and Leiba Rodman. A generalized numerical range: the range of a constrained sesquilinear form. Linear and Multilinear Algebra, 37(1–3):25–49, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1994:LOP

[LMR94b] Chi-Kwong Li, Paras P. Mehta, and Leiba Rodman. Linear operators preserving the inner and outer c-spectral radii. Linear and Multilinear Algebra, 36(3):195–204, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Livshits:2000:CTM

[LMR00] L. Livshits, G. Macdonald, and H. Radjavi. Cone-transitive matrix semigroups. Linear and Multilinear Algebra, 47(4):313– 350, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leahey:1980:NTL

[LN80] W. J. Leahey and J. E. Nymann. A note on a theorem of Lindström. Linear and Multilinear Algebra, 8(3):259–260, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1998:SRN

[LN98] Chi-Kwong Li and Hiroshi Nakazato. Some results on the qnumerical range. Linear and Multilinear Algebra, 43(4):385– 409, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lopez:2010:OWD

[LN10] Eloy A. Lopez and Michael G. Neubauer. D-optimal (0, 1)- weighing designs for 10 objects. Linear Multilinear Algebra, 58(2):151–171, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2014:LPD

[LN14] Wen Li and Michael K. Ng. On the limiting probability distribution of a transition probability tensor. Linear Multilinear Algebra, 62(3):362–385, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2016:SAE

[LN16] Binlong Li and Bo Ning. Spectral analogues of erdős’ and Moon–Moser’s theorems on Hamilton cycles. Linear Multilinear Algebra, 64(11):2252–2269, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:1973:MEL

[Loe73] Raphael Loewy. The minimal eigenvalue of the Lyapunov transform. Linear and Multilinear Algebra, 1(3):213–229, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:1974:RLT

[Loe74] Raphael Loewy. On ranges of Lyapunov transformations II. Linear and Multilinear Algebra, 2(3):227–237, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:1987:DNS

[Loe87] R. Loewy. On determinants of nonprincipal submatrices of Hermitian matrices. Linear and Multilinear Algebra, 20(3): 219–228, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:1992:CIP

[Loe92] Raphael Loewy. Chapter 3: inertia preservers. Linear and Multilinear Algebra, 33(1–2):23–30, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:1993:LMWa

[Loe93a] Raphael Loewy. Linear mappings which are rank-k nonincreasing. Linear and Multilinear Algebra, 34(1):21–32, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:1993:LMWb

[Loe93b] Raphael Loewy. Linear mappings which are rank-k nonincreasing, II. Linear and Multilinear Algebra, 36(2):115–123, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:2001:LSS

[Loe01] Raphael Loewy. Large spaces of symmetric matrices of bounded rank are decomposable. Linear and Multilinear Algebra, 48(4):355–382, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

London:1973:CBP

[Lon73] David London. On a connection between permanent function and polynomials. Linear and Multilinear Algebra, 1(3):231– 240, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

London:1976:DAD

[Lon76] David London. On derivations arising in differential equations. Linear and Multilinear Algebra, 4(3):179–189, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

London:1980:VWC

[Lon80] David London. On the van der Waerden conjecture for matrices of rank two. Linear and Multilinear Algebra, 8(4):281–289, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

London:1981:DOC

[Lon81] David London. On the – Doković conjecture for matrices of rank two. Linear and Multilinear Algebra, 9(4):317–327, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

London:1983:ISN

[Lon83] David London. An inequality for the spectral norm of certain matrices. Linear and Multilinear Algebra, 14(1):37–44, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

London:1990:MSS

[Lon90] David London. On matrices stochastically similar to matrices with equal diagonal elements. Linear and Multilinear Algebra, 26(1–2):107–117, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

London:1997:DIP

[Lon97] David London. A determinantal inequality and its permanental counterpart. Linear and Multilinear Algebra, 42(4): 281–290, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:SCO

[LOO22] Saihua Liu, Chen Ouyang, and Jianping Ou. Sufficient conditions for optimally and super m-restricted edge-connected graphs with given girth. Linear Multilinear Algebra, 70(6): 1146–1158, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lopatin:2011:MGS

[Lop11a] A. A. Lopatin. On minimal generating systems for matrix O (3)-invariants. Linear Multilinear Algebra, 59(1):87–99, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lopatin:2011:OIS

[Lop11b] A. A. Lopatin. Orthogonal invariants of skew-symmetric matrices. Linear Multilinear Algebra, 59(8):851–862, 2011. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lopatin:2018:MSG

[Lop18] Artem Lopatin. Minimal system of generators for O(4)- invariants of two skew-symmetric matrices. Linear Multilinear Algebra, 66(2):347–356, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lorenzini:2000:APL

[Lor00] Dino Lorenzini. Arithmetical properties of Laplacians of graphs. Linear and Multilinear Algebra, 47(4):281–306, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lorch:2014:LMR

[Lor14] John Lorch. Linear magic rectangles. Linear Multilinear Algebra, 62(4):530–537, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lovitz:2021:DCM

[Lov21] Benjamin Lovitz. On decomposable correlation matrices. Linear Multilinear Algebra, 69(11):2115–2129, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1993:LOPb

[LP93] Chi-Kwong Li and Stephen Pierce. Linear operators preserving certain singular matrix sets. Linear and Multilinear Algebra, 36(1):19–25, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2000:NSL

[LP00] Jiong-Sheng Li and Yong-Liang Pan. A note on the second largest eigenvalue of the Laplacian matrix of a graph. Linear and Multilinear Algebra, 48(2):117–121, 2000. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2002:ADR

[LP02] Chi-Kwong Li and Edward Poon. Additive decomposition of real matrices. Linear Multilinear Algebra, 50(4):321–326, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2003:PSH

[LP03] Chi-Kwong Li and Yiu-Tung Poon. Principal submatrices of a Hermitian matrix. Linear Multilinear Algebra, 51(2):199–208, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lal:2007:MLS

[LP07] A. K. Lal and K. L. Patra. Maximizing Laplacian spectral radius over trees with fixed diameter. Linear Multilinear Algebra, 55(5):457–461, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2011:ICP

[LP11] Chi-Kwong Li and Yiu-Tung Poon. Interpolation by completely positive maps. Linear Multilinear Algebra, 59(10): 1159–1170, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lins:2016:CSI

[LP16] Brian Lins and Parth Parihar. Continuous selections of the inverse numerical range map. Linear Multilinear Algebra, 64 (1):87–99, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:OTI

[LP22a] Yongtao Li and Yuejian Peng. An Oppenheim type inequality for positive definite block matrices. Linear Multilinear Algebra, 70(19):4455–4466, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loreaux:2022:COCb

[LP22b] Jireh Loreaux and Sasmita Patnaik. Closedness of the orbitclosed C-numerical range and submajorization. Linear Multilinear Algebra, 70(21):7204–7223, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loreaux:2022:COCa

[LP22c] Jireh Loreaux and Sasmita Patnaik. Convexity of the orbitclosed C-numerical range and majorization. Linear Multilinear Algebra, 70(19):4707–4750, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2023:SSC

[LP23] Yongtao Li and Yuejian Peng. Sufficient spectral conditions for graphs being k-edge-Hamiltonian or k-Hamiltonian. Linear Multilinear Algebra, 71(13):2093–2113, 2023. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2017:ILS

[LPH17] Jia-Bao Liu, Xiang-Feng Pan, and Fu-Tao Hu. The 1-inverse of the Laplacian of subdivision-vertex and subdivision-edge coronae with applications. Linear Multilinear Algebra, 65(1): 178–191, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2008:RDS

[LPS08] Chi-Kwong Li, Yiu-Tung Poon, and Nung-Sing Sze. Ranks and determinants of the sum of matrices from unitary orbits. Linear Multilinear Algebra, 56(1–2):105–130, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2009:CHR

[LPS09] Chi-Kwong Li, Yiu-Tung Poon, and Nung-Sing Sze. Condition for the higher rank numerical range to be non-empty. Linear Multilinear Algebra, 57(4):365–368, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2012:GII

[LPS12] Chi-Kwong Li, Yiu-Tung Poon, and Nung-Sing Sze. Generalized interlacing inequalities. Linear Multilinear Algebra, 60 (11–12):1245–1254, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2013:JJH

[LPS13] Jiankui Li, Zhidong Pan, and Qihua Shen. Jordan and Jordan higher all-derivable points of some algebras. Linear Multilinear Algebra, 61(6):831–845, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lopez-Permouth:2021:GMA

[LPS21] Sergio R. López-Permouth and Benjamin Stanley. Graph Magma Algebras have no projective bases. Linear Multilinear Algebra, 69(11):1997–2005, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2011:SNN

[LPT11] Chi-Kwong Li, Yiu-Tung Poon, and Masaru Tominaga. Spectra, norms and numerical ranges of generalized quadratic operators. Linear Multilinear Algebra, 59(10):1077–1104, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lebtahi:2013:SER

[LPT13] Leila Lebtahi, Pedro Patrício, and Néstor Thome. Special elements in a ring related to Drazin inverses. Linear Multilinear Algebra, 61(8):1017–1027, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lebtahi:2014:DPO

[LPT14] Leila Lebtahi, Pedro Patrício, and Néstor Thome. The diamond partial order in rings. Linear Multilinear Algebra, 62(3): 386–395, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2014:P

[LPW14] Chi-Kwong Li, Yiu-Tung Poon, and Ngai-Ching Wong. Preface. Linear Multilinear Algebra, 62(5):553, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2017:GHT

[LPW17] Jun-Tong Liu, Yiu-Tung Poon, and Qing-Wen Wang. A generalized Hölder type eigenvalue inequality. Linear Multilinear Algebra, 65(10):2145–2151, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2010:NUC

[LPWX10] Hao Liang, Yong-Liang Pan, Jian Wang, and Jun-Ming Xu. A note on unimodular congruence of the Laplacian matrix of a graph. Linear Multilinear Algebra, 58(4):497–501, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2023:CSW

[LPX23] Junming Liu, Saminathan Ponnusamy, and Huayou Xie. Complex symmetric weighted composition-differentiation operators. Linear Multilinear Algebra, 71(5):737–755, 2023. CO-

Lavrauw:2013:RT

[LPZ13] Michel Lavrauw, Andrea Pavan, and Corrado Zanella. On the rank of 3 × 3×3-tensors. Linear Multilinear Algebra, 61(5): 646–652, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liping:1995:MES

[LQ95] Huang Liping and Zeng Qingguang. The matrix equation AXB + CY D = E over a simple artinian ring. Linear and Multilinear Algebra, 38(3):225–232, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Longstaff:1989:ECM

[LR89] W. E. Longstaff and H. Radjavi. On the equation AB = BX in collections of matrices. Linear and Multilinear Algebra, 25(3): 173–184, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1991:ISG

[– DLR91] Dragomir Ž. – Doković, Chi-Kwong Li, and Leiba Rodman. Isometries of symmetric gauge functions. Linear and Multilinear Algebra, 30(1–2):81–92, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lerer:1996:SFL

[LR96] Leonid Lerer and Leiba Rodman. Symmetric factorizations and localization of zeroes of rational matrix functions. Linear and Multilinear Algebra, 40(3):259–281, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2004:MPN

[LR04] Chi-Kwong Li and Leiba Rodman. Multiplicative preservers of C-numerical ranges and radii. Linear Multilinear Algebra, 52(3–4):265–279, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laliena:2013:SSW

[LR13] Jesús Laliena and Roberto Rizzo. Superalgebras with superinvolution whose symmetric or skewsymmetric elements are regular. Linear Multilinear Algebra, 61(9):1280–1286, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ladra:2017:ACM

[LR17] M. Ladra and U. A. Rozikov. Algebras of cubic matrices. Linear Multilinear Algebra, 65(7):1316–1328, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ljubenovic:2021:LPD

[LRD21] Martin Z. Ljubenović, Dragan S. Rakić, and Dragan S. Djordjević. Linear preservers of DSS-weak majorization on discrete Lebesgue space, when I is an infinite set. Linear Multilinear Algebra, 69(14):2657–2673, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lancaster:1987:EMT

[LRR87] P. Lancaster, A. C. M. Ran, and L. Rodman. An existence and monotonicity theorem for the discrete algebraic matrix Riccati equation. Linear and Multilinear Algebra, 20(4):353– 361, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lerer:1991:ITM

[LRT91] Leonid Lerer, Leiba Rodman, and Miron Tismenetsky. Inertia theorems for matrix polynomials. Linear and Multilinear Algebra, 30(3):157–182, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2008:PCR

[LRT08] Chi-Kwong Li, Leiba Rodman, and Christiane Tretter. Preface and conference report. Linear Multilinear Algebra, 56(1–2): 1–2, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lesnjak:1996:QAOa

[LŠ96a] Gorazd Lesnjak and Peter Šemrl. Quasidirect addition of operators. Linear and Multilinear Algebra, 40(4):373–377, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lesnjak:1996:QAOb

[LŠ96b] Gorazd Lesnjak and Peter Šemrl. Quasidirect addition of operators. Linear and Multilinear Algebra, 41(4):377–381, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2002:NRI

[LŠ02] Chi-Kwong Li and Peter Šemrl. Numerical radius isometries. Linear Multilinear Algebra, 50(4):307–314, 2002. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2004:IRN

[LS04] Chi-Kwong Li and Nung-Sing Sze. Inclusion regions for numerical ranges and linear preservers. Linear Multilinear Algebra, 52(3–4):235–249, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2009:GEE

[LS09] Chi-Kwong Li and Wasin So. Graphs equienergetic with edgedeleted subgraphs. Linear Multilinear Algebra, 57(7):683–693, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Laffey:2010:NRS

[LS10] Thomas J. Laffey and Helena Smigoc. Nonnegatively realizable spectra with two positive eigenvalues. Linear Multilinear Algebra, 58(8):1053–1069, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2012:SBD

[LS12] Huiqiu Lin and Jinlong Shu. Sharp bounds on distance spectral radius of graphs. Linear Multilinear Algebra, 60(9):1115, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2013:SBD

[LS13] Huiqiu Lin and Jinlong Shu. Sharp bounds on distance spectral radius of graphs. Linear Multilinear Algebra, 61(4):442– 447, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2015:RHO

[LS15] Jin Liang and Guanghua Shi. Refinements of the Heinz operator inequalities. Linear Multilinear Algebra, 63(7):1337–1344, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:2016:EAA

[LS16] Doohann Lee and Youngkwon Song. Extension of algebras and its applications to the real Clifford algebras. Linear Multilinear Algebra, 64(11):2241–2251, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2017:PGM

[LS17] Minghua Lin and Fangfang Sun. A property of the geometric mean of accretive operators. Linear Multilinear Algebra, 65(3): 433–437, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2021:MNC

[LS21a] Shuchao Li and Wanting Sun. Matching number, connectivity and eigenvalues of distance signless Laplacians. Linear Multilinear Algebra, 69(1):74–92, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Loewy:2021:SNS

[LS21b] Raphael Loewy and Oren Spector. Some notes on the spectra of nonnegative symmetric 5 × 5 matrices. Linear Multilinear Algebra, 69(7):1363–1382, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:SBS

[LS22a] Jinlu Li and Jiancai Sun. Super-biderivations on the super Heisenberg–Virasoro algebra. Linear Multilinear Algebra, 70 (21):6500–6510, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:SBI

[LS22b] S. C. Li and W. T. Sun. Some bounds on the A_{α}-index of connected graphs with fixed order and size. Linear Multilinear Algebra, 70(20):5859–5878, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:SDI

[LS22c] Yan Liu and Yuming Shi. Stability of deficiency indices of Hermitian subspaces under relatively bounded perturbations. Linear Multilinear Algebra, 70(7):1356–1368, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1995:MPD

[LSOvdD95] Chi-Kwong Li, David P. Stanford, D. D. Olesky, and P. van den Driessche. Minimum positive determinant of integer matrices with constant row and column sums. Linear and Multilinear Algebra, 40(2):163–170, 1995. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2001:LOPa

[LŠS01] Chi-Kwong Li, Peter Šemrl, and Graca Soares. Linear operators preserving the numerical range (radius) on triangular matrices. Linear and Multilinear Algebra, 48(4):281–292, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:2016:P

[LSS16] Lek-Heng Lim, Jiayu Shao, and Nung-Sing Sze. Preface. Linear Multilinear Algebra, 64(4):573, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1989:CMC

[LST89] Chi-Kwong Li, Chen-Han Sung, and Nam-Kiu Tsing. c-convex matrices: characterizations, inclusion relations and normality. Linear and Multilinear Algebra, 25(4):275–287, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lazzarin:2023:LEE

[LST23] João Lazzarin, Oscar F. Sosa, and Fernando C. Tura. Laplacian eigenvalues of equivalent cographs. Linear Multilinear Algebra, 71(6):1003–1014, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Luo:2019:PEP

[LSX19] Wei Luo, Chuanning Song, and Qingxiang Xu. Perturbation estimation for the parallel sum of Hermitian positive semidefinite matrices. Linear Multilinear Algebra, 67(10):1971–

London:1986:BSN

[LSZ86] David London, Binyamin Schwarz, and Abraham Zaks. Bounds for the spectral norm of certain matrices. Linear and Multilinear Algebra, 19(1):57–66, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1987:UIN

[LT87] Chi-Kwong Li and Nam-Kiu Tsing. On the unitarily invariant norms and some related results. Linear and Multilinear Algebra, 20(2):107–119, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1988:DBS

[LT88a] Chi-Kwong Li and Nam-Kiu Tsing. Duality between some linear preservers problems: the invariance of the c-numerical range, the c-numerical radius and certain matrix sets. Linear and Multilinear Algebra, 23(4):353–362, 1988. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1988:LOPb

[LT88b] Chi-Kwong Li and Nam-Kiu Tsing. Linear operators preserving the decomposable numerical radius. Linear and Multilinear Algebra, 23(4):333–341, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1988:LOPa

[LT88c] Chi-Kwong Li and Nam-Kiu Tsing. Linear operators that preserve the c-numerical range or radius of matrices. Linear and Multilinear Algebra, 23(1):27–46, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1988:SIR

[LT88d] Chi-Kwong Li and Nam-Kiu Tsing. Some isometries of rectangular complex matrices. Linear and Multilinear Algebra, 23 (1):47–53, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1989:DCH

[LT89a] Chi-Kwong Li and Nam-Kiu Tsing. Distance to the convex hull of the unitary orbit with respect to unitary similarity invariant norms. Linear and Multilinear Algebra, 25(2):93– 103, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1989:NIU

[LT89b] Chi-Kwong Li and Nam Kiu Tsing. Norms that are invariant under unitary similarities and the C-numerical radii. Linear and Multilinear Algebra, 24(3):209–222, 1989. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1990:LOPb

[LT90a] Chi-Kwong Li and Nam-Kiu Tsing. Linear operators preserving certain functions on singular values of matrices. Linear and Multilinear Algebra, 26(1–2):133–143, 1990. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1990:LOPa

[LT90b] Chi-Kwong Li and Nam-Kiu Tsing. Linear operators preserving unitarily invariant norms of matrices. Linear and Multilinear Algebra, 26(1–2):119–132, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1990:LOPc

[LT90c] Chi-Kwong Li and Nam Kiu Tsing. Linear operators preserving unitary similarity invariant norms. Linear and Multilinear Algebra, 27(3):213–224, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1991:TMN

[LT91] Chi-Kwong Li and Nam-Kiu Tsing. On the k th matrix numerical range. Linear and Multilinear Algebra, 28(4):229–239, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei:1992:PCN

[LT92a] Tian-Gang Lei and Tin-Yau Tam. On the permanental congruence numerical range. Linear and Multilinear Algebra, 32 (1):51–59, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1992:CLPa

[LT92b] Chi-Kwong Li and Nam-Kiu Tsing. Chapter 5: linear preservers on functions of singular values. Linear and Multilinear Algebra, 33(1–2):53–61, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1992:CLPb

[LT92c] Chi-Kwong Li and Nam-Kiu Tsing. Chapter 6: linear preservers on numerical ranges, numerical radii and unitary similarity invariant norms. Linear and Multilinear Algebra, 33(1– 2):63–73, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1992:CLPc

[LT92d] Chi-Kwong Li and Nam-Kiu Tsing. Chapter 7: linear preservers of relations on matrix spaces. Linear and Multilinear Algebra, 33(1–2):75–84, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1993:LOL

[LT93] Chi-Kwong Li and Nam-Kiu Tsing. Linear operators leaving a class of matrices with fixed singular values invariant. Linear and Multilinear Algebra, 34(1):41–49, 1993. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2004:E

[LT04] C. K. Li and T. Y. Tam. Editorial. Linear Multilinear Algebra, 52(3–4):157, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lourenco:2007:ZCH

[LT07] Mary Lillian Lourenço and Neusa Nogas Tocha. Zeros of complex homogeneous polynomials. Linear Multilinear Algebra, 55(5):463–469, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2009:PSI

[LT09] C.-K. Li and T.-Y. Tam. Preface to the special issue on the numerical range and numerical radius. Linear Multilinear Algebra, 57(5):429, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2010:HUR

[LT10] Yonghui Liu and Yongge Tian. How to use RSVD to solve the matrix equation A = BXC. Linear Multilinear Algebra, 58(5):

Li:2011:SRW

[LT11] Shuchao Li and Yi Tian. On the spectral radius of weighted unicyclic graphs with a positive weight set. Linear Multilinear Algebra, 59(12):1399–1407, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2012:SRW

[LT12a] Shuchao Li and Yi Tian. On the spectral radius of weighted trees with given number of pendant vertices and a positive weight set. Linear Multilinear Algebra, 60(8):955–965, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2012:DSM

[LT12b] Zongbing Lin and Qianrong Tan. Determinants of Smith matrices on three coprime divisor chains and divisibility. Linear Multilinear Algebra, 60(4):475–486, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2013:GFM

[LT13] Xuhua Liu and Tin-Yau Tam. Gradient flows for the minimum distance to the sum of adjoint orbits. Linear Multilinear Algebra, 61(5):678–696, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lim:2014:SPT

[LT14] Ming-Huat Lim and Sin-Chee Tan. Semimodules and preservers of traceless symmetric Boolean matrices of factor rank two. Linear Multilinear Algebra, 62(2):159–179, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:DAC

[LT22a] Cheng-Kai Liu and Yuan-Tsung Tsai. Derivations with annihilator conditions on multilinear polynomials. Linear Multilinear Algebra, 70(12):2397–2413, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lyubimtsev:2022:LCE

[LT22b] O. V. Lyubimtsev and A. A. Tuganbaev. Local centrally essential subalgebras of triangular algebras. Linear Multilinear

Li:1984:GSR

[LTT84] Chi-Kwong Li, Tin-Yau Tam, and Nam-Kiu Tsing. The generalized spectral radius, numerical radius and spectral norm. Linear and Multilinear Algebra, 16(1–4):215–237, 1984. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2011:RAS

[LTT11] Xuhua Liu, Dawit G. Tadesse, and Tin-Yau Tam. A range associated with skew symmetric matrix. Linear Multilinear Algebra, 59(10):1105–1120, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lu:2022:SGS

[Lu22] Yong Lu. Some generalizations of spectral conditions for 2shamiltonicity and 2s-traceability of bipartite graphs. Linear Multilinear Algebra, 70(10):1907–1927, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ludkowski:2019:ADN

[Lud19] S. V. Ludkowski. Automorphisms and derivations of nonassociative C∗-algebras. Linear Multilinear Algebra, 67(8):1531– 1538, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ludkowski:2021:CHI

[Lud21a] S. V. Ludkowski. Completions and homomorphisms of infinite dimensional Cayley–Dickson algebras. Linear Multilinear Algebra, 69(11):2040–2049, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ludkowski:2021:WSA

[Lud21b] S. V. Ludkowski. Weakly separable algebras. Linear Multilinear Algebra, 69(7):1236–1247, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Luks:1974:NLL

[Luk74] Eugene M. Luks. Normalizers of linear Lie algebras. Linear and Multilinear Algebra, 2(2):151–160, 1974. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Luor:2019:SNE

[Luo19] Dah-Chin Luor. Sharp norm estimates for a class of integral operators. Linear Multilinear Algebra, 67(12):2504–2519, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Luque:2008:HS

[Luq08] Jean-Gabriel Luque. Hyperdeterminants on semilattices. Linear Multilinear Algebra, 56(3):333–344, 2008. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Langberg:2018:CCG

[LV18] Michael Langberg and Dan Vilenchik. Constructing cospectral graphs via a new form of graph product. Linear Multilinear Algebra, 66(9):1838–1852, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Leipnik:1978:BR

[LW78] R. B. Leipnik and Robert Lee Wilson. Book reviews. Linear and Multilinear Algebra, 6(4):331–338, 1978. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:1996:RP

[LW96] Chi-Kwong Li and Hugo J. Woerdeman. A research problem. Linear and Multilinear Algebra, 41(1):41–47, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2008:ROL

[LW08] Qiaohua Liu and Musheng Wei. Reverse order law for least squares g-inverses of multiple matrix products. Linear Multilinear Algebra, 56(5):491–506, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2010:DEM

[LW10] Qiaohua Liu and Musheng Wei. On direct elimination methods for solving the equality constrained least squares problem. Linear Multilinear Algebra, 58(2):173–184, 2010. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2011:GLT

[LW11a] Hailing Li and Ying Wang. Generalized Lie triple derivations. Linear Multilinear Algebra, 59(3):237–247, 2011. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2011:RHM

[LW11b] Minghua Lin and Qing-Wen Wang. Remarks on Hua’s matrix equality involving generalized inverses. Linear Multilinear Algebra, 59(9):1059–1067, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2012:EMI

[LW12] Minghua Lin and Henry Wolkowicz. An eigenvalue majorization inequality for positive semidefinite block matrices. Linear Multilinear Algebra, 60(11–12):1365–1368, 2012. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03081087.2011.651723.

Li:2017:ICT

[LW17] Cui-Xia Li and Shi-Liang Wu. Improved convergence theorems for new Hermitian and skew-Hermitian splitting methods. Linear Multilinear Algebra, 65(3):510–513, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2020:RBO

[LW20] Linlin Liu and Shuanhong Wang. Rota–Baxter H-operators and pre-Lie H-pseudoalgebras over a cocommutative Hopf algebra H. Linear Multilinear Algebra, 68(11):2170–2184, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:DSR

[LW22a] Xiangxiang Liu and Ligong Wang. Distance spectral radii of k-uniform bicyclic hypergraphs. Linear Multilinear Algebra, 70(21):6190–6210, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:GDS

[LW22b] Xiaogang Liu and Tingzeng Wu. Graphs determined by the (signless) Laplacian permanental polynomials. Linear Multilinear Algebra, 70(18):3599–3615, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2023:OMB

[LW23] Xiaogang Liu and Tingzeng Wu. One-matching bi-Cayley graphs and homogeneous bi-Cayley graphs over finite cyclic groups. Linear Multilinear Algebra, 71(1):88–99, 2023. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2011:LCG

[LWB11a] Xiaoji Liu, Lingling Wu, and Julio Benítez. On linear combinations of generalized involutive matrices. Linear Multilinear Algebra, 59(11):1221–1236, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2011:NCT

[LWB11b] Xiaoji Liu, Shuxia Wu, and Julio Benítez. On nonsingularity of combinations of two group invertible matrices and two tripotent matrices. Linear Multilinear Algebra, 59(12):1409–1417, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:GQN

[LWC22] Rongfang Li, Deyu Wu, and Alatancang Chen. On the geometry of the quadratic numerical range for 2 × 2 block operator matrices. Linear Multilinear Algebra, 70(7):1315–1328, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2021:DSR

[LWL21] Xiangxiang Liu, Ligong Wang, and Xihe Li. Distance spectral radii of k-uniform hypertrees with given parameters. Linear Multilinear Algebra, 69(14):2558–2571, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2014:EON

[lWpW14] Hui li Wang and Xue ping Wang. Eigenproblem for optimalnode matrices in max-plus algebra. Linear Multilinear Algebra, 62(8):1105–1113, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2015:CIM

[lWW15] Hui li Wang and Xue-Ping Wang. The characterizations of irreducible matrices with proper supereigenvectors. Linear Mul-

Liang:2015:CTL

[LWXF15] Xinfeng Liang, Feng Wei, Zhankui Xiao, and Ajda Fošner. Centralizing traces and Lie triple isomorphisms on generalized matrix algebras. Linear Multilinear Algebra, 63(9):1786–1816, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2011:GIC

[LWY11] Xiaoji Liu, Lingling Wu, and Yaoming Yu. The group inverse of the combinations of two idempotent matrices. Linear Multilinear Algebra, 59(1):101–115, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2019:NPU

[lWY19] Hui li Wang and Yan Yang. A note on the paper ‘Ultra discrete permanent and the consistency of max plus linear equations’. Linear Multilinear Algebra, 67(6):1103–1108, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:2015:SJD

[LWZ15] Tsiu-Kwen Lee, Tsai-Lien Wong, and Yiqiang Zhou. The structure of Jordan *-derivations of prime rings. Linear Multilinear Algebra, 63(2):411–422, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2017:SLD

[LXG17] Ruifang Liu, Jie Xue, and Litao Guo. On the second largest distance eigenvalue of a graph. Linear Multilinear Algebra, 65(5):1011–1021, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lei-Xiao-Qiang:2021:ISG

[LXQ21] Léi-Xiăo-Qiáng. α_{β}Ω-inclusion sets for general C-eigenvalues of a general piezoelectric-type tensor. Linear Multilinear Algebra, 69(16):3043–3068, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2021:SG

[LXS21] Huiqiu Lin, Jie Xue, and Jinlong Shu. On the D_{α}-spectra of graphs. Linear Multilinear Algebra, 69(6):997–1019, 2021.

Lam:2022:SIA

[LY22a] Ching Hung Lam and Hiromichi Yamada. Sigma involutions associated with parafermion vertex operator algebra K(∫↕_{∈}, ∥). Linear Multilinear Algebra, 70(21):6780–6819, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:RBA

[LY22b] Shuchao Li and Ting Yang. On the relation between the adjacency rank of a complex unit gain graph and the matching number of its underlying graph. Linear Multilinear Algebra, 70(9):1768–1787, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2022:EEM

[LY22c] Shuchao Li and Yuantian Yu. The effect on A_{α}-eigenvalues of mixed graphs and unit gain graphs by adding edges in clusters. Linear Multilinear Algebra, 70(20):5732–5749, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2022:NDS

[LY22d] Yan Liu and Jin Yan. Note on distance signless Laplacian spectral radius under given matching number. Linear Multilinear Algebra, 70(6):1185–1190, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2015:PAS

[LYS15] Hanyu Li, Hu Yang, and Hua Shao. Perturbation analysis for the symplectic QR factorization. Linear Multilinear Algebra, 63(1):78–96, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2012:ILR

[LYZW12] Xiaoji Liu, Yaoming Yu, Jin Zhong, and Yimin Wei. Integral and limit representations of the outer inverse in Banach space. Linear Multilinear Algebra, 60(3):333–347, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2001:LOPb

[LZ01] Chi-Kwong Li and Alexandru Zaharia. Linear operators preserving decomposable numerical radii on orthonormal tensors. Linear and Multilinear Algebra, 49(1):25–43, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lee:2009:ITS

[LZ09] Tsiu-Kwen Lee and Yiqiang Zhou. On irreducible and transitive subalgebras in matrix algebras. Linear Multilinear Algebra, 57(7):659–672, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2011:CJD

[LZ11a] Jiankui Li and Jiren Zhou. Characterizations of Jordan derivations and Jordan homomorphisms. Linear Multilinear Algebra, 59(2):193–204, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2011:PBS

[LZ11b] Shuchao Li and Li Zhang. Permanental bounds for the signless Laplacian matrix of bipartite graphs and unicyclic graphs. Linear Multilinear Algebra, 59(2):145–158, 2011. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2012:ELA

[LZ12] Wenli Liu and Zhixue Zhang. T^{∗}-extension of a 3-Lie algebra. Linear Multilinear Algebra, 60(5):583–594, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2013:ELA

[LZ13] Wenli Liu and Zhixue Zhang. T^{∗}-extension of n-Lie algebras. Linear Multilinear Algebra, 61(4):527–542, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2014:CPP

[LZ14a] Shunyi Liu and Heping Zhang. Characterizing properties of permanental polynomials of lollipop graphs. Linear Multilinear Algebra, 62(4):419–444, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2014:SNC

[LZ14b] Xiaogang Liu and Sanming Zhou. Spectra of the neighbourhood corona of two graphs. Linear Multilinear Algebra, 62(9): 1205–1219, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lan:2015:SGO

[LZ15a] Jie Lan and Bo Zhou. Spectra of graph operations based on R-graph. Linear Multilinear Algebra, 63(7):1401–1422, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2015:GMO

[LZ15b] Hongying Lin and Bo Zhou. Graphs with at most one signless Laplacian eigenvalue exceeding three. Linear Multilinear Algebra, 63(2):377–383, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2016:PMB

[LZ16] Wen Li and Hua Zheng. A preconditioned modulus-based iteration method for solving linear complementarity problems of H-matrices. Linear Multilinear Algebra, 64(7):1390–1403, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2017:SBS

[LZ17a] Hongying Lin and Bo Zhou. On sharp bounds for spectral radius of nonnegative matrices. Linear Multilinear Algebra, 65(8):1554–1565, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2017:EHT

[LZ17b] Minghua Lin and Fuzhen Zhang. An extension of Harnack type determinantal inequality. Linear Multilinear Algebra, 65 (10):2024–2030, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2019:SAL

[LZ19a] Li Liang and Baodong Zheng. Sensitivity analysis of the Lyapunov tensor equation. Linear Multilinear Algebra, 67(3):555– 572, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liao:2019:GMS

[LZ19b] Li-Dan Liao and Guo-Feng Zhang. The generalized C-to-R method for solving complex symmetric indefinite linear systems. Linear Multilinear Algebra, 67(9):1727–1735, 2019. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2019:DSR

[LZ19c] Hongying Lin and Bo Zhou. The distance spectral radius of trees. Linear Multilinear Algebra, 67(2):370–390, 2019. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lv:2022:PAF

[LZ22] Peng Lv and Bing Zheng. Perturbation analysis for the QX factorization for centrosymmetric matrices. Linear Multilinear Algebra, 70(3):557–580, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2023:ALW

[LZ23] Tingting Li and Mengmeng Zhou. The absorption laws for the weighted core inverse in rings. Linear Multilinear Algebra, 71(3):480–495, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liu:2023:HOS

[LZB23] Chunmeng Liu, Jiang Zhou, and Changjiang Bu. The high order spectrum of a graph and its applications in graph colouring and clique counting. Linear Multilinear Algebra, 71(14): 2354–2365, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2016:NBT

[LZL16] Chaoqian Li, Jianjun Zhou, and Yaotang Li. A new Brauertype eigenvalue localization set for tensors. Linear Multilinear Algebra, 64(4):727–736, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lin:2018:DSR

[LZL18] Hongying Lin, Bo Zhou, and Yaduan Li. On distance spectral radius of uniform hypergraphs. Linear Multilinear Alge-

Liu:2023:CNM

[LZS23] Qiaohua Liu, Qian Zhang, and Dongmei Shen. Condition numbers of the mixed least squares-total least squares problem revisited. Linear Multilinear Algebra, 71(13):2144–2162, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Liang:2018:ALS

[LZT18] Li Liang, Baodong Zheng, and Yunbo Tian. Algebraic Lyapunov and Stein stability results for tensors. Linear Multilinear Algebra, 66(4):731–741, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2019:RBR

[LZX19] Shuchao Li, Siqi Zhang, and Baogen Xu. The relation between the H-rank of a mixed graph and the independence number of its underlying graph. Linear Multilinear Algebra, 67(11): 2230–2245, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Li:2017:NVS

[LZZ17] Zhongshan Li, Fuzhen Zhang, and Xiao-Dong Zhang. On the number of vertices of the stochastic tensor polytope. Linear Multilinear Algebra, 65(10):2064–2075, 2017. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1979:PEP

[– DM79] Dragomir Ž. – Doković and Jerry Malzan. Perturbation of eigenvalues in U(n) by pseudo-reflections. Linear and Multilinear Algebra, 8(1):35–40, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1975:CMD

[MA75] Marvin Marcus and M. Shafqat Ali. Congruence maps and derivations. Linear and Multilinear Algebra, 3(1–2):115–133, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2015:ISP

[Ma15a] Chao Ma. Idempotent sign patterns of special types. Linear Multilinear Algebra, 63(5):1002–1008, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mau:2015:CCL

[Ma’15b] Sione Ma’u. Chebyshev constants, linear algebra and computation on algebraic curves. Linear Multilinear Algebra, 63(10): 2043–2060, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mourad:2015:SDS

[MA15c] Bassam Mourad and Hassan Abbas. On the symmetric doubly stochastic matrices that are determined by their spectra and their connection with spectral graph theory. Linear Multilinear Algebra, 63(5):869–881, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2022:CRC

[Ma22a] Haifeng Ma. Characterizations and representations for the CMP inverse and its application. Linear Multilinear Algebra, 70(20):5157–5172, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2022:DSC

[Ma22b] Haifeng Ma. Displacement structure of the core inverse. Linear Multilinear Algebra, 70(2):203–214, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Machado:1995:M

[Mac95] Joaquim F. Machado. The maximum of i(XAX^{−1} + B). Linear and Multilinear Algebra, 40(1):33–35, 1995. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Machado:2009:CEB

[Mac09] Armando Machado. Complex eigenvalues before the fundamental theorem of algebra. Linear Multilinear Algebra, 57(6): 561–566, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Magnus:1983:SML

[Mag83] Jan R. Magnus. L-structured matrices and linear matrix equations. Linear and Multilinear Algebra, 14(1):67–88, 1983. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Magajna:2013:BRD

[Mag13] Bojan Magajna. Bicommutants and ranges of derivations. Linear Multilinear Algebra, 61(9):1161–1180, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Magajna:2014:BRD

[Mag14] Bojan Magajna. Bicommutants and ranges of derivations. Linear Multilinear Algebra, 62(9):1272–1273, 2014. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Magalhaes:2022:SPE

[Mag22] Luis T. Magalhães. Simple proof of existence of a complex eigenvalue of a complex square matrix … and yet another proof of the fundamental theorem of algebra with linear algebra. Linear Multilinear Algebra, 70(20):5329–5333, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mahoney:1982:DIG

[Mah82] Michael K. Mahoney. Determinants of integral group matrices for some nonabelian 2-generator groups. Linear and Multilinear Algebra, 11(2):189–201, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Motlaghian:2018:TPO

[MAH18] Sara M. Motlaghian, Ali Armandnejad, and Frank J. Hall. Topological properties of J -orthogonal matrices. Linear Multilinear Algebra, 66(12):2524–2533, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Motlaghian:2021:TPO

[MAH21] Sara M. Motlaghian, Ali Armandnejad, and Frank J. Hall. Topological properties of J-orthogonal matrices, part II. Linear Multilinear Algebra, 69(3):438–447, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maiti:2011:CPS

[Mai11] Saran Ishika Maiti. Construction, properties and statistical applications of positive definite intraclass matrix. Linear Multilinear Algebra, 59(8):891–904, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Majidinya:2022:ALF

[Maj22] Ali Majidinya. Automorphisms of linear functional graphs over vector spaces. Linear Multilinear Algebra, 70(20):5681–5697, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malek:1986:M

[Mal86] Massoud Malek. On the maximum of per(I − A). Linear and Multilinear Algebra, 19(4):347–355, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malek:1989:NPC

[Mal89] Massoud Malek. A note on a permanental conjecture of M. Marcus and H. Minc. Linear and Multilinear Algebra, 25(1): 71–73, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malek:1991:MSS

[Mal91] Massoud Malek. On the monotonicity of the sum of subpermanents of doubly stochastic matrices. Linear and Multilinear Algebra, 29(3–4):291–297, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maliakas:1992:NSL

[Mal92] Mihalis Maliakas. A note on the straightening law for the symplectic group. Linear and Multilinear Algebra, 33(3–4): 131–136, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malek:1994:HM

[Mal94] Massoud Malek. Hypernormal matrices. Linear and Multilinear Algebra, 36(4):255–266, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maliakas:1995:IMT

[Mal95] Mihalis Maliakas. On induced modules and tableaux for the special orthogonal groups. Linear and Multilinear Algebra, 38(4):283–294, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mourad:2015:NIS

[MAM15] Bassam Mourad, Hassan Abbas, and Mohammad Sal Moslehian. A note on the inverse spectral problem for symmetric doubly stochastic matrices. Linear Multilinear Algebra, 63 (12):2537–2545, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Man:1987:CGN

[Man87] Wai-Yip Man. The convexity of the generalized numerical range. Linear and Multilinear Algebra, 20(3):229–245, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Man:1991:INR

[Man91] Wai Yip Man. The invariance of C-numerical range Cnumerical radius, and their dual problems. Linear and Multilinear Algebra, 30(1–2):117–128, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Man:1992:NRN

[Man92] Wai Ypi Man. C-numerical ranges and normality of matrices. Linear and Multilinear Algebra, 32(3–4):237–247, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Manjegani:2017:YHI

[Man17] S. Mahmoud Manjegani. On Young and Heinz inequalities for faithful tracial states. Linear Multilinear Algebra, 65(12): 2432–2456, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1975:ANR

[Mar75] Marvin Marcus. Adjoints and the numerical range. Linear and Multilinear Algebra, 3(1–2):81–89, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1976:CFI

[Mar76] Marvin Marcus. Converses of the Fischer inequality and an inequality of A. Ostrowski. Linear and Multilinear Algebra, 4(2): 137–148, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1978:DSTb

[Mar78] Marvin Marcus. Decomposable symmetrized tensors and an extended LR decomposition theorem. Linear and Multilinear Algebra, 6(4):327–330, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1982:EN

[Mar82a] Marvin Marcus. Editorial note. Linear and Multilinear Algebra, 11(2):105–106, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1982:RPP

[Mar82b] Marvin Marcus. A remark on the preceding paper. Linear and Multilinear Algebra, 11(3):287, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Cho82].

Marcus:1984:EG

[Mar84] Marvin Marcus. An exponential group. Linear and Multilinear Algebra, 16(1–4):295–296, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1989:UES

[Mar89] Marvin Marcus. A unified exposition of some classical matrix theorems. Linear and Multilinear Algebra, 25(2):137–147, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

MarquesDeSa:1990:IDC

[Mar90] Eduardo Marques De Sá. Interlacing and degree conditions for invariant polynomials. Linear and Multilinear Algebra, 27(4): 303–316, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

MarquesDeSa:1992:SNO

[Mar92] Eduardo Marques De Sá. Some notes on orthant-monotonic norms. Linear and Multilinear Algebra, 32(2):167–175, 1992.

Marcus:1997:LBP

[Mar97] Marvin Marcus. A lower bound for the product of linear forms. Linear and Multilinear Algebra, 43(1–3):115–120, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcuson:2000:SSK

[Mar00] Richard Marcuson. On the simplicity of some Kac–Moody groups. Linear and Multilinear Algebra, 47(1):49–55, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:2003:HMB

[Mar03] Marvin Marcus. Henryk Minc — a biography. Linear Multilinear Algebra, 51(1):1–10, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martin:2011:RT

[Mar11] Carla D. Martin. The rank of a 2 × 2× 2 tensor. Linear Multilinear Algebra, 59(8):943–950, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martin:2012:SSN

[Mar12] Antonio J. Calderón Martín. On the structure of split noncommutative Poisson algebras. Linear Multilinear Algebra, 60(7):775–785, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marciniak:2013:RPE

[Mar13] Marcin Marciniak. Rank properties of exposed positive maps. Linear Multilinear Algebra, 61(7):970–975, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marovt:2015:POR

[Mar15] Janko Marovt. On partial orders in Rickart rings. Linear Multilinear Algebra, 63(9):1707–1723, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marovt:2018:ORB

[Mar18] Janko Marovt. Orders in rings based on the core-nilpotent decomposition. Linear Multilinear Algebra, 66(4):803–820, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marrero:2020:EIF

[Mar20] Jesús Abderramán Marrero. Explicit inversion formulas for properly Hessenberg and Toeplitz–Hessenberg matrices. Linear Multilinear Algebra, 68(6):1239–1256, 2020. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Massey:2010:NCS

[Mas10] Pedro G. Massey. Non-commutative Schur–Horn theorems and extended majorization for Hermitian matrices. Linear Multilinear Algebra, 58(4):465–480, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mathias:1990:CMM

[Mat90] Roy Mathias. Concavity of monotone matrix functions of finite order. Linear and Multilinear Algebra, 27(2):129–138, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mathias:1992:SVH

[Mat92] Roy Mathias. The singular values of the Hadamard product of a positive semidefinite and a skew-symmetric matrix. Linear and Multilinear Algebra, 31(1–4):57–70, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mattarei:2011:ECS

[Mat11] Sandro Mattarei. Engel conditions and symmetric tensors. Linear Multilinear Algebra, 59(4):441–449, 2011. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moosavianfard:2022:IGF

[MAT22] Z. Moosavianfard, M. Abolghasemi, and Y. Tolooei. Invertibility of generalized g-frame multipliers in Hilbert spaces. Linear Multilinear Algebra, 70(8):1590–1609, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maurer:1978:ITF

[Mau78] Donald Maurer. Invariants of the trace-form of a number field. Linear and Multilinear Algebra, 6(1):33–36, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Miers:1976:BR

[MB76] C. R. Miers and George Phillip Barker. Book reviews. Linear and Multilinear Algebra, 4(3):229–232, 1976. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mondal:2022:CQP

[MB22a] Chiranjit Mondal and R. Balaji. Characterization of qproperty for cone automorphisms in second-order cone linear complementarity problems. Linear Multilinear Algebra, 70(21):6155–6175, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mwafise:2022:ERA

[MB22b] Arnauld Mesinga Mwafise and Paul Barry. Exponential Riordan arrays and Jacobi elliptic functions. Linear Multilinear Algebra, 70(20):5770–5789, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moulaharabbi:2022:NRQ

[MBB22] Somayya Moulaharabbi, Mohamed Barraa, and El Hassan Benabdi. Numerical range of quaternionic right linear bounded operators. Linear Multilinear Algebra, 70(22):7599– 7610, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moyls:1974:BR

[MBSL74] B. N. Moyls, Robert E. Beck, Hans Schneider, and Roy B. Letpnik. Book reviews. Linear and Multilinear Algebra, 2(3): 293–298, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1978:DSTa

[MC78] Marvin Marcus and John Chollet. Decomposable symmetrized tensors. Linear and Multilinear Algebra, 6(4):317–326, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1980:LGD

[MC80] Marvin Marcus and John Chollet. Linear groups defined by decomposable tensor equalities. Linear and Multilinear Algebra, 8(3):207–212, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1982:ISC

[MC82] Marvin Marcus and John Chollet. The index of a symmetry class of tensors. Linear and Multilinear Algebra, 11(3):277– 281, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1983:EDS

[MC83] Marvin Marcus and John Chollet. On the equality of decomposable symmetrized tensors. Linear and Multilinear Algebra, 13(3):253–266, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1986:COB

[MC86] Marvin Marcus and John Chollet. Construction of orthonormal bases in higher symmetry classes of tensors. Linear and Multilinear Algebra, 19(2):133–140, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1989:LBN

[MC89] Marvin Marcus and John Chollet. Lower bounds for the norms of decomposable symmetrized tensors. Linear and Multilinear Algebra, 25(4):269–274, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2017:JLT

[MC17] Lili Ma and Liangyun Chen. On δ-jordan Lie triple systems. Linear Multilinear Algebra, 65(4):731–751, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Masque:2019:NLC

[MC19] J. Muñoz Masqué and L. M. Pozo Coronado. A new look at the classification of the tri-covectors of a 6-dimensional symplectic space. Linear Multilinear Algebra, 67(5):939–952, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McCloskey:1985:PNP

[McC85] Joseph P. McCloskey. Properties of normal r-potent matrices. Linear and Multilinear Algebra, 17(3–4):325–336, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McDonald:1982:CD

[McD82] B. R. McDonald. A characterization of the determinant. Linear and Multilinear Algebra, 12(1):31–36, 1982. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McDonald:1992:CIP

[McD92a] Bernard R. McDonald. Chapter 8: invariant preserving linear matrix mappings over rings and related topics. Linear and Multilinear Algebra, 33(1–2):85–108, 1992. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McDonald:1992:PZE

[McD92b] Judith J. McDonald. Partly zero eigenvectors and matrices that satisfy Au = b. Linear and Multilinear Algebra, 33(3–4): 163–170, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mcfall:1979:EDT

[Mcf79a] J. D. Mcfall. Elementary divisors of transformations related to tensor powers. Linear and Multilinear Algebra, 7(1):13–25, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McFall:1979:HCE

[McF79b] John D. McFall. How to compute the elementary divisors of the tensor product of two matrices. Linear and Multilinear Algebra, 7(3):193–201, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McDaniel:2021:FEL

[MCIM21] Chris McDaniel, Shujian Chen, Anthony Iarrobino, and Pedro Macias Marques. Free extensions and Lefschetz properties, with an application to rings of relative coinvariants. Linear Multilinear Algebra, 69(2):305–330, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mahmoud:2020:QIO

[MCL20] Sid Ahmed Ould Ahmed Mahmoud, Muneo Cho, and Ji Eun Lee. On n-quasi-(m, C)-isometric operators. Linear Multilinear Algebra, 68(5):1001–1020, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mal:2022:CSO

[MDP22] Arpita Mal, Subhrajit Dey, and Kallol Paul. Characterization of k-smoothness of operators defined between infinitedimensional spaces. Linear Multilinear Algebra, 70(18):3477– 3489, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2011:SZD

[MDW11] Xiaobin Ma, Genhong Ding, and Long Wang. Square-zero determined matrix algebras. Linear Multilinear Algebra, 59 (11):1311–1317, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mukhamedov:2018:NLM

[ME18a] Farrukh Mukhamedov and Ahmad Fadillah Embong. On nonlinear Markov operators: surjectivity vs orthogonal preserving property. Linear Multilinear Algebra, 66(11):2183–2190, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mukhamedov:2018:SBQ

[ME18b] Farrukh Mukhamedov and Ahmad Fadillah Embong. On stable b-bistochastic quadratic stochastic operators and associated non-homogeneous Markov chains. Linear Multilinear Algebra, 66(1):1–21, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mukhamedov:2021:IDO

[ME21] Farrukh Mukhamedov and Ahmad Fadillah Embong. Infinite dimensional orthogonality preserving nonlinear Markov operators. Linear Multilinear Algebra, 69(3):526–550, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meenakshi:1980:MEA

[Mee80] Ar. Meenakshi. On EP_{r} matrices with entries from an arbitrary field. Linear and Multilinear Algebra, 9(2):159–164, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mehrmann:1984:SCS

[Meh84] Volker Mehrmann. On some conjectures on the spectra of τ-matrices. Linear and Multilinear Algebra, 16(1–4):101–112, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mejjaoli:2019:WMA

[Mej19] Hatem Mejjaoli. Wavelet-multipliers analysis in the framework of the k-Laguerre theory. Linear Multilinear Algebra, 67(1): 70–93, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Melman:2010:ABS

[Mel10] A. Melman. An alternative to the Brauer set. Linear Multilinear Algebra, 58(3):377–385, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Melman:2012:OCT

[Mel12] A. Melman. Ovals of Cassini for Toeplitz matrices. Linear Multilinear Algebra, 60(2):189–199, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Melman:2013:SOC

[Mel13a] A. Melman. A single oval of Cassini for the zeros of a polynomial. Linear Multilinear Algebra, 61(2):183–195, 2013. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Melman:2013:ULB

[Mel13b] A. Melman. Upper and lower bounds for the Perron root of a nonnegative matrix. Linear Multilinear Algebra, 61(2): 171–181, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Melman:2018:APP

[Mel18] A. Melman. An alternative proof of Pellet’s theorem for matrix polynomials. Linear Multilinear Algebra, 66(4):785–791, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Melman:2019:PEB

[Mel19] A. Melman. Polynomial eigenvalue bounds from companion matrix polynomials. Linear Multilinear Algebra, 67(3):598– 612, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Melman:2020:RPR

[Mel20] A. Melman. Refinement of Pellet radii for matrix polynomials. Linear Multilinear Algebra, 68(11):2185–2200, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Melman:2022:PEE

[Mel22] Aaron Melman. Polynomial eigenvalue estimation: numerical radii versus norms. Linear Multilinear Algebra, 70(22):7656– 7671, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meng:2022:NAP

[Men22] Lingsheng Meng. New additive perturbation bounds of the Moore–Penrose inverse. Linear Multilinear Algebra, 70(6): 1096–1101, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1974:IIG

[Mer74] Russell Merris. An identity involving generalized matrix functions. Linear and Multilinear Algebra, 2(2):123–125, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1975:I

[Mer75] Russell Merris. The irreducibility of K(A). Linear and Multilinear Algebra, 2(4):299–303, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1977:VDS

[Mer77] Russell Merris. On vanishing decomposable symmetrized tensors. Linear and Multilinear Algebra, 5(2):79–86, 1977. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merikoski:1978:SNH

[Mer78a] Jorma Kaarlo Merikoski. On the submultiplicativity of norms of Hölder and Minkowski type. Linear and Multilinear Algebra, 6(1):51–54, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1978:LIP

[Mer78b] Russell Merris. Linearly independent permutation characters afforded by action on tensor spaces. Linear and Multilinear Algebra, 6(4):263–267, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1978:SHD

[Mer78c] Russell Merris. The structure of higher degree symmetry classes of tensors. II. Linear and Multilinear Algebra, 6(3): 171–178, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1979:GMF

[Mer79a] Russell Merris. Generalized matrix functions: a research problem. Linear and Multilinear Algebra, 8(1):83–86, 1979. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1979:PPK

[Mer79b] Russell Merris. On the pseudocharacteristic polynomial of K. Bogart and J. Gordon. Linear and Multilinear Algebra, 7(3): 237–239, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1979:RAS

[Mer79c] Russell Merris. Recent advances in symmetry classes of tensors. Linear and Multilinear Algebra, 7(4):317–328, 1979. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merriell:1980:MP

[Mer80] David Merriell. The maximum permanent in. Linear and Multilinear Algebra, 9(2):81–91, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merikoski:1981:LBH

[Mer81] Jorma Kaarlo Merikoski. Lower bounds for ||A||_{2} by Hölder norms. Linear and Multilinear Algebra, 9(4):341–344, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1982:RGM

[Mer82] Russell Merris. Representations of GL(n, R) and generalized matrix functions of class MPW. Linear and Multilinear Algebra, 11(2):133–141, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1983:GMF

[Mer83a] Russell Merris. Generalized matrix functions and pattern inventory. Linear and Multilinear Algebra, 12(4):315–327, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1983:SHC

[Mer83b] Russell Merris. Single-hook characters and Hamiltonian circuits. Linear and Multilinear Algebra, 14(1):21–35, 1983. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1989:EVM

[Mer89] Russell Merris. An edge version of the matrix-tree theorem and the Wiener index. Linear and Multilinear Algebra, 25(4): 291–296, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1991:RP

[Mer91] Russell Merris. Research problems. Linear and Multilinear Algebra, 29(3–4):315–317, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1992:AMA

[Mer92] Russell Merris. Applications of multlinear algebra. Linear and Multilinear Algebra, 32(3–4):211–224, 1992. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1994:IES

[Mer94] Russell Merris. An inequality for eigenvalues of symmetric matrices with applications to max-cuts and graph energy. Linear and Multilinear Algebra, 36(3):225–229, 1994. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1995:SGL

[Mer95] Russell Merris. A survey of graph Laplacians. Linear and Multilinear Algebra, 39(1–2):19–31, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1997:LFL

[Mer97] Russell Merris. Large families of Laplacian isospectral graphs. Linear and Multilinear Algebra, 43(1–3):201–205, 1997. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1998:BR

[Mer98a] Russell Merris. Book review. Linear and Multilinear Algebra, 45(2–3):287–289, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1998:DSG

[Mer98b] Russell Merris. Doubly stochastic graph matrices, II. Linear and Multilinear Algebra, 45(2–3):275–285, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:2017:MMM

[Mer17] Russell Merris. Marvin Marcus — the man and his legacy. Linear Multilinear Algebra, 65(10):1927–1933, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meshulam:1990:SRM

[Mes90] Roy Meshulam. On k-spaces of real matrices. Linear and Multilinear Algebra, 26(1–2):39–41, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mesyan:2013:PSD

[Mes13] Zachary Mesyan. Polynomials of small degree evaluated on matrices. Linear Multilinear Algebra, 61(11):1487–1495, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meyer:2015:CPR

[Mey15] Carl D. Meyer. Continuity of the Perron root. Linear Multilinear Algebra, 63(7):1332–1336, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1977:INV

[MF77] Marvin Marcus and Ivan Filippenko. Invariance of the nonvanishing specializations of polynomials. Linear and Multilinear Algebra, 5(2):99–105, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1979:LOP

[MF79] Marvin Marcus and Ivan Filippenko. Linear operators preserving the decomposable numerical range. Linear and Multilinear Algebra, 7(1):27–36, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1992:RIC

[MF92] Marvin Marcus and Susan Franklin. Rank and inertia of a complex Hadamard power. Linear and Multilinear Algebra, 32(2):149–166, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2020:ILB

[MF20a] Xiaobin Ma and Xianwen Fang. An improved lower bound for the nullity of a graph in terms of matching number. Linear Multilinear Algebra, 68(10):1983–1989, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moradi:2020:CII

[MF20b] Hamid Reza Moradi and Shigeru Furuichi. A complementary inequality to the information monotonicity for Tsallis relative operator entropy. Linear Multilinear Algebra, 68(8): 1557–1567, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Munoz-Fernandez:2012:EAB

[MFPSS12] G. A. Muñoz-Fernández, D. Pellegrino, and J. B. Seoane-Sepúlveda. Estimates for the asymptotic behaviour of the constants in the Bohnenblust–Hille inequality. Linear Multilinear Algebra, 60(5):573–582, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moradi:2022:SOI

[MFS22] Hamid Reza Moradi, Shigeru Furuichi, and Mohammad Sababheh. Some operator inequalities via convexity. Linear Multilinear Algebra, 70(22):7740–7752, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martin:2022:FDA

[MG22] Antonio J. Calderón Martín and Babacar Gaye. Fine decompositions of algebraic systems induced by bases. Linear Multilinear Algebra, 70(14):2804–2817, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mehranian:2017:SPG

[MGA17] Z. Mehranian, A. Gholami, and A. R. Ashrafi. The spectra of power graphs of certain finite groups. Linear Multilinear Algebra, 65(5):1003–1010, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martinez:2019:SHP

[MGAH19] Noé Martínez, Luis E. Garza, and Baltazar Aguirre-Hernández. On sequences of Hurwitz polynomials related to orthogonal polynomials. Linear Multilinear Algebra, 67(11): 2191–2208, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moradi:2019:GOK

[MGH19] Hamid Reza Moradi, Ibrahim Halil Gümüs, and Zahra Heydarbeygi. A glimpse at the operator Kantorovich inequality. Linear Multilinear Algebra, 67(5):1031–1036, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Murdoch:1974:BR

[MGK^{+}74] D. C. Murdoch, Larry J. Gerstein, David A. Klarner, Stephen Pierce, and Richard Jacobson. Book reviews. Linear and

Mahato:2023:SRE

[MGKA23] Iswar Mahato, R. Gurusamy, M. Rajesh Kannan, and S. Arockiaraj. On the spectral radius and the energy of eccentricity matrices of graphs. Linear Multilinear Algebra, 71 (1):5–15, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mattila:2012:SPR

[MH12] Mika Mattila and Pentti Haukkanen. Some properties of rowadjusted meet and join matrices. Linear Multilinear Algebra, 60(10):1211–1221, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2016:SLS

[MH16] Xiaoling Ma and Qiongxiang Huang. Signless Laplacian spectral characterization of 4-rose graphs. Linear Multilinear Algebra, 64(12):2474–2485, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marsli:2019:BEM

[MH19] Rachid Marsli and Frank J. Hall. On bounding the eigenvalues of matrices with constant row-sums. Linear Multilinear Algebra, 67(4):672–684, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marsli:2020:ECM

[MH20] Rachid Marsli and Frank J. Hall. Equivalence classes of ematrices and associated eigenvalue localization regions. Linear Multilinear Algebra, 68(5):915–930, 2020. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See typographical error correction [Ano22a].

Marsli:2022:SPE

[MH22] Rachid Marsli and Frank J. Hall. Some properties of ergodicity coefficients with applications in spectral graph theory. Linear Multilinear Algebra, 70(10):1886–1906, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martin:2017:CSL

[MHA17] A. J. Calderon Martin, A. S. Hegazi, and Hani Abdelwahab. A characterization of the semisimplity of Lie-type algebras through the existence of certain linear bases. Linear Multilinear Algebra, 65(9):1781–1792, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Manthey:1996:FFT

[MHH96] W. Manthey, D. Hinrichsen, and U. Helmke. On Fischer– Frobenius transformations and the structure of rectangular block Hankel matrices. Linear and Multilinear Algebra, 41(3): 255–288, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meng:2022:SBE

[MHM22] Lingsheng Meng, Yuwen He, and Shu-Xin Miao. Structured backward errors for two kinds of generalized saddle point systems. Linear Multilinear Algebra, 70(7):1345–1355, 2022. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Miao:1993:AMI

[Mia93a] Jianming Miao. Analog of the Minkowski inequality for the difference of two M-matrices. Linear and Multilinear Algebra, 34(2):133–139, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Miao:1993:RGI

[Mia93b] Jianming Miao. Reflexive generalized inverses and their minors. Linear and Multilinear Algebra, 35(2):153–163, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martin:2017:MLS

[MID17] Antonio J. Calderón Martín, Francisco J. Navarro Izquierdo, and José M. Sánchez Delgado. Modules over linear spaces admitting a multiplicative basis. Linear Multilinear Algebra, 65(1):156–165, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Militaru:2015:GEP

[Mil15] G. Militaru. The global extension problem, co-flag and metabelian Leibniz algebras. Linear Multilinear Algebra, 63 (3):601–621, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Milosevic:2022:ACS

[Mil22] Jovana Milosević. Algebraic conditions for the solvability of system of three linear equations in a ring. Linear Multilinear Algebra, 70(4):672–688, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1974:SIM

[Min74a] Henry Minc. The structure of irreducible matrices. Linear and Multilinear Algebra, 2(1):85–90, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1974:IM

[Min74b] Henryk Minc. Irreducible matrices. Linear and Multilinear Algebra, 1(4):337–342, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1974:SLA

[Min74c] Henryk Minc. Six letters from Alexander C. Aitken. Linear and Multilinear Algebra, 2(1):1–12, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1974:UCP

[Min74d] Henryk Minc. An unresolved conjecture on permanents of (0, 1)-matrices. Linear and Multilinear Algebra, 2(2):97–113, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). The conjecture appears in the author’s paper in Bull. Amer. Math. Soc. 69 (1963), 788–791, MR0155843.

Minc:1975:SDS

[Min75] Henryk Minc. Subpermanents of doubly stochastic matrices. Linear and Multilinear Algebra, 3(1–2):91–94, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1976:IES

[Min76] Henryk Minc. The invariance of elementary symmetric functions. Linear and Multilinear Algebra, 4(3):209–215, 1976.

Minc:1977:LTM

[Min77] Henryk Minc. Linear transformations on matrices: rank 1 preservers and determinant preservers. Linear and Multilinear Algebra, 4(4):265–272, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1980:ASM

[Min80a] Henryk Minc. An asymptotic solution of the multidimensional dimer problem. Linear and Multilinear Algebra, 8(3):235–239, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1980:BPD

[Min80b] Henryk Minc. Bounds for permanents and determinants. Linear and Multilinear Algebra, 9(1):5–16, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1982:IED

[Min82a] Henryk Minc. Inverse elementary divisors problem for doubly stochastic matrices. Linear and Multilinear Algebra, 11(2): 121–131, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1982:NEP

[Min82b] Henryk Minc. A note on Egoryčev’s proof of the van der Waerden conjecture. Linear and Multilinear Algebra, 11(4): 367–371, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1983:TP

[Min83] Henryk Minc. Theory of permanents 1978–1981. Linear and Multilinear Algebra, 12(4):227–263, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1984:MPD

[Min84] Henryk Minc. Minimum permanents of doubly stochastic matrices with prescribed zero entries. Linear and Multilinear Algebra, 15(3–4):225–243, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1987:TP

[Min87] Henryk Minc. Theory of permanents 1982–1985. Linear and Multilinear Algebra, 21(2):109–148, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Miranda:1981:NCV

[Mir81] M. E. Fernandes Miranda. Note on the characteristic values and the three types of singular values of a complex matrix. Linear and Multilinear Algebra, 10(2):155–161, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See note [Li84].

Miranda:1992:PDE

[Mir92] Maria Emília F. Miranda. On the product of the diagonal elements of Hermitian matrices. Linear and Multilinear Algebra, 31(1–4):225–234, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mirman:2009:PPF

[Mir09] Boris Mirman. Poncelet’s porism in the finite real plane. Linear Multilinear Algebra, 57(5):439–458, 2009. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mirzaei:2018:IEP

[Mir18] Hanif Mirzaei. Inverse eigenvalue problem for pseudosymmetric Jacobi matrices with two spectra. Linear Multilinear Algebra, 66(4):759–768, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mirafzal:2023:LGC

[Mir23] S. Morteza Mirafzal. The line graph of the crown graph is distance integral. Linear Multilinear Algebra, 71(4):662–672, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mishra:2017:FSA

[Mis17] Debasisha Mishra. Further study of alternating iterations for rectangular matrices. Linear Multilinear Algebra, 65(8):1566– 1580, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mitra:1992:BES

[Mit92] Sujit Kumar Mitra. On G-based extensions of the sharp order. Linear and Multilinear Algebra, 31(1–4):147–151, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mitsuhashi:2010:SFF

[Mit10] Hideo Mitsuhashi. A super Frobenius formula for the characters of Iwahori–Hecke algebras. Linear Multilinear Algebra, 58(8):941–955, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Miyajima:2022:VCM

[Miy22] Shinya Miyajima. Verified computation of matrix gamma function. Linear Multilinear Algebra, 70(7):1207–1229, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2007:GJD

[MJ07] Fei Ma and Guoxing Ji. Generalized Jordan derivations on triangular matrix algebras. Linear Multilinear Algebra, 55(4): 355–363, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meng:2016:PDS

[MK16] Jie Meng and Hyun-Min Kim. The positive definite solution to a nonlinear matrix equation. Linear Multilinear Algebra, 64(4):653–666, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1984:PED

[MKS84a] Marvin Marcus, Kent Kidman, and Markus Sandy. Products of elementary doubly stochastic matrices. Linear and Multilinear Algebra, 15(3–4):331–340, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1984:UIG

[MKS84b] Marvin Marcus, Kent Kidman, and Markus Sandy. Unitarily invariant generalized matrix norms and Hadamard products. Linear and Multilinear Algebra, 16(1–4):197–213, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mehatari:2022:AMC

[MKS22] Ranjit Mehatari, M. Rajesh Kannan, and Aniruddha Samanta. On the adjacency matrix of a complex unit gain graph. Linear Multilinear Algebra, 70(9):1798–1813, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Makar-Limanov:2012:CBH

[ML12] Leonid Makar-Limanov. A conjecture of Bavula on homomorphisms of the Weyl algebras. Linear Multilinear Algebra, 60(7):787–796, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2016:RBC

[ML16] Tianshui Ma and Linlin Liu. Rota–Baxter coalgebras and Rota–Baxter bialgebras. Linear Multilinear Algebra, 64(5): 968–979, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2021:CRC

[ML21a] Haifeng Ma and Tingting Li. Characterizations and representations of the core inverse and its applications. Linear Multilinear Algebra, 69(1):93–103, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mao:2021:SIA

[ML21b] Yanling Mao and Xilan Liu. On some inequalities for accretivedissipative matrices. Linear Multilinear Algebra, 69(9):1657– 1664, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1976:DCS

[MM76] Rueeell Merris and Marvin Marcus. The dimensions of certain symmetry classes of tensors. II. Linear and Multilinear Algebra, 4(3):205–207, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malek:1990:CCM

[MM90] Massoud Malek and H. Minc. On the convex combinations of matrices and the sum of subpermanents. Linear and Multilinear Algebra, 26(4):293–297, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mallol:1992:ABI

[MM92] Cristián Mallol and Artibano Micali. Sur les algèbres de Bernstein III. (French) [On Bernstein algebras III]. Linear and Multilinear Algebra, 31(1–4):85–91, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcelo:2005:IIF

[MM05] Agustín Marcelo and J. Muñoz Masqué. Irreducibility, integrality and factoriality of symmetric algebras. Linear Multilinear Algebra, 53(2):147–152, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malik:2006:SGA

[MM06] Madhur Malik and S. R. Mohan. Some geometrical aspects of semidefinite linear complementarity problems. Linear Multilinear Algebra, 54(1):55–70, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martin:2007:NIS

[MM07] Miguel Martín and Javier Merí. Numerical index of some polyhedral norms on the plane. Linear Multilinear Algebra, 55(2): 175–190, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martin:2009:NNI

[MM09] Miguel Martín and Javier Merí. A note on the numerical index of the L_{p} space of dimension two. Linear Multilinear Algebra, 57(2):201–204, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2011:INP

[MM11] Hongping Ma and Zhengke Miao. Imprimitive non-powerful sign pattern matrices with maximum base. Linear Multilinear Algebra, 59(4):371–390, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meng:2020:GFD

[MM20a] Jian Meng and Liquan Mei. A generalized fractional difference operator and its applications. Linear Multilinear Algebra, 68(9):1848–1860, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meng:2020:SGD

[MM20b] Jian Meng and Liquan Mei. The spectrum of the generalized difference operator B_{v}^{(m)} over the sequence space l_{p}(1 < p < ∞). Linear Multilinear Algebra, 68(7):1329–1339, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mao:2022:PTI

[MM22a] Yanling Mao and Yaping Mao. Popoviciu type inequalities for determinants. Linear Multilinear Algebra, 70(20):5032–5039, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mosic:2022:WWC

[MM22b] Dijana Mosić and Janko Marovt. Weighted weak core inverse of operators. Linear Multilinear Algebra, 70(20):4991–5013, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mukherjee:2023:GLF

[MM23] Sajal Kumar Mukherjee and Sanjay Mukherjee. Generalized Larcombe–Fenessey invariants of matrix powers. Linear Multilinear Algebra, 71(4):545–551, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moghaddam:2015:GIE

[MMG15] Mahsa R. Moghaddam, Hanif Mirzaei, and Kazem Ghanbari. On the generalized inverse eigenvalue problem of constructing symmetric pentadiagonal matrices from three mixed eigendata. Linear Multilinear Algebra, 63(6):1154–1166, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcoux:2007:AMV

[MMR07] L. W. Marcoux, M. Mastnak, and H. Radjavi. An approximate, multivariable version of Specht’s theorem. Linear Multilinear Algebra, 55(2):159–173, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mehl:2009:PAL

[MMRR09] Christian Mehl, Volker Mehrmann, André C. M. Ran, and Leiba Rodman. Perturbation analysis of Lagrangian invariant subspaces of symplectic matrices. Linear Multilinear Algebra, 57(2):141–184, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mehl:2016:EPT

[MMRR16] Christian Mehl, Volker Mehrmann, André C. M. Ran, and Leiba Rodman. Eigenvalue perturbation theory of structured real matrices and their sign characteristics under generic structured rank-one perturbations. Linear Multilinear Algebra, 64(3):527–556, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mitsuhashi:2021:NGB

[MMS21] Hideo Mitsuhashi, Hideaki Morita, and Iwao Sato. A new generalized Bartholdi zeta function for a digraph. Linear Multilinear Algebra, 69(6):1045–1060, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mahoney:1980:DAG

[MN80] Michael K. Mahoney and Morris Newman. Determinants of abelian group matrices. Linear and Multilinear Algebra, 9(2): 121–132, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mohan:1996:RRS

[MN96] S. R. Mohan and S. K. Neogy. The role of representative submatrices in vertical linear complementarity theory. Linear and Multilinear Algebra, 41(2):175–187, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Molitierno:2004:OWG

[MN04] Jason J. Molitierno and Michael Neumann. On operations in which graphs are appended to trees. Linear Multilinear Algebra, 52(2):145–156, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Molnar:2012:IRE

[MN12] Lajos Molnár and Gergő Nagy. Isometries and relative entropy preserving maps on density operators. Linear Multilinear Algebra, 60(1):93–108, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Midoune:2013:TAF

[MN13] N. Midoune and L. Noui. Trilinear alternating forms on a vector space of dimension 8 over a finite field. Linear Multilinear Algebra, 61(1):15–21, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McDonald:1998:ITM

[MNN^{+}98] J. J. McDonald, R. Nabben, M. Neumann, H. Scheider, and M. J. Tsatsomeros. Inverse tridiagonal Z-martices. Linear and Multilinear Algebra, 45(1):75–97, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merlet:2022:TBT

[MNS22] Glenn Merlet, Thomas Nowak, and Sergei Sergeev. On the tightness of bounds for transients of weak CSR expansions and periodicity transients of critical rows and columns of tropical matrix powers. Linear Multilinear Algebra, 70(19):4320–4350, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meisters:1991:SNH

[MO91] Gary H. Meisters and Czeslaw Olech. Strong nilpotence holds in dimensions up to five only. Linear and Multilinear Algebra, 30(4):231–255, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meisters:1993:PEN

[MO93] Gary H. Meisters and Czeslaw Olech. Power-exact, nilpotent, homogeneous matrices. Linear and Multilinear Algebra, 35(3– 4):225–236, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mathias:1994:INS

[MO94] R. Mathias and K. Okubo. The induced norm of the Schur multiplication operator with respect to the operator radius. Linear and Multilinear Algebra, 37(1–3):111–124, 1994. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mbekhta:2013:AMP

[MO13] Mostafa Mbekhta and Mourad Oudghiri. On additive maps preserving certain semi-Fredholm subsets. Linear Multilinear

Maniscalco:2022:FSF

[MO22] Sophia Maniscalco and Vignon Oussa. Full-spark frames arising from one-parameter groups. Linear Multilinear Algebra, 70(16):3031–3053, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moradi:2018:NSI

[MOGN18] Hamid Reza Moradi, Mohsen Erfanian Omidvar, Ibrahim Halil Gümüs, and Razieh Naseri. A note on some inequalities for positive linear maps. Linear Multilinear Algebra, 66(7):1449– 1460, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mohindru:2015:DJL

[Moh15] Preeti Mohindru. The Drew–Johnson–Loewy conjecture for matrices over max-min semirings. Linear Multilinear Algebra, 63(5):914–926, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mohammadian:2016:GTR

[Moh16a] A. Mohammadian. Graphs and their real eigenvectors. Linear Multilinear Algebra, 64(2):136–142, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mohammadian:2016:TAM

[Moh16b] Ali Mohammadian. Trees and acyclic matrices over arbitrary fields. Linear Multilinear Algebra, 64(3):466–476, 2016. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Molnar:2009:LMM

[Mol09] Lajos Molnár. Linear maps on matrices preserving commutativity up to a factor. Linear Multilinear Algebra, 57(1):13–18, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Molnar:2015:JTE

[Mol15] Lajos Molnár. Jordan triple endomorphisms and isometries of spaces of positive definite matrices. Linear Multilinear Alge-

Molitierno:2018:EGI

[Mol18] Jason J. Molitierno. Entries of the group inverse of the Laplacian matrix for generalized Johnson graphs. Linear Multilinear Algebra, 66(6):1153–1170, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moon:1976:NMG

[Moo76] J. W. Moon. A note on matrices with given diagonal entries. Linear and Multilinear Algebra, 4(3):177–178, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moon:1995:AMA

[Moo95] J. W. Moon. On the adjoint of a matrix associated with trees. Linear and Multilinear Algebra, 39(1–2):191–194, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moslehian:2012:KFI

[Mos12] Mohammad Sal Moslehian. Ky Fan inequalities. Linear Multilinear Algebra, 60(11–12):1313–1325, 2012. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mosic:2015:NCF

[Mos15] Dijana Mosić. A note on Cline’s formula for the generalized Drazin inverse. Linear Multilinear Algebra, 63(6):1106–1110, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mosic:2019:WCE

[Mos19] Dijana Mosić. Weighted core-EP inverse of an operator between Hilbert spaces. Linear Multilinear Algebra, 67(2):278– 298, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mosic:2020:MCO

[Mos20] Dijana Mosić. Maximal classes of operators determining some weighted generalized inverses. Linear Multilinear Algebra, 68 (11):2201–2220, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mosic:2021:CEI

[Mos21a] Dijana Mosić. Core-EP inverses in Banach algebras. Linear Multilinear Algebra, 69(16):2976–2989, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mosic:2021:GPS

[Mos21b] Dijana Mosić. The generalized and pseudo n-strong Drazin inverses in rings. Linear Multilinear Algebra, 69(2):361–375, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McDonald:2003:SSS

[MOTvdD03] J. J. McDonald, D. D. Olesky, M. J. Tsatsomeros, and P. van den Driessche. On the spectra of striped sign patterns. Linear Multilinear Algebra, 51(1):39–48, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mourad:2004:LTA

[Mou04] Bassam Mourad. On a Lie-theoretic approach to generalized doubly stochastic matrices and applications. Linear Multilinear Algebra, 52(2):99–113, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mourad:2013:GSR

[Mou13] Bassam Mourad. Generalization of some results concerning eigenvalues of a certain class of matrices and some applications. Linear Multilinear Algebra, 61(9):1234–1243, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mousavi:2020:TTS

[Mou20] Ramin Mousavi. Thin trees in some families of distanceregular graphs. Linear Multilinear Algebra, 68(7):1374–1383, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moucouf:2022:APP

[Mou22] Mohammed Mouçouf. Arbitrary positive powers of semicirculant and r-circulant matrices. Linear Multilinear Algebra, 70(21):6682–6697, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maybee:1990:PZE

[MOvdD90] John S. Maybee, D. D. Olesky, and P. van den Driessche. Partly zero eigenvectors. Linear and Multilinear Algebra, 28(1–2):83–92, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1973:EDH

[MP73] Russell Merris and Stephen Pierce. Elementary divisors of higher degree associated transformations. Linear and Multilinear Algebra, 1(3):241–250, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1987:CGN

[MP87a] Marvin Marcus and Claire Pesce. Computer generated numerical ranges and some resulting theorems. Linear and Multilinear Algebra, 20(2):121–157, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maxson:1987:SSM

[MP87b] Carlton J. Maxson and Günter F. Pilz. Simple subrings of matrix rings. Linear and Multilinear Algebra, 21(3):271–275, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maybee:1990:TMT

[MP90] John S. Maybee and Norman J. Pullman. Tournament matrices and their generalizations, I. Linear and Multilinear Algebra, 28(1–2):57–70, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mond:1994:MVK

[MP94] B. Mond and J. D. Pecarić. A matrix version of the Ky Fan generalization of the Kantorovich inequality. Linear and Multilinear Algebra, 36(3):217–221, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mond:1995:MVK

[MP95] B. Mond and J. E. Pecarić. A matrix version of the Ky Fan generalization of the Kantorovich inequality II. Linear and Multilinear Algebra, 38(4):309–313, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mono:1996:ISR

[MP96] B. Mono and J. E. Pecaric. On an inequality for spectral radius. Linear and Multilinear Algebra, 40(3):203–206, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maroulas:1998:CBN

[MP98] J. Maroulas and P. Psarrakos. A connection between numerical ranges of selfadjoint matrix polynomials. Linear and Multilinear Algebra, 44(4):327–340, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mary:2013:GIM

[MP13a] Xavier Mary and Pedro Patrício. Generalized inverses modulo h in semigroups and rings. Linear Multilinear Algebra, 61(8): 1130–1135, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Morris:2013:TFA

[MP13b] Karli Morris and Robert Perlis. Trace forms of abelian extensions of number fields of type (1, 0). Linear Multilinear Algebra, 61(2):233–253, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mary:2016:GIP

[MP16a] Xavier Mary and Pedro Patrício. The group inverse of a product. Linear Multilinear Algebra, 64(9):1776–1784, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mohindru:2016:OSC

[MP16b] Preeti Mohindru and Rajesh Pereira. Orderings on semirings and completely positive matrices. Linear Multilinear Algebra, 64(5):818–833, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mecheri:2018:QIO

[MP18] Salah Mecheri and S. M. Patel. On quasi-2-isometric operators. Linear Multilinear Algebra, 66(5):1019–1025, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mecheri:2019:TIO

[MP19] S. Mecheri and T. Prasad. Triangular n-isometric operators. Linear Multilinear Algebra, 67(6):1132–1145, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Montaner:2021:PTA

[MP21] Fernando Montaner and Irene Paniello. PI theory for associative pairs. Linear Multilinear Algebra, 69(14):2586–2606, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mal:2022:CEC

[MPD22] Arpita Mal, Kallol Paul, and Subhrajit Dey. Characterization of extreme contractions through k-smoothness of operators. Linear Multilinear Algebra, 70(20):5301–5315, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Maletesta:2021:RLA

[MPGS21] Nicola Maletesta, Pilar Páez-Guillán, and Salvatore Siciliano. Restricted Lie algebras having a distributive lattice of restricted subalgebras. Linear Multilinear Algebra, 69(16):3112– 3120, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Madariaga:2017:LRF

[MPI17] Sara Madariaga and José M. Pérez-Izquierdo. Linear representations of formal loops. Linear Multilinear Algebra, 65(1): 79–112, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meri:2021:NIA

[MQ21] Javier Merí and Alicia Quero. On the numerical index of absolute symmetric norms on the plane. Linear Multilinear Algebra, 69(5):971–979, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merris:1974:DCS

[MR74] Russell Merris and Muneer A. Rashid. The dimensions of certain symmetry classes of tensors. Linear and Multilinear Algebra, 2(3):245–248, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1975:BFG

[MR75] Marvin Marcus and Herbert Robinson. Bilinear functionals on the Grassmannian manifold. Linear and Multilinear Algebra, 3(3):215–225, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Markus:1997:SRN

[MR97] Alexander S. Markus and Leiba Rodman. Some results on numerical ranges and factorizations of matrix polynomials. Linear and Multilinear Algebra, 42(2):169–185, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mashreghi:2007:CAE

[MR07] Javad Mashreghi and Roland Rivard. On a conjecture about the eigenvalues of doubly stochastic matrices. Linear Multilinear Algebra, 55(5):491–498, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mazorchuk:2008:MMO

[MR08] Volodymyr Mazorchuk and Slavik Rabanovich. Multicommutators and multianticommutators of orthogonal projections. Linear Multilinear Algebra, 56(6):639–646, 2008. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moslehian:2010:GBI

[MR10] Mohammad Sal Moslehian and Rajna Rajić. Generalizations of Bohr’s inequality in Hilbert C^{∗}-modules. Linear Multilinear Algebra, 58(3):323–331, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marrero:2012:ACF

[MR12] J. Abderramán Marrero and M. Rachidi. Application of the companion factorization to linear non-autonomous areapreserving maps. Linear Multilinear Algebra, 60(2):201–217, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marrero:2013:NRI

[MR13] J. Abderramán Marrero and M. Rachidi. A note on representations for the inverses of tridiagonal matrices. Linear Multilinear Algebra, 61(9):1181–1191, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marrero:2015:IGH

[MR15] J. Abderramán Marrero and M. Rachidi. Inverses of generalized Hessenberg matrices. Linear Multilinear Algebra, 63(3): 559–570, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martinez-Rivera:2017:CFP

[MR17] Xavier Martínez-Rivera. Classification of families of pr- and epr-sequences. Linear Multilinear Algebra, 65(8):1581–1599, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martinez-Rivera:2018:SEP

[MR18] Xavier Martínez-Rivera. The signed enhanced principal rank characteristic sequence. Linear Multilinear Algebra, 66(7): 1484–1503, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Monsalve:2019:OBG

[MR19] Juan Monsalve and Juan Rada. Oriented bipartite graphs with minimal trace norm. Linear Multilinear Algebra, 67(6): 1121–1131, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marovt:2015:SLS

[MRD15] Janko Marovt, Dragan S. Rakić, and Dragan S. Djordjević. Star, left-star, and right-star partial orders in Rickart ∗-rings. Linear Multilinear Algebra, 63(2):343–365, 2015. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mortini:2010:MTR

[MRS10] Raymond Mortini, Rudolf Rupp, and Amol Sasane. Matricial topological ranks for two algebras of bounded holomorphic functions. Linear Multilinear Algebra, 58(6):741–752, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marrero:2012:NSO

[MRT12] J. Abderramán Marrero, M. Rachidi, and V. Tomeo. On the nested sums and orthogonal polynomials. Linear Multilinear Algebra, 60(8):995–1007, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malik:2014:FPC

[MRT14] Saroj B. Malik, Laura Rueda, and Néstor Thome. Further properties on the core partial order and other matrix partial orders. Linear Multilinear Algebra, 62(12):1629–1648, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malik:2016:CEN

[MRT16] Saroj B. Malik, Laura Rueda, and Néstor Thome. The class of m–EP and m-normal matrices. Linear Multilinear Algebra, 64(11):2119–2132, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marsaglia:1974:EIR

[MS74] George Marsaglia and George P. H. Styan. Equalities and inequalities for ranks of matrices. Linear and Multilinear Algebra, 2(3):269–292, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1979:NRC

[MS79] Marvin Marcus and B. N. Shure. The numerical range of certain 0, 1-matrices. Linear and Multilinear Algebra, 7(2): 111–120, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mirsky:1980:BR

[MS80] Leon Mirsky and Richard Sinkhorn. Book reviews. Linear and Multilinear Algebra, 9(1):77–79, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McMath:1981:MVT

[MS81] John S. McMath and S. E. Sims. Mean value theorems for a matrix valued derivative. Linear and Multilinear Algebra, 9(4): 255–261, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1982:BR

[MS82a] Marvin Marcus and Markus Sandy. The G-bilinear range. Linear and Multilinear Algebra, 11(4):317–332, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1982:TEP

[MS82b] Marvin Marcus and Markus Sandy. Three elementary proofs of the Goldberg–Straus theorem on numerical radii. Linear and Multilinear Algebra, 11(3):243–252, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malek-Shahmirzadi:1983:CCC

[MS83] Massoud Malek-Shahmirzadi. A characterization of certain classes of matrix norms. Linear and Multilinear Algebra, 13 (2):97–99, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1985:RPI

[MS85a] Marvin Marcus and Markus Sandy. Ryser’s permanent identity in the symmetric algebra. Linear and Multilinear Algebra, 18(3):183–196, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1985:SVN

[MS85b] Marvin Marcus and Markus Sandy. Singular values and numerical radii. Linear and Multilinear Algebra, 18(4):337–353, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1987:VPN

[MS87] Marvin Marcus and Markus Sandy. Vertex points in the numerical range of a derivation. Linear and Multilinear Algebra, 21(4):385–394, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1988:BIT

[MS88] Marvin Marcus and Markus Sandy. Bessel’s inequality in tensor space. Linear and Multilinear Algebra, 23(3):233–249, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1989:NDE

[MS89] Marvin Marcus and Terence R. Smith. A note on the determinants and eigenvalues of distance matrices. Linear and Multilinear Algebra, 25(3):219–230, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martins:1995:EMC

[MS95] Enide Andrade Martins and Fernando C. Silva. Eigenvalues of matrix commutators. Linear and Multilinear Algebra, 39(4): 375–389, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Molnar:1998:SLP

[MŠ98] Lajos Molnár and Peter Šemrl. Some linear preserver problems on upper triangular matrices. Linear and Multilinear Algebra, 45(2–3):189–206, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Molnar:2002:EOS

[MŠ02] Lajos Molnár and Peter Šemrl. Elementary operators on standard operator algebras. Linear Multilinear Algebra, 50(4):315– 319, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Madadi:2008:SCT

[MS08] A. R. Madadi and M. Shahryari. Symmetry classes of tensors as ∫↕_{\}(C)-modules. Linear Multilinear Algebra, 56(5):517– 541, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mbekhta:2009:LMP

[MŠ09] M. Mbekhta and P. Šemrl. Linear maps preserving semi-Fredholm operators and generalized invertibility. Linear Multilinear Algebra, 57(1):55–64, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2010:DCR

[MS10a] Wen-Xiu Ma and Boris Shekhtman. Do the chain rules for matrix functions hold without commutativity? Linear Multilinear Algebra, 58(1):79–87, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McKinley:2010:SBD

[MS10b] Tom McKinley and Boris Shekhtman. On simultaneous blockdiagonalization of cyclic sequences of commuting matrices. Linear Multilinear Algebra, 58(2):245–256, 2010. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mizuno:2010:WSM

[MS10c] Hirobumi Mizuno and Iwao Sato. Weighted scattering matrices of regular coverings of graphs. Linear Multilinear Algebra, 58(7):927–940, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mishra:2012:DNS

[MS12] Debasisha Mishra and K. C. Sivakumar. A dominance notion for singular matrices with applications to nonnegative generalized inverses. Linear Multilinear Algebra, 60(8):911–920, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Monfared:2015:SIE

[MS15a] Keivan Hassani Monfared and Bryan L. Shader. The λ − τ structured inverse eigenvalue problem. Linear Multilinear Algebra, 63(11):2275–2300, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mousavi:2015:JHR

[MS15b] Sayyed Ahmad Mousavi and Abbas Salemi. Joint higher rank numerical range of Pauli group. Linear Multilinear Algebra, 63(3):439–454, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mallik:2016:GMS

[MS16a] Sudipta Mallik and Bryan L. Shader. On graphs of minimum skew rank 4. Linear Multilinear Algebra, 64(2):279–289, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mousavi:2016:QEC

[MS16b] Sayyed Ahmad Mousavi and Abbas Salemi. Quantum error correction for a subgroup of the Pauli group using higher rank numerical range. Linear Multilinear Algebra, 64(1):68–77, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mehdipour:2018:NGI

[MS18] Mahboobeh Mehdipour and Abbas Salemi. On a new generalized inverse of matrices. Linear Multilinear Algebra, 66(5):

Moradi:2021:EIT

[MS21a] Hamid Reza Moradi and Mohammad Sababheh. Eigenvalue inequalities for n-tuple of matrices. Linear Multilinear Algebra, 69(12):2192–2203, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Moradi:2021:MAN

[MS21b] Hamid Reza Moradi and Mohammad Sababheh. More accurate numerical radius inequalities (II). Linear Multilinear Algebra, 69(5):921–933, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Malek-Shahmirzadi:1983:CCM

[MSO83] Massoud Malek-Shahmirzadi and Kunio Oshima. On certain classes of matrix norms. Linear and Multilinear Algebra, 14(4): 357–363, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Majdak:2013:EPO

[MSS13] Witold Majdak, Nicolae-Adrian Secelean, and Laurian Suciu. Ergodic properties of operators in some semi-Hilbertian spaces. Linear Multilinear Algebra, 61(2):139–159, 2013. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Miyazaki:2018:TRC

[MSS18] Mitsuhiro Miyazaki, Toshio Sumi, and Toshio Sakata. Typical ranks of certain 3-tensors and absolutely full column rank tensors. Linear Multilinear Algebra, 66(1):193–205, 2018. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mohtashami:2019:LPC

[MSS19] Shiva Mohtashami, Abbas Salemi, and Mohammad Soleymani. Linear preservers of circulant majorization on rectangular matrices. Linear Multilinear Algebra, 67(8):1554–1560, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Martinez:2004:CMW

[MSZ04] Servet Martínez, Jaime San Martín, and Xiao-Dong Zhang. A class of M-matrices whose graphs are trees. Linear Multilinear Algebra, 52(5):303–319, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1983:SBR

[MT83] M. Marcus and R. C. Thompson. Short book review. Linear and Multilinear Algebra, 14(3):305, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1984:LMA

[MT84] Marvin Marcus and Robert C. Thompson. Linear and Multilinear Algebra. Linear and Multilinear Algebra, 15(1):2, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Markus:2011:CSF

[MT11] Alexander Markus and Tin-Yau Tam. Connectedness of some fibres on a compact connected Lie group. Linear Multilinear Algebra, 59(10):1121–1126, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2022:MAG

[MT22a] Rong Ma and Yongge Tian. A matrix approach to a general partitioned linear model with partial parameter restrictions. Linear Multilinear Algebra, 70(13):2513–2532, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Medina:2022:BRD

[MT22b] Luis Medina and Macarena Trigo. Bounds on the Reciprocal distance energy and Reciprocal distance Laplacian energies of a graph. Linear Multilinear Algebra, 70(16):3097–3118, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mohammadian:2011:SCI

[MTR11] A. Mohammadian and B. Tayfeh-Rezaie. Some constructions of integral graphs. Linear Multilinear Algebra, 59(11):1269– 1279, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mohammadian:2019:HMF

[MTR19] A. Mohammadian and B. Tayfeh-Rezaie. Hadamard matrices with few distinct types. Linear Multilinear Algebra, 67(8): 1596–1605, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mukherjee:2020:RFT

[Muk20] Sajal Kumar Mukherjee. On the rooted forests in triangulated closed manifolds. Linear Multilinear Algebra, 68(10): 2034–2043, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mummy:1986:FTM

[Mum86a] Mark Stephen Mummy. Full transversal matroids, strict gammoids, and the matroid components problem. Linear and Multilinear Algebra, 19(2):97–116, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mummy:1986:GCI

[Mum86b] Mark Stephen Mummy. The generalized column incidence graph and a matroid base-listing algorithm. Linear and Multilinear Algebra, 19(1):33–49, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mummy:1986:SMM

[Mum86c] Mark Stephen Mummy. Strict matching matroids and matroid algorithms. Linear and Multilinear Algebra, 19(3):221–247, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Murota:1989:ILM

[Mur89] Kazuo Murota. On the irreducibility of layered mixed matrices. Linear and Multilinear Algebra, 24(4):273–288, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mann:1975:NRE

[MV75] Henry B. Mann and William Yslas Vélez. On normal radical extensions of the rationals. Linear and Multilinear Algebra, 3(1–2):73–80, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Merikoski:1991:LHI

[MV91] Jorma Kaarlo Merikoski and Ari Virtanen. On the London– Hoffman inequality for the sum of elements of powers of nonnegative matrices. Linear and Multilinear Algebra, 30(4):257– 259, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Macias-Virgos:2014:TAL

[MVPS14] E. Macías-Virgós and M. J. Pereira-Sáez. A topological approach to left eigenvalues of quaternionic matrices. Linear Multilinear Algebra, 62(2):139–158, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Macias-Virgos:2023:RQL

[MVPSTT23] E. Macías-Virgós, M. J. Pereira-Sáez, and A. D. Tarrío-Tobar. Rayleigh quotient and left eigenvalues of quaternionic matrices. Linear Multilinear Algebra, 71(13):2163–2179, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1974:APS

[MW74a] Marvin Marcus and William Watkins. Annihilating polynomials for similarity and commutator mappings. Linear and Multilinear Algebra, 2(1):81–84, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1974:EDD

[MW74b] Marvin Marcus and William Watkins. Elementary divisors of derivations. Linear and Multilinear Algebra, 2(1):65–80, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcus:1980:SVN

[MW80] Marvin Marcus and Bo-Ying Wang. Some variations on the numerical range. Linear and Multilinear Algebra, 9(2):111– 120, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2015:NCG

[MW15] Xiaobin Ma and Dein Wong. The nullity of k-cyclic graphs of ∞-type. Linear Multilinear Algebra, 63(11):2200–2211, 2015.

Ma:2016:AGI

[MW16] Xiaobin Ma and Dein Wong. Automorphism group of an idealrelation graph over a matrix ring. Linear Multilinear Algebra, 64(2):309–320, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ma:2018:INS

[MW18] Xiaobin Ma and Dengyin Wang. Independence number of subspace inclusion graph and subspace sum graph of a vector space. Linear Multilinear Algebra, 66(12):2430–2437, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meng:2019:WHP

[MW19] Qing Meng and Liguang Wang. Weak Haagerup property of dynamical systems. Linear Multilinear Algebra, 67(7):1294– 1307, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meng:2021:HAP

[MW21] Qing Meng and Liguang Wang. The Haagerup approximation property for arbitrary C∗-algebras. Linear Multilinear Algebra, 69(7):1275–1285, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mosic:2022:LRO

[MW22] Dijana Mosić and Long Wang. Left and right G-outer inverses. Linear Multilinear Algebra, 70(17):3319–3344, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mao:2023:GSC

[MW23] Lihuan Mao and Wei Wang. Generalized spectral characterization of rooted product graphs. Linear Multilinear Algebra, 71(14):2310–2324, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Miao:2022:FOT

[MWC22] Yun Miao, Yimin Wei, and Zhen Chen. Fourth-order tensor Riccati equations with the Einstein product. Linear Multi-

linear Algebra, 70(10):1831–1853, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Miao:2022:TLI

[MWLW22] Feifei Miao, Xiao Wang, Lei Li, and Liguang Wang. Two-local isometries on vector-valued differentiable functions. Linear Multilinear Algebra, 70(13):2505–2512, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wei:2011:NAR

[mWyD11] Yi min Wei and Chun yuan Deng. A note on additive results for the Drazin inverse. Linear Multilinear Algebra, 59(12): 1319–1329, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McDonald:2012:CSA

[MY12] J. J. McDonald and A. A. Yielding. Complex spectrally arbitrary zero-nonzero patterns. Linear Multilinear Algebra, 60 (1):11–26, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

McDaniel:1991:COW

[MZ91] Andrew McDaniel and Stanley M. Zoltek. Curvature operators which preserve normal spaces. Linear and Multilinear Algebra, 30(3):183–194, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marcaida:2004:IFW

[MZ04] S. Marcaida and I. Zaballa. Invariant factors, Wiener–Hopf factorization indices, and invariant factors at infinity of rational matrices. Linear Multilinear Algebra, 52(6):427–439, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Meng:2015:NMP

[MZ15] Lingsheng Meng and Bing Zheng. New multiplicative perturbation bounds of the Moore–Penrose inverse. Linear Multilinear Algebra, 63(5):1037–1048, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Marko:2017:MDI

[MZ17] Frantisek Marko and Alexandr N. Zubkov. Minimal degrees of invariants of (super)groups — a connection to cryptology.

Moucouf:2022:NAC

[MZ22] Mohammed Mouçouf and Said Zriaa. A new approach for computing the inverse of confluent Vandermonde matrices via Taylor’s expansion. Linear Multilinear Algebra, 70(20):5973– 5986, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Monika:2023:NI

[MZ23] Monika and Bentuo Zheng. The numerical index of ℓp^{2}. Linear Multilinear Algebra, 71(5):768–773, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nabben:1998:MWI

[Nab98] Reinhard Nabben. Matrices whose inverses are generalized Mmatrices. Linear and Multilinear Algebra, 44(1):67–84, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nagy:2017:DPM

[Nag17] Gergő Nagy. Determinant preserving maps: an infinite dimensional version of a theorem of Frobenius. Linear Multilinear Algebra, 65(2):351–360, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nakazato:1995:BRC

[Nak95] Hiroshi Nakazato. The boundary of the range of a constrained sesquilinear form. Linear and Multilinear Algebra, 40(1):37– 43, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nassopoulos:1974:ECC

[Nas74] George F. Nassopoulos. On the existence of coproducts in the category of noncommutative algebras. Linear and Multilinear Algebra, 2(1):39–54, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nash:1975:GEY

[Nas75] David Nash. The generalized Eckart–Young principal axis theorem. Linear and Multilinear Algebra, 2(4):353–355, 1975.

Nunez-Alarcon:2019:BCM

[NASR19] Daniel Núñez-Alarcón and Diana Marcela Serrano-Rodríguez. The best constants in the multiple Khintchine inequality. Linear Multilinear Algebra, 67(11):2325–2344, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nath:2013:MCB

[Nat13] Milan Nath. Minimal configuration bicyclic graphs. Linear Multilinear Algebra, 61(2):255–264, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Naveen:2022:SLE

[NC22] Palanivel Naveen and A. V. Chithra. Signless Laplacian energy, distance Laplacian energy and distance signless Laplacian spectrum of unitary addition Cayley graphs. Linear Multilinear Algebra, 70(22):7514–7535, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neogy:2005:ATC

[ND05] S. K. Neogy and A. K. Das. On almost type classes of matrices with Q-property. Linear Multilinear Algebra, 53(4):243–257, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neogy:2013:WGP

[ND13] S. K. Neogy and A. K. Das. On weak generalized positive subdefinite matrices and the linear complementarity problem. Linear Multilinear Algebra, 61(7):945–953, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nedoma:1998:ILE

[Ned98] Josef Nedoma. Inaccurate linear equation system with a restricted-rank error matrix. Linear and Multilinear Algebra, 44(1):29–44, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neretin:2022:DCG

[Ner22] Yury A. Neretin. On double cosets of groups GL(n) with respect to subgroups of block strictly triangular matrices. Linear

Neto:2022:MCO

[Net22] Antônio Francisco Neto. Matrix computations with the Omega calculus. Linear Multilinear Algebra, 70(20):5075–5106, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neuberger:1974:NSP

[Neu74] J. W. Neuberger. Norm of symmetric product compared with norm of tensor product. Linear and Multilinear Algebra, 2(2): 115–121, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:1979:WSM

[Neu79] M. Neumann. Weak stability for matrices. Linear and Multilinear Algebra, 7(3):257–262, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:1981:SCL

[Neu81] M. Neumann. On the Schur complement and the LUfactorization of a matrix. Linear and Multilinear Algebra, 9(4): 241–254, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neudecker:1983:JTS

[Neu83] H. Neudecker. On Jacobians of transformations with skewsymmetric, strictly (lower) triangular or diagonal matrix arguments. Linear and Multilinear Algebra, 14(3):271–295, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:1984:BCS

[Neu84] Michael Neumann. On bounds for the convergence of the SSOR method for H-matrices. Linear and Multilinear Algebra, 15(1):13–21, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1974:RIS

[New74] Morris Newman. A result on integral symmetric matrices. Linear and Multilinear Algebra, 2(2):143–145, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1975:SRS

[New75] Morris Newman. Some results on symplectic matrices. Linear and Multilinear Algebra, 3(1–2):95–98, 1975. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1976:TC

[New76a] Morris Newman. On a theorem of Cebotarev. Linear and Multilinear Algebra, 3(4):259–262, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1976:PM

[New76b] Morris Newman. Polynomials in a matrix. Linear and Multilinear Algebra, 3(4):263–266, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1978:NME

[New78] Morris Newman. A note on matrix equivalence. Linear and Multilinear Algebra, 5(4):265–266, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1979:EDD

[New79] Morris Newman. Equivalence without determinantal divisors. Linear and Multilinear Algebra, 7(2):107–109, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1980:DCP

[New80a] Morris Newman. Determinants of circulants of prime power order. Linear and Multilinear Algebra, 9(3):187–191, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1980:MFP

[New80b] Morris Newman. Matrices of finite period and some special linear equations. Linear and Multilinear Algebra, 8(3):189– 195, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1982:RAD

[New82a] Morris Newman. A result about determinantal divisors. Linear and Multilinear Algebra, 11(4):363–366, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1982:S

[New82b] Morris Newman. Similarity over SL(n, F). Linear and Multilinear Algebra, 12(3):223–226, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1983:PHJ

[New83] Morris Newman. On a problem of H. J. Ryser. Linear and Multilinear Algebra, 12(4):291–293, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1985:GSR

[New85a] Morris Newman. Generating sets for rings of algebraic integers. Linear and Multilinear Algebra, 18(3):207–212, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1985:S

[New85b] Morris Newman. Similarity over Sp(2n, Z). Linear and Multilinear Algebra, 17(3–4):377–378, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1993:UDR

[New93] Morris Newman. Units differing by rationals in a cyclotomic field. Linear and Multilinear Algebra, 34(1):55–57, 1993. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1994:MSN

[New94] Morris Newman. Multiplicativity of the Smith normal form. Linear and Multilinear Algebra, 36(3):223–224, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1997:WRC

[New97] Morris Newman. The work of Robert C. Thompson. Linear and Multilinear Algebra, 43(1–3):1–6, 1997. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ntogramatzidis:2019:GGA

[NF19] Lorenzo Ntogramatzidis and Augusto Ferrante. The geometry of the generalized algebraic Riccati equation and of the singular Hamiltonian system. Linear Multilinear Algebra, 67(1):

Ng:2014:AIP

[Ng14] Chi-Keung Ng. An application of infinite projective tensor products of normed *-algebras. Linear Multilinear Algebra, 62(7):875–883, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ning:2015:SRH

[NG15] Bo Ning and Jun Ge. Spectral radius and Hamiltonian properties of graphs. Linear Multilinear Algebra, 63(8):1520–1530, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ng:2022:REG

[Ng22] P. W. Ng. Remarks on the extension group for purely infinite corona algebras. Linear Multilinear Algebra, 70(13): 2459–2504, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Niezgoda:2007:BMI

[Nie07] Marek Niezgoda. G-majorization inequalities for linear maps, II. Linear Multilinear Algebra, 55(1):1–17, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Niezgoda:2010:CSR

[Nie10] Marek Niezgoda. Complements to some results on eigenvalues and singular values of matrices. Linear Multilinear Algebra, 58 (1):1–13, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Niemiec:2014:FCD

[Nie14] Piotr Niemiec. Functional calculus for diagonalizable matrices. Linear Multilinear Algebra, 62(3):297–321, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nien:2022:CLH

[Nie22a] Chufeng Nien. Characters of 2-layered Heisenberg groups. Linear Multilinear Algebra, 70(19):3633–3642, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Niezgoda:2022:MIG

[Nie22b] Marek Niezgoda. On G -majorization inequalities for gradients of G -increasing functions. Linear Multilinear Algebra, 70(22): 7307–7317, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Niezgoda:2022:MLI

[Nie22c] Marek Niezgoda. On Maligranda like inequalities for Gmajorization. Linear Multilinear Algebra, 70(2):264–279, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nikoufar:2022:NCO

[Nik22] Ismail Nikoufar. A new characterization of the operator perspective. Linear Multilinear Algebra, 70(19):4297–4319, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Niroomand:2011:TSN

[Nir11] Peyman Niroomand. On the tensor square of non-abelian nilpotent finite-dimensional Lie algebras. Linear Multilinear Algebra, 59(8):831–836, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Niroomand:2019:CLA

[NJP19] Peyman Niroomand, Farangis Johari, and Mohsen Parvizi. Capable Lie algebras with the derived subalgebra of dimension 2 over an arbitrary field. Linear Multilinear Algebra, 67(3): 542–554, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nader:2022:PSD

[NMAB22] Rafic Nader, Bassam Mourad, Hassan Abbas, and Alain Bretto. On the positive semi-definite p th roots of positive semi-definite doubly stochastic matrices. Linear Multilinear Algebra, 70(15):2836–2855, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nakazato:1995:BNR

[NNT95] Hiroshi Nakazato, Yasutaka Nishikawa, and Makoto Takaguchi. On the boundary of the C-numerical range of a matrix.

Linear and Multilinear Algebra, 39(3):231–240, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Noferini:2015:WHM

[Nof15] Vanni Noferini. When is a Hamiltonian matrix the commutator of two skew-Hamiltonian matrices? Linear Multilinear Algebra, 63(8):1531–1552, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nomizu:1973:CRV

[Nom73] Katsumi Nomizu. Characteristic roots and vectors of a differentiable family of symmetric matrices. Linear and Multilinear Algebra, 1(2):159–162, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Norman:1995:JFT

[Nor95] C. W. Norman. On the Jordan form of the tensor product over fields of prime characteristic. Linear and Multilinear Algebra, 38(4):351–371, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Norman:2008:JBT

[Nor08] C. W. Norman. On Jordan bases for the tensor product and Kronecker sum and their elementary divisors over fields of prime characteristic. Linear Multilinear Algebra, 56(4):415– 451, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1973:BMG

[NP73] M. Newman and S. Pierce. Bounded matrix groups. Linear and Multilinear Algebra, 1(3):251–256, 1973. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:1980:MCI

[NP80] M. Neumann and R. J. Plemmons. M-matrix characterization II: General M-matrices. Linear and Multilinear Algebra, 9(3): 211–225, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nomizu:1991:LGR

[NP91] Katsumi Nomizu and Ulrich Pinkall. Lorentzian geometry for 2 × 2 real matrices. Linear and Multilinear Algebra, 28(4):

Neumann:2009:LSG

[NP09] M. Neumann and S. Pati. The Laplacian spectra of graphs with a tree structure. Linear Multilinear Algebra, 57(3):267– 291, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nath:2013:DSR

[NP13] Milan Nath and Somnath Paul. On the distance spectral radius of trees. Linear Multilinear Algebra, 61(7):847–855, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nath:2015:SGE

[NP15] Milan Nath and Somnath Paul. On the spectra of graphs with edge-pockets. Linear Multilinear Algebra, 63(3):509–522, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Narkuziyev:2023:CCT

[NR23] B. A. Narkuziyev and U. A. Rozikov. Classification in chains of three-dimensional real evolution algebras. Linear Multilinear Algebra, 71(2):265–300, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:1985:BRP

[NS85] Michael Neumann and Ronald J. Stern. Boundary results for positively invariant cones and their reachability cones. Linear and Multilinear Algebra, 17(2):143–154, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:1992:PCM

[NS92] Michael Neumann and Hans Schneider. Principal components of minus M-matrices. Linear and Multilinear Algebra, 32(2): 131–148, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrections and additions [NS93].

Neumann:1993:CAP

[NS93] Michael Neumann and Hans Schneider. Corrections and additions to: “Principal components of minus m-matrices”. Linear and Multilinear Algebra, 36(2):147–149, 1993. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [NS92].

Neto:1999:SRD

[NS99] O. Neto and F. C. Silva. Singular regular differential equations and eigenvalues of products of matrices. Linear and Multilinear Algebra, 46(1–2):145–164, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nath:2009:MCU

[NS09] Milan Nath and Bhaba Kumar Sarma. Minimal configuration unicyclic graphs. Linear Multilinear Algebra, 57(1):19–27, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Narayan:2015:BMS

[NS15] Sivaram K. Narayan and Yousra Sharawi. Bounds on minimum semidefinite rank of graphs. Linear Multilinear Algebra, 63(4):774–787, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nikoufar:2018:CLH

[NS18] Ismail Nikoufar and Moosa Shamohammadi. The converse of the Loewner–Heinz inequality via perspective. Linear Multilinear Algebra, 66(2):243–249, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Numazawa:2022:CNK

[NS22] Yuki Numazawa and Takashi Sano. Conditional negativity of Kwong matrices II. Linear Multilinear Algebra, 70(3):511–516, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1989:CDC

[NST89] Morris Newman, Wasin So, and Robert C. Thompson. Convergence domains for the Campbell–Baker–Hausdorff formula. Linear and Multilinear Algebra, 24(4):301–310, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:1991:RCE

[NST91] Michael Neumann, Ronald J. Stern, and Michael J. Tsatsomeros. The reachability cones of essentially nonnegative matrices. Linear and Multilinear Algebra, 28(4):213–224, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nishida:2021:MPA

[NSW21] Yuki Nishida, Kohei Sato, and Sennosuke Watanabe. A minplus analogue of the Jordan canonical form associated with the basis of the generalized eigenspace. Linear Multilinear Algebra, 69(15):2933–2943, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Newman:1986:IFG

[NT86] Morris Newman and Robert C. Thompson. Invariant factors and generalized Schur matrices. Linear and Multilinear Algebra, 19(4):399–404, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:1991:SPL

[NT91] Michael Neumann and Michael J. Tsatsomeros. Symbiosis points for linear differential systems. Linear and Multilinear Algebra, 30(1–2):49–59, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nylen:1994:SHM

[NT94] Peter Nylen and Tin-Yau Tam. On the spread of a Hermitian matrix and a conjecture of Thompson. Linear and Multilinear Algebra, 37(1–3):3–11, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Naves:2022:GI

[NT22] Fernando Augusto Naves and Humberto Luiz Talpo. Graded A-identities for M_{1,1}(E). Linear Multilinear Algebra, 70(20): 5373–5391, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nylen:1993:EPS

[NTU93] Peter Nylen, Tin-Yau Tam, and Frank Uhlig. On the eigenvalues of principal submatrices of normal, Hermitian and symmetric matrices. Linear and Multilinear Algebra, 36(1):69–78, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nam:2023:SGM

[NTW23] Taeuk Nam, Cindy Tan, and Ben Williams. Spaces of generators for matrix algebras with involution. Linear Multilinear Algebra, 71(8):1407–1430, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Nylen:1994:RIT

[NU94] Peter Nylen and Frank Uhlig. Realizations of interlacing by tree-patterned matrics. Linear and Multilinear Algebra, 38(1– 2):13–37, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neubauer:2002:ODS

[NW02] Michael Neubauer and William Watkins. D-optimal designs for seven objects and a large number of weighings. Linear Multilinear Algebra, 50(1):61–74, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ng:2021:PNU

[NY21] Chi-Keung Ng and ZhaoLin Yao. Property T for non-unital C *-crossed products. Linear Multilinear Algebra, 69(14):2629– 2640, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Neumann:2023:LTC

[NY23] Hans Henrich Neumann and Makoto Yamashita. Lieb type convexity for positive operator monotone decreasing functions. Linear Multilinear Algebra, 71(5):728–736, 2023. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ning:2021:BLD

[NZK21] Tong Ning, Jianhua Zhang, and Liang Kong. Bilocal Lie derivations on factor von Neumann algebras. Linear Multilinear Algebra, 69(8):1468–1483, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oblak:2012:NCN

[Obl12] Polona Oblak. On the nilpotent commutator of a nilpotent matrix. Linear Multilinear Algebra, 60(5):599–612, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oboudi:2016:CCB

[Obo16] Mohammad Reza Oboudi. Cospectrality of complete bipartite graphs. Linear Multilinear Algebra, 64(12):2491–2497, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oboudi:2022:CRS

[Obo22] Mohammad Reza Oboudi. On a conjecture related to the smallest signless Laplacian eigenvalue of graphs. Linear Multilinear Algebra, 70(19):4425–4431, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oboudi:2023:RWS

[Obo23] Mohammad Reza Oboudi. Reverse Wiener spectral radius of trees. Linear Multilinear Algebra, 71(2):256–264, 2023. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Odoni:1983:SRV

[Odo83] R. W. K. Odoni. Some remarks on the values taken by abelian group determinants. Linear and Multilinear Algebra, 14(3): 297–303, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oger:1994:IEE

[Oge94] Francis Oger. Isomorphism and elementary equivalence of multilinear maps. Linear and Multilinear Algebra, 36(3):151–174, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ogur:2020:SOS

[Ogu20] Oguz Ogur. Superposition operators on sequence spaces ℓ_{p}(F) derived by using matrix of Fibonacci numbers. Linear Multilinear Algebra, 68(10):2087–2098, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oi:2020:HOI

[Oi20] Shiho Oi. Hermitian operators and isometries on algebras of matrix-valued Lipschitz maps. Linear Multilinear Algebra, 68 (6):1096–1112, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1973:CML

[– Dok73] D. Ž. – Doković. A characterization of minimal left or right ideals of matrix algebras. Linear and Multilinear Algebra, 1(2): 139–147, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1975:NIS

[– Dok75] D. Ž. – Doković. On Neuberger’s identity for symmetric tensors. Linear and Multilinear Algebra, 2(4):305–309, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1981:PSS

[– Dok81] Dragomir Ž. – Doković. Point-sequences on spheres and the spin groups. Linear and Multilinear Algebra, 9(4):263–270, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1983:CTE

[– Dok83] Dragomir Ž. – Doković. Classification of trivectors of an eightdimensional real vector space. Linear and Multilinear Algebra, 13(1):3–39, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1990:IMS

[– Dok90] Dragomir Ž. – Doković. On isometries of matrix spaces. Linear and Multilinear Algebra, 27(2):73–78, 1990. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:1993:CHM

[– Dok93] Dragomir Ž. – Doković. Correction to “Hermitian matrices of order 2 over the ring of finite Fourier series with complex coefficients”. Linear and Multilinear Algebra, 34(3–4):403–406, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [– DL92].

Orel:2007:AMH

[OK07] M. Orel and B. Kuzma. Additive maps on Hermitian matrices. Linear Multilinear Algebra, 55(6):599–617, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:2009:OSO

[– Dok09] Dragomir Ž. – Doković. On orthogonal and special orthogonal invariants of a single matrix of small order. Linear Multilinear Algebra, 57(4):345–354, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Dokovic:2017:GST

[– Dok17] Dragomir Z. – Doković. Generalization of Scarpis’ theorem on Hadamard matrices. Linear Multilinear Algebra, 65(10):1985– 1987, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Okubo:1991:CST

[Oku91] K. Okubo. A Cauchy–Schwarz type norm inequality. Linear and Multilinear Algebra, 30(1–2):109–115, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Okubo:1993:SEC

[Oku93] K. Okubo. Some equality conditions with respect to the dual norm of the numerical radius. Linear and Multilinear Algebra, 34(3–4):365–376, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Omidvar:2021:NEN

[OM21] Mohsen Erfanian Omidvar and Hamid Reza Moradi. New estimates for the numerical radius of Hilbert space operators. Linear Multilinear Algebra, 69(5):946–956, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Omladic:2014:MNM

[Oml14] Matjaž Omladič. On maximal nilspaces of matrices. Linear Multilinear Algebra, 62(9):1258–1265, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

OMeara:2013:SDM

[OP13] Corey O’Meara and Rajesh Pereira. Self-dual maps and symmetric bistochastic matrices. Linear Multilinear Algebra, 61 (1):23–34, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Omladic:2009:NCR

[OR09] Matjaž Omladič and Heydar Radjavi. Nilpotent commutators and reducibility of semigroups. Linear Multilinear Algebra, 57(3):307–317, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

ORyan:2015:PHD

[O’R15] Manuel O’Ryan. Products of higher degree forms. Linear Multilinear Algebra, 63(1):34–45, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Orel:2019:NRO

[Ore19] Marko Orel. Nonstandard rank-one nonincreasing maps on symmetric matrices. Linear Multilinear Algebra, 67(2):391– 432, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Orlitzky:2018:LRP

[Orl18] Michael Orlitzky. Lyapunov rank of polyhedral positive operators. Linear Multilinear Algebra, 66(5):992–1000, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Omirov:2015:RCE

[ORT15] B. A. Omirov, U. A. Rozikov, and K. M. Tulenbayev. On real chains of evolution algebras. Linear Multilinear Algebra, 63(3): 586–600, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Orzech:1984:SVR

[Orz84] Grace Orzech. Several versions of the resultant of two polynomials. Linear and Multilinear Algebra, 16(1–4):275–282, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oussa:2018:DGF

[OS18] Vignon Oussa and Brian Sheehan. Dihedral group frames with the Haar property. Linear Multilinear Algebra, 66(6):1121– 1132, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ostrand:1974:CNG

[Ost74] Phillip A. Ostrand. Cutting numbers in graphs. Linear and Multilinear Algebra, 2(1):25–38, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oliveira:2020:SRO

[OST20] Elismar R. Oliveira, Dragan Stevanović, and Vilmar Trevisan. Spectral radius ordering of starlike trees. Linear Multilinear Algebra, 68(5):991–1000, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Otake:2013:CND

[Ota13] Shuichi Otake. Counting the number of distinct real roots of certain polynomials by Bezoutian and the Galois groups over the rational number field. Linear Multilinear Algebra, 61(4): 429–441, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Otachel:2020:RSB

[Ota20] Zdzisław Otachel. On reverse Schwarz, Bessel, grüss and Hadamard inequalities in inner product spaces. Linear Multilinear Algebra, 68(4):805–827, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Osaka:2022:IFI

[OU22] Hiroyuki Osaka and Yoichi Udagawa. Interpolation functions and inequalities. Linear Multilinear Algebra, 70(20):4897– 4906, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oussa:2015:CVP

[Ous15a] Vignon Oussa. Computing vergne polarizing subalgebras. Linear Multilinear Algebra, 63(3):578–585, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Oussa:2015:DGF

[Ous15b] Vignon Oussa. Dihedral group frames which are maximally robust to erasures. Linear Multilinear Algebra, 63(12):2494– 2510, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See correction of communicating editor [Ano16c].

Olesky:1977:NMN

[OvdD77] D. D. Olesky and P. van den Driessche. A note on minimal norms and the logarithmic derivative of a matrix. Linear and Multilinear Algebra, 5(3):229–233, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Okubo:2002:RRM

[OW02] K. Okubo and H. J. Woerdeman. Rank reducing matrix norms. Linear Multilinear Algebra, 50(2):185–197, 2002. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ou:2021:FNM

[OWTZ21] Shikun Ou, Dein Wong, Fenglei Tian, and Qi Zhou. Fixing number and metric dimension of a zero-divisor graph associated with a ring. Linear Multilinear Algebra, 69(10):1789– 1802, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ou:2009:ANL

[OWX09] Shikun Ou, Dengyin Wang, and Chunguang Xia. Automorphisms of a nilpotent Lie algebra over a commutative ring. Linear Multilinear Algebra, 57(1):75–85, 2009. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pless:1979:BR

[PA79] Vera Pless and Charles A. Akemann. Book reviews. Linear and Multilinear Algebra, 8(1):87–88, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pisonero:1994:DMF

[PA94] Miriam Pisonero and G. Almkvist. Determinants on modules that have a finite free resolution. Linear and Multilinear Algebra, 38(1–2):97–107, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Patricio:2010:MPI

[PA10] Pedro Patrício and C. Mendes Araújo. Moore–Penrose invertibility in involutory rings: the case aa^{+} = bb^{+}. Linear Multilinear Algebra, 58(4):445–452, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

PablosRomo:2022:DIF

[Pab22] Fernando Pablos Romo. On G-Drazin inverses of finite potent endomorphisms and arbitrary square matrices. Linear Multilinear Algebra, 70(12):2227–2247, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Palmer:1979:SCM

[Pal79] Kenneth J. Palmer. A sufficient condition that a matrix have eigenvalues with non-zero real parts. Linear and Multilinear Algebra, 7(1):43–45, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pal:2021:QST

[Pal21] Hiranmoy Pal. Quantum state transfer on a class of circulant graphs. Linear Multilinear Algebra, 69(13):2527–2538, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pamfilos:1987:DIM

[Pam87] Paris Pamfilos. On the dimension of the image of a multilinear function. Linear and Multilinear Algebra, 20(4):333–336, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pankov:2013:CSS

[Pan13] Mark Pankov. Characterizations of strong semilinear embeddings in terms of general linear and projective linear groups. Linear Multilinear Algebra, 61(11):1555–1567, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pankov:2015:GCS

[Pan15] Mark Pankov. Geometrical characterization of semilinear isomorphisms of vector spaces and semilinear homeomorphisms of normed spaces. Linear Multilinear Algebra, 63(4):695–701, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Paniello:2017:BGI

[Pan17] I. Paniello. Backwards genetic inheritance through coalgebragraphs. Linear Multilinear Algebra, 65(5):943–961, 2017. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Paniello:2019:EC

[Pan19] I. Paniello. Evolution coalgebras. Linear Multilinear Algebra, 67(8):1539–1553, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Paniello:2020:ECC

[Pan20] I. Paniello. Evolution coalgebras on chicken populations. Linear Multilinear Algebra, 68(3):528–540, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Paniello:2022:MEA

[Pan22] Irene Paniello. Markov evolution algebras. Linear Multilinear Algebra, 70(19):4633–4653, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pappas:2013:RLC

[Pap13] Dimitrios Pappas. Restricted linear constrained minimization of quadratic functionals. Linear Multilinear Algebra, 61(10): 1394–1407, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Parrish:1978:MUC

[Par78] Christopher Parrish. Multivariate Umbral calculus. Linear and Multilinear Algebra, 6(2):93–109, 1978. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Passer:2015:AVJ

[Pas15] Benjamin Passer. An approximate version of the Jordan von Neumann theorem for finite-dimensional real normed spaces. Linear Multilinear Algebra, 63(1):68–77, 2015. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pasini:2016:EPG

[Pas16] Antonio Pasini. Equations for polar Grassmannians. Linear Multilinear Algebra, 64(11):2224–2240, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1977:ISC

[Pat77] Thomas H. Pate. Invariant subspaces for certain linear transformations on a space of tensors. Linear and Multilinear Algebra, 5(1):11–18, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1981:EII

[Pat81] T. H. Pate. An extension of an inequality involving symmetric products with an application to permanents. Linear and Multilinear Algebra, 10(2):103–105, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1982:IRG

[Pat82] Thomas H. Pate. Inequalities relating groups of diagonal products in a gram matrix. Linear and Multilinear Algebra, 11(1): 1–17, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1983:SMF

[Pat83] Thomas Pate. On those symmetric multilinear functions that are expressible as a linear combination of a small number of linearly independent powers. Linear and Multilinear Algebra, 12(4):265–272, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1984:HCF

[Pat84] Thomas H. Pate. Hypersurfaces containing flat subspaces. Linear and Multilinear Algebra, 16(1–4):253–259, 1984. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1986:AVS

[Pat86] Thomas H. Pate. Algebraic varieties in the symmetric algebra. Linear and Multilinear Algebra, 20(1):63–74, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1987:LBP

[Pat87] Thomas H. Pate. Lower bounds for polar and tensor ranks. Linear and Multilinear Algebra, 21(3):277–287, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1989:PDS

[Pat89] Thomas H. Pate. Permanental dominance and the Soules conjecture for certain right ideals in the group algebra. Linear and Multilinear Algebra, 24(2):135–149, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1990:IDT

[Pat90] Thomas H. Pate. Immanants and decomposable tensors that symmetrize to zero. Linear and Multilinear Algebra, 28(3): 175–184, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1992:TPS

[Pat92] Thomas H. Pate. Tensor products, symmetric products, and permanents of positive semi-definite Hermitian matrices. Linear and Multilinear Algebra, 31(1–4):27–36, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1993:PFI

[Pat93] Thomas H. Pate. Psi-functions, irreducible characters and a conjecture of Merris and Watkins. Linear and Multilinear Algebra, 35(3–4):195–212, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1995:HIS

[Pat95] Thomas H. Pate. Hook immanants and symmetry-antisymmetry modules. Linear and Multilinear Algebra, 38(4):295–307, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:1996:GAM

[Pat96] Thomas H. Pate. Group algebras, monotonicity, and the Lieb permanent inequality. Linear and Multilinear Algebra, 40(3): 207–220, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:2003:BLB

[Pat03a] Thomas H. Pate. The best lower bound for the permanent of a correlation matrix of rank two. Linear Multilinear Algebra, 51(3):263–278, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:2003:ERL

[Pat03b] Thomas H. Pate. Extreme rays for the large cone of Hermitian generalized matrix functions. Linear Multilinear Algebra, 51(4):361–391, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Patra:2007:MDB

[Pat07] K. L. Patra. Maximizing the distance between center, centroid and characteristic set of a tree. Linear Multilinear Algebra, 55(4):381–397, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:2008:GNI

[Pat08] Thomas H. Pate. A generalized Neuberger identity for the inner product of symmetric products of tensors. Linear Multilinear Algebra, 56(5):555–563, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:2009:NLB

[Pat09] Thomas H. Pate. A new lower bound inequality for the norm of the symmetric product. Linear Multilinear Algebra, 57(1): 87–102, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pate:2017:SFP

[Pat17] Thomas H. Pate. Subspace filters and polar bases for spaces of symmetric multilinear functions. Linear Multilinear Algebra, 65(10):2076–2100, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Paz:1984:ACH

[Paz84] Azaria Paz. An application of the Cayley–Hamilton theorem to matrix polynomials in several variables. Linear and Multilinear Algebra, 15(2):161–170, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pal:2017:PST

[PB17] Hiranmoy Pal and Bikash Bhattacharjya. Perfect state transfer on gcd-graphs. Linear Multilinear Algebra, 65(11):2245– 2256, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Panigrahy:2020:ROL

[PBM20] Krushnachandra Panigrahy, Ratikanta Behera, and Debasisha Mishra. Reverse-order law for the Moore–Penrose inverses of tensors. Linear Multilinear Algebra, 68(2):246–264, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pearl:1984:ATA

[Pea84] Martin H. Pearl. Automorphic transformations of an arbitrary matrix. Linear and Multilinear Algebra, 15(3–4):245– 256, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pena:2002:SRM

[Peñ02] J. M. Peña. Sign regular matrices of order two. Linear Multilinear Algebra, 50(1):91–97, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Peng:2015:SLS

[Pen15] Zhuohua Peng. The (R, S)-symmetric least squares solutions of the general coupled matrix equations. Linear Multilinear Algebra, 63(6):1086–1105, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Peperko:2011:FIS

[Pep11] Aljosa Peperko. On the functional inequality for the spectral radius of compact operators. Linear Multilinear Algebra, 59(4):357–364, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Peperko:2019:IJG

[Pep19a] Aljosa Peperko. Inequalities on the joint and generalized spectral and essential spectral radius of the Hadamard geometric mean of bounded sets of positive kernel operators. Linear Multilinear Algebra, 67(11):2159–2172, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Peperko:2019:ISR

[Pep19b] Aljosa Peperko. Inequalities on the spectral radius, operator norm and numerical radius of the Hadamard weighted geometric mean of positive kernel operators. Linear Multilinear Algebra, 67(8):1637–1652, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

PereiradaSilvaeSilva:2015:PPC

[Per15] Diogo Diniz Pereira da Silva e Silva. A primeness property for central polynomials of verbally prime P.I. algebras. Linear Multilinear Algebra, 63(11):2151–2158, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Peralta:2017:OFO

[Per17] Antonio M. Peralta. Orthogonal forms and orthogonality preservers on real function algebras revisited. Linear Multilinear Algebra, 65(2):361–374, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pervova:2019:DCA

[Per19] Ekaterina Pervova. Diffeological Clifford algebras and pseudobundles of Clifford modules. Linear Multilinear Algebra, 67(9): 1785–1828, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petek:1997:AMP

[Pet97] Tatjana Petek. Additive mappings preserving commutativity. Linear and Multilinear Algebra, 42(3):205–211, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petrychkovych:2000:GEP

[Pet00] Vasyl Petrychkovych. Generalized equivalence of pairs of matrices. Linear and Multilinear Algebra, 48(2):179–188, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petrovic:2009:BSM

[Pet09] Zoran Z. Petrović. Bases of spaces of matrices satisfying rank conditions. Linear Multilinear Algebra, 57(6):625–631, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petek:2010:MPI

[Pet10] Tatjana Petek. Mappings preserving the idempotency of products of operators. Linear Multilinear Algebra, 58(7):903–925, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pauwelyn:2022:PND

[PG22] P. J. Pauwelyn and M. A. Guerry. Perturbations of nondiagonalizable stochastic matrices with preservation of spectral properties. Linear Multilinear Algebra, 70(20):5115–5145, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pirzada:2022:SSG

[PGAB22] S. Pirzada, H. A. Ganie, A. Alhevaz, and M. Baghipur. On spectral spread of generalized distance matrix of a graph. Linear Multilinear Algebra, 70(15):2819–2835, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petrovic:2000:BGS

[PGLM00] Miroslav Petrović, Ivan Gutman, Mirko Lepović, and Bojana Milekić. On bipartite graphs with small number of Laplacian eigenvalues greater than two and three. Linear and Multilinear Algebra, 47(3):205–215, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Puystjens:1997:GIC

[PH97] R. Puystjens and R. E. Hartwig. The group inverse of a companion matrix. Linear and Multilinear Algebra, 43(1–3):137– 150, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pan:2003:TNC

[PH03] Yong-Liang Pan and Yao-Ping Hou. Two necessary conditions for λ2(G) = d2(G). Linear Multilinear Algebra, 51(1):31–38, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Patricio:2015:SRS

[PH15] P. Patrício and R. E. Hartwig. Some regular sums. Linear Multilinear Algebra, 63(1):185–200, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Phuong:1980:BPS

[Phu80] Vu Anh Phuong. Behavior of the permanent of a special class of doubly stochastic matrices. Linear and Multilinear Algebra, 9(3):227–229, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pierce:1974:AMF

[Pie74] Stephen Pierce. Anisotropic multilinear forms. Linear and Multilinear Algebra, 2(2):127–133, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Piercet:1979:DPL

[Pie79] Stephen Piercet. Discriminant preserving linear maps. Linear and Multilinear Algebra, 8(2):101–114, 1979. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pierce:1985:ESF

[Pie85] S. Pierce. Elementary symmetric functions. Linear and Multilinear Algebra, 17(3–4):375, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pierce:1992:BR

[Pie92a] Stephen Pierce. Book review. Linear and Multilinear Algebra, 33(3–4):305–306, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pierce:1992:CGI

[Pie92b] Stephen Pierce. Chapter 1: general introduction. Linear and Multilinear Algebra, 33(1–2):3–5, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pierce:1992:CAS

[Pie92c] Stephen Pierce. Chapter 4: algebraic sets, polynomials, and other functions. Linear and Multilinear Algebra, 33(1–2):31– 52, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Piff:1987:DDS

[Pif87] M. J. Piff. Diagonals of doubly stochastic and equiponderant matrices. Linear and Multilinear Algebra, 20(2):159–165, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pizarro:1987:BMV

[Piz87] Antonio Pizarro. Block matrices over a valuation ring. Linear and Multilinear Algebra, 21(3):239–241, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pavlikova:1990:IDI

[PJ90] Sonăa Pavlíková and Ján Kr’c Jediný. On the inverse and the dual index of a tree. Linear and Multilinear Algebra, 28(1–2): 93–109, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pan:2001:LBT

[PLH01] Yong-Liang Pan, Jiong-Sheng Li, and Yao-Ping Hou. Lower bounds on the third smallest Laplacian eigenvalue of a graph. Linear and Multilinear Algebra, 49(3):209–218, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pan:2020:NLS

[PLLL20] Yingui Pan, Jianping Li, Shuchao Li, and Wenjun Luo. On the normalized Laplacians with some classical parameters involving graph transformations. Linear Multilinear Algebra, 68(8): 1534–1556, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Peng:2019:NIM

[PLPC19] Jingjing Peng, Anping Liao, Zhenyun Peng, and Zhencheng Chen. Newton’s iterative method to solve a nonlinear matrix equation. Linear Multilinear Algebra, 67(9):1867–1878, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Porter:1978:RSS

[PM78] A. Duane Porter and Nick Mousouris. Ranked solutions of some matric equations. Linear and Multilinear Algebra, 6(2): 145–151, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Porter:1981:PSM

[PM81] A. Duane Porter and Nick Mousouris. Partitions of a symmetric matrix over a finite field. Linear and Multilinear Algebra, 10(4):329–341, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Porter:1985:PSM

[PM85] A. Duane Porter and Nick Mousouris. Partitions of a skew matrix over a finite field. Linear and Multilinear Algebra, 17(3– 4):309–323, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Prasad:2014:CEI

[PM14] K. Manjunatha Prasad and K. S. Mohana. Core-EP inverse. Linear Multilinear Algebra, 62(6):792–802, 2014. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Panigrahy:2022:EMP

[PM22] Krushnachandra Panigrahy and Debasisha Mishra. Extension of Moore–Penrose inverse of tensor via Einstein product. Linear Multilinear Algebra, 70(4):750–773, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Perovic:2023:QRP

[PM23] Vasilije Perović and D. Steven Mackey. Quadratic realizability of palindromic matrix polynomials: the real case. Linear Multilinear Algebra, 71(5):797–841, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Panigrahy:2022:TAE

[PMP22] Krushnachandra Panigrahy, Debasisha Mishra, and Juan Manuel Peña. On C-tensor and its application to eigenvalue localization. Linear Multilinear Algebra, 70(21):6279–6296, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Padhan:2022:CSM

[PN22] Rudra Narayan Padhan and Saudamini Nayak. On capability and the Schur multiplier of some nilpotent Lie superalgebras. Linear Multilinear Algebra, 70(8):1467–1478, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pokrzywa:1987:UIC

[Pok87] Andrzej Pokrzywa. Unitarily invariant class of matrices with powers without vanishing diagonal entries (with two figures). Linear and Multilinear Algebra, 21(4):373–384, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Poon:1980:APR

[Poo80a] Yiu-Tung Poon. Another proof of a result of Westwick. Linear and Multilinear Algebra, 9(1):35–37, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Poon:1980:GNR

[Poo80b] Yiu-Tung Poon. The generalized K — numerical range. Linear and Multilinear Algebra, 9(3):181–186, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Poon:1994:CHM

[Poo94] Yiu Tung Poon. On the convex hull of the multiform numerical range. Linear and Multilinear Algebra, 37(1–3):221–223, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Poon:2017:CMM

[Poo17] Edward Poon. Continuous multiplicative maps of Toeplitz algebras. Linear Multilinear Algebra, 65(10):2011–2023, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Popa:2014:SSC

[Pop14] N. Popa. Some special classes of infinite matrices. Linear Multilinear Algebra, 62(4):518–529, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Popa:2017:DLO

[Pop17] Dumitru Popa. Dominated n-linear operators with respect to a system of n bounded linear operators. Linear Multilinear Algebra, 65(6):1097–1107, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Popa:2020:EMF

[Pop20a] Dumitru Popa. An extension of the maurey factorization theorem. Linear Multilinear Algebra, 68(2):304–322, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Popa:2020:OMP

[Pop20b] Dumitru Popa. Operators with the Maurey–Pietsch multiple splitting property. Linear Multilinear Algebra, 68(6):1201– 1217, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Popa:2022:MVM

[Pop22] Dumitru Popa. Multilinear variants of the Maurey factorization theorem. Linear Multilinear Algebra, 70(19):3722–3733, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Powers:1988:SGT

[Pow88] David L. Powers. Semiregular graphs and their algebra. Linear and Multilinear Algebra, 24(1):27–37, 1988. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pozhidaev:2022:ASP

[Poz22] Aleksandr P. Pozhidaev. Ω-algebras with split products. Linear Multilinear Algebra, 70(16):3054–3069, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Panda:2016:SGW

[PP16] S. K. Panda and S. Pati. On some graphs which possess inverses. Linear Multilinear Algebra, 64(7):1445–1459, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pandey:2022:DCP

[PP22] D. Pandey and K. L. Patra. Different central parts of trees and their pairwise distances. Linear Multilinear Algebra, 70 (19):3790–3802, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pena:2008:ED

[PR08] Ismael Peña and Juan Rada. Energy of digraphs. Linear Multilinear Algebra, 56(5):565–579, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pellegrino:2012:ASP

[PR12] Daniel Pellegrino and Joilson Ribeiro. On almost summing polynomials and multilinear mappings. Linear Multilinear Algebra, 60(4):397–413, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petek:2020:NEP

[PR20] Tatjana Petek and Gordana Radić. A note on equivalence preserving maps. Linear Multilinear Algebra, 68(11):2289–2297, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pourgholamhossein:2022:FPN

[PRA22] Mahmood Pourgholamhossein, Mohammad Rouzbehani, and Massoud Amini. On finiteness properties of Noetherian (artinian) C^{∗}-algebras. Linear Multilinear Algebra, 70(3):419– 430, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pirzada:2022:DSL

[PRAC22] S. Pirzada, Bilal A. Rather, M. Aijaz, and T. A. Chishti. On distance signless Laplacian spectrum of graphs and spectrum of zero divisor graphs of Z_{n}. Linear Multilinear Algebra, 70 (17):3354–3369, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Preissmann:1985:NM

[Pre85] Emmanuel Preissmann. On the norm of the matrix (e(x_{r}y_{s})). Linear and Multilinear Algebra, 17(1):39–40, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Preusser:2022:GLG

[Pre22] Raimund Preusser. On general linear groups over exchange rings. Linear Multilinear Algebra, 70(4):705–713, 2022. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Price:1978:DDP

[Pri78] G. Baley Price. Determinants with dominant principal diagonal: a personal historical note. Linear and Multilinear Algebra, 6(4):247–250, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See comments [TT78].

Prokip:1995:MSN

[Pro95] V. Prokip. The multiplicativity of the Smith normal form. Linear and Multilinear Algebra, 38(3):189–192, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Protasov:2022:ACD

[Pro22] Vladimir Yu Protasov. Antinorms on cones: duality and applications. Linear Multilinear Algebra, 70(22):7387–7413, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petrovic:2003:GCS

[PRS03] M. Petrović, Z. Radosavljević, and S. Simić. A graph and its complement with specified spectral properties. Linear Multi-

Przeslawski:1992:LAC

[Prz92] Krzysztof Przeslawski. Linear algebra of convex sets and the Euler characteristic. Linear and Multilinear Algebra, 31(1–4): 153–191, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Poljak:1991:CGJ

[PS91] Svatopluk Poljak and Milos Schlegel. Computing the generic Jordan canonical form. Linear and Multilinear Algebra, 28(4): 241–249, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Patra:2020:SFS

[PS20] Arnab Patra and P. D. Srivastava. Spectrum and fine spectrum of generalized lower triangular triple band matrices over the sequence space l_{p}. Linear Multilinear Algebra, 68(9):1797– 1816, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petrullo:2021:ERG

[PS21] Pasquale Petrullo and Domenico Senato. Endomorphisms of the Riordan group. Linear Multilinear Algebra, 69(9):1683– 1693, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Parulekar:2022:SRR

[PS22] Tushar D. Parulekar and Sharad S. Sane. Some results on the Ryser design conjecture — II. Linear Multilinear Algebra, 70 (1):81–90, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pal:2014:GOV

[PSA14] Rajinder Pal, Mandeep Singh, and Jaspal Singh Aujla. Generalized operator version of Bernoulli’s inequality. Linear Multilinear Algebra, 62(2):267–273, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pirzada:2023:NCW

[PSB23] S. Pirzada, Tahir Shamsher, and Mushtaq A. Bhat. A note on the class of weighted graphs with strong anti-reciprocal eigenvalue property. Linear Multilinear Algebra, 71(1):1–4, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pal:2016:NCO

[PSMA16] Rajinder Pal, Mandeep Singh, Mohammad Sal Moslehian, and Jaspal Singh Aujla. A new class of operator monotone functions via operator means. Linear Multilinear Algebra, 64(12): 2463–2473, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pellegrino:2015:GTA

[PSS15] D. Pellegrino and J. B. Seoane-Sepúlveda. Grothendieck’s theorem for absolutely summing multilinear operators is optimal. Linear Multilinear Algebra, 63(3):554–558, 2015. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Panario:2018:NHC

[PSW18] Daniel Panario, Murat Sahin, and Qiang Wang. Nonhomogeneous conditional recurrences. Linear Multilinear Algebra, 66(10):2089–2099, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Poon:1987:IRB

[PT87] Yiu-Tung Poon and Nam-Kiu Tsing. Inclusion relations between orthostochastic matrices and products of pinching matrices. Linear and Multilinear Algebra, 21(3):253–259, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Psarrakos:2002:SRM

[PT02] P. J. Psarrakos and M. J. Tsatsomeros. On the stability radius of matrix polynomials. Linear Multilinear Algebra, 50(2):151– 165, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Piccione:2010:ATG

[PT10] Paolo Piccione and Daniel V. Tausk. An algebraic theory for generalized Jordan chains and partial signatures in the Lagrangian Grassmannian. Linear Multilinear Algebra, 58(1): 89–103, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pullman:1975:NTM

[Pul75] Norman J. Pullman. A note on a theorem of Minc on irreducible non-negative matrices. Linear and Multilinear Algebra, 2(4):335–336, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Petik:2015:GQL

[PUÖ15] T. Petik, M. Uç, and H. Özdemir. Generalized quadraticity of linear combination of two generalized quadratic matrices. Linear Multilinear Algebra, 63(12):2430–2439, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Putcha:1982:GRC

[Put82a] Mohan S. Putcha. Green’s relations on a connected algebraic monoid. Linear and Multilinear Algebra, 12(3):205–214, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Putcha:1982:GUC

[Put82b] Mohan S. Putcha. The group of units of a connected algebraic monoid. Linear and Multilinear Algebra, 12(1):37–50, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Psarrakos:2000:NRM

[PV00] Panayiotis J. Psarrakos and Panayiotis M. Vilamos. The qnumerical range of matrix polynomials. Linear and Multilinear Algebra, 47(1):1–9, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Prakash:2020:CMC

[PV20] Anupam Prakash and Antonios Varvitsiotis. Correlation matrices, Clifford algebras, and completely positive semidefinite rank. Linear Multilinear Algebra, 68(5):1039–1056, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Podesta:2023:RGE

[PV23] Ricardo A. Podestá and Denis E. Videla. On regular graphs equienergetic with their complements. Linear Multilinear Algebra, 71(3):422–456, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Pierce:1978:ILM

[PW78] Stephen Pierce and William Watkins. Invariants of linear maps on matrix algebras. Linear and Multilinear Algebra, 6(3):185– 200, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Putcha:1982:MSC

[PY82] Mohan S. Putcha and Adil Yaqub. Matrix semigroups closed under scalar translations. Linear and Multilinear Algebra, 11(2):183–187, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Peng:2022:UEM

[PZGL22] Xiao-Song Peng, Yi Zhang, Xing Gao, and Yan-Feng Luo. Universal enveloping of (modified) λ-differential Lie algebras. Linear Multilinear Algebra, 70(6):1102–1127, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qin:2023:NMP

[QC23] Zijie Qin and Chaoqun Chen. Nonlinear maps preserving similarity on B(X). Linear Multilinear Algebra, 71(3):363–376, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Queiro:1985:CDM

[QD85] Joáo Filipe Queiró and António Leal Duarte. On the Cartesian decomposition of a matrix. Linear and Multilinear Algebra, 18 (1):77–85, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qi:2011:ALM

[QH11] Xiaofei Qi and Jinchuan Hou. Additivity of Lie multiplicative maps on triangular algebras. Linear Multilinear Algebra, 59(4):391–397, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qi:2017:SCP

[QH17] Xiaofei Qi and Jinchuan Hou. Strong 3-commutativity preserving maps on prime rings. Linear Multilinear Algebra, 65 (11):2153–2169, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qiao:2022:NRN

[QHB22] Hongwei Qiao, Guojun Hai, and Eburilitu Bai. A-numerical radius and A-norm inequalities for semi-Hilbertian space operators. Linear Multilinear Algebra, 70(21):6891–6907, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qi:2015:CLD

[Qi15] Xiaofei Qi. Characterizing Lie n-derivations for reflexive algebras. Linear Multilinear Algebra, 63(8):1693–1706, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Queiro:1992:BDS

[QK92] João Filipe Queiró and Alexander Kovacec. A bound for the determinant of the sum of two normal matrices. Linear and Multilinear Algebra, 33(3–4):171–173, 1992. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qi:2022:OTM

[QLK22] Yaru Qi, Yuying Li, and Yihui Kong. The Ostrowski theorem for matrices of operators. Linear Multilinear Algebra, 70(22): 7339–7346, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qaralleh:2021:VEA

[QM21] Izzat Qaralleh and Farrukh Mukhamedov. Volterra evolution algebras and their graphs. Linear Multilinear Algebra, 69(12): 2228–2244, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Queiro:2008:CNP

[Que08] João Filipe Queiró. Composite norms and perfect conditioning. Linear Multilinear Algebra, 56(5):589–596, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qin:2023:SLG

[QX23] Zhengxin Qin and Xinping Xu. The spectrum of a line graph of a Cayley graph over an abelian group. Linear Multilinear Algebra, 71(13):2204–2211, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Qi:2016:MJD

[QZ16] Xiaofei Qi and Feifei Zhang. Multiplicative Jordan *- derivations on rings with involution. Linear Multilinear Algebra, 64(6):1145–1162, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rajkumar:2021:LSR

[RA21] R. Rajkumar and T. Anitha. Laplacian spectrum of reduced power graph of certain finite groups. Linear Multilinear Algebra, 69(9):1716–1733, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Radjavi:1980:PBC

[Rad80] Heydar Radjavi. Proof of Ballantine’s conjecture on simplic cosquares. Linear and Multilinear Algebra, 9(3):193–194, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rada:2012:BEN

[Rad12] Juan Rada. Bounds for the energy of normal digraphs. Linear Multilinear Algebra, 60(3):323–332, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Radenkovic:2016:ROL

[Rad16] Jovana Nikolov Radenković. Reverse order law for generalized inverses of multiple operator product. Linear Multilinear Algebra, 64(7):1266–1282, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rais:1987:UGP

[Rai87] M. Rais. The unitary group preserving maps (the infinite dimensional case). Linear and Multilinear Algebra, 20(4):337– 345, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Raissouli:2014:CSS

[Raï14a] Mustapha Raïssouli. Corrigendum: Stable and stabilizable means involving linear operator arguments. Linear Multilinear Algebra, 62(9):i, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Raï14b].

Raissouli:2014:SSM

[Raï14b] Mustapha Raïssouli. Stable and stabilizable means involving linear operator arguments. Linear Multilinear Algebra, 62(9): 1153–1168, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [Raï14a].

Raissouli:2019:SII

[Raï19] Mustapha Raïssouli. Some inequalities involving quadratic forms of operator means. Linear Multilinear Algebra, 67(2): 213–220, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rajeswari:2009:UBS

[Raj09] K. N. Rajeswari. An upper bound on separable index of certain SL(3)-modules. Linear Multilinear Algebra, 57(6):633– 640, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ramer:2002:LE

[Ram02] Arthur Ramer. Letter to the Editor. Linear Multilinear Algebra, 50(1):99–100, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ramezani:2022:SRS

[Ram22] Farzaneh Ramezani. Some regular signed graphs with only two distinct eigenvalues. Linear Multilinear Algebra, 70(3): 517–530, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rezagholibeigi:2022:SCU

[RAN22] Meysam Rezagholibeigi, Ghodratollah Aalipour, and Ali Reza Naghipour. On the spectrum of the closed unit graphs. Linear Multilinear Algebra, 70(10):1871–1885, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rao:1981:GIM

[Rao81] K. P. S. Bhaskara Rao. On generalized inverses of matrices over principal ideal rings. Linear and Multilinear Algebra, 10(2):145–154, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rao:2018:AOS

[Rao18] T. S. S. R. K. Rao. Adjoints of operators as smooth points in spaces of compact operators. Linear Multilinear Algebra, 66(4):668–670, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rashid:2017:QFP

[Ras17] M. H. M. Rashid. Quasinormality and Fuglede–Putnam theorem for (s, p) − w-hyponormal operators. Linear Multilinear Algebra, 65(8):1600–1616, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rashid:2019:SIS

[Ras19] M. H. M. Rashid. Some invariant subspaces for w-hyponormal operators. Linear Multilinear Algebra, 67(7):1460–1470, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rau:2018:PMN

[Rau18] Christian Rau. On Procrustes matching of non-negative matrices and an application to random tomography. Linear Multilinear Algebra, 66(1):96–106, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rawlings:1981:SNP

[Raw81] Don Rawlings. The (qr)-Simon Newcomb problem. Linear and Multilinear Algebra, 10(3):253–260, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Roy:2022:OSF

[RB22] Saikat Roy and Satya Bagchi. Orthogonality of sesquilinear forms and spaces of operators. Linear Multilinear Algebra, 70 (19):4416–4424, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodriguez:2022:PAP

[RBB22] José L. Vilca Rodríguez, Tauan L. A. Brandão, and Victor M. O. Batista. Principal angles and pairs of totally geodesic submanifolds of the real hyperbolic space. Linear Multilinear Algebra, 70(20):5932–5947, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ramamurthy:2023:RMW

[RBG23] Balaji Ramamurthy, Ravindra B. Bapat, and Shivani Goel. On resistance matrices of weighted balanced digraphs. Linear Multilinear Algebra, 71(13):2222–2248, 2023. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodtes:2018:SCT

[RC18] Kijti Rodtes and Kunlathida Chimla. Symmetry classes of tensors and semi-direct product of finite abelian groups. Linear Multilinear Algebra, 66(11):2191–2198, 2018. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Reis:2022:DEA

[RC22] Tiago Reis and Paula Cadavid. Derivations of evolution algebras associated to graphs over a field of any characteristic. Linear Multilinear Algebra, 70(15):2884–2897, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Roberts:2012:FRS

[RD12] Mark L. Roberts and Michael P. Drazin. Faithfully ⋆- representing semigroups and groupoids with involution by infinite matrices. Linear Multilinear Algebra, 60(5):511–524, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodrigo-Escudero:2022:FGA

[RE22] Adrián Rodrigo-Escudero. Fine gradings and automorphism groups on associative algebras. Linear Multilinear Algebra, 70 (17):3402–3418, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Reams:1996:INM

[Rea96] Robert Reams. An inequality for nonnegative matrices and the inverse eigenvalue problem. Linear and Multilinear Algebra, 41(4):367–375, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Redmond:2002:PPR

[Red02] Timothy Redmond. Proving polynomial Riemann hypotheses through eigenfunctions. Linear Multilinear Algebra, 50(1):49– 59, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Regev:1987:CPI

[Reg87a] Amitai Regev. A combinatorial proof for an identity of Procesi of S_{n} characters. Linear and Multilinear Algebra, 21(1):29–39, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Regev:1987:STI

[Reg87b] Amitai Regev. Sign trace identities. Linear and Multilinear Algebra, 21(1):1–28, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Regev:1989:PFI

[Reg89] Amitai Regev. From polynomial to formal identities. Linear and Multilinear Algebra, 24(4):231–234, 1989. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Regev:1993:CIS

[Reg93] Amitai Regev. Capelli identities for some modular P.I. algebras. Linear and Multilinear Algebra, 35(2):131–134, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodriguez:2007:FCG

[REG07] J. A. Rodríguez, E. Estrada, and A. Gutiérrez. Functional centrality in graphs. Linear Multilinear Algebra, 55(3):293– 302, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Reiner:1975:PCG

[Rei75] Irving Reiner. Projective class groups of symmetric and alternating groups. Linear and Multilinear Algebra, 3(1–2):147– 153, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Reiner:1977:SPF

[Rei77] D. L. Reiner. Sequences of polynomials of fractional binomial type. Linear and Multilinear Algebra, 5(3):175–179, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Reich:1983:LTP

[Rei83] Simeon Reich. A limit theorem for projections. Linear and Multilinear Algebra, 13(3):281–290, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Remmel:1982:BPF

[Rem82] J. B. Remmel. Bijective proofs of formulae for the number of standard yound tableaux. Linear and Multilinear Algebra, 11 (1):45–100, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Remmel:1988:CCS

[Rem88] J. B. Remmel. Computing cocharacters of sign trace identities. Linear and Multilinear Algebra, 23(1):1–14, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Remmel:1990:BPF

[Rem90] J. B. Remmel. A bijective proof of a factorization theorem for (k, l)-hook Schur functions. Linear and Multilinear Algebra, 28(3):119–154, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Revoy:1981:RFT

[Rev81] Philippe Revoy. Remarques sur la forme trace. Linear and Multilinear Algebra, 10(3):223–233, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Roopaei:2019:NMO

[RF19a] H. Roopaei and D. Foroutannia. The norm of matrix operators on Cesàro weighted sequence space. Linear Multilinear Algebra, 67(1):175–185, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Roopaei:2019:NCM

[RF19b] H. Roopaei and D. Foroutannia. The norms of certain matrix operators from l_{p} spaces into l_{p}(Δ^{n}) spaces. Linear Multilinear Algebra, 67(4):767–776, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rakhimov:2011:CSL

[RH11a] I. S. Rakhimov and S. K. Said Husain. Classification of a subclass of low-dimensional complex filiform Leibniz algebras.

Rakhimov:2011:ICI

[RH11b] I. S. Rakhimov and S. K. Said Husain. On isomorphism classes and invariants of a subclass of low-dimensional complex filiform Leibniz algebras. Linear Multilinear Algebra, 59(2):205– 220, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rhoades:1987:LBS

[Rho87] B. E. Rhoades. Lower bounds for some matrices. Linear and Multilinear Algebra, 20(4):347–352, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rhoades:1990:LBS

[Rho90] B. C. Rhoades. Lower bounds for some matrices, II. Linear and Multilinear Algebra, 26(1–2):49–58, 1990. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rhoades:1994:SUB

[Rho94] B. E. Rhoades. Sharp upper bounds for the norms of certain bounded operators on l^{p}. Linear and Multilinear Algebra, 36(4):251–253, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Richman:1978:SGL

[Ric78] Daniel J. Richman. The singular graph of lower triangular, nilpotent matrices. Linear and Multilinear Algebra, 6(1):37– 49, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Richman:1981:MWC

[Ric81] David Ross Richman. Matrices which commute with their transposes over a field of characteristic two. Linear and Multilinear Algebra, 10(2):131–140, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Richman:1985:MSS

[Ric85] David R. Richman. Matrices as sums of squares: a conjecture of Griffin and Krusemeyer. Linear and Multilinear Algebra, 17(3–4):289–294, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See addendum [Ric87a].

Richman:1987:AMS

[Ric87a] David R. Richman. Addendum to “Matrices as sums of squares: a conjecture of Griffin and Krusemeyer”. Linear and Multilinear Algebra, 20(4):389, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Ric85].

Richman:1987:RAB

[Ric87b] David R. Richman. A result about block Vandermonde matrices. Linear and Multilinear Algebra, 21(2):181–189, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ren:2004:CSS

[RJ04] Bin Ren and Guoxing Ji. Certain structure of similaritypreserving linear maps. Linear Multilinear Algebra, 52(1): 61–68, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rohn:1994:LII

[RK94] Jiří Rohn and Jana Kreslová. Linear interval inequalities. Linear and Multilinear Algebra, 38(1–2):79–82, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rattanakangwanwong:2022:EZD

[RM22] Jitsupat Rattanakangwanwong and Yotsanan Meemark. Eigenvalues of zero divisor graphs of principal ideal rings. Linear Multilinear Algebra, 70(20):5445–5459, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Robinson:1975:LBD

[Rob75] Donald W. Robinson. Lower bounds on the dimension of the space of matrices transposed by a similarity transformation. Linear and Multilinear Algebra, 3(1–2):135–141, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Robinson:1977:MZT

[Rob77a] Donald W. Robinson. Matrices with zero trace as commutators of nilpotents. Linear and Multilinear Algebra, 5(1):45–51, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Robinson:1977:QFS

[Rob77b] Herbert A. Robinson. Quadratic forms on symmetry classes of tensors. Linear and Multilinear Algebra, 4(4):233–241, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodman:1981:NIS

[Rod81] Leiba Rodman. On nonnegative invariant subspaces in indefinite scalar product spaces. Linear and Multilinear Algebra, 10 (1):1–14, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodman:1997:PSS

[Rod97] Leiba Rodman. Products of symmetric and skew-symmetric matrices. Linear and Multilinear Algebra, 43(1–3):19–34, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodriguez:2002:LEM

[Rod02] J. A. Rodríguez. On the Laplacian eigenvalues and metric parameters of hypergraphs. Linear Multilinear Algebra, 50 (1):1–14, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodriguez:2003:LSW

[Rod03] J. A. Rodríguez. On the Laplacian spectrum and walk-regular hypergraphs. Linear Multilinear Algebra, 51(3):285–297, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodman:2008:PQS

[Rod08] Leiba Rodman. Pairs of quaternionic selfadjoint operators with numerical range in a halfplane. Linear Multilinear Algebra, 56(1–2):179–184, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodriguez:2022:FTH

[Rod22] R. González Rodríguez. The fundamental theorem of Hopf modules for Hopf braces. Linear Multilinear Algebra, 70(20): 5146–5156, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rohn:1991:ETS

[Roh91a] Jiri Rohn. An existence theorem for systems of linear equations. Linear and Multilinear Algebra, 29(2):141–144, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rohn:1991:TM

[Roh91b] Jiri Rohn. A theorem on p-matrices. Linear and Multilinear Algebra, 30(3):209–211, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rohn:2000:CNN

[Roh00] Jirí Rohn. Computing the norm ||A||_{∞}1 is NP-hard. Linear and Multilinear Algebra, 47(3):195–204, 2000. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rohn:2004:TAE

[Roh04] Jiri Rohn. A theorem of the alternatives for the equation Ax + B|x| = b. Linear Multilinear Algebra, 52(6):421–426, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rohn:2006:PVI

[Roh06] Jiri Rohn. Perron vectors of an irreducible nonnegative interval matrix. Linear Multilinear Algebra, 54(6):399–404, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rohn:2013:LE

[Roh13] Jiri Rohn. Letter to the editor. Linear Multilinear Algebra, 61(5):697–698, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Roitman:1999:SPJ

[Roi99] Moshe Roitman. A short proof of the Jordan decomposition theorem. Linear and Multilinear Algebra, 46(3):245–247, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Roman:1982:OF

[Rom82] Steven Roman. Operational formulas. Linear and Multilinear Algebra, 12(1):1–20, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Romdhani:1989:CRC

[Rom89] Mustapha Romdhani. Classification of real and complex nilpotent Lie algebras of dimension 7. Linear and Multilinear Algebra, 24(3):167–189, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Romo:2007:LPT

[Rom07] Fernando Pablos Romo. On the linearity property of Tate’s trace. Linear Multilinear Algebra, 55(4):323–326, 2007. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Romo:2019:DIF

[Rom19] Fernando Pablos Romo. On the Drazin inverse of finite potent endomorphisms. Linear Multilinear Algebra, 67(10):2135– 2146, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Romo:2020:CND

[Rom20] Fernando Pablos Romo. Core-nilpotent decomposition and new generalized inverses of finite potent endomorphisms. Linear Multilinear Algebra, 68(11):2254–2275, 2020. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Romo:2021:DMP

[Rom21] Fernando Pablos Romo. On Drazin-Moore–Penrose inverses of finite potent endomorphisms. Linear Multilinear Algebra, 69(4):627–647, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Romo:2022:ESN

[Rom22a] Fernando Pablos Romo. Explicit solutions of non-homogeneous difference equations from finite potent endomorphisms. Linear Multilinear Algebra, 70(20):5346–5361, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Romo:2022:CDI

[Rom22b] Fernando Pablos Romo. On the coincidence of the Drazin inverse and the Drazin–Moore–Penrose inverses. Linear Multilinear Algebra, 70(19):3617–3632, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Roopaei:2020:FCH

[Roo20] H. Roopaei. Factorization of Cesàro and Hilbert matrices based on generalized Cesàro matrix. Linear Multilinear Algebra, 68(1):193–204, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rousseau:1998:SCN

[Rou98] G. Rousseau. A sufficient condition for a non-negative matrix to have non-negative determinant. Linear and Multilinear Algebra, 44(2):179–185, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rowlinson:1995:I

[Row95] P. Rowlinson. Introduction. Linear and Multilinear Algebra, 39(1–2):1, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ruzieh:1990:DSP

[RP90] Subhi N. Ruzieh and David L. Powers. The distance spectrum of the path P_{n} and the first distance eigenvector of connected graphs. Linear and Multilinear Algebra, 28(1–2):75–81, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodman:2004:RPL

[RP04] Leiba Rodman and Stephen J. Pierce. Research problem: Lipschitz properties of matrix group actions. Linear Multilinear Algebra, 52(6):461–466, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rajkumar:2022:SRP

[RP22] R. Rajkumar and R. Pavithra. Spectra of M-rooted product of graphs. Linear Multilinear Algebra, 70(1):1–26, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ringh:2022:FGS

[RQ22] A. Ringh and L. Qiu. Finsler geometries on strictly accretive matrices. Linear Multilinear Algebra, 70(21):6753–6771, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ran:1991:SFR

[RR91] André C. M. Ran and Leiba Rodman. On symmetric factorizations of rational matrix functions. Linear and Multilinear Algebra, 29(3–4):243–261, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rex:1995:NCR

[RR95] G. Rex and J. Rohn. A note on checking regularity of interval matrices. Linear and Multilinear Algebra, 39(3):259–262, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Radjavi:2002:CI

[RR02] Heydar Radjavi and Peter Rosenthal. On commutators of idempotents. Linear Multilinear Algebra, 50(2):121–124, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Radjabalipour:2004:RMD

[RR04] Mehdi Radjabalipour and Heydar Radjavi. On rings of matrices over division rings. Linear Multilinear Algebra, 52(5): 375–380, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rebei:2023:HDL

[RRC23] Habib Rebei, Anis Riahi, and Wathek Chammam. Hankel determinant of linear combination of the shifted Catalan numbers. Linear Multilinear Algebra, 71(11):1829–1848, 2023. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ramesh:2022:CDN

[RRN22] G. Ramesh, B. Sudip Ranjan, and D. Venku Naidu. Cartesian decomposition of C-normal operators. Linear Multilinear Algebra, 70(21):6640–6647, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ramesh:2023:RCN

[RRN23] G. Ramesh, B. Sudip Ranjan, and D. Venku Naidu. A representation of compact C-normal operators. Linear Multilinear Algebra, 71(9):1565–1577, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ran:1993:SII

[RRR93] André C. M. Ran, Leiba Rodman, and Arthur L. Rubin. Stability index of invariant subspaced of matrices. Linear and Multilinear Algebra, 36(1):27–39, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Richard:2022:BHT

[RRR^{+}22] Logan Richard, Franziska Riepl, Leonel Robert, George Turcu, and Lucas Walls. Building homotheties through pinches. Linear Multilinear Algebra, 70(22):7469–7483, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Richman:1974:PSN

[RS74] Daniel J. Richman and Hans Scheuder. Primes in the semigroup of non-negative matrices. Linear and Multilinear Algebra, 2(2):135–140, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ramanujan:1985:OIS

[RS85] M. S. Ramanujan and E. Schock. Operator ideals and spaces of bilinear operators. Linear and Multilinear Algebra, 18(4): 307–318, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodman:2006:OPB

[RŠ06] Leiba Rodman and Peter Šemrl. Orthogonality preserving bijective maps on real and complex projective spaces. Linear Multilinear Algebra, 54(5):355–367, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Radjavi:2007:SGU

[RS07] H. Radjavi and A. R. Sourour. The semigroup generated by a unitary orbit of a singular matrix. Linear Multilinear Algebra, 55(5):417–428, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodman:2008:OPB

[RŠ08a] Leiba Rodman and Peter Šemrl. Orthogonality preserving bijective maps on finite dimensional projective spaces over division rings. Linear Multilinear Algebra, 56(6):647–664, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rose:2008:NRB

[RS08b] David E. V. Rose and Ilya M. Spitkovsky. On the numerical range behavior under the generalized Aluthge transform. Linear Multilinear Algebra, 56(1–2):163–177, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rodman:2009:PGO

[RŠ09] L. Rodman and P. Šemrl. Preservers of generalized orthogonality on finite dimensional real vector and projective spaces. Linear Multilinear Algebra, 57(8):839–858, 2009. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rhoades:2010:KLI

[RS10] Brendon Rhoades and Mark Skandera. Kazhdan–Lusztig immanants and products of matrix minors, II. Linear Multilinear Algebra, 58(2):137–150, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Robeva:2017:SVO

[RS17] Elina Robeva and Anna Seigal. Singular vectors of orthogonally decomposable tensors. Linear Multilinear Algebra, 65 (12):2457–2471, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ramirez:2018:GSM

[RS18] José L. Ramírez and Víctor F. Sirvent. Generalized Schröder matrix and its combinatorial interpretation. Linear Multilinear Algebra, 66(2):418–433, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Roy:2023:AOC

[RS23] Saikat Roy and Debmalya Sain. Approximate orthogonality in complex normed spaces: a directional approach. Linear Multilinear Algebra, 71(5):711–727, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rout:2021:SNR

[RSM21] Nirmal Chandra Rout, Satyajit Sahoo, and Debasisha Mishra. Some A-numerical radius inequalities for semi-Hilbertian space operators. Linear Multilinear Algebra, 69(5):980–996, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rout:2022:NRI

[RSM22] Nirmal Chandra Rout, Satyajit Sahoo, and Debasisha Mishra. On A-numerical radius inequalities for 2×2 operator matrices. Linear Multilinear Algebra, 70(14):2672–2692, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rueda:2017:TTN

[RSPT17] P. Rueda, E. A. Sánchez-Pérez, and A. Tallab. Traced tensor norms and multiple summing multilinear operators. Linear Multilinear Algebra, 65(4):768–786, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ribeiro:2022:CCS

[RST22] Joilson Ribeiro, Fabricio Santos, and Ewerton R. Torres. Coherence and compatibility: a stronger approach. Linear Multilinear Algebra, 70(1):66–80, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Radjavi:2009:SSO

[RT09] Heydar Radjavi and Vladimir G. Troitsky. Semitransitive spaces of operators. Linear Multilinear Algebra, 57(1):1–11, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rublein:1989:CCJ

[Rub89] G. T. Rublein. On a conjecture of C. Johnson. Linear and Multilinear Algebra, 25(4):257–267, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rubei:2020:GRT

[Rub20] Elena Rubei. A generalization of Rohn’s theorem on full-rank interval matrices. Linear Multilinear Algebra, 68(5):931–939, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ruch:1989:CCB

[Ruc89] David Ruch. Characterizations of compact bilinear maps. Linear and Multilinear Algebra, 25(4):297–307, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rump:1997:ASB

[Rum97] S. M. Rump. Almost sharp bounds for the componentwise distance to the nearest singular matrix. Linear and Multilinear Algebra, 42(2):93–107, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rusek:2016:LTQ

[Rus16] Kamil Rusek. Linearly triangularizable quadratic endomorphisms. Linear Multilinear Algebra, 64(8):1527–1537, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Remmel:1984:BPG

[RW84] Jeffrey B. Remmel and Roger Whitney. A bijective proof of the generating function for the number of reverse plane partitions via lattice paths. Linear and Multilinear Algebra, 16(1–4): 75–91, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Remmel:1996:TMK

[RW96] Jeffrey B. Remmel and Tamsen Whitehead. Transition matrices and Kronecker product expansions of symmetric functions. Linear and Multilinear Algebra, 40(4):337–352, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rosa:2021:LRV

[RW21] Kennett L. Dela Rosa and Hugo J. Woerdeman. Location of Ritz values in the numerical range of normal matrices. Linear Multilinear Algebra, 69(15):2749–2778, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ransford:2022:SSE

[RW22] Thomas Ransford and Nathan Walsh. Spectral sets, extremal functions and exceptional matrices. Linear Multilinear Algebra, 70(17):3310–3318, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Robinson:1973:BR

[RY73] Donald W. Robinson and Adil Yaqub. Book reviews. Linear and Multilinear Algebra, 1(3):275–277, 1973. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Rowlinson:1993:THG

[RY93] Peter Rowlinson and Yang Yuansheng. Tricylic Hamiltonian graphs with minimal index. Linear and Multilinear Algebra, 34(3–4):187–196, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ryser:1973:IIM

[Rys73] H. J. Ryser. Indeterminates and incidence matrices. Linear and Multilinear Algebra, 1(2):149–157, 1973. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ryser:1975:FIM

[Rys75] H. J. Ryser. The formal incidence matrix. Linear and Multilinear Algebra, 3(1–2):99–104, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Repovs:2017:IGS

[RZ17] Dusan Repovs and Mikhail Zaicev. Identities of graded simple algebras. Linear Multilinear Algebra, 65(1):44–57, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shively:1977:BR

[SA77] Ralph L. Shively and Charles A. Akemann. Book reviews. Linear and Multilinear Algebra, 4(4):307–309, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Salemi:2007:PNH

[SA07] A. Salemi and Gh. R. Aghamollaei. Polynomial numerical hulls of matrix polynomials. Linear Multilinear Algebra, 55(3): 219–228, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shebrawi:2010:NIA

[SA10] Khalid Shebrawi and Hussien Albadawi. Norm inequalities for the absolute value of Hilbert space operators. Linear Multilin-

Soleymani:2014:LPE

[SA14] M. Soleymani and A. Armandnejad. Linear preservers of even majorization on m_{n,m}. Linear Multilinear Algebra, 62(11): 1437–1449, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shamsabadi:2017:IFF

[SA17] Mitra Shamsabadi and Ali Akbar Arefijamaal. The invertibility of fusion frame multipliers. Linear Multilinear Algebra, 65(5):1062–1072, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shamsabadi:2021:CFT

[SA21] Mitra Shamsabadi and Ali Akbar Arefijamaal. On C–D controlled frames and their duality. Linear Multilinear Algebra, 69(9):1761–1770, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sababheh:2019:HTN

[Sab19] M. Sababheh. Heinz-type numerical radii inequalities. Linear Multilinear Algebra, 67(5):953–964, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sadkane:2023:SSV

[Sad23] Miloud Sadkane. A smallest singular value method for nonlinear eigenvalue problems. Linear Multilinear Algebra, 71(1): 16–28, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sahoo:2021:CAS

[Sah21] Gopinath Sahoo. Complex adjacency spectra of digraphs. Linear Multilinear Algebra, 69(2):193–207, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sain:2021:SNA

[Sai21] Debmalya Sain. Smoothness and norm attainment of bounded bilinear operators between Banach spaces. Linear Multilinear Algebra, 69(16):3010–3018, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Singh:2022:AMG

[SAK22] Jitendra Kumar Singh, Purshottam Narain Agrawal, and Arun Kajla. Approximation by modified q-gamma type operators via A-statistical convergence and power series method. Linear Multilinear Algebra, 70(21):6548–6567, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Salazarf:1979:AMM

[Sal79] Tomas Eufracio Salazarf. An abstract matrix model:the column model. Linear and Multilinear Algebra, 8(2):153–170, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Salem:2012:GBM

[Sal12] Ahmed Salem. On a q-gamma and a q-beta matrix functions. Linear Multilinear Algebra, 60(6):683–696, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Salem:2014:BGH

[Sal14] Ahmed Salem. The basic Gauss hypergeometric matrix function and its matrix q-difference equation. Linear Multilinear Algebra, 62(3):347–361, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Samelson:1978:NLH

[Sam78] Hans Samelson. Note on the Lefschetz–Hodge operators. Linear and Multilinear Algebra, 6(4):253–255, 1978. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sano:2014:NDM

[San14] Mari Sano. A note on determinants, minors and a boundary map. Linear Multilinear Algebra, 62(4):474–485, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sandovici:2018:NTL

[San18] Adrian Sandovici. Von Neumann’s theorem for linear relations. Linear Multilinear Algebra, 66(9):1750–1756, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sarma:2020:MHQ

[Sar20] Deepak Sarma. Minors of Hermitian (quasi-) Laplacian matrix of a mixed graph. Linear Multilinear Algebra, 68(11):2157– 2169, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sato:2011:BZF

[Sat11] Iwao Sato. Bartholdi zeta functions of periodic graphs. Linear Multilinear Algebra, 59(1):11–24, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Singh:2001:IHP

[SAV01] Mandeep Singh, Jaspal Singh Aujla, and H. L. Vasudeva. Inequalities for Hadamard product and unitarily invariant norms of matrices. Linear and Multilinear Algebra, 48(3):247–262, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seidenberg:1975:BR

[SB75] A. Seidenberg and M. Boon. Book reviews. Linear and Multilinear Algebra, 3(3):245–247, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2003:LOP

[SB03] Seok-Zun Song and LeRoy B. Beasley. Linear operators that preserve commuting pairs of nonnegative real matrices. Linear Multilinear Algebra, 51(3):279–283, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2016:PCP

[SBMP16] Seok-Zun Song, LeRoy B. Beasley, Preeti Mohindru, and Rajesh Pereira. Preservers of completely positive matrix rank. Linear Multilinear Algebra, 64(7):1258–1265, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shi:2020:SMW

[SC20] Dongwei Shi and Jianbing Cao. The Sherman–Morrison– Woodbury formula for the Moore–Penrose metric generalized inverse. Linear Multilinear Algebra, 68(5):972–982, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sang:2021:NEB

[SC21] Caili Sang and Zhen Chen. A new error bound for linear complementarity problems of weakly chained diagonally dominant B-matrices. Linear Multilinear Algebra, 69(10):1909– 1921, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2022:ZEU

[SC22] Zhiwei Song and Lin Chen. On the zero entries in a unitary matrix. Linear Multilinear Algebra, 70(7):1271–1280, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schwerdtfeger:1975:PIG

[Sch75] Hans Schwerdtfeger. On (n + 1)-point invariants of the group GL(n, F). Linear and Multilinear Algebra, 3(1–2):105–109, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schneider:1977:OTT

[Sch77] Hans Schneider. Olga Taussky-Todd’s influence on matrix theory and matrix theorists. Linear and Multilinear Algebra, 5(3): 197–224, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). A discursive personal tribute.

Schwerdtfeger:1978:PIG

[Sch78] Hans Schwerdtfeger. On (n + 1)-point invariants of the group GL(n, F). II. Linear and Multilinear Algebra, 6(1):63–64, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Scharlemann:1979:SFG

[Sch79] Martin Scharlemann. Subgroups of SL(2, R) freely generated by three parabolic elements. Linear and Multilinear Algebra, 7(3):177–191, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schmidt:1991:RP

[Sch91] Frank W. Schmidt. Research problems. Linear and Multilinear Algebra, 30(3):229–230, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Scheinerman:1992:BR

[Sch92a] Edward R. Scheinerman. Book review. Linear and Multilinear Algebra, 32(1):85, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schwarz:1992:BCD

[Sch92b] Binyamin Schwarz. Bounds for the Carathéodory distance of the unit ball. Linear and Multilinear Algebra, 32(2):93–101, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schellwat:1995:ICG

[Sch95] Holger Schellwat. Induced Cayley graphs and double covers. Linear and Multilinear Algebra, 39(1–2):161–164, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schmale:2015:LE

[Sch15] Wiland Schmale. Letter to the editor. Linear Multilinear Algebra, 63(3):650, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schmitt:2022:EDR

[Sch22] Alexander H. W. Schmitt. An elementary discussion of the representation and geometric invariant theory of equioriented quivers of type D with an application to quiver bundles. Linear Multilinear Algebra, 70(21):7142–7175, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2020:OIT

[SĆK^{+}20] Predrag S. Stanimirović, Miroslav Ćirić, Vasilios N. Katsikis, Chaoqian Li, and Haifeng Ma. Outer and (b, c) inverses of tensors. Linear Multilinear Algebra, 68(5):940–971, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2022:RGP

[SCL^{+}22] Predrag S. Stanimirović, Miroslav ’Cirić, Alberto Lastra, J. Rafael Sendra, and Juana Sendra. Representations and geometrical properties of generalized inverses over fields. Linear Multilinear Algebra, 70(22):7318–7338, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2020:NLE

[SD20] Shaowei Sun and Kinkar Ch. Das. Normalized Laplacian eigenvalues with chromatic number and independence number of graphs. Linear Multilinear Algebra, 68(1):63–80, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sciriha:2012:RSG

[SdF12] Irene Sciriha and C. M. da Fonseca. On the rank spread of graphs. Linear Multilinear Algebra, 60(1):73–92, 2012. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2010:NSM

[SDL10] Xiaosong Sun, Xiankun Du, and Dayan Liu. On non-singular multilinear maps. Linear Multilinear Algebra, 58(3):297–303, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Salazar-Diaz:2016:CDT

[SDVWT16] O. P. Salazar-Díaz, R. Velásquez, and L. A. Wills-Toro. Construction of dialgebras through bimodules over algebras. Linear Multilinear Algebra, 64(10):1980–2001, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seddik:2004:NRN

[Sed04] Ameur Seddik. On the numerical range and norm of elementary operators. Linear Multilinear Algebra, 52(3–4):293–302, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seddik:2011:CCU

[Sed11] Ameur Seddik. Characterization of the class of unitary operators by operator inequalities. Linear Multilinear Algebra, 59(9):1069–1074, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seidel:1995:GRG

[Sei95] J. J. Seidel. Geometric representations of graphs. Linear and Multilinear Algebra, 39(1–2):45–57, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sell:1981:CFD

[Sel81] George R. Sell. A characterization of finite dimensional translation invariant spaces. Linear and Multilinear Algebra, 10 (1):53–57, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Semrl:1997:CMH

[Šem97] Peter Šemrl. Characterization of matrices having rank k. Linear and Multilinear Algebra, 42(3):233–238, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seneta:1981:BR

[Sen81] E. Seneta. Book reviews. Linear and Multilinear Algebra, 9(4): 347–349, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seneta:1983:SLE

[Sen83] E. Seneta. Spectrum localization by ergodicity coefficients for stochastic matrices. Linear and Multilinear Algebra, 14(4): 343–347, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sawano:2021:SGD

[SES21] Yoshihiro Sawano and Saad R. El-Shabrawy. The spectra of the generalized difference operators on the spaces of convergent series. Linear Multilinear Algebra, 69(4):732–751, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stegeman:2017:BRR

[SF17] Alwin Stegeman and Shmuel Friedland. On best rank-2 and rank-(2, 2, 2) approximations of order-3 tensors. Linear Multilinear Algebra, 65(7):1289–1310, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shen:1996:SEI

[SG96] Jian Shen and D. A. Gregory. Some eigenvalue inequalities for matrices with restricted entries. Linear and Multilinear Algebra, 41(2):107–112, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sciriha:2006:MCT

[SG06] Irene Sciriha and Ivan Gutman. Minimal configuration trees. Linear Multilinear Algebra, 54(2):141–145, 2006. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2009:ECL

[SG09a] Yanling Shao and Yubin Gao. Exponents of 2-colorings of loopless, symmetric digraphs. Linear Multilinear Algebra, 57 (1):65–74, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2009:ETC

[SG09b] Yanling Shao and Yubin Gao. On the exponents of two-colored digraphs with two cycles. Linear Multilinear Algebra, 57(2): 185–199, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shafiei:2018:NRP

[SG18] M. Shafiei and A. G. Ghazanfari. Numerous refinements of Pólya and Heinz operator inequalities. Linear Multilinear Algebra, 66(4):852–860, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shekhar:2022:NDW

[SGM22] Vaibhav Shekhar, Chinmay Kumar Giri, and Debasisha Mishra. A note on double weak splittings of type II. Linear Multilinear Algebra, 70(12):2297–2317, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2005:ECT

[SGS05] Yanling Shao, Yubin Gao, and Liang Sun. Exponents of a class of two-colored digraphs. Linear Multilinear Algebra, 53(3): 175–188, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:1992:DOM

[SH92] Hung-Min Sun and Tzonelih Hwang. On double-one matrices and double-zero matrices. Linear and Multilinear Algebra, 31(1–4):47–55, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Smith:1998:IMP

[SH98] Ronald L. Smith and Shu-An Hu. Inequalites for monotonic pairs of Z-matrices. Linear and Multilinear Algebra, 44(1): 57–65, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Safarizadeh:2021:MPQ

[SH21] M. Safarizadeh and S. Hejazian. Maps preserving quadratic product between positive cones of JBW -algebras. Linear Multilinear Algebra, 69(3):471–488, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shapiro:1991:UBN

[Sha91] Alexander Shapiro. Upper bounds for nearly optimal diagonal scaling of matrices. Linear and Multilinear Algebra, 29(2): 145–147, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shahriari:1994:M

[Sha94] Shahriar Shahriari. On maximizing det X^{t}X. Linear and Multilinear Algebra, 36(4):275–278, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shahriari:1995:BR

[Sha95] Shahriar Shahriari. Book review. Linear and Multilinear Algebra, 39(4):397–400, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shang:2015:DEI

[Sha15] Yilun Shang. Distance Estrada index of random graphs. Linear Multilinear Algebra, 63(3):466–471, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2018:NIC

[Sha18] Jingjing Shao. Norm inequalities for commutators of operators in noncommutative-spaces. Linear Multilinear Algebra, 66(10):2100–2106, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2020:GRY

[Sha20] Jingjing Shao. Generalizations of refined Young inequalities and reverse inequalities for τ-measurable operators. Linear

Song:2002:MPF

[SHC02] Seok-Zun Song, Suk-Geun Hwang, and Gi-Sang Cheon. Minimum permanents on a face of the polytope of doubly stochastic matrices II. Linear Multilinear Algebra, 50(2):101–111, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shchegolev:2021:BPM

[Shc21] Alexander Shchegolev. Bijective PC-maps of the unipotent radical of the Borel subgroup of the classical symplectic group. Linear Multilinear Algebra, 69(7):1302–1337, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shen:2022:PSR

[SHC22] Runshuan Shen, Guolin Hou, and Alatancang Chen. On the pseudo-spectra and the related properties of infinitedimensional Hamiltonian operators. Linear Multilinear Algebra, 70(22):7728–7739, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schulte-herbruggen:2008:SBN

[ShDHG08] Thomas Schulte-herbrüggen, Gunther Dirr, Uwe Helmke, and Steffen J. Glaser. The significance of the C-numerical range and the local C-numerical range in quantum control and quantum information. Linear Multilinear Algebra, 56(1–2):3–26, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shemesh:1977:TO

[She77] Dan Shemesh. On a theorem of A. Ostrowski. Linear and Multilinear Algebra, 5(2):139–143, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shitov:2015:NRM

[Shi15a] Yaroslav Shitov. Nonnegative rank of a matrix with one negative eigenvalue. Linear Multilinear Algebra, 63(2):221, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shitov:2015:TCS

[Shi15b] Yaroslav Shitov. Two concepts of singularity for matrices over semirings. Linear Multilinear Algebra, 63(1):201–220, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shitov:2016:MMT

[Shi16] Yaroslav Shitov. On the max-min and tropical CP-rank conjectures. Linear Multilinear Algebra, 64(2):219–220, 2016. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shirokov:2018:CLA

[Shi18] D. S. Shirokov. Classification of Lie algebras of specific type in complexified Clifford algebras. Linear Multilinear Algebra, 66(9):1870–1887, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shkarin:2012:HTO

[Shk12] Stanislav Shkarin. Hypercyclic tuples of operators on C^{n} and R^{n}. Linear Multilinear Algebra, 60(8):885–896, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shreve:1990:QFC

[Shr90] W. E. Shreve. A quadratic form characterization of complete N-partite graphs. Linear and Multilinear Algebra, 28(1–2): 71–73, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shiu:2017:SLS

[SHS17] Wai Chee Shiu, Peng Huang, and Pak Kiu Sun. The signless Laplacian spectral radius of k-connected irregular graphs. Linear Multilinear Algebra, 65(4):830–839, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sierksma:1999:LPP

[Sie99] Gerard Sierksma. Limiting polytopes and periodic Markov chains. Linear and Multilinear Algebra, 46(4):281–298, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1984:CAR

[Sil84] F. Conceição Silva. On a conjecture about rational valued characters. Linear and Multilinear Algebra, 15(3–4):257–263, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1988:ESM

[Sil88a] Fernando C. Silva. The eigenvalues of the sum of matrices with prescribed invariant polynomials, II. Linear and Multilinear Algebra, 24(1):45–49, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1988:RDMb

[Sil88b] Fernando C. Silva. The rank of the difference of matrices with prescribed invariant factors. Linear and Multilinear Algebra, 24(1):59–63, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1988:RDMa

[Sil88c] Fernando C. Silva. The rank of the difference of matrices with prescribed similarity classes. Linear and Multilinear Algebra, 24(1):51–58, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1990:SPM

[Sil90] Fernando C. Silva. Sums and products of matrices with prescribed similarity classes. Linear and Multilinear Algebra, 27(4):317–323, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1993:EPM

[Sil93] Fernando C. Silva. The eigenvalues of the product of matrices with prescribed similarity classes. Linear and Multilinear Algebra, 34(3–4):269–277, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1995:CPP

[Sil95] Fernando C. Silva. On controllability of partially prescribed matrices. Linear and Multilinear Algebra, 40(2):95–102, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1998:IPM

[Sil98] Fernando C. Silva. Invariant polynomials of matrices with prescribed entries. Linear and Multilinear Algebra, 44(4):357– 372, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1999:EMP

[Sil99] Fernando C. Silva. On the eigenvalues of matrices with prescribed upper triangular part. Linear and Multilinear Algebra, 46(1–2):117–138, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Simons:1983:MRP

[Sim83] S. Simons. Measurable relations, pentagonal arrays and the determinants of probabilistic integrals. Linear and Multilinear Algebra, 13(1):41–66, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Simic:1995:SNG

[Sim95] Slobodan K. Simić. Some notes on graphs whose second largest eigenvalue is less than (\sqrt{5} − 1)/2. Linear and Multilinear Algebra, 39(1–2):59–71, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1973:DWS

[Sin73] Richard Sinkhorn. Doubly whose stochastic matrices squares leave the permanent invariant. Linear and Multilinear Algebra, 1(2):103–118, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1976:DSM

[Sin76a] Richard Sinkhorn. Doubly stochastic matrices whose squares decrease the permanent. Linear and Multilinear Algebra, 4(2): 123–128, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1976:MWC

[Sin76b] Richard Sinkhorn. Matrices which commute with doubly stochastic matrices. Linear and Multilinear Algebra, 4(3):201– 203, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1977:DSM

[Sin77] Richard Sinkhorn. Doubly stochastic matrices with dominant p-minors. Linear and Multilinear Algebra, 5(2):107–117, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1980:CQM

[Sin80] Richard Sinkhorn. Concerning the question of monotonicity of the permanent on the doubly stochastic matrices. Linear and Multilinear Algebra, 8(4):323–328, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1981:CME

[Sin81a] Richard Sinkhorn. Concerning the magnitude of the entries in a doubly stochastic matrix. Linear and Multilinear Algebra, 10(2):107–112, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1981:NWV

[Sin81b] Richard Sinkhorn. A neighborhood in which the van der Waerden permanent conjecture is valid. Linear and Multilinear Algebra, 10(3):217–221, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1982:LAP

[Sin82a] Richard Sinkhorn. Linear adjugate preservers on the complex matrices. Linear and Multilinear Algebra, 12(3):215–222, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1982:SPR

[Sin82b] Richard Sinkhorn. Some problems and results concerning the spectrums of stochastic matrices. Linear and Multilinear Algebra, 11(4):357–361, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Singer:1983:APM

[Sin83] Ivan Singer. Abstract Pontryagin maximum principles for linear systems. Linear and Multilinear Algebra, 13(3):203–219, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1984:GSA

[Sin84a] Richard Sinkhorn. Generalized self adjointness in the complex doubly stochastic matrices. Linear and Multilinear Algebra, 16(1–4):67–73, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1984:PRV

[Sin84b] Richard Sinkhorn. A problem related to the van der Waerden permanent theorem. Linear and Multilinear Algebra, 16(1–4): 167–173, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sing:1990:STM

[Sin90] Lee Yew Sing. On the sum of two N_{0}-matrices. Linear and Multilinear Algebra, 26(3):215–221, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinkhorn:1993:CHF

[Sin93] Richard Sinkhorn. Concerning Hermite forms for matrices and their transposes. Linear and Multilinear Algebra, 34(1):75–84, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Singh:2021:BST

[Sin21] Gurmail Singh. Bialgebra structures on table algebras. Linear Multilinear Algebra, 69(12):2288–2314, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sivakumar:2006:GII

[Siv06a] K. C. Sivakumar. Generalized inverses of an invertible infinite matrix. Linear Multilinear Algebra, 54(2):113–122, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sivakumar:2006:MPI

[Siv06b] K. C. Sivakumar. Moore–Penrose inverse of an invertible infinite matrix. Linear Multilinear Algebra, 54(1):71–77, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2002:MPT

[SJ02] Seok-Zun Song and Young-Bae Jun. Minimum permanents of tridiagonal doubly stochastic matrices. Linear Multilinear Algebra, 50(4):301–306, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sjogren:1993:CBR

[Sjo93] Jon A. Sjogren. Circulant blocks and rotational graphs. Linear and Multilinear Algebra, 34(2):141–176, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2003:CRT

[SK03] Seok-Zun Song and Kyung-Tae Kang. Column ranks and their preservers of matrices over max algebra. Linear Multilinear Algebra, 51(3):311–318, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Suyama:2015:BEB

[SK15a] Koichi Suyama and Nobuko Kosugi. Bezout equations over bivariate polynomial matrices related by an entire function. Linear Multilinear Algebra, 63(6):1138–1153, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Suyama:2015:NAB

[SK15b] Koichi Suyama and Nobuko Kosugi. A new approach to Bezout equations derived from multivariate polynomial matrices and real entire functions. Linear Multilinear Algebra, 63(11): 2318–2331, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Skalna:2017:SRP

[Ska17] Iwona Skalna. Strong regularity of parametric interval matrices. Linear Multilinear Algebra, 65(12):2472–2482, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Su:2020:LSR

[SKLS20] Li Su, Liying Kang, Honghai Li, and Erfang Shan. The largest spectral radius of uniform hypertrees with a given size of matching. Linear Multilinear Algebra, 68(9):1779–1791, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Su:2021:SLS

[SKLS21] Li Su, Liying Kang, Honghai Li, and Erfang Shan. The second largest spectral radii of uniform hypertrees with given size of matching. Linear Multilinear Algebra, 69(14):2674–2701, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2017:RPW

[SKM17] Predrag S. Stanimirović, Vasilios N. Katsikis, and Haifeng Ma. Representations and properties of the W -weighted Drazin inverse. Linear Multilinear Algebra, 65(6):1080–1096, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Skorohod:1985:BR

[Sko85] A. V. Skorohod. Book review. Linear and Multilinear Algebra, 18(3):267–269, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2018:CIB

[SKP18] Predrag S. Stanimirović, Vasilios N. Katsikis, and Dimitrios Pappas. Computation of {2, 4} and {2, 3}-inverses based on rank-one updates. Linear Multilinear Algebra, 66(1):147–166, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sohn:1995:CPC

[SL95] Moo Young Sohn and Jaeun Lee. Characteristic polynomials of complement of graph bundles and their applications. Linear and Multilinear Algebra, 39(1–2):179–190, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2015:VGR

[SL15] Youngkwon Song and Doohann Lee. Vector generators of the real Clifford algebras. Linear Multilinear Algebra, 63(12): 2448–2460, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2017:EPS

[SL17] Xiuhong Sun and Yuan Li. Extension properties of some completely positive maps. Linear Multilinear Algebra, 65(7):1374–

Sang:2019:ESE

[SL19a] Caili Sang and Chaoqian Li. Exclusion sets in eigenvalue localization sets for tensors. Linear Multilinear Algebra, 67(12): 2399–2409, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lau:2019:POB

[sL19b] Pan shun Lau. Pre-order between diagonal elements and singular values of real matrices. Linear Multilinear Algebra, 67(9): 1856–1866, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stavrou:2021:MRT

[SL21] Stavros Georgios Stavrou and Richard M. Low. The maximum rank of 2 × ⋯ × 2 tensors over F_{2}. Linear Multilinear Algebra, 69(3):394–402, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slanina:2016:GFP

[Sla16] Piotr Slanina. Generalizations of Fibonacci polynomials and free linear groups. Linear Multilinear Algebra, 64(2):187–195, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stavrou:2016:RCT

[SLH16] Stavros G. Stavrou, Richard M. Low, and Nicholas J. Hernandez. Rank classification of 2 × 2 × 2 × 2 × 2 tensors over F_{2}. Linear Multilinear Algebra, 64(11):2297–2312, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [SLH18].

Stavrou:2018:CRC

[SLH18] Stavros G. Stavrou, Richard M. Low, and Nicholas J. Hernandez. Corrigendum to ‘Rank classification of 2 × 2 × 2 × 2 × 2 tensors over F_{2}’. Linear Multilinear Algebra, 66(3):514–515, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [SLH16].

Shao:2020:BFS

[SLH20] Changpeng Shao, Hongbo Li, and Lei Huang. Basis-free solution to general linear quaternionic equation. Linear Multilinear Algebra, 68(3):435–457, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slowik:2013:BMI

[Slo13a] Roksana Slowik. Bijective maps of infinite triangular and unitriangular matrices preserving commutators. Linear Multilinear Algebra, 61(8):1028–1040, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slowik:2013:ITG

[Slo13b] Roksana Slowik. Involutions in triangular groups. Linear Multilinear Algebra, 61(7):909–916, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slowik:2013:RE

[Slo13c] Roksana Slowik. Real elements in t_{n} ( K ). Linear Multilinear Algebra, 61(5):667–677, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slowik:2014:HCT

[Slo14a] Roksana Slowik. How to construct a triangular matrix of a given order. Linear Multilinear Algebra, 62(1):28–38, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slowik:2014:MIT

[Slo14b] Roksana Slowik. Maps on infinite triangular matrices preserving idempotents. Linear Multilinear Algebra, 62(7):938–964, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slowik:2015:TLP

[Slo15a] R. Slowik. Three linear preservers on infinite triangular matrices. Linear Multilinear Algebra, 63(4):672–694, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slowik:2015:ERI

[Slo15b] Roksana Slowik. Epimorphisms of the ring of infinite triangular matrices. Linear Multilinear Algebra, 63(7):1372–1378,

Slowik:2016:CMT

[Slo16a] Roksana Slowik. Continuous maps on triangular matrices that preserve commutativity. Linear Multilinear Algebra, 64(9): 1716–1730, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slowik:2016:SSZ

[Slo16b] Roksana Slowik. Sums of square-zero infinite matrices. Linear Multilinear Algebra, 64(9):1760–1768, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [Sło18b].

Slowik:2017:EIT

[Slo17] Roksana Slowik. Every infinite triangular matrix is similar to a generalized infinite Jordan matrix. Linear Multilinear Algebra, 65(7):1362–1373, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See corrigendum [Sło18a].

Slowik:2018:CEI

[Sło18a] Roksana Słowik. Corrigendum to ‘Every infinite triangular matrix is similar to a generalized infinite Jordan matrix’ (Linear Multilinear Algebra 65 (2017), 1362–1373). Linear Multilinear Algebra, 66(6):1278, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Slo17].

Slowik:2018:CSS

[Sło18b] Roksana Słowik. Corrigendum to ‘Sums of square-zero infinite matrices’ (Linear Multilinear Algebra 64 (2016), 1760–1768). Linear Multilinear Algebra, 66(6):1277, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Slo16b].

Shi:2004:TKN

[SLUG04] Xiquan Shi, Fengshan Liu, Hanson Umoh, and Frank Gibson. Two kinds of nonlinear matrix equations and their corresponding matrix sequences. Linear Multilinear Algebra, 52(1):1–15, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Samar:2021:CNW

[SLW21] Mahvish Samar, Hanyu Li, and Yimin Wei. Condition numbers for the K-weighted pseudoinverse and their statistical estimation. Linear Multilinear Algebra, 69(4):752–770, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shaked-Monderer:1994:EPS

[SM94] Naomi Shaked-Monderer. Extermal positive semidefinite matrices with given sparsity pattern. Linear and Multilinear Algebra, 36(4):287–292, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sirisuk:2019:GSG

[SM19] Siripong Sirisuk and Yotsanan Meemark. Generalized symplectic graphs and generalized orthogonal graphs over finite commutative rings. Linear Multilinear Algebra, 67(12):2427– 2450, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shaked-Monderer:2020:GIC

[SM20] Naomi Shaked-Monderer. On the generalized inverse of a copositive matrix. Linear Multilinear Algebra, 68(7):1384– 1394, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sababheh:2021:MAN

[SM21] Mohammad Sababheh and Hamid Reza Moradi. More accurate numerical radius inequalities (i). Linear Multilinear Algebra, 69(10):1964–1973, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shaked-Monderer:2022:NCF

[SM22a] Naomi Shaked-Monderer. On the number of CP factorizations of a completely positive matrix. Linear Multilinear Algebra, 70(19):3887–3904, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shaked-Monderer:2022:PCR

[SM22b] Naomi Shaked-Monderer. Preservers of the cp-rank. Linear Multilinear Algebra, 70(21):6211–6222, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Saddi:2023:PIS

[SM23] Adel Saddi and Fatma Mahmoudi. (A, m)-partial isometries in semi-Hilbertian spaces. Linear Multilinear Algebra, 71(10): 1640–1656, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Smaga:2017:OHE

[Sma17] Ł. Smaga. E-optimal and highly E-efficient designs with (mod 4) correlated observations. Linear Multilinear Algebra, 65(11): 2293–2305, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shaked-Monderer:2015:NRC

[SMBB^{+}15] Naomi Shaked-Monderer, Abraham Berman, Immanuel M. Bomze, Florian Jarre, and Werner Schachinger. New results on the cp-rank and related properties of co(mpletely )positive matrices. Linear Multilinear Algebra, 63(2):384–396, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shiu:1995:RCL

[SMF95] W. C. Shiu, S. L. Ma, and K. T. Fang. On the rank of cyclic latin squares. Linear and Multilinear Algebra, 40(2):183–188, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Smith:1974:GTN

[Smi74] John Howard Smith. A geometric treatment of non-negative generalized inverses. Linear and Multilinear Algebra, 2(2):179– 184, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Smith:1976:CNM

[Smi76] John Howard Smith. Commutators of nilpotent matrices. Linear and Multilinear Algebra, 4(1):17–19, 1976. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Smith:1983:SCL

[Smi83] John Howard Smith. Schur complements and LDU decomposition in an additive category. Linear and Multilinear Algebra, 14(1):11–19, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Smith:1994:FPM

[Smi94] Ronald L. Smith. On fan products of Z-matrices. Linear and Multilinear Algebra, 37(4):297–302, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Smigoc:2005:CNM

[Smi05] Helena Smigoc. Construction of nonnegative matrices and the inverse eigenvalue problem. Linear Multilinear Algebra, 53(2): 85–96, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sato:2014:MWZ

[SMM14] Iwao Sato, Hideo Mitsuhashi, and Hideaki Morita. A matrixweighted zeta function of a graph. Linear Multilinear Algebra, 62(1):114–125, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sato:2016:GBZ

[SMM16] Iwao Sato, Hideo Mitsuhashi, and Hideaki Morita. A generalized Bartholdi zeta function for a general graph. Linear Multilinear Algebra, 64(5):991–1008, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2022:NCM

[SMM22] Predrag S. Stanimirović, Dijana Mosić, and Haifeng Ma. New classes of more general weighted outer inverses. Linear Multilinear Algebra, 70(1):122–147, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sumi:2015:TRT

[SMS15] Toshio Sumi, Mitsuhiro Miyazaki, and Toshio Sakata. Typical ranks of m × n × (m − 1)n tensors with 3 ≤ m ≤ n over the real number field. Linear Multilinear Algebra, 63(5):940–955, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sumi:2021:TRS

[SMS21] Toshio Sumi, Mitsuhiro Miyazaki, and Toshio Sakata. Typical ranks of semi-tall real 3-tensors. Linear Multilinear Algebra, 69(9):1734–1746, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shamsaki:2022:TTP

[SN22a] Afsaneh Shamsaki and Peyman Niroomand. On the triple tensor product of nilpotent Lie algebras. Linear Multilinear Algebra, 70(20):5879–5887, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sheng:2022:GIP

[SN22b] Zhou Sheng and Qin Ni. A generalization of inverse power method for computing eigenpairs of symmetric tensors. Linear Multilinear Algebra, 70(20):4973–4990, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shateri:2023:BSI

[SN23] T. L. Shateri and B. Mohebbi Najmabad. On the Bohr and Selberg inequalities in 2-∗-inner product spaces. Linear Multilinear Algebra, 71(3):348–362, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

So:1990:FSS

[So90] W. So. Facial structures of Schatten p-norms. Linear and Multilinear Algebra, 27(3):207–212, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

So:1999:ROP

[So99] Wasin So. Rank one perturbation and its application to the Laplacian spectrum of a graph. Linear and Multilinear Algebra, 46(3):193–198, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Soares:2008:ZSC

[Soa08] Carlos A. S. Soares. Zero-sets of complex homogeneous polynomials. Linear Multilinear Algebra, 56(3):293–298, 2008. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Solow:1980:SCI

[Sol80] Anita E. Solow. The square class invariant for quadratic forms and the classification problem. Linear and Multilinear Algebra, 9(1):39–50, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Solues:1997:NP

[Sol97] George W. Solues. Nonconvexity of the permanent. Linear and Multilinear Algebra, 42(1):89–91, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Solues:2000:EMB

[Sol00] George W. Solues. Extending the Minc–Brègman upper bound for the permanent. Linear and Multilinear Algebra, 47(1):77– 91, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Solak:2006:RPN

[Sol06] Süleyman Solak. Research problem: on the norms of infinite Cauchy–Toeplitz-plus-Cauchy–Hankel matrices. Linear Multilinear Algebra, 54(6):397–398, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Soleymani:2014:FRA

[Sol14] Fazlollah Soleymani. On finding robust approximate inverses for large sparse matrices. Linear Multilinear Algebra, 62(10): 1314–1334, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2014:CSS

[Son14] Guang-Jing Song. Common solutions to some operator equations over Hilbert C^{∗}-modules and applications. Linear Multilinear Algebra, 62(7):895–912, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2022:CCN

[Son22] Youngkwon Song. Construction of commutative number systems. Linear Multilinear Algebra, 70(12):2382–2396, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Soules:1983:CSN

[Sou83] George W. Soules. Constructing symmetric nonnegative matrices. Linear and Multilinear Algebra, 13(3):241–251, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sourour:1986:FTM

[Sou86] A. R. Sourour. A factorization theorem for matrices. Linear and Multilinear Algebra, 19(2):141–147, 1986. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sourour:1992:NFM

[Sou92] A. R. Sourour. Nilpotent factorization of matrices. Linear and Multilinear Algebra, 31(1–4):303–308, 1992. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Soules:2003:NPU

[Sou03] George W. Soules. New permanental upper bounds for nonnegative matrices. Linear Multilinear Algebra, 51(4):319–337, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Simoes-Pereira:1987:NCS

[SP87] J. M. S. Simões-Pereira. A note on convexity and submatrices of distance matrices. Linear and Multilinear Algebra, 20(4): 363–366, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Spiegel:1976:PGS

[Spi76] Eugene Spiegel. Polynomials of generating series type. Linear and Multilinear Algebra, 4(2):115–121, 1976. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Spitkovsky:2000:NRP

[Spi00] Ilya M. Spitkovsky. On the non-round points of the boundary of the numerical range. Linear and Multilinear Algebra, 47 (1):29–33, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Spiegel:2004:LPI

[Spi04] E. Spiegel. Lie products in incidence algebras. Linear Multilinear Algebra, 52(2):115–120, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2015:GIR

[SPK15] Predrag S. Stanimirović, Dimitrios Pappas, and Vasilios N. Katsikis. Generalized inverse restricted by the normal Drazin equation. Linear Multilinear Algebra, 63(5):893–913, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2015:OIR

[SPKC15] Predrag S. Stanimirović, Dimitrios Pappas, Vasilios N. Katsikis, and Marija S. Cvetković. Outer inverse restricted by a linear system. Linear Multilinear Algebra, 63(12):2461–2493, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2014:MQF

[SPM14] Predrag S. Stanimirović, Dimitrios Pappas, and Sladjana Miljković. Minimization of quadratic forms using the Drazininverse solution. Linear Multilinear Algebra, 62(2):252–266, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sain:2020:CCS

[SPMR20] Debmalya Sain, Kallol Paul, Arpita Mal, and Anubhab Ray. A complete characterization of smoothness in the space of bounded linear operators. Linear Multilinear Algebra, 68(12): 2484–2494, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sushama:2015:EPF

[SPS15] Agrawal N. Sushama, K. Premakumari, and K. C. Sivakumar. Extensions of Perron–Frobenius splittings and relationships with nonnegative Moore–Penrose inverses. Linear Multilinear Algebra, 63(1):1–11, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2015:NSC

[SQ15] Yisheng Song and Liqun Qi. Necessary and sufficient conditions for copositive tensors. Linear Multilinear Algebra, 63(1): 120–131, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2015:SNT

[SQH15] Jia-Yu Shao, Liqun Qi, and Shenglong Hu. Some new trace formulas of tensors with applications in spectral hypergraph theory. Linear Multilinear Algebra, 63(5):971–992, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Santana:1999:GRM

[SQM99] Ana Paula Santana, João Filipe Queiró, and Educardo Marques De Sá. Group representations and matrix spectral problems. Linear and Multilinear Algebra, 46(1–2):1–23, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Selvan:2017:ITO

[SR17] A. Antony Selvan and R. Radha. Invertibility of a tridiagonal operator with an application to a non-uniform sampling problem. Linear Multilinear Algebra, 65(5):973–990, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinha:2022:ESR

[SR22] Deepa Sinha and Anita Kumari Rao. Embedding of signregular signed graphs and its spectral analysis. Linear Multilinear Algebra, 70(8):1496–1512, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sra:2020:NCC

[Sra20] Suvrit Sra. New concavity and convexity results for symmetric polynomials and their ratios. Linear Multilinear Algebra, 68(5):1031–1038, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sain:2022:SSP

[SRBB22] Debmalya Sain, Saikat Roy, Satya Bagchi, and Vitor Balestro. A study of symmetric points in Banach spaces. Linear Multilinear Algebra, 70(5):888–898, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sinha:2019:SAP

[SRD19] Deepa Sinha, Anita Kumari Rao, and Ayushi Dhama. Spectral analysis of t-path signed graphs. Linear Multilinear Algebra, 67(9):1879–1897, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sriwongsa:2020:OGM

[Sri20] Songpon Sriwongsa. Orthogonal graphs modulo power of 2. Linear Multilinear Algebra, 68(11):2337–2345, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Saha:2022:NLS

[SRK22] Tanay Saha, Suman Rakshit, and Swanand R. Khare. Nearest linearly structured polynomial matrix with some prescribed distinct eigenvalues. Linear Multilinear Algebra, 70(21):7000– 7026, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sahoo:2021:SEN

[SRS21] Satyajit Sahoo, Nirmal Chandra Rout, and Mohammad Sababheh. Some extended numerical radius inequalities. Linear Multilinear Algebra, 69(5):907–920, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sain:2022:LNB

[SRT22] Debmalya Sain, Saikat Roy, and Ryotaro Tanaka. Level numbers of a bounded linear operator between normed linear spaces and singular value decomposition revisited. Linear Multilinear Algebra, 70(19):3905–3922, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2007:CAI

[SRW07] Jia-Yu Shao, Ling-Zhi Ren, and Qun Wu. Characterizations and applications of the isolated sets of permutations. Linear Multilinear Algebra, 55(1):45–56, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Saunders:1979:CGO

[SS79] B. David Saunders and Hans Schneider. Cones, graphs and optimal scalings of matrices. Linear and Multilinear Algebra, 8(2):121–135, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1996:PNI

[SS96] Fernando C. Silva and Wasin So. Possibilities for the number of invariant polynomials of the difference of two similarity classes. Linear and Multilinear Algebra, 41(4):289–302, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shahryari:1998:PC

[SS98] M. Shahryari and M. A. Shahabi. On a permutation character of S_{m}. Linear and Multilinear Algebra, 44(1):45–52, 1998.

Simic:2007:PRU

[SS07] Slobodan K. Simić and Zoran Stanić. The polynomial reconstruction of unicyclic graphs is unique. Linear Multilinear Algebra, 55(1):35–43, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:2012:RCR

[SS12] Fernando C. Silva and Rita Simões. Row and column rank partitions under small perturbations. Linear Multilinear Algebra, 60(1):1–9, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seeger:2022:VRV

[SS22] Alberto Seeger and David Sossa. Vertex-removal, vertexaddition and different notions of similarity for vertices of a graph. Linear Multilinear Algebra, 70(20):5173–5192, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seeger:2023:SRF

[SS23] Alberto Seeger and David Sossa. Spectral radii of friendship graphs and their connected induced subgraphs. Linear Multilinear Algebra, 71(1):63–87, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2015:SSP

[SSfW15] Jia-Yu Shao, Hai-Ying Shan, and Bao feng Wu. Some spectral properties and characterizations of connected odd-bipartite uniform hypergraphs. Linear Multilinear Algebra, 63(12): 2359–2372, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shao:2012:ULB

[SSG12] Yanling Shao, Jian Shen, and Yubin Gao. The kth upper and lower bases of primitive nonpowerful minimally strong signed digraphs. Linear Multilinear Algebra, 60(9):1093–1113, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sain:2023:BCS

[SSGP23] Debmalya Sain, Shamim Sohel, Souvik Ghosh, and Kallol Paul. On best coapproximations in subspaces of diagonal matrices. Linear Multilinear Algebra, 71(1):47–62, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sivakumar:2022:MMT

[SST22] K. C. Sivakumar, P. Sushmitha, and M. Tsatsomeros. Q_{#}- matrices and Q_{+}-matrices: two extensions of the Q-matrix concept. Linear Multilinear Algebra, 70(21):6947–6964, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sierksma:1993:ICS

[SSZ93] Gerard Sierksma, Valeriu Soltan, and Tudor Zamfirescu. Invariance of convex sets under linear transformations. Linear and Multilinear Algebra, 35(1):37–47, 1993. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sanna:2022:LEI

[SSZ22] Carlo Sanna, Jeffrey Shallit, and Shun Zhang. The largest entry in the inverse of a Vandermonde matrix. Linear Multilinear Algebra, 70(20):5634–5641, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schneider:1973:MHA

[ST73] Hans Schneider and Robert E. L. Turner. Matrices Hermitian for an absolute norm. Linear and Multilinear Algebra, 1(1): 9–31, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shapiro:1980:APT

[ST80] H. Shapiro and O. Taussky. Alternative proofs of a theorem of Moyls and Marcus on the numerical range of a square matrix. Linear and Multilinear Algebra, 8(4):337–340, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shih:1986:CTS

[ST86] Mau-Hsiang Shih and Kok-Keong Tan. Covering theorems of simplexes and systems of linear inequalities. Linear and Multilinear Algebra, 19(4):309–320, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

So:1991:PEH

[ST91] W. So and Robert C. Thompson. Products of exponentials of Hermitian and complex symmetric matrices. Linear and Multilinear Algebra, 29(3–4):225–233, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sagae:1994:ULB

[ST94] Masahiko Sagae and Kunio Tanabe. Upper and lower bounds for the arithmetic–geometric–harmonic means of positive definite matrices. Linear and Multilinear Algebra, 37(4):279–282, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

So:1996:CUC

[ST96] Wasin So and Robert Thompson. Convexity of the upper complex plane part of the numerical range of a quaternionic matrix. Linear and Multilinear Algebra, 41(4):303–365, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Silva:1997:RPU

[ST97a] Fernando C. Silva and Maria M. Torres. Rank partitions under small perturbations. Linear and Multilinear Algebra, 43(1–3): 299–313, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Szlgeti:1997:SSL

[ST97b] Jenö Szlgeti and Zsolt Tuza. Solving systems of linear equations over Lie nilpotent rings. Linear and Multilinear Algebra, 42(1):43–51, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stitzinger:1999:CDL

[ST99] Ernest L. Stitzinger and Rudolph M. Turner. Concerning derivations of Lie algebras. Linear and Multilinear Algebra, 45(4):329–331, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

So:2000:SVM

[ST00] Wasin So and Robert C. Thompson. Singular values of matrix exponentials. Linear and Multilinear Algebra, 47(3):249–258,

Slyusareva:2008:MPP

[ST08] Olga Slyusareva and Michael Tsatsomeros. Mapping and preserver properties of the principal pivot transform. Linear Multilinear Algebra, 56(3):279–292, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Seltzer:2015:PBE

[ST15] Kevin Seltzer and J. Tao. On perturbation bounds of eigenvalues in Euclidean Jordan algebras. Linear Multilinear Algebra, 63(7):1379–1389, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sebestyen:2019:AHS

[ST19] Zoltán Sebestyén and Zsigmond Tarcsay. On the adjoint of Hilbert space operators. Linear Multilinear Algebra, 67(3): 625–645, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sady:2022:AJS

[ST22] Fereshteh Sady and Masoumeh Najafi Tavani. Additive jointly separating maps and ring homomorphisms. Linear Multilinear Algebra, 70(11):2162–2177, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanley:1984:SBS

[Sta84] Richard P. Stanley. The stable behavior of some characters of SL. Linear and Multilinear Algebra, 16(1–4):3–27, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanton:1994:BR

[Sta94] Dennis Stanton. Book review. Linear and Multilinear Algebra, 36(4):315–316, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanic:2010:RGC

[Sta10] Zoran Stanić. On regular graphs and coronas whose second largest eigenvalue does not exceed 1. Linear Multilinear Algebra, 58(5):545–554, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stavrou:2014:CFS

[Sta14] Stavros G. Stavrou. Canonical forms of 2 × 2× 2 and 2 × 2 × 2 × 2 symmetric tensors over prime fields. Linear Multilinear Algebra, 62(9):1169–1186, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanimirovic:2015:CIH

[Sta15a] Ivan P. Stanimirović. Computing A_{T,S}^{(2)} inverses of Hermitian matrices via LDL^{∗} decomposition for a square matrix A. Linear Multilinear Algebra, 63(8):1553–1567, 2015. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stavrou:2015:CFO

[Sta15b] Stavros Georgios Stavrou. Canonical forms of order-k (k = 2, 3, 4) symmetric tensors of format 3×… ×3 over prime fields. Linear Multilinear Algebra, 63(6):1111–1124, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanic:2020:SCS

[Sta20] Z. Stanić. Star complementary strongly regular decompositions of strongly regular graphs. Linear Multilinear Algebra, 68(12):2448–2461, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stanic:2022:COG

[Sta22] Z. Stanić. On cospectral oriented graphs and cospectral signed graphs. Linear Multilinear Algebra, 70(19):3689–3701, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Starcic:2023:IGA

[Sta23] Tadej Starcic. Isotropy groups of the action of orthogonal similarity on symmetric matrices. Linear Multilinear Algebra, 71(5):842–866, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stein:1982:ICI

[Ste82] Paul R. Stein. On an identity from classical invariant theory. Linear and Multilinear Algebra, 11(1):39–44, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stephens:1994:MAC

[Ste94a] Peter W. Stephens. A matrix analysis of carrier posets of biconnected graphs. Linear and Multilinear Algebra, 38(1–2): 45–58, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stephens:1994:MCI

[Ste94b] Peter W. Stephens. A matrix characterization of induced paths in bridge graphs. Linear and Multilinear Algebra, 36(3): 175–188, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stevanovic:2005:ENG

[Ste05] Dragan Stevanović. Energy and NEPS of graphs. Linear Multilinear Algebra, 53(1):67–74, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stitzinger:1987:PLS

[Sti87] Ernest L. Stitzinger. The probability that a linear system is consistent. Linear and Multilinear Algebra, 21(4):367–371, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Salemkar:2010:NAT

[STME10] Ali Reza Salemkar, Hamid Tavallaee, Hamid Mohammadzadeh, and Behrouz Edalatzadeh. On the non-abelian tensor product of Lie algebras. Linear Multilinear Algebra, 58(3):333–341, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stolarsky:1979:QDD

[Sto79] Kenneth B. Stolarsky. Quintuple diagonal determinants related to a digital distribution problem. Linear and Multilinear Algebra, 8(2):145–152, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stuart:1989:MWP

[Stu89] Jeffrey L. Stuart. Matrices whose powers are completely reducible Z-matrices. Linear and Multilinear Algebra, 25(1): 75–83, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sturtivant:1992:TEA

[Stu92] Carl Sturtivant. Are there elimination algorithms for the permanent? Linear and Multilinear Algebra, 33(3–4):145–162, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

So:1994:NRN

[STZ94] Wasin So, Robert C. Thompson, and Fuzhen Zhang. The numerical range of normal matrices with quaternion entries. Linear and Multilinear Algebra, 37(1–3):175–195, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Semrl:2015:P

[ŠTZ15] Peter Šemrl, Tin-Yau Tam, and Fuzhen Zhang. Preface. Linear Multilinear Algebra, 63(10):1915, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schlegelmilch:1977:MSR

[SU77] Rolf Schlegelmilch and Frank Uhlig. Über Matrizen mit Semidefinitem Realteil, Imaginärteil oder Polarem Defekt und Ihre Eigenwerte. (German) [On matrices with semidefinite real part, imaginary or polar defect, and their eigenvalues]. Linear and Multilinear Algebra, 5(3):225–228, 1977. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Suciu:2022:LEO

[Suc22] Laurian Suciu. Liftings and extensions of operators in Brownian setting. Linear Multilinear Algebra, 70(15):2948–2965, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sullivan:2008:PNM

[Sul08] R. P. Sullivan. Products of nilpotent matrices. Linear Multilinear Algebra, 56(3):311–317, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sundheim:2015:CH

[Sun15] Paul Sundheim. Construction of hyperdeterminants. Linear Multilinear Algebra, 63(9):1872–1900, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2020:NMB

[Sun20] Deshu Sun. New methods based on h-tensors for identifying positive definiteness of homogeneous polynomial forms. Linear Multilinear Algebra, 68(12):2473–2483, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2021:STS

[Sun21] Yujiao Sun. Supercharacter theories for Sylow p-subgroups of Lie type G_{2}. Linear Multilinear Algebra, 69(6):1102–1136, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sury:1998:ASC

[Sur98] B. Sury. Arithmetic of subgroup counting in some free products. Linear and Multilinear Algebra, 44(4):347–355, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Saeedi:2014:STC

[SV14] F. Saeedi and B. Veisi. On Schur’s theorem and its converses for n-Lie algebras. Linear Multilinear Algebra, 62(9):1139– 1145, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Szigeti:2014:DMS

[SvW14] J. Szigeti and L. van Wyk. Determinants for n × n matrices and the symmetric Newton formula in the 3 × 3 case. Linear Multilinear Algebra, 62(8):1076–1090, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Steeba:1981:EFK

[SW81] W.-H Steeba and F. Wilhelm. Exponential functions of Kronecker products and trace calculation. Linear and Multilinear Algebra, 9(4):345–346, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Steidl:2022:NCO

[SW22] Gabriele Steidl and Maximilian Winkler. A new constrained optimization model for solving the nonsymmetric stochastic inverse eigenvalue problem. Linear Multilinear Algebra, 70 (18):3419–3448, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shan:2022:SSE

[SWH22] Haiying Shan, Feifei Wang, and Changxiang He. Some αspectral extremal results for some digraphs. Linear Multilinear Algebra, 70(22):7493–7513, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shi:2013:BEP

[SWL13] Xinghua Shi, Yimin Wei, and Shuyang Ling. Backward error and perturbation bounds for high order Sylvester tensor equation. Linear Multilinear Algebra, 61(10):1436–1446, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2015:SRR

[SWZB15] Lizhu Sun, Wenzhe Wang, Jiang Zhou, and Changjiang Bu. Some results on resistance distances and resistance matrices. Linear Multilinear Algebra, 63(3):523–533, 2015. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Szeto:2012:MRA

[SX12] George Szeto and Lianyong Xue. A matrix representation of an Azumaya group ring. Linear Multilinear Algebra, 60(7): 769–774, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sheng:2015:HLA

[SX15] Yunhe Sheng and Zhen Xiong. On Hom–Lie algebras. Linear Multilinear Algebra, 63(12):2379–2395, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sheng:2022:MGJ

[SX22] Xingping Sheng and Dawei Xin. Methods of Gauss–Jordan elimination to compute core inverse and dual core inverse. Linear Multilinear Algebra, 70(12):2354–2366, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sendov:2023:CMM

[SX23] Hristo Sendov and Junquan Xiao. Coninvolutory matrices, multi-affine polynomials, and invariant circles. Linear Multilinear Algebra, 71(6):911–944, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shi:2018:BCL

[SXR18] Yuming Shi, Guixin Xu, and Guojing Ren. Boundedness and closedness of linear relations. Linear Multilinear Algebra, 66(2):309–333, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2023:GAG

[SY23] Shuoxi Sun and Cheng Yan. A generalized arithmeticgeometric mean-type inequality of measurable operator. Linear Multilinear Algebra, 71(1):120–132, 2023. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sababheh:2015:IYH

[SYK15] M. Sababheh, A. Yousef, and R. Khalil. Interpolated Young and Heinz inequalities. Linear Multilinear Algebra, 63(11): 2232–2244, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Su:2022:LSL

[SYZ22] Yucai Su, Xiaoqing Yue, and Xiaoyu Zhu. 2-local superderivations on the Lie superalgebra of Block type. Linear Multilinear Algebra, 70(21):6425–6440, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shen:2017:CSL

[SYZL17] Yun Shen, Lihua You, Minjie Zhang, and Shuchao Li. On a conjecture for the signless Laplacian spectral radius of cacti with given matching number. Linear Multilinear Algebra, 65(3):457–474, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schwarz:1985:EPS

[SZ85] Binyamin Schwarz and Abraham Zaks. On the embeddings of projective spaces in lines. Linear and Multilinear Algebra, 18(4):319–336, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sriwongsa:2019:OAC

[SZ19] Songpon Sriwongsa and Yi Ming Zou. Orthogonal abelian Cartan subalgebra decomposition of ∫↕_{\} over a finite commutative ring. Linear Multilinear Algebra, 67(4):817–829, 2019.

Szafraniec:2009:DNR

[Sza09] Franciszek Hugon Szafraniec. Decomposing numerical ranges along with spectral sets. Linear Multilinear Algebra, 57(5): 431–437, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2016:MPI

[SZBW16] Lizhu Sun, Baodong Zheng, Changjiang Bu, and Yimin Wei. Moore–Penrose inverse of tensors via Einstein product. Linear Multilinear Algebra, 64(4):686–698, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Szwarc:2018:CWA

[Szw18] Ryszard Szwarc. Convergence of weighted averages of relaxed projections. Linear Multilinear Algebra, 66(10):1991–2002, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2016:CSR

[SZWB16] Lizhu Sun, Baodong Zheng, Yimin Wei, and Changjiang Bu. Characterizations of the spectral radius of nonnegative weakly irreducible tensors via a digraph. Linear Multilinear Algebra, 64(4):737–744, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Sun:2018:SIH

[SZZY18] Lizhu Sun, Baodong Zheng, Jiang Zhou, and Hong Yan. Some inequalities for the Hadamard product of tensors. Linear Multilinear Algebra, 66(6):1199–1214, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taghavi:2012:AMA

[Tag12] Ali Taghavi. Additive mappings on C^{∗}-algebras preserving absolute values. Linear Multilinear Algebra, 60(1):33–38, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Torabian:2022:MOP

[TAK22] M. Torabian, M. Amyari, and M. Moradian Khibary. More on ω-orthogonalities and ω-parallelism. Linear Multilinear Alge-

Talebi:2017:ONF

[Tal17] Gholamreza Talebi. Operator norms of four-dimensional Hausdorff matrices on the double Euler sequence spaces. Linear Multilinear Algebra, 65(11):2257–2267, 2017. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1977:SRP

[Tam77] Bit-Shun Shun Tam. Some results of polyhedral cones and simplicial cones. Linear and Multilinear Algebra, 4(4):281–284, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1985:NPT

[Tam85] Tin-Yau Tam. Note on a paper of Thompson: the congruence numerical range. Linear and Multilinear Algebra, 17(2):107– 115, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Tho80].

Tam:1986:GRM

[Tam86] Tin-Yau Tam. On the generalized radial matrices and a conjecture of Marcus and Sandy. Linear and Multilinear Algebra, 19 (1):11–20, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1988:NRI

[Tam88] Tin-Yau Tam. On the numerical range of an induced power. Linear and Multilinear Algebra, 23(3):207–211, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1991:APR

[Tam91] Tin-Yau Tam. Another proof of a result of Marcus without plücker relations. Linear and Multilinear Algebra, 29(3–4): 313–314, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1994:CNR

[Tam94] Bit-Shun Tam. Circularity of numerical ranges and blockshift matrices. Linear and Multilinear Algebra, 37(1–3):93– 109, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1997:KCT

[Tam97] Tin-Yau Tam. Kostant’s convexity theorem and the compact classical groups. Linear and Multilinear Algebra, 43(1–3):87– 113, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1998:ENR

[Tam98] Bit-Shun Tam. Exterior numerical radii and antiinvariant subspaces. Linear and Multilinear Algebra, 44(3):195–245, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:2002:GSC

[Tam02] Tin-Yau Tam. Generalized Schur-concave functions and Eaton triples. Linear Multilinear Algebra, 50(2):113–120, 2002. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2010:DAP

[Tan10] Qianrong Tan. Divisibility among power GCD matrices and among power LCM matrices on two coprime divisor chains. Linear Multilinear Algebra, 58(5):659–671, 2010. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2012:LCL

[Tan12] Shang-Wang Tan. On the Laplacian coefficients and Laplacian-like energy of bicyclic graphs. Linear Multilinear Algebra, 60(9):1071–1092, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2013:IMC

[Tan13a] Yi-Jia Tan. On invertible matrices over commutative semirings. Linear Multilinear Algebra, 61(6):710–724, 2013. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tang:2013:RST

[Tan13b] Yuehan Tang. Random spectral theorems of self-adjoint random linear operators on complete complex random inner product modules. Linear Multilinear Algebra, 61(3):409–416, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2014:DMS

[Tan14] Yi-Jia Tan. Determinants of matrices over semirings. Linear Multilinear Algebra, 62(4):498–517, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2016:IPS

[Tan16] Yi-Jia Tan. Inner products on semimodules. Linear Multilinear Algebra, 64(8):1595–1616, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2018:DLI

[Tan18a] Yi-Jia Tan. Dedekind’s linear independence theorem over commutative semirings. Linear Multilinear Algebra, 66(1):107– 116, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2018:SIM

[Tan18b] Yi-Jia Tan. On strongly invertible matrices over semirings. Linear Multilinear Algebra, 66(12):2501–2511, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tang:2018:BFD

[Tan18c] Xiaomin Tang. Biderivations of finite-dimensional complex simple Lie algebras. Linear Multilinear Algebra, 66(2):250– 259, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2020:DMC

[Tan20a] Yi-Jia Tan. Diagonability of matrices over commutative semirings. Linear Multilinear Algebra, 68(9):1743–1752, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2020:FTM

[Tan20b] Yi-Jia Tan. Fillmore’s theorem for matrices over factorial rings. Linear Multilinear Algebra, 68(3):563–567, 2020. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1973:HTM

[Tau73a] Olga Taussky. Hilbert’s theorem 90 in matrix rings. Linear and Multilinear Algebra, 1(1):5–8, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1973:RP

[Tau73b] Olga Taussky. Research problem. Linear and Multilinear Algebra, 1(3):273, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1974:SRC

[Tau74] Olga Taussky. Some results concerning the transition from the L- to the P -property for pairs of finite matrices II. Linear and Multilinear Algebra, 2(3):195–202, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1981:CBI

[Tau81] Olga Taussky. Composition of binary integral quadratic forms via integral 2 × 2 matrices and composition of matrix classes. Linear and Multilinear Algebra, 10(4):309–318, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1976:BR

[TB76] R. C. Thompson and Richard A. Brualdi. Book reviews. Linear and Multilinear Algebra, 4(1):71–75, 1976. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1992:GIC

[TB92] Bit-Shun Tam and George Phillip Barker. Graphs and irreducible cone preserving maps. Linear and Multilinear Algebra, 31(1–4):19–25, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Teimoori:2001:PEF

[TB01] H. Teimoori and M. Bayat. Pascal k-eliminated functional matrix and Eulerian numbers. Linear and Multilinear Algebra, 49(3):183–194, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2004:SIM

[TC04] Yongge Tian and Shizhen Cheng. Some identities for Moore– Penrose inverses of matrix products. Linear Multilinear Alge-

Tosic:2013:IDS

[TCI13] Marina Tosić and Dragana S. Cvetković-Ilić. The invertibility of the difference and the sum of commuting generalized and hypergeneralized projectors. Linear Multilinear Algebra, 61(4): 482–493, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Talebi:2014:UBO

[TD14] Gholamreza Talebi and Mohammad Ali Dehghan. Upper bounds for the operator norms of Hausdorff matrices and nörlund matrices on the Euler weighted sequence space. Linear Multilinear Algebra, 62(10):1275–1284, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Talebi:2016:AUB

[TD16] Gholamreza Talebi and Mohammad Ali Dehghan. Approximation of upper bound for matrix operators on the Fibonacci weighted sequence spaces. Linear Multilinear Algebra, 64(2): 196–207, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Talebi:2019:AUB

[TD19] Gholamreza Talebi and Mohammad Ali Dehghan. Approximation of upper bound for matrix operators on the Fibonacci weighted sequence spaces. Linear Multilinear Algebra, 67(8): 1717–1718, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2018:MFN

[TDL18] Yan Tian, Xiankun Du, and Yueyue Li. Matrices with few nonzero principal minors. Linear Multilinear Algebra, 66(7): 1362–1379, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Teranishi:2001:MEG

[Ter01] Yasuo Teranishi. Main eigenvalues of a graph. Linear and Multilinear Algebra, 49(4):289–303, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Teranishi:2006:MEA

[Ter06] Yasuo Teranishi. Main eigenvalues and automorphisms of a graph. Linear Multilinear Algebra, 54(3):211–217, 2006. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Teranishi:2009:EAG

[Ter09] Yasuo Teranishi. Eigenvalues and automorphisms of a graph. Linear Multilinear Algebra, 57(6):577–585, 2009. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Teretenkov:2022:SVD

[Ter22a] A. E. Teretenkov. Singular value decomposition for skew-Takagi factorization with quantum applications. Linear Multilinear Algebra, 70(22):7762–7769, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Teretenkov:2022:SAP

[Ter22b] A. E. Teretenkov. Symplectic analogs of polar decomposition and their applications to bosonic Gaussian channels. Linear Multilinear Algebra, 70(9):1673–1681, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tarazaga:2009:EDM

[TG09] Pablo Tarazaga and Juan E. Gallardo. Euclidean distance matrices: new characterization and boundary properties. Linear Multilinear Algebra, 57(7):651–658, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tao:2023:DIT

[TG23] Zhengyu Tao and Xuejun Guo. On determinants involving tangent functions. Linear Multilinear Algebra, 71(13):2212– 2221, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tsai:2014:RNR

[TGW14] Ming-Cheng Tsai, Hwa-Long Gau, and Han-Chun Wang. Refinements for numerical ranges of weighted shift matrices. Linear Multilinear Algebra, 62(5):568–578, 2014. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1980:RCC

[TH80] Olga Taussky and J. Hsia. Results concerning composition of sums of three squares. Linear and Multilinear Algebra, 8(3): 231–233, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taghavi:2014:MPD

[TH14] Ali Taghavi and Roja Hosseinzadeh. Maps preserving the dimension of fixed points of products of operators. Linear Multilinear Algebra, 62(10):1285–1292, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taghavi:2018:MAP

[TH18] Ali Taghavi and Roja Hosseinzadeh. Mappings on C∗-algebras preserving the sum of absolute values. Linear Multilinear Algebra, 66(2):217–223, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

terHorst:2017:GSO

[tHK17] S. ter Horst and E. M. Klem. Graphs with sparsity order at most two: the complex case. Linear Multilinear Algebra, 65 (12):2367–2386, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1974:EPU

[Tho74] R. C. Thompson. On the eigenvalues of a product of unitary matrices I. Linear and Multilinear Algebra, 2(1):13–24, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1975:EDC

[Tho75a] R. C. Thompson. Extreme determinants on the convex hull of matrices with prescribed singular values. Linear and Multilinear Algebra, 3(1–2):15–17, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1975:ICH

[Tho75b] R. C. Thompson. The intersection of the convex hulls of proper and improper real matrices with prescribed singular values. Linear and Multilinear Algebra, 3(1–2):155–160, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1978:MTM

[Tho78] R. C. Thompson. Matrix type metric inequalities. Linear and Multilinear Algebra, 5(4):303–319, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1980:CNR

[Tho80] Robert C. Thompson. The congruence numerical range. Linear and Multilinear Algebra, 8(3):197–206, 1980. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See note [Tam85].

Thompson:1981:SBRa

[Tho81a] R. C. Thompson. Short book reviews. Linear and Multilinear Algebra, 10(2):163–164, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1981:SBRb

[Tho81b] R. C. Thompson. Short book reviews. Linear and Multilinear Algebra, 10(4):353–354, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1982:RP

[Tho82a] R. C. Thompson. Research problem. Linear and Multilinear Algebra, 11(4):373, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1982:SBRb

[Tho82b] R. C. Thompson. Short book review. Linear and Multilinear Algebra, 11(4):375, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1982:SBRc

[Tho82c] R. C. Thompson. Short book review. Linear and Multilinear Algebra, 12(1):73–74, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1982:SBRd

[Tho82d] R. C. Thompson. Short book review. Linear and Multilinear Algebra, 12(2):155–156, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1982:SBRa

[Tho82e] R. C. Thompson. Short book reviews. Linear and Multilinear Algebra, 11(1):103–104, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thomson:1982:SBR

[Tho82f] R. C. Thomson. Short book review. Linear and Multilinear Algebra, 11(3):303, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1983:SBR

[Tho83] R. C. Thompson. Short book review. Linear and Multilinear Algebra, 13(1):83–85, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1985:RP

[Tho85] R. C. Thompson. Research problem. Linear and Multilinear Algebra, 18(1):87–88, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1986:LMR

[Tho86a] Robert C. Thompson. Left multiples and right divisors of integral matrices. Linear and Multilinear Algebra, 19(3):287– 295, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1986:PCE

[Tho86b] Robert C. Thompson. Proof of a conjectured exponential formula. Linear and Multilinear Algebra, 19(2):187–197, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1986:SIM

[Tho86c] Robert C. Thompson. Sums of integral matrices. Linear and Multilinear Algebra, 19(2):173–186, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1987:IFI

[Tho87a] Robert C. Thompson. Invariant factors of integral quaternion matrices. Linear and Multilinear Algebra, 21(4):395–417, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1987:RPM

[Tho87b] Robert C. Thompson. Research problem the matrix numerical range. Linear and Multilinear Algebra, 21(3):321–323, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1988:DNS

[Tho88] Robert C. Thompson. A divisibility nonrelation for the Smith invariants of a product of integral matrices. Linear and Multilinear Algebra, 23(3):277–282, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1989:MVT

[Tho89] Robert C. Thompson. The matrix valued triangle inequality:quaternion version. Linear and Multilinear Algebra, 25(1): 85–91, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1990:EAA

[Tho90] Robert C. Thompson. Editorial announcement: Algebraic graph theory. Linear and Multilinear Algebra, 28(1–2):1, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1991:RP

[Tho91] R. C. Thompson. Research problem. Linear and Multilinear Algebra, 30(4):315–316, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1992:D

[Tho92a] Robert C. Thompson. Dedication. Linear and Multilinear Algebra, 32(2):87, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1992:ESH

[Tho92b] Robert C. Thompson. The eigenvalue spreads of a Hermitian matrix and its principal submatrices. Linear and Multilinear Algebra, 32(3–4):327–333, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thomsen:1992:HBF

[Tho92c] Klaus Thomsen. Homomorphisms between finite direct sums of circle algebras. Linear and Multilinear Algebra, 32(1):33–50, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1977:CHS

[TI77] R. C. Thompson and M. Ishaq. The convex hull of signed rearrangements with prescribed sign parity. Linear and Multilinear Algebra, 5(3):235–241, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2000:STL

[Tia00] Yongge Tian. The solvability of two linear matrix equations. Linear and Multilinear Algebra, 48(2):123–147, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2001:RER

[Tia01] Yongge Tian. Rank equalities related to outer inverses of matrices and applications. Linear and Multilinear Algebra, 49(4): 269–288, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2002:MRP

[Tia02] Yongge Tian. The minimum rank of a 3 × 3 partial block matrix. Linear Multilinear Algebra, 50(2):125–131, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2003:RSM

[Tia03] Yongge Tian. Ranks of solutions of the matrix equation AXB = C. Linear Multilinear Algebra, 51(2):111–125, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2005:MPI

[Tia05a] Yongge Tian. The Moore–Penrose inverse for sums of matrices under rank additivity conditions. Linear Multilinear Algebra, 53(1):45–65, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2005:SRE

[Tia05b] Yongge Tian. Some rank equalities and inequalities for Kronecker products of matrices. Linear Multilinear Algebra, 53(6): 445–454, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2010:EEG

[Tia10] Yongge Tian. Eight expressions for generalized inverses of a bordered matrix. Linear Multilinear Algebra, 58(2):203–220, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2011:DID

[Tia11a] Yongge Tian. A disjoint idempotent decomposition for linear combinations produced from two commutative tripotent matrices and its applications. Linear Multilinear Algebra, 59(11): 1237–1246, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2011:ERQ

[Tia11b] Yongge Tian. Extremal ranks of a quadratic matrix expression with applications. Linear Multilinear Algebra, 59(6):627–644, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2011:TUS

[Tia11c] Yongge Tian. Two universal similarity factorization equalities for commutative involutory and idempotent matrices and their applications. Linear Multilinear Algebra, 59(2):129–144, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2017:ULB

[Tia17] Yongge Tian. Upper and lower bounds for the rank of the matrix-valued function when X has a fixed rank. Linear Multilinear Algebra, 65(10):2101–2113, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2018:HEE

[Tia18] Yongge Tian. How to establish exact formulas for calculating the max–min ranks of products of two matrices and their generalized inverses. Linear Multilinear Algebra, 66(1):22–73,

Tian:2021:CAE

[Tia21] Yongge Tian. Classification analysis to the equalities a^{(i,...,j)} = b^{(k,...,l)} for generalized inverses of two matrices. Linear Multilinear Algebra, 69(8):1383–1406, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tie:2016:ECR

[Tie16] Lin Tie. On the existence of complex roots of real polynomials. Linear Multilinear Algebra, 64(5):980–990, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tikhonov:1998:MSF

[Tik98] O. E. Tikhonov. On matrix-subadditive functions and a relevant trace inequality. Linear and Multilinear Algebra, 44(1): 25–28, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tikaradze:2022:ISP

[Tik22] Akaki Tikaradze. Induced subgraphs of powers of oriented cycles. Linear Multilinear Algebra, 70(17):3301–3303, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tilgner:1977:GTR

[Til77] Hans Tilgner. Generalized tangent representations for groups and symmetric spaces of matrices. Linear and Multilinear Algebra, 5(1):27–31, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tilli:1998:NSD

[Til98] Paolo Tilli. A note on the spectral distribution of Toeplitz matrices. Linear and Multilinear Algebra, 45(2–3):147–159, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tismenetsky:1993:IHB

[Tis93] Miron Tismenetsky. On inversion of Hermitian block Toeplitz matrices. Linear and Multilinear Algebra, 35(2):165–171, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2011:LSR

[TJ11] Shang-Wang Tan and Jing-Jing Jiang. On the Laplacian spectral radius of weighted trees with fixed diameter and weight set. Linear Multilinear Algebra, 59(2):173–192, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2016:ASB

[TJ16] Yongge Tian and Bo Jiang. An algebraic study of BLUPs under two linear random-effects models with correlated covariance matrices. Linear Multilinear Algebra, 64(12):2351–2367, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Toyonaga:2021:CEG

[TJ21] Kenji Toyonaga and Charles R. Johnson. Classification of edges in a general graph associated with the change in multiplicity of an eigenvalue. Linear Multilinear Algebra, 69(10): 1803–1812, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tao:2022:SLW

[TJG22] J. Tao, J. Jeong, and M. Seetharama Gowda. Some log and weak majorization inequalities in Euclidean Jordan algebras. Linear Multilinear Algebra, 70(16):3189–3206, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taghavi:2019:NSS

[TKT19] Ali Taghavi, Farzaneh Kolivand, and Ensiyeh Tavakoli. A note on strong (skew) η-Lie products preserving maps on some algebras. Linear Multilinear Algebra, 67(5):886–895, 2019. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2012:DEI

[TL12] Yongge Tian and Ying Li. Distributions of eigenvalues and inertias of some block Hermitian matrices consisting of orthogonal projectors. Linear Multilinear Algebra, 60(9):1027–1069, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2013:CCB

[TL13] Xuezhong Tan and Bolian Liu. Complete characterization of bicyclic minimal configurations. Linear Multilinear Algebra, 61(7):976–985, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2016:DAP

[TLC16] Qianrong Tan, Zongbing Lin, and Long Chen. Divisibility among power matrices associated with arithmetic functions on finitely many quasi-coprime divisor chains. Linear Multilinear Algebra, 64(10):2030–2048, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2022:CSE

[TLC22] Gui-Xian Tian, Yuan Li, and Shu-Yu Cui. The change of Seidel energy of tripartite Turán graph due to edge deletion. Linear Multilinear Algebra, 70(19):4597–4614, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tan:2011:DAP

[TLL11] Qianrong Tan, Zongbing Lin, and Liu Liu. Divisibility among power GCD matrices and among power LCM matrices on two coprime divisor chains II. Linear Multilinear Algebra, 59(9): 969–983, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thongsomnuk:2019:PST

[TM19] Issaraporn Thongsomnuk and Yotsanan Meemark. Perfect state transfer in unitary Cayley graphs and gcd-graphs. Linear Multilinear Algebra, 67(1):39–50, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:2021:BCS

[TMH21] Derek Thompson, Thaddeus McClatchey, and Caleb Holleman. Binormal, complex symmetric operators. Linear Multilinear Algebra, 69(9):1705–1715, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tsatsomeros:1993:NMT

[TMOvdD93] Michael Tsatsomeros, John S. Maybee, D. D. Olesky, and P. van den Driessche. Nullspaces of matrices and their compounds. Linear and Multilinear Algebra, 34(3–4):291–300,

Taki:2022:SBV

[TN22] Zakaria Taki and Abdelkrim Nokrane. Spectrum of Banach valued holomorphic functions and polar sets. Linear Multilinear Algebra, 70(22):7650–7655, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tvalavadze:2022:NOF

[TNJ22] Marina Tvalavadze, Ekaterina Napedenina, and Qingzi Jiang. Note on one family of real division algebras. Linear Multilinear Algebra, 70(19):4351–4359, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Teresa:1988:HBW

[TOM88] M. Teresa and F. Oliveira-Martins. On homomorphisms between Weyl modules for hook partitions. Linear and Multilinear Algebra, 23(4):305–323, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tomaz:2019:LSC

[Tom19] Daniel Tomaz. Linear structure in certain subsets of quasi-Banach sequence spaces. Linear Multilinear Algebra, 67(8): 1561–1566, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tominaga:2022:EAH

[Tom22] Masaru Tominaga. Estimates of Ando–Hiai-type inequalities on operator power means. Linear Multilinear Algebra, 70(16): 3207–3219, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tong:1987:SRM

[Ton87] Wen-Ting Tong. On the spectral radius of matrices. Linear and Multilinear Algebra, 20(2):175–182, 1987. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tosio:1979:SOM

[Tos79] Kato Tosio. Spectral order and a matrix limit theorem. Linear and Multilinear Algebra, 8(1):15–19, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Towers:2022:ASI

[Tow22] David A. Towers. Abelian subalgebras and ideals of maximal dimension in supersolvable and nilpotent lie algebras. Linear Multilinear Algebra, 70(13):2551–2568, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taghavi:2021:GAO

[TR21] Ali Taghavi and Tahere Azimi Roushan. On a generalization of an approximation operator defined by A. Guessab. Linear Multilinear Algebra, 69(12):2340–2352, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Trapp:1984:HSM

[Tra84] George E. Trapp. Hermitian semidefinite matrix means and related matrix inequalities — an introduction. Linear and Multilinear Algebra, 16(1–4):113–123, 1984. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tran:1989:CPS

[Tra89] Cam Van Tran. Combinatorial properties of some classes of matrices over GF(2). Linear and Multilinear Algebra, 24(3): 153–166, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See correction [Tra90].

Tran:1990:CCP

[Tra90] Cam Van Tran. Correction to “Combinatiorial properties of some classes of matrices over GF(2)”. Linear and Multilinear Algebra, 26(1–2):145, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Tra89].

Taghavi:2016:NLJ

[TRD16] Ali Taghavi, Hamid Rohi, and Vahid Darvish. Non-linear ∗- Jordan derivations on von Neumann algebras. Linear Multilinear Algebra, 64(3):426–439, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Trench:1985:EPT

[Tre85] Willian F. Trench. On the eigenvalue problem for Toeplitz matrices generated by rational functions. Linear and Multilinear Algebra, 17(3–4):337–353, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Trench:1986:EPT

[Tre86] William F. Trench. On the eigenvalue problem for Toeplitz matrices generated by rational functions. Linear and Multilinear Algebra, 20(1):87, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Trench:1987:CPS

[Tre87] William F. Trench. Characteristic polynomials of symmetric rationally generated Toeplitz matrices. Linear and Multilinear Algebra, 21(3):289–296, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tug:2022:DGD

[TRM22] Orhan Tuğ, Viladimir Rakocević, and Eberhard Malkowsky. Domain of generalized difference operator Δ_{i}^{3} of order three on the Hahn sequence space h and matrix transformations. Linear Multilinear Algebra, 70(22):7433–7451, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2005:CST

[TS05] Yongge Tian and George P. H. Styan. Cochran’s statistical theorem for outer inverses of matrices and matrix quadratic forms. Linear Multilinear Algebra, 53(5):387–392, 2005. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tavakolipour:2019:NRT

[TS19] Hanieh Tavakolipour and Fatemeh Shakeri. On the numerical range in tropical algebra. Linear Multilinear Algebra, 67(10): 1985–1998, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taghavi:2021:CMP

[TS21] Ali Taghavi and Sadegh Salehi. Continuous maps preserving Jordan triple products from GL_{1} to GL_{2} to GL_{3}. Linear Multilinear Algebra, 69(2):208–223, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tsing:1981:SGN

[Tsi81] Nam-Kiu Tsing. On the shape of the generalized numerical ranges. Linear and Multilinear Algebra, 10(3):173–182, 1981.

Tsin:1983:DMW

[Tsi83] Nam-Kiu Tsin. Diameter and minimal width of the numerical range. Linear and Multilinear Algebra, 14(2):179–185, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tsay:1982:BPF

[TSYB82] Yih T. Tsay, Leang S. Shieh, R. E. Yates, and S. Barnett. Block partial fraction expansion of a rational matrix. Linear and Multilinear Algebra, 11(3):225–241, 1982. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1974:CBP

[TT74] R. C. Thompson and S. Therianos. On a construction of B. P. Zwahlen. Linear and Multilinear Algebra, 1(4):309–325, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky-Todd:1978:SCC

[TT78] Olga Taussky-Todd. Some comments concerning G. B. Price’s note on: “Determinants with dominant principal diagonal: a personal historical note” (Linear and Multilinear Algebra 6 (1978/79), no. 4, 247–250). Linear and Multilinear Algebra, 6(4):251, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Pri78].

Taussky-Todd:1985:RP

[TT85] Olga Taussky-Todd. Research problems. Linear and Multilinear Algebra, 17(1):85, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2006:CGI

[TT06] Yongge Tian and Yoshio Takane. On common generalized inverses of a pair of matrices. Linear Multilinear Algebra, 54(3): 195–209, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:2012:DST

[TT12] Tin-Yau Tam and Mary Clair Thompson. Determinants of sums of two real matrices and their extensions. Linear Multilinear Algebra, 60(11–12):1409–1431, 2012. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:2015:RAB

[TT15] Mary Clair Thompson and Tin-Yau Tam. Rutishauser’s LR algorithm and Bruhat iteration. Linear Multilinear Algebra, 63(10):2061–2070, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Torres:2018:SCH

[TT18] Patrick K. Torres and Michael J. Tsatsomeros. Stability and convex hulls of matrix powers. Linear Multilinear Algebra, 66(4):769–775, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Turcotte:2002:SRA

[Tur02] Daniel Turcotte. Spectral representation of analytic diagonalizable matrix-valued functions that commute with their derivative. Linear Multilinear Algebra, 50(2):181–183, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Turnsek:2016:CTS

[Tur16] Aleksej Turnsek. Circular two-sided multiplications. Linear Multilinear Algebra, 64(11):2169–2176, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tura:2018:CEC

[Tur18] Fernando Tura. Q-cospectrality and L-energy in cographs. Linear Multilinear Algebra, 66(2):398–409, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2017:GAG

[TW17] Fenglei Tian and Dein Wong. Generators of the automorphism group of a regular graph over a ring. Linear Multilinear Algebra, 65(5):1045–1052, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2019:UBE

[TW19] Fenglei Tian and Dein Wong. Upper bounds of the energy of triangle-free graphs in terms of matching number. Linear Multilinear Algebra, 67(1):20–28, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tao:2022:ALT

[TWK22] J. Tao, G. Q. Wang, and L. Kong. The Araki–Lieb–Thirring inequality and the Golden–Thompson inequality in Euclidean Jordan algebras. Linear Multilinear Algebra, 70(19):4228– 4243, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2016:CSP

[TWZ16] Fenglei Tian, Dengyin Wang, and Min Zhu. A characterization of signed planar graphs with rank at most 4. Linear Multilinear Algebra, 64(5):807–817, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tang:2013:DDU

[TX13] Peipei Tang and Aimin Xu. Divided differences and a unified approach to evaluating determinants involving generalized factorials. Linear Multilinear Algebra, 61(1):59–71, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tang:2008:NIP

[TXC08] Xiao-Min Tang, Jin-Li Xu, and Chong-Guang Cao. A note on idempotence-preserving maps. Linear Multilinear Algebra, 56(4):399–414, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tang:2021:NRP

[TXL21] Ling Tang, Chuanfu Xiao, and Hanyu Li. A new random perturbation interval of symmetric eigenvalue problem. Linear Multilinear Algebra, 69(1):147–154, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tian:2021:SLS

[TYC21] Gui-Xian Tian, Ping-Kang Yu, and Shu-Yu Cui. The signless Laplacian state transfer in coronas. Linear Multilinear Algebra, 69(2):278–295, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tang:2006:AAP

[TZ06] Xiao-Min Tang and Xian Zhang. Additive adjoint preservers between matrix spaces. Linear Multilinear Algebra, 54(4):285–

Tang:2017:ETT

[TZ17] Gaohua Tang and Yiqiang Zhou. An embedding theorem on triangular matrix rings. Linear Multilinear Algebra, 65(5): 882–890, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:2019:SDS

[TZ19a] Tin-Yau Tam and Pingping Zhang. Spectral decomposition of selfadjoint matrices in positive semidefinite inner product spaces and its applications. Linear Multilinear Algebra, 67(9): 1829–1838, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tang:2019:BPG

[TZ19b] Xiaomin Tang and Yongyue Zhong. Biderivations of the planar Galilean conformal algebra and their applications. Linear Multilinear Algebra, 67(4):649–659, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tsatsomeros:2021:FMR

[TZ21] Michael J. Tsatsomeros and Yueqiao Faith Zhang. The fibre of P -matrices: the recursive construction of all matrices with positive principal minors. Linear Multilinear Algebra, 69(2): 224–232, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tang:2019:MCR

[TZS19] Gaohua Tang, Yiqiang Zhou, and Huadong Su. Matrices over a commutative ring as sums of three idempotents or three involutions. Linear Multilinear Algebra, 67(2):267–277, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Usta:2022:NMG

[UB22] Fuat Usta and Ömür Betus. A new modification of Gamma operators with a better error estimation. Linear Multilinear Algebra, 70(11):2198–2209, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Urure:2022:LII

[UF22] Ronald Ismael Quispe Urure and Willian Franca. Lie identities and images of Lie polynomials for the skew-symmetric elements of UT_{m}. Linear Multilinear Algebra, 70(19):3643–3659, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Urure:2021:CPI

[UG21] Ronald Ismael Quispe Urure and Dimas José Gonçalves. Central polynomials with involution for the algebra of 2 × 2 upper triangular matrices. Linear Multilinear Algebra, 69(10): 1889–1901, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Uhlig:1979:RCP

[Uhl79] Frank Uhlig. A rational canonical pair form for a pair of symmetric matrices over an arbitrary field F with charF ≠ 2 and applications to finest simultaneous block diagonalizations. Linear and Multilinear Algebra, 8(1):41–67, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Uhlig:1982:NSP

[Uhl82] Frank Uhlig. A note on semibounded pairs of inner products. Linear and Multilinear Algebra, 11(2):179–181, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Uhlig:1999:GPR

[Uhl99] Frank Uhlig. General polynomial roots and their multiplicities in O(N) memory and O(N^{2}) time. Linear and Multilinear Algebra, 46(4):327–359, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Uhlig:2013:CMS

[Uhl13] Frank Uhlig. Computing matrix symmetrizers, finally possible via the Huang and nong algorithm. Linear Multilinear Algebra, 61(7):954–969, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Uhlig:2014:FMA

[Uhl14] Frank Uhlig. Faster and more accurate computation of the field of values boundary for n by n matrices. Linear Multilinear

Uhlig:2020:ZNN

[Uhl20] Frank Uhlig. Zhang Neural Networks for fast and accurate computations of the field of values. Linear Multilinear Algebra, 68(9):1894–1910, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Uc:2015:QLC

[UÖÖ15] M. Uç, H. Özdemir, and A. Y. Özban. On the quadraticity of linear combinations of quadratic matrices. Linear Multilinear Algebra, 63(6):1125–1137, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Uc:2016:GQL

[UPÖ16] Mahmut Uç, Tuğba Petik, and Halim Özdemir. The generalized quadraticity of linear combinations of two commuting quadratic matrices. Linear Multilinear Algebra, 64(9):1696– 1715, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Urschel:2016:MES

[UZ16] John C. Urschel and Ludmil T. Zikatanov. On the maximal error of spectral approximation of graph bisection. Linear Multilinear Algebra, 64(10):1972–1979, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vanstone:1995:CHF

[Van95] Ray Vanstone. Cayley–Hamilton formulae. Linear and Multilinear Algebra, 38(3):163–170, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vaserstein:1986:EIM

[Vas86] Leonid N. Vaserstein. Every integral matrix is the sum of three squares. Linear and Multilinear Algebra, 20(1):1–4, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vaserstein:1987:SPM

[Vas87] L. N. Vaserstein. On the sum powers of matrices. Linear and Multilinear Algebra, 21(3):261–270, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

vonBelow:1995:NPR

[vBW95] Joachim von Below and Renate Wagner. A note on the Perron root of a weighted t-th order mean of nonnegative matrices. Linear and Multilinear Algebra, 40(2):103–109, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

vanDorsselaer:1996:ACS

[vD96] J. L. M. van Dorsselaer. An algorithm to compute sets enclosing M-numerical ranges with applications in numerical analysis. Linear and Multilinear Algebra, 41(1):49–61, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vo:1985:RUP

[VDW85] Kiem-Phong Vo, Wayne E. Dick, and S. Gill Williamson. Ranking and unranking planar embeddings. Linear and Multilinear Algebra, 18(1):35–65, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

vomEnde:2022:SPM

[vE22] Frederik vom Ende. Strict positivity and D-majorization. Linear Multilinear Algebra, 70(19):4023–4048, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vein:1982:PDB

[Vei82a] P. R. Vein. Persymmetric determinants 3: a basic determinant. Linear and Multilinear Algebra, 11(4):305–315, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vein:1982:PDF

[Vei82b] P. R. Vein. Persymmetric determinants: Families of distinct submatrices with non-distinct determinants. Linear and Multilinear Algebra, 11(3):267–276, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vein:1982:PDD

[Vei82c] P. R. Vein. Persymmetric determinants: The derivatives of determinants with Appell function elements. Linear and Multilinear Algebra, 11(3):253–265, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vein:1983:PDA

[Vei83a] P. R. Vein. Persymmetric determinants 4. An alternative form of the Yamazaki–Hori determinantal solution of the Ernst equation. Linear and Multilinear Algebra, 12(4):329– 339, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vein:1983:PDF

[Vei83b] P. R. Vein. Persymmetric determinants 5. Families of overlapping coaxial equivalent determinants. Linear and Multilinear Algebra, 14(2):131–141, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Velanga:2018:IPB

[Vel18] T. Velanga. Ideals of polynomials between Banach spaces revisited. Linear Multilinear Algebra, 66(11):2328–2348, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Velanga:2021:MMV

[Vel21] T. Velanga. Multilinear mappings versus homogeneous polynomials and a multipolynomial polarization formula. Linear Multilinear Algebra, 69(9):1625–1634, 2021. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ventura:2019:RRB

[Ven19a] Emanuele Ventura. Real rank boundaries and loci of forms. Linear Multilinear Algebra, 67(7):1404–1419, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Venturelli:2019:PTS

[Ven19b] Federico Venturelli. Prehomogeneous tensor spaces. Linear Multilinear Algebra, 67(3):510–526, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Veselic:1980:OLQ

[Ves80] K. Veselić. On optimal linearisations of a quadratic eigenvalue problem. Linear and Multilinear Algebra, 8(3):253–258, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

VanDam:1995:EDG

[VH95] Edwin R. Van Dam and Willem H. Haemers. Eigenvalues and the diameter of graphs. Linear and Multilinear Algebra, 39(1– 2):33–44, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vinroot:2004:F

[Vin04] C. Ryan Vinroot. A factorization in GSp(V). Linear Multilinear Algebra, 52(6):385–403, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vinroot:2006:NOS

[Vin06] C. Ryan Vinroot. A note on orthogonal similitude groups. Linear Multilinear Algebra, 54(6):391–396, 2006. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vladeva:2022:DUT

[Vla22] D. I. Vladeva. Derivations of upper triangular matrix semirings. Linear Multilinear Algebra, 70(4):625–641, 2022. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vladeva:2023:DTM

[Vla23] D. I. Vladeva. Derivations of triangular matrix rings. Linear Multilinear Algebra, 71(9):1450–1461, 2023. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vecchio:2007:BIN

[VM07] Antonia Vecchio and Ranjan K. Mallik. Bounds on the inverses of nonnegative triangular matrices with monotonicity properties. Linear Multilinear Algebra, 55(4):365–379, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vosough:2018:SMI

[VM18] Mehdi Vosough and Mohammad Sal Moslehian. Solvability of the matrix inequality. Linear Multilinear Algebra, 66(9): 1799–1818, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vo:1983:DOU

[Vo83a] Kiem-Phong Vo. Determining outerplanarity using segment graphs. Linear and Multilinear Algebra, 13(4):333–343, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vo:1983:FTC

[Vo83b] Kiem-Phong Vo. Finding triconnected components of graphs. Linear and Multilinear Algebra, 13(2):143–165, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vo:1983:SGD

[Vo83c] Kiem-Phong Vo. Segment graphs, depth-first cycle bases, 3connectivity, and planarity of graphs. Linear and Multilinear Algebra, 13(2):119–141, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Volkov:2022:AEA

[Vol22] Yury Volkov. Anticommutative Engel algebras of the first five levels. Linear Multilinear Algebra, 70(1):148–175, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

vonRosen:2011:ICP

[vR11] Tatjana von Rosen. On the inverse of certain patterned sums of matrices with Kronecker product structures. Linear Multilinear Algebra, 59(5):595–606, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Vrba:1986:PSL

[Vrb86] Antonín Vrba. Principal subpermanents of the Laplacian matrix. Linear and Multilinear Algebra, 19(4):335–346, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

vanWyk:2019:LSM

[vWZ19] Leon van Wyk and Michał Ziembowski. Lie solvability in matrix algebras. Linear Multilinear Algebra, 67(4):777–798, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Verma:2021:ISA

[VY21] Ashish Verma and Sarasvati Yadav. On the incomplete second Appell hypergeometric matrix functions. Linear Multilinear Algebra, 69(9):1747–1760, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wadsworth:1985:DCT

[Wad85] Adrian R. Wadsworth. Discriminants in characteristic two. Linear and Multilinear Algebra, 17(3–4):235–263, 1985. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wadsworth:1990:AGT

[Wad90] Adrian R. Wadsworth. The algebra generated by two commuting matrices. Linear and Multilinear Algebra, 27(3):159–162, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wagner:1987:CSB

[Wag87] Richard C. Wagner. Cancellation of symmetric bilinear forms over Dedekind domains. Linear and Multilinear Algebra, 20(3): 183–189, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wallis:1973:CHM

[Wal73] Jennifer Wallis. Complex Hadamard matrices. Linear and Multilinear Algebra, 1(3):257–272, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wall:1975:DSM

[Wal75a] J. R. Wall. The doubly stochastic matrix equations XA^{t}Y = A and XA^{t}Y = A^{t}. Linear and Multilinear Algebra, 2(4):337– 340, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wallis:1975:CWT

[Wal75b] Jennifer Seberry Wallis. Construction of Williamson type matrices. Linear and Multilinear Algebra, 3(3):197–207, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wales:1979:CBF

[Wal79a] David B. Wales. Connections between finite linear groups and linear algebra. Linear and Multilinear Algebra, 7(4):267–279, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wall:1979:RFP

[Wal79b] J. R. Wall. Rank factorizations of positive operators. Linear and Multilinear Algebra, 8(2):137–144, 1979. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wallis:1986:EMH

[Wal86] W. D. Wallis. Embedding matrices in Hadamard matrices. Linear and Multilinear Algebra, 19(4):387–388, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Walker:2011:BSL

[Wal11] Stephen G. Walker. Bounds for the second largest eigenvalue of a transition matrix. Linear Multilinear Algebra, 59(7):755– 760, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:1976:DSD

[Wan76] Edward T. H. Wang. Diagonal sums of doubly stochastic matrices. Linear and Multilinear Algebra, 4(3):217–228, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:1977:CMM

[Wan77] Edward T. H. Wang. On a conjecture of M. Marcus and H. Minc. Linear and Multilinear Algebra, 5(2):145–148, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). The conjecture appears in Proc. Cambridge Philos. Soc. 63, 305–309 (1967).

Wang:1979:CAF

[Wan79] Edward T. H. Wang. A combinatorial application of the Frobenius inequality on rank function to maximum set of commuting nilpotent matrices. Linear and Multilinear Algebra, 8(1): 79–82, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:1980:KPP

[Wan80] Kai Wang. Kronecker products of permutations and its applications to group matrices. Linear and Multilinear Algebra, 9 (1):69–76, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:1982:SNS

[Wan82] Bo-Ying Wang. Some necessary and sufficient conditions satisfied by decomposable tensors in Grassmann spaces. Linear and Multilinear Algebra, 11(3):289–299, 1982. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wanless:2005:PMS

[Wan05] Ian M. Wanless. Permanents of matrices of signed ones. Linear Multilinear Algebra, 53(6):427–433, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wanless:2007:MSC

[Wan07] Ian M. Wanless. On Minc’s sixth conjecture. Linear Multilinear Algebra, 55(1):57–63, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2015:LSS

[Wan15] Hongxing Wang. On least squares solutions subject to a rank restriction. Linear Multilinear Algebra, 63(2):264–273, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:ASR

[Wan16a] Long Wang. Automorphisms of the subgroup-relation graph over the group of unit upper triangular matrices. Linear Multilinear Algebra, 64(5):955–967, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:BTR

[Wan16b] Yu Wang. Biderivations of triangular rings. Linear Multilinear Algebra, 64(10):1952–1959, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:LSS

[Wan16c] Yu Wang. Lie superderivations of superalgebras. Linear Multilinear Algebra, 64(8):1518–1526, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:NCT

[Wan16d] Yu Wang. Notes on centralizing traces and Lie triple isomorphisms on triangular algebras. Linear Multilinear Algebra, 64(5):863–869, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2017:NMF

[Wan17a] Long Wang. Nonlinear mappings on full matrices derivable at zero point. Linear Multilinear Algebra, 65(11):2324–2332, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2017:TLS

[Wan17b] Xue-Feng Wang. Total least squares problem with the arbitrary unitarily invariant norms. Linear Multilinear Algebra, 65(3):438–456, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2018:NJD

[Wan18] Yu Wang. A note on Jordan σ-derivations of triangular algebras. Linear Multilinear Algebra, 66(3):639–644, 2018. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2019:RBR

[Wan19a] Shujing Wang. Relation between the rank of a signed graph and the rank of its underlying graph. Linear Multilinear Algebra, 67(12):2520–2539, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2019:FIU

[Wan19b] Yu Wang. Functional identities in upper triangular matrix rings revisited. Linear Multilinear Algebra, 67(2):348–359, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2019:IMPb

[Wan19c] Yu Wang. The images of multilinear polynomials on 2 × 2 upper triangular matrix algebras. Linear Multilinear Algebra, 67(11):2366–2372, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2020:LDT

[Wan20] Yu Wang. Lie derivations of triangular algebras without assuming unity. Linear Multilinear Algebra, 68(10):2069–2086, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2021:NGT

[Wan21] Long Wang. Nullity of a graph in terms of path cover number. Linear Multilinear Algebra, 69(10):1902–1908, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:EBS

[Wan22a] Jie Wang. Error bounds for the solution sets of generalized polynomial complementarity problems. Linear Multilinear Algebra, 70(19):3982–3993, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:MSR

[Wan22b] Wen-Huan Wang. The minimum spectral radius of the runiform supertree having two vertices of maximum degree. Linear Multilinear Algebra, 70(15):2898–2918, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Watkins:1977:PFP

[Wat77] William Watkins. Polynomial functions that preserve commuting pairs of matrices. Linear and Multilinear Algebra, 5 (2):87–90, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Watkins:1982:LTP

[Wat82] William Watkins. Linear transformations that preserve a similarity class of matrices. Linear and Multilinear Algebra, 11 (1):19–22, 1982. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Waterhouse:1983:ILM

[Wat83] William C. Waterhouse. Invertibility of linear maps preserving matrix invariants. Linear and Multilinear Algebra, 13(2):105– 113, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Waterhouse:1984:BR

[Wat84a] William C. Waterhouse. Book review. Linear and Multilinear Algebra, 15(3–4):281–283, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Waterhouse:1984:SMU

[Wat84b] William C. Waterhouse. Similarity of matrices under SL(n, K). Linear and Multilinear Algebra, 16(1–4):141–145, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Waterhouse:1985:CSM

[Wat85a] William C. Waterhouse. Circulant-style matrices closed under multiplication. Linear and Multilinear Algebra, 18(3):197–206, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Waterhouse:1985:LTP

[Wat85b] William C. Waterhouse. On linear transformations preserving rank one matrices over commutative rings. Linear and Multilinear Algebra, 17(2):101–106, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Watkins:1986:LCB

[Wat86] William Watkins. Littlewood’s correspondence between Schur polynomials and immanants. Linear and Multilinear Algebra, 20(1):75–86, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Waterhouse:1989:TGG

[Wat89] William C. Waterhouse. Two generators for the general linear groups over finite fields. Linear and Multilinear Algebra, 24(4): 227–230, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Watkins:1990:LMG

[Wat90] William Watkins. The Laplacian matrix of a graph: unimodular congruence. Linear and Multilinear Algebra, 28(1–2):35–43, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2014:SLS

[WBZ14] JianFeng Wang, Francesco Belardo, and QiangLong Zhang. Signless Laplacian spectral characterization of line graphs of T -shape trees. Linear Multilinear Algebra, 62(11):1529–1545, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2012:SPN

[WC12] Deyu Wu and Alatancang Chen. Spectral properties and numerical range of off-diagonal infinite-dimensional Hamiltonian operators. Linear Multilinear Algebra, 60(5):613–619, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2015:FRP

[WC15] Long Wang and Jianlong Chen. Further results on partial ordering and the generalized inverses. Linear Multilinear Algebra, 63(12):2419–2429, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:IPU

[WC16] Xingping Wang and Zhaolin Cheng. Infinite products of uniformly paracontracting matrices. Linear Multilinear Algebra, 64(5):856–862, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:PGS

[WC22a] Dandan Wang and Xiwang Cao. Pretty good state transfer on Cayley graphs over semi-dihedral groups. Linear Multilinear Algebra, 70(20):5716–5731, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:DME

[WC22b] Xiaoyuan Wang and Wenchang Chu. Determinants with matrix entries consisting of polynomial differences. Linear Mul-

Wang:2021:CUP

[WCS^{+}21] Kai Wang, Lin Chen, Yi Shen, Yize Sun, and Li-Jun Zhao. Constructing 2×2×4 and 4×4 unextendible product bases and positive-partial-transpose entangled states. Linear Multilinear Algebra, 69(1):131–146, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2017:PEO

[WD17] Xiunan Wang and Chunyuan Deng. Properties of m-EP operators. Linear Multilinear Algebra, 65(7):1349–1361, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2010:SRC

[WDJ10] Jichun Wu, Hanyuan Deng, and Qin Jiang. On the spectral radius of cacti with k-pendant vertices. Linear Multilinear Algebra, 58(3):391–398, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Weaver:1983:EEC

[Wea83] James R. Weaver. Eigenvalues and eigenvectors of a certain stochastic matrix II. Linear and Multilinear Algebra, 13(4): 351–355, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Webster:1983:HMD

[Web83] R. J. Webster. The harmonic means of diagonals of doublystochastic matrices. Linear and Multilinear Algebra, 13(4): 367–369, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Weintraub:1993:NLT

[Wei93] Steven H. Weintraub. A note on linear transformations over integral domains. Linear and Multilinear Algebra, 35(3–4): 295–297, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wei:1998:EDI

[Wei98] Yimin Wei. Expressions for the Drazin inverse of a 2 × 2 block matrix. Linear and Multilinear Algebra, 45(2–3):131–

Wenchang:2006:FBF

[Wen06] Chu Wenchang. The Faà di Bruno formula and determinant identities. Linear Multilinear Algebra, 54(1):1–25, 2006. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Werner:1984:ECR

[Wer84] Hans Joachim Werner. On extensions of Cramer’s rule for solutions of restricted linear systems. Linear and Multilinear Algebra, 15(3–4):319–330, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Werner:1986:GIW

[Wer86] Hans Joachim Werner. Generalized inversion and weak bicomplementarity. Linear and Multilinear Algebra, 19(4):357– 372, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Werner:1987:SRR

[Wer87] Hans Joachim Werner. Some recent results on Drazin monotonicty of property n martices. Linear and Multilinear Algebra, 21(3):243–251, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Werner:1994:CMS

[Wer94] Hans Joachim Werner. Characterizations of minimal semipositivity. Linear and Multilinear Algebra, 37(4):273–278, 1994. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1974:LTG

[Wes74] R. Westwick. Linear transformations on Grassmann spaces III. Linear and Multilinear Algebra, 2(3):257–268, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1975:TNR

[Wes75] R. Westwick. A theorem on numerical range. Linear and Multilinear Algebra, 2(4):311–315, 1975. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1980:MLR

[Wes80] R. Westwick. The maximal length of real trivectors of rank seven. Linear and Multilinear Algebra, 9(1):1–4, 1980. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1981:RTRb

[Wes81a] R. Westwick. Real trivectors of rank seven. Linear and Multilinear Algebra, 10(3):183–204, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1981:RTRa

[Wes81b] R. Westwick. A representation theorem for real or complex bivectors. Linear and Multilinear Algebra, 10(2):89–91, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1987:SMF

[Wes87] R. Westwick. Spaces of matrices of fixed rank. Linear and Multilinear Algebra, 20(2):171–174, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1990:ECR

[Wes90] R. Westwick. Examples of constant rank spaces. Linear and Multilinear Algebra, 28(3):155–174, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1995:TTS

[Wes95] R. Westwick. Transformations on tensor spaces II. Linear and Multilinear Algebra, 40(1):81–92, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Westwick:1996:DUF

[Wes96] R. Westwick. Decomposability under field extensions. Linear and Multilinear Algebra, 41(3):251–253, 1996. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:NPB

[WfMR16] Xue-Zhong Wang, Hai feng Ma, and Jovana Nikolov Radenković. A note on the perturbation bounds of W - weighted Drazin inverse. Linear Multilinear Algebra, 64(10):

Wang:1991:HSC

[WG91a] Bo-Ying Wang and Ming-Peng Gong. A high symmetry class of tensors with an orthogonal basis of decomposable symmetrized tensors. Linear and Multilinear Algebra, 30(1–2): 61–64, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:1991:SOB

[WG91b] Bo-Ying Wang and Ming-Peng Gong. The subspaces and orthonormal bases of symmetry classes of tensors. Linear and Multilinear Algebra, 30(3):195–204, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2011:MSL

[WH11] Jianfeng Wang and Qiongxiang Huang. Maximizing the signless Laplacian spectral radius of graphs with given diameter or cut vertices. Linear Multilinear Algebra, 59(7):733–744, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2012:SME

[WH12] Qing-Wen Wang and Zhuo-Heng He. Some matrix equations with applications. Linear Multilinear Algebra, 60(11–12): 1327–1353, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2022:LRI

[WHC22] Xiufeng Wu, Junjie Huang, and Alatancang Chen. On the left (right) invertibility of operator matrices. Linear Multilinear Algebra, 70(22):7836–7855, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2017:TAB

[WHG17] Yan Wang, Bo Hou, and Suogang Gao. Totally almost bipartite Leonard pairs and Leonard triples of q-Racah type. Linear Multilinear Algebra, 65(2):235–255, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

White:1974:LMA

[Whi74] Dennis E. White. Linear and multilinear aspects of isomorph rejection. Linear and Multilinear Algebra, 2(3):211–226, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

White:1975:MET

[Whi75] Dennis E. White. Multilinear enumerative techniques. Linear and Multilinear Algebra, 2(4):341–352, 1975. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

White:1979:MTP

[Whi79] Dennis E. White. Multilinear techniques in Pólya enumeration theory. Linear and Multilinear Algebra, 7(4):299–315, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

White:1980:ABT

[Whi80] Dennis E. White. Applications of a binomial-type identity. Linear and Multilinear Algebra, 8(4):341–345, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:HMS

[WHL16] Bangyan Wang, Shaofang Hong, and Mao Li. Hyperdeterminants on meet-semilattices. Linear Multilinear Algebra, 64 (10):1899–1915, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:SOE

[WHL22] Hua Wang, Junjie Huang, and Mengran Li. On the solutions of the operator equation XAX = BX. Linear Multilinear Algebra, 70(22):7753–7761, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2008:JDC

[WHX08] Dengyin Wang, Qian Hu, and Chunguang Xia. Jordan derivation of certain Jordan matrix algebras. Linear Multilinear Algebra, 56(5):581–588, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Williamson:1973:TCL

[Wil73] S. G. Williamson. Tensor compositions and lists of combinatorial structures. Linear and Multilinear Algebra, 1(2):119–138, 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Williamson:1981:SOP

[Wil81] S. G. Williamson. Symmetry operators, polarizations, and a generalized Capelli identity. Linear and Multilinear Algebra, 10(2):93–102, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wilson:1983:TEP

[Wil83] Gregory L. Wilson. Three enumeration problems for a congruence equation. Linear and Multilinear Algebra, 14(1):97–110, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Williamson:1987:GCM

[Wil87] S. Gill Williamson. Generic common minor expansions. Linear and Multilinear Algebra, 20(3):253–279, 1987. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Williamson:1993:CFC

[Wil93] S. Gill Williamson. Canonical forms for cycles in bridge graphs. Linear and Multilinear Algebra, 34(3–4):301–341, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Williamson:1999:SBT

[Wil99] S. Gill Williamson. Straightening bases for tensor products. Linear and Multilinear Algebra, 45(4):375–392, 1999. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wimmer:1985:WMI

[Wim85] Harald K. Wimmer. On a weighted mean inequality for nonnegative symmetric matrices. Linear and Multilinear Algebra, 17(1):25–27, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wimmer:1988:FGB

[Wim88] Harald K. Wimmer. A factorization of the generalized Bezoutian of polynomial matrices. Linear and Multilinear Algebra, 23(3):255–261, 1988. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wisbauer:2012:CSM

[Wis12] Robert Wisbauer. Coalgebraic structures in module theory. Linear Multilinear Algebra, 60(7):829–853, 2012. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Witowicz:2011:DMM

[Wit11] Pawel Witowicz. Divisibility of multilinear mappings with applications to affine differential geometry. Linear Multilinear Algebra, 59(8):879–890, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:MPL

[WJ16] Meili Wang and Guoxing Ji. Maps preserving ∗-Lie product on factor von Neumann algebras. Linear Multilinear Algebra, 64(11):2159–2168, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2015:CCI

[WL15] Hongxing Wang and Xiaoji Liu. Characterizations of the core inverse and the core partial ordering. Linear Multilinear Algebra, 63(9):1829–1836, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:ASM

[WL16a] Shujuan Wang and Wende Liu. The abelian subalgebras of maximal dimensions for general linear Lie superalgebras. Linear Multilinear Algebra, 64(10):2081–2089, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:MDF

[WL16b] Shujuan Wang and Wende Liu. Minimal dimensions of faithful representations for abelian Lie superalgebras. Linear Multilinear Algebra, 64(7):1354–1361, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2016:NMM

[WL16c] Shi-Liang Wu and Cui-Xia Li. A note on the MSMAOR method for linear complementarity problems. Linear Multilinear Algebra, 64(4):795–800, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2018:POS

[WL18a] Hongxing Wang and Xiaoji Liu. A partial order on the set of complex matrices with index one. Linear Multilinear Algebra, 66(1):206–216, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2018:PNM

[WL18b] Tingzeng Wu and Hong-Jian Lai. On the permanental nullity and matching number of graphs. Linear Multilinear Algebra, 66(3):516–524, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2020:END

[WL20a] Hongxing Wang and Xiaoji Liu. EP-nilpotent decomposition and its applications. Linear Multilinear Algebra, 68(8):1682– 1694, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wei:2020:RBH

[WL20b] Wei Wei and Shuchao Li. Relation between the Hermitian energy of a mixed graph and the matching number of its underlying graph. Linear Multilinear Algebra, 68(7):1395–1410, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2022:PTS

[WL22] Shi-Liang Wu and Cui-Xia Li. A preconditioned two-sweep shift splitting method for non-Hermitian positive definite linear systems. Linear Multilinear Algebra, 70(21):7075–7092, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2023:IIC

[WL23] Qi Wu and Yong Lu. Inertia indices of a complex unit gain graph in terms of matching number. Linear Multilinear Alge-

Wang:2019:IMPa

[WLB19] Yu Wang, Panpan Liu, and Jie Bai. The images of multilinear polynomials on 2 × 2 upper triangular matrix algebras. Linear Multilinear Algebra, 67(11):i–vi, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:SMI

[WLF22] Yi Wang, Wang Le, and Yi-Zheng Fan. A spectral method to incidence balance of oriented hypergraphs and induced signed hypergraphs. Linear Multilinear Algebra, 70(19):4804–4818, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2019:IBM

[WLL19] Yanfeng Wang, Feng Lü, and Weiran Lü. The inverse of bisymmetric matrices. Linear Multilinear Algebra, 67(3):479–489, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2021:CIA

[WLL21] Hongxing Wang, Ning Li, and Xiaoji Liu. The ⇕-core inverse and its applications. Linear Multilinear Algebra, 69(13):2491– 2509, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wei:2021:BNR

[WLM21] Wei Wei, Shuchao Li, and Hongping Ma. Bounds on the nullity, the H-rank and the Hermitian energy of a mixed graph. Linear Multilinear Algebra, 69(13):2469–2490, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2023:LPS

[WLW23] Wei Wang, Xiaogang Liu, and Jing Wang. Laplacian pair state transfer in vertex coronas. Linear Multilinear Algebra, 71(14): 2282–2297, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:PVT

[WLX16] Mingyue Wang, Xian Li, and Changqing Xu. Positivities of Vandermonde tensors. Linear Multilinear Algebra, 64(4):767–

Wang:2022:LSS

[WLZ22] Qing-Wen Wang, Ru-Yuan Lv, and Yang Zhang. The leastsquares solution with the least norm to a system of tensor equations over the quaternion algebra. Linear Multilinear Algebra, 70(10):1942–1962, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wong:1984:ODM

[WM84] Chi Song Wong and Joseph C. Masaro. A-optimal design matrices X = (x_{ij})_{N×n} with x_{ij} = −1, 0, 1. Linear and Multilinear Algebra, 15(1):23–46, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2017:CSL

[WM17] Long Wang and Xiaobin Ma. Characterization of sub-long graphs of arbitrary rank. Linear Multilinear Algebra, 65(7): 1427–1445, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2021:OSI

[WM21] Long Wang and Dijana Mosić. The one-sided inverse along two elements in rings. Linear Multilinear Algebra, 69(13):2410– 2422, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2021:BSR

[WMFW21] Zhao Wang, Yaping Mao, Boris Furtula, and Xu Wang. Bounds for the spectral radius and energy of extended adjacency matrix of graphs. Linear Multilinear Algebra, 69(10): 1813–1824, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wilson:1974:BR

[WNM74] Elizabeth Wilson, Morris Newman, and Russell Merris. Book reviews. Linear and Multilinear Algebra, 2(1):91–95, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Woerdeman:1993:MRC

[Woe93] Hugo J. Woerdeman. Minimal rank completions of partial banded matrices. Linear and Multilinear Algebra, 36(1):59–68, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Woerdeman:1997:HNC

[Woe97] Hugo J. Woerdeman. Hermitian and normal completions. Linear and Multilinear Algebra, 42(3):239–280, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Woerdeman:2008:HRN

[Woe08] Hugo J. Woerdeman. The higher rank numerical range is convex. Linear Multilinear Algebra, 56(1–2):65–67, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Woodroofe:2022:AGP

[Woo22] Russ Woodroofe. An algebraic groups perspective on Erdős– Ko–Rado. Linear Multilinear Algebra, 70(22):7825–7835, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Worz:2020:NBV

[Wör20] Sascha Wörz. A note on Bernstein–Vandermonde matrices. Linear Multilinear Algebra, 68(3):458–471, 2020. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2009:ALA

[WOW09] Dengyin Wang, Shikun Ou, and Xuansheng Wang. Algebraic Lie algebras with complete Borel subalgebras. Linear Multilinear Algebra, 57(8):817–821, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2006:DIL

[WOY06] Dengyin Wang, Shikun Ou, and Qiu Yu. Derivations of the intermediate Lie algebras between the Lie algebra of diagonal matrices and that of upper triangular matrices over a commutative ring. Linear Multilinear Algebra, 54(5):369–377, 2006. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wright:1993:JCL

[Wri93] David Wright. The Jacobian conjecture: linear triangularizatlon for cubics in dimension three. Linear and Multilinear Algebra, 34(2):85–97, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wright:2022:CCR

[Wri22a] Stephen E. Wright. Constructing and counting regular triangular matrices. Linear Multilinear Algebra, 70(19):3936–3943, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wright:2022:TIM

[Wri22b] Stephen E. Wright. Triangular idempotent matrices over a general ring. Linear Multilinear Algebra, 70(9):1717–1731, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2018:NEB

[WS18] Feng Wang and Deshu Sun. New error bound for linear complementarity problems for B-matrices. Linear Multilinear Algebra, 66(11):2156–2167, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2023:APN

[WS23] Hua-Ming Wang and Hongyan Sun. Asymptotics of product of nonnegative 2-by-2 matrices with applications to random walks with asymptotically zero drifts. Linear Multilinear Algebra, 71(2):150–177, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2011:LSC

[WSH^{+}11] J. F. Wang, S. K. Simić, Q. X. Huang, F. Belardo, and E. M. Li Marzi. Laplacian spectral characterization of disjoint union of paths and cycles. Linear Multilinear Algebra, 59(5):531–539, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2017:NPC

[WSZL17] Feng Wang, Deshu Sun, Jianxing Zhao, and Chaoqian Li. New practical criteria for h-tensors and its application. Linear

Wang:2016:SCG

[WT16] Dengyin Wang and Fenglei Tian. The signature of k-cyclic graphs of ∞-type. Linear Multilinear Algebra, 64(3):375–382, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2023:DBS

[WT23] Qingyan Wu and Xiaomin Tang. Derivations and biderivations of the Schrödinger algebra in (n + 1)-dimensional space-time. Linear Multilinear Algebra, 71(7):1073–1097, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:PAS

[WTD22] Shuaijie Wang, Xiaoyi Tian, and Chunyuan Deng. On the parallel addition and subtraction of operators on a Hilbert space. Linear Multilinear Algebra, 70(19):3660–3688, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Williamson:1978:BR

[WTM78] S. G. Williamson, R. C. Thompson, and Kenneth Millett. Book reviews. Linear and Multilinear Algebra, 6(3):243–245, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2005:ACD

[Wu05] Chai Wah Wu. Algebraic connectivity of directed graphs. Linear Multilinear Algebra, 53(3):203–223, 2005. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2009:PNO

[Wu09] Honglin Wu. Possible numbers of ones in the squares of 0–1 matrices with a given rank. Linear Multilinear Algebra, 57(6): 609–623, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2022:GSI

[Wu22] Chai Wah Wu. On generalizations of Schur’s inequality on sums of products of differences of real numbers. Linear Multilinear Algebra, 70(21):6452–6468, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

White:1976:PAP

[WW76] Dennis E. White and S. G. Williamson. Probabilistic analogs of Pólya’s enumeration theory. Linear and Multilinear Algebra, 4(1):21–30, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

White:1980:CMR

[WW80] D. E. White and S. G. Williamson. Construction of minimal representative systems. Linear and Multilinear Algebra, 9(3): 167–180, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

White:2001:RPC

[WW01] Dennis E. White and S. Gill Williamson. Research problem: combinatorial and multilinear aspects of sign-balanced posets. Linear and Multilinear Algebra, 49(2):169–181, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2012:RBC

[WW12] Gang Wu and Yimin Wei. Relationship between the characteristic polynomial and the spectrum of a diagonalizable matrix and those of its low-rank update. Linear Multilinear Algebra, 60(8):967–978, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2015:LSC

[WW15] Luhua Wang and Ligong Wang. Laplacian spectral characterization of clover graphs. Linear Multilinear Algebra, 63(12): 2396–2405, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2020:RBN

[WW20a] Yunjie Wang and Gang Wu. Refined bounds on the number of distinct eigenvalues of a matrix after low-rank update. Linear Multilinear Algebra, 68(5):903–914, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wong:2020:NSA

[WW20b] K. L. Wong and M. W. Wong. Normality, self-adjointness, spectral invariance, groups and determinants of pseudodifferential operators on finite abelian groups. Linear Multi-

Wang:2019:ADN

[WWS19] Xinlei Wang, Dein Wong, and Dongqin Sun. Automorphisms and domination numbers of transformation graphs over vector spaces. Linear Multilinear Algebra, 67(7):1350–1363, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2020:SOT

[WWW20] Gang Wang, Yanan Wang, and Yiju Wang. Some Ostrowskitype bound estimations of spectral radius for weakly irreducible nonnegative tensors. Linear Multilinear Algebra, 68(9): 1817–1834, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:MET

[WWWT22] Xinlei Wang, Dein Wong, Liangli Wei, and Fenglei Tian. On the multiplicity of −1 as an eigenvalue of a tree with given number of pendant vertices. Linear Multilinear Algebra, 70 (17):3345–3353, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2020:ZFN

[WWZ20] Xinlei Wang, Dein Wong, and Yuanshuai Zhang. Zero forcing number of a graph in terms of the number of pendant vertices. Linear Multilinear Algebra, 68(7):1424–1433, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2019:IAS

[WX19] Qing-Wen Wang and Xiangjian Xu. Iterative algorithms for solving some tensor equations. Linear Multilinear Algebra, 67(7):1325–1349, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2021:LSS

[WXD21] Qing-Wen Wang, Xiangjian Xu, and Xuefeng Duan. Least squares solution of the quaternion Sylvester tensor equation. Linear Multilinear Algebra, 69(1):104–130, 2021. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2011:LTD

[WY11] Dengyin Wang and Xiaoxiang Yu. Lie triple derivations on the parabolic subalgebras of simple Lie algebras. Linear Multilinear Algebra, 59(8):837–840, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2018:CRR

[WY18] Yali Wu and Yichuan Yang. Cramer’s rule over residue class rings of Bézout domains. Linear Multilinear Algebra, 66(6): 1268–1276, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wei:2020:SRT

[WY20] Jia Wei and Zhifu You. Spectral radius and traceability of graphs with large minimum degree. Linear Multilinear Algebra, 68(1):161–176, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2021:CTE

[WY21] Shaoxin Wang and Hu Yang. Conditioning theory of the equality constrained quadratic programming and its applications. Linear Multilinear Algebra, 69(6):1161–1183, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2011:PNP

[WYM11] Longqin Wang, Lihua You, and Hongping Ma. Primitive nonpowerful sign pattern matrices with base 2. Linear Multilinear Algebra, 59(6):693–700, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wei:2022:SEC

[WYSL22] Jia Wei, Zhifu You, Sulin Song, and Hong-Jian Lai. Spectral and extremal conditions for supereulerian graphs. Linear Multilinear Algebra, 70(20):5995–6017, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:1995:WNM

[WZ95] Bo-Ying Wang and Fuzhen Zhang. Words and normality of matrices. Linear and Multilinear Algebra, 40(2):111–118, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:1997:SCM

[WZ97] Bo-Ying Wang and Fuzhen Zhang. Schur complements and matrix inequalities of Hadamard products. Linear and Multilinear Algebra, 43(1–3):315–326, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2013:SLS

[WZ13] Shilin Wang and Bo Zhou. The signless Laplacian spectra of the corona and edge corona of two graphs. Linear Multilinear Algebra, 61(2):197–204, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2014:OIR

[WZ14] J. Wu and J. Zhao. Operator inequalities and reverse inequalities related to the Kittaneh–Manasrah inequalities. Linear Multilinear Algebra, 62(7):884–894, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2018:DSR

[WZ18] Yanna Wang and Bo Zhou. On distance spectral radius of hypergraphs. Linear Multilinear Algebra, 66(11):2232–2246, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wojcik:2022:OHM

[WZ22] Paweł Wójcik and Ali Zamani. Orthogonality Hilbert Amodules and operators preserving multi-A-linearity. Linear Multilinear Algebra, 70(21):6176–6189, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wojcik:2023:NDO

[WZ23] Paweł Wójcik and Ali Zamani. From norm derivatives to orthogonalities in Hilbert C^{∗}-modules. Linear Multilinear Algebra, 71(6):875–888, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2011:BMU

[WZC11] Dengyin Wang, Huixiang Zhai, and Meilin Chen. Bijective maps on unit upper triangular matrices preserving commutators. Linear Multilinear Algebra, 59(1):25–40, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2016:CSF

[WZD16] Yueqing Wang, Ning Zuo, and Hongke Du. Characterizations of the support function of the numerical range of the product of positive contractions. Linear Multilinear Algebra, 64(10): 2068–2080, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2013:HSSa

[WZH13a] Qing-Wen Wang, Xiang Zhang, and Zhuo-Heng He. On the Hermitian structures of the solution to a pair of matrix equations. Linear Multilinear Algebra, 61(1):73–90, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2013:HSSb

[WZH13b] Qing-Wen Wang, Xiang Zhang, and Zhuo-Heng He. On the Hermitian structures of the solution to a pair of matrix equations. Linear Multilinear Algebra, 61(8):1158, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2013:GCP

[WZL13] Dengyin Wang, Min Zhu, and Wenping Lv. The group of commutativity preserving maps on upper triangular matrices over a commutative ring. Linear Multilinear Algebra, 61(6): 775–783, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wu:2022:MBI

[WZY22] Yu-Jiang Wu, Wei-Hong Zhang, and Ai-Li Yang. Modulusbased inexact non-alternating preconditioned splitting method for linear complementarity problems. Linear Multilinear Algebra, 70(22):7414–7432, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Wang:2022:RSN

[WZZ22] Cun Wang, Jiayin Zhao, and Sen Zhu. Remarks on the structure of C-normal operators. Linear Multilinear Algebra, 70(9): 1682–1696, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2018:PGI

[XCB18] Sanzhang Xu, Jianlong Chen, and Julio Benítez. Projections for generalized inverses. Linear Multilinear Algebra, 66(8):

Xian:1999:AMP

[XCC99] Zhang Xian, Cao Chongguang, and Bu Changjiang. Additive maps preserving M–P inverses of matrices over fields. Linear and Multilinear Algebra, 46(3):199–211, 1999. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2020:NCC

[XCM20] Sanzhang Xu, Jianlong Chen, and Dijana Mosić. New characterizations of the CMP inverse of matrices. Linear Multilinear Algebra, 68(4):790–804, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2014:PIS

[XFG14] Xiao-Ming Xu, Xiaochun Fang, and Fugen Gao. Principal invariant subspaces theorems. Linear Multilinear Algebra, 62 (11):1428–1436, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2020:CMC

[XHL20] Changqing Xu, Lingling He, and Zerong Lin. Commutation matrices and commutation tensors. Linear Multilinear Algebra, 68(9):1721–1742, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2023:GED

[XHL23] Xiao-Ming Xu, Yi Huang, and Yuan Li. General explicit descriptions for two classes of symmetries involving idempotents. Linear Multilinear Algebra, 71(3):457–471, 2023. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xiang:2020:IFM

[Xia20] Zhong-Qi Xiang. Invertibility of g-frame multipliers and Bessel multipliers for unitary systems in Hilbert C∗-modules. Linear Multilinear Algebra, 68(8):1663–1681, 2020. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xiong:2023:MUL

[Xio23] Jin-Song Xiong. Modified upper and lower triangular splitting iterative method for a class of block two-by-two linear systems.

Xie:2020:OPT

[XJS20] Ze-Jia Xie, Xiao-Qing Jin, and Vai-Kuong Sin. An optimal preconditioner for tensor equations involving Einstein product. Linear Multilinear Algebra, 68(5):886–902, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xie:2017:FAS

[XJW17] Ze-Jia Xie, Xiao-Qing Jin, and Yi-Min Wei. A fast algorithm for solving circulant tensor systems. Linear Multilinear Algebra, 65(9):1894–1904, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2011:DAP

[XL11] Jiehong Xu and Mao Li. Divisibility among power GCD matrices and among power LCM matrices on three coprime divisor chains. Linear Multilinear Algebra, 59(7):773–788, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2017:AMR

[XL17] Xiaowei Xu and Hanchao Liu. Additive maps on rank-s matrices. Linear Multilinear Algebra, 65(4):806–812, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xia:2015:FTT

[XLL15] Mengxue Xia, Wei Li, and Haohao Li. Farkas-type theorems for interval linear systems. Linear Multilinear Algebra, 63(7): 1390–1400, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xue:2020:MDS

[XLS20] Jie Xue, Shuting Liu, and Jinlong Shu. On the multiplicity of distance signless Laplacian eigenvalues of graphs. Linear Multilinear Algebra, 68(11):2276–2288, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2016:AMI

[XPY16] Xiaowei Xu, Yue Pei, and Xiaofei Yi. Additive maps on invertible matrices. Linear Multilinear Algebra, 64(7):1283–1294,

Xiong:2015:IPR

[XQ15] Zhiping Xiong and Yingying Qin. An invariance property related to the mixed-type reverse order laws. Linear Multilinear Algebra, 63(8):1621–1634, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xie:2016:SDH

[XQ16a] Jinshan Xie and Liqun Qi. Spectral directed hypergraph theory via tensors. Linear Multilinear Algebra, 64(4):780–794, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xiong:2016:NRO

[XQ16b] Zhiping Xiong and Yingying Qin. A note on the reverse order law for least square g-inverse of operator product. Linear Multilinear Algebra, 64(7):1404–1414, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xiong:2013:LSI

[XQZ13] Zhiping Xiong, Yingying Qin, and Bing Zheng. The least squares g-inverses for sum of matrices. Linear Multilinear Algebra, 61(4):448–462, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2018:ESS

[XS18] Guixin Xu and Yuming Shi. Essential spectra of self-adjoint relations under relatively compact perturbations. Linear Multilinear Algebra, 66(12):2438–2467, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2018:RDI

[XSH18] Qingxiang Xu, Chuanning Song, and Lili He. Representations for the Drazin inverse of an anti-triangular block operator matrix E with ind(E) ≤ 2. Linear Multilinear Algebra, 66(5):1026–1045, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2015:GES

[XSL15] Qingxiang Xu, Chuanning Song, and Xiaofang Liu. General exact solutions of the second-order homogeneous algebraic differential equations. Linear Multilinear Algebra, 63(2):244–263, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xi:2021:DSR

[XSW21] Weige Xi, Wasin So, and Ligong Wang. On the distance spectral radius of digraphs with given diameter. Linear Multilinear Algebra, 69(14):2547–2557, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xi:2022:SRD

[XSW22] Weige Xi, Wasin So, and Ligong Wang. On the A_{α} spectral radius of digraphs with given parameters. Linear Multilinear Algebra, 70(12):2248–2263, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:1993:ISP

[Xu93] Song Xu. On inverse spectrum problems for normal nonnegative matrices. Linear and Multilinear Algebra, 34(3–4):353– 364, 1993. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2015:IIN

[Xu15] Chuang Xu. On the idempotency, involution and nilpotency of a linear combination of two matrices. Linear Multilinear Algebra, 63(8):1664–1677, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2018:CHO

[Xu18] Yangyang Xu. On the convergence of higher-order orthogonal iteration. Linear Multilinear Algebra, 66(11):2247–2265, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2022:NUB

[Xu22] Xuefeng Xu. New upper bounds for the spectral variation of a general matrix. Linear Multilinear Algebra, 70(16):3070–3080, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xiao:2012:NLH

[XW12] Zhankui Xiao and Feng Wei. Nonlinear Lie higher derivations on triangular algebras. Linear Multilinear Algebra, 60(8):979–

Xiao:2019:MSR

[XW19] Peng Xiao and Ligong Wang. The maximum spectral radius of uniform hypergraphs with given number of pendant edges. Linear Multilinear Algebra, 67(7):1392–1403, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xia:2022:TSC

[XW22a] Chao Xia and Xi Wang. Total strong common multiplicity of two matrices and Sylvester resultant. Linear Multilinear Algebra, 70(9):1788–1797, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xia:2022:TVM

[XW22b] Chao Xia and Xi Wang. Tropical Vandermonde matrices. Linear Multilinear Algebra, 70(2):238–249, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2022:SCT

[XW22c] Xiangjian Xu and Qing-Wen Wang. On the solutions of a class of tensor equations. Linear Multilinear Algebra, 70(21): 6141–6154, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xiao:2015:CTL

[XWF15] Zhankui Xiao, Feng Wei, and Ajda Fošner. Centralizing traces and Lie triple isomorphisms on triangular algebras. Linear Multilinear Algebra, 63(7):1309–1331, 2015. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2022:AGI

[XWT22] Feng Xu, Dein Wong, and Fenglei Tian. Automorphism group of the intersection graph of ideals over a matrix ring. Linear Multilinear Algebra, 70(2):322–330, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2022:MPT

[XXF22] Jinli Xu, Yuan Xue, and Ajda Fosner. Maps preserving transition probability from pure product states to pure states. Linear Multilinear Algebra, 70(19):4177–4190, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2011:CMH

[XXZ11] Guanghui Xu, Changqing Xu, and Fuzhen Zhang. Contractive matrices of hua type. Linear Multilinear Algebra, 59(2):159– 172, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2022:DMU

[XY22] Jiao Xu and Yongxin Yuan. A direct method for updating mass and stiffness matrices with submatrix constraints. Linear Multilinear Algebra, 70(19):4266–4281, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2022:DLS

[XYH22] Nannan Xu, Aimei Yu, and Rong-Xia Hao. On the distance Laplacian spectral radius of bicyclic graphs. Linear Multilinear Algebra, 70(19):4654–4674, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xing:2016:SLD

[XZ16] Rundan Xing and Bo Zhou. On the second largest distance eigenvalue. Linear Multilinear Algebra, 64(9):1887–1898, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2018:CTM

[XZ18] Xiaowei Xu and Jianwei Zhu. Central traces of multiadditive maps on invertible matrices. Linear Multilinear Algebra, 66(7): 1442–1448, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xie:2022:IAG

[XZ22] Mengyan Xie and Jiao Zhang. Inequalities for arithmeticgeometric-harmonic means of sector matrices. Linear Multilinear Algebra, 70(20):5922–5931, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xing:2014:DSL

[XZL14] Rundan Xing, Bo Zhou, and Jianping Li. On the distance signless Laplacian spectral radius of graphs. Linear Multilinear Algebra, 62(10):1377–1387, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xue:2019:FMN

[XZS19] Jie Xue, Mingqing Zhai, and Jinlong Shu. Fractional matching number and eigenvalues of a graph. Linear Multilinear Algebra, 67(12):2565–2574, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Xu:2022:ISC

[XZW22] Yi Xu, Mingqing Zhai, and Bing Wang. An improvement of spectral conditions for Hamilton-connected graphs. Linear Multilinear Algebra, 70(21):6096–6107, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yamaguchi:2018:CEG

[Yam18] Naoya Yamaguchi. Capelli elements of the group algebra. Linear Multilinear Algebra, 66(10):2003–2010, 2018. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2021:NLD

[Yan21] Yuping Yang. Nonlinear Lie derivations of incidence algebras of finite rank. Linear Multilinear Algebra, 69(9):1665–1682, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yao:2022:LDW

[Yao22] Yu-Feng Yao. Local derivations on the Witt algebra in prime characteristic. Linear Multilinear Algebra, 70(15):2919–2933, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yaqub:1980:BR

[Yaq80] Adil Yaqub. Book review. Linear and Multilinear Algebra, 8(3):269–308, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:1996:GSN

[YBCW96] Jie-Tai Yu, H. Bass, E. H. Connell, and D. Wright. On generalized strongly nilpotent matrices. Linear and Multilinear Algebra, 41(1):19–22, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2019:BNR

[YC19] Jiahui Yu and Alatancang Chen. Block numerical range and estimable total decompositions of normal operators. Linear Multilinear Algebra, 67(9):1750–1756, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2021:SSO

[YCZ21] Yongxin Yuan, Yinlan Chen, and Lei Zhou. Solutions of a second-order conjugate matrix equation. Linear Multilinear Algebra, 69(9):1645–1656, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yan:2014:CTS

[YD14] Zaili Yan and Shaoqiang Deng. Completable two step nilpotent Lie algebras of type (2, p). Linear Multilinear Algebra, 62(4):445–452, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Ye:2011:KIS

[Ye11] Luzhen Ye. On the Kirchhoff index of some toroidal lattices. Linear Multilinear Algebra, 59(6):645–650, 2011. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2018:WLD

[YF18] Bing Yang and Xiaochun Fang. Weak-2-local derivations on finite von Neumann algebras. Linear Multilinear Algebra, 66(8): 1520–1529, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yamanci:2019:HHT

[YGÇ19] Ulas Yamanci, Mubariz T. Garayev, and Ceren Çelik. Hardy– Hilbert type inequality in reproducing kernel Hilbert space: its applications and related results. Linear Multilinear Algebra, 67(4):830–842, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2015:MIO

[YGW15] Guanglong Yu, Shu-Guang Guo, and Yarong Wu. Maxima of the Q-index for outer-planar graphs. Linear Multilinear Algebra, 63(9):1837–1848, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2022:BCI

[YGZC22] Chuan Yang, Zhihua Guo, Chengyang Zhang, and Huaixin Cao. Broadcasting coherence via incoherent operations. Linear Multilinear Algebra, 70(21):6469–6477, 2022. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yopp:2005:EEF

[YH05] David A. Yopp and Richard D. Hill. Extremals and exposed faces of the cone of positive maps. Linear Multilinear Algebra, 53(3):167–174, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2022:BDT

[YH22] Lamei Yuan and Qianyi Hua. -(bi)derivations and transposed Poisson algebra structures on Lie algebras. Linear Multilinear Algebra, 70(22):7672–7701, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2014:EPT

[YHCW14] Ber-Lin Yu, Ting-Zhu Huang, Hong Cheng, and Dongdong Wang. Eventual positivity of tridiagonal sign patterns. Linear Multilinear Algebra, 62(7):853–859, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2013:BSR

[YHS13] Guanglong Yu, Yuan Hong, and Jinlong Shu. Bounds of spectral radii on edge-most outer-planar bipartite graphs. Linear Multilinear Algebra, 61(6):791–800, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yin:2021:MER

[Yin21] Daijun Yin. On the multiple eigenvalue of Randić matrix of trees. Linear Multilinear Algebra, 69(6):1137–1150, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yizheng:2002:SIV

[Yiz02] Fan Yizheng. On spectral integral variations of graphs. Linear Multilinear Algebra, 50(2):133–142, 2002. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yizheng:2003:SIV

[Yiz03] Fan Yizheng. Spectral integral variations of degree maximal graphs. Linear Multilinear Algebra, 51(2):147–154, 2003. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2021:BDP

[YJ21] Lili Yang and Guoxing Ji. Bounds for the diamond partial order in B(H). Linear Multilinear Algebra, 69(8):1415–1421, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yamanci:2022:FRB

[YK22] Ulas Yamanci and Ismail Murat Karli. Further refinements of the Berezin number inequalities on operators. Linear Multilinear Algebra, 70(20):5237–5246, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2011:LSS

[YL11] Shifang Yuan and Anping Liao. Least squares solution of the quaternion matrix equation with the least norm. Linear Multilinear Algebra, 59(9):985–998, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2019:RRT

[YLC19a] Junjian Yang, Linzhang Lu, and Zhen Chen. A refinement of Rotfel’d type inequality for partitioned matrices with numerical ranges in a sector. Linear Multilinear Algebra, 67(9): 1719–1726, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2019:SND

[YLC19b] Junjian Yang, Linzhang Lu, and Zhen Chen. Schatten qnorms and determinantal inequalities for matrices with numerical ranges in a sector. Linear Multilinear Algebra, 67(2): 221–227, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yeng:1990:DPT

[YLL90] Shwu Yeng, T. Lin, and You-Feng Lin. The n-dimensional Pythagorean theorem. Linear and Multilinear Algebra, 26(1– 2):9–13, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2010:LIW

[YLS10] Hu Yang, Hanyu Li, and Hua Shao. Lavoie inequalities for weighted generalized inverses of matrices. Linear Multilinear Algebra, 58(3):285–295, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2015:LSB

[YLW15] Shi-Fang Yuan, An-Ping Liao, and Peng Wang. Least squares η-bi-Hermitian problems of the quaternion matrix equation (AXB, CXD) = (E, F). Linear Multilinear Algebra, 63(9): 1849–1863, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2022:ECU

[YLW22] Jinxing Yang, Xiaogang Liu, and Ligong Wang. Energies of complements of unitary one-matching bi-Cayley graphs over commutative rings. Linear Multilinear Algebra, 70(20):5618– 5633, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2020:IHO

[YLX20] Ming Yang, Wen Li, and Mingqing Xiao. On identifiability of higher order block term tensor decompositions of rank L_{r}⊗ rank-1. Linear Multilinear Algebra, 68(2):223–245, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2013:GBS

[YMYS13] Guanglong Yu, Zhengke Miao, Chao Yan, and Jinlong Shu. Gaps in the base set of primitive nonpowerful sign patterns. Linear Multilinear Algebra, 61(6):801–810, 2013. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Young:1980:FCM

[You80] N. J. Young. Functions of canonical matrices. Linear and Multilinear Algebra, 9(2):141–149, 1980. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

You:2015:PRW

[You15] Chao You. The parallel refinement of weakening idempotent pair. Linear Multilinear Algebra, 63(12):2440–2447, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yun:1990:DTS

[YQ90] Fan Yun and Huang Qiang. Decomposition of tensor spaces and block idempotents of group rings. Linear and Multilinear Algebra, 26(4):299–305, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2016:SSL

[YqS16] Juan Yu and Shu qian Shen. Solvability of systems of linear matrix equations subject to a matrix inequality. Linear Multilinear Algebra, 64(12):2446–2462, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2023:MSR

[YS23] Xiying Yuan and Zhenan Shao. On the maximal α-spectral radius of graphs with given matching number. Linear Multilinear Algebra, 71(10):1681–1690, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yousef:2015:ONI

[YSK15] A. Yousef, M. Sababheh, and R. Khalil. Operator-norm inequalities and interpolated trace inequalities. Linear Multilinear Algebra, 63(3):455–465, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shu:2014:DFS

[ySpW14] Qian yu Shu and Xue ping Wang. Dimensional formulas of semilinear subspaces over semirings. Linear Multilinear Algebra, 62(11):1475–1490, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shu:2017:ABM

[ySpW17] Qian yu Shu and Xue ping Wang. The applications of the bideterminant of a matrix over commutative semirings. Linear Multilinear Algebra, 65(7):1462–1478, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shu:2022:OCE

[ySpW22] Qian yu Shu and Xue ping Wang. Orthogonal complements and extending orthogonal subsets of semimodules. Linear Multilinear Algebra, 70(21):6393–6411, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shu:2021:EFS

[ySpWQ21] Qian yu Shu, Xue ping Wang, and Lei Qiao. Extension of free sets over commutative semirings. Linear Multilinear Algebra, 69(16):3019–3030, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

You:2017:SUL

[YSY17] Lihua You, Yujie Shu, and Pingzhi Yuan. Sharp upper and lower bounds for the spectral radius of a nonnegative irreducible matrix and its applications. Linear Multilinear Algebra, 65(1):113–128, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2022:CGM

[YSZW22] Guanglong Yu, Lin Sun, Hailiang Zhang, and Yarong Wu. 2connected graphs with the minimum algebraic connectivity. Linear Multilinear Algebra, 70(21):6108–6113, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:1986:GTD

[yT86] Tin yau Tam. On the generalized m th decomposable numerical radius on symmetry classes of tensors. Linear and Multilinear Algebra, 19(2):117–132, 1986. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tam:2000:GME

[yT00] Tin yau Tam. Group majorization, Eaton triples and numerical range. Linear and Multilinear Algebra, 47(1):11–28, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yan:2015:SRL

[YT15] Dan Yan and Guoping Tang. Some remarks on linear triangularizability. Linear Multilinear Algebra, 63(1):132–140, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2020:HSR

[YTL20] Shi-Fang Yuan, Yong Tian, and Ming-Zhao Li. On Hermitian solutions of the reduced biquaternion matrix equation (AXB, CXD) = (E, G). Linear Multilinear Algebra, 68(7): 1355–1373, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:1996:NNM

[Yu96] Jie-Tai Yu. Noncommutative-nilpotent matrices and the Jacobian conjecture. Linear and Multilinear Algebra, 41(2):189– 192, 1996. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2014:SMS

[Yu14] Tao Yu. Samuel multiplicity of a semi-Fredholm operator. Linear Multilinear Algebra, 62(10):1361–1364, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuanwu:1998:NRC

[Yua98] Pei Yuanwu. A numerical range characterization of Jordan blocks. Linear and Multilinear Algebra, 43(4):351–361, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuzvinsky:1984:SMP

[Yuz84] Sergey Yuzvinsky. A series of monomial pairings. Linear and Multilinear Algebra, 15(2):109–119, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2005:GIC

[YW05] Yaoming Yu and Guorong Wang. The generalized inverse over commutative rings. Linear Multilinear Algebra, 53(4):293–302, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2009:DRG

[YW09] Yaoming Yu and Y. Wei. Determinantal representation of the generalized inverse A_{T,S}^{(2)} over integral domains and its applications. Linear Multilinear Algebra, 57(6):547–559, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yongfeng:2010:DLD

[YW10] Pang Yongfeng and Yang Wei. Derivations and local derivations on strongly double triangle subspace lattice algebras. Linear Multilinear Algebra, 58(7):855–862, 2010. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2016:SQM

[YW16] Shi-Fang Yuan and Qing-Wen Wang. L-structured quaternion matrices and quaternion linear matrix equations. Linear Multilinear Algebra, 64(2):321–339, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yan:2019:SGU

[YW19] Juan Yan and Xinxia Wang. Spectra of graphs under an operation. Linear Multilinear Algebra, 67(11):2246–2252, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2007:ASB

[YWO07] Qiu Yu, Dengyin Wang, and Shikun Ou. Automorphisms of the standard Borel subalgebra of Lie algebra of C_{m} type over a commutative ring. Linear Multilinear Algebra, 55(6):545–550, 2007. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yu:2021:FRL

[YWZ21] Guanglong Yu, Yarong Wu, and Mingqing Zhai. Further results on the least Q-eigenvalue of a graph with fixed domination number. Linear Multilinear Algebra, 69(12):2274–2287, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2013:CAG

[YZ13] Wenlei Yang and Jun Zhu. Characterizations of additive (generalized) ξ-Lie (α, β) derivations on triangular algebras. Linear Multilinear Algebra, 61(6):811–830, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2014:NMP

[YZ14] Hu Yang and Pingping Zhang. A note on multiplicative perturbation bounds for the Moore–Penrose inverse. Linear Multi-

Yuan:2016:CPS

[YZ16] Yongxin Yuan and Kezheng Zuo. Compute lim_{λ→0} X(λI_{p} + Y AX)^{−1}Y by the product singular value decomposition. Linear Multilinear Algebra, 64(2):269–278, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2020:NMP

[YZ20] Zhujun Yang and Jianhua Zhang. Nonlinear maps preserving the second mixed Lie triple products on factor von Neumann algebras. Linear Multilinear Algebra, 68(2):377–390, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yang:2022:EPS

[YZB22] Yang Yang, Jiang Zhou, and Changjiang Bu. Equitable partition and star set formulas for the subgraph centrality of graphs. Linear Multilinear Algebra, 70(16):3142–3150, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2021:IQE

[YZC21] Yongxin Yuan, Lei Zhou, and Jinghua Chen. An inverse quadratic eigenvalue problem for damped gyroscopic systems. Linear Multilinear Algebra, 69(14):2572–2585, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yue:2019:LCI

[YZH19] Junjie Yue, Liping Zhang, and Zhaohui Huang. A linearly convergent iterative method for identifying H-matrices. Linear Multilinear Algebra, 67(12):2488–2503, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2017:LST

[YZL17] Yongxin Yuan, Wenhua Zhao, and Hao Liu. Least squares Toeplitz matrix solutions of the matrix equation. Linear Multilinear Algebra, 65(9):1867–1877, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yuan:2022:RNR

[YZL22] Yongxin Yuan, Huiting Zhang, and Lina Liu. The Re-nnd and Re-pd solutions to the matrix equations AX = C, XB = D. Linear Multilinear Algebra, 70(18):3543–3552, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yan:2020:GJL

[YZZ20] Kai Yan, Qingping Zeng, and Yucan Zhu. Generalized Jacobson’s lemma for Drazin inverses and its applications. Linear Multilinear Algebra, 68(1):81–93, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zaballa:1989:MPI

[Zab89] Ion Zaballa. Matrices with prescribed invariant factors and off-diagonal submatrices. Linear and Multilinear Algebra, 25 (1):39–54, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zaballa:1990:MPI

[Zab90] Ion Zaballa. Matrices with prescribed invariant factors. Linear and Multilinear Algebra, 27(4):325–343, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zheng:2022:LAP

[ZAK22] De-Yin Zheng, Ilker Akkus, and Gonca Kizilaslan. The linear algebra of a Pascal-like matrix. Linear Multilinear Algebra, 70(14):2629–2641, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zarnowski:2001:GIT

[Zar01] Roger E. Zarnowski. Generalized inverses and the total stopping times of Collatz sequences. Linear and Multilinear Algebra, 49(2):115–130, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2013:SCL

[ZB13] Jiang Zhou and Changjiang Bu. Spectral characterization of line graphs of starlike trees. Linear Multilinear Algebra, 61(8): 1041–1050, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2012:GIB

[ZBW12] Jiang Zhou, Changjiang Bu, and Yimin Wei. Group inverse for block matrices and some related sign analysis. Linear Multilinear Algebra, 60(6):669–681, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2016:LBF

[ZBW16] Lin Zhang, Kaifeng Bu, and Junde Wu. A lower bound on the fidelity between two states in terms of their trace-distance and max-relative entropy. Linear Multilinear Algebra, 64(5): 801–806, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2022:WCI

[ZC22] Yukun Zhou and Jianlong Chen. Weak core inverses and pseudo core inverses in a ring with involution. Linear Multilinear Algebra, 70(21):6876–6890, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhuang:2012:APD

[ZCCIW12] Guifen Zhuang, Jianlong Chen, Dragana S. Cvetković-Ilić, and Yimin Wei. Additive property of Drazin invertibility of elements in a ring. Linear Multilinear Algebra, 60(8):903–910, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zheng:2022:IIE

[ZCLW22] Sai-Nan Zheng, Xi Chen, Lily Li Liu, and Yi Wang. Inertia indices and eigenvalue inequalities for Hermitian matrices. Linear Multilinear Algebra, 70(8):1543–1552, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2016:FRI

[ZCP16] Huihui Zhu, Jianlong Chen, and Pedro Patrício. Further results on the inverse along an element in semigroups and rings. Linear Multilinear Algebra, 64(3):393–403, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2017:ROL

[ZCP17] Huihui Zhu, Jianlong Chen, and Pedro Patrício. Reverse order law for the inverse along an element. Linear Multilinear Alge-

Zou:2018:CEE

[ZCP18] Honglin Zou, Jianlong Chen, and Pedro Patrício. Characterizations of m-EP elements in rings. Linear Multilinear Algebra, 66(6):1244–1256, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2021:CIL

[ZCW21] Mengmeng Zhou, Jianlong Chen, and Dingguo Wang. The core inverses of linear combinations of two core invertible matrices. Linear Multilinear Algebra, 69(4):702–718, 2021. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2022:NDG

[ZCY22] Zerui Zhang, Yuqun Chen, and Bing Yu. No dialgebra has Gelfand–Kirillov dimension strictly between 1 and 2. Linear Multilinear Algebra, 70(1):42–52, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2012:IPM

[ZCZ12] Yang Zhang, Jun Cao, and Bao-Dong Zheng. Idempotence preserving maps between matrix spaces. Linear Multilinear Algebra, 60(3):349–375, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2022:WGI

[ZCZT22] Mengmeng Zhou, Jianlong Chen, Yukun Zhou, and Néstor Thome. Weak group inverses and partial isometries in proper ∗-rings. Linear Multilinear Algebra, 70(19):4528–4543, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2018:MBG

[ZD18] Hui Zhou and Qi Ding. On mixed block graphs. Linear Multilinear Algebra, 66(8):1668–1680, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zelmanowitz:1975:IMA

[Zel75] Julius Zelmanowitz. Invertibility of matrices with almostcommuting entries. Linear and Multilinear Algebra, 3(1–2):

Zellini:1985:ACP

[Zel85] P. Zellini. Algebraic and computational properties of a set of (0, 1) matrices with prescribed sum. Linear and Multilinear Algebra, 17(3–4):355–373, 1985. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zeng:2022:SLE

[Zen22] Xingyuan Zeng. Spectrum of large Euclidean random matrices generated from l_{p} ellipsoids. Linear Multilinear Algebra, 70(8): 1403–1422, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zelmanowitz:1978:BR

[ZF78] J. Zelmanowitz and Joel Franklin. Book reviews. Linear and Multilinear Algebra, 6(1):91–92, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2017:RNE

[ZFSX17] Xiaobo Zhang, Xuxu Fang, Chuanning Song, and Qingxiang Xu. Representations and norm estimations for the Moore– Penrose inverse of multiplicative perturbations of matrices. Linear Multilinear Algebra, 65(3):555–571, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2008:CBO

[ZG08] Bo Zhou and Ivan Gutman. A connection between ordinary and Laplacian spectra of bipartite graphs. Linear Multilinear Algebra, 56(3):305–310, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2011:CMD

[ZG11] Xiaoling Zhang and Chris Godsil. Connectivity and minimal distance spectral radius of graphs. Linear Multilinear Algebra, 59(7):745–754, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zaman:2022:RBI

[ZH22] Shahid Zaman and Xiaocong He. Relation between the inertia indices of a complex unit gain graph and those of its underlying graph. Linear Multilinear Algebra, 70(5):843–877, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:1989:NHP

[Zha89] Fuzhen Zhang. Notes on Hadamard products of matrices. Linear and Multilinear Algebra, 25(3):237–242, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:1992:RP

[Zha92] Fuzhen Zhang. Research problem. Linear and Multilinear Algebra, 33(3–4):301–303, 1992. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhan:1995:QPR

[Zha95] Xingzhi Zhan. On a question of P. R. Halmos. Linear and Multilinear Algebra, 39(3):255–258, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:1997:WLA

[Zha97] Hechun Zhang. Witt-like algebras and their q-analogues. Linear and Multilinear Algebra, 42(3):221–231, 1997. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhan:2000:SRP

[Zha00] Xingzhi Zhan. Some research problems on the Hadamard product and singular values of matrices. Linear and Multilinear Algebra, 47(2):191–194, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhan:2001:SOO

[Zha01a] Xingzhi Zhan. Span of the orthogonal orbit of real matrices. Linear and Multilinear Algebra, 49(4):337–346, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2001:EWI

[Zha01b] Fuzhen Zhang. Equivalence of the Wielandt inequality and the Kantorovich inequality. Linear and Multilinear Algebra, 48(3): 275–279, 2001. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2004:MIS

[Zha04a] Fuzhen Zhang. A matrix identity on the Schur complement. Linear Multilinear Algebra, 52(5):367–373, 2004. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2004:AMP

[Zha04b] Xian Zhang. Additive maps preserving Moore–Penrose inverses of matrices on symmetric matrix spaces. Linear Multilinear Algebra, 52(5):349–358, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2004:GCH

[Zha04c] Xian Zhang. The general common Hermitian nonnegativedefinite solution to the matrix equations AXA^{∗} = B and CXC^{∗} = D. Linear Multilinear Algebra, 52(1):49–60, 2004. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2005:LPR

[Zha05] Xian Zhang. Linear preservers of rank-sum-maximum, rank, rank-subtractivity, and rank-sum-minimum on symmetric matrices. Linear Multilinear Algebra, 53(3):153–165, 2005. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2013:AMA

[Zha13] Yun Zhang. On the angle and the minimal angle between subspaces. Linear Multilinear Algebra, 61(12):1660–1667, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2014:ESM

[Zha14a] Lin Zhang. Entropy, stochastic matrices, and quantum operations. Linear Multilinear Algebra, 62(3):396–405, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2014:DPL

[Zha14b] Jianrong Zhao. Divisibility of power LCM matrices by power GCD matrices on gcd-closed sets. Linear Multilinear Algebra, 62(6):735–748, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2015:MDA

[Zha15a] Fuzhen Zhang. A matrix decomposition and its applications. Linear Multilinear Algebra, 63(10):2033–2042, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2015:FER

[Zha15b] Pingping Zhang. A further extension of Rotfel’d theorem. Linear Multilinear Algebra, 63(12):2511–2517, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2015:CDL

[Zha15c] Tao Zhang. Cohomology and deformations of 3-Lie colour algebras. Linear Multilinear Algebra, 63(4):651–671, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2016:NSJ

[Zha16a] Fangjuan Zhang. Nonlinear skew Jordan derivable maps on factor von Neumann algebras. Linear Multilinear Algebra, 64 (10):2090–2103, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2016:NPD

[Zha16b] Guan-Lin Zhang. A note on positive definite solution of matrix equation X + M^{∗}X^{−1}M − N^{∗}X^{−1}N = 1. Linear Multilinear Algebra, 64(5):951–954, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2019:TNS

[Zha19] Jianxing Zhao. Two new singular value inclusion sets for rectangular tensors. Linear Multilinear Algebra, 67(12):2451–2470, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2020:MBP

[Zha20a] Li-Tao Zhang. Modified block preconditioner for generalized saddle point matrices with highly singular (1, 1) blocks. Linear Multilinear Algebra, 68(1):152–160, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2020:FCP

[Zha20b] Qinghong Zhang. Faces of the 5 × 5 completely positive cone. Linear Multilinear Algebra, 68(12):2523–2540, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2021:SFF

[Zha21] Deke Zhao. A super Frobenius formula for the characters of cyclotomic Hecke algebras. Linear Multilinear Algebra, 69(12): 2169–2186, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2022:LCP

[Zha22] Lijun Zhao. Linear constraint problem of Hermitian unitary symplectic matrices. Linear Multilinear Algebra, 70(8):1423– 1441, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2009:MPN

[ZHH09] Xiuling Zhang, Jinchuan Hou, and Kan He. Maps preserving numerical radius and cross norms of operator products. Linear Multilinear Algebra, 57(5):523–534, 2009. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2013:SSA

[ZHLZ13] Ling Zhang, Ting-Zhu Huang, Zhongshan Li, and Jing-Yue Zhang. Several spectrally arbitrary ray patterns. Linear Multilinear Algebra, 61(4):543–564, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhong:2016:UBM

[Zho16] Jin Zhong. On upper bounds for the minimum rank of regular classes of (0, 1)-matrices. Linear Multilinear Algebra, 64(8): 1550–1561, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2022:RWM

[Zho22] Yiqiang Zhou. Rings over which matrices are products of qpotents. Linear Multilinear Algebra, 70(21):6375–6392, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2015:BLE

[Zhu15] Bao-Xuan Zhu. Bound on the least eigenvalue of a graph with cut vertices. Linear Multilinear Algebra, 63(5):993–1001, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2018:SCG

[Zhu18] Huihui Zhu. Several characterizations for generalized inverses in a ring. Linear Multilinear Algebra, 66(7):1351–1361, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2019:DIE

[Zhu19] Huihui Zhu. On DMP inverses and m-EP elements in rings. Linear Multilinear Algebra, 67(4):756–766, 2019. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2022:CBT

[Zhu22] Haixing Zhu. The construction of braided tensor categories from Hopf braces. Linear Multilinear Algebra, 70(16):3171– 3188, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2011:IEP

[ZHZ11] Lijun Zhao, Xiyan Hu, and Lei Zhang. Inverse eigenvalue problems for bisymmetric matrices under a central principal submatrix constraint. Linear Multilinear Algebra, 59(2):117– 128, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zuo:2022:DCD

[ZIC22] Kezheng Zuo, Dragana Cvetković Ilić, and Yingjie Cheng. Different characterizations of DMP-inverse of matrices. Linear Multilinear Algebra, 70(3):411–418, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zimmerman:1989:BTP

[Zim89] Dale L. Zimmerman. Block Toeplitz products of block Toeplitz matrices. Linear and Multilinear Algebra, 25(3):185–190, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zimmermann:1998:MCC

[Zim98] Georg Zimmermann. A minimax-condition for the characteristic center of a tree. Linear and Multilinear Algebra, 45(2–3): 161–187, 1998. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zheng:2022:MSC

[ZJC^{+}22] Yanpeng Zheng, Xiaoyu Jiang, Xiaoting Chen, Ancai Zhang, and Fawaz Alsaadi. Means and the Schur complement of sector matrices. Linear Multilinear Algebra, 70(13):2611–2617, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2015:DEL

[ZL15] Tao Zhang and Juan Li. Deformations and extensions of Lie– Yamaguti algebras. Linear Multilinear Algebra, 63(11):2212– 2231, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2017:LMM

[ZL17] Jun Zhu and Pan Li. Local multivariate majorization preserving linear mappings. Linear Multilinear Algebra, 65(3): 600–612, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2018:SVI

[ZL18] Jianxing Zhao and Chaoqian Li. Singular value inclusion sets for rectangular tensors. Linear Multilinear Algebra, 66(7): 1333–1350, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2019:MSR

[ZL19] Jianbin Zhang and Jianping Li. The maximum spectral radius of k-uniform hypergraphs with r pendent vertices. Linear Multilinear Algebra, 67(5):1062–1073, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2021:HPD

[ZL21] Juan Zhang and Shifeng Li. On the Hermitian positive definite solution and Newton’s method for a nonlinear matrix equation. Linear Multilinear Algebra, 69(11):2093–2114, 2021.

Zhang:2022:EHT

[ZL22] Huihui Zhang and Shuchao Li. Extremal hitting times of trees with some given parameters. Linear Multilinear Algebra, 70 (11):2127–2149, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhai:2011:EGT

[ZLS11] Mingqing Zhai, Ruifang Liu, and Jinlong Shu. An edgegrafting theorem on Laplacian spectra of graphs and its application. Linear Multilinear Algebra, 59(3):303–315, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2022:CRR

[ZLS22] Jia Zhao, Jiefeng Liu, and Yunhe Sheng. Cohomologies and relative Rota–Baxter–Nijenhuis structures of 3-LieRep pairs. Linear Multilinear Algebra, 70(21):6240–6264, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zheng:2020:PST

[ZLZ20] Shasha Zheng, Xiaogang Liu, and Shenggui Zhang. Perfect state transfer in NEPS of some graphs. Linear Multilinear Algebra, 68(8):1518–1533, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2022:LSB

[ZLZ^{+}22] Yanzhen Zhang, Ying Li, Hong Zhao, Jianli Zhao, and Gang Wang. Least-squares bihermitian and skew bihermitian solutions of the quaternion matrix equation AXB = C. Linear Multilinear Algebra, 70(6):1081–1095, 2022. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zamani:2019:NPD

[ZMCN19] Ali Zamani, Mohammad Sal Moslehian, Mao-Ting Chien, and Hiroshi Nakazato. Norm-parallelism and the Davis–Wielandt radius of Hilbert space operators. Linear Multilinear Algebra, 67(11):2147–2158, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2020:DIS

[ZMG20] Daochang Zhang, Dijana Mosić, and Li Guo. The Drazin inverse of the sum of four matrices and its applications. Linear Multilinear Algebra, 68(1):133–151, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2018:SER

[ZN18] Jianguo Zhao and Keshe Ni. Some extensions of Rotfel’d theorem. Linear Multilinear Algebra, 66(2):410–417, 2018. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zou:2012:CCF

[Zou12] Limin Zou. On a conjecture concerning the Frobenius norm of matrices. Linear Multilinear Algebra, 60(1):27–31, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zou:2020:NSV

[Zou20] Limin Zou. A note on singular value inequalities related to the Audenaert–Zhan inequality. Linear Multilinear Algebra, 68(9):1792–1796, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2019:CPC

[ZP19] Huihui Zhu and Pedro Patrício. Characterizations for pseudo core inverses in a ring with involution. Linear Multilinear Algebra, 67(6):1109–1120, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2016:IAP

[ZPCZ16] Huihui Zhu, Pedro Patrício, Jianlong Chen, and Yulin Zhang. The inverse along a product and its applications. Linear Multilinear Algebra, 64(5):834–841, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhan:2018:EZS

[ZQC18] Fuqin Zhan, Youfu Qiao, and Junliang Cai. On edge-Zagreb spectral radius and edge-Zagreb energy of graphs. Linear Multilinear Algebra, 66(12):2512–2523, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2021:CSC

[ZQZ21] Mu-Zheng Zhu, Ya-E Qi, and Guo-Feng Zhang. On circulant and skew-circulant preconditioned Krylov methods for steadystate Riesz spatial fractional diffusion equations. Linear Multilinear Algebra, 69(4):719–731, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2005:ARO

[ZS05] Xian Zhang and Nung-Sing Sze. Additive rank-one preservers between spaces of rectangular matrices. Linear Multilinear Algebra, 53(6):417–425, 2005. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Song:2016:UBN

[zSqSZ16] Ya zhi Song, Xiao qiu Song, and Mingcui Zhang. An upper bound for the nullity of a bipartite graph in terms of its maximum degree. Linear Multilinear Algebra, 64(6):1107–1112, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2012:URA

[ZT12] Xiufu Zhang and Shaobin Tan. θ-Unitary representations for the W -algebra W (2, 2). Linear Multilinear Algebra, 60(5):533– 543, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zusmanovich:2022:CBT

[Zus22] Pasha Zusmanovich. On contact brackets on the tensor product. Linear Multilinear Algebra, 70(19):4695–4706, 2022. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zheng:2019:ICT

[ZV19] Hua Zheng and Seakweng Vong. Improved convergence theorems of the two-step modulus-based matrix splitting and synchronous multisplitting iteration methods for solving linear complementarity problems. Linear Multilinear Algebra, 67(9): 1773–1784, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zheng:2018:PBL

[ZVL18] Hua Zheng, Seakweng Vong, and Wen Li. On perturbation bounds of the linear complementarity problem. Linear Multilinear Algebra, 66(3):625–638, 2018. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2014:LMM

[ZW14] Jinming Zhou and Dengyin Wang. Linear maps on matrix algebra Jordan derivable at involutory matrices. Linear Multilinear Algebra, 62(7):913–917, 2014. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2017:DSL

[ZW17a] Qiannan Zhou and Ligong Wang. Distance signless Laplacian spectral radius and Hamiltonian properties of graphs. Linear Multilinear Algebra, 65(11):2316–2323, 2017. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2017:SSS

[ZW17b] Qiannan Zhou and Ligong Wang. Some sufficient spectral conditions on Hamilton-connected and traceable graphs. Linear Multilinear Algebra, 65(2):224–234, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2019:SMI

[ZW19] Jianguo Zhao and Junliang Wu. Some matrix inequalities related to Kwong functions. Linear Multilinear Algebra, 67(5): 995–1005, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2020:WPC

[ZW20] Huihui Zhu and Qing-Wen Wang. Weighted pseudo core inverses in rings. Linear Multilinear Algebra, 68(12):2434–2447, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2015:PSC

[ZWL15] Heping Zhang, Tingzeng Wu, and Hong-Jian Lai. Per-spectral characterizations of some edge-deleted subgraphs of a complete graph. Linear Multilinear Algebra, 63(2):397–410, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2022:EML

[ZWLZ22] Fengxia Zhang, Musheng Wei, Ying Li, and Jianli Zhao. An efficient method for least-squares problem of the quaternion matrix equation X − AX̂B = C. Linear Multilinear Algebra, 70(13):2569–2581, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2017:AGT

[ZWM17a] Jinming Zhou, Dein Wong, and Xiaobin Ma. Automorphism group of the total graph over a matrix ring. Linear Multilinear Algebra, 65(3):572–581, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2017:AZD

[ZWM17b] Jinming Zhou, Dein Wong, and Xiaobin Ma. Automorphisms of the zero-divisor graph of the full matrix ring. Linear Multilinear Algebra, 65(5):991–1002, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2023:OSC

[ZWM23] Huihui Zhu, Liyun Wu, and Dijana Mosić. One-sided w-core inverses in rings with an involution. Linear Multilinear Algebra, 71(4):528–544, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhou:2020:LBG

[ZWS20] Qi Zhou, Dein Wong, and Dongqin Sun. A lower bound for graph energy. Linear Multilinear Algebra, 68(8):1624–1632, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2010:MPS

[ZWW10] Yanxia Zhao, Dengyin Wang, and Xuansheng Wang. Maps preserving square-zero matrices on the algebra consisting of all upper triangular matrices. Linear Multilinear Algebra, 58(1): 121–136, 2010. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2022:OSC

[ZWW22] Liang Zhao, Cang Wu, and Yao Wang. One-sided central Drazin inverses. Linear Multilinear Algebra, 70(7):1193–1206,

Zhou:2010:SZD

[ZWZ10] Jinming Zhou, Dengyin Wang, and Qinghua Zhang. Squarezero derivations of matrix algebras over commutative rings. Linear Multilinear Algebra, 58(2):239–243, 2010. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhao:2023:RIM

[ZWZ23] Jing Zhao, Xiang Wang, and Jianhua Zhang. A randomised iterative method for solving factorised linear systems. Linear Multilinear Algebra, 71(2):242–255, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zheng:2010:ROL

[ZX10] Bing Zheng and Zhi-Ping Xiong. The reverse order laws for {1, 2, 3}- and {1, 2, 4}-inverses of multiple matrix products. Linear Multilinear Algebra, 58(6):765–782, 2010. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2016:FPT

[ZX16] Haiyan Zhang and Mingzhi Xue. Fixed points of trace preserving completely positive maps. Linear Multilinear Algebra, 64(3):404–411, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zheng:2015:LMP

[ZXF15] Baodong Zheng, Jinli Xu, and Ajda Fošner. Linear maps preserving rank of tensor products of matrices. Linear Multilinear Algebra, 63(2):366–376, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2016:CAD

[ZXL16] Jun Zhu, Changping Xiong, and Pan Li. Characterizations of all-derivable points in B(H). Linear Multilinear Algebra, 64(8):1461–1473, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhai:2022:EPS

[ZXL22] Mingqing Zhai, Jie Xue, and Ruifang Liu. An extremal problem on Q-spectral radii of graphs with given size and matching number. Linear Multilinear Algebra, 70(20):5334–5345, 2022. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2008:MTE

[ZY08] Zhanmin Zhu and Jinxiang Yu. On GC_{2} modules and their endomorphism rings. Linear Multilinear Algebra, 56(5):511– 515, 2008. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Yan:2011:CKV

[zY11] Zi zong Yan. A Cauchy–Khinchin–van Dam matrix inequality. Linear Multilinear Algebra, 59(7):825–829, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2013:PAT

[ZY13] Lei-Hong Zhang and Wei Hong Yang. Perturbation analysis for the trace quotient problem. Linear Multilinear Algebra, 61(12):1629–1640, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2017:RWG

[ZY17] W. J. Zhang and A. M. Yu. On the rank of weighted graphs. Linear Multilinear Algebra, 65(3):635–652, 2017. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2013:NPB

[ZYL13] Pingping Zhang, Hu Yang, and Hanyu Li. New perturbation bounds for the subunitary polar factors. Linear Multilinear Algebra, 61(4):517–526, 2013. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zobin:2000:GSO

[ZZ00] Nahum Zobin and Veronica Zobin. Geometric structure of b_{2,2}-orbihedra and interpolation of operators. Linear and Multilinear Algebra, 48(1):67–91, 2000. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2015:SLC

[ZZ15] Jie Zhang and Xiao-Dong Zhang. Signless Laplacian coefficients and incidence energy of unicyclic graphs with the matching number. Linear Multilinear Algebra, 63(10):1981–2008,

Zariouh:2016:PWO

[ZZ16] H. Zariouh and H. Zguitti. On pseudo B-Weyl operators and generalized Drazin invertibility for operator matrices. Linear Multilinear Algebra, 64(7):1245–1257, 2016. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhu:2021:HTR

[ZZ21] Xiaomin Zhu and Xiao-Dong Zhang. The hitting time of random walk on unicyclic graphs. Linear Multilinear Algebra, 69(4):573–592, 2021. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2023:DFD

[ZZ23] Ying Zhang and Wenlong Zhang. The q-derivative formula and determinant identities. Linear Multilinear Algebra, 71(14): 2325–2336, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zivkovic-Zlatanovic:2017:GKR

[ZZC17] Snezana C. Zivković-Zlatanović and Milos D. Cvetković. Generalized Kato–Riesz decomposition and generalized Drazin– Riesz invertible operators. Linear Multilinear Algebra, 65(6): 1171–1193, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2023:MGS

[ZZL23] Jia-Lin Zhang, Guo-Feng Zhang, and Zhao-Zheng Liang. A modified generalized SOR-like method for solving an absolute value equation. Linear Multilinear Algebra, 71(9):1578–1595, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zong:2023:MSI

[ZZT23] Chunxiang Zong, Guofeng Zhang, and Yuchao Tang. Multistep inertial forward-backward-half forward algorithm for solving monotone inclusion. Linear Multilinear Algebra, 71(4): 631–661, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zeng:2016:ERD

[ZZY16] Qingping Zeng, Huaijie Zhong, and Kai Yan. An extension of a result of Djordjević and its applications. Linear Multilinear Algebra, 64(2):247–257, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2012:FKT

[ZZZ12] Guang-Jun Zhang, Jie Zhang, and Xiao-Dong Zhang. Faber– Krahn type inequality for unicyclic graphs. Linear Multilinear Algebra, 60(11–12):1355–1364, 2012. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2016:PBS

[ZZZ16] Shifang Zhang, Lin Zhang, and Huaijie Zhong. Perturbation of Browder spectrum of upper triangular operator matrices. Linear Multilinear Algebra, 64(3):502–511, 2016. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Zhang:2023:MLE

[ZZZ23] Haixia Zhang, Ning Zhang, and Zhuolin Zhang. The maximum Laplacian Estrada index of connected graphs. Linear Multilinear Algebra, 71(2):178–189, 2023. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).